A Complete Bibliography of the Journal of Statistical Physics: 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

31 May 2024 Version 1.31

Title word cross-reference

#SAT [BMT15].

(2 + 1) [XTpXpH12, CTH^{+}11]. + [Zuc11b]. 0 [Fed17]. 1 [BELP15, CAS11, Cor16, Fed17, GDL10, GBL16, Hau16, JV19, KT12, KM19c, Li19, MN14b, Nak17, Pal11, Pan14, RT14, RBS16b, RY12, SS18c, Sug10, dOP18]. 1 + 1 [Sak18, CP15b]. 1/2 [MD10]. 1/f [FDR12]. 1/n [Per17]. 1/|x − y|^{2} [MSV10, MSV13]. 13 [DFL17]. 1 ≤ p ≤ ∞ [Dud13]. 1/f^{α} [HPF15]. 2 [AB19, ADS^{+}19, BF12, BNT13, DSS15, EKD12, Her13, Ily12, Lan10, Li12, Li19, LZ11, Ny13, Ost16, PSS16, ST14, Sch13b, TJ15, WPB15, dWL10]. 2 + 1 [dWL14]. 2.5 [BC15a]. 2R [WLEC17]. 2 × 2 [CLTT13]. 3 [BCF19, BLS17, ESPP^{+}14, Kar18, SH16, SWKS14, dCCS19]. 3/2 [DK10]. 38 [Cam13]. 4 [BBS14, Zha14]. 4 × 4 [LN19a]. 5/2 [DK10, EKD12]. 6 [EC11]. 8 [Zha14]. 90^{∘} [YM11]. ^{3} [Afz12]. _{1−x} [EFO11]. _{13} [CDCL18]. _{2} [ML15, QR13, ST11c]. _{4} [HBB10]. _{6} [BCL10a, BCL10b, EFO11]. _{x} [EFO11].

A + A → ⊘ [KCB13]. α [CP18, DDN14, DXZ14, LT10b, ZGL13, ZGL15]. b [She15]. d̄ [GGP10]. β [DHR18, DDN14, HF12b, Koz17, Kum19, MMA15, SW12]. β = 1 [AP14a]. β = 4 [AP14a]. c [ESPP^{+}14]. C^{∗} [ADR18]. C^{1} [Tia14]. C^{1+α} [Tia14]. d [BHNY15, BNY16, Bis19, CMS13, CvELR18, Ein12, Far15, GJ15a, SS16, WX15]. d = 1 [Far15]. d = 2 [TV12]. d_{2} [Tos17]. d ≥= 2 [MW12a]. ϵ [BC15b, Sou18]. Γ [PST12]. H [BJM15, SB15a, AU15]. SLE_{6} [Ken19b]. J_{1} [MD10]. J_{2} [MD10]. K [PRD11, TVP13]. k = O(n) [dMPTW16]. k · p [BF10]. L^{1} [ASA15]. L^{2} [HX15b, MC11, TV16]. L^{p} [Dud13]. λ_{d} [Fed13]. m [KID^{+}11]. N [BZ13, GGJR14, Mar15]. R_{v}^{l} [JL16]. Z [LMM16]. Z^{2} [ABMP16, AD15b, DP19]. Z^{d} [DEK^{+}15, AFGL15]. Z^{d+s} [SS17, SS18b]. Z^{d}, d ≥ 3 [CG15]. Z_{p} [CAS11]. H [Kha19]. U_{q}[osp(1|2)] [LS16c]. }↕ (n|m) [MM14]. U_{q}(∫⊓(1, 1)) [CGRS16]. N [CD14a, CB16, Cha14a, Kie14, KMS14, Leh13, NBK14, RE13, Sch12a, SMCF13, BBS14, CCG14b, DRCV19, XY13]. N^{3/4} [MPSS19, Sch13b]. O(N) [DTW19, KN16, LSW17]. p [ALS18, AJ19, CP17b, Pan14]. P (φ)_{2} [FZ11b]. ±Z^{d} [Tyo12]. ±Z^{d}, d ≥ 3 [CG11a]. q [ALS18, BDP19, BCF10, CS15, CP15b, GN13, MSS15, MLCPS13]. R [KS13]. R^{d} [IPS10a]. s [GJ15a]. SLE(2) [AKM13]. SLE(κ; ρ) [Zha10a]. SLE_{8/3} [Ken12]. SO(3) [BJGL^{+}17]. SO(m + 1) [Mac10]. T [ISZ16]. → [Gal13]. φ [AvB16]. XY [CR16]. Z [Kel19]. |φ|^{4} [BBS14].

-adic [ALS18]. -ary [DRCV19]. -Body [Kie14]. -Clique [TVP13]. -Colour [dCCS19]. -Component [BBS14]. -Conjecture [WLEC17]. -Cycle [Cha14a]. -d [ADS^{+}19, BLS17, PSS16, KT12, Pal11, RBS16b, RY12]. -Deformed [GN13]. -Dimensional [BELP15, BBS14, BNY16, Bis19, CTH^{+}11, CMS13, WX15, XTpXpH12, dCCS19]. -Dimensions [BHNY15]. -distance [GGP10]. -Divergences [AvB16]. -Ensemble [SW12]. -Ensembles [DHR18, Koz17]. -Erlang [PRD11]. -Exponential [ISZ16]. -Invariance [Kel19]. -Jacobi [HF12b]. -Limit [PST12]. -Matrix [KS13]. -Minimization [ESPP^{+}14]. -Model [CP18]. -Moment [DFL17]. -Norm [MC11]. -Norms [TV16]. -Particle [RE13]. -Point [KID^{+}11, NBK14]. -Potts [BCF10]. -Processes [AK19]. -PushASEP [CP15b]. -Quadratic [ADR18]. -RSB [Pan14]. -Spin [AJ19, CP17b, Pan14]. -Stability [HX15b]. -Stable [DXZ14, ZGL15, ZGL13]. -State [ALS18, CS15, EC11, LN19a]. -Theorem [Kha19, SB15a]. -Transducer [BC15b]. -Transform [AU15]. -Voter [MLCPS13]. -Wishart [Kum19]. -Zero [BDP19].

1 [FM18, Ker13, SS16]. 103rd [Leb10]. 105th [Leb12a, Leb12c]. 106th [Ano12a]. 107th [Ano12b]. 108th [Leb13b]. 109th [Leb13c]. 110th [Ano15a]. 111th [Ano16a, Leb14]. 113th [Ano16a]. 119th [Ano18a]. 120th [Ano19a]. 121st [Ano19b]. 16-18 [Ano19a]. 18-20 [Ano12a]. 1D [IS11, IS13].

2011 [Ano12a, Leb12a, Leb12c]. 2012 [Ano12b, Leb13b]. 2013 [Leb13c]. 2014 [Leb14].

7th [NBK14].

80th [Gal17a].

ABC [ACL^{+}11, BLS11, BCP11, BB13, BD11a]. Abelian [JRS15, RRS12, Tyo12]. Abrupt [BJ16]. Absence [BC11, CvELR18, Ito18, PS18]. Absolute [AG12b, BK15, FAB16]. Absorbers [GTT14]. Absorbing [GW15a, WPB15]. Absorption [BPF^{+}14]. Abstract [MP13, Tem14]. Abundance [HM13b]. Accelerated [BT12, Bur17]. Accelerating [GXL12]. Acceleration [Bur19, Els12, Fil16, KJZ17, WHC14, FLS12]. Acceleration/ [FLS12]. Accessibility [DRCV19, FK12, SK14]. According [LPS12]. Account [QLCL16]. Accumulated [CB16]. Accuracy [AG15b, HNZ16, TLC13]. Accurate [Sug10, Wil11, WZIG14]. Achieving [GGP10]. Acquisition [BFFS16]. Action [LK18, LE15, PSVG18]. Actions [CRV17]. Activated [CRS14, DRS10]. Active [AG15a, BCMP15, DUU15, RBGV12, WG19]. Activity [NKR15, Vid15]. Adaptation [LW18]. Adaptive [ABT10, Chu16, CM12b, FT18a, FT18b, FJLS18, HZS11, KR16, PPK11, ST11b, TSS13]. Adapts [SKT16]. Additional [Her13]. Additive [BCW13, BKK15, IY14]. Adiabatic [AFG12, BK11, GPP12, IiS15b, Kha19]. adic [ALS18]. Adjacent [YKS16]. Adjustment [KNPF19]. Adsorption [BIP11, Cie17, DvLM16, DvLM18, RW11]. Advected [GM13, PW13]. Advection [CJW17, GdK10, GW15a, HS10, KCB13, RPPF15, Zha10c]. Advection-Diffusion [GW15a]. Advection-Mediated [HS10]. Adventures [Kie13]. Adversarial [MSB13]. Adversity [HZS11]. Affect [MC10]. Affiliation [BG14]. Affine [Fer14]. Affinity [BC14a]. After [Buc16, Gal17a, Gin14a, Gin14b, Ras12, Zia10]. After-Dinner [Gin14a, Gin14b]. Again [AWE16]. Against [HZS11]. Age [CG16, FC13, MLS16]. Age-Dependent [FC13]. Age-Structured [CG16]. Agents [BELP15, DUU15, DM11]. Ages [SV16]. Agglomeration [KPR18]. Aggregate [KOT11, KO15]. Aggregated [JYZ11]. Aggregation [EM10, FC17, LLJH10, SB14b]. Aggregations [FW15]. Aging [DWTW16, FMAG11, GvdHW17, Gre16, NVL11]. Air [JMSW13]. Airfoil [PM17]. Airy [AP14a, CB16, HDP17, Pim18, QR13, ST11c]. Aizenman [Ano18a]. AKLT [LSY19]. Alcohol [DT19]. Alex [VL12]. Algebra [BS15a, Gro19, KS13]. Algebraic [Ohk14]. Algebras [BS15d]. Algorithm [Ark10, Cli10, CM12b, GOPS11, KBSM16, Moh11, RBS16b, TM10]. Algorithmic [AP14b, Cha15b, HiS19, RMS19]. Algorithms [CVE14, CF19a, EV14, Har11, HSZ19, SST14, Wu14, Yed11]. Alignment [BCMP15, DFR14, DMP17]. Alignments [HM13a, HM16]. Allen [RBS16b]. Alliances [Gal13]. Allocation [HZ13]. Allowing [Pir18b]. Alloy [CE12]. Alloy-Type [CE12]. Almost [Bra14, CYZ18, Hay15, IY14, LR15, Ny13, ST16b]. Almost-Additive [IY14]. Almost-Periodic [Bra14]. Along [Mar16, MG17, Sam16, Zha13]. Alphabet [MU15]. Alphabets [DMP17]. also [Lep15]. Alternate [SSR12]. Alternated [Dym15]. Alternating [CMM14, DMM14, Lee10, Sam13]. Alternations [PRD11]. Alternative [KNPF19]. Altmann [Ero14]. Amenable [Fid15]. Amistad [Kad14]. Among [DUU15, FAM13, GT15b, Kos13, ORS16, QLCL16, SR19]. Amorphous [DKKP14]. Amphiphilic [Mül11]. Amplifier [AFS^{+}13, MC10, MC11]. Amplitude [FL12, RBS16b, Sam16]. Amplitudes [CG13, DG14]. Analog [Gou15]. Analogical [BGG10]. Analogues [LTM16]. Analogy [RSY14]. Analyses [CLMK18]. Analysis [AGR19, AM12, AK14b, Art19a, ACGM18, AKL18, Aza11, BCW13, BP18, BT11, BBS15a, BC19, BM12a, BS15f, BP15, CLS11a, CGN16, ES13, EC11, Fid15, Fil16, GRT17, HA13, HSZ19, HT15, He14, HIK^{+}18, KN13, Leh13, Lim16, LO17, LY16, MdG13, MLCPS13, Mor12, OK14, Osi16, RST16, RGL11, She15, SCSS18, SV10, VAY^{+}12, WGLE11, WYG16, Zha13, dAPdA^{+}13, vKSZ18]. Analytic [MS19c, Neu14, O’C12a, She15]. Analytical [AFCA16, CE14, GLML16, Ily12, KSSH15, KT12, Mar16, Mas13, PR19, SB14a, TFES19, del12, SB14b]. Analyticity [BdlL13, CdlL10, KKN12, SLM15]. Analyzing [LH17, MR13]. and/or [CL14, OC12c]. Andersen [Fan17a]. Anderson [BR10, BN15, CE12, Chu16, CGG^{+}11, EFO11, EH12, HY19b, IM16, KN13, MW12a, PdOC17, Tau11, Wre12]. Anderson-like [MW12a]. Anderson-Witting [HY19b]. Anecdotes [Gin14a, Gin14b]. Angle [ADH12]. Angles [LK18]. Angular [MWY16]. Anharmonic [BO11, BM16, KT12, MS17, Spo14]. Animal [LNP19]. Anisotropic [Bal17, Bal18, CCP16, CdLS13, CCFR18, OK14, QD12, SSR12, SS17, SS18b, vEF12]. Annealed [Can17, DGGvdH18, OR16, dAPS11]. Annealing [LTM16, ME11]. Annihilation [LPS19, NWL19, KCB13]. Annihilations [CF16]. Annuli [GGD16]. Annulus [WYG16, Zha12a]. Anomalies [HJR^{+}11, Sep13]. Anomalous [AT12, BL15, BFNZ11, CDS19, CTB10, DSPC14, ED15, GDL10, Gre16, LL19, MMA15, MBWC16, NKK19, Shi13, TFES19, ZLL13, ZL19]. Anomaly [Bir18, RF18]. Ansatz [BS19, Lee11, Sim11]. Antagonistic [ZB13]. Anti [JJB14, Ras12, SST14]. Anti-Ferromagnetic [SST14, JJB14]. Anti-norms [Ras12]. Antialigning [BC11]. Antibodies [HZS11]. Anticonformity [NSW13]. Antiferromagnetic [BCPS18, MD10, QD12, SS11a]. Antiquadrupolar [LN11, Lim16]. Antiquadrupolar-Disordered [Lim16]. Antishocks [BS13a]. Antisymmetry [ABF16]. Antoine [Mon12a]. Anyons [Ark13, LR15]. Aoyama [Sta11]. Apollonian [HDS15]. Appearance [Ber14]. Applicability [Rab11]. Application [AG12a, Afz12, Bae11, BL15, CVE14, CN13, DHK11, EFO11, Fre15, KKA19, KD19, KiMM13, LS14a, LW19, MW10, PS19, Sch12a, Tak09, YBF^{+}17, Cam13]. Applications [ABF16, BK11, CCR17, FP11, FBR19, Fun14, GJS17, GT10, Han14, MS19b, Ost16, PCM15, Shi15, DMY13, Dyk14]. Applied [Ark10, CA19, Ero14, GLM^{+}15, OC12b]. Appreciation [Fis11, Fis12]. Approach [AP10, Alb16, AF14b, BS11, Bec10, BPP18, Bon15, BMR10, CFL17, CEL^{+}18, CSC14, Cha14b, Cha15b, CAG^{+}13, CL15, CLP17, Col14, DG14, EKD12, FLTV11, GSSV11, HMO12, HA19, JLMG11, KLM13, KA17, KK15, LHZ^{+}19, Lia13, LO17, Luc12, LLS13, MSS^{+}11a, NT17, NP16, NG10, PCMM18, RKGZ12, RBR11, San18, SS16, Shi16, TGP12, WL13, ZJ14]. Approaches [dMS17a]. Approaching [JRS15]. Approximate [ART15, ZW12]. Approximation [ABMP16, BGN16a, BGG10, BS15e, CSC14, DD15, II13, JK12, LT10a, LLM12, LSS19, LR15, QLCL16, Sch12c, SS18c, SST14, SSBS14, Vid17, Vie16, Zuc11a]. Approximations [AvB16, BCJP19, BCL10b, GLO10, KTJ10, OO18, Pav11, TFES19]. Aqueous [FD11]. Arak [Thä11]. Arbitrarily [KMTC10]. Arbitrary [Aum15, Dor16, HVW12, KS14, KT17, Mar16, TKK15, YM11]. Archimedean [CS19]. Architecture [MC17a]. Arcs [XZ17]. Arctic [CS16c, CPS19, DG19b, TW17]. Area [CB16, Cha15a, KMB14, KD19]. Argument [BS17]. Arguments [LKR^{+}11]. Arising [CCG14b, HM13b, KV16a]. Arithmetic [OW11]. Arm [DHS18, Wu18]. Arnold [dlLL11]. Arrested [LE17]. Arrhenius [SW11]. Arrival [DT18]. Artificial [Cug17]. ary [DRCV19]. ASEP [BS13a, BS15a, IS11, Lee10, PS11, TW10]. Asexual [PSK10]. Aspects [BGP10, CSV11, DM10, LVE19]. Assembled [Ara11b, DK09]. assembly [Mül11]. Assess [RBM^{+}18]. Assessment [AB17]. Asset [BJM15]. Assets [DPT17]. Assignment [CDS19, PSS11]. Assisted [PS16]. Associated [BK18, DKY19, Fed17, GMT17b, GMT17a, Kat12d, Kra16, MBS16, PPS16, Tid16, WP11a, ZGL13]. Associating [FBE^{+}11]. Association [Dzu11]. Associative [DHL^{+}17, GHLV16]. Asymmetric [AFR19, BW17, BS19, CCG14a, CGRS16, CL16, Chu18, Dai17, DG17b, DLLX16, GS11b, Gon14, GG11a, HM16, Kan14, KL15, KMO16, Lee12, LN19a, LN19b, Mat15, Mor11, Pro15, SS10, SJHW11, Sim11, TW13b, TW13a]. Asymmetrically [Gup16, ST14]. Asymmetry [BCS18, FLS12, Hen12]. Asymptotic [BCPS18, Các14, CLS11a, CFTW15, CTM13, CC14, CL18, CMM14, DHR18, DF16, DMM14, FS14a, FF11a, Fed13, Fre17, He14, JV19, KSSH15, Kol17, Kuo17, Lan16, LH11, Pan14, Pan16, PMC15, PR19, Sam16, Sch12b, Sch13b, TH12, Tou18, TC11a, WYG16, Zhu17]. Asymptotically [BCW13, RS15a]. Asymptotics [BFKP10, BDES19, BCNS12, BL11a, BS16, BP19, CF11, DG15a, DM10, DG15b, EL12, GW15b, Kie17b, KM13, KN16, Kua13, LN19b, MP18a, MNV11, NS12, NZB16, Pro15, ST11c, WX15, WXX16, Web11]. atom [Bru14]. Atomic [Zho17, Zho18]. Atoms [Hog11, KT12, SH12]. Attachment [BG14, DvdHH10, EMO18, GvdHW17, PPS16]. Attacks [Sha18]. Attaining [WHC14]. Attention [Hub13]. Attilio [Mon12a]. Attraction [Mov16, BSW17]. Attractive [HKW11, MPS14]. Attractor [BM18, TLD18]. Attractors [BFT10, GNP18, Lia19]. Aubry [LN15]. Autocatalysis [WGLE11]. Autocatalytic [GS13]. Autocorrelation [MCG12]. Automata [CSAS17, CAS11, CT10, EJ10, KT15, PSS16, RL17, SM15, Tag15, WPB15, dMP12, dMS17a]. Automaton [GG11b]. Autonomous [Ply15]. Autoreactivity [KKC12]. Autoregressive [SP13]. Auxiliary [BSM^{+}16]. Availability [XJZY13]. Avalanche [HTX^{+}12, JRS15]. Avalanches [BDL16]. Average [LR15, OR19, Sod11]. Averages [Abr17, ASA15, FRT15, Fre17, FN15b, GLBP12, LRL17]. Averaging [FK18c, Gao18, Ito19a, MBC14]. Aversion [CA18]. Avoidance [PF17]. Avoiding [BBC15, CS16b, Gil15, Ken15a, Ken15b, SH16, BSW17, Cli10, Cli18, DGK^{+}11, Ghe10, GOdSS11, Ken12, RW11]. Axiom [Luc16]. Axis [Bur17]. Aztec [DG19b, FF11c].

B [EFO11, MvS19]. Bachelier [MOW11]. Back [SCSS18]. Bäcklund [Liu15b]. Backtracking [Klu11, FvdH13]. Backward [Lu12]. Bacterial [BNXL11, CSV11, Man11, RESA10]. Bak [BBGS19, BAS18, MS12b, Sch12b]. Balance [HT11, KJZ17, RCV16, Rue16]. Balancing [DA18]. Ball [BGN^{+}17]. Ballistic [CWD17, CDdS14, EM10, MRR19]. Balls [HY17]. Banach [Bal18, Bal17]. Band [BF10, BYYY19, EH12, HM19, JS17, Mar18, Mie18, Shc14, Shc15, SS16, SS18c, Sod11, TT16, Tan18]. Bandlimited [Fra17b]. Bandwagon [GNPS13]. Banking [BA14]. Barnes [Ost16]. Barotropic [BNT13]. barrier [CCR19]. Barshad [SM12a, SM12b]. Barycentric [CMVW11]. Based [AvB16, BCS18, CFM14, Cha14b, CGN16, DARM^{+}13, DJRZ11, Fil16, FD16, GRT17, HFWT15, Kuo17, LO17, LZHS19, Mas16, PT14a, PF17, Rue17, RF18, SCSS18, XZ19, LLM12]. Basic [DM12, Mih11, Thä11]. Bath [BT12, BR11, CEGW18, Los17, VV17, vHL13]. Baths [Mae14]. Battery [BDL10]. Bautin [CF19b]. Baxter [KS13]. Bayesian [CSC14, dAPdA^{+}13]. BBGKY [CIM14, GS12]. Bead [VV17]. Beautiful [WLL11]. Becker [HY19a, Sun18]. BECs [Buc16]. Before [Ras12]. BEG [LN11]. Behavior [ASA15, AM12, BCKY19, BP12a, BCP11, Can17, CC14, Cha15b, CL18, CMM16, FW15, GOPS11, GLM^{+}15, HK19, HR18, Kuo17, MGAPQH13, Mih11, MiS18, MT11b, PK11, PR19, RMS19, TH12, Wil11, ZWGM15, Zoc18, dCFC11]. Behaviour [BBS14, BCPS18, CG11a, CG15, Fre17, JSV19, LFW12, MS12c, Pim18, RW11, SV15, Thr18, XTpXpH12]. Behind [GdHR18]. Belies [Lee18]. Belonging [YYZ11]. Below [LSW17]. Ben [Leb11]. Ben-Naim [Leb11]. Benard [Bha15, CDS17]. Beneficial [BPR13]. Benjamini [ACH15]. Bernard [Ano15b]. Bernoulli [BCM10, BT11, BW12b, BBW15, CG11a, Kos13, LNT13, TW10, dLPS15]. Bessel [KT11a, KMB14]. Bessis [CDTA10, LS12]. Best [BB11]. Beta [BMSS13, Hay15, Nak14, NT18, Ost16]. Beta-Mixing [Hay15]. Bethe [BS11, Lee11, Mor11, Par17, Pro15, RZ17, Sim11]. Better [AV16]. Betti [GTT19]. Between [AKD19, BHS13, Bar16, BCFS17, BL11b, BCY16, BPRT14, BV15, FF10, FS11a, GGP10, GNPS13, Jia14, Kol14, NP14, RCV16, RC17b, ST14, SS19, Tas18b, Tay16, Tia14, WFK11a, WFK11b, Zho17]. Beyond [LM17, Mas13]. BGK [BCNS12, HHM17, Pir18b, Zha10b]. Bias [HM13a, LP16]. Biased [GMM18, Sta13]. Biasing [FJLS18].

Bidimensional [GPMEA17, GPMEA18]. Bifurcation [BP18, CF19b, FS11b, Li12, Tak16]. Bifurcations [MG19]. Big [TGGS13]. Bigger [SLdEC11]. Bijection [NS10]. Billiard [FZ10, LS17]. Billiards [Art19b, Bun14, CS10c, CZZ13, FZ10, Ste10]. Billion [Cli18]. Binary [AG15a, BLT11, BA19, DMP17, DL17, FCLK14, GGP10, HM16, JM14, KR15, LPS12, LN19b, Moh17, OEA18, PK11, RELV11, RRS12, RG17, YY10]. Binary-Tree [YY10]. Binding [MSS11b, SKT16]. Binomial [CP12, CGR12, KPR18, SSBS14]. Biochemical [FP11, KYA16, MMW16, MLS16, Qia10, ST11b]. Biological [GL13, JP18a, MB11, MV16, Pos16, Sch12b, WLJH18]. Biology [MLS16, Qia10]. Biomolecular [EAL12]. Biorthogonal [Zha15]. Bipartite [ADP14, AC14b, Col14, Fed14, NMV11, SCSS18, del12]. Biquadratic [EKD12]. Birds [HH15a]. Birth [CG16, DDC18, FK18c, HIK^{+}18, LKD12, Ohk14, SV15, ZHRB16]. Birth-and-Death [FK18c, SV15, ZHRB16]. Birthday [Gal17a]. Births [MPTV12]. Bistable [GG11a, Shi16, XJZY13, XLL^{+}15, ZYZ^{+}18]. Black [LWL^{+}12]. Blackwell [BPS12, BK15]. Blends [KT11b]. Blob [HNZ16]. Bloch [Wat20, Wat19]. Block [BDES19, BA19, EPS17, FLTV11, GO11, KMM11, Shc14]. Block-Maxima [FLTV11]. Block-rectangular [GO11]. Blockage [SLM15]. Blocks [Shc14]. Blocs [Lee18]. Blotto [HZ13]. Blow [FG12, Ngu17]. Blow-Up [FG12, Ngu17]. Blowing [HMRW13]. Bluff [FCK15]. Bluff-Bodies [FCK15]. Blume [AB14, ACL14, GV12b, KOT11, LHZ^{+}19, LL16, LLM19, Lim16]. Board [GNP16]. Bodies [FCK15]. Body [BFM10, CL14, FK11, Imb16, Kie14, Koi18, MAPS11, SVRL11, dVO15]. Bohmian [Gal15]. Boiling [PKBW19]. Boltzmann [Sha10, Tak10a, AvB16, Afz12, APZ19, AG11, ACR18, BCNS12, Ber12, BD16b, BJM15, BBR^{+}19, BE16, CG10a, CL19, CE14, CIM14, CR11a, Che13, CLMK18, CH14, CE19, DP17, DFF18, DL12, DL17, Dud13, FL15a, Fow19, FCK15, GdK10, GJS17, HX15b, Han18, He14, HJ17b, HY19b, KNPF19, KT17, Lan17, LWW18, LZ11, Lu12, Lu13, Lu14, MN16a, MM16, MM17a, MCK15, MWY16, NE16, PM17, PB11, ST18a, Sug10, Tak09, Tak10b, Tri14, WZ12, Wu15, Yan11, YH14, ZSHL15]. Boltzmann-Type [BD16b]. Bond [BCL10a, CG11a, CS12, Cla19, OR16, VAY^{+}12, dLPS15]. Bond-Triangular [BCL10a]. Bonds [LMC11]. Bone [LKR^{+}11]. Book [Kie17a]. Boolean [ABLT17, CMM16]. Bootstrap [Ami10, AF14a, BW12a, CG14c, ESPP^{+}14, KN13, STBT10, SH16, TV15b, vEF12]. Bootstrapping [FCLK14, MBC^{+}13]. Borel [CH11]. Boron [HW13]. Bose [ADU17, Ark13, AN15, BC12, CL19, CG14a, Fid15, HT19, KPS19, LZ11, Lu13, Lu14, Lu16, NiS19c, NP12, RSY14, TKK15, TT17b, TW13a, Yin10]. Boson [Ngu17, Ohk10]. Bosonic [Hog11, LR15, SSE15]. Bosons [BE16, Lee19, Mie18, RK12]. Bottleneck [ABR18, KBB19]. Bottlenecks [LW18]. Bound [Bou13b, JL17b, Kur18, LU19, LJN18, MN12, O’C12a, Ste10, TS19, Yin10]. Boundaries [BW17, CG19, EV14, RW19, TM10, TT17a, WZIG14]. Boundary [ABH^{+}17, BN14, BHNY15, BMNS17, BL10b, BGL14, Bis19, BS19, BKP13, CLS11a, CD14a, CL16, Che13, CDS17, CAS11, Coq15, CPRY11, DLS10, Eri18, ELX18, FSV10, FGP15, GRR17, GDMS19, GW15a, Kan14, KC18, Kua13, Kuo15, LH11, MdG13, Mar16, SS11a, ST18a, Sim10, TH12, Yan11]. Boundary-Driven [BKP13]. Boundary-Induced [KC18]. Bounded [BFL18, CTM13, DP19, KD19, SST14, Tak10b, ZYZ^{+}18]. Bounds [ABR18, ANSW14, BKS12, BJS17, CBG14, FS14b, GL16a, IST12, LO18, Mal12, NT16, NC10, Pro17, Tat13, Tem14, TV12, dLP14]. Bow [MGAPQH13]. Bow-Shaped [MGAPQH13]. Bowen [CLP17]. Braking [FLS12]. Branch [Hal17]. Branched [NKK19]. Branching [ADGPP17, BBS11, Ber19, BDL16, BJR10, BD11b, CG13, DS16, FZ12, FC13, HL16, Zuc11a]. Breakdown [BdlL13, CdlL10, MC10, MC11]. Breaking [Aum15, BSS19, Con18, Fer14, GdHR18, Ito17, Ito18, KJZ17, Mor18, Tas19]. Breather [TV15a]. Breiman [BMSS13]. Brézin [Kar11]. Bricklayers [BKS12]. Bricmont [Kie17a]. Bridge [ADC10, Gil15, Tay16]. Bridges [MP18a]. Broadwell [Ily12]. Brownian [BCJ17, BBS11, BBC15, CB16, CWD17, CD14b, DS16, FZ12, FL13, GW12, GDMS19, Git12, GOdSS11, HWZ19, HKW11, JL16, Kat12b, Kat12d, Kat12c, Kim12, KK15, KLACTL19, LPK13, LM13, Lia18, Lie12, LWL^{+}18, MP18a, Mol17, Mol18, Ngu19a, OW11, PP14, PCM15, Ply15, RS13, Tay17, TV12]. Bruijn [Osi16]. Bubble [LE17]. Bulk [BKLL12, NT18, NP12, NP14, RRW11, SW12, VCT11]. Bundle [HTX^{+}12, RG18]. Burgers’ [Liu15b, Abr13, BL18, DXZ14, ED15, MS10, Mol17]. Busch [Ano12a, Ano12b, Ano18a, Ano19a, Ano19b, Leb12a, Leb12c, Leb13b, Leb13c, Leb14]. Business [Sta11, AFI^{+}10]. Butterflies [Lep15]. Butterfly [AEG14].

Ca [EFO11]. Cahn [RBS16b]. Calculation [Bax11, Bax12, BGN16b, KBSM16]. Calibration [OW11]. Callan [Ano19b]. Calogero [AKD19]. Cambered [PM17]. Campus [Ano12a, Ano12b, Ano18a, Ano19a, Ano19b, Leb12a, Leb12c, Leb13b, Leb13c, Leb14]. Can [DEF12, FN13, Hol11, HBC^{+}15]. Candidates [PT14a]. Canonical [CO17, KJZ18, KM18, KM19c, MP18b, MP13, Pav11, PT15, TKK15, Tas18b]. Capability [KNPF19]. Capacity [AM19, DHL^{+}17, DDC18]. Capel [AB14, GV12b, KOT11, LHZ^{+}19, LL16, LLM19]. Capillary [CP10b]. Capital [HH15b]. Capturing [NE16]. Carbide [HW13]. Cards [Hui17]. Cardy [BCL10b]. Caretakers [NBB13]. Cargo [Gol10, KBLL13]. Carlo [BBD^{+}11, CF19a, Cli18, HG11, Koi10, LL13, ME11, MMSY11, UK16]. Carpet [Mis19]. Carrying [DDC18]. Cascade [FG14, HHT10, JLLP17]. Cascades [ES12, TGGS13]. Case [BL12, BHJ^{+}12, BDY17, BS19, FAB16, FDR12, PWZ16, SS15]. Cases [Pos16]. Casimir [CSC11, NP12, NP14, TT17b]. Catastrophe [DPT17]. Catastrophes [DD15]. Cauchy [AG11, BE16, WZ12, Wei18, WWKK16, Zha10b]. Causal [DJW10, MC17a, NT16, SYZ13]. Caused [TAG10]. Causes [GPGA17]. Cavity [GSSV11]. Cayley [AMS14, ALS18, ART11, EHR12, GRS12, GRRR13, GHRR13, GRR17, KRK14, MBS16, Pah10, Pat17, RVY18]. Celebration [Zia10]. Cell [GNP16, KKC12, KD19, LH17, RELV11, SK19, ZW10]. Cell-Board [GNP16]. Cells [FMAG11, FAB16, Kos11, RELV11]. Cellular [AKH13, CSAS17, CAS11, CT10, EJ10, GG11b, KT15, PSS16, Qia10, RL17, SM15, Tag15, WPB15, dMP12, dMS17a, tWBOM16]. Center [Ano12a, Ano12b, Ano18a, Ano19a, Ano19b, Leb12a, Leb12c, Leb13b, Leb13c, Leb14]. Centered [Asa13]. Central [BPZ13, BSS14, CCEF10, FL14, GGvdHP15, HNVZ13, Hor16, Kle13, KKSY19, LLM12, NSV12, PB11, RVY18, Sun18, TWT14, Tos16, Tzi18, Cha19]. Centred [JD11]. Century [Bak10]. Cercignani [Ein12]. Certain [BCL10a, Kac13, LNY16, LY10, PCM15, Ras12]. CFT [GLM18]. Chaikin [Ano18a]. Chain [AS16, BB11, BFKP10, BO11, DDN14, DKLS19, DPT17, GL16a, IPS^{+}10b, KMS19a, KT12, LSY19, LMN16, MMA15, QD12, SSR12, DS19a]. Chains [AMS14, ADF18, BL10a, BL12, BCFS17, BM11, BLM13, BK11, CG12a, DDHS17, DKY19, FP11, FKR12, Fre14, FK17, GGP10, GL13, GM15, Gas16, GPP12, Hag13, HS14a, HIK^{+}18, Imb16, KJZ17, KJZ18, Li19, LNT13, LN15, Luk14, MP18b, MM14, MS17, MDP^{+}18, MBS16, RY12, Spo14]. Chalker [ABJ12]. Challenges [Bou13a, DRS10]. Chandler [Fan17a]. Change [Che14, DG15b, HM13a, LRL17, OMC11]. Changed [D’O14b]. Changes [BC16]. Channel [AM19, BBR^{+}19, Bur11, HNZ16, KK15, HA13, RSB10, SSB15]. Channels [FH13, KU11, LS15b, LN19b]. Chaos [ALS18, AK17, ADU17, BCE^{+}14, CMW15, CHHS15, CGY17, CP17b, CRV13, Cor16, GG11b, Kau11, Kie13, Kur18, Lep15, Mal12, MAPS11, Sin10a, Kla11]. Chaotic [GMT17a, LK14, LK18, LFWK14, Naz18, TLD18, Ven14]. Chapman [CLMK18, GPMSBBSV15]. Characteristic [Afa16, Afa19, FN15b, FL16b, Sha12, TA16]. Characteristics [ADH12, PK11]. Characterization [GR12, PR15a, dHO13]. Chargaff [HMO12]. Charge [AF16, BGLL13, CR14, CR11b, DLLX16, Miy16a, SJ10, Sam19, TT12, TS19, WBL11, Zho17]. Charge-Symmetric [TS19]. Charged [Ban10, BG12, CF11, FW12, HKW11, RW19, ST14, Sam15, Sam16, Vie16]. Charges [Sam13]. Charles [Mac13]. Chayes [Ano18a, GOPS11]. Cheap [LTR17]. Chebyshev [Tak16]. Chemical [ADE18, Bai10, BD15b, GQ17, Gor18, LL16, LLM19, Qia10]. Chemically [SB15a]. Chemistry [Bai10]. Chemotactic [HCLR11, RESA10]. Chemotaxis [CSV11, Man11]. Chiral [IPS^{+}10b]. Chirality [FY16]. Choice [OEA18]. Cholesteric [SGL15]. Chopping [CP18]. Chordal [HMN15, Zha10a]. Chromatin [CM11]. Circle [BHJ^{+}12, DMY13, Fer14, SB16]. Circuit [Ham11]. Circulations [KW12]. Circumstances [Hol11]. Citations [GS13]. City [LLS13]. City-Size [LLS13]. Class [ABF16, BFKR10, BCHM12, BS15a, BD15a, CS10b, CT10, CFG13, CQR15, CH11, GN19, HKR17, HKR23a, He19, LS15b, LM17, LW19, LMC19, MW12a, MU13, QS15, Sco11, Ven14]. Classes [HSFK18, LSBS13, PSS15, RMS19]. Classical [AKD19, AF16, AG11, AF14b, BDDH14, BM16, CH14, Cra11, CMO18, Dal11, DKLS19, FBE^{+}11, FS11a, FR17, GS17a, Gou15, KC19, LTM16, LLS17, LC19, MSS15, Mon12b, RRW11, Sch13c, Wre12]. Classification [DF17, GJMS10, OC12c]. Clausius [MN14a, ST11a]. Clifford [Thä11]. Climate [LRL17]. Climatology [TW17]. Climbing [BLL^{+}13]. Clique [GSSV11, TVP13]. Cliques [JvLS19, PT14a]. Clock [KS18, MS11a]. Clock- [MS11a]. Close [AN15, BvE11, CLTT13, GBTL17, NV14, PKDK13]. Close-packed [BvE11]. Closed [LH11]. Closure [AvB16, KTJ10, vdHvLS18]. Clouds [AG14]. CLT [Ste10]. Clumping [BKM15]. Cluster [Alb16, BFP10, Cam13, CVE14, CG15, CDCL18, CM12b, CEGW18, FBR19, Fed17, FR17, FD16, Gan18, GOPS11, Ham11, JK12, Jan15b, KiMM13, MW10, Mis15, Mis16, Mis19, RLCMRT10, vdHKvL18]. Clustered [RF18]. Clustering [AKL18, FL18, GO13, MS10, SBK10b, dVO15, FL20]. Clusterings [Sha12]. Clusters [AM10, Car11, ELO11, FL16a, HvdHH14, PSW17, RT17b, Xue16b]. CMV [Ong14]. Coagulation [Aza11, Lau18, MNV11, NV14, NTV16, Thr18]. Coalescence [CRTZ13, DS16]. Coalescent [BCL19, dMS17a]. Coalescing [BCL19]. Coarse [Ark10, BHS13, BTV14, Mül11]. Coarse-Grained [Ark10, BHS13, Mül11]. Coarsening [DEK^{+}15, IPP14, NC10, Sme18]. Cocktail [HHT15]. Cocycles [BGLZ19]. Coddington [ABJ12]. Coefficient [YH14]. Coefficients [BGN16a, BM12a, Coh09, Coh10, DMY13, Gal17b, HHV16, JPV18]. Coexistence [AT18, BNTT16, EL12, LN11, Car11]. Cohen [BCHM12, For11, FP11]. Cohen-Type [FP11]. Coherent [AP11, BBH11, BGTV11]. Cohesion [LNP13a]. Coin [Yam13]. Coin-Dividing [Yam13]. Coincidences [KLACTL19]. Coins [XY13]. Collaborative [YK13]. Collapse [BDG^{+}14, KM19a, MRCJ18, VCT11]. Collective [Bou13a, CDG^{+}15, Cha15b, FW15, Gup16, MMST13, MG17, MPM17, OV15, PSVG18]. Collective-Density [MMST13]. Colliding [CLT19]. collinear [SSR12]. Collision [BV16, Ber18, BG12, CE14, Moh17, PF17]. Collisional [LZ10a, Tri17]. Collisions [BLT11, HKR16, He14, HR18, Huv12, LSY18, MCK15, Tak15]. Colloids [DOGK16, San18, SGL15]. Colonel [HZ13]. Colonies [BNXL11]. Colonization [MRCJ18]. Color [FL18, FL20]. Colored [AEG14, GG11a, HMW19, LL19, NML^{+}11, ZW10]. Colored-Noise [GG11a]. Colorings [CS10a, SX10a, SX10b]. Colour [dCCS19]. Colouring [Tho11]. Column [CCFR18, SLM12]. Columnar [GHS17, NKR15]. Comb [Cla13].

Combination [dlLOP11]. Combinatorial [BS11, Bar12, LO18]. Combinatorics [BK18]. Combined [BH11]. Combining [OMC11]. Comment [MGZ14, Tho12]. Commentary [MSV13]. Comments [Sha10, Tak10a]. Commitment [Mob13]. Committee [Nei12]. Common [CLTC15]. Communication [PKDK13, dHNT11]. Communities [AFS^{+}13, ARS17b, GPSB18, KP18]. Community [FW17, LN19a, Mon15]. Compact [CMO18, ELO11, DMS12]. Compacting [SLM12]. Companies [Sta11, AFI^{+}10]. Comparative [GHLV16]. Comparing [LNP11]. Comparison [Ark10, Kra16, RvH14]. Compensation [DK10, KE10]. Competing [AEWI14, FL12]. Competition [BPF^{+}14, FAM13, Hal17]. Competitions [BNHRV13, Rad17]. Competitive [LYZ11]. Complete [ABFP15, GR12, XP17]. Complex [BAC13, BNTT16, BVL16, BCF10, BDY17, BLS17, CVE14, For13, FW17, GvdHdHM18, HHT10, Kat12d, Kie17b, LBW^{+}13, LT10b, Mac10, MBC^{+}13, PPK11, Rab11, SZ18, Sta11, TWT14, Tho12, AFI^{+}10, For11]. Complexity [AC14b, Bon15, HBC^{+}15, KBLL13, MB18, WLL11, SN13, Mac13]. Component [AF16, AN15, BBS14, CFTW15, CP10b, CMV16, Fan17a, GH16a, GH16b, GT15b, PSS15, ST16a, Sam13, Sam18, TF12, TT12, TT17b, Kel19]. Components [EMO18, Yag16]. Composed [Hon19]. Composite [FK18b, LWL^{+}12]. Composition [Wil11]. Compositions [OP12]. Compound [CNZ17]. Compounds [Wid17]. Compressible [ABH^{+}17, GM13]. Compressional [CL13a]. Compton [CE19]. Computable [RW12]. Computation [Aur18, CdlL10, JPV18, Lan11, ST16a, ZL13]. Computational [Ano11d, Ano11e, BC15b, LVE19]. Computations [DOGKP19]. Computer [GT10, PPK11, San13]. Computing [Cam13, CAG^{+}13, GT15b, Hub13, RBS16b]. Concentration [BC16, CCR17, Krü12, MC11]. Concerning [Cha12]. condensable [Ber12]. Condensates [Ark13, AN15, TT17b]. Condensation [ADU17, CCG14a, CG10b, CG14b, Fid15, GRV11, HT19, JS11, KPS19, Lu13, Lu14, Lu16, Mor11, NP12, TKK15]. Condensed [CG14a]. Condensing [Ber12, DSZ17, Sta15]. Condition [DLS10, Kuo15, Lee10, Pic10, TW10]. Conditional [EL12, HL14, LS15a, MS19b, OEA18, TV15a]. Conditioned [APdM^{+}18, BS13a, DS19b, PS11, DS19a]. Conditioning [Fun14]. Conditions [ABH^{+}17, BN18, CLS11a, CTM13, CD14a, Coq15, CPRY11, GRR17, Klu11, Kor16, KOSV18, Lan15, PG10, SS11a, ST18a, Tem14]. Conductance [BLRVR13, GRT17, KMS19b, Sep13]. Conductance-Based [GRT17]. Conductances [FGN14]. Conducting [KNiST11, ST11a]. Conduction [DDN14, DI13, DG17b, LY13, MS11b, NiS19a]. Conductivity [ILOS10, Mas16]. Conductors [CS16b]. Cone [BBGS19, BDV12]. Conference [Ano12a, Ano12b, Ano18a, Ano19a, Ano19b, Leb12a, Leb12c, Leb13b, Leb13c, Leb14, Ano15a, Ano16a, Leb10]. Confidence [Bal14, CTM13]. Configuration [AK18a, BDL11, DO18, SB15b, vdHvLS18].

Configurations [ADS^{+}19, CGS15, CNS15, CCFR18, NBK14, dlLSZ16, dHNT11]. Confined [AKH13, BHF^{+}12, BPDH10, Bur11, HNZ16, KK15, KT11b, Mar11a, MdSB18, Per10, SdlPRA16, SK19, TT17b, VCT11]. Confinement [CE14]. Confining [DL12, AD15a]. Conformal [BCL10a, CTH^{+}11, DTW19, ESPP^{+}14, Ken15a, Ken19a, MU16, PM17, SH16]. Congestion [DNBS10]. Conjecture [Bon15, CDTA10, DD10, Ein12, Lia13, LS12, PC19, WLEC17]. Conjectures [AT12, DRS10, dCCS19]. Conjugacy [GXL12]. Connected [HBB10, Lan17, LR18, LR19, LMN18]. Connecting [BD15b, WL13]. Connection [FF10]. Connections [AKD19, Ber19, CDV17, TV15b]. Connectivities [CdLS13]. Connectivity [BH11, CDG12, Det18, GGD16, Koi10, KD19, MQW18, Tou14, UK16]. Consciousness [Hep18]. Consensus [APZ19, Lee18, WTM19, LBW^{+}13]. Consequences [BC19]. Conservation [BBC18, GS11c, GW15b]. Conservative [BO14, BFNZ11, HHM17, ILOS10]. Conserved [FSV10, Sim10, SS15]. Conserving [BD16a, GDL10, Lau18, RMN15]. Considering [HMRW13]. Consistency [TP19]. Consistent [Bha15, MOT14a, Sch10, KT12]. Constant [Bot18, MTVU18, Mot14b, NV14, Yuh15]. Constant-Length [MTVU18]. Constants [CS19, GP10, GLBP12]. Constrained [BLT12, CM11, KRRS17, Mob13, Mou15, OiS14, STBT10, WK18, Zhu17]. Constraint [HMU13]. Constraints [Lan13]. Constructing [CLP17]. Construction [BBH11, GN19, Tid16]. Constructive [BMR10]. Consumption [DT19, MLS16]. Contact [BBGS19, BSW17, CGL17, CGL18, CMS10, KPZ16, LH13a, SdlPRA16, Sch13a, Tzi13, Xue16a, Xue16b, XP17, vHL13]. Contacts [FL16a]. Contagion [BRSW15, PS19]. Contagious [GS15]. Content [LKR^{+}11]. Continued [OP12]. Continuity [ABT^{+}14, AT12, BK15, BC18]. Continuous [AM19, AGJP19, BN14, BL10a, BL12, BK17a, BC11, BN15, CCG14b, CTM13, CMM14, DD15, DMM14, FKK10, GI19, HTX^{+}12, Kar10, KPZ16, LNT13, LN15, LH13b, MC17a, MC17b, MN14b, MP13, MGMMP13, Ong14, OO18, PPS16, Pel14, RM16b, SWB10, TM10, WZL^{+}14]. Continuous-Opinion [CTM13]. Continuous-Spin [BC11]. Continuous-Time [DD15, Kar10, MC17b, MGMMP13, MC17a]. Continuum [ANS18, AKQ14, BG17a, BCJP19, BKK15, Bla10, BM12b, CP16, DH19, FPR11, FK18c, Hel16]. Contours [VFT12]. Contracting [GNP18, PT10]. Contractions [BMT15]. Contractivity [CCH^{+}14, Hau16]. Contrast [PZ17]. Contribute [HBC^{+}15]. Contributions [SvHM^{+}11]. Control [DM11, HH15a, HHT10, KR16, KU11]. Controllability [BT19]. Controlled [MGMMP13, ZGL13]. Contucci [Gue13]. Convection [CDS17]. Convective [Bha15, BBR^{+}19]. Convergence [ACH15, AKM13, AF14b, BJ16, BT19, BCL10a, BCL10b, BFP10, BMR10, BM12b, CL19, CLS11b, CM12b, DHX19, Els12, FLTV11, FRT15, FJLS18, GR12, HY19a, KJZ17, KMS19a, Lee19, LNP13b, LX17, LT10b, Lu16, Luc16, Mar18, PR15a, Pro17, RL17, RBS16a, Tag15, Tri14, Tzi13, Xue16a, YB14, dLP14]. Convex [GT16, AMT18, CFG13]. Convexity [KM18]. Cooperation [CAG^{+}13]. Cooperative [BC16]. Coordinate [DG19a]. Coordinates [D’O14b]. Copepod [FW15]. coplanar [FBR19]. Copolymer [dHO13]. Copula [Naz18]. Copy [GV12a]. Copying [PS16]. Core [AP14b, HL18, Mie18, MPS14, NZB16]. Corner [YM11]. Corners [AB18a, CDG12]. Cornuéjols [Mon12a]. Correct [MM16, MM17a]. Corrected [DvLM18]. Correction [Bal18, BS20, CLTC23, CGL18, CL16, FT18b, FL20, GPMEA18, HKR23a, HR23, HKR23b, Hal19a, Ken16, MC21, Pro17, Sam18, SS18b, TN21, Wat20]. Corrections [CPV10, DR13, DM18b, FF11b, KSH11, LPK13, PT15]. Correlated [AEK16, ADGPP17, AB18b, Ber14, BCJP19, DD10, FL16b, Leh13, LNS^{+}12b, LMC11, MK19, O’C12a, RdAB18, Web11, WWKK16, dWL14]. Correlation [Afa16, Afa19, AVW17, Bao17, BLU16, BL15, Gan18, GRV10, HMW19, Lan13, Lim16, LMN18, Mat12, NP12, RKGZ12, Sch10, Sim14, Ste10, VW18]. Correlations [AF16, Bab12, BMC17, CT10, GNP18, GDL10, HH15a, HFWT15, IS13, IL11, JV19, Kat15, KM18, KM19c, Los17, MN14b, MG17, NP14, PPS11, PT15, Sam16, WFK11a, WCX^{+}11, dMP12, WFK11b]. Correlators [AW18]. Correspondence [Yam14]. Corresponding [DWTW16]. Corrupted [Fyo19]. Cosmic [RPPF15]. Cosmic-Ray [RPPF15]. Cost [CDS19]. Couette [RM11]. Coulomb [AB19, CP14, CF19a, CFLV19, DP14, FK11, Leb16, PRSS17, Sam16, Sam17, Ser14]. Countable [AFFR17, Kem11, MU15]. Counter [Ein12, ST14, Sam15]. Counter-Ions [ST14, Sam15]. Counterparty [BA14]. Counting [BGL10, FDR12, JL17a, JvLS19, Pah10, Rue18]. Coupled [BLV14, Dor16, Eva16, FBR19, Fer14, FSS13, Gup16, HMU13, LMN18, MM13, MG19, SJ10, SWB10, dMP12]. Coupling [BGTV11, CEGW18, EKD12, GGP10, GLO10, KOT11, KO15, OR15, ST16a, SB16, SJHW11]. Couplings [AS16, MM17b]. Course [Täu10, Rei09]. Court [LBB15, Lee18]. Covariance [GdHR18, MM13, MR13, ST11c]. Covariances [BM12b]. Covariant [WW16]. Coverage [GG18]. Coverages [SM12b]. Covers [AVW17]. Creation [BE16, GW15b]. Credit [DPT17, PS19]. Creep [AGGL^{+}16]. Crises [Bou13a]. Crisis [TLD18]. Cristian [Gue13]. Criteria [Tau11]. Criterion [Lep12, MN15]. Critical [Afz12, AB14, AP12, AB17, ACL14, Bab12, Bai10, Bak10, BCM10, Bal14, BFKP10, BBS14, BBS11, BBLP12, BU18, BPRT14, Can17, CG14a, Cen13, Chh12, CS16b, CNS15, CPRY11, CG13, DG14, EMO18, ESPP^{+}14, Fed14, FN15a, Fil16, FMM^{+}15, FPR11, GOPS11, GLM^{+}15, GLML16, GLM^{+}16, GL16b, JV19, JJ12, KLM13, Ker10, KPZ16, KY12, LU19, Li12, LSW17, Mac10, NiS19c, NT16, Neu14, RMS19, Sta15, Tag15, TKK15, TN13, TVP13, WZIG14, Wu18, XTL14, Xue12, Yin13, dHNT12, vdHKvL18, DSZ17, TGGS13]. Criticality [ADS^{+}19, BBS^{+}15b, CRS14, CGL17, CGL18, CDS10, JRS15, MS12b, MB11]. Crooks [WES11]. Crossing [DRCV19, JZ10, Kos13, RMN15, GDMS19]. Crossing-Twisted [JZ10]. Crossover [AB19, BCJ15, BPRT14, GLM^{+}15, GLML16, GLM^{+}16, SS17, SS18b, SS10, dCFC11]. Crow [AP11]. Crowd [CAG^{+}13, DARM^{+}13]. Crowded [Gol10]. Crystal [Alb16, Bae11, FY16, GV12b, KKSY19, Lan19, TS12]. Crystalline [BCY16, DG15c, Wid17]. Crystals [BG17a, BM16, KWZ14, KK16, Shl17, SGL15, Wen12]. Cube [BW12a]. Cubic [Asa13, BDY17, CE12, JD11, MPSS19, MD10, WGLE11]. Culture [SK19]. Cumulant [Art19a]. Cumulative [PSCD13]. Curie [AT18, Col14, CF19b, EK10, Fed14, GNS18, GV15, LM12b, OK14, OEA18, PFR13, SZ18]. Curious [Kie13]. Curl [TV12]. Current [BS13a, BL10c, CZZ13, CGP17, CDP17, DM18a, GS11c, IS11, KMKT11, Lee12, MPTV11, MS17, PPS11, SJ10, SLM15, SS19, Sim11]. Currents [CCGT10, HEdPG14, KJZ18, MMM15]. Curtis [Ano19b]. Curvature [CE14, CFS18, DM11, Els15, FT19b, HT15]. Curve [CPS19, DG19b, LPS12]. Curves [CS16c, DG15c]. Cusp [BKP13]. Customer [FM18]. Cut [BDY17, BJR10, Fer18, HJ17b, Jin18, Tri14]. Cut-Off [Fer18, Jin18, Tri14, HJ17b]. Cutoff [LNS12a, MWY16, FL15a]. Cycle [AFGL15, Cha14a, FP11, FD16, GXL12, GLU12, Ker10]. Cycle-Based [FD16]. Cycles [AT18, Cam13, JvLS19, LL10]. Cyclic [FL18, FL20]. Cyclotron [Ban10]. Cylindrical [HNZ16]. Cyril [Fis11, Fis12].

d [ADS^{+}19, BLS17, KM19c, Ny13, PSS16, SS18c, AB19, BF12, BCF19, BNT13, CAS11, Cor16, DSS15, ESPP^{+}14, GDL10, Hau16, Her13, Ily12, JV19, Kar18, KT12, Lan10, Li12, LZ11, MN14b, Ost16, Pal11, RT14, RBS16b, RY12, ST14, Sch13b, SH16, SWKS14, Sug10, TJ15, WPB15, dOP18, dWL10, Nak17]. d-Lattice [KM19c]. Damage [HTX^{+}12]. Damp [ELO11]. Damped [BCFS17]. Damping [CF15, CLL18, LWL^{+}18, Tri17]. Daniel [Ano19a, Mac13]. Darwinian [BBK17]. Data [Abr13, ART15, Cro12, Mol17, RBM^{+}18, SWK^{+}18]. Data-Driven [SWK^{+}18]. David [Ano19a, Gal17a]. Deactivation [TSS13]. Death [CG16, DM18a, DDC18, FK18c, HIK^{+}18, LKD12, Ohk14, OPS10, SV15, Sta11, ZHRB16, AFI^{+}10]. Death-Immigration [DDC18]. Deaths [MPTV12]. Debts [LR19]. Decay [CJN18, CLL19, CT10, FL15a, GNP18, KM18, KM19c, LWY18, MN14b, PT15, SM14, TA12, dMP12]. Decaying [CLP19, GL16a, TJ15]. December [Ano12a, Ano19a, Leb13b]. Decidability [BCHW19]. Decimated [Ny13]. Decision [FAB16, TLC13]. Decoherence [BdSPMS14, GS17a, GS19, XY13]. Decomposition [ADC10, BBS^{+}15b, Mit16, Mit17a, Mit17b]. Decompositions [FP11]. Deconstructing [RRW11]. Decorrelation [Shi15]. Decoupling [PR15b, dBGP19]. Decrease [MP14]. Decreasing [GPS13]. Dedicated [Ano18d, Leb19]. Deep [FW17, LTR17, NC10]. Default [DPT17]. Defect [IvRM15]. Defects [Aum15, DG19b]. Defense [HHT10]. Definitions [HS14b, LS14b]. Deformed [GN13, Han15, HDP17, Shc11]. Degeneracy [DLY18]. Degenerate [BS13b, BS20, CA19, LKD12, Yin16]. Degradation [NVL11]. Degree [GVJ^{+}18, Kie17b, LN19b, Sha12, SC10, SST14, vdHvLS18, vdHLK18]. Degrees [BvdHK19, DGGvdH18, OR19, XY13]. Delaunay [AE16]. Delay [AGO18, BGTV11, Các14, DG19a, Hol11, HMW19, KKV^{+}11]. Delay-Coordinate [DG19a]. Delayed [GG11a, LK19, MMW16, MG19, Vid15]. Delayed-Time [GG11a]. Delays [GMT17a, Tou12]. Delbrück [KL15]. Delocalized [BYYY19]. Delone [HHLP17]. Demand [GNPS13]. Deme [HCI14]. Demise [GPGA17]. Democratic [Gal13]. Demographic [BML12]. Demonstration [DP17]. Dense [Coh09, Coh10, Els12, HNT18, KLS19, Klu11, SSB15, SK19, vEdG11]. Densities [GLBP12, Tos16]. Density [AZ11, BBC18, BDG^{+}14, CFTW15, Coh09, Coh10, CE19, DHR18, DDHS17, DR13, DJRZ11, DOGK16, EF13, FBE^{+}11, FG12, GS17a, Ito19a, Ito19b, JK12, KBSM16, KKN12, KNPF19, Kol17, KS14, Kum19, Mac10, Man11, MMST13, Moh17, MFLA15, Nak17, PVCG11, Sam19, Sch12c, SS16, Xue12]. Density-Dependent [Man11]. Dependence [AG12b, CO17, CRL15, Lee19, PZ15, Shi13, WBL11]. Dependency [AI12, BH11]. Dependent [AF14c, AN15, BM12a, Car11, Cro12, FC13, FK11, GC17, HMW19, JP18b, KKA19, LQR12, MGZ14, Man11, MNS12, NSV12, Pic10, RESA10, TH12, TS12, ZLL13, ZGL15]. Depending [ZL19]. Dephasing [BCM16]. Depinning [DR14]. Deposition [CRTZ13, FMAG11, MRR19, MS12c]. Depth [RBM^{+}18]. Derivation [BC12, DF16, FK11, Gri19, IiS15a, JP18b, JLMG11, KSSH15, MN16a, Pic10, RS13]. Derivative [Zha12c]. Derivatives [RESA10]. Derived [Han16, Sug10]. Deriving [KLMP18, Tur13]. Derrida [HS18]. Desai [Ara11b]. Descendants [Fra17a]. Descent [BW18]. Described [Pav11]. Describing [BCJ15, MiS18, MS11b, SWK^{+}18, TFES19]. Description [BCKY19, FdHM14, GT12, GT15a, Kha19, KRK14, MPL^{+}16, MOT14a, NWL19, Ny13, Wu15]. Descriptions [MBGK12]. Descriptors [LZHS19]. Detailed [KJZ17, Rue16, Yin16]. Detecting [BC16, FW17]. Detection [BTV14, RF18]. Determinant [CS10b]. Determinantal [FL14, IS16, Kat12a, Kat15, Lam19, OO18]. Determinants [IL11]. Determination [AB10, GLML16]. Determine [DRT18, DPT17]. Deterministic [AK17, Bec10, Coq15, LSS19, MBGK12, Mou15, Naz18, WTM19, WG18, dSRT15]. Deterministically [Art19a]. Developed [KBB19]. Development [DM12, KKC12, Wei13]. Developments [BB15, FMMPS18, Har11, SL12, ZDS11]. Deviation [DR13, Fre15, IiS15a, LS15a, MK19, Sch13b, Shi13]. Deviations [ADE18, AGO18, BC15a, BG11, BBC18, BDT17, BL10c, BLZ14, BGTVE16, CT13, CFL17, CF15, CLL18, CGP17, CG17, CMV16, DS16, DS19b, DGGvdH18, FSV10, FZ11b, FT18a, FT18b, FD14, Jan15a, KO15, Kra16, LNT13, LN15, LM12b, Luç17, OR19, Pat17, RW12, Ren18, Var12, DS19a]. Devoted [GvdHdHM18]. Diagnosing [BG19]. Diagonal [NS10, Mia11]. Diagram [AM14, ACL^{+}11, AFR19, BNTT16, BLS11, CRTZ13, IvRM15, dHO13]. Diagrams [AT12, EKD12, Fun14, LHZ^{+}19, RMN15, SM15, VB11, ZST18]. Diameter [Cha12, Mis15, Mis19, Moh17]. Diameters [DvdHH10]. Diamond [FF11c]. Dice [LYT16]. Dichotomies [BFVZ10]. Dichotomous [GS11b, GG11a, LNS^{+}12b, XJZY13]. Dicing [MSS^{+}11a]. Dictates [Dzu11]. Difference [Sug10]. Different [CEL^{+}18]. Differentiability [BC19, dlLL11]. Differential [GHMR17, HVW12, HMW19, LLL17, SCY^{+}12, WXX16]. Diffraction [AI12, BW10, BvE11, BKM15, KMSS19, Mol14, ST16b, Ter13, YM11]. Diffusing [AKH13]. Diffusion [AG12a, Afz12, ACCG19, Ark13, BGN16a, BNP14, BO14, BGL^{+}11, BD16a, BGJ^{+}15b, BP11b, BL15, BL17, BM12a, BKLO19, CTB10, CWD17, CPSV10, CLT19, Det12, DD15, FGP15, FH13, FK18b, GdK10, GN19, GTT14, GPD16, GW15a, HHLP17, HM19, HVW12, HS10, JPV18, Kos11, KMS15, Lee12, Lef13, LNP13b, LK18, LWL^{+}18, LLS13, LQR12, Man11, Mar18, MBWC16, MR12, PGS12, RESA10, RPPF15, RNDS13, RDNS15, RBGV12, Sor18, Sug10, TFES19, TV12, Vid17, WGLE11, WHC14, Zha10c, MGZ14]. Diffusionless [MW10]. Diffusions [ABT10, ART15, AGO18, BGN16b, CF16, CGN16, Cro12, DEF12, GW15b, Luç17, OR19]. Diffusive [ABC10, BF10, BN18, BDL10, BKP13, CDdS14, HR18, HEdPG14, LWL^{+}12, OC12c, OVC14, PSS15, Tel10, Vau10]. Diffusivity [SvHM^{+}11]. Dilatancy [AR11a]. Dilemma [CA18]. Dilute [BP12b, BCPS18, GPMEA17, GPMEA18, GNS18, KCB13, PSAPR12, Yin10]. Diluted [KL19, LO18, Pan14, Pan16, VAY^{+}12]. Dimension [ALAF18, BRSW15, BBC15, CTT11, CP15b, DJW10, Far15, FSS13, FL18, KS12, LSY18, LSW17, MU18, MSV13, Mol17, MHD17, PS18, PC19, RS14, SLdEC11, Shi15, SS15, SZ12, Tzi13, FL20]. Dimensional [AB14, BELP15, BELP18, BJ16, BBS14, BSW17, BCLL16, BW12b, BNY16, Bis19, Bot18, BLZ14, BD15b, BCM12, CG10a, CL14, CDS19, CNZ17, Car11, CMV11, CD14a, CG14a, CTH^{+}11, CL18, CL15, Cli18, CMS13, CLP19, CFG13, CCFR18, CMV16, DHS18, DP15, DF16, DR13, Ein12, Fan16a, Fan16b, Fan17b, FS17, FRT15, FLS12, GG18, GMM18, GMM19, GTT14, Gra13, GG11b, GT15b, HI18, Har11, Hel16, HvdHH14, HTZ12, HA19, HCO15, JP18b, KP12, Ken15a, Kir12, Kol17, KM19a, KNSS18, KT17, KM19b, LL16, LLM19, LS17, Miy12b, MS19c, MFLA15, MS16, Nak14, Ngu18, ORW15, PF17, PR15a, Pos16, PB11, QLCL16, QS15, RNDS13, RDNS15, RBGV15, SML19, Sam17, Sam19, SS17, Tas18a, TF12, TT12, TM18, TW14, Vau10, VB11, VFT12]. Dimensional [WX15, WG18, WZIG14, Wu14, XTpXpH12, Xue16b, You17, dCCS19, vEKRS19, ASA15, BBBP11, BL19, CMP17, LLJH10, LSBS13, MSV10, SS18b]. Dimensions [AP12, BHNY15, Bal14, BELP16, CFTW15, Cha15a, GOPS11, HMN15, Leh13, MW12a, MGAPQH13, OCM15, Sak18, SH16, Sou18, Tzi18, Yam17, dWL14, vEF12]. Dimer [ACH15, AVW17, ACM15, AM18, AS16, BEP18, Chh12, DG15b, FF11a, Fed13, GJL16, Per17]. Dimers [BvE11, YYZ11]. Dinner [Gin14a, Gin14b]. Diophantine [dlLL11]. Dipolar [KL11, LK11]. Dirac [BC12, Cla13, PdOC17, RT14]. Direct [FD16, RM16a]. Directed [AKQ14, ALAF18, BL18, Cal15, CC14, CS19, Cla19, CFN15, ELO11, IS16, KMTC10, MT17, Sch12a, SB15b, Thi16, Wei16, Wei18, XZ17]. Direction [PRD11]. Directional [MMST13, MR12]. Directions [Gre12, PRD12]. Dirichlet [BBP17, Caë10, Caë11, CPRY11, GLU12]. Disagreement [SLST13]. Disc [CPRY11, Mar16, Sam15, Sch13b]. Discharge [KMKT11]. Discontinuity [KC18]. Discontinuous [Cal15]. Discovery [MSS11b]. Discrete [AM19, AY10, Ast12, Ast13, BBGS19, BK17a, Ber12, Ber15, BV16, Ber18, BCL10b, BW12b, Bla10, BP15, CCG14b, CFM14, CG19, DDF15, FS14a, GGJR14, GG18, Ghe10, HIK^{+}18, Kua13, Lie12, MC17b, MR13, OO18, PPS16, PdOC17, RM16b, Shi15, WQ10, WGLE11, Zuc11a, dOP18, vEdG11]. Discrete-Event [MC17b]. Discrete-Time [CFM14, HIK^{+}18, Kua13, Zuc11a]. Discretization [RM16b, ST18a]. Discrimination [FAB16]. Discussions [Ohk14]. Disease [NBB13]. Disentangling [SvHM^{+}11]. Disk [Gri19, Yar14]. Dismantling [BW12a]. Disorder [AW18, ARBJ15, BL11a, Ber14, BV11, CHHS15, Cla19, CTT11, DM18a, DR14, EFO11, Fra17a, GHS17, HTX^{+}12, Ito17, KS19, Mia11, NSS12, dAPS11, SZS15]. Disordered [ACL14, BBR12, BF11, CT13, EPS17, GTT14, Ito16, Lim16, LSBS13, PS18, RdAB18, RSB10, Sep13, WW16]. Disparate [DL17]. Dispersing [LS17, CS10c]. Dispersion [CJW17, Cla13, HS10, JMRC16, Sch15]. Disperson [ABT^{+}14]. Displacements [Chu18, ZSHL15]. Dissipation [AT18, ED15, FGJ14, GH16a, GH16b, JRS15, LLL17, Mae14, PFR13]. Dissipative [ABS12, Ban10, BL10c, CF19b]. Dissolution [Gas16, GW15b]. Distance [CLSW17, JSJ10, Kol14, KMS14, GGP10]. Distances [AK18a, BvdHK19, BHJ^{+}12, CCH^{+}14]. Distant [RW19]. Distinct [Sha12, She15]. Distinguishable [Pet10]. Distortion [MC16, MC21]. Distributed [BBW15, Các14, Caë10, Gre12, ISZ16, ISZ17, PRD11, PRD12, WCX^{+}11]. Distribution [Ari11, BNP14, Bec11, Bur11, CCG14b, CB16, CGR12, Dai17, DLR14, DMP17, FLTV11, Fer18, GV12a, GLU12, GI19, HY16, HY17, JS17, Ken16, KMB14, KD19, KLACTL19, Lan17, LR18, LR19, Lee10, Lee12, LPS19, MT17, Mar16, NMV11, OCM15, Ost16, PT14b, PSW17, PR19, RW14, RZ17, RBGV15, Sch12b, Sha12, SY12, Sim11, TWT14, TW17, ZGH18, vEdG11, vdHLK18]. Distributional [ASA15, DS17]. Distributions [ABA14, AJ19, CTB10, DDC18, DNP17, FC13, FL15b, FK17, HM13b, JK12, LS17, MS14, Ost16, Pir14, RELV11, RT12, ST16b, WK18, YB14, vGRS16, Ast12]. Diverge [LK18]. Divergence [CLL19, TT17a]. Divergences [AvB16, Gal14]. Diverging [OR19]. Diversity [DDC18, HZS11]. Dividing [Yam13]. Division [RELV11]. DLA [PZ19]. DLR [Coq15]. Dmitry [Con13]. DMPK [BR10, BBR12]. DNA [MMM15, SS11b, XZ19]. DNLS [IPP14]. Do [SBK10a]. Dobrushin [CvELR18]. Does [HiS18, JLLP17, LTR17, MC10, Sch15, DRT18]. Dollar [CC18]. Domain [CPS19, GI19, Her13, PM17, Tak10b]. Domains [BGL^{+}11, BCL10b, CS16c, HBB10, RBGV12, RBGV15, TP19]. Domb [Fis12, Fis11]. Dominating [ZHZ15]. Domino [NS10]. Doob [AU15]. Döring [HY19a, Sun18]. Dots [GMT17a]. Double [FS11b]. Doubly [HBB10]. Down [CC19]. Drag [Bir18]. Drastic [Gal13]. Drift [BB13, CP10b, CDdS14, HVW12, Kat12b, LNP13b, LQR12, MGZ14]. Drifts [FF14]. Driven [AGGL^{+}16, AM14, ABFP15, ADP14, BL10b, BGL14, BL10c, BDL10, BKP13, BP15, CL16, DLR14, DXZ14, Eri18, ELX18, FSV10, GKW12, Grm17, GW15b, HS14a, Hui17, JPS17, KP11, KNSS18, LNS12a, LL19, MS11a, MdG13, Mar15, MC10, Pir14, Ply15, PSS15, PST13, PR19, QD12, SWK^{+}18, SZS15, Sim10, SH12, VP14, VP15, XLL^{+}15, YBF^{+}17, DNBS10, LH11]. Driving [DM18a, RC17b]. Drop [Aza11]. Droplet [Pat11, SM12a]. Droplets [MCK15, dHNT12]. Drude [BM11, MP18b]. Dual [Thi16, YB14, BCF10, Wan12]. Duality [ACR18, CGGR13, GRV10, Gro19, IS11, LM12a, Ohk10, RS18, vGRS16]. Due [BA14]. Duffing [XLL^{+}15]. Dugesia [CLTC15]. During [OV15]. Dwell [GV12a]. Dyck [HDP17]. Dynamic [AT12, ABF16, ByChL^{+}19, BL10d, DK10, EKD12, FMAG11, GOPS11, GV12b, KWZ14, LZ15, MM14, PS19, PS11, Sha12, TT17a]. Dynamical [Abr17, ADF18, ABJ12, AFFR17, BW10, BvE11, BJ16, BGP10, BCP11, BP11b, BLT12, BP16, CR11a, CEB^{+}15, Col14, DGL16, DS17, DOGK16, EP14, FSV10, FdHM14, Fil16, FKKO15, Fra11, Fre14, FFT11, GJMS10, GXL12, GPGA17, GKLT11, Hag13, HY17, Hor16, HŚ16, IPS10a, JPW14, JM10, Luc12, LFW12, LFWK14, MCG12, Mal12, MNS12, Mih11, MV16, NSV12, ORW15, PdOC17, Sor18, VP14, VP15, VV19, WG19, WCX^{+}11, WG18, WL13, ZP15, ZP16, dCFC11]. Dynamically [MSB18, MM13]. Dynamics [ABC10, APZ19, AWM13, AF14b, ABS12, AGJP19, Ano15d, AFG12, BBMW10, Bao17, BR11, BCP13, BPP18, BD16a, BKK15, Bir18, BL17, BNT13, BLZ14, BIM18, BBK17, BSS14, Bur19, BM16, CT13, Các14, Cam13, CCG14a, CFL17, CL13b, Cen13, CLS11b, CDCL18, CFS18, CNS15, CDS10, CM18, CDL^{+}12, DSS15, DF18, DEK^{+}15, DFF18, DARM^{+}13, DDF15, DK09, DRS18, Dor16, DG15c, DNP17, Eri18, ELX18, EK10, FKK10, FH11, Fri17, FT19b, GvdHW17, GS19, GNS18, Grm17, Gup16, HKR17, HKR18, HKR23a, HKR23b, HZ13, HHV16, HE17, Hui17, HT19, IPP14, JMH13, JPS15, Kim12, KO15, KM19b, LL13, LN13, LPS12, LH11, Lee19, LZ10a, Lef13, LNP11, LE15, LO17, LL19, ML15, MBGK12, MW12b, MS12a, MG17, MPM17, MT11b, NS12, Naz18, NKK19, NBB13]. Dynamics [OC12b, PL13, PT14a, Per10, PST13, Qia10, Rad17, RT11a, RdAB18, RST16, Riv16, RT16, RMN15, RT15, RNDS13, RDNS15, SML19, SLM12, SvHM^{+}11, SGS19, Sim14, SLST13, SG17, Ste19, SK19, SSE15, TZ16, TLC13, Tou12, Tur13, WTM19, dHNT11, dHNT12, Ara11b]. Dyson [GDMS19, LM13]. Dzyaloshinskii [QD12].

Each [CLTC15, CLTC23]. Early [Mer14]. Earthquakes [TGP12]. East [ABF16]. Ecological [BML12]. Ecology [CG11c, PCMM18, SCSS18]. Economic [Bar14, Bou13a, DLR14]. Economy [Hub13]. Econophysics [AFI^{+}10, Sta11]. Ecosystem [IK17]. Edge [BNY16, CFTW15, CP14, DLY18, EP14, FK18a, HF12b, Jan18, PWZ16, Shc11, Thä11, XP17]. Edge-Weighted [DLY18]. Edges [CDG12, EH12]. Edinburgh [BKPW14]. Editorial [Ano18b, FMR13]. Edouard [Kar11]. Edwards [CGG^{+}11, GOdSS11, Wre12]. Effect [ADH12, BH11, BGJ15a, BL11a, CE14, CSC11, DKLS19, FY16, GJ15a, HL16, KNK15, Koi18, Kuo15, LNS^{+}12b, RESA10, SM12a, WBL11]. Effective [Cha14a, CFS18, Hag15, JSJ10, KA17, KCB13, LL13, LS14b, MiS18, PE19, TT12, VAY^{+}12, Zho17]. Effective-Field [VAY^{+}12]. Effects [AMS19, BP11a, CCD15, CG10b, CR11b, CF19b, DWTW16, DDC18, GV12b, HT15, HRW14, HS10, KT12, LN19a, LLS13, MSB13, NE16, Sam17, SSB15, Sha18, TT12, TP15]. Efficiencies [LNP11]. Efficiency [FJLS18, KYA16, LL10, SB15a]. Efficient [APdM^{+}18, Cli10, Moh11, Wu14]. Efficiently [DNP17]. Ehrenfest [DD15, Tro10]. Eigen [AP11]. Eigenfunction [KK14]. Eigenfunctions [TV16]. Eigenspace [Lan15]. Eigenstates [Pro15]. Eigenvalue [BPZ13, EF13, Kum19, Mia11, MP14, Mov16, NT18, Shi16, Wan12]. Eigenvalues [Fre17, GMT17b, GMT17a, Kar12, Krü12, PWZ16, PS14, Shc15, dMPTW16]. Eigenvector [PSC18]. Eigenvectors [BP12a]. Eight [HF12a, KS13]. Eight-Vertex [HF12a, KS13]. Einstein [ADU17, Ark13, BC12, CL19, Fid15, GMM19, GJ15b, KPS19, LPK13, LZ11, Lu13, Lu14, Lu16, NP12, TKK15, TT17b, UV16]. Elapsed [MQW18]. Elastic [MdSB18]. Elasticity [CS16a, Raz18]. Electric [UV16, dOP18]. Electrical [CZZ13, KMKT11]. Electrodynamics [BDDH14, FS14b]. Electrolytes [DLLX16]. Electron [DP14, GM10, Miy19]. Electrons [CE19, DI13, Sit11, UV16]. Electrostatic [Zho17]. Elementary [Jus10]. Elements [CG11b, IPS^{+}10b]. Elephant [BL19, KT19]. Eli [Leb11]. Elliott [Kie11]. Ellipses [TU17]. Elliptic [AP14a, BK18, IL11, ORSV15]. Embedding [LK19, MNBC19]. Embryology [Per19]. Emergence [BHMGM13, HR16, HR23, OV15, Pir14, Sch11, Yam13]. Emergencies [AFS^{+}13]. Emergent [BJGL^{+}17, HKR17, HK19, Qia10, VV19, HKR23a]. Emery [ACL14, Lim16]. Empirical [GLT15, RW14, Shi13]. Employing [KBSM16]. Emulsions [SSB15]. Encounter [AKH13]. Encrypted [Fyo19]. Endomorphisms [MT16]. Endpoints [Bur11, Jia14]. Energetic [Bec11, TFES19]. Energies [ANSW14, Bec11, BBD^{+}11, GRRR13, Ser14, SST15, Yin10]. Energy [ACH15, AM14, AF14b, AK14b, AC14b, BL16a, BGN^{+}17, BFKP10, BO14, BJP17, BL11a, BGJ^{+}15b, BW12b, BLT12, CA19, CBG14, Cha19, CN13, CFN15, CGG^{+}11, CG12b, EAL12, GT17, Hog11, HA19, HS18, HRW14, HW13, Huv12, Ily16, Jan15a, JLLP17, KV15, KT12, KM18, Leb16, ME11, MMSY11, MLS16, MMA15, Miy13, MS11b, PGS12, RVB16, Sak12, ST16a, SW11, SKT16, SvHM^{+}11, WBL11, Web11, Wei18, Wre12, WZIG14, ZBVE11, Zha12c, ZW12]. Energy-Driven [AM14]. Engine [Tay17]. Enhanced [Mül16, NiS19c, ZW10]. Ensemble [BL11b, CO17, GdHR18, Gil15, HF12b, KW12, Kum19, KM18, KM19c, May13, MP13, Nak14, PT15, Shc11, SW12, TKK15, VB11, dMPTW16, vKSZ18]. Ensembles [AGR19, AP14a, Bar14, BGP15a, BMSS13, CO17, CCG14b, DHR18, FS18, FS11a, Fun14, GL14, GO13, HH13, HT19, KO15, Koz17, LM13, LZ10b, NT18, SWKS14, Tas18b, Tou15, VV19, WWKK16, Zha15]. Enskog [ARS17a, CLMK18, GPMSBBSV15, MdSB18]. Entangled [AM10]. Entanglement [ADP14, BCFS17, EPS17, EV14, FL16a, GNPS13, HA19, NMV11, PS18]. Entire [BS17]. entrant [MMR18]. Entrepreneurship [ZGH18]. Entries [Afa19, AEK16, PRS12]. Entropic [FCK15, HHM17, IST15, JPW14, JPS17, MM17b, MCK15, Pav11]. Entropies [HL14, Ras11]. Entropy [ADU17, AG12b, AP18, BCJ17, BCC^{+}16, BCS18, BL10b, BGL14, BSV18, Bir18, BC14b, BMSS13, CL10, DJRZ11, DHX19, EPS17, Fra17b, Fra11, FL16a, Han16, HL14, HS14b, HRW14, Kie17b, KNiST11, LNS12a, Lia19, LN11, LNT13, LH13b, MS19a, ME14, MN16b, MU13, MDP^{+}18, MHD17, NT17, OCM15, PS18, Per17, RS15a, RGL11, ST11a, SS19, SCSS18, SP18, Tay16, Tia14, Tos16, WX15, WXX16, YSSH13, ZST18, Zha14, ZP15, ZP16, vdHLK18]. Entropy-Based [BCS18, SCSS18]. Entropy-Driven [LNS12a]. Entry [HY16, HY17, LH17]. Environment [AEWCD19, ADGPP17, Ber14, BZ13, BCLL16, CFN15, DP19, DLR14, Hag15, HL16, Ko19, PZ17, YK13]. Environmental [BMC17, DKS18, Kan12]. Environments [AT12, ABF16, BR13, CG17, JM10, RL11, SGS19, SSE15]. Epidemic [LLS13, PSC18]. Epidemics [BDL11, Gra13, TSS13, WQ10]. Epistasis [WLJH18]. Epitaxial [XTpXpH12]. Equal [SC10]. Equalities [HH15b, KNiST15]. Equality [HSZ19]. Equation [AvB16, AB10, Afz12, AFCA16, AF12, AF14c, BR10, BL18, BCJ15, Ber12, Ber15, BDL16, BGP15b, BGZ17, BP19, BJM15, BBS^{+}15b, BE16, BD15a, CL19, CJW17, CMW15, CCH^{+}14, CF15, CES19, CIM14, CP10a, Che13, CLMK18, CLL18, CM12a, CL18, CE19, CG17, CHHK19, DLBK11, DWTW16, DRS18, DL12, EM10, ED15, Fow19, FK11, Gao18, GLM^{+}15, GJ15b, GVJ^{+}18, GW15a, HX15a, HX15b, Hag15, He14, HY19b, IS13, JP18b, KP12, KMTC10, KSSH15, KV15, KV16a, KK15, Kra16, KT17, KM17, Laf19, Lau18, LWL^{+}12, LWW18, LZ11, Lu12, Lu13, Lu14, LY16, LQR12, MGZ14, MU18, MN16a, MM16, MM17a, MN12, MdSB18, MiS18, Mol17, Mor11, MWY16, NT17, Ngu18, NTV16, PSW17, Pic10, QS15, RT14, RPPF15, RBS16b, ST18a, SWB10, Sha10, Shi16, SVRL11, Tak09, Tak10b]. Equation [Thr18, Tor12, Tri14, Tri17, VV17, WGLE11, WZ12, WYG16, XTL14, Yan11, Yan15, ZLL13, ZWGM13, dZS11, dZS13, DS19a, Tak10a]. Equations [AG12a, ABH^{+}17, BC12, BFVZ10, BCF19, BLS17, BL17, BNT13, BLZ14, BM12a, CCG14b, CLL19, CN14, DGL16, DM10, DXZ14, FKLL16, FSS13, FRT15, FK18b, FS11b, FS14c, GOP14, GHMR17, HJ17b, Her13, HR18, HVW12, HMW19, HP11, HT11, JLMG11, KTJ10, Kel19, KS13, Kha19, KK14, KOSV18, LT10a, Li12, LLL17, LLS17, LWY18, LW19, LT10b, Liu15b, Liu15a, LK19, Mar11b, MNS12, MNV11, MS10, MM17b, Neu14, NV14, Pav11, Pel14, PR15a, QLCL16, RS13, SCY^{+}12, SSE15, Sug10, WXX16, XTpXpH12, Zha10c, ZGL13, ZGL15]. Equidistribution [DMS12]. Equilateral [ACM11]. Equilibrated [Mar11a]. Equilibrating [Kuo15]. Equilibration [BP11b, DP17, Sor18, Tos17]. Equilibria [DLR14, Her13]. Equilibrium [AN15, BVL16, BP18, BFL18, BD15b, CL19, CCGT10, Dal11, GR12, GS12, GdK10, HiS18, HY19b, IiS17, KJZ17, Kem11, KW15, LS13, LNP13b, LNY16, LO17, Mor15, Nán16, PT14b, RM16a, RVB16, Rue14, San13, SY12, Shi16, SVRL11, Tas16, Tri14, TAG10, Wre17, dlLSZ16, BGL14, BGN16b, CRS14, CSL19, CRTZ13, Cro12, DF18, Dav11, DG17a, Eva16, Fri17, GL14, GS17b, GS17c, Gor18, Kim12, Kle15, MPTV12, MV16, NE16, PL13, Rue17, SGS19, Ste19, VL12, Zia10, Gou14]. Equipartition [BP11a, BP11b]. Equivalence [CO17, Far15, Fun14, GL14, GdHR18, HT19, LS14b, Tas18b, Tou15, Tou18, ZGL15, VV19]. Equivalent [HS14b]. Erased [AKM13, AO19, Ken19a]. Erdos [ALS14, DvLM16, KLS19, LV11]. Ergodic [BMR10, CS10c, CT10, DM13, GT15a, LM12a, SLdEC11, SM14, GS11a]. Ergodicity [BHF^{+}12, BFNZ11, DXZ14, Fer14, GHMR17, LS16b, Mor18, dAPS11, VV19]. Erlang [PRD11]. Eroders [dSRT15]. Erratum [ACL^{+}11, Coh10, FVV15, Fan16a, Fis12, GL17a, Gin14a, GH16a, GS17b, GKLS17a, Gut12, Leb12a, Mit17a, QTDS12, SB14b, Sod17, VP15, WFK11b, ZP16]. Error [MM13, RST16]. Errors [AV16, Aur18]. Escape [CNS15, LH17, Mar16]. Estimate [Aum15, BS13b, BS20, LWW18, Sod11, TV15a, Wid17, Yuh15]. Estimates [Aur18, BYYY19, KOSV18, MZ19, Mih19, Shi15, Wil11]. Estimating [BBD^{+}11, CGS15, DG15a]. Estimation [ABT10, ABT11, ART15, BG11, Cro12, GLT15, HA13, HLZ17, MM13, RM16b, ZST18]. Estimators [AG15b]. Eu [EFO11]. Euclidean [Kol14]. Eukaryotic [HCLR11]. Euler [BCF19, BLZ14, Her13, Li12, LT10b]. Euler- [LT10b]. Evaluating [Luc18]. Evaluation [Asa13, JSJ10, JD11, SWB10]. Evanescent [BGJ^{+}15b, HR18]. Evans [BCHM12]. Evaporation [CRTZ13]. Even [ST16a]. Event [AGJP19, MC17b, RBS16b]. Events [BLM13, DHS18, GKLT11, OMC11, WFK11a, WFK11b]. Evidence [TS12]. Evidences [FZ11a]. Evolution [AG12a, BAS16, BKK15, BCM12, CCM16, CM18, DKS18, DLR14, Ghe10, GN13, IiS17, JP18a, Kan12, KPR18, Ken15b, KL15, KSH11, Lan16, Los17, PSK10, PRSS17, RBM^{+}18, Sch12b, Sim14, Yag16]. Evolutionary [CG11c, CAG^{+}13, FK12, HZ13, Hui17, MSB13, OTNN11, Riv16, SK14]. Evolutions [FK18c, JL18, KP11]. Evolves [HZS11]. Evolving [LPS12, MSB18]. Exact [AG12a, APdM^{+}18, ACM15, Asa13, CFTW15, CS12, CS15, CEB^{+}15, DDC18, ES13, Fan16b, Ghe10, Ily16, JD11, JJB14, Jus10, KNiST15, LO17, Liu15b, MT17, MS19b, MR19, MG19, MS19c, RKGZ12, ST16a, SGU13, Sam15, Sco11, SX10a, TC11b, Van17, Web11, ZST18, Fan16a]. Exactly [BD15a, GG11b]. Exactness [Mor12]. Example [BGN16b, Bra14, Ein12]. Examples [BML12, Coq15, RW12, Ter13]. Excess [NT17, YSSH13]. Exchange [BJM15, vGRS16]. Exchangeable [Jan18]. Exchanges [BGN^{+}17]. Excitable [BGTV11, DT18]. Excitation [ZYZ^{+}18]. Excitations [AN15, BNY16]. Excitatory [Vid15]. Excluded [Cha15a]. Exclusion [AFR19, BMNS17, BL10b, Bla10, BW17, BS19, CL16, DR13, Eri18, ELX18, FGN14, Kan14, LP16, MPTV11, MPTV12, Mat15, NS11, NVL11, Ngu19b, Pro15, SS10, SJHW11, Sim11, TW13b, TW13a]. Exclusive [AY10]. Excursion [BK18, KMB14, PR15b]. Exist [JLLP17]. Existence [BR13, Bra14, BE16, CH14, Coh09, Coh10, LY13, LWY18, SST15, Zha10b]. Exit [BBP17, Dai17, IPS10a, Ken16, Mar16, RS15b, RBGV15, Ryt12]. Exotic [MMST13]. Expanded [TF12]. Expanding [BGN16a, CRV17, DP17, Fer14, GLBP12, JPV18, SB16]. Expansion [Alb16, BDY17, CM12b, Coh09, Coh10, FF11a, Fed17, KK14, KiMM13, MCG12, NKR15, Per17, SGU13, Sou18, Tat13, TW14, ZW12]. Expansion-Modification [SGU13]. Expansions [BFP10, Fed13, GXL12, Jan15b, WZIG14, dlLSZ16]. Expectation [Fed17]. Experimental [KBB19, LKR^{+}11]. Experiments [AK16, YBF^{+}17]. Explanation [CdlL10]. Explanations [dlLL11]. Explicit [GLBP12]. Exploitation [RCV16]. Exploiting [CM18]. Exploration [BJS17, RCV16]. Explorations [BdlL13]. Explorative [Tem14]. Explorer [Ken19a]. Explosion [BPR14]. Explosions [FLP10, FG12]. Exponent [CLTT13, JL17b, Kar14b, ZLL13]. Exponential [BKS12, BCM16, BCCD18, BMR10, BSM^{+}16, Các14, CJN18, CLL19, CS19, CG11b, Chu16, CGN16, DLY18, DXZ14, Goo12, ISZ16, Tri14, Yin13, Yin16, Zhu17, dMP12, Yar14]. Exponentially [RVY18]. Exponentials [Han15]. Exponents [AP12, AK19, BPP18, Bou13b, BPRT14, ESPP^{+}14, For13, GM13, HL16, Mol18, ORW15, PSAPR12, Wu18, XTL14, dlLOP11]. Expressing [HL14]. Expression [MBGK12]. Expressions [NiS19b]. Extended [BCJ15, BD16a, BFM10, CCM16, Hag15, Kel19, LH17, Lia13, LR15, Miy16a, ST11a, Tou14]. Extending [Luc18]. Extension [AW18, HT11, MN14a]. Extensions [GKLS17a, GKLS17b, MSZ12]. Extensive [GPP12]. External [AP10, BBLP12, BBLP13, BC10, CEB^{+}15, CV16, DZ15, FKR12, FR17, GNP16, KMTC10, MM12, Ngu17, PL13, RdAB18, Sak12, SdlL18, SX10a, Wan12, WZL^{+}14, WZ12]. Externally [Grm17]. Extinction [FC13, LPS19, PK10, SL12, Zuc11a]. Extra [Hol11, SS11a]. Extra-Vertex [SS11a]. Extraction [BCL10b]. Extremal [Ast12, Ast13, AFFR17, FS19, Fre17]. Extremality [KR15]. Extreme [AFFR17, FLTV11, FFT11, FDR12, LFW12, PS14]. Extremes [CCH16, LFWK14, Sch12a].

Face [Asa13, JD11, Kel19]. Face-Centered [Asa13]. Face-Centred [JD11]. Faces [CDG12]. Facet [EP14]. Factor [FO18, MSS11b, Yoo10]. Factorization [CMS13]. Factorized [RS18]. Factors [SKT16]. Facts [DRS10]. Fading [Ber19]. Failure [TGGS13]. Fall [HJR^{+}11]. False [Sha18]. Fame [SR13]. Families [BDV12, BCC^{+}16, Thi16]. Family [BCF10, BIP11, Caë11, CCD15, FZ10, Kie17b, LK18, LN15, RL17]. Fan [Gon14]. Fans [BNTT16, CFP10]. Far [SVRL11, ABJ12]. Far-from-Equilibrium [SVRL11]. Farey [BFKP10]. Fast [AGMM^{+}12, BC14b, BGTVE16, CM18, DSS15, Gao18, Har11, LL13, LKD12, Pir18b, SS18a, TM10, dSLPV17]. Fast-Variables [CM18]. Fate [FG14]. Faults [TGP12]. Features [CFTW15, LKD12]. Fecund [EF18]. Feedback [Chu16, FM18, KYA16, Vid15, XZ17]. Fermi [AMS19, BC12, BP11a, BCP13, BPP18, CR11a, CSC11, DDN14]. Fermion [CPS19, Mor18, ST14]. Fermionic [BGJ15a, HJ17a, SSE15]. Fermions [AG14, BF11, CMO18, EPS17, LSBS13, PS18, PRSS17, RK12, dWL10, dWL14]. Ferrari [EP14]. Ferrimagnetic [EKD12]. Ferroelectricity [QD12]. Ferromagnet [CG12b, DKKP14, Har11]. Ferromagnet-Spin [Har11]. Ferromagnetic [Afz12, BLW18, BC10, BCPS18, GNP16, LN11, O’C12a, SST14, JJB14]. Ferromagnetism [GM10, Miy19, TT16, Tan18]. Ferromagnets [FR17, GT16, WW16]. Feynman [Aur18, ZST18, ZLL13]. Fiber [HTX^{+}12, RG18]. Fibonacci [DMY13, Lan11]. Fick [ACCG19, DOP19, Ngu19a]. Field [AP10, ACM15, AM18, AM12, AK14b, AKL18, BG17a, BT12, Bar16, BGP10, BCF17, BCF19, BR16, Bha15, BS16, Bir18, BDL10, BP16, BGN16b, BIM18, BV15, CJN18, CF11, CS15, Cha19, CHHS15, CGY17, CDS10, CR16, CCL19, Cra11, CDL^{+}12, DM12, DF16, DvLM18, FC11, FMM^{+}15, FKR12, FW12, FR17, FMMPS18, GRT17, GL16a, GNP16, GV12b, HBB10, HKN16, Hon19, HMRW13, Ito18, KMTC10, Kar11, Kar18, KE10, KM13, KN16, KOT11, Kra16, LS16a, LL16, LLM19, Lan15, LM12b, LR15, MN16a, MW12b, MNS12, MM12, MLCPS13, Mor12, MC10, MC11, NP16, OK14, Ost16, PL13, PR15b, PRSS17, SML19, ST18b, SB16, SX10a, TV12, UV16, VAY^{+}12, VL12, WGLE11, Wu14, Bré10, CR17, IK10]. Field-Theoretic [NP16]. Fields [AI12, ALM18, BC10, BG12, CGL12, CAS11, CV16, CF19b, D’O14b, DZ15, EH12, FVV14, FKR12, Gan18, GL17b, GM13, Hon19, Jaf15, Kar14a, Lan16, LMT15, LM17, LMN18, MPL^{+}16, Mol18, Mou15, Sch10, SS15, WZL^{+}14, ZGL15, dOP18, FVV15, GL17a]. Filament [Raz18]. Filaments [BCF17, BCF19, MG17]. File [Bur19, KMS15]. Films [CCD15, DOGKP19]. Filtering [ST11b]. Finance [SCSS18]. Financial [BCKY19, BPR13, KiMM13]. Finding [Mon15]. Fine [Bun14, Sch13b]. Finite [AG12a, AGR19, AFGL15, BCHW19, BP11a, BLT12, BM12b, CPV10, CCG14b, CSAS17, Cha12, CP10a, CG10b, CAS11, CPSV10, Coq15, DR13, DM18b, ELO11, FF11b, Fer18, GLO10, GLM18, GKLS17a, GKLS17b, GG11b, HCI14, IK10, IP12, KPS19, KSH11, Kra16, LHZ^{+}19, LMM16, LS17, Lu13, Luc16, MP18a, MU15, Mit16, Mit17a, Mit17b, Pan16, PRD11, dAPS11, PT15, RBS16b, Sam17, Sam18, Shc14, Sug10, Tau11, Thi16, TP19, TC11a, TC11b, YS13, Zoc18]. Finite-Amplitude [RBS16b]. Finite-Diameter [Cha12]. Finite-Dimensional [LS17]. Finite-RSB [Pan16]. Finite-Size [BP11a, IP12, Sam17, Sam18]. Finite-State [TC11a, TC11b]. Finite-Time [GLM18]. Finite-type [IK10]. Finiteness [HKR16]. Fire [Vid17]. First [AB18a, AD15b, BGL^{+}11, Bur17, CP16, CDCL18, CPSV10, DO18, DT18, FZ11a, Fra11, Han16, HKW11, IPS10a, Ken16, KV16b, MS19b, Nak19, PS14, Pet10, SM12a, SM12b, TM10, Tak16, VYH11, vKSZ18]. First-Order [FZ11a, HKW11, SM12b]. First-Passage [BGL^{+}11, DO18, Nak19, TM10]. First-Principle [CDCL18]. Fisher [Ano19a, BD15a]. Fission [CG16]. Fitness [FK12, GvdHW17, Hui17, HSFK18, KSH11, MBC^{+}13, OTNN11, Sch12b, SK14, TM18, WLJH18]. FitzHugh [HR10]. Five [Bur17, SP18, Zia10]. Five-Moment [SP18]. Fixation [HCI14]. Fixed [CQR15, CCFR18, FZ11a, Gil15, Jia14, Ken12, Koi10, Tak10b, UK16]. Fixed-Connectivity [Koi10]. Fixed-Length [Ken12]. FK [Wu18]. Flat [FN15a, Koi10, Mie18, Tan18]. Flatness [Lep12]. Fleming [Mar15]. Flexible [HNZ16, Mar11a, Sch13c, KMS19a, LZ15]. Flight [AM14, Kol17]. Flights [BP15, DO15a, Gre12, Kol14]. Flip [BO11, BGJ^{+}15b, FdHM14]. Flips [Luk14, LMN16]. Flocking [CDG^{+}15, HFWT15, Jin18, MPT19, MT11b]. Flocks [BAC13, HH15a]. Flooding [CP17a]. Flop [CR16]. Flory [WCX^{+}11]. Flory-Distributed [WCX^{+}11]. Flow [AGGL^{+}16, AM12, Bar16, Ber12, BBR^{+}19, Che14, CGY17, CLMK18, CFS18, ES13, ES12, Els15, HT15, Mar11b, Mat12, MFLA15, MvS19, NE16, PM17, RM11, SBK10a, TLD18, dZS11, dZS13]. Flows [BM18, CVE14, DM10, DMS17b, Esl17, GL14, HPF15, NKK19, PT10, SSB15, VB11, VFT12, YKS16]. Fluctuate [HiS18]. Fluctuating [AN19, CDS17, DS19b, Spo14, SS15, dZS11, dZS13]. Fluctuation [ANSW14, BKS12, BC15a, CF16, CG11b, CM12a, FGJ14, FY16, GH16a, GH16b, HSZ19, Kan12, KLMP18, KT19, Lan16, LLL17, Mae14, Mal12, Nak17, PE19, Ren18, Van17, WES11, YBF^{+}17]. Fluctuations [ANSW16, AT12, BL16a, BvdHK19, BC14a, BB13, BD11a, BS17, CO17, CSL19, CC19, Chh12, CDS10, Dav11, DSPC14, DOGKP19, EP14, EAL12, EL12, FS14a, FN15a, FDR12, GV15, HPF15, HMRW13, JPW14, JPS17, KLS19, KMKT11, MP18a, MS17, NT18, OTNN11, PWZ16, PVCG11, PdS17, PRS12, RT14, RC17b, RM11, SB14a, SB14b, SMCF13, Shc15, TS12, TA16, dZS11, dZS13, vHL13, vdHvLS18]. Fluid [BT14, BC12, CLMK18, DDF15, FBE^{+}11, Gri19, Hon19, KA17, LT10b, Mat12, MvS19, ROS19, RM16a, TT17a, TN18, Xue15, ZL19, dZS11, dZS13, TN21]. Fluid-Dynamic [TT17a]. Fluids [BCJ15, Dal11, Fan16a, Fan16b, Fan17b, FS17, GLML16, GLM^{+}16, HP11, HT11, KID^{+}11, KL11, LK11, MS19c, Per10, PPK11, Sam16, Sam17, SM14]. Fluorescence [BL15]. Flux [GPMEA17, GPMEA18, PS11, Ren18, ScCdP11]. Fluxes [HHT10]. Fock [CR11a, FK11]. Focusing [JP18b]. Fokker [HX15a, HR18, JLMG11, MGZ14, ACCG19, AFCA16, AF12, AF14c, BBS^{+}15b, BL17, CLL19, DLBK11, DGL16, DWTW16, KP12, LWL^{+}12, LWY18, LK19, LQR12, MM16, MM17a, Shi16, Wu15, Yan15, ZGL13, ZGL15]. Fold [Osi16]. Folded [Mih11]. Folding [Ark10, VCT11]. Foliation [Gal15]. Force [Bur11, SdlL18, TT17b, WZL^{+}14, WZ12]. Forced [LZ15]. Forces [BC11, BCM12, CIM14, CZZ13, DL12, KJZ18, NP12, NP14]. Forcing [AB10, BdSPMS14, JPS15, WG19]. Foresight [PSVG18]. Forests [ACGM18]. Foreword [Ano14a]. Form [IiS17, Mou15]. Formal [BC15a, DF16]. Formalism [Aur18, GPMSBBSV15, GQ17, GTZ12, Kle15, LNT13, MMSY11, PT10, RMN15]. Formation [GVJ^{+}18, Mie18, Per19, PKBW19]. Forming [KID^{+}11]. Formula [AW18, BSV18, BCL10b, CFK13, DHK11, GJL16, IiS15a, JL17a, LV19, LWL^{+}12, Mas16, Tia14, Zha12c, dBP15]. Formulae [Hui17]. Formulas [Mat12, TW10]. Formulation [DM12, EN18, JJ12, Kan12, Ker13]. Fortuin [CEGW18]. Foster [JL17a]. Foundations [KLM13, Kie14, DP14]. Four [BSW17, Cha14a, Cli18, FL18, FL20, HKW11, Sch12b]. Four-Dimensional [BSW17, Cli18]. Four-Point [HKW11]. Fourier [GDL10]. Fourth [AF16, Sam19]. FPU [CP18, MBC14, MMA15]. FPU- [MMA15]. Fractal [DD10, EN18, RS14, SH16, VFT12]. Fraction [BFKP10, ZB13]. Fractional [ALM18, Ara11a, BK17b, BBC15, BL17, CP12, CJW17, CWD17, GOP14, GLT15, GOdSS11, Kaw16, LLL17, LWL^{+}12, LCZW15, LQR12, MS12a, Mit16, Mit17a, Mit17b, Mol17, Mol18, OPS10, OD11, OP12, RESA10, TFES19, XTL14, Zha10c, ZLL13, ZGL15, ZWGM13, ZWGM15, MGZ14]. Fractionally [EN18]. Fractions [OP12]. Fragment [KLM17]. Fragmentation [BDL16, KLM17]. Framework [AL18, BD16b]. Fréchet [LPS19]. Free [ACH15, AK18a, ARS17b, ANSW14, AC14b, ALS14, BL16a, BFKP10, Bec11, BL11a, BBD^{+}11, BLT12, CA19, CBG14, Cha19, CD16, CPS19, CFN15, CG12b, CMO18, EPS17, EAL12, GRRR13, HBB10, HH15a, HS18, HFWT15, HW13, Jan15a, JvLS19, KT12, KM18, LH13a, LE15, ME11, MS14, MMSY11, NiS19c, Ost16, PS18, PXX15, PR15b, RK12, Sak12, ST14, SW11, SvHM^{+}11, TV12, TAG10, TA12, WBL11, Wei18, Wen12, WCX^{+}11, WZIG14, Yin10, ZBVE11, Zha12c, ZW12, vdHK17]. Free-Energy [BL11a, BLT12, CBG14, MMSY11]. Free-Fermion [CPS19, ST14]. Free-Molecular [TAG10, TA12]. Freedom [XY13]. Freely [DP17, KMS19a]. Freeway [LLJH10]. Freezing [HRW14, Mot14b, RM16a, Web11]. Freidlin [Cam13, dSLPV17]. Frenkel [BdlL13, SdlL18]. Frequency [AWM13, CE14, HPF15, LCZW15, Wid17].

Fresnel [OD11]. Friction [FGJ14, GL14, HHV16, HVW12]. Frictionless [BJGL^{+}17, GP10]. Frobenius [LYX17]. Frog [LU19]. Fröhlich [Ano18d]. Front [FK18b]. Fronts [CQR13, Man11, MK19]. Frozen [DEK^{+}15, FGP15]. Frustrated [LMC19, MD10]. Fujiwara [Sta11]. Full [AK14a, Con18, CDG12, PK11]. Fully [KBB19, LMC19, MS19c]. Function [ADE18, AW18, AK16, BM18, BDG^{+}14, BCF10, Bot18, BS17, CB16, CGNP11, CL18, CG13, DR13, DM18b, FDR12, GLT15, GT15b, HM16, JL18, Jin18, JD11, LV19, LSS19, LSW17, MM12, MU16, Moh17, PS11, SS11a, Sam16, TS19, VW18, Zha12c, ZW12]. Functional [CSV11, DOGK16, Fan16a, Fan16b, FBE^{+}11, Sit11, Sun18, WES11]. Functionals [CTB10, KOSV18, PCM15]. Functions [Afa16, Afa19, AVW17, Ari11, BDV12, CS15, CFG13, DJRZ11, GLML16, GLM^{+}16, GJL16, GJS17, Gro19, Han14, Kad13, Lan13, Lim16, LMN18, Mat12, MU16, MP14, PT14b, PRS12, PPS11, RS18, Sch10, She15, Shi13, SX10a, Sim14, Tid16]. Fundamental [BHS^{+}12, Per13, QLCL16, Rab11, SS19, Tho12, tWBOM16]. Furstenberg [BPS12]. Further [LS12]. Fusion [Sch11].

G [Gin14a, Gin14b, Szá12, Weg14]. Gallavotti [BCHM12, FP11]. Galton [BGMS14, BZ13, SL12]. Gambler [LS16b]. Gambling [HH15b]. Game [BLL^{+}13, GPD16, HZ13, MSB13, SC10, CA18]. Games [APZ19]. Gamow [CH11]. Gap [BS13b, BS20, BNY16, CCG14b, CLSW17, GVJ^{+}18, NS11, PS14]. Gapless [Tan18]. Gapped [BN14, JV19, LSY19, dWL10]. Gaps [Bis19]. Gas [AKD19, AP14b, BGN^{+}17, Ber12, BDY17, BCKL12, CLS11a, CG10a, CG14a, CR11a, CSC11, CP10b, Coh09, CPSV10, CFLV19, DP17, DM10, Det12, DFL17, FSV10, GPMEA17, GM10, Jan15b, Ken10, Koi18, KKA19, KM19a, KM19b, Kuo15, MS19a, Moh17, MP13, NiS19c, NKR15, NWL19, Pir18b, PR19, RNDS13, Sch13c, ST18b, Sim10, SH12, Sor18, TH12, TKK15, TW13a, TAG10, TA12, YKS16, ZDG19, dLP14, Coh10, GPMEA18]. Gas-Particles [DM10]. Gaseous [NE16]. Gases [AB19, Bec10, BP11b, BFP10, CP14, CF19a, CIM14, Far15, KNSS18, MS14, MM17a, PSW17, RSY14, Sam13, Ser14, TT17a, Vie16, YM11, Yin10]. Gasket [JY18, Mis15, Mis16, Yam13]. Gate [Hep18]. Gaudin [MR19]. Gauge [dWL14, Yam14]. Gauge/ [Yam14]. Gauss [TM10]. Gaussian [ABLT17, AEK16, ADGPP17, AB18b, ABT10, ART15, AGO18, BF12, BS15d, BS17, CCG14b, CRV13, CCH16, Dod15, EL12, Els12, FS14a, For13, FL16b, HBB10, Ito17, Jav15, KSSH15, Kar14b, KT19, LT10a, Lan16, Lie12, LZ10b, Mol18, MC10, MC11, Mou15, Nak14, NT18, Ost16, Pir14, PR15b, Sak12, ST18b, Shc11, SW12, TV12, WHC14]. Gaussianity [BTV14]. Gel [dCFC11]. Gel-like [dCFC11]. Gene [CG11c, KSH11, MBGK12]. Genealogical [EF18]. General [AB19, ABA14, AF12, BCL10b, CZZ13, CV16, CNS15, EF13, GQ17, Ito16, JJ12, KL15, KS13, LS15a, LFWK14, MW10, MQW18, Miy13, Pav11, PR15a, Ras11, Tas18a, YBF^{+}17]. Generalised [BL17, CB16, CTT10, RS15b, TWT14]. Generalization [CGR12, GC17, KMO16, RP12]. Generalizations [You17, vGRS16]. Generalized [AT18, AFCA16, BLS11, BBH11, BP11b, BYYY19, CGS15, CCEF10, CS15, CC19, CL18, DvLM16, FLTV11, FS11a, FH11, GH16a, GH16b, HKR17, HKR23a, HT15, Hel16, JL17b, KK15, KLMP18, KS18, KM17, Lan11, LPS16, LW19, Liu15b, LZHS19, Mis19, Ngu18, PPS11, Rue16, SX10a, SVRL11, VV17]. Generalizing [DO14a]. Generate [Cie17, Lep15]. Generated [FKR12, Jav15, Ong14]. Generating [Moh11, RS18]. Generation [ACM11, MS11b, TJ15]. Generator [Gou15]. Generators [RT17b]. Generic [AJ19, BJS17, BSV18, CS16c, FL12, Far15, KLMP18]. Genetic [APR15]. Genetics [MSS^{+}11a]. Gennes [KWZ14]. Genome [KSH11, MSS^{+}11a]. Geodesics [AB18a]. Geometric [CLP17, Det18, DvLM18, GW12, GGD16, KA17, Lia19, PK11, PSAPR12]. Geometrical [BTV14, YSSH13]. Geometries [BPDH10, Sam15]. Geometry [ADE18, Dzu11, EV11, Mih11, MB18, TS12, TT17b, dlLOP11]. Geometry-Dependent [TS12]. Geophysical [VB11]. Geostrophic [BLZ14, Li12]. GI [FM18, FM18]. GI/GI/1 [FM18]. Giant [DRT18, EMO18, GVJ^{+}18, SB14a, SB14b]. Giardinà [Gue13]. Gibbs [ALS18, ART11, AFGL15, AP18, ACR18, CCR17, Coq15, CRL15, Cra13, EHR12, FS18, FdHM14, FO18, GLO10, Gan18, GRS12, GRR17, GJS17, HMO12, HM13b, IY14, Kad14, KSY13, Kem11, KO15, KRK14, KM19c, Lan17, MTVU18, ME14, MSLT16, Pan14, Pan16, Pet10, Pir18a, PSS12, RvH14, RW12, SL12, Var12, Yoo10]. Gibbs-Non-Gibbs [RW12]. Gibbsian [DG17a]. Gibbsianness [EK10, Ny13]. Ginibre [AP14a, HH13, Kar18, dMPTW16]. Ginzburg [RBS16b, Ser14]. Giordana [Mon12a]. Gitterman [Bai10, Dyk14]. Given [MC11, Tid16, vdHLK18]. Glansdorff [MN15]. Glass [BLW18, Con18, Fyo19, Har11, Pan16, Par17, Wre12]. Glasses [AD11, ANSW14, ANSW16, BS16, Cas14, CMS13, Har11, Mac13, SN13, Gue13]. Glassy [AB17, HT19, dCFC11]. Glassy-like [dCFC11]. Glauber [CDL^{+}12, EKD12, FT19b, Jus10, KO15, MW12b, SG15]. Global [Aur18, CDCL18, CH14, CN14, DFL17, FL15a, FN13, HX15b, LWY18, LRL17, NiS19a, Zha10b]. Globalization [KL13]. Globally [Fer14, KT17]. Go [TK16]. GOE [FN15b]. Golden [Han15]. Goldenfeld [Ano19b]. Goldstein [BS20, BS13b]. Gompertz [LPS12]. Good [AEWCD19]. Goods [GNPS13, TBD13]. Gordon [GOP14]. Gould [bA10]. Governing [MK19]. Grad [CG10a, Mar11b]. Grad-Type [Mar11b]. Grade [LT10b]. Gradient [Bir18, HCLR11, MM17b, TT17a]. Grain [SWKS14]. Grained [Ark10, BHS13, Mül11]. Graining [BTV14]. Grand [TKK15]. Granular [AR11a, MPL^{+}16, RS19, RC17a, SLM12]. Granular-Like [MPL^{+}16]. Graph [AW18, AZ11, AKL18, BT19, BCCD18, BA19, BSM^{+}16, Cha14a, CS10a, Con18, FK18a, KPR18, LV11, MSB18, Mon15, Rue18, SX10a, SX10b, TV15b, XZ17, ZW12, Zhu17, vdHK17]. Graph-Counting [Rue18]. Graphene [BN18, CR14, Dod15, MP14]. Graphical [CV16]. Graphs [APS12, AHDV17, AF14a, ADS^{+}19, AP18, BvdHK19, BCF10, BR13, BJR10, CP17a, Can17, CS15, Con11, DLY18, DGL16, Det18, DvLM16, DvLM18, DGvdH10, DGGvdH18, GdHR18, GSSV11, GGvdHP15, GGD16, GS15, HNT18, Jan18, JvLS19, KLS19, KRRS17, Lan17, LR18, LR19, OR19, PPS16, Per17, RS15a, Rad17, RVY18, Rue18, SH16, SST14, SB15b, Tau11, WTM19, XP17, Yin13, Yin16, Zha12b, vEdLV16, vdHKvL18, vdHLK18, vBKBS13, vBSB15]. Gravitons [Art19b]. Gray [CLMK18]. Grazing [He14, Tak15]. Green [IiS15a, JD11, LV19, Mat12, PPS11]. Gregariousness [DNBS10]. GREM [DM18b]. Grid [Chh12, OF18]. Grids [NZB16]. Griffiths [ACL14, Lim16, Miy16b]. Gross [AU15, CPRY11, Pic10]. Ground [AF14b, AS16, BW12b, BCPS18, GT12, GT15a, Har11, Hog11, HRW14, MMST13, Miy12a, Ngu17, Sch13b, SST15, Wre12]. Group [ACL14, AM12, Bat17, BBS15a, BM12a, Bré14, BS15d, BS15e, BS15f, BS15g, CdlL10, CCEF10, ES13, Gal14, HK19, Ken10, LHZ^{+}19, LS14a, LNP19, OC12c, OC12b, ZJ14, dlLL11, OEA18]. Groups [CMO18, HKR17, HKR23a, PSVG18]. Growing [RELV11, WK18]. Grows [KS19]. Growth [AB10, AMW17, Ano15d, CS19, CC19, CMSV18, FF11b, FLP10, Gas16, GS13, Goo12, Gre16, HH15b, Lan11, LPS12, LYZ11, Man11, PZ15, RESA10, SY12, SYZ13, TWT14, XTpXpH12, ZW10, ZL13]. GRW [GTZ12]. GTW [Tos17]. GUE [FF14, Fer18, FL16b]. Guest [Sam19]. Guests [Ano18a, Ano19a, Ano19b]. Gulari [SM12a, SM12b].

H [Kie11, Sta11]. H. [Ano18d, bA10]. Haff [Gri19]. Hagen [Mar11b]. Half [Bak10, Ber12, Ber15, BNY16, BS15b, Bur17, FN15b, JL17a, LLS17, Lie12, PZ19, QR13, TW13a, Yam17, Yan11]. Half-Century [Bak10]. Half-Integer [FN15b]. Half-Line [JL17a, Lie12, TW13a]. Half-Plane [BS15b]. Half-Space [Ber12, Ber15, LLS17, Yan11]. Halfplane [ANS18]. Hall [RSY14, Yam17]. Hamburger [SP18]. Hamilton [KK14, Kra16, Lef13, MN12]. Hamilton-like [Lef13]. Hamiltonian [AU15, CFG13, Dym15, Grm10, HT19, Ito19a, Ito19b, Kle15, LK18, LY10, LQY17, Mia11, MAPS11, RW14, RNDS13, RDNS15, SZS15, Tur13]. Hamiltonians [KC19, Sch12c, SW15, dVO15]. Hammersley [CG19, PdS17]. Hard [AP14b, Bae11, BT14, CFLV19, Fan17a, HL18, MGAPQH13, MdSB18, Mie18, MPS14, NZB16, NKR15, QLCL16, RG17, RM16a, Sim14, Tri14]. Hard-Core [AP14b, HL18, MPS14, NZB16]. Hard-Spheres [Fan17a]. Hardcore [MS12c, RK12]. Hardy [FVV15, FVV14]. Harems [OV15]. Harmonic [AGR19, Art19a, BO14, BCFS17, BJP17, BKLL12, BR13, Fid15, GS11b, GLBP12, JPS17, Luk14, LMN16, NSS12, PZ19, TZ16]. Harmonically [Bur11]. Harnesses [BW17]. Harper [JL17b]. Harris [CG13]. Hartree [CR11a, CLS11b, DRS18, FK11, Laf19, Lee19]. Havlin [For11]. Heat [Bar16, BBR^{+}19, CP10b, CDS17, CEGW18, DDN14, DG17b, FGP15, GPMEA17, KNiST11, KOSV18, LZ10a, LY13, LNY16, Los17, MN14a, MGMMP13, MS11b, NiS19a, Nán16, ST11a, SS19, Tay17, WZIG14, vHL13, GPMEA18]. Heat-Bath [CEGW18]. Heat-Type [KOSV18].

Heated [FCK15]. Heats [KKA19]. Heavy [AHDV17, Ast12, FM18, Kaw16, Nán11, Pir14]. Heavy-Tailed [Nán11, Pir14]. Hedging [CC18]. Hegselmann [WLEC17]. Height [CC19, Chh12, CN13, DG15b]. Heights [dHNT11]. Heilmann [Alb16]. Heisenberg [AGR19, BU18, CG12b, FL16a, HA19, KM13, MD10, OK14, QD12]. Heitmann [DF17]. Helmholtz [MMSY11]. Helmholtz-Potential [MMSY11]. Help [Sch15]. Hénon [Tak16]. Hénon-Like [Tak16]. Herbert [Ber19]. Herding [BCKY19, ByChL^{+}19]. Herds [OV15]. Hermite [ST18a]. Hermitian [Afa16, Afa19, BBLP12, BBLP13, Koz17, PMC15, PS14, Wan12, YB14]. Heston [AG15b]. Heteroclinic [GBTL17]. Heterogeneities [dCFC11]. Heterogeneity [JMH13, LLS13, Zho17]. Heterogeneous [APZ19, ARBJ15, BPF^{+}14, KA17, LKR^{+}11, PSVG18, RT17a, RG18, Tou12, WZL^{+}14]. Heteropolygonal [CS19]. Heteropolymers [Ark10]. Heuristic [DARM^{+}13, Kie17b]. Heuristic-Based [DARM^{+}13]. Hexagonal [ADS^{+}19, FS19, GL16b, LSY19]. Hidden [AG15a, HA13, MSS11b]. Hierarchical [ACL14, BGJ15a, CBG14, Cas14, CG16, Cla19, CG13, GJ15a, GdHMZ19, GO11, Ker13, Pos16, SSE15]. Hierarchy [CIM14, DARM^{+}13, Fre14, GS12, Mas16, OV15]. High [AB19, AD14, BCM10, BS16, Cla19, FBR19, HvdHH14, KNPF19, Mar18, MFLA15, NT18, NKR15, OCM15, PZ17, RBM^{+}18, SH16, TM18, Wei18, WG18]. High-Activity [NKR15]. High-Contrast [PZ17]. High-Density [KNPF19]. High-Dimensional [HvdHH14, TM18, WG18]. High-Order [FBR19]. High-Temperature [Cla19, SH16]. Higher [CM11, Gao18, HDP17, Kaw16, KOSV18, MU18, Sch10, SLdEC11, WLJH18]. Higher-Order [CM11, HDP17, KOSV18, WLJH18]. Highlights [TM18]. Highly [CCP16, CdLS13, EF18, Kuo15, Per10, RF18]. Highly-Clustered [RF18]. Highway [EK19]. Hilbert [WK18]. Hill [Ano12a, Ano12b, Ano18a, Ano19a, Ano19b, Leb12a, Leb12c, Leb13b, Leb13c, Leb14, AB10]. Histogram [Koi10]. Histories [EK10]. Hitting [ACG15, BLM13, CNZ17, DMS17b, LS16b, MS19b, NZB16, RT15]. Hofstadter [AEG14]. Hohenberg [Hal19a, Bri19, Gar19, Hal19b, Leb19]. Hölder [Pir18a, TM10]. Holes [LNP13a, Mar16]. Holographic [GLBP14]. Holomorphic [DMS12]. Holstein [Miy12b, Miy16a]. Homoclinic [MT16]. Homogeneous [BB15, BE16, CL19, CP15a, CR11a, DFR14, He14, Kar10, KLM17, LU19, Lu12, MRCJ18, MWY16, Tri14, Xue12]. Homogenization [Bou18, DP17, LW19]. Homopolymers [WCX^{+}11]. Honeycomb [AEG14, MS16, YYZ11]. Honor [Ano18a, Ano19a, Ano19b]. Hoover [FG11]. Hopfield [AMT18, LV11]. Hopping [Gol10, Lee12, NiS19c]. Horizons [Det12]. Horizontal [CG11c]. Horng [Ano19b]. Horng-Tzer [Ano19b]. Host [Ara11a]. Hot [LZ10a]. House [Hui17]. House-of-Cards [Hui17]. Howard [Ano16b]. Hubbard [HT19, Miy12b, Miy16a, TT16, Tan18]. Huge [DHL^{+}17]. Hulthén [AFCA16]. Human [AFS^{+}13, HSUG13, HBC^{+}15].

Hurwitz [Nov15]. Huse [Ano19a]. Husimi [JJ12, JJB14]. Huygens [MK19]. Hybrid [CF19a]. Hydration [Bec11, WBL11]. Hydrodynamic [BIM18, CMGP14, CL16, DGLP15, DO15b, DSZ17, DOR15, DOGK16, Eri18, FT19a, HK18, HT11, MPL^{+}16, OR15, ROS19, RG17, Rue17, Sim10, Sta15, YBF^{+}17, dZS11, dZS13]. Hydrodynamical [CR14, Mas16]. Hydrodynamics [BCNS12, DUU15, DS19b, Hon19, MOT14a, OR19, Spo14, SS15, Yam17]. Hydrogen [LMC11]. Hydrophobic [Dzu11, VCT11]. Hyperbolic [BNTT16, BBC18, BK17a, BKLO19, CNZ17, CYZ18, FG14, KT17, Mih10, Mih19, ORS16, STBT10, dSLPV17]. Hyperbolicity [DFL17]. Hypercube [Jov17, Sco11]. Hyperdiffusion [Kaw16]. Hypergraphs [DA18, LN13]. Hyperparabolic [KP11]. Hyperscaling [Sak18]. Hyperuniform [CLM15]. Hypocoercivity [DL12]. Hypothesis [Det12]. Hysteresis [EKD12].

Ice [CR11b, Cug17, Kir12, Tho11, TW17]. Ideal [CIM14, FDR12, JK12, KT11b, Moh17, TKK15]. Identical [SB16, Sha12]. Identifiability [FC13]. Identification [Fil16]. Identifying [AWM13, OVC14]. Identities [SB12]. Identity [CLTT13, OR16]. Ih [CR11b]. II [Ast12, BBMW10, Bax12, BL12, BCL10b, BA19, BYYY19, BS15e, DMY13, DS19b, ESPP^{+}14, GMT17a, HT15, LK11, Mit17b, Miy13, RDNS15, SJ10, dZS13, vBSB15, Szá12]. III [Ano14b, AMS14, Ast13, BBS15a, Ham11]. Ikeda [Sta11]. Imitation [DP15]. Immediate [vGRS16]. Immersed [BFM10]. Immigration [DDC18, Ohk14]. Immortality [GS13]. Immune [FAB16, HZS11, ZW10]. Impact [BG12, LW18, LL10]. Impacts [ZGH18]. Imperfectly [AL18]. Implementation [BR11, Cli10]. Implications [AKL18, Bat17, EP14, GPMSBBSV15, LS12, MV16, PSC18, Ren18, VCT11]. Implies [ALS18, Ste10]. Importance [FJLS18, KR16]. Improved [Bae11, dBGP19, dLP14]. Improvement [TM18]. Improving [RBS16a, ZYZ^{+}18]. Impulsive [AC14a, BFVZ10]. Impurities [EV14, MS16, NS10]. Incipient [HvdHH14]. Including [BM18, DOGK16, JMH13, KT12, Lia18]. Inclusion [CCG14a, GRV11, OR15]. Incoherent [HX15a]. Incomplete [Lam19]. Incorporated [QLCL16]. Incorporation [HS10]. Increment [BS17]. Independence [NSW13]. Independent [Afa19, ORSV15, Ren18]. Index [BCS18, Zha14]. Indirect [ABT11, BFFS16, PVCG11]. Individual [CFM14]. Indoor [JMSW13]. Induced [ByChL^{+}19, BT12, BCMP15, BGTV11, CP10b, Cra11, GVJ^{+}18, HVW12, KTJ10, KC18, NML^{+}11, Sam19, XJZY13, ZWGM15]. Inducing [SB15a, SZ19]. Inelastic [BPR14, CH14, GR12, Ily16, PR19, WZ12]. Inequalities [APS12, BLU16, CCR17, Dav11, FVV14, FVV15, FF11a, FPR11, Gan18, GRV10, Miy16b, PR15b, Tos16]. Inequality [Han15, TV16, dBGP19]. Inequivalence [VB11]. Inertia [DOGK16]. Inertial [AM12, BBMW10, CFL17, LL19]. Inertial-Range [AM12].

Infection [Gra13]. Infections [MMR18]. Inference [AG15a, CSC14, FD16, KR16, Yed11, Zha12b]. Infinite [ASA15, AK17, APS12, And19, Asa13, AP18, BvdHK19, BCFS17, Ber14, Bha15, BS15c, BIP11, BM16, Car11, Chu18, CG12a, DP15, DP19, FGGL18, GGP10, GLO10, GL13, HvdHH14, Ily16, ISZ17, KM19c, LN19a, LVE19, MC16, MC21, Ngu18, OC12c, ORW15, PL13, PC19, SP13, SYZ13, SZ12, Thr18, Yag16, You17]. Infinite-Dimensional [ORW15]. Infinite-Order [MC16, MC21]. Infinite-Variance [BvdHK19]. Infinite-Volume [KM19c]. Infinitely [BFM10, BCM12, CCM16, SS18a]. Infinitely-Fast [SS18a]. Infinity [KV15]. Influence [BGTV11, CF11, GG11a, KKC12, LK11, Los17, RW19, WLJH18, dAPdA^{+}13]. Influences [FW15]. Influential [BHMGM13, HH15a]. Infomax [RCV16]. Information [BFFS16, BHS13, BCS18, CGL17, CGL18, DUU15, FAB16, Han13, HL16, HBC^{+}15, HH15b, HCLR11, LO17, LN19a, MMW16, MLS16, MPR16, MHD17, PSCD13, RL11, RBR11, SKT16, SLST13, SG17, TN13]. Information-Theoretic [LO17]. Informational [MC17a]. Informations [Riv16]. Inheritability [Qia10]. Inhomogeneities [Wei13]. Inhomogeneous [ACCG19, BT11, BDL11, DGGvdH18, FZ12, Fid15, Fun14, GN19, IvRM15, JvLS19, KS13, KMO16, LWL^{+}18, MSLT16, Nán16, Ohk14, RZ17, Sch13a, She15, vdHKvL18]. Initial [Abr13, CTM13, DSZ17, KOSV18, Lee10, LSY18, Los17, Mol17, Sta15, TW10]. Innovation [TM18]. Input [BC15b]. Insights [FBE^{+}11, NSW13]. Instabilities [HKN16, MvS19]. Instability [CSL19, DOGKP19, FZ11a, GPGA17, HX15a]. Instantons [BLZ14]. Institutional [Lee18]. Insurgent [JMH13]. Integer [FN15b, Kar10]. Integers [ALS14, Kos13]. Integrability [HCO15, YS13]. Integrable [BCP13, BPP18, CP15b, DMS17b, Kel19, LK14, Van17, Yam14, MP18b]. Integral [CSC14, EN18, HP11, Kle15]. Integrals [Fra11, Nig15, Zuc11b]. Integrate [Vid17]. Integrate-and-Fire [Vid17]. Integrated [AZ11, AB18b, Sch12c]. Integration [BS15d, Fan16a, Fan16b]. Intensities [BKK15]. Intensity [Det18]. Intensive [GP10]. Interacting [BGN^{+}17, BCJP19, BCF17, BCF19, BFNZ11, BCM12, CSL19, CMGP14, CF16, CG16, DF18, DGLP15, Dor16, DOR15, FF10, GL13, GRT17, GRV10, Hag15, JSV19, Lee11, Lee19, LSY18, Luç17, Mas14, Mor12, Ohk10, OR19, PT14a, RE13, SV15, Sit11, SH12, TD19, Yag16, LSBS13, Pal11]. Interaction [AN15, BBGS19, CC18, CA18, CCM16, Chu18, CMSV18, DM18a, FdHM14, FK11, GS11a, HP11, HSFK18, IST15, Jin18, JJB14, KPS19, KA17, Kra16, MN14b, Ono11, OEA18, PSW17, Pic10, PRSS17, PSS12, PST13, QD12, SGS19, SST15, TLC13, TAG10, CMVW11]. Interactions [AEWI14, ADP14, BD16b, BD17, BCPS18, BPRT14, CR11a, CR11b, CDdS14, CDS10, CLP19, DOGK16, EAL12, FS17, GJMS10, Gar19, HFWT15, KU11, MN16b, MG17, MS19c, MPT19, NZB16, Pes14, VCT11, VW18, Yar14, ZB13, Zho17]. Interactive [Gup16]. Interchange [SS18a]. Interdependent [ZB13]. Interface [ADMS11, ANSW14, BD16a, BDT17, CGG^{+}11, Dzu11, HMN15, Lim16]. Interfaces [AGGL^{+}16, AN19, BCY16, CCH16, CDV17, EV14, RRW11, Sch12a, SvHM^{+}11, TS12]. Interfacial [BBD^{+}11]. Interlaced [FF11c]. Interlacements [dBP15, dBGP19]. Intermediate [BV15, PM17]. Intermittency [BA19, GLST16, SB14a, SB14b]. Intermittent [ASA15, GVJ^{+}18, Kos11, MNBC19]. Internal [Ply15, WZIG14, XD18]. Interparticle [CIM14]. Interpolating [AP14a]. Interpolation [FS11a, LO18]. Interpretation [Ero14, SCY^{+}12, Sit11]. Interpreting [EK19]. intersecting [Ken19b]. Intersections [MP14]. Interval [AK14a, BLS11, CG12a, DJRZ11, GLBP12, KM19a, ACL^{+}11]. Intervals [Bal14, Kol17]. Introduction [Bré10, Gut11, WP11a, Gut12, Kar11]. Intuition [Lee18]. Invariance [BCL10a, CTH^{+}11, CYZ18, CMM16, DTW19, Hay15, Kel19, Ken15a, Ken19a, dWL14]. Invariant [ASA15, APR15, AC14a, BFVZ10, BF12, CL15, CT10, CFG13, GRR17, GKLS17a, GKLS17b, GLBP12, KC19, KR15, KRK14, LM13, MR13, RM16b, RS13, SVW12]. Invariants [BV16, Ber18, GPP12, Her13, MP14]. Invasion [DHS18, Goo12]. Inverse [CSC14, CM12b, FG14, FD16, KM19a, Lau18, Mas13, MP10, Mih10, MU13, Ter13]. Inversely [CL14]. Inversely-Quadratic [CL14]. Inversion [Han18]. Investigation [LK18, SdlL18, Yin16]. Inviscid [BL18, LT10b, Mol17]. Invisibility [IK10]. Involvement [RM16a]. Involving [FN15b]. Ionic [HP11, MMM15, MW10, VCT11]. Ionic-Interactions [VCT11]. Ionization [CCL19]. Ions [QLCL16, ST14, Sam15]. Irreducible [BCHW19, Gil15]. Irregular [RZ17]. Irreversibility [BDDH14, HiS19]. Irreversible [DSS15, Des11, KJZ18, MN15, Tay17]. Ising [ANS18, AMS14, ART11, ADMS10, ACL14, ADS^{+}19, AS16, BCM10, Bax11, Bax12, BC10, Bot18, Bra14, CA19, CJN18, Can17, CSL19, CMP17, CGNP11, Cha19, CV16, CM12b, CGHT16, Con18, CvELR18, CLP19, CDV17, DSS15, DK10, DGvdH10, DGGvdH18, ESPP^{+}14, EKD12, FMMPS18, GRS12, GRRR13, GHRR13, GLML16, GGvdHP15, GNP16, HI18, Har11, HTZ12, IST15, IL11, Ito18, Jov17, JJ12, JJB14, KLS19, KLM13, Ken10, KE10, KY12, LS13, LO17, Lis17, LSS19, MW12b, Mas13, MS16, MBS16, NT16, Nig15, Ny13, dAPS11, PSS16, RS15b, SH16, SSBS14, TW14, VAY^{+}12, WW16, WZIG14, Wu14, Wu18, Zha12b, vEKRS19]. Ising-Like [GLML16]. Iso [VFT12]. Iso-Contours [VFT12]. Isogenetic [Qia10]. Isoheight [CTH^{+}11]. Isolated [DI13, JYZ11, Tas16]. Isolation [Det18]. isometries [CMM16]. Isoperimetric [APS12]. Isotherm [RM16a]. Isothermal [BC12, LNP11]. Isotropic [BB15, CRV13, PRD11, SSR12, Sch13c]. Isotropically [ISZ16, ISZ17]. Issue [Ano11d, Ano11e, Ano15d, Ano18d, GvdHdHM18, JP18a, Leb19, MPR16, WP11a]. Iterated [MU16]. Iterates [Mih10]. Iterations [Cie17]. Iterative [Han18]. Ito [SCY^{+}12]. IV [BS15f]. Iyetomi [Sta11].

J [bA10]. J. [Ano18d]. Jacobi [HF12b, Kra16, MN12, WWKK16]. Jacobs [San12]. Jamming [BJGL^{+}17, OiS14]. Jan [Ano14b]. Jancovici [Ano15b]. japonica [CLTC15]. Jarzynski [HSZ19, HH15b, MT11a]. Jarzynski-Type [HH15b]. Jaynes [ME14]. Jean [Kie17a]. Jellium [Kie14, Sam19]. Jellium-Like [Kie14]. Jennifer [Ano18a]. Jet [BNT13]. Joel [Ano18c]. Joined [FL16a]. Joint [KBSM16, PGS12]. Jones [Cam13, CVE14, FC17, JZ10, MPS14, ROS19, Yuh15, dLP14]. Josephson [NML^{+}11]. Journal [GvdHdHM18]. Journey [Kha17]. JSP [JP18a, MPR16]. Jump [BCHM12, BKK15, DD15, Kra16, MGMMP13, ZL19, vKSZ18]. Jump-Diffusion [DD15]. Jumps [CFN15, LPS19, PS10a]. Junction [NML^{+}11].

Kac [AW18, ALS14, BLV14, BPR14, CMW15, CMV11, Cor16, DP17, Ein12, Eva16, FdHM14, GR12, HI18, Hau16, Ono11, Tos17, ZLL13]. Kac-Type [FdHM14]. Kadanoff [AB17, Bat17, Weg17]. Kamenev [VL12]. Kampen [FvB14]. Kapral [Ara11b]. Kardar [Ano15d, Ano19a, CG17, FSS13, TS12, TA16, XTL14]. Karlin [Hui18, JL17a]. Kasteleyn [CEGW18]. Kaufman [Bax11, Bax12]. Kautz [Kla11]. Kawasaki [FT19b, MdG13, dHNT11, dHNT12]. Ken [Mer14, Pes14]. Kenneth [BKPW14, Gin14a, Gin14b, Uka15, Weg14]. Kernel [CL18, MNV11]. Kernels [AP14a, Lau18, NV14]. Kesten [LJN18]. Key [GI19]. Kicked [Cla13]. KIF1A [Zha13]. Killing [Kat12c]. Kimura [AP11]. Kind [Pet10]. Kinematics [CG11b]. Kinesin [Zha13]. Kinetic [Ark13, BCKY19, BLT11, Bec10, BCM16, Ber15, BD17, BGP15b, BGZ17, BP19, BCE^{+}14, BNT13, CEL^{+}18, CCH^{+}14, CIM14, CG16, Dai17, Dal11, DM12, DM13, DFR14, DT19, FS14a, FKLL16, HKLN19, Ily12, KTJ10, Ken15b, Kha19, KV16a, KKA19, Kuo17, LL13, Leb11, LMN16, MS10, NWL19, PSW17, Pav11, PR15a, dAPS11, RS15b, Sch13c, SVRL11, Sit11, Tak15, TN18, TN21, TBD13, VAY^{+}12, Vau10, Zha12b, KRBN10]. Kinetically [BLT12, OiS14, STBT10]. Kinetics [AKH13, Gup16, MOT14a, RBGV12, SH12]. Kingman [BCL19]. Kinks [Raz18]. Kirkpatrick [Con13, BL16a, BSS19, Pan12]. Kirkwood [IiS15a]. Klein [GOP14, TFES19]. Klimontovich [GRT17]. KLS [Zia10]. Knot [RMN15]. Knots [dCCS19]. Know [AWE16]. Knudsen [CPSV10, HT15]. Knudsen-Layer [HT15]. Kohn [Gar19]. Kolmogorov [MDP^{+}18]. Kondo [BGJ15a, GJ15a]. Kontorova [BdlL13]. Kontrova [SdlL18]. Koopman [DG19a]. Kosterlitz [EC11]. KPP [BD15a]. KPZ [CES19, CQR15, FF11b, GJ15b, HHT15, IS13, KM17, MU18, MP18a, QS15, Sch12a]. Kramers [Bir18, HVW12, KP12, Ryt12, TFES19]. Krapivsky [Leb11]. Kratky [KMS19a]. Krause [WLEC17]. Kubo [GM13, IiS15a, Mat12]. Kuramoto [BCM16, BGP10, BC17, CCH^{+}14, DeV19, HX15a, HKR16, MG19, MiS18, dAPdA^{+}13]. Kurt [San12].

L [Leb11, CPS19]. L-Shaped [CPS19]. L. [bA12]. Ladders [Sep13]. Lagrangian [EN18, Mol17, RW14]. Laguerre [Kum19, Zha15]. Lamellar [KT11b]. Laminar [Che14, dZS11, dZS13]. Laminar-Turbulent [Che14]. Landau [RBS16b, BGP15b, BGZ17, BP19, CFK13, FH11, KBSM16, KWZ14, LY16, MN16a, Ser14, Tri17, WLL11]. Landau-de [KWZ14]. Landauer [MLS16]. Landscape [BBK17, WLJH18]. Landscapes [CN13, Cro12, FK12, HSFK18, ME11, MMSY11, SW11, SK14, ZW12]. Lane [Gup16]. Lanford [Ano14b]. Langevin [Bir18, BLZ14, Bur19, Cam13, CL13b, CF15, CWD17, CLL18, CL18, DS19a, DLP16, DNP17, FG11, Hag15, HHV16, JPS15, KSSH15, KK15, LW19, MSZ12, Ngu18, RST16, VV17, ZWGM13]. Langevin-like [KSSH15]. Langmuir [Gup16]. Languages [FAM13]. Laplacian [AMW17, AZ11, FZ10, Fre17, LYT16, LYZ11, Mit16, Mit17a, Mit17b]. Large [ADE18, ARS17b, AGO18, BP12a, BC15a, BG11, BBC18, BGP15a, BCCD18, BDT17, BL10c, BLZ14, BGTVE16, CT13, CFL17, CF15, CGP17, CF19b, CG12b, CG17, CMV16, DF16, DR13, DS16, DS19a, DS19b, DGGvdH18, EPS17, FSV10, FZ11b, FT18a, FT18b, FG12, FD14, GRT17, GTT19, GM13, HR18, HŚ16, IiS15a, Jan15a, KRRS17, KL15, Kie17b, KS12, KO15, Kra16, LS15a, LLM19, LN19b, LNT13, LN15, Luç17, MK19, MPM17, MC11, Mou15, OR19, PSK10, Pat17, RS15a, RW12, Ren18, Sch12a, SMCF13, Shi13, SGL15, TGP12, Var12, Xue16b, YB14]. Large-Degree [Kie17b]. Large-Deviation [MK19]. Large-Dimensional [Xue16b]. Large-Scale [GRT17, MPM17]. Large-Time [HR18]. Largest [Kar14b, KLM17, Wan12]. Laser [HV14]. Last [Bak10, Cal15, CQR13]. Late [MP18a]. Lattice [ABMP16, AEG14, AK14a, AP14b, ACL14, BCM10, BS11, Bet14, Bis19, BEP18, BBR^{+}19, BM12b, CE12, CP15a, CC14, CS15, CIM14, Cha14b, CCR17, CLMK18, CL18, CCR19, Cla19, CSS15, CLT19, DO14a, FSV10, FS19, FCK15, GdHMZ19, GdK10, GLU12, HF12a, Hag13, HL14, HA19, ILOS10, IST12, JD11, JJB14, Kac13, Ken10, Ken19b, KNPF19, Kor16, KNSS18, KSM16, KM19b, KM19c, LYT16, LMN18, MPSS19, MD10, MPL^{+}16, Mas14, MM14, MN14b, Mia11, MFLA15, MCK15, Mor18, Nig15, NE16, O’C12a, Par17, PM17, PB11, RZ17, SS11a, ST18b, SC10, Sim10, Sug10, Tas19, Thi16, Uka15, WGLE11, Wat19, Wat20, WLL11, YH14, ZDG19, Zha14, ZSHL15, dVO15, Ark10]. Lattice-Based [Cha14b]. Lattice-Gas [Ken10]. Lattices [BF11, CS19, JJ12, Kar10, KKSY19, MS16, STBT10, TW11, Xue16a, YYZ11, Zoc18]. Laughlin [LRY18]. Law [AM19, AF14a, ACCG19, AGMM^{+}12, BBR12, Ban10, CNZ17, DGvdH10, Ero14, Far15, FG12, GNP18, GS11a, GQ17, GDL10, GTT19, HŚ16, JS17, Kim12, Kol14, Lau18, MPSS19, MS14, PK11, PSCD13, Sch13b, Sod09, Sod17, SW12, vdHKvL18, vdHLK18, SM14, DOP19, Gri19, IiS15a, Lee18, Ngu19a]. Laws [AFFR17, BBC18, FFT11, GS11c, GW15b, LS14a, dlLOP11, GvdHW17]. Lax [MS10]. Layer [HT15, KP18, Yan11, CES19]. Layered [FMM^{+}14, FMM^{+}15, NKK19]. Layering [ADMS10, ADMS11]. Layers [CDS17]. Lead [FN13]. Leadership [OV15]. Leading [Abr17, DM10]. Leaky [Vid17]. Learning [AK16, DFF18, LTR17, Nei12, OCM15, SGC11, Mon12a]. Lebesgue [BW10]. Lebowitz [Ano18c]. Left [AFR19, Hen12, IP12]. Left-Passage [IP12]. Left-Permeable [AFR19]. Left-Right [Hen12]. Leidenfrost [PKBW19]. Lemma [AP14b]. Length [CFMT17, CG12a, ELO11, GL13, Ken12, KSH11, MTVU18, MS14, NiS19b, NP12, Sep13, Thä11, WBL11, ZL19, Zho17, Zho18]. Length-Scale [WBL11]. Lengths [BFL18, Caë10, GLU12, OBX11, Sha12, Wen12]. Lennard [Cam13, CVE14, FC17, MPS14, ROS19, Yuh15, dLP14]. Lennard-Jones [CVE14, FC17, MPS14, ROS19, Yuh15, dLP14]. Lennard-Jones- [Cam13]. Leo [Weg17]. Leonard [Gou14]. Letter [Ano18c, HM13a]. Level [BC15a, IM16, Nak14, Per13, Sug10, WL13]. Levels [Tou15, YS13]. Levelset [BM17]. Leveraging [BMC17]. Lévy [Abr13, AM14, BCLL16, BP15, CTT11, GW15b, IPS10a, LL19, MSZ12, XLL^{+}15, XD18]. Lie [Gro19, HKR17, HKR23a, Ohk14]. Lieb [Kie11, Alb16, Buc16, GL16a, IST12, LS16c, MR19, Tas18a]. Life [SC10, Sta11, AFI^{+}10]. Lifetime [LMM16]. Lifetimes [SW11]. Lifschitz [BBW15, CN14]. Lifshits [KKR18]. Lifshitz [BS11, BN15, FH11]. Lifts [Mih19]. Ligand [FAB16]. Light [BBGS19, CP10b, EK19, Lia18]. Like [AP18, BCHM12, Bra14, CN14, Fre16, GLML16, Kac13, Kie14, MPL^{+}16, SL12, CGR12, Han13, KSSH15, Lef13, MW12a, dCFC11, Tak16]. Likelihood [AG15b]. Limit [AM19, AF14b, AT18, ABA14, AWE16, BG17a, BL18, BPZ13, BF10, BNP14, BP11a, BGMS14, BGJ^{+}15b, BCF17, BCF19, Bha15, BIM18, BSS14, CG10a, CCEF10, CMGP14, Cha19, CL16, CR11a, CF16, CGY17, Chu16, Cla19, Col14, CFP10, CGR12, DG17b, DGLP15, DF16, DSZ17, DS17, DOR15, Els12, Eri18, FS18, FKKO15, FL14, FT19a, FG11, FS11b, GGvdHP15, HKN16, HNVZ13, He14, HR18, HHV16, HNT18, Hor16, HK18, HVW12, HMW19, JRS15, KL15, KKSY19, KS12, KY13, Kur18, Laf19, LLM12, LK18, LS14a, LWL^{+}18, LVE19, MU18, MBC14, NSV12, Ngu18, OR15, PC19, PZ17, PST12, RVY18, RE13, Sha12, Sim10, Sta15, Sun18, TWT14, Tos16, Tzi18, Vau10, Wre12, Xue15]. Limitation [Rab11, Tho12]. Limitations [Hal19b, LKR^{+}11, TM18, Hal19a]. Limited [CC19]. Limiting [ART11, BP12a, BGN^{+}17, CFN15, DMP17, FK18a]. Limits [BBS14, BJS17, BCCD18, BP16, CCP16, Coq15, DO15b, FC11, Fra17b, HNZ16, IY14, Kem11, KV17, Tel10, Tou12, tWBOM16]. Limping [ZF11]. Linda [Täu10]. Lindblad [AFG12, CM12a, Gou15, MM17b, Pel14, RS13]. Line [BvE11, BBD^{+}11, CDH15, JL17a, Lie12, LN11, NT16, QR13, TW13a]. Linear [ADF18, BBMW10, BPZ13, BC19, Des11, DJRZ11, DL12, GMT17b, GMT17a, GLT15, HLZ17, He19, KK16, KS12, LNP13b, LNS^{+}12b, Luc18, MT17, MN16a, MC11, Ohk14, OPS10, PWZ16, PMC15, PZ15, SG15, TA12, WFK11a, WFK11b, WG18, ZWGM13, BCF17, GS13, IPS10a, LWY18]. Linearised [ST18a]. Linearized [Ber15, Che13, Dud13, LWW18, LY16, Sha10, Tak09, Tak10a, Tak10b, Tri17, Wu15]. Lines [BW18, CSL19, CTH^{+}11]. Linguistic [Ero14]. Liniger [Buc16, MR19]. Link [Luc16, TSS13]. Linking [Mar11a]. Links [AK14a, BH11, JZ10]. Linnik [GW15b]. Lipofuscin [FMAG11]. Liquid [Alb16, BG17a, DOGKP19, FY16, GPGA17, KWZ14, LRY18, Ono11, SGL15, TS12]. Liquid-Crystal [TS12]. Liquids [HJR^{+}11, MOT14a, WP11a]. Lists [Ano16a]. Lived [BCFS17]. Lives [HBC^{+}15]. Living [Ami10, LKD12, MPR16, Rab11]. Load [DA18]. Loading [RG18]. Local [AEK16, AP14b, ABT^{+}14, AKL18, BAS18, BIM18, BE16, BS15e, CES19, CS10c, CF19b, DLR14, DFR14, DO18, EF13, FL14, FN13, HM13a, HNT18, KL13, KD19, Lan13, LNY16, Mih11, Nán16, Pim18, QLCL16, Shc11, Shc14, Shi13, SW12, Tas18b, Tas18a, TV15b, Zha15, BBGS19, RVY18]. Localisation [CMSV18]. Localization [ABJ12, CE12, CWD17, Chu18, DHK11, EFO11, EH12, GPS13, GPD16, HKW11, Imb16, JM10, Mor18, MC10, PSC18, PdOC17, Sep13, Tau11, Ast13]. Localized [AI12, DDHS17, KW15]. Locally [AP18, PS10a, SV15, Tau11]. Location [HRW14]. Locked [HR16, HKR16, HR23]. Loewner [Ghe10, GN13, JL18, Ken15b]. Log [AKD19, BDY17, CF19a, DG14, FL16b, MK19, Sam13]. Log-Correlated [FL16b, MK19]. Log-Gas [AKD19, BDY17]. Log-Gases [CF19a, Sam13]. Log-periodic [DG14]. Logarithm [GNP18]. Logarithmic [DLBK11, KNSS18, Leb16, Sam18]. Logarithmically [Web11]. Logical [XJZY13]. Lohe [HK19, HR16, HKR17, HKR18, HKR23a, HR23, HKR23b]. Long [AHDV17, BCFS17, BC11, BP19, BPRT14, CMV11, CMP17, CF11, CDdS14, CDH15, CvELR18, CDV17, GDL10, Gra13, JSV19, KS12, KL19, KiMM13, LSW17, Lu16, MN14b, Mis15, Mis16, Mis19, Miy16a, Mor12, Sch12c, SV15, Tas19, VW18, Wei16, vEdLV16, vEKRS19]. Long-Lived [BCFS17]. Long-Range [AHDV17, BC11, BPRT14, CDdS14, CvELR18, CDV17, KL19, LSW17, MN14b, Mis15, Mis16, Mis19, Miy16a, Mor12, Sch12c, Tas19, vEdLV16, vEKRS19]. Long-Time [CF11]. Longevity [LKD12]. Longitudinal [Ito18]. Loop [AKM13, AO19, FL13, HWZ19, KLM13, Ken19a, PSAPR12]. Loop-Erased [AKM13, AO19, Ken19a]. Loops [BU18, EKD12]. Looptrees [BS15c]. Lorentz [BP11b, BCKL12, CG10a, Det12, Koi18, KS14, KM19b, MN16a, MS14, SLdEC11, Sor18, TA12, WWKK16]. Lorentzian [KSY13, KY12]. Lorenz [BM18, GNP18, GL14, Lia19, PT10, TLD18]. Lorenzo [Mon12a]. Loss [EK10, Koi18, dAPS11]. Losses [BA14]. Lovász [AP14b]. Low [AM19, ABS12, Ark13, BS13a, BL11a, BV11, BD15b, CFMT17, Con11, CRL15, CE19, DSS15, DUU15, EK10, HPF15, IST15, JK12, Jan12, Lim16, MdG13, PSS16, RW11, SSR12, Zoc18]. Low-Noise [AM19]. Low-Temperature [Con11, EK10, PSS16, Zoc18]. Lower [Bou13b, CGP17, CFG13, JL17b, Pro17, TS19]. Lower-Dimensional [CFG13]. Lozenge [Nov15]. Lucky [Bri19]. Łukasiewicz [BIP11]. Luria [KL15]. Luttinger [KPS19, Lan10]. Lyapounov [Bou13b]. Lyapunov [ADE18, BDV12, BPP18, CLTT13, For13, Gal17b, GM13, HL16, JL17b, Kar14b, LSBS13, ORW15, RSB10, RF18]. Lyapunov-Based [RF18]. Lying [BV11].

M [Gou14]. Machine [Mon12a, SGC11]. Machines [DFF18, Nei12]. Machta [GOPS11]. Macro [Ste19]. Macro- [Ste19]. Macromolecular [Gas16]. Macroscopic [BDT17, BD15b, CL13a, Col14, DNBS10, DM11, Grm17, Hen12, HiS19, Lef13, MM17b, NS16, Ren18, SGS19, Tas16, Tay16, Van17]. Macrostate [Tou15]. Made [OV15]. Magic [NBK14]. Magnet [SSR12]. Magnetic [ACL14, AK14b, Bir18, BC10, BG12, CA19, CCM17, CF11, CS15, DF16, EH12, FVV14, FVV15, FW12, FR17, KE10, LL16, LLM19, MN16a, OK14, SSR12]. Magnetism [BC11, FBR19]. Magnetization [Bax11, Bax12, Fed14, KLS19, MSLT16, NT16]. Magneto [Yam17]. Magneto-hydrodynamics [Yam17]. Majorana [AG14]. Majority [BV15, LN13, SM15]. Maki [APRT17]. Making [Kie17a]. Managing [MMSY11]. Manhattan [Ken19b]. Manifold [CFL17, GRS12]. Manifolds [BFVZ10, BC17, HKR18, HKR23b]. MANOVA [EF13]. Many [BCM12, CL14, FSV10, Imb16, LH17, ML15, Miy19, MAPS11, SVRL11, Sim10, dVO15]. Many-Body [CL14, Imb16, MAPS11, SVRL11]. Many-Electron [Miy19]. Map [GBL16]. Mapping [PM17]. Mappings [DJRZ11, LK18]. Maps [ASA15, BGN16a, CRV17, CG12a, DG19a, Fer14, FO18, GLBP12, JPV18, Ken10, SB16, Tak16, Yoo10]. Marchenko [JS17]. Marginal [BM12a]. Marginally [Par17]. Market [BCKY19, ByChL^{+}19, KiMM13]. Markets [CC18, GNPS13]. Markets-Phase [CC18]. Markov [MC21, AMS14, AG15a, ADF18, BBMW10, BB11, BGN^{+}17, BCHM12, BL10a, BL12, BLM13, BKK15, BA19, BK11, CGL12, DS19a, DKY19, DPT17, FP11, FKR12, Fre14, FK17, Gas16, GKW12, HA13, HS14a, KJZ17, KJZ18, Kem11, KMKT11, Kol17, Kra16, Leh13, Li19, LNT13, LN15, Luc16, MC16, MSS11b, MDP^{+}18, MBS16, MGMMP13, PS19, RW14, RT17a, Sch10, TM10, TGP12, Vid15, Yoo10]. Markovian [Bao17, BHS13, Cra13, LPS12, LPS19, LK19, Pel14, PZ15, SML19, SGS19]. Martingale [BGMS14, CG11b]. Martingales [ALM18, Kat15]. maser [Bru14]. Mass [BD16a, Cen13, CF11, DL17, GS11b, Git14, HHV16, HCO15, Lau18, LS14b, LWL^{+}18, LCZW15, Ngu18, RT14, Thr18, ZWGM15, dWL14]. Mass-Conserving [BD16a, Lau18]. Massive [Lia18]. Massively [AK16]. Massless [DP14]. Master [CMW15, CM12a, KMTC10, LT10a, NT17, Pel14]. Matching [ACH15, BT19, del12]. Matchings [Per17]. Materials [CS16a]. Matérn [LM17]. Mathematical [GQ17, GLBP14, HSZ19, Kha17, KiMM13, SR13]. Mathematics [Dia11]. Mather [LN15]. Matrices [AB10, Afa16, Afa19, AEK16, AD15a, AO19, BP12a, BPZ13, BDES19, BBLP12, BBLP13, BYYY19, CRV13, CTT10, CLTT13, CFLV19, EF13, Fed17, FF10, For13, Gal17b, GMT17b, GMT17a, GO11, GO13, HM19, He19, ISZ16, ISZ17, JS17, Kac13, Kar14b, Koz17, May13, ORSV15, PWZ16, PMC15, PS14, PRS12, RKGZ12, SS11a, Shc14, Shc15, SS16, SS18c, Sod09, Sod11, YB14, Sod17]. Matrix [ABA14, BR10, BGP15a, Bir18, BN15, BS19, BMSS13, CLSW17, DZ15, FS11a, FL15b, FN15b, GS17a, HK19, IPS^{+}10b, KS13, Lam19, LM13, Mar18, MC19, MM13, O’C12a, PC19, PRS12, RT14, SS16, SW15, Wan12, ZJ14]. Matrix-Valued [BN15]. Matter [BDG^{+}14, Kie11, RS19, SB15a, LS10]. Matters [TQS12]. Mattis [Tas18a]. Maxima [FLTV11]. Maximal [HL18, KMS14, PST12, RGL11]. Maximizing [Lep12]. Maximum [AG15b, DJRZ11, FL16b, HA13, MN16b, MDP^{+}18, OCM15, Ost16, PCM15, SP18, Tay16, ZST18, vdHLK18]. Maximum-Entropy [ZST18]. Maxwelian [Lu12]. Maxwell [FL15a, CE14, Ily16, Kuo15, Liu15a, Moh11]. Maxwell-Type [Kuo15]. Maxwellian [LWY18]. Maxwellians [HX15b]. May [Ano12b, Ano18a, Ano19b, Leb12a, Leb12c, Leb13c, Leb14]. Mayer [Jan12, KM19a, dLP14]. McGregor [Hui18, JL17a]. McKean [CP10a]. MCMC [BSM^{+}16]. McMillan [BMSS13]. Mean [ACM15, AM18, AKL18, BG17a, BGL^{+}11, BGP10, BCF17, BCF19, BR16, BS16, BP16, BGN16b, BIM18, BV15, CHHS15, CGY17, CFS18, CDS10, CR16, CDL^{+}12, DvLM18, DPT17, ELO11, FC11, FMM^{+}15, FKR12, FT19b, GRT17, HKN16, IK10, KSH11, KM13, KN16, KOT11, Kra16, LS16a, Lan15, MNS12, MLCPS13, Mor12, PK10, PRSS17, ST18b, SB16, Sta15, VYH11, WGLE11, ZDG19]. Mean-Field [AM18, BG17a, BR16, BS16, BGN16b, BIM18, CHHS15, CR16, CDL^{+}12, DvLM18, FC11, FKR12, KM13, KN16, KOT11, Kra16, LS16a, MNS12, MLCPS13, Mor12, SB16, IK10]. Mean-Field-Type [WGLE11]. Means [BVL16, LS13, Mol14, PSS12]. Measure [ACH15, BK15, BR13, EHR12, FL13, Fra11, HWZ19, KM19c, MWY16, Per13, PSS12, PZ19, QLCL16, RW14, Shi13, Sod11, Tou15]. Measurement [AV16, Leh13]. Measures [ALS18, ASA15, ART11, APR15, ADC10, AC14a, AF14b, ADS^{+}19, AFGL15, AP18, ALS14, BFKR10, BPS12, BL10b, BGL14, BF12, CCR17, CG14b, CLP17, CT10, Coq15, CRL15, Det18, FO18, GLO10, GRR17, GKLS17b, GBL16, Han13, Hay15, HS14a, Jia14, KMSS19, Kie17b, KRK14, LNP13a, Lep12, MTVU18, MT16, Mih10, MU13, MU16, MSLT16, Pan14, Pan16, RvH14, RS18, SL12, She15, Thi16, Tid16, Tou18, Var12, Yoo10, You17, GKLS17a]. Measuring [IK17, ZF11]. Mechanic [Kie13]. Mechanical [Ero14, GDL10, Lia18, Luc18, NiS19b, RY12, WG19]. Mechanics [Ano11d, Ano11e, Ano12a, Ano12b, Ano15a, Ano16a, Ano18a, Ano19a, Ano19b, AN19, Bae11, BFFS16, BC15b, BG19, BG17b, CdlL10, CM11, CG11c, CTT11, Dym15, FC11, FCLK14, Gal15, GS17c, GTT14, IiS15b, KT15, Leb10, Leb12a, Leb12c, Leb13b, Leb13c, Leb14, LBB15, MSS11b, RVB16, Rue14, Rue17, TW17, XZ19, ZHZ15, Zia10, bA12, del12, LS10, Pel11, GS17b, Kie11, Kie17a]. Mechanism [BG12, BD15b, Gol10, KKV^{+}11]. Mechanisms [CA18]. Media [ACCG19, BdlL13, BBR^{+}19, CLMK18, GHS17, HS10, Luç17, MBWC16, NG10, RT17a, SJ10, ZSHL15, dlLSZ16]. Mediated [BGL^{+}11, HS10, RBGV12]. mediterranea [CLTC15]. Medium [BFM10, CR11a, SH12]. Meets [BB17]. Mehran [Ano19a]. Melt [HMRW13]. Melt-Blowing [HMRW13]. Melts [WCX^{+}11]. Membrane [FT19a, Sak12, Sch11, Yag16]. Membranes [AN19, CF16, UK16]. Memoriam [Ano15b, Ano16b, FvB14]. Memories [GHLV16, Mer14, Mer19]. Memorized [WZL^{+}14]. Memory [AG12a, AMS19, Ber19, CG12a, DP15, DHL^{+}17, GL13, GM15, Koi18, KiMM13, Leb19, RESA10, SP13, Weg14, Weg17, ZL13]. Menzerath [Ero14]. Merger [GVJ^{+}18]. Merging [LNP19]. Merit [LLM12]. Mermin [Hal19a, Hal19b, KSY13]. Merrier [QTDS11, QTDS12]. Merrifield [Zha14]. Meso [Ste19]. Meso- [Ste19]. Mesoscopic [AEWCD19, BC14a, BCFS17, BKK15, FKKO15, He19, Kha19, SSB15]. Message [Yed11]. Message-Passing [Yed11]. Metadynamics [ScCdP11]. Metal [DI13, Kor16]. Metallic [TT16]. Metapopulation [LLS13]. Metastability [And19, BRSV18, BL10a, BL12, CG10b, CNS15, Jov17, LMT15, LS16a, LL16, LLM19, MS19b, vEF12]. Metastable [AF14c, BD16a, BP15, CN13, FK17, dHNT11]. Metastate [WW16]. Metastates [FKR12, IK10]. Method [AG12a, AK14a, Ark10, BBS15a, BS15d, BS15e, BS15f, BS15g, CPV10, CJW17, CIM14, CLMK18, CS16c, DG19b, DJRZ11, FBR19, Fil16, FCK15, FD16, GRT17, HLZ17, LS14a, MW10, MCK15, OVC14, PM17, PB11, RLCMRT10, SSBS14, Sug10, YH14, ZSHL15]. Methodological [LVE19]. Methods [BBD^{+}11, CA19, GT15b, Jus10, KO15, LNP11, LVE19, ME11, Rab11, Tay16]. Metric [Bla10, BJR10, Hau16, HY17, Tos17]. Michael [Ano18a]. Micro [NE16, SSB15]. Micro-channel [SSB15]. Micro/Nano [NE16]. Microbial [LW18]. Microcanonical [AK14b, CO17, Fra11, Mor12, OK14, Tas18b]. Microdomains [AKH13]. Microscopic [BBR12, BS13a, BL17, BD15b, CD14b, FK11, HiS19, Kie14, Mon12b, RG17, Ste19, Tay16]. Microstructure [FBE^{+}11]. Microstructures [SWK^{+}18, SWKS14]. Microtubule [HE17, Zha13]. Microtubules [Gol10]. Midpoint [MT17]. Migdal [AB17]. Migration [FN13, HCI14]. Mineral [LKR^{+}11]. Mini [Per13]. Mini-review [Per13]. Minicircle [ADH12]. Minimal [BG17b, CDV17, FS14b, GS15]. Minimalistic [Kor16]. Minimization [DHX19, ESPP^{+}14]. Minimizer [DF17]. Minimizers [Bra14]. Minimum [FD14, MDP^{+}18, XZ17, ZHZ15]. Minkowski [JL16]. Minor [FF14]. Minority [BELP18, GPD16]. Minus [QR13]. Mirror [KS14]. Miscible [ZSHL15]. Mixed [CD14a, CP17b, DK10, EKD12, He19, KE10, LHZ^{+}19]. Mixed-Spin [KE10]. Mixing [ABR18, Bru14, CFMT17, CGHT16, DSS15, Dia11, Hay15, HY16, Hol11, Ker13, KOT11, KO15, LP16, Li19, MDP^{+}18, Mis16, SS18a, Yar14]. Mixture [DL17, Har11, JK12, RG17]. Mixtures [BV16, BCPS18, HCO15, Pir18b, Wil11]. Mobile [BRSV18, PKDK13]. Mobility [CC19, HSUG13, Los17]. Model [ANS18, AMS14, AKD19, APRT17, AGGL^{+}16, AEWCD19, ALS18, AI12, ART11, ACM15, Alb16, AM18, ADMS10, AMS19, AF14b, AT18, AM12, ALAF18, AK14b, AR11a, AN15, ABJ12, ABF16, ACL^{+}11, Aza11, BR10, ByChL^{+}19, BAC13, BL16a, BCM10, BF10, BELP18, BLS11, BSS19, BBS14, Bax11, Bax12, BCNS12, BAS16, BAS18, BLU16, BCM16, BGJ15a, BL11a, Ber14, BK17b, BS13b, BGP10, BCP11, BB13, Ber19, BLRVR13, BDY17, BEP18, BL10c, BD11a, BLT12, BJM15, BLV14, BCE^{+}14, BPR14, BPF^{+}14, Bot18, BBR^{+}19, BLZ14, BV15, BFL18, BDL11, BBK17, CA19, CJN18, CR14, CG15, CC18, Can17, CE12, CSL19, CP18, CTM13, CMV11, CMP17, CDG^{+}15, Cen13, CS15, CCD15, Cha19, CHHS15, CGY17, Che18, Chh12, CL18, Chu16, CRS12, CMM16, CDS10, Col14]. Model [CR16, CF19b, CGHT16, CEGW18, CS16c, CPS19, CRTZ13, CGG^{+}11, CLP19, CP15b, CMSV18, CdLS13, Cra11, CDL^{+}12, DSS15, DTW19, DNBS10, DM11, DFR14, DHL^{+}17, DO18, DH19, DR14, DvLM18, DD15, DPT17, DLLX16, Dod15, DGGvdH18, DP14, Ein12, ESPP^{+}14, EC11, EK10, Eva16, FKLL16, FSV10, FBR19, FKKO15, FL16a, Fri17, FW17, FAM13, Fyo19, FMMPS18, GR12, GL14, GJ15a, GL13, GRS12, GRRR13, GHRR13, GRR17, GMM18, GMM19, GOPS11, GDL10, GNS18, GGvdHP15, GNP16, Gor18, GV15, GL16b, GOdSS11, GV12b, HR16, HKR16, HKR17, HK19, HKR23a, HR23, HHM17, HF12a, HI18, Hal17, HTX^{+}12, Hau16, HMN15, HNZ16, HE17, HS18, HMRW13, HW13, HY19b, Ily12, Ily16, IS16, Ito18, IPP14, JRS15, JYZ11, Jin18, JL17b, Jov17, JJ12, JJB14, Jus10]. Model [KLM13, KID^{+}11, KPS19, KE10, KNPF19, KMS19a, KM13, KN13, KK16, KV16b, KPZ16, Kor16, KKA19, KOT11, KS14, KY12, KT17, KR15, KRK14, KS18, KiMM13, LHZ^{+}19, LL13, LN13, LS16a, LL16, LLM19, Lan10, LU19, Lee12, LSY19, LY13, Li19, Lia18, LS16c, LN11, Lim16, LMC19, LO17, LT10b, Lis17, LJN18, LN19a, LNP19, LYZ11, LV11, MOW11, MD10, Man11, MN16a, Mar18, MW12b, Mas13, MM16, MM17a, MSS15, MM12, MS12b, MG19, MiS18, MQW18, Miy12a, Miy12b, Miy16a, Mob13, MLCPS13, Mor18, MT11b, MR12, MS11b, MBC^{+}13, NT16, Ngu19b, Nig15, NE16, Ny13, OK14, OEA18, PL13, PT14a, Pan12, PK10, PF17, Pat11, Pat17, PC19, Pir14, Pir18b, Pos16, PFR13, dAPS11, PSS16, PW13, RRS12, RVB16, RESA10, RS15b, RG18, SML19]. Model [SG15, Sak12, SdlL18, SdlPRA16, Sch13a, Sch12b, Sch13b, Sch13c, SLM12, SY12, SS18c, SX10a, Sim10, SM12b, SB15b, SSBS14, Sun18, SV10, Tak15, TN18, TN21, TT16, Tan18, TN13, TM18, Tos17, TW14, Tro10, UK16, VAY^{+}12, WGLE11, Web11, WK18, Wei16, Wei18, WZIG14, Wu14, Wu15, Wu18, WLL11, XZ19, Yag16, YY10, Zha12b, ZMD^{+}19, Zoc18, vGRS16, AB14, SZ18, BS20, SB14b, Con13]. Modeling [IK17, DNBS10, DT19, FD11, FC17, FAB16, Grm10, JMSW13, JMH13, Naz18, OV15, RBM^{+}18, SZS15, SK19, WQ10, ZDS11]. Modelling [GLM^{+}16, KR10]. Models [AE16, ABMP16, AK18a, AK14a, AMT18, AG15a, AP11, ACL14, Ark10, AC14b, AJ19, BHNY15, BELP15, Bar12, BLT11, BDG^{+}14, BV16, Ber18, BFT10, BCCD18, BR16, BCL10a, BNY16, Bis19, BC10, BU18, BD17, BdlL13, BGZ17, BHMGM13, Bou13a, BN15, BIP11, BG17b, BSM^{+}16, Các14, CdlL10, CGGR13, CGRS16, CEL^{+}18, CBG14, Cas14, CC19, CP17b, CFMT17, Chu18, CV16, CDP17, CvELR18, CG13, CDV17, Cra13, CLT19, Cug17, DG17b, DP15, DARM^{+}13, DGL16, DF10, DvdHH10, DGvdH10, DG15b, DHK11, EHR12, FS14a, FC11, Fed14, FF11b, FKR12, FK18a, GLM^{+}15, GT17, GdK10, GRV11, HKLN19, HA13, Ham11, HM13b, IK10, IM16, II13, KLS19, Kel19, Ken10, KSH11, KN16, Kir12, Koi10, KL19, LR19, LS13, LPS16, LBW^{+}13, LNY16]. Models [LXHAA19, LY10, LSW17, LC19, LM12b, MS11a, MPL^{+}16, MSV10, MW12a, MR19, Mas16, MSS11b, MN16b, MM17b, MS16, MBS16, Mül11, Nán16, OVC14, OiS14, OR16, PPS16, Pan14, Pan16, PCMM18, PSAPR12, RVY18, RRW11, RP12, RT17a, Riv16, SS11a, SP13, STBT10, SWK^{+}18, SCSS18, Tau11, Thi16, TBD13, Tur13, Wan12, Wat19, Xue12, Xue15, Yam14, ZDG19, Zhu17, vEKRS19, vdHvLS18, MSV13, Wat20]. Moderate [CLL18, Fre15, LM12b]. Moderately [Coh09, Coh10]. Modern [Dyk14, Täu10, Rei09]. Modes [BDES19, LL10, MPM17, RG17]. Modification [SGU13]. Modified [BDES19, BGLL13, BHMGM13, CG14c, KPR18, MS12b, RST16, SM12b]. Modular [OV15, SK14]. Modularity [TVP13]. Modulated [PS19, SLST13]. Modulation [MMM15]. Molecular [BR11, CSV11, CDCL18, FS14a, Gol10, HV14, KL11, KBLL13, KU11, LNP11, LL10, LK11, ML15, MG17, ROS19, TAG10, TA12, LK11]. Molecule [Zho18]. Molecules [Ber18, Lu12, Pir18b, SdlPRA16]. Moment [AvB16, AF16, DFL17, KT17, OC12c, PB11, Sam19, SP18, Vie16]. Moments [BE16, Bur17, Che13, FL16b, IS11, KMB14, Laf19, RT12, Zha10c]. Momentum [CLT19, NS16, GDL10]. Monads [RT11a, RT11b, RT12]. Monday [Leb12a, Leb12c, Leb13b]. Money [Lan17, LR18, LR19]. Monodisperse [WCX^{+}11]. Monomer [ACH15, ACM15, AM18, AS16, FF11a, Fed13, GJL16, NKK19, Per17]. Monomer-Dimer [AS16, GJL16]. Monotone [IM16]. monotonic [RL17]. Monotonicity [Miy13, dLPS15]. Monte [BBD^{+}11, CF19a, Cli18, HG11, Koi10, LL13, ME11, MMSY11, UK16]. More-Dimensional [CL14]. Mori [WL13]. Moriya [QD12]. Morriss [BCHM12]. Morse [AFCA16, LQY17]. Moser [AKD19, AP14b]. Moshe [Bai10, Dyk14]. Most [MiS18]. Motility [BCMP15]. Motility-Induced [BCMP15]. Motion [BCJ17, Ban10, BBS11, BDT17, BBC15, BCY16, Bur19, CDG^{+}15, CWD17, DSPC14, DS16, FZ12, FW12, FT19b, GW12, GP10, GDMS19, Git12, GLT15, GOdSS11, JL16, Kat12b, Kaw16, KK15, Koi18, LPK13, LM13, Lia18, LWL^{+}18, Mol18, Nig15, ORS16, OW11, OV15, PP14, PCM15, PRD11, PRD12, RT17a, SP16, Vie16, Kau11, Kla11]. Motions [BFM10, Kat12d, Kat12c, Lie12, Ngu19a]. Motor [Ply15, ZF11]. Motors [FS14a, Gol10, KBLL13, LL10, MG17]. Motsch [Jin18]. Moussa [CDTA10, LS12]. Moving [GDMS19, KM19b, MBWC16]. mRNA [NVL11]. Multi [AMT18, BSS19, BL19, CES19, CCR19, CHHK19, FC11, FCK15, HA13, Jan15b, Kel19, Kie17b, KN13, Mat15, NS11, OEA18, PT14a, RSB10, Sug10, TT16, Tou14, Van17, WL13, ZL13]. Multi-band [TT16]. Multi-barrier [CCR19]. Multi-channel [HA13, RSB10]. Multi-component [Kel19]. Multi-dimensional [BL19]. Multi-group [OEA18]. Multi-layer [CES19]. Multi-level [Sug10, WL13]. Multi-particle [KN13]. Multi-scale [Tou14]. Multi-species [AMT18, BSS19, CHHK19, FC11, Jan15b, NS11, Van17]. Multi-Speed [FCK15]. Multi-State [PT14a, Mat15]. Multi-step [ZL13]. Multi-variate [Kie17b]. Multicanonical [Ark10]. Multicarrier [GI19]. Multicomponent [Wid17, Zoc18]. Multicritical [EKD12]. Multidimensional [BBC18, BHJ^{+}12, CSAS17, Det12]. Multifractal [BCW13, BT11, BC19, FDR12, IPS10a, She15]. Multilayer [FK10]. Multinomial [She15]. Multiparticle [Lee12, dVO15]. Multipartite [GT16, GT17]. Multiphase [CS16a, DM10, MFLA15]. Multiple [BBR^{+}19, CN13, EK10, HK18, JL18, JZ10, LV19, LJN18, MS12a, RELV11, TW13b, XD18]. Multiple-Relaxation-Time [BBR^{+}19]. Multiplex [KP18, Mül16]. Multiplicative [CRV13, Pir14]. Multipodal [KRRS17]. Multipoint [MZ19]. Multiscale [GW15b, KN13]. Multisite [JJB14]. Multispecies [AFR19, DDC18, HHM17, KMO16]. Multivalued [Mih10]. Multivariate [BSS14]. Mushroom [Bun14]. Mutational [Hui18]. Mutations [DDC18, Fri17, Hui17, LJN18]. Mutual [BHS13, CCM16]. Mutually [BCM12, Gup16, Kat12c]. Mysterious [Fed17].

Nagumo [HR10]. Naim [Leb11]. Nano [HNZ16, NE16, YKS16]. Nano-Channel [HNZ16]. Nano-scale [YKS16]. Nanochannels [TP15]. Nanoislands [Kor16]. Nanoparticles [ML15, OMC11]. Nanopore [BHF^{+}12, LZ15]. Nanopores [MMM15]. Nanosphere [DI13]. Narrow [FH13, LH17, LE17]. Nash [DLR14]. Natural [AKM13, SP16, Sme18]. Nature [Ito17]. Navier [ABH^{+}17, BLS17, BNT13, DLS10, FRT15, JLMG11]. Near [Afz12, Bai10, Bar16, BBS11, BBLP12, DLY18, EMO18, ELO11, HX15b, HY19b, Kua13, LWY18, Mih10, PS14, Sam15, Tak16, Yin16]. Near-Critical [BBS11, BBLP12]. Near-Degeneracy [DLY18]. Near-Extreme [PS14]. Near-Field [Bar16]. Nearest [BD17, BHJ^{+}12, Cha14a, FS17, Rad17]. Nearest-Neighbor [BD17, FS17]. Nearest-Neighbour [Rad17]. Nearly [Tan18]. Nearly-Flat [Tan18]. Needed [Cie17]. Negative [GdK10, MS19a, Sha18]. Neighbor [BD17, BHJ^{+}12, FS17]. Neighborhood [CGL12, OR16]. Neighbors [Cha14a, DF18]. Neighbour [Rad17]. Neighbourhoods [DO18]. Neighbours [HH15a]. Nelson [AF14b, DP14]. Nematic [BG17a, KWZ14]. Nernst [QLCL16]. Nernst-Planck [QLCL16]. Net [NML^{+}11]. Nets [GL13]. Network [APRT17, ARS17b, Asa13, BAC13, BGG10, BCCD18, BMT15, Che18, EV11, HRW14, JY18, KID^{+}11, KBLL13, LXHAA19, MNS12, MQW18, MB18, PT14a, Sha18, dAPdA^{+}13]. Networks [ADE18, ABLT17, AL18, Ami10, ALAF18, ADH12, ACGM18, BRSV18, Bar14, BVL16, BCS18, BH11, BJP17, BG14, BL10d, CT13, CVE14, CD16, CCGT10, CEB^{+}15, EMO18, Fid15, FC17, FW17, FCLK14, GvdHdHM18, GVJ^{+}18, HHT10, JSJ10, JPS17, Jav15, KKV^{+}11, KP18, Kle13, KYA16, LBW^{+}13, LH13a, LZHS19, MMW16, MSB13, MLS16, MN16b, MMR18, MLCPS13, Mül16, MHD17, MV16, MBC^{+}13, PSC18, RdAB18, RT16, RF18, ST11b, Sha12, SCSS18, Tou12, Tou14, TSS13, WQ10, ZHRB16, ZB13, AFI^{+}10, For11, Sta11]. Neural [ABLT17, Ami10, BGG10, CT13, Che18, GL13, MNS12, MPM17]. Neuron [MQW18, Vid15, Vid17]. Neuronal [RT16, Tou12]. Neurons [And19, DGLP15, DOR15, FGGL18, GRT17, Yag16]. Neutral [BPF^{+}14, TS19, Zho17, Zho18, Esl17, Sam17, Sam18]. Neutron [CHHK19]. Newman [Mac13, KV16b]. Newtonian [DM13, Kie14]. Nico [FvB14]. Nicolai [Mor18]. Nigel [Ano19b]. NK [HSFK18]. NLS [JP18b]. No [PS10b, TK16, CE14, DLS10]. No-Go [TK16]. Noble [Mol14]. Node [GVJ^{+}18]. Nodes [Kle13]. Noëlle [Des11]. Noise [AM19, Abr13, AEWCD19, BO14, BGJ^{+}15b, BML12, BGTV11, CTM13, DP15, FDR12, Fyo19, GS11b, GG11a, Gou15, HVW12, HMW19, ILOS10, IPS10a, KSSH15, Leh13, LL19, Mal12, Ngu18, NML^{+}11, RBM^{+}18, RBS16b, Ryt12, ST11b, Ste19, Wil10, XJZY13, XLL^{+}15, YBF^{+}17, ZW10, ZYZ^{+}18, ZWGM15]. Noise-Induced [HVW12]. Noises [BKLL12, LNS^{+}12b, MNBC19]. Noisy [Các14, FW12, GBTL17, MM13, MiS18, WLEC17]. Non [ABMP16, Afa19, AMT18, ASA15, AD15a, AC14a, ART15, Bao17, BHS13, BM11, BGL14, Ber12, BCF17, BTV14, BC10, BGN16b, Bra14, CRS14, CSL19, Car11, CIM14, CYZ18, CCGT10, Coh09, CRTZ13, CFG13, CE19, Cro12, DF18, Dav11, DG17a, DMS12, DMP17, EHR12, Esl17, Eva16, FL15a, FBR19, FS14b, FdHM14, FvdH13, FFT11, Fri17, GL14, GLML16, GDMS19, GS11a, GT16, GDL10, GS17b, GS17c, GS13, GBL16, Gor18, Gre12, Ily16, IM16, IPS10a, KSSH15, Ken19b, Kim12, Kle15, Koz17, KS18, LS13, LS16a, LNP13b, LBW^{+}13, LWY18, LPS19, LSBS13, MPTV12, MP18b, Moh17, MV16, NE16, OPS10, PL13, Pel14, RL17, RW12, Rue14, Rue17, SML19, SSR12, Sam17, Sam18, SLM15, SZS15, SGS19, Ste19, TP15, Var12, Vid15, VL12, VV19]. Non [Wre17, Zia10, Coh10]. Non- [ASA15]. Non-Analytical [GLML16]. Non-analyticity [SLM15]. Non-backtracking [FvdH13]. Non-coexistence [Car11]. Non-collinear [SSR12]. Non-compact [DMS12]. Non-condensable [Ber12]. Non-confining [AD15a]. Non-consensus [LBW^{+}13]. Non-Convex [GT16, AMT18, CFG13]. Non-coplanar [FBR19]. Non-crossing [GDMS19]. Non-cutoff [FL15a]. Non-Equilibrium [CCGT10, Rue14, BGL14, BGN16b, CRS14, CSL19, CRTZ13, Cro12, DF18, Dav11, DG17a, Eva16, Fri17, GL14, GS17b, GS17c, Gor18, Kim12, Kle15, LS13, MPTV12, MV16, NE16, PL13, Rue17, SGS19, Ste19, VL12, Wre17, Zia10]. Non-equivalence [VV19]. Non-ergodic [GS11a]. Non-ergodicity [VV19]. Non-Existence [Bra14]. Non-Hamiltonian [SZS15]. Non-Hermitian [Koz17]. Non-Ideal [CIM14, Moh17]. Non-integrable [MP18b]. Non-interacting [LSBS13]. Non-intersecting [Ken19b]. Non-Linear [OPS10, BCF17, GS13, IPS10a, LWY18]. Non-Markovian [Bao17, SGS19]. Non-momentum [GDL10]. Non-Monotone [IM16]. Non-monotonic [RL17]. Non-neutral [Esl17, Sam17, Sam18]. Non-normal [DMP17].

Non-relativistic [FS14b]. Non-reversible [LS16a, LNP13b]. Non-robust [KS18]. Non-smooth [FFT11]. Non-stationary [CYZ18, Ily16]. Non-Uniform [TP15, BC10, Var12]. Non-uniformly [Gre12]. Non-uniqueness [ABMP16, EHR12]. Non-wandering [AC14a]. Nonadditivity [Mor15]. Nonadiabatic [HS14b, Ply15]. Nonautonomous [You17]. Noncolliding [KT11a, Kat12b, Kat15]. Nonconcentrating [WK18]. Nonconventional [KV17]. Nonequilibrium [BBMW10, BL10b, BGJLL12, BCKL12, CA19, Des11, Dym15, EWSR16, GLM18, GQ17, Hag15, HEdPG14, IiS15b, JPS15, KNiST11, KNiST15, LY13, LE15, LY10, MN14a, Mae14, MOT14a, Mon12b, RC17b, RBR11, RY12, ST11a, SB15a, TN13, Tur13, VV17, WES11, YSSH13, ZBVE11]. Nonequivalence [Tou15]. Nonergodic [DM13, SM14]. Nonergodicity [Bao17]. Nongeneric [JS11]. Nonintegrable [MAPS11]. Nonintersecting [Lie12]. Nonlinear [BL15, BM12a, CFK13, FS11b, FS14c, Gao18, GQ17, GPP12, GHMR17, GH16a, GH16b, HX15a, KV15, KV16a, KNiST11, LPS16, Pol17, Qia10, RBR11, Spo14, SS15, XJZY13]. Nonlinearly [Fyo19]. Nonlocal [CLS11a, CDV17, KKV^{+}11, SST15]. Nonlocal-Interaction [SST15]. Nonmonotonic [AG12b]. Nonphysical [BV16, Ber18]. Nonpolar [Zho18]. Nonrelativistic [Hon19]. Nonreversible [BL12, DLP16, Eri18, ELX18]. Nonspecific [EAL12]. Nonstationary [OC12c]. Nonuniform [BFVZ10, SK19]. Nonuniversality [XTL14]. Nordheim [BE16]. Noria [BC14a]. Norm [MC11, Pro15]. Normal [BGJ^{+}15b, Hon19, Kie17b, DMP17]. Normalized [GV15]. Normed [BS15d]. Norms [Ras12, TV16]. Nosé [FG11]. Nosé-Hoover [FG11]. Not-So-Fundamental [Tho12]. Note [AHDV17, Bot18, CJN18, CP16, CP14, DOP19, DGL16, KKN12, LS17, Miy12b, MR12, Tyo12]. Notes [SB12]. Nov [Ano14b]. Novel [Lan11, NE16, Sch13c]. Nucleation [Gor18, vEKRS19]. Nucleosomes [CM11]. Nucleus [LH17]. Null [SCSS18]. Number [AM10, Bha15, CGS15, Che14, Cie17, DF18, FC13, GL17b, GV12a, MM16, MM17a, Per17, Shc14, TW11, XZ17, YS13, GL17a]. Numbers [DF17, FG12, GTT19, GM13, HŚ16, Mar11a, Mih19, NBK14, Nov15]. Numerical [BdlL13, CdlL10, FLTV11, HSZ19, LM13, LX17, SdlL18, Sit11, SV10, ZW12, dCCS19, dlLL11]. Numerics [AT12]. nutshell [Pel11, bA12].

Objects [DSPC14]. Observability [ABT11]. Observable [Wer12]. Observables [ASA15, LFW12, LFWK14, Luc18, OCM15]. Observations [BD15b, FFT11]. Observed [AL18, CGL12]. Observing [ORW15]. Obstacle [NC10]. Obstacles [ORS16]. Occupancy [HM13b, Hui18, TP15]. Occupation [Bur17, FGN14]. O’Connell [IS16, Jan15a, Kat12d, Kat12c]. Off [Ark10, Chh12, EK19, Fer18, Jin18, Mia11, Tri14, HJ17b]. Off-Critical [Chh12]. Off-diagonal [Mia11]. Off-lattice [Ark10]. Offs [MMW16]. Ohm [BBR12]. Oil [CC18]. Oldroyd [MvS19]. Oldroyd-B [MvS19]. Olsen [She15]. On/Off [EK19]. One [ASA15, ALAF18, BHS^{+}12, BELP18, BJ16, BRSW15, BC18, BCLL16, BW12b, BDY17, BBC15, CL14, CFTW15, CDS19, CMV11, CMP17, CD14a, CP10b, CL18, CL15, Cli18, CTT11, CLP19, CMV16, DP15, DR13, Fan16a, Fan16b, Fan17a, Fan17b, FS17, FSS13, FL18, FL20, FLS12, GG18, GMM18, GMM19, GH16a, GH16b, GTT14, GG11b, GT15b, HA19, HCO15, KP12, KS12, Koz17, KT17, KM19b, LSY18, Lia18, LLJH10, MSV10, MSV13, Miy12b, Mon15, MS19c, Nak14, PS18, PR15a, QS15, RNDS13, RDNS15, SML19, ST16a, Shi15, SS15, Tas18a, TGP12, TF12, TT12, Tzi13, Wan12, Xue16a, vEKRS19, Bru14, LSBS13]. One- [CL14]. one-atom [Bru14]. One-Component [CP10b, CMV16, Fan17a, GH16b, GT15b, TF12]. One-Cut [BDY17]. One-Dimension [ALAF18, LSY18, MSV13]. One-Dimensional [BELP18, BW12b, CDS19, CMV11, CD14a, CL18, CL15, DP15, Fan17b, FS17, FLS12, GG18, GMM18, GMM19, GTT14, GG11b, HCO15, KP12, KT17, Miy12b, MS19c, Nak14, PR15a, QS15, RNDS13, SML19, Tas18a, vEKRS19, ASA15, CMP17, LLJH10, MSV10]. One-Sided [BC18]. One-Way [TGP12]. Ones [ABLT17]. Onsager [Bax11, Bax12, KW12]. Open [ABFP15, APSS12, BBP17, BJP17, BW17, BS19, CP15a, DKY19, Fil16, JL17a, Kan14, KKSY19, KY13, KS14, LS15b, LS16b, MC19, Mon12b, NS16, Pel14, RM16a, SSE15, TZ16, Ter13, Van17, XY13, Xue16b, Yar14]. Open-Boundary [Kan14]. Operation [LL10, XJZY13]. Operations [LYX17]. Operator [Ast12, Ast13, BW12b, BBW15, Bou13b, CL15, Dud13, FKLL16, IPS^{+}10b, MU15, Miy13, Pol17, RMN15]. Operators [BN14, DG19a, DTW19, Fra17a, Ito16, KMM11, KP11, Krü12, Luc16, Nak14, Nak17, OC12b, Ong14, PdOC17, RST11, Shi15, TV16, dOP18, dlLOP11]. Opinion [CTM13, LBW^{+}13, RdAB18, SLST13, TLC13]. Optical [SH12]. Optimal [ADF18, BB11, BT14, BS13b, CDS19, DMP17, GdK10, HM13a, HM16, LNP13b, Nak19, NBK14, RCV16, SG15, SP16, WHC14, XZ17, BS20]. Optimization [CDCL18, DZ15, FD14, HA19, Tur13, Yed11]. Optimizations [LC19]. Option [MOW11]. Orbit [GBTL17]. Orbital [HR16, HR23]. Orbits [BC17, CNZ17, EK10, GO13, LQY17]. Order [Abr17, AW18, AK17, ARBJ15, CP16, CP14, CM11, CG19, CG12a, CFLV19, FZ11a, FBR19, GGP10, Gao18, GJS17, HK19, HT15, HDP17, HKW11, KL11, Kaw16, Ken16, Kie13, KL19, KOSV18, KT17, KS19, LK11, Liu15b, MC16, MC21, Miy16a, Pah10, Sch10, SM12a, SM12b, Tas19, WLJH18]. Order-Disorder [ARBJ15]. Ordered [ACL14, DF10, GGP10]. Ordering [BGL10, BGLL13, CLP19, Cra11, GC17, OR16, RT17b]. Ordering-Dependent [GC17]. organisation [dMS17a]. organised [MS12b]. Organisms [Hen12, Rab11]. Organization [Ero14]. Organized [Ara11b, CDG^{+}15, DK09, MT11b, OF18]. Orientable [DOGK16]. Orientation [Caë10, RT17b]. Orientations [YYZ11]. Oriented [AHDV17, MSV10, Rue18, Sak18, Tzi18, Zhe13, vEdLV16, MSV13]. Original [Hau16]. Originates [KS19]. Origins [BBR12, Hen12]. Ornstein [CG11a, CG15, GLST16, RE13, WX15, Wil10, WP11b]. Orr [dZS11]. Orthogonal [ACR18, BS16, DMY13, Gro19, Kat12a, Lie12, LZ10b]. Orthogonality [KJZ18]. Oscar [Ano14b]. Oscillating [CSC11, CCL19, KE10, SM12b]. Oscillation [DLLX16, Gao18]. Oscillations [BGTV11, CL13a, CF19b]. Oscillator [Art19a, Ban10, Fre16, GS11b, Git14, Kim12, LCZW15, NSS12, Sch13c, ZWGM15]. Oscillators [Bar16, BCM16, BO11, Dor16, EN18, GdHMZ19, HKR18, HKR23b, Ko19, KS12, MG19, SZS15, dAPdA^{+}13]. Oscillatory [CG13]. Other [BHS^{+}12, Rab11, BBGS19]. Outcomes [Gal13]. Output [BC15b, Vid17]. Overall [Zho17]. Overcome [LKR^{+}11]. Overdamped [Cam13]. Overlap [AD11, DM18b, GT17, Ito16, MU16, Mih19, SB12, vEdG11]. Overlapping [TVP13]. Overview [Pan12]. Own [CLTC15, CLTC23].

P [Weg17]. packed [BvE11]. Packing [Cie17]. Packings [HDS15]. PageRank [AKL18]. Painlevé [CCG14b, Liu15b, RK12]. Pair [GJMS10, MMST13, Mie18, PSS12, Ste10, Tid16]. Pairing [Mor11]. Pairs [GGP10]. Pairwise [MN16b]. Palis [DD10]. Panchenko [Con13]. Paper [Pro17]. Paperfolding [vEdG11]. Papers [GS13]. Parabola [QR13]. Paradigm [Sch11]. Paradigmatic [Bré14, MiS18]. Paradox [Pet10]. Parafermionic [Wer12]. Parallel [Ark10, AK16, DSS15, LVE19, ME11, MG17, NS12, SLM15, SJHW11, TT17b, Tor12, UK16]. Paramagnetic [BLW18]. Parameter [AG15b, BSM^{+}16, GLML16, HLZ17, ST16a]. Parameters [AB10, BP18, GLT15, Ply15]. Parametric [ABT10, ABT11, ART15, KBLL13, SWKS14]. Parametrization [AKM13]. Parasitoid [Ara11a]. Parcel [Ker13]. Parcel-Swapping [Ker13]. Parent [SW15]. Parisi [Ano15d, CG17, FSS13, TS12, TA16, XTL14, DM18b]. Parity [HMO12]. Parking [FK10]. Parsimonious [Ny13]. Part [Ker13, Wan12, dSRT15]. Partial [AKM13, Ari11, BPZ13, CF16, FF10, FJLS18, GHMR17, Ras12, GMT17a]. Partially [CGL12, DF10, SB15b, Vau10, dSLPV17, dWL10]. Particle [ByChL^{+}19, Bae11, BNTT16, BNP14, BHF^{+}12, Bur17, Bur19, CJW17, CMGP14, CF11, CG14b, Chu16, Cla13, CD14b, CDP17, Cor16, DF18, DF17, FF10, FF11c, FL18, FL20, FG12, FW12, FLS12, GRV10, GS11c, HKLN19, Hag15, HL18, KK15, KMS15, Kua13, KM19b, Lee11, Lia18, Luk14, Mar15, MSS15, MBWC16, Moh17, Ngu19b, Ohk10, Per10, Ren18, RE13, RS13, RNDS13, RDNS15, VW18, Vie16, Yar14, KN13, Lee10]. Particles [ARBJ15, BFKR10, BCMP15, BDDH14, BS15a, BCJP19, BG12, BFNZ11, BCM12, CL19, CS10b, CLT19, DHR18, DNBS10, DM10, Els12, FG12, GM13, HKLN19, LSY18, Lia18, LZ11, Lu13, Lu14, MGAPQH13, MP13, Pal11, Pet10, RW19, TFES19, Vau10, WG19, dMS17a, dHNT11, dHNT12, FC17]. Partisan [Lee18]. Partition [BCF10, CS15, Cra13, GJL16, GT15b, JL18, LSS19, MM12, RVB16, SS11a, SX10a, SVW12, Tem14, TS19, ZW12]. Partition-Function [SS11a]. Partitioning [Bec11, DSPC14]. Partitions [DG15a, FS18]. Party [Gal13, Mob13]. Passage [AB18a, AD15b, BBP17, BGL^{+}11, Các14, Cal15, CPSV10, CQR13, DO18, IP12, KV16b, Nak19, TM10, vKSZ18]. Passing [Yed11]. Passive [AM12, VFT12, WG19]. Pasta [AMS19, BP11a, BCP13, BPP18, DDN14].

Pastur [JS17]. Patchiness [BML12]. Path [BL11b, CSC14, CR17, EN18, JL17a, Kle15, KOT11, KO15, MS14, OBX11, Sha12, Wen12, vKSZ18]. Path-Integral [CSC14]. Pathogen [KKC12]. Paths [BIP11, CTB10, CB16, DP19, GDMS19, Ghe10, HDP17, KS14, Nak19, PST12]. Pathwise [HY19a]. Pattern [Per19, PKBW19, RM11]. Patterned [CCD15]. Patterns [ABLT17, BCY16, FDR12, HSUG13, LRL17, TD19, Tou14]. Paul [Ano18a]. Pausing [Klu11]. Pavel [Leb11]. PCA [BB17, BV15, LS13, PSS12]. PDE [CFL17]. PDEs [BCF17]. Pearson [Caë10, Caë11]. Pedestrian [CGY17]. Peliti [bA12]. Pendulum [TA12]. Penetrable [Cha14b, Cha15a]. Peng [BCJ15]. Penrose [AP14b, Tel10]. Percolation [AB18a, AD15b, APS12, AHDV17, Ami10, AF14a, Aza11, BCM10, BW12a, BH11, BS15b, BP12b, BM17, Cal15, CG11a, Car11, CS12, CC14, Cha12, Cha14b, Cha15a, CS16a, CG14c, CMM16, CQR13, CdLS13, DHS18, DD10, DO18, DRCV19, ELO11, FPR11, GMM18, GMM19, Goo12, GHS17, HTZ12, IP12, IvRM15, Ken19a, KV16b, MSV10, MSV13, Mis15, Mis16, Mis19, Mot14b, Nak19, PR15b, RZ17, RT17b, Sak18, SS17, SS18b, STBT10, Sch12c, Tag15, TU17, TVP13, TV15b, Tzi18, ZB13, Zhe13, dCCS19, dLPS15, vEF12, vEdLV16]. Percolations [RS14]. Percolative [AP18]. Perfect [DFR14, GGP10, GLO10, SL12, SP13]. Periodic [ALS18, BdSPMS14, BCKL12, Bra14, BC17, CG10a, CL13a, CL13b, CNZ17, CAS11, CD14b, DM18a, EK10, FS14c, Gri19, GO13, JPS15, LPK13, MBWC16, PZ17, Pro15, SdlL18, ST16b, WG19, dOP18, BdlL13, DG14, KMS19b, dlLSZ16]. Periodicity [GG11b, RT15]. PERM [HG11]. Permanent [Fed17]. Permanental [Cra13]. Permeable [AFR19]. Permutations [AFGL15, Bet14, GLU12, Ker10]. Perron [LYX17]. Persistence [AB18b, AK19, CP19, LPS19, Mol18]. Personal [Bak10, BKPW14, Fis11, Fis12, Per13]. Personalized [AKL18]. Perspective [Bak10, BA19]. Perspectives [Con10, CMS13, Gue13]. Persuasion [Mob13]. Perturbation [AK19, CMV11, Fan17a, Ito19a, Ito19b, LK14, O’C12a, SS18a, SW15, TZ16, Wil10]. Perturbations [Abr17, BJ16, BM12a, CF11, CEB^{+}15, CFG13, Dod15, FH11, FW12, Fre14, Fre16, Koz17, Luc12, RdAB18, dOP18]. Perturbative [BBS15a, BGN16b, DG14, MR19, dlLSZ16]. Perturbed [ADF18, BLT11, DNP17, EN18, FF14, GM15, IPS10a, Ito19a, Ito19b, PS10a]. Pesin [BSV18, Tia14]. Pfaffian [AVW17, BEP18, GJL16, Kar14a, Kat12a]. Phase [AGR19, AE16, APRT17, AM14, ARBJ15, AT12, AFR19, BG17a, BNTT16, BCMP15, BLS11, BL11a, BB13, BPDH10, BC10, BD11a, BYYY19, BM17, CT13, CRS14, CC18, CP16, CMP17, CL15, CRTZ13, CLP19, CFLV19, DKKP14, DOP19, DLY18, DH19, DK10, EL12, Els15, EKD12, FZ11a, FGGL18, FMM^{+}14, Fow19, GHRR13, GdHMZ19, GSSV11, GS19, GNP16, HR16, HKR16, HR23, Ham11, Hep18, HY19a, IvRM15, JYZ11, KP12, KMTC10, KRRS17, KY12, KS18, LHZ^{+}19, Lep15, LH13b, MGAPQH13, Mas14, MMR18, MCK15, MS16, MBS16, NKR15, OMC11, RMS19, RT11a, RL17, RM16a, RSB10, SGC11, SZ18, Sim11, SM12a, SM12b, SM15, SG17, Tak16, TN18, TN21, VB11, WLEC17, Wil11, XLL^{+}15, Xue16b, dHO13, ACL^{+}11, Dyk14, Mon12a]. Phase-Coexistence [EL12]. Phase-Locked [HR16, HKR16, HR23]. Phase-Ordering [CLP19]. Phases [BN14, JV19, KR15, MD10, MV16, RS19, RSY14]. Phenomena [ByChL^{+}19, Bak10, Bou13a, CCFR18, HY19a, LNS12a, NiS19c, PS10b, Yin13]. Phenomenological [BFT10]. Phenomenon [KiMM13, LX17, SdlL18]. Phenotypic [Kan12]. Philosophical [Bat17]. Phosphorylation [RGL11]. Photon [FS14b, dWL14]. Photons [CCL19]. Physical [BBH11, FD11, Mih10]. Physicists [San12]. Physics [Ano11d, Ano11e, Bai10, CCGT10, Des11, FW15, GT10, HA13, Han14, Kad14, LLM12, MSS^{+}11a, MP10, NSW13, PCMM18, PPK11, San18, KRBN10, Rei09, San13, Gou14, Leb11, Täu10, bA10, GvdHdHM18]. Picture [MSZ12]. Piecewise [DJRZ11, Fer14, GLBP12]. Pierluigi [Gue13]. Pierre [Bri19, Leb19, Mer19]. Pinning [BL11a, Ber14, DKLS19, SdlL18, dHO13]. Piston [BGN^{+}17, IiS15b, Kha19]. Pitaevskii [AU15, CPRY11, Pic10]. Pivot [Cli10]. Planar [AVW17, BW10, CS10c, EC11, Kol14, Lis17, NT16, RMN15, Sam15, VYH11, Wu18]. Planarian [CLTC15]. Planck [MGZ14, Wu15, ACCG19, AFCA16, AF12, AF14c, BBS^{+}15b, BL17, CLL19, DLBK11, DGL16, DWTW16, HR18, JLMG11, KP12, LWL^{+}12, LWY18, LK19, LQR12, MM16, MM17a, QLCL16, Shi16, Yan15, ZGL13, ZGL15]. Plane [BK17a, BGP10, BS15b, Bra14, CP14, FG14, IvRM15, MZ19, MvS19, PZ19, RM11, Sch10, YYZ11, dSRT15]. Plane-Like [Bra14]. Planelike [CDV17]. Plank [HX15a]. Plaquettes [CFMT17]. Plasma [AF16, CFTW15, CCM17, CMV16, FS11a, GH16a, GH16b, GT15b, KMKT11, RSY14, ST16a, Sam18, TF12, TT12]. Plasticity [Kan12, Lan19, MNS12]. Plate [TT17b]. Plates [OD11]. Plus [Mal12]. Point [BKM15, BFKP10, BDDH14, CG14a, CS16b, CTT10, CQR15, Esl17, FL14, FL16a, GL17a, GL17b, HKW11, JY18, KLM13, KID^{+}11, Kar14a, Kar18, KW12, Leb16, LSW17, MS19a, NBK14, OO18, ST14, Ter13, TVP13, dOP18, dSRT15]. Points [Bab12, BELP15, FZ11a, Li12, Mac10, Mol17, Xue12, Bai10]. Pointwise [LWW18, Wu15]. Poised [MB11]. Poiseuille [Mar11b, MvS19]. Poisson [BP19, CH14, DL17, HR18, ALM18, BGP15a, CCM17, CNZ17, CR17, CGR12, DF16, GLU12, HM13b, KK16, KD19, Lam19, OP12, QLCL16, Tri17, Vid15, Zha10b]. Poisson-BGK [Zha10b]. Poisson-like [CGR12]. Polar [HP11]. Polarization [TT12]. Polaron [LS14b, Los17, Miy13]. Poly [ABMP16]. Poly-time [ABMP16]. Polyanalytic [HH13]. Polyatomic [Ber18, KKA19, MM17a, Pir18b]. Polychromatic [Wu18]. Polycrystalline [SWK^{+}18, SWKS14]. Polydisperse [Cha12, Ono11]. Polydispersity [BT14]. Polygon [GL16b, Pah10]. Polygonal [Sch10]. Polygons [ACM11]. Polymer [AKQ14, AKH13, BFP10, Bur11, CSS15, HNZ16, IS16, Jan15a, KT11b, LZ15, MP13, NKK19, RW11, Sch12a, Tem14, VYH11, Wei16, Wei18]. Polymerization [PS16]. Polymerized [UK16]. Polymers [BL18, CDdS14, Cla19, CFN15, Hel16, HG11, HKW11, IST15, KMS19a, MT17, Mar11a, Thi16, TV12, WLL11]. Polynomial [ACR18, BFVZ10, CP19, CD16, GL16a, GC17, GJS17, JZ10, LX17, Mac10, PXX15, SB12].

Polynomial-Speed [LX17]. Polynomials [Afa16, Afa19, BS15e, CDTA10, DMY13, FZ11b, FN15b, FL16b, Kie17b, KV17, LXHAA19, Lie12, Rue18]. Ponds [Goo12]. Population [BELP18, BO16, CM18, KL15, LW18, ME11, MSS^{+}11a, Yag16]. Populations [AM18, ARBJ15, BL10d, CG16, EF18, GRT17, Gre16, HCI14, PSK10, RL11]. Pore [ZSHL15]. Pore-Scale [ZSHL15]. Porod [KMS19a]. Porous [BBR^{+}19, CLMK18, HS10, NG10, ZSHL15]. Portable [BR11]. Posedness [Han18, HJ17b]. Position [FL16b, HFWT15, Lee10, NS16]. Position-Based [HFWT15]. Position-Momentum [NS16]. Positions [LSY18]. Positive [Bur17, Fan17b, HHV16, KR10, MU15, Sha18, GdK10]. Positive/ [GdK10]. Positivity [ABFP15, Jaf15, Lis17, Pic10]. Possible [Hen12]. Posteriori [HA13]. Potent [HZS11]. Potential [ACG15, AF14c, Bec11, BW12b, BGN16b, CL13b, CD14b, CDH15, Cro12, DLBK11, DKLS19, Far15, FS11b, GPS13, HJ17b, KNK15, LMT15, LL16, LLM19, LS16b, LPK13, MMSY11, Ngu17, TV15a, TT12, Yuh15, KNPF19]. Potentials [AD15a, AG11, AFCA16, BN18, BBW15, CL14, Fan17b, FS14c, IY14, Kos13, MMST13, MFLA15, MPS14, Pir18a, Sak12, Tri14, Yag16, dVO15]. Pottier [Des11]. Potts [AE16, ALS18, ALAF18, BCF10, CS15, CV16, CDL^{+}12, EC11, GRR17, IPS^{+}10b, KRK14, LS16a, MM12, SS11a, SX10a]. Power [AF14a, DGvdH10, Far15, GvdHW17, GBTL17, Lau18, MS14, Mit16, Mit17a, Mit17b, PK11, PSCD13, SM14, vdHKvL18, vdHLK18]. Power-Law [AF14a, DGvdH10, MS14, PK11, PSCD13, vdHKvL18, vdHLK18, SM14]. Powers [FN15b, TF12]. Practical [Ko19]. Prähofer [EP14]. Prandtl [Bha15, MM16, MM17a]. Pre [Mis15, Mis16, Mis19]. Pre-Sierpinski [Mis15, Mis16, Mis19]. Preceding [ZL19]. Precise [ESPP^{+}14]. Predator [PK10]. Predictable [Kla11, Kau11]. Predictands [Luc18]. Predicting [LRL17]. Prediction [SP16]. Predictions [Ken15a]. Predictive [MC16, TN13, MC21]. Predictors [Luc18]. Preface [Ano10, Ano11d, Ano11e, Ano11a, Ano11b, Ano11c, Ano14c, Ano15c, Ano17a, Ano17b, Ano18d, Leb12b, Leb13a]. Preferential [DvdHH10, EMO18, GvdHW17, PPS16]. Preferred [BG14, Gal15]. Prescribed [Wil11]. Presence [BCMP15, Ber12, DR14, DS16, FM18]. Preservation [GW15b]. Preserving [CFS18]. Pressure [ABMP16, HiS18, KW15, LNT13]. Prey [PK10]. Pricing [MOW11]. Prigogine [MN15]. Primitive [DLLX16]. Princeton [BKPW14]. Principle [Ari11, CDCL18, CYZ18, FK18c, Gao18, GJS17, GLBP14, Hay15, Kra16, LS15a, Tur13]. Principles [ACR18, Han16, Kie17b]. Prisoner [CA18]. Probabilistic [BPR14, CT10, PSS16, RL17, SM15, Tag15, dMP12]. Probabilities [AB18b, CD14a, CC19, CS10b, FC13, KSH11, Lee11, LN15, NG10, OCM15, PS11, Tag15]. Probability [AC14a, CP19, CCG14b, Dai17, DNP17, IP12, Kat12c, Kol14, KS14, LU19, Mar18, PR15a, PST12, RS15b, SMCF13, Tou18]. Probability-Valued [PR15a]. Problem [AG11, BW12a, BP11a, BCP13, BPP18, Ber12, Bot18, Bou18, BE16, CGS15, CL14, CDS19, CSAS17, CAG^{+}13, CM12b, FF11a, GPMSBBSV15, GSSV11, GS17b, GS17c, GO11, HNVZ13, HRW14, Hui18, IiS15b, Kac13, Kuo17, Mar16, MP10, NBK14, NC10, PSS11, PSVG18, Ryt12, SBK10b, SLM15, Ter13, Tho11, WZ12, Yan11, Zha10b, ZHZ15, del12]. Problems [AM19, BBP17, Ber15, BP12b, CSC14, Fed13, HMU13, PPK11, Yam13]. Procacci [Pro17]. Procedure [Sit11]. Process [ARS17a, AY10, AFR19, BKS12, BMNS17, BGN^{+}17, BW17, BS19, CP12, CCG14a, CGL17, CGL18, CG19, CGHT16, CEGW18, CMS10, DR13, FF14, GG18, GRV11, HTX^{+}12, HY19a, Kan14, Kat12a, Kat12d, Kat12c, KL15, KA17, KCB13, KMO16, KLM17, LP16, LH13a, MdG13, MRR19, MPTV11, Ngu19b, OR15, PS19, PdS17, Pir14, Pro15, QR13, RE13, SS18a, SS10, Sch13a, SCY^{+}12, ST11c, Sim11, TGP12, TP19, TW13b, TW13a, Tzi13, Wil10, Xue16b]. Processes [ACG15, APR15, AGMM^{+}12, ADGPP17, AK19, ABT11, BKM15, BK18, BCHM12, BC15b, BO16, BJS17, BL10b, BZ13, Ber19, BDL16, Bla10, BJR10, BG17b, CFM14, CL16, CGP17, CFP10, CG13, Des11, DDC18, DSZ17, DXZ14, Eri18, ELX18, FC13, FC17, FL14, FGN14, FLS12, FL16b, GGJR14, GOP14, GKLT11, GN19, GLST16, GKW12, HL16, HL18, KT11a, KNK15, Kra16, Lam19, Leb16, LS17, LKD12, LPS19, LC19, Luc16, MOT14a, MC16, MC17a, MC17b, Mat15, MNBC19, MC19, MR13, MS12c, MHD17, MGMMP13, NS11, NVL11, OC12c, Ohk14, OPS10, OD11, OP12, OO18, Pim18, RW14, RM16b, San12, SL12, SV15, SJHW11, Sta15, TM10, Tou18, WX15, WP11b, XTpXpH12, Xue16a, XP17, ZGL13, ZGL15, vKSZ18, BBGS19, MC21]. Processing [FAB16, MLS16, MPR16]. Processive [ZF11]. Product [ABA14, BHNY15, BFKR10, Bis19, BS19, CG14b, MNV11, MC19, RS18, SW15]. Production [BC14b, HS14b, LH13b, MU13, MHD17, NT17, SS19, WX15, WXX16, YK13, YSSH13]. Products [AK17, BT11, CTT10, CLTT13, For13, ISZ17, Kar14b, ORSV15, RT14]. Profile [CFTW15, HM19, Ngu17]. Profiles [DDHS17, DSZ17, NTV16, Sta15, Thr18, YKS16, ZBVE11]. Program [Leb10, Leb12c, Leb13b, Leb13c, Leb14, Leb12a]. Projection [BC18, FKLL16]. Projections [BL11b, Kar12, Pir18a, RS14]. Projective [BPS12]. Projectors [TV16]. Proof [CE14, HM19, Lan17, PC19, Wat19, Wat20]. Proofreading [PS16]. Propagate [DEF12]. Propagation [BCE^{+}14, CP17a, CMW15, CGY17, Cor16, FK18b, KC18, Laf19, Liu15a, MBWC16, SR19]. propelled [ARBJ15]. Proper [GMT17a]. Properties [AS16, Art19a, Ast13, BM18, BF11, BCJ15, BO11, BBH11, BC18, Bla10, BGP15b, BBC15, CR11b, Cie17, CPSV10, CMS13, DHR18, DI13, DKLS19, Dud13, EF18, EPS17, FBE^{+}11, FK18a, FR17, GLT15, HKLN19, ISZ16, Ito16, JYZ11, JM14, Ker13, KMKT11, KSM16, LZ10b, Miy12a, MC11, Mou15, MHD17, MBC^{+}13, Nak19, Nán11, NT16, Pos16, Qia10, Ras11, RM16a, SSR12, Sch13b, SLM12, SLdEC11, SH12, Thä11, Wre17, Zha12a, dAPdA^{+}13, Bru14]. Property [FS19, LZHS19, PS18, RSB10, Zhe13]. Protein [BHS^{+}12, BCC^{+}16, GV12a, JYZ11, ZDS11]. Proteins [VCT11, WLL11, ZF11]. Protocols [PS16]. Prototype [PP14]. Prove [Fre15]. Pseudo [JSJ10, JL16, KNPF19, MFLA15, OD11]. Pseudo-Distance-Regular [JSJ10]. Pseudo-potential [KNPF19]. Pseudo-Potentials [MFLA15]. Pseudo-Processes [OD11]. Pseudofractal [PXX15]. Pseudospectral [Shi16]. Pt [CDCL18]. Publication [Ano16a]. Puiseux [GT12]. Pulse [HV14]. Punishment [PSVG18]. Pure [ANS18, BW10, BCJ15, GRS12, JJ12, JJB14, Ter13]. Purely [Ong14]. Push [Aza11]. PushASEP [CP15b]. PVBS [BNY16]. Python [San13].

QCD [Uka15]. Quad [Kel19]. Quadrangulations [BS15b]. Quadrant [Zhe13]. Quadratic [ADR18, BW17, CL14, Mou15, PSS11, Tak16]. Qualitative [dMS17a]. Quantified [dAPdA^{+}13]. Quantifying [HCLR11, PVCG11]. Quantitative [AK16, ACR18, GNP18, LWW18, Ven14, dMS17a]. Quantities [BC19, FSV10, Sim10, TT17a]. Quantization [ADR18, Jaf15, KMS19b]. Quantized [Vau10]. Quantum [AMS14, ABFP15, APSS12, ADP14, Aur18, BBR12, BFFS16, Ban10, BBP17, BC12, BN18, BR11, Bar16, BBBP11, BS15a, BGLZ19, BLU16, BM11, Ber15, BV11, BG19, BSS14, CP15a, CG14a, CM12a, CCR19, Cla13, CTT10, CTT11, CG12b, CH11, DeV19, DKY19, DFL17, FBR19, Far15, FS14b, Fra17b, GS19, GTZ12, GTT14, GMT17a, Grm10, GI19, HJ17a, Han14, Han16, HIK^{+}18, Hon19, HS14b, Imb16, IPS^{+}10b, Ito16, Ito17, JL17a, Jaf15, JPW14, JV19, JM10, JM14, JLMG11, Kar10, Kie11, Kie17a, Kim12, KKR18, KKSY19, KMS19b, KY13, KL19, Kur18, LTM16, LYX17, LS15b, LS16b, LWL^{+}18, LC19, Mar18, MW12b, MP18b, Mat12, MP10, MM17b, Miy16b, MT11a, MBS16, Mül16, NSS12, NMV11, NS16, NT17, OK14, Pel14, RS13, RSY14, ST11a, SJ10]. Quantum [SGS19, Tas16, Tas18b, Tas18a, Tas19, Wre12, XY13, Yan15, YSSH13, YS13, dVO15, LS10]. Quantum-Oscillator [Ban10]. Quasi [BdlL13, BLZ14, BGN16b, CG14a, CL16, CMM16, DO15b, Gor18, KMS19b, Li12, LSBS13, MS19b, Mor15, Mou15, RVY18, SdlL18, Wen12, dlLSZ16]. Quasi-Chemical [Gor18]. Quasi-Condensed [CG14a]. Quasi-Deterministic [Mou15]. Quasi-Equilibrium [Mor15]. Quasi-Geostrophic [BLZ14, Li12]. Quasi-isometries [CMM16]. Quasi-local [RVY18]. Quasi-one-dimensional [LSBS13]. Quasi-Periodic [SdlL18, BdlL13, KMS19b, dlLSZ16]. Quasi-Potential [BGN16b]. Quasi-Static [CL16, DO15b]. Quasi-Stationary [MS19b]. Quasicrystals [KR10]. Quasifree [BGLZ19]. Quasipotential [Ryt12]. Quasispecies [AP11, BAS16, KPZ16]. Quaternion [May13, YB14]. Quench [Buc16, NC10]. Quenched [ACL14, CS16b, DKKP14, GGvdHP15, GV12b, Luç17, OR16, dAPS11, SLdEC11, XTL14, Zho17]. Quenching [CDCL18]. Quest [Bal17, Bal18]. Questions [Ter13]. Queue [BFL18, FM18]. Queueing [AY10]. Queues [MSB18]. Queuing [BRSV18, CCGT10].

Rabin [Tho12]. Radial [CMV16, FL13, Ken12, MZ19]. Radiation [BDDH14, Fra17b, SJ10, SB15a]. Radiative [KC18]. Radii [DH19]. Radin [DF17]. Radius [Luc16, Pro17, dLP14]. Raman [HV14]. Ramp [MS19c]. Ramps [EK19]. Random [AB10, Afa16, Afa19, AI12, AKM13, AKQ14, AD15a, ACM11, AF14a, ABA14, AM12, AS16, AD11, AK14b, AO19, AZ11, Ast12, Ast13, APSS12, AGMM^{+}12, ADGPP17, AT12, ABF16, ACGM18, Bab12, BR10, BFKR10, BDP19, BvdHK19, BBBP11, BNP14, BO16, BGLZ19, BCL19, BL19, BL11a, Ber14, BO11, BZ13, BBLP12, BBLP13, Bet14, BCCD18, BCLL16, BPR13, BW12b, BBW15, BLRVR13, BS15c, BS15b, BU18, BR13, BJR10, BYYY19, Bou13b, BL10d, BKLO19, BD11b, BMSS13, Bur19, Bus18, BSM^{+}16, CRS14, Caë10, Caë11, CJN18, CG15, Can17, CP19, CDS19, CP15a, CLM15, CGL12, CD14a, Cha19, CR11a, CS16a, CCH16, CS16b, Chu18, Cie17, CDdS14, CCR19, CDH15, CR16, CEGW18, CPSV10, CMVW11, CFN15, CTT10, CLTT13, Con18, CG17, Cra11, CMM14, CCFR18]. Random [CFLV19, D’O14b, DO15a, DLY18, DP19, DGL16, DZ15, DH19, Det18, DKY19, DvLM16, DvLM18, DMM14, DRS10, DGvdH10, DGGvdH18, EH12, FZ11b, FZ10, FF10, FvdH13, FK10, FKR12, FS11a, For13, FL15b, FK18a, FH13, Fun14, FD14, FN15b, FMMPS18, Gal17b, GPS13, Gan18, GMM18, GMM19, Gao18, GdHR18, GOPS11, GSSV11, GL16a, GT17, GL17a, GL17b, GGvdHP15, GGD16, GS11b, Git14, GTT14, GMT17b, GMT17a, Gre12, GS15, GV12b, GM13, Ham11, HL16, HM19, Hor16, HŚ16, HM16, HMRW13, Hui17, IK10, ISZ16, ISZ17, IS16, Ito18, JL17a, JS17, Jan18, JSV19, JvLS19, Jav15, JM10, KLS19, KPR18, Kar12, Kar14a, Kat15, Ken16, Ken19b, Ker10, KA17, KW12, Kie17b, KV17, KMM11, Kla11, KKSY19, Ko19, KP11, Kol14, Kol17, KY13, Kos13]. Random [KOSV18, Koz17, KT19, Lam19, Lan11, LMT15, LS15b, LS16b, LMM16, LSY18, LM17, LCZW15, LM12b, Luç17, Mac10, MTVU18, MW12a, Mar18, Mar15, MN12, MU15, May13, MBWC16, Mis16, Moh11, Mol18, Mol14, MS16, Mou15, NMV11, Nán11, NT16, NKK19, OC12b, ORSV15, OCM15, OR19, OP12, ORS16, OR16, PPS16, PMC15, PSS11, PS10a, Pel14, Per13, PC19, PS14, PZ17, Pir14, PRD11, PRD12, RT14, RT11a, RT11b, RT12, RRS12, RZ17, RT16, RSB10, RT15, SS18a, SWB10, Sch15, Sco11, SLdEC11, SC10, Shc15, SS16, SS18c, Shi15, Sod09, Sod11, Sod17, SB12, SV16, TV15a, Tou12, TV15b, VFT12, Wan12, WZL^{+}14, Web11, WFK11a, WFK11b, XY13, Xue16b, XP17, YY10, Yin13, YB14, Yin16, You17, ZL13, ZHRB16, ZL19, Zhe13, ZWGM15, Zhu17, ZJ14]. Random [Zuc11a, dBP15, dBGP19, dCCS19, dSLPV17, vdHKvL18, vdHLK18, vBKBS13, vBSB15, Kau11]. Random-Acceleration [Bur19]. Random-Cluster [CG15, CEGW18, Gan18, GOPS11]. Random-Field [FMMPS18]. Random-Oriented [Zhe13]. Randomly [BLT11, Bur17, Cla13, KKR18]. Randomness [ACM15, BNHRV13, HiS19, MC17b]. Raney [FL15b]. Range [AHDV17, AM12, AD11, ANSW16, BDP19, BC11, BPRT14, BM12b, CMV11, CMP17, CG10b, CGP17, Chu18, CDdS14, CDH15, CvELR18, CDV17, DSZ17, GJMS10, GLO10, GDL10, Gon14, Gra13, KL19, KMO16, LSW17, MC19, MN14b, Mis15, Mis16, Mis19, Mit16, Mit17a, Mit17b, Miy16a, MPT19, Mor12, NP12, PL13, Sch12c, Sta15, Tas19, VW18, Wei16, vEdLV16, vEKRS19]. Rank [BBLP12, BBLP13, GKW12, Koz17, VP14, VP15, Wan12]. Rank-Driven [GKW12]. Ranking [BCS18]. Rapid [KO15]. Rapidly [CGHT16]. Rare [AGJP19, BLM13, CNS15, FK17, GKLT11, HCI14, Huv12, OMC11, RBS16b]. Rarefaction [BNTT16, CFP10, Gon14, MS17]. Rarefied [CLS11a, Kuo15, PSW17, TH12, TT17a]. Rarely [BGN^{+}17]. Rashmi [Ara11b]. Rate [BHS13, BGZ17, BFNZ11, CL19, CLS11b, GQ17, Lan11, Lee19, LT10b, MC16, MC21, PZ15, TWT14, WX15, WXX16]. Rate-Distortion [MC16, MC21]. Rates [AB10, BMR10, CMM14, DMM14, FLP10, Jus10, KCB13, LWY18, RBS16b, Tzi13, Xue16a, Xue16b, XP17, YB14, ZL13]. Rather [TM18]. Ratio [BT19, FL12, KNPF19]. Ratios [MFLA15]. Ray [RPPF15]. Rayleigh [Bha15, CDS17, HV14, KK15, Kuo17, NWL19]. Raymond [Ara11b]. Re [MMR18]. Re-entrant [MMR18]. Reactant [SM12b]. Reacting [Ren18]. Reaction [AG12a, ADE18, AFS^{+}13, FK18b, GQ17, Kor16, KCB13, LLS13, RESA10, TP19, WGLE11]. Reaction-Telegraph [TP19]. Reactions [Bai10]. Reactive [ZL19]. Real [AB17, BLS17, Mac10, May13, PSC18, Wan12, dMPTW16]. Realizable [Lan13]. Realization [Kan14]. Rearrangement [CVE14]. Rebellious [SV10]. Receptor [AWE16]. Receptors [BC16]. Reciprocal [Kat12b]. Recombination [APR15]. Reconstructing [Fyo19, ZBVE11]. Reconstruction [LJN18, LN19a, LN19b, RMS19, SP18]. Record [WFK11a, WFK11b]. Records [PKDK13, SV16]. Recovery [CG14c, DJRZ11, Xue16b, XP17]. Rectangle [NKR15]. Rectangles [Cha15a, DG19b]. rectangular [GO11]. Recurrence [GNP18, LS15b, Nán11, SLdEC11, Tro10]. Recurrent [DP19]. Recursion [Kum19, PZ15]. Recursive [FF11a, HS18]. Redner [Leb11]. Reduced [BC17, CFL17]. Reduction [AGGL^{+}16, DLP16, FKLL16, HSZ19, Hel16, KP12]. Refined [AGMM^{+}12, EL12]. Reflected [CGN16]. Reflecting [DO15a, Kua13]. Reflection [Jaf15, OiS14, YH14]. Reflections [AN19, BKPW14, Kad14]. Refreshing [CG11b]. Regime [Ber14, BCPS18, CG11a, Con11, GTT19, HKW11, KPZ16, LWL^{+}12, MT17, SS10, Shc11, Shc14, Tri17]. Regimes [BBLP12, BBLP13, MQW18]. Region [ABMP16, BPRT14, EL12, TKK15, ZW12, dCFC11]. Regression [dAPdA^{+}13]. Regular [AP18, Bet14, BU18, Can17, Con18, Goo12, GS15, JSJ10, MW12b, Mis19, Mol17, Per17, PRS12]. Regularity [BK17b, Gal17b, GMM18, Ham11, Lia19, Lu14]. Regularization [LLS17]. Regulation [KYA16]. Regulatory [IK17]. Reichl [Täu10]. Reinforced [PP14]. Reiss [Ano16b]. Related [BW12a, BDL16, BGZ17, BLS17, CDTA10, DFF18, FP11, GOP14, GRV11, HIK^{+}18, MS16]. Relating [MHD17]. Relation [BL11b, FGJ14, GMM19, GH16a, GH16b, GJ15b, Kac13, Kat12b, Kol17, KNiST11, LPK13, MT16, NS16, NP14, Rue16, ST11a, SCY^{+}12, SS19, UV16, YBF^{+}17, dlLOP11]. Relations [BC15a, CM12a, GT15b, HTZ12, KS13, KNiST15, MT11a, Ras12]. Relationship [Jia14]. Relationships [AK16, PKDK13]. Relative [BC17, GPD16, Kie17b, ME14]. Relativistic [BDG^{+}14, BCNS12, DRS18, FVV14, Gal15, GPMSBBSV15, GPMEA17, GPMEA18, HY19b, KT17, LY16, CE19, FS14b, FVV15]. Relaxation [BT14, BGZ17, BBR^{+}19, CN13, GdK10, HKR18, HKR23b, HCO15, Pir18b, PSCD13, TAG10, ZBVE11, BL16b]. Release [MGMMP13]. Remark [Cha12, Hog11, Pro17, dLPS15]. Remarkable [EP14]. Remarks [CL14, Per19]. remembrance [Bri19]. Removal [Tak16]. Renewal [ACG15, GGJR14, GG18, MC17a]. Renewals [TA16]. Renormalisation [AB17, BBS15a, BS15d, BS15e, BS15f, BS15g, WPB15]. Renormalization [AEWI14, ACL14, AM12, Bar12, Bat17, BM12a, Bré14, CdlL10, CCEF10, CG14a, CQR15, DLLX16, ES13, EJ10, EV14, Gal14, Ken10, KM17, LHZ^{+}19, LS14a, MPM17, OC12c, OC12b, ZJ14, dlLOP11, dlLL11]. Renormalized [Dal11, Gin14a, Gin14b, Ser14]. Renyi [LV11, CLL19, DvLM16, KLS19]. Repeat [TGP12]. Repeated [ADP14, TZ16]. repellent [TV12]. Repellors [Mih10]. Replica [BGG10, BSS19, CPV10, Con18, Ito17, Ito18, Lan15, LO18, RLCMRT10]. Replication [GG11b]. Reply [Tak10a]. Representation [ABMP16, AP11, ABA14, CS10b, GL14, Ker13, Neu14, Ohk10, Sch10, SWKS14]. Representations [CV16, Gro19, OO18]. Reproductive [CLTC15]. Repulsion [BKM15, IST15, QLCL16]. Repulsive [CDdS14, EKD12]. Requisite [BMC17]. Rescaled [Fed14]. Reservoirs [CEL^{+}18, Dor16, MPTV11, Ngu19b]. Reshaping [Kor16]. Residence [ZDG19]. Residual [LN11]. Resistance [Asa13, Cha14a, TW11]. Resistances [JSJ10, JY18]. Resistor [JSJ10]. Resolution [Bon15, Pet10, PPK11]. Resolve [SKT16]. Resolvent [BYYY19, Mit16, Mit17a, Mit17b]. Resonance [DT18, GS11b, HV14, KNK15, LNS^{+}12b, LCZW15, ZYZ^{+}18, ZWGM13, ZWGM15]. Resonances [CCL19, TLD18]. Resonant [dlLSZ16]. Resource [HZ13]. Response [Abr17, AFG12, BBMW10, BC19, Bur11, CEB^{+}15, HLZ17, HCLR11, Luc12, Luc16, LRL17, Luc18, MT17, PE19, RC17a, WG19, WG18, WL13, YBF^{+}17, ZW10]. Responses [ADF18]. Restoration [IK10]. Restricted [BIP11, DFF18, LZ10b]. Restriction [ADC10, ADH12, Zha12a]. Result [And19, MT17, NTV16, SL12]. Results [ACM15, AKM13, AB17, Bur17, CdlL10, Cas14, CS12, CEB^{+}15, CDTA10, CFN15, CMS13, CMM14, DMM14, EP14, Fan16a, Fan16b, LR18, LR19, MS19b, MR19, MG19, PR19, RKGZ12, Sam15, SX10a, ZW12, dCCS19]. Resurrection [DM18a]. Retaux [HS18]. Retina [SP16]. Retrieving [ABLT17]. Retrospective [CL14]. Return [KMSS19]. Reunion [CD14a, SMCF13]. Reuven [For11]. Reversed [FL13]. Reversibility [BCJ17, CSAS17, CAS11, HiS19, Mon12b, Zha10a]. Reversible [ABF16, JSV19, Kor16, LMT15, LO17, RBS16a, SY12, LS16a, LNP13b]. Review [Con10, FMMPS18, HG11, Kie17a, Sta13, SCSS18, Per13]. Revising [Luc18]. Revision [AG11]. Revisited [HvdHH14, Hui18, LMC19, NTV16, Par17]. Revisiting [HX15b, MN15]. Reweighted [BL11b]. Reynolds [Che14, DUU15]. RG [BB17]. Ribbon [DO14a]. Ribbons [DO14a]. Ribosome [GV12a]. Ribosomes [NVL11]. Richard [Kla11]. Riemann [Bab12, Det12]. Riemannian [HKR18, HKR23b]. Riesz [Leb16]. Right [Hen12]. Rigid [Koi18, Tor12]. Rigidity [AR11b, Aum15, Bar12, GL17a, GL17b, LRY18]. Rigorous [ACM15, BGN16a, Bal14, Cas14, Gri19, JPV18, Lan17, LR18, LR19, MS12b, Ono11, CMS13]. Rigourous [DP17]. Ring [BL16b, CGP17, DP17, Kan14, SLM15]. Rise [HJR^{+}11, LE17]. Risk [BA14, CA18, MBC^{+}13, PS19]. Risky [DPT17]. RNA [SW11]. Robert [Kie11]. Robinson [BCJ15, GL16a, IST12]. robust [KS18]. Robustness [Kan12]. Rods [Cha12, OD11]. Role [BPR13, CTM13, DUU15, HH15b, MG17, NBB13, Ste19, TLC13]. Roof [BM18]. Room [Ano12a, Ano12b, Ano18a, Ano19a, Ano19b, Leb12a, Leb12c, Leb13b]. Rooted [JM14, LPS12]. Rotating [Buc16, KMS19a, MS11a, RSY14]. Rotation [VYH11]. Rotational [Aum15, BCJ17, CPRY11, Esl17]. Rotations [dlLL11]. Rotators [BGP10, Dym15]. Rough [Fre14, ME11, TJ15]. Roughness [SSB15, Zho18]. Rouse [VV17]. Rowlinson [DH19, FKKO15, Fri17]. Rowlison [Zoc18]. RSB [Pan14, Pan16]. Ruelle [CLP17, CL15, Gal17a, WL13]. Rugged [BBK17, FK12, MMSY11]. Ruin [LS16b]. Rule [AF16, HMO12, LN13, NS12]. Rules [SJ10, TF12]. Rumor [BZ13, BHMGM13, CP17a, CRS12, GGJR14]. Rumour [APRT17, SR19]. Rupture [DOGKP19]. Russo [dBP15]. Rutgers [Ano12a, Ano12b, Ano18a, Ano19a, Ano19b, Leb12a, Leb12c, Leb13b, Leb13c, Leb14]. RWRE [Fre15].

S [Sha10]. Sacred [MSB13]. Sagawa [MT11a]. Saitta [Mon12a]. Sakaguchi [HX15a]. Sale [BJM15]. Salesman [SBK10b]. Salpeter [Mor11]. Sample [DNP17, WXX16]. Sampled [BO16]. Samplers [DLP16, RBS16a]. Sampling [APdM^{+}18, FT18a, FT18b, FJLS18, Ghe10, HMRW13, HM13b, Hui17, KR16, LM13, PVCG11, PSS12, TV16, WES11, WLL11, ABT10]. Sander [Gou14]. Sandpile [RRS12, Tyo12]. Sandpiles [FLP10]. Satisfaction [HMU13]. Satisfied [Lan13]. Satisfying [LLL17]. Saturated [Cie17]. Saturation [HMU13]. Sausage [NP16]. Saving [HBC^{+}15]. Scalar [AM12, PR19, TM18, VFT12]. Scale [AK18a, ARS17b, BL16b, CL16, CD16, GRT17, HH15a, HFWT15, JvLS19, KTJ10, KS12, LH13a, MR13, MPM17, NE16, PXX15, RM16b, WBL11, WCX^{+}11, ZSHL15, Zho17, Zho18, vdHK17, Tou14, YKS16]. Scale-Free [AK18a, ARS17b, CD16, HH15a, HFWT15, JvLS19, LH13a, PXX15, WCX^{+}11, vdHK17]. Scale-Induced [KTJ10]. Scaled [Bab12, Bae11, ZP15, ZP16]. Scales [AEWI14, BP11a, BC14b, CFMT17, LL13]. Scaling [AM19, BBS14, BJS17, BKK15, BBC15, BLT12, Bot18, CDS19, CCP16, CGNP11, Chu16, Cie17, Cla19, EP14, FMAG11, FC11, FN15a, FKK10, FL16a, GLML16, GLT15, HDP17, HTZ12, HF12b, HKW11, Ker13, KL15, KWZ14, LZ15, MU18, Mar11a, PZ17, RP12, SGU13, TW11, dlLOP11]. Scalings [EM10]. Scatterers [BNP14, CTT10, LZ10a]. Scattering [CE19, DP14, HV14, HJ17b]. Scenario [DP17]. Sceptics [SR19]. Schelling [BELP15, BELP16, BELP18]. Scheme [LM13, Sug10]. Schemes [Tem14]. Schloegl [WGLE11]. Schmidtea [CLTC15]. Scholes [LWL^{+}12]. Schonmann [BC18]. Schramm [ACH15, Ghe10, JL18, Ken15b, LV19]. Schrödinger [Ast12, Ast13, BW12b, BBW15, Bou13b, FS11b, FS14c, Gao18, Hon19, KV15, KV16a, KK16, Krü12, MC10, Nak14, Nak17, Shi15, TV16, dOP18]. Schrödinger-Type [MC10]. Schultz [Tas18a]. Science [HBC^{+}15, Kla11, Kau11]. Scientific [GS13]. Scope [BCCD18]. Score [HM16]. Scores [DMP17, GPD16]. Scoring [HM16]. Screened [CFLV19]. Screening [AG14, CA18, FK10]. SDE [AG15b]. Sea [ABLT17, TW17]. Search [OTNN11]. Searchability [Kle13]. Second [AGMM^{+}12, BFKR10, BS15a, CP14, CS10b, FS17, GJS17, HK19, HMO12, HT15, Kim12, LT10b, Mae14, OC12c, WGLE11, Vie16]. Second-Class [BS15a]. Second-Grade [LT10b]. Second-Order [HK19, HT15]. Secondary [MMR18]. Sections [Bab12, DMS12]. Secure [KD19]. Seeds [GG11b]. Seen [BD11b]. Segmentation [XZ17]. Segregation [BELP16, BELP18, OF18]. Seiringer [Kie11]. Selecta [Sin10b, Szá12]. Selected [LFW12]. Selection [BW18, CR17, DS16, FN13, Kos11, Lep12, LKD12, SWK^{+}18, Sme18]. Selective [TP15]. Self [Ara11b, ARBJ15, BSW17, Bha15, BCF10, BBC15, CDG^{+}15, CS16b, Cli10, Cli18, CDP17, CF19b, CMVW11, DNBS10, Det18, DGK^{+}11, EJ10, Els15, Ghe10, Gil15, GV15, GOdSS11, Ito19a, Ken12, Ken15a, Ken15b, KV16a, KT12, Lau18, LH11, LSY18, LXHAA19, LNP19, LPS19, MNV11, MS12b, MT11b, Mül11, NV14, NTV16, OF18, Pal11, PG10, RW11, SH16, SM12b, TW11, Thr18, TV12, Wan12, YY10, YB14, dMS17a, DK09]. Self-Assembled [Ara11b, DK09]. Self-assembly [Mül11]. Self-attraction [BSW17]. Self-Averaging [Ito19a]. Self-Avoiding [BBC15, CS16b, Gil15, Ken15a, Ken15b, SH16, BSW17, Cli10, Cli18, DGK^{+}11, Ghe10, GOdSS11, Ken12, RW11]. Self-Consistent [Bha15, KT12]. Self-driven [DNBS10, LH11]. Self-Dual [YB14, BCF10, Wan12]. Self-interacting [Pal11]. Self-interaction [CMVW11]. Self-Normalized [GV15]. Self-organisation [dMS17a]. Self-organised [MS12b]. Self-Organized [Ara11b, CDG^{+}15, MT11b, OF18, DK09]. Self-oscillating [SM12b]. Self-propelled [ARBJ15]. Self-repellent [TV12]. Self-Similar [KV16a, LPS19, NTV16, TW11, Thr18, Det18, Lau18, LXHAA19, LNP19, MNV11, NV14]. Self-Similarity [EJ10, Els15, YY10]. Self-Sorting [LSY18]. Self-sustained [CF19b]. Self-tuning [PG10]. Selfaveraging [PS18]. Semi [CS10c, KMS19a, LZ15, RT17a]. Semi-dispersing [CS10c]. Semi-flexible [KMS19a, LZ15]. Semi-Markov [RT17a]. Semicircle [SW12]. Semiclassical [CSC14, FS11b, Kur18, Laf19]. Semiconductors [ABS12, BF10, MS11b]. Semiflexible [Bur11]. Semiflows [AC14a]. Semigroup [CRV17]. Semigroups [JPW14]. Semimetallic [Mas14]. Send [Yoo10]. Sense [Kie17a]. Sensible [Yuh15]. Sensing [AWE16, HCLR11, tWBOM16]. Sensitive [CRL15]. Sensitivity [Che18]. Separate [Dor16]. Separation [BCMP15, BPDH10, CMP17, Ham11]. Septuagenarians [AK18b]. Sequence [AK16, BHS^{+}12, HM13a]. Sequences [ABR18, AB18b, BCW13, Lan11, Osi16, TV16, WFK11a, WFK11b, ZL13, vEdG11]. Sequencing [RBM^{+}18]. Sequential [Cie17, DvLM16, DvLM18, FO18]. Series [Ara11a, EC11, Jan12, Kie13, KM19a, Leh13, MCG12, MM13, dLP14, Pro17]. Servers [BRSV18]. Set [ALS18, ART11, EHR12, SX10a, SX10b, ZHZ15]. Sets [AI12, AC14a, AFFR17, Bun14, GS15, Mih11, PR15b, HHLP17]. Setting [DMS12]. Several [GS11c]. Shaken [HHT15]. Shallow [RVB16]. Shannon [BMSS13]. Shape [EAL12, KNK15, Sch13b]. Shaped [CPS19, DG15c, MGAPQH13]. Shapes [FS18, PVCG11, PW13]. Sharipov [Tak10a]. Shark [Bus18]. Sharp [CG14c, NS12, YM11]. Shear [AM12, BJGL^{+}17, ES13, Mat12]. Shell [BFT10, Sch13c]. Sherrington [Con13, BL16a, BSS19, Pan12]. Shield [CCM17]. Shift [Bré14, Ker10, Krü12]. Shifts [BCHW19, Kem11]. Shlomo [For11]. Shock [BFKR10, FN15a, Fer18, GW15b, KKA19, MS10, PdS17, YM11]. Shock-Wave [KKA19]. Shocks [Abr13, BNTT16, BDP19, BS13a, BRSW15, CFP10, MS17, TGGS13]. Short [AD11, ANSW16, BB15, Bot18, BPRT14, CLM15, HV14, HM19, MPT19]. Short-Range [AD11, ANSW16, BPRT14, MPT19]. Shrinking [HNVZ13]. Si [OMC11]. Sided [AB18b, BC18, TW10]. Sidney [Leb11]. Sierpinski [JY18, Mis15, Mis16, Mis19, LZHS19, Yam13]. Sigma [SS18c]. Sigma-Model [SS18c]. Sign [BEP18, TA16]. Signaling [RGL11, ST11b]. Signals [ByChL^{+}19, Fyo19]. Signature [Dzu11]. Signed [ALS14, KLM13]. Silicon [AG12b, MS11b]. Similar [KV16a, LPS19, NTV16, TW11, Thr18, Det18, Lau18, LXHAA19, LNP19, MNV11, NV14]. Similarity [EJ10, Els15, GPSB18, YY10]. Simmons [Zha14]. Simon [Ber19]. Simple [AM12, BNTT16, BL10b, Bou13a, BW17, BS19, Cen13, CDS10, DR13, GLML16, GLM^{+}16, HT11, II13, JYZ11, Kan14, LY13, MD10, MOT14a, MPTV11, MPTV12, Mat15, Ngu19b, RS15a, RY12, SS10, Sim11, SM14, TN18, TN21, TW13b, TW13a, WLL11]. Simplest [BG17b, FD14, JJB14]. Simplicity [KBLL13]. Simplified [HMRW13, IPP14, del12]. Simulated [BR16, CCD15, LTM16]. Simulating [CF19a, GKLT11]. Simulation [AD15b, AGJP19, DDF15, GGP10, GLO10, HTX^{+}12, HEdPG14, II13, LX17, ML15, MCK15, PM17, SL12, SP13, SSB15, Wil11, YM11]. Simulations [BR11, FCK15, HG11, Koi10, LKR^{+}11, MFLA15, PPK11, UK16, San13]. Sinai [Szá12, CZZ13, CLP17, Fre15, GTT14, Kha17, MDP^{+}18, Ste10]. Single [AWE16, BHF^{+}12, BS15g, Bur19, Chu16, KMS15, Pah10, SSR12, TT16]. Single-File [Bur19, KMS15]. Single-Particle [Chu16]. Single-Receptor [AWE16]. Singular [BLS17, BS19, BCM12, CCM16, GBL16, JS17, Ong14, Tou14]. Singularities [BKP13, CYZ18, RS15a, TW14]. Singularity [Che13]. Sink [BB11]. Sintering [ML15]. SIR [Gra13]. Site [BS15b, Car11, HP11, MSS11b, RZ17, Wre12]. Sivashinsky [BC17, MiS18]. Six [CS16c, CPS19]. Six-Vertex [CS16c, CPS19]. Size [BP11a, BLT12, CG10b, DRT18, DR13, DM18b, IP12, JK12, LNP19, LLS13, RELV11, Sam17, Sam18, Shi13, SM12a, SK19, TQS12]. Sizes [PVCG11]. Skew [ABT^{+}14, Han13, Krü12]. Skew-Shift [Krü12]. SLE [BCL10a, BCL10b, FL13, HBB10, HK18, Kem10, LV19, MZ19, Wer12, Zha12a]. Sleep [Hep18]. Slicing [MSS^{+}11a]. Slightly [CLS11a, TH12]. Slip [ABH^{+}17, DLS10, HT15, LL10, NiS19b]. Slip-Flow [HT15]. Slippery [Kad13]. Slow [BMNS17, BC14b, BGTVE16, FT19a, IiS17, LL13, Lia18, Pir18b, RT11a, RT12, dSLPV17]. Slowdown [FZ12]. Slowly [Ber19, GPS13]. Slyozov [CN14]. Smale [NBK14]. Small [Abr17, APRT17, AEWI14, ALAF18, BJ16, BP18, BBLP12, BBLP13, CF11, CZZ13, EMO18, GM13, HHV16, HMW19, Kol17, KV16b, LL16, LLM19, LP16, LWL^{+}18, Ngu18, OTNN11, Sha12, vdHK17]. Small-Mass [HHV16, Ngu18]. Small-Rank [BBLP12, BBLP13]. Small-World [APRT17, ALAF18, Sha12]. Smallest [Kum19]. Smart [Dai17, Ken15b]. Smoluchowski [GVJ^{+}18, HVW12, Lau18, NV14, NTV16, Thr18]. Smooth [TH12, TT17a, VFT12, FFT11]. Smoothing [BL11a]. Smoothly [BDL10]. Smoothness [JL18]. Sneppen [BBGS19, BAS18, MS12b, Sch12b]. SOC [GV15]. Social [ABC10, AFS^{+}13, BLL^{+}13, BPR13, CDS10, Hub13, LNP13a, OEA18, TSS13]. Socio [Bou13a]. Socio-Economic [Bou13a]. Sociophysics [Sta13]. Soft [AG11, GPSB18, GGD16, HJ17b, Mül11, Pos16]. Sojourn [FM18]. Solid [DG15a, KA17, RM16a, TH12]. Solids [AR11b, MW10, SBK10a]. Soliton [Bon15]. Solute [WBL11]. Solutes [SvHM^{+}11]. Solution [ACM15, AF12, CTM13, CES19, DLBK11, JJB14, Jus10, KSSH15, KL15, KT12, LM13, LWW18, LO17, Mas13, MW10, MC19, MS19c, RPPF15, Sco11, Shi16, Van17]. Solutions [AM19, AG12a, AG11, AFCA16, Ara11a, BL18, BLS17, CH14, FL15a, FRT15, HR18, HR10, Ily12, Ily16, KS13, KV16a, KK14, KOSV18, Lau18, Li12, LWY18, Liu15b, Lu12, Mar16, Mar11b, MN12, MNV11, MWY16, NV14, PR15a, Tho11, VAY^{+}12, Yan11, Zha10b]. Solvable [AK14a, BM11, BD15a, Caë11, Lee11, TM18, VB11]. Solvation [FD11, RW19]. Solving [CJW17, ESPP^{+}14, Pes14]. Some [AKD19, AK17, AKM13, BBH11, BGP15b, CL14, CS12, CS10b, CP17b, CCH16, Con10, CMS13, DM10, GT12, HKR17, HKR18, HKR23a, HKR23b, Hol11, JY18, Ken16, Kie13, LZ10b, PPS16, Per19, Ras11, RS13, RY12, She15, Sod09, Sod17, Zha10a, ZW12]. Somitogenesis [ZMD^{+}19]. Sommerfeld [dZS11]. Sorting [LSY18]. SOS [ADMS11, KR15]. SOS-Model [KR15]. Souma [Sta11]. Source [BB11, BBLP12, BBLP13, HRW14, Wan12]. Sources [TC11a, TC11b]. Space [AGR19, AB17, AN15, Bal17, BK17b, Ber12, Ber15, BC17, Con11, CLP19, Cro12, FL15a, GPSB18, JL16, JPS15, KP12, KMTC10, LLS17, LNT13, Luc16, LH13b, LQR12, MGZ14, MU18, ORS16, Sak18, WK18, Yan11, ZLL13, ZGL15, vEdG11, Bal18]. Space-Dependent [AN15, Cro12, ZLL13]. Spaces [BK17b, BNY16, CL14, HY17, OO18]. Spacing [IM16]. Spanning [DGK^{+}11, HNT18, TW11]. Sparse [Afa16, GHLV16, Kos13, LV11, MW12a, Mon15, RMS19, Sod09, Sod17, Zha12b, vdHLK18]. Spatial [BCKL12, Bur11, CRS12, DP17, EP14, FK18c, FN13, Ker10, KT11b, LN13, LRL17, PCMM18, TD19, TLC13, Wei13]. Spatial-Temporal [KT11b]. Spatially [BD16a, CL19, DFR14, DOR15, FGJ14, Hag15, HMU13, He14, Lan16, Lu12, MSLT16, MWY16, Tou14]. Spatiotemporal [HSUG13]. SPDEs [BK17b]. Special [Ano11d, Ano11e, Ano15d, Ano18d, GPMSBBSV15, GvdHdHM18, GN19, JP18a, Koi18, KT17, Leb19, MPR16, Nán11, TA12, WP11a]. Speciation [Sch13a]. Species [Bis19, DSZ17, MC19, Sch12b, TW13b, AMT18, BSS19, CLTC15, CHHK19, FC11, Jan15b, NS11, Van17]. Specific [KKA19, WZIG14]. Specification [Var12]. Specificity [KKC12]. Spectra [BBLP12, BBLP13, DG19a, LSBS13, MR13, RM16b]. Spectral [AEK16, BF11, BS13b, BS20, BNY16, Bis19, BBS^{+}15b, BP15, CLSW17, Dud13, DJW10, GO11, HIK^{+}18, JM14, KSM16, MHD17, NS11, NG10, PMC15, SLM12, Shi15, Sod11, SZ12, TV16, YB14]. Spectroscopy [BL15]. Spectrum [ACG15, Ast12, Ast13, BW10, BvE11, BG11, BV11, Els12, FZ10, GBTL17, LYT16, LY16, O’C12a, Ong14, RSB10, Shi16, dOP18]. Specular [AN19]. Speculative [CC18]. Speed [DKS18, FCK15, GMM18, HH15a, LX17, PSK10, PRD11]. Speeding [OTNN11]. Speeds [CPRY11, Kos13]. Speedy [BFM10]. Spencer [Ano18d]. Sphere [Bae11, BT14, Cha14b, D’O14b, HKLN19, JL16, KW12, NBK14, QLCL16, RG17, RM16a, ST16a, Sam18, Sim14]. Spheres [Bae11, Fan17a, MdSB18]. Spherical [AC14b, BL16a, BLW18, BGL^{+}11, BKLO19, CHHS15, CP17b, Her13, May13, Pat11, Pat17, UK16, WP11b]. Spherically [Mar11a, RBGV12, RBGV15]. Spherocylinders [Tor12]. Spike [MNS12]. Spiked [PMC15]. Spiking [And19, Che18, FGGL18, GRT17]. Spiky [Kos13]. Spin [Afz12, AP11, AD11, ANSW14, ANSW16, Ark13, AJ19, BLW18, BFKP10, BBS14, BM11, BBH11, BS16, BC11, BP12b, CBG14, CP17b, CFMT17, CR16, Con18, CMS13, CG12b, Cug17, DK10, EFO11, EL12, EKD12, EHR12, FdHM14, Fyo19, Gar19, Gue13, GV12b, Hag13, Har11, Imb16, IPS^{+}10b, Ito16, Ito19a, Ito19b, KMTC10, KE10, KL19, LHZ^{+}19, LO18, MD10, MP18b, MM14, O’C12a, Pan14, Pan16, PT14b, Par17, RVY18, SST14, SN13, Tas18b, WK18, Wre12, Mac13]. Spin- [DK10, EKD12, MD10]. Spin-1 [GV12b, LHZ^{+}19]. Spin-2 [LHZ^{+}19]. Spin-Flip [FdHM14]. Spin-Flop [CR16]. Spin-Ice [Cug17]. Spine [SZ12]. Spinless [Hon19]. Spins [BK17a, BIM18, DGGvdH18, Dym15, LS16a, MN14b]. Spiral [DG15c]. Spiral-Shaped [DG15c]. Splitting [KRK14, LNP19]. Spohn [Ano18d, EP14]. Spontaneous [Aum15, Bax11, Bax12, DHK11, ED15]. Spreaders [BHMGM13]. Spreading [LNP19, LLS13, PSC18]. Square [ADS^{+}19, ALS14, CLSW17, Chh12, DLS10, GBL16, KID^{+}11, MC10, Thi16, Zoc18, SS11a]. Square-Free [ALS14]. Square-Well [KID^{+}11]. Squared [KT11a]. Squares [CDTA10]. Squeezing [SZS15]. Squire [dZS13]. SRB [GBL16, Lia19, MT16, MU13, You17]. SSEP [FT19a, Van17]. SSH [Miy12a]. Stability [AV16, BP18, BS15f, Cen13, CN14, Con10, DeV19, DFR14, FS11b, GPMSBBSV15, GdK10, Git14, HX15b, HR16, HKLN19, HR23, Kie11, KK16, Lan15, MN15, Miy19, ROS19, Yuh15, ZW10, dVO15, LS10]. Stabilization [HR10]. Stabilized [MW10]. Stable [AT18, CGS15, DXZ14, GW12, GLT15, GW15b, Kaw16, KV17, LS14a, OW11, Par17, ST18a, Tos16, ZGL15, ZGL13, dHNT11]. Stable/ [dHNT11]. Stage [LH13a]. Staggered [KL19]. Standard [BAC13, TV15a]. Stars [Ngu17]. Starting [Kat12a]. State [ALS18, AP10, AF14b, AP11, AS16, BHNY15, BT12, BCJ15, BKK15, BW12b, Buc16, BC17, CEL^{+}18, CG14a, CS15, Che14, DDC18, EC11, GLM^{+}15, GS17a, HX15a, Hog11, HMW19, HRW14, Kra16, KR15, LY13, LN19a, Luc16, MC19, Miy12a, NMV11, Neu14, PT14a, SB15a, SST14, Tor12, TC11a, TC11b, Tzi13, VAY^{+}12, WPB15, Wre12, vEdG11, Mat15]. State-Dependent [HMW19]. States [ANS18, AZ11, BP18, BBH11, BGJLL12, Bis19, BCKL12, BCPS18, CN13, Coq15, CvELR18, DG17a, ELX18, Eva16, EWSR16, EV11, FBR19, Fil16, FS14c, GRS12, GT12, GT15a, GS12, HR16, HKR16, HR23, Har11, IY14, KBSM16, KKN12, KSY13, Kem11, KNiST11, KNiST15, KW15, Lia13, LY10, MS11a, MS19a, MMST13, MPTV12, Mat12, Nak17, Ngu17, O’C12a, Pir18a, RC17b, RY12, ST11a, Sch13b, Sch12c, SS16, Shi16, SST15, SW15, Tas19, XD18, dSRT15]. Static [CL16, DO15b, WCX^{+}11]. Stationarity [Dor16, Kem10]. Stationary [AP10, AB18b, ABT11, BL18, BPS12, BT12, BL10b, BGL14, BCKL12, Buc16, CG14b, DG17a, DJRZ11, ELX18, FS14c, GS12, HDS15, HL18, Ily12, Ily16, IS13, KC18, LN15, MT17, MS19b, MPTV12, MC19, PZ19, RELV11, RS18, Thi16, VAY^{+}12, vGRS16, CYZ18]. Statistical [Abr17, Ano11d, Ano11e, Ano12a, Ano12b, Ano15a, Ano16a, Ano18a, Ano19a, Ano19b, Art19a, AN19, Bae11, BM18, Bar14, BVL16, BC19, CdlL10, CM11, CCGT10, CG11c, DI13, Des11, Dym15, Ero14, Far15, FC11, FCLK14, GvdHdHM18, GS17b, GS17c, GT10, Gou14, HA13, Her13, ISZ16, IiS15b, JP18a, Kad14, Kar11, KT15, Kie13, KWZ14, Kle15, Lan19, Leb10, Leb11, Leb12a, Leb12c, Leb13b, Leb13c, Leb14, LBB15, LLM12, LLJH10, Luc18, MSS^{+}11a, MP10, MSS11b, NMV11, NiS19b, NSW13, Pel11, PCMM18, PPK11, RBM^{+}18, Rab11, RM16a, RVB16, Rue14, Rue17, San18, SdlPRA16, Sit11, Sta11, Täu10, TW17, Tur13, ZHZ15, Zia10, del12, vKSZ18, AFI^{+}10, Bré10, KRBN10, Rei09, San13, bA10, bA12]. Statistics [AEK16, BGL10, BPZ13, BC12, BGLL13, BGP15a, BJP17, Bur19, Các14, CMV16, DMS17b, Esl17, GMT17b, GMT17a, HLZ17, He19, Lam19, LFW12, MK19, Mia11, Nak14, NT18, PWZ16, PMC15, Pet10, RVY18, SP16, Shi15, TKK15, Tho12, YM11]. Steady [CEL^{+}18, DDC18, Eva16, EWSR16, KNiST11, KNiST15, LY13, LY10, Mat12, RC17b, RY12, ST11a]. Steady-State [DDC18]. Steep [BN18]. Stein [Mac13]. Step [BS15g, DDHS17, ZL13]. Steps [Caë10, Cli18]. Steric [MMM15]. Sticky [Bun14, Fan17a, Ngu19a, Sch13b]. Stiffness [CS16a]. Stigum [LJN18]. Stillinger [AG12b]. Stimulated [HV14, Vid15]. Stimulus [Che18]. Stirred [HHT15]. Stochastic [Abr17, AMW17, AGJP19, ABT11, AG15b, BL18, BCJ17, BBC18, BJ16, BGLZ19, BAS16, BFT10, BDL16, BC14b, BNT13, CT13, CGGR13, CGRS16, CG14b, CRS12, Con10, DXZ14, DT18, DHK11, ES12, Els12, FSV10, Fil16, FKK10, FC17, Fre14, FK18c, FW17, GL13, GRT17, GOP14, GS11b, GHMR17, Gro19, HVW12, HMW19, HE17, IiS17, II13, Jaf15, JMSW13, KNK15, KL15, LK14, LN13, Leh13, LNP11, LNS^{+}12b, LNY16, LLL17, LCZW15, Liu15a, LC19, Luc12, LKR^{+}11, MBGK12, MHD17, Nán16, NS12, OC12b, OVC14, PT14a, PCMM18, Pir14, PZ15, Qia10, RBS16b, SCY^{+}12, Sim10, Tou12, Tou18, WQ10, WXX16, WTM19, Wil10, XTpXpH12, Yag16, Yam17, ZDG19, ZYZ^{+}18, ZWGM13, San12]. Stochastically [ADF18, GM15, RE13]. Stochasticity [DKS18, ED15]. Stoke [IiS15a]. Stokes [ABH^{+}17, BLS17, BNT13, DLS10, FRT15, JLMG11]. Storage [DHL^{+}17]. Stored [AM14, CLT19]. Strategic [ABC10, PST13]. Strategies [BPR13, CLTC15, RCV16, ST11b]. Strategy [GPD16, JMRC16]. Stream [Vid15, Vid17]. Strength [KPS19]. Strengthened [Tos16]. Stress [RC17a]. Stretching [BSV18, Bur11, NKK19]. Strict [CSS15, FPR11, KM18]. Strictly [TS19]. Strings [DMP17]. Strip [BR10, Bou13b, CS15, DGK^{+}11, MRR19, Nig15]. Striped [MS16]. Strong [BG12, CL19, CF15, CLL18, GTT19, HTX^{+}12, Lee18, Leh13, LC19, MS19b, MQW18, MAPS11, Pes14, PS11, SJHW11, Ste10]. Stronger [FN13]. Strongly [SH12, dMP12]. Structural [Cas14, LNP13a, LZHS19]. Structure [Abr13, BAC13, BHS^{+}12, BS13a, BCKL12, Bun14, CM11, GdHR18, GLT15, KJZ18, KRRS17, KL13, KKA19, MC17b, Mas13, MM17b, MR13, Pan14, Pan16, PS11, TKK15, Zhu17]. Structured [BC15b, CG16, DOR15, HCI14]. Structures [AD11, BK17b, Cra13, FK18b, HSFK18, IS16, KT11b, SW11, SB12, SVW12, DK09, Ara11b]. Studies [dlLL11]. Study [AD15b, ACL14, CMV11, CS19, CIM14, Cli18, FDR12, GHLV16, HTX^{+}12, HP11, JYZ11, KBB19, PKDK13, Rab11, RM16a, SB14a, SB14b, SSR12, SSB15, Tak15, TFES19, Tid16, ZSHL15]. Studying [Mül11, Tho12]. Stylized [BA14, OEA18]. Sub [ABT10, BL16b, DSZ17, Sta15, Yar14]. Sub-Critical [Sta15, DSZ17]. Sub-exponential [Yar14]. Sub-relaxation [BL16b]. Sub-sampling [ABT10]. Subcarrier [GI19]. Subcritical [BBLP13, CMM16, MvS19, Xue16a]. Subdiffusion [MBWC16, Naz18].

Subdiffusive [MOW11, OW11]. Subexponentially [WK18]. Subject [GS11b, GG11a]. Sublinear [MN12, OPS10]. Subordinated [ZGL13, ZGL15]. Subsamples [EF18]. Subshifts [MU15, Ong14]. Substitution [RST11]. Substitutions [MTVU18]. Substrate [CCD15, Zho18]. Subsystem [Los17]. Success [CLTC15]. Sudden [MR12]. Sufficiency [MN16b]. Sufficient [Tem14]. SULE [dOP18]. Sum [AF16, CDTA10, GV15, Kar12, SJ10, TF12]. Sum-of-Squares [CDTA10]. Summability [CH11, LMN18]. Sums [ABA14, AB18b, DP19, GMT17a, HL14, KV17, OP12]. Sunday [Ano12a, Ano12b, Ano18a, Ano19a, Ano19b, Leb12a, Leb12c, Leb13b, Leb13c, Leb14]. Super [CMGP14]. Super-Hydrodynamic [CMGP14]. Superconductivity [AEWI14, TK16]. Superconductors [Wre17]. Supercooled [AG12b, GPGA17]. Supercritical [BBLP13, CG11a, CG15, CMM16, CdLS13, Tzi18]. Superdiffusion [BGJ^{+}15b, CLT19, KNSS18]. Superdiffusive [KT19, TV12]. Superfluid [Afz12]. Superfluids [Wre17]. Superhomogeneous [GL17a, GL17b]. Superintegrable [IPS^{+}10b]. Supermarket [BFL18]. Superpositions [GLST16]. Supersymmetric [CTT11, Mor18]. Supersymmetry [HF12a, Hag13, MM14, RKGZ12]. Supply [GNPS13]. Supported [Kor16, TT16]. Supports [BD15b]. Suppressed [Cla13]. Suppression [BdSPMS14]. Supreme [LBB15, Lee18]. Supremum [QR13]. Sure [CYZ18, Hay15]. Surface [BGL^{+}11, GL16b, KNPF19, Koi10, NiS19c, RBGV12, Wu14]. Surface-Mediated [BGL^{+}11, RBGV12]. Surfaces [Bab12, CTH^{+}11, CDV17, KA17, Zho17]. Surgailis [Thä11]. Surjective [KT15]. Survey [BB15]. Survival [BBS11, CC19, JMRC16, Kat12c, KMS14, MMR18, NG10, Sch15, Zuc11a]. Susceptibility [CGNP11, FL12, TW14]. Sustainable [Wei13]. Sustained [CDP17, CF19b]. Swapping [BR16, Ker13, LVE19]. Swarming [DM11, HKLN19, vBKBS13, vBSB15]. Swept [CB16]. Swim [Bus18]. Switches [PS10b]. Switching [BELP15, PS10b]. Sy [KPS19]. Symbolic [Osi16]. Symmetric [AG15a, BGG10, BL10b, CRV13, Dai17, DR13, LJN18, MPTV12, Ngu19b, PZ17, Ras12, RBGV12, RBGV15, TS19, Wan12]. Symmetries [Esl17, GS11c, Liu15b]. Symmetrized [Mat12]. Symmetry [Aum15, BN14, BJGL^{+}17, BCHM12, BSS19, BS15a, BC17, CGRS16, CC19, Con18, FP11, Fre14, He19, IK10, Ito17, Ito18, Kat12a, KLMP18, LSBS13, OiS14, Tak09, Tak10b, Tas19, Sha10, Tak10a]. Symmetry-Reduced [BC17]. Symplectic [Gou15, LZ10b]. Synchronisation [GT17]. Synchronization [CR16, DeV19, GdHMZ19, KKV^{+}11, Ko19, TC11a, TC11b, dAPdA^{+}13]. Synthesis [ZDS11]. Synthetic [MLS16]. System [Abr17, And19, ABFP15, BNTT16, BO14, Bha15, BA14, BDL10, Bra14, BFNZ11, BC17, BP15, CL13a, CCM16, CH14, Cor16, DM11, DF16, DK10, DFL17, DLS10, DL17, EKD12, FL15a, FG12, GBTL17, GG11a, HZS11, HR10, JSV19, Kat12d, KS12, Kua13, LNS^{+}12b, Mar15, Mas14, MR12, NS16, NT17, Ono11, PS10b, Qia10, RM16a, RE13, SGU13, Shi13, SSE15, TD19, TZ16, TS19, Vau10, XJZY13, XLL^{+}15, Yag16, Yam17, ZW10, Zha10b, ZYZ^{+}18, Zho18, vHL13]. Systemic [BA14, MBC^{+}13, TGGS13]. Systems [ABC10, AEWCD19, AB17, ACL14, ADF18, Ara11a, ADP14, AFFR17, BW10, BKM15, BJ16, BP18, BCJP19, BKLL12, BPR13, BC11, BC14b, BP16, BGTVE16, BP12b, BKP13, BGTV11, BCM12, CA19, CMGP14, CNZ17, Cha12, Cha14b, CCR17, CYZ18, CL10, CG14b, CL15, CCR19, CH11, DS17, DT18, FGJ14, FL12, Fer14, FdHM14, FF10, FGGL18, Fil16, FKK10, FF11c, FMM^{+}14, FMM^{+}15, FL18, FL20, Fra11, Fre14, FFT11, GL13, GQ17, GRV10, GS11c, Grm17, HY16, HBC^{+}15, HL14, Hor16, HŚ16, HFWT15, HEdPG14, IPS10a, IiS17, Ito16, Ito17, Ito19a, Ito19b, JV19, Kie14, Kle15, KCB13, KM19c, LK14, LMM16, Lee11, LO18, LK19, LL19, Luc12, LFW12, LFWK14, Luc16, Luc18, LSBS13, LH13b, Mal12, Mat12, MN14b, Mie18, MU16, MM17b, Miy19, Mon12b]. Systems [MB11, MPR16, Mor12, MAPS11, NSS12, NiS19a, NSV12, O’C12a, Ohk10, ORW15, Pav11, PSS15, Rab11, RC17a, Ren18, RSB10, RDNS15, ST11a, SGS19, SVRL11, SST14, Ste19, TWT14, Tas16, Tas18b, Tas18a, Tas19, Tem14, Tho12, Tia14, VP14, VP15, VW18, Ven14, VL12, WLEC17, WG18, WL13, Yar14, You17, YSSH13, YS13, ZP15, dSLPV17, ZP16]. Szego [BDES19].

T [FAB16, Kos11, KKC12]. T. [Ano18d]. Tackling [BCS18]. Tadmor [Jin18]. Tagged [Bur19, KMS15, NKK19, Per10, RNDS13, RDNS15]. Tagged-Particle [Bur19]. Tail [DRT18, MPS14, OO18, Thr18]. Tailed [Nán11, Pir14]. Tails [AHDV17, Ast12, BS11, BBW15, BN15, FM18, Kaw16, KV17, KKR18, vdHKvL18]. Takata [Sha10]. Tale [LNP13a]. Tangent [CS16c, DG19b]. Tardos [AP14b]. Target [HNVZ13]. TASEP [BL16b, BGL14, CS10b, CFP10, EK19, FN15a, Gup16, PPS11, SLM15, ZDS11]. Taylor [BS20, BS13b]. Technique [CDCL18]. Techniques [FJLS18, Fre15, GT12, OMC11]. Telegraph [AEWCD19, TP19]. Teleportation [AK18b]. Temperature [Afz12, AB19, AD14, ABS12, AG12b, ANSW16, Ark13, BL18, BCM10, BGP15a, BL11a, BS16, Bir18, BU18, BP12b, Che14, CP17b, CFMT17, Cla19, Con11, DSS15, EK10, FMM^{+}15, Fra11, Gar19, GNS18, Hag15, IST15, IY14, JMSW13, Jan12, Kem11, Ker10, KE10, KR10, KKA19, KM19a, LHZ^{+}19, Lim16, MdG13, MS19a, MD10, Moh17, NT18, PSS16, RW11, SSR12, SLM12, SZ18, SH16, Thi16, WW16, Wei18, Zoc18, vHL13]. Temperature-Dependent [KKA19]. Temperatures [CEL^{+}18, CRL15, DK10, JJ12, Pir18b]. Tempered [CWD17, DWTW16, GW12, OW11, ScCdP11, Zha10c, ZGL13, ZGL15]. Tempering [Ark10, BR16, LVE19, ME11, UK16]. Temperley [LS16c]. Temperleyan [DG15b]. Template [PS16, XZ19]. Template-Assisted [PS16]. Temporal [BL17, DS17, KT11b, SB14a, SB14b]. Temporary [TSS13]. Ten [LSBS13]. Tension [KNPF19]. Tensions [BBD^{+}11]. Tensor [BMT15, CS16a, EV11, LM17]. Tensor-Valued [LM17]. Term [JSV19, SV15]. Terms [Kan12, Qia10]. Ternary [Wil11]. Tessellation [SWK^{+}18]. Tessellations [KD19, Thä11]. Test [BVL16, BNP14, Liu15b, TT12]. Tests [CM12b]. Tethered [MMSY11]. Tetrahedral [GPGA17, HJR^{+}11, KS13]. Their [AI12, BF11, CCR17, Far15, Fra17a, GLO10, Kac13, KMS14, MSZ12, PCM15]. Theorem [BPZ13, Bec11, BEP18, BJM15, BMR10, BSS14, BMSS13, CCEF10, CS10c, FL14, Fre15, Hal19a, Hal19b, Hor16, KSY13, Kha19, LYX17, LLM12, LLL17, MN14a, Mae14, MBC14, NS10, NSV12, Pol17, RVY18, SB15a, Sun18, Tas18a, Tzi18, Wat19, Wat20, WES11, Cha19, JL17a]. Theorem-based [LLM12]. Theorems [ALS14, CG11b, DS17, Gar19, GGvdHP15, HSZ19, HNVZ13, KKSY19, KY13, LM12a, RvH14, TK16, TWT14, Tos16]. Theoretic [DM12, LO17, NP16]. Theoretical [XZ19, Zha13]. Theories [Lan15, ZJ14]. Theory [AMW17, AP11, Ari11, Ark13, Ast12, Ast13, AK19, BKM15, BR10, BBR12, Bae11, BVL16, Bec10, Bha15, BL15, BNT13, CVE14, CG14a, CFM14, CG16, CPRY11, Dal11, DM12, DM13, DOGK16, DOGKP19, Fan17a, FBE^{+}11, FL15b, HX15b, HT15, Hon19, HEdPG14, IiS15a, JP18a, KA17, KV15, KV16a, Kie14, KWZ14, Kuo17, LZ15, LS16b, LLJH10, LN15, Luc12, LFWK14, Luc16, LRL17, Luc18, LMN16, MS10, Miy13, Mor12, O’C12a, Per13, PC19, PE19, QLCL16, Ren18, RBR11, Rue17, Sch13c, SR13, Sin10a, ST16b, SW15, TH12, Tak15, Tay16, Van17, VAY^{+}12, Vie16, VL12, Wil10, WG18, WL13, ZW12, Bré10, Kar11]. There [CE14]. Thermal [BR11, CEL^{+}18, CDS17, GS17a, GT10, ILOS10, LNY16, Mas16, PT14b, SZS15, Ste19, Tas16, bA10]. Thermalization [HJ17a, Luk14, LMN16, Tas16]. Thermally [JPS17]. Thermodynamic [BP18, BCJ15, BP11a, BGJLL12, Dzu11, GTT19, Grm10, KNiST11, KNiST15, LNT13, MBC14, NP12, PT10, SB15a, Tou15, Wre12]. Thermodynamics [ABFP15, AGMM^{+}12, BMC17, DFF18, Far15, GPMSBBSV15, GS11a, GQ17, Grm17, HEdPG14, Kim12, Lan19, MN15, MSLT16, NiS19a, RM16a, RBR11, Sam13, SZ19, Tak16, TKK15]. Thermophoresis [LW19, San18]. Thermostat [BLV14, Eva16, FG11, Fre16]. Thermostat-Like [Fre16]. Thermostated [Art19a, BCE^{+}14]. Thermostats [GP10]. Thermostatted [BCKL12, CMW15]. Thickness [TW17]. Thin [CCD15, DOGKP19]. Things [Dia11]. Think [AWE16]. Thinning [SVW12]. Thinnings [HL18]. Third [Ban10, CFLV19, Liu15b]. Third-Order [CFLV19, Liu15b]. Thompson [APRT17, Han15]. Thomson [CGS15]. Thouless [EC11]. Three [BELP16, BML12, CLTC15, DF16, FRT15, FL18, FL20, Gal13, Ken15a, Kol17, KR15, LS16a, Mob13, MFLA15, PF17, Pos16, PB11, QLCL16, Sou18, Tzi13, Yam17, YYZ11, dSRT15, vEF12]. Three-Dimensional [FRT15, Ken15a, Kol17, MFLA15, Pos16, PB11]. Three-Party [Gal13, Mob13]. Three-State [KR15]. Threshold [CG14c, FDR12, HMU13, LN19b, Xue12, Xue15, Xue16a]. Threshold-One [Xue16a]. Thresholds [Bal14, Cha12, Cha15a, vEF12]. Thue [LQY17]. Thymic [Kos11]. Tight [BvdHK19]. Tightness [HI18, LJN18]. tigrina [CLTC15]. Tiling [Tel10]. Tilings [FF11c, NS10, Nov15]. Till [KMS14]. Tilted [CL13b, LPK13]. Time [ACG15, ABR18, ASA15, AGJP19, ABT^{+}14, Ara11a, AF14c, AY10, BGL10, BL16b, Bao17, BGLL13, BL10a, BL12, BP11a, BGL^{+}11, Bla10, BC14b, BPF^{+}14, BBR^{+}19, BM12a, Bur17, BCM12, Các14, CP17a, CJW17, CCM16, CP18, CF11, CFM14, CL16, CNS15, CEGW18, CMM14, CGN16, DMS17b, DD15, DPT17, DMM14, DT18, FL15a, FZ12, FF11b, FRT15, FM18, FG12, FGN14, FK11, GLM18, GN19, GG11a, GMT17a, HR18, HIK^{+}18, HMW19, JP18b, Jia14, JPS15, KKV^{+}11, Kar10, Kat12b, KMSS19, Kol17, KS12, Kua13, LL13, Lan16, Lee19, Leh13, Li19, LPS19, LK19, LNT13, LN15, Lu13, Lu16, LQR12, MGZ14, MCG12, MP18a, MC17b, MNBC19, MM13, MG19, MDP^{+}18, MQW18, Mis16, MGMMP13, NSV12, NZB16, Ohk14, PPS16, PK10, Pel14, Pic10, RL17, RM16b, RBGV15]. Time [Sak18, SWB10, Tag15, TH12, VYH11, WZL^{+}14, WG19, Yag16, ZDG19, ZGL15, ZL19, dOP18, ABMP16, MC17a, Zuc11a]. Time-Asymptotic [Lan16]. Time-Based [CGN16]. Time-Delay [Các14]. Time-Delayed [MG19]. Time-Dependent [FK11, NSV12, TH12]. Time-Fractional [Ara11a, CJW17]. Time-Inhomogeneous [GN19, Ohk14]. Time-Periodic [WG19, dOP18]. Time-Scales [BP11a, BC14b]. Times [BBP17, BO16, BLM13, BK11, CNZ17, GW12, GdK10, GV12a, HCI14, HY16, HY17, IPS10a, KOT11, LS16b, LSY18, MS19b, MS12a, OW11, RT15, TM10, TGP12, ZGL15]. Timing [MNS12]. TiO [ML15]. Tip [BD11b, dlLL11]. Tipping [BELP15]. Tissue [Pos16]. Tobochnik [bA10]. Toda [DKLS19, ILOS10, IST12]. Today [Sin10a]. Toeplitz [BDES19, Kac13]. Toeplitz-Like [Kac13]. Tomorrow [Sin10a]. Tongues [dlLL11]. Tonks [Jan15b]. Too [TGGS13]. Toom [Jus10, dMP12]. Top [GMT17b]. Topography [WLJH18]. Topological [ADH12, BD16b, BD17, BDY17, JV19, LZHS19, MP14, MBC^{+}13, YY10]. Topologically [Mar11a]. Topology [FD14, RMN15]. Tops [SS11b]. Tori [CFG13, Xue15]. Toroidal [DG15b]. Torsional [XZ19]. Torus [BCL19, HI18, LLM19, Mit17b, Pal11]. Total [CP17a, Lis17, Thä11]. Totally [CCG14a, DG17b, Kan14, KMO16, Pro15]. Tower [Tas19]. Toy [Art19b, DR14, Tak15]. Toy-Model [Tak15]. Trace [Han14, LZ10b, LQY17, Ras12]. Tracer [BT12]. Tracers [LZ10a]. Traces [HCO15, Zha10a]. Tracking [CJW17]. Tracks [NVL11]. Tracy [Sod17, AD14, Sod09]. Trade [MMW16]. Trade-Offs [MMW16]. Tradeoff [SKT16]. Trading [TBD13]. Traffic [CP15b, EK19, Klu11, LLJH10]. Trajectories [OVC14, RT11b, RT12, ZBVE11]. Trajectory [BCJ17, RW14]. Transcription [Klu11, MSS11b, SKT16, XZ19]. Transcription-Factor [MSS11b]. Transducer [BC15b]. Transfer [BBR^{+}19, CP10b, CG11c, CDS17, CR11b, DUU15, KC18, KV15, LZ10a, MU15, O’C12a, PGS12, Pol17, SS11a, SS16]. Transform [AU15, KS12]. Transformation [CGN16]. Transformations [BPS12, BC15b, BGJLL12, Liu15b]. Transforming [Ken12]. Transition [APRT17, ARBJ15, AR11a, ABJ12, BLW18, BJGL^{+}17, BL11a, BC10, BM17, CC18, CS10b, Che14, CDdS14, CFLV19, Dai17, DKKP14, DM13, DH19, DR14, EC11, FGGL18, Fow19, GS13, GNP16, Hep18, HY19a, HKW11, JYZ11, Kol17, KY12, Lee11, MW12a, MMR18, MS16, OiS14, Ono11, OR16, PST12, PS11, dAPS11, RMS19, RL17, RM16a, RBS16b, Sim11, SM12a, SM12b, SG17, SM14, Tak16, TN18, TN21, TD19, WLEC17, Web11, XLL^{+}15, Xue16b, Zho18]. Transitional [RM11]. Transitions [AE16, ADMS11, BG17a, CT13, CRS14, CP16, CL15, CNS15, CDH15, CR16, DOP19, DLY18, DK10, Dyk14, FZ11a, FdHM14, FMM^{+}14, FK17, GHRR13, GSSV11, GS19, KRRS17, KS18, Lep15, MW10, Mon12a, MBS16, RT11a, RW12, SZ18, SGC11]. Translation [CL15, GRR17, GKLS17b, KR15, KRK14, NVL11, RS13, WW16, ZDS11, GKLS17a]. Translation-Covariant [WW16]. Translation-Invariant [GRR17, KRK14, RS13]. Translational [Esl17]. Translationally [KC19]. Translocation [LZ15, MMM15]. Transmission [BN18, CGL17, CGL18, CRS12, HCLR11, MMW16, YH14]. Transmission-Reflection [YH14]. Transport [AG12a, BN18, BGLL13, BO11, BG12, BG17b, CR14, CGGR13, CGRS16, CCH^{+}14, CCR19, Coh09, Coh10, CPSV10, CHHK19, CCFR18, DKLS19, ES13, Gol10, Huv12, KBLL13, KU11, LNY16, LM12a, MMA15, Mül16, MS11b, Nán16, RPPF15, SWB10, TP15, WYG16, Zha13, ZL19]. Transverse [Ito18, MW12b, SML19]. Transverse-Field [SML19]. Trap [TKK15]. Trapped [CSC11, Far15, RSY14]. Trapping [BLRVR13]. Traps [Kos11]. Traveling [CQR13, Man11, SBK10b]. Travelled [KMS14]. Travelling [BD15a]. Treatment [Ono11, del12]. Tree [AMS14, ART11, AP18, BCHW19, EHR12, FK10, GRS12, GRRR13, HNT18, KR15, KRK14, LN19a, MW12b, Mis19, MBS16, OO18, Pah10, Pat17, RRS12, Tro10, YY10]. Tree-Like [AP18]. Trees [ALS18, AP14b, AP18, BGMS14, BZ13, BU18, DRCV19, GHRR13, GRR17, GvdHW17, Goo12, JS11, JM14, Kar10, KS18, LPS12, LU19, MRCJ18, PK11, SZ12, TW11, Xue12]. Trend [PST13, WFK11a, WFK11b]. Trend-Driven [PST13]. Trends [AWM13]. Triadic [vdHvLS18]. Triangle [MS19c]. Triangle-Well [MS19c]. Triangles [PW13]. Triangular [BCM10, BCL10a, BEP18, CC14]. Triangulated [Koi10]. Triangulation [KY12, NT16]. Triangulations [DJW10, KSY13, KSM16, SYZ13]. Trichotomous [ZWGM15]. Tricritical [AB14]. Triggered [EFO11]. Trigonometric [AK17, Kie13]. Triviality [OO18]. Trivialization [FD14]. Trotter [DHK11]. Trouble [BG19]. Truncated [AHDV17, CdLS13, GMT17b, GMT17a, GO13, vEdLV16]. Tube [CPSV10, LE17]. Tubes [SLdEC11]. Tuesday [Ano12a, Ano12b, Ano18a, Ano19a, Ano19b, Leb12a, Leb12c, Leb13b, Leb13c, Leb14]. Tumor [RBM^{+}18, ZW10]. Tunably [FK12]. tuning [PG10]. Tunneling [BL10a, BL12, LMC11]. Turbulence [Abr13, BB15, BFT10, BF12, Bha15, Esl17, FG14, GS17c, Grm10, JLLP17, KV15, KV16a, KBB19, MK19, MS10, OBX11, RP12, Rue14, Rue17, TS12, TJ15, GS17b]. Turbulent [Che14, CDS17, ES13, GL14, HPF15, HMRW13, Ker13, VB11]. Turing [CSL19, TD19]. Turning [CDG^{+}15]. Tutorial [ZDS11]. Tutte [CD16, GC17, LXHAA19, PXX15]. Twenty [Zia10]. Twist [DO14a, Tas18a]. Twisted [JZ10, KKR18]. Two [AF16, AB14, AM18, AN15, AB18b, Bao17, BF10, BELP16, Bar16, BBBP11, BCFS17, Bis19, Bot18, BLZ14, BCM12, CG10a, CFTW15, CSL19, CNZ17, Car11, CG14a, Cha15a, CG11b, CCFR18, CMV16, DHS18, DSZ17, Dor16, DP14, Els15, Fra11, FK11, FAM13, GdK10, Gra13, Gri19, GT15b, Gup16, HI18, Har11, HTZ12, HMN15, HCO15, JP18b, JY18, Kar12, KBLL13, Kir12, Kol14, KM19a, KNSS18, LL16, LLM19, LNP13a, LH13a, LXHAA19, LS14b, LSW17, MGAPQH13, MC19, MCK15, MS16, Osi16, PSS15, Pos16, RBGV15, Ryt12, Sam13, Sam17, Sam18, Sam19, SLdEC11, SJHW11, SST14, SS15, TK16, TF12, TT12, Thi16, TT17b, TW10, TW14, Tzi18, Vau10, VW18, VB11, VFT12, WZIG14, Wu14, Yam17, Zho17, dHNT11, dHNT12, GOPS11]. Two-Band [BF10]. Two-Body [FK11]. Two-Component [AF16, AN15, PSS15, Sam13, Sam18, TT17b]. Two-Dimensional [AB14, Bot18, BLZ14, CNZ17, Car11, CCFR18, CMV16, DHS18, Gra13, GT15b, HI18, Har11, HTZ12, Kir12, KM19a, LLM19, MS16, Pos16, RBGV15, Sam19, TF12, TT12, TW14, VB11, VFT12, WZIG14, Wu14, BBBP11]. Two-Disk [Gri19]. Two-Electron [DP14]. Two-Fold [Osi16]. Two-Mass [HCO15]. Two-Particle [VW18]. Two-Phase [Els15, MCK15]. Two-Point [JY18, LSW17]. Two-Sided [AB18b, TW10]. Two-Species [Bis19, DSZ17, MC19]. Two-Stage [LH13a]. Two-State [SST14]. Two-Time [Bao17]. Two-Variable [Ryt12]. Type [ALS14, BCHW19, BF12, BCL10a, BD16b, Bla10, CE12, FP11, FdHM14, Fri17, Git12, GLST16, HR10, HH15b, KOSV18, Kuo15, LYX17, Mar11b, MU15, Mol18, MPS14, MBS16, Nán11, WGLE11, ZGL13, IK10, BA19, MC10]. Type-II [BA19]. Types [MSS15, dHNT11, dHNT12]. Typical [SMCF13, Tro10]. Typicality [EWSR16, Tas16]. Tzer [Ano19b].

U.S. [IK17, CC18]. Ueda [MT11a]. Uhlenbeck [GLST16, RE13, WX15, Wil10, WP11b]. Ulam [AMS19, BP11a, BCP13, BPP18, DDN14]. Ultimate [Bal17, Bal18]. Ultra [vdHK17]. Ultra-Small [vdHK17]. Ultradistributions [RPPF15]. Ultrametric [Nei12, vEdG11]. Umbrella [PVCG11, WES11]. Unbounded [BM18, BIM18, CFN15, DF18, HHV16, LQY17, PS10a, vdHvLS18]. Uncertainty [NS16]. Uncountable [EHR12]. Undamped [CL13a]. Underdamped [DNP17, NML^{+}11]. Underlying [BCP13, BPP18]. Understand [CM18, II13]. Understanding [KL11]. Undirected [FCLK14]. Unified [AG12a, Far15, Ras11]. Uniform [Caë10, Cor16, Hau16, HNT18, KMTC10, Lan16, Sch12c, SYZ13, Tem14, TP15, Vie16, BC10, Var12]. Uniformly [PRD12, Gre12]. Unifying [PC19]. Unimodular [KSM16]. Unique [DF17, GHMR17, SZ12]. Uniqueness [BE16, BM17, GLM^{+}16, HM13a, KM19c, Lau18, NV14, NTV16, WW16, ABMP16, EHR12]. Unit [DMY13, DLS10, GBL16]. Unitary [CCG14b, GO13, HK19, LZ10b, Shc11]. Units [SSR12]. Universal [DMS17b, IiS17, Ito17, KC19, LZ10b, LFW12, MiS18, TS12, TKK15, XTpXpH12]. Universality [BNXL11, CP14, CQR15, DLY18, GLM18, HSFK18, LMC19, PG10, PSS15, QS15, RMS19, Shc11, Shc14, SS18c, Sou18, Ven14, Zha15]. University [Ano12a, Ano12b, Ano18a, Ano19a, Ano19b, Leb12a, Leb12c, Leb13b, Leb13c, Leb14]. Unperturbed [BELP16]. Unraveling [WLL11]. Unstable [BC17]. Unsteady [Tak10b]. Updates [Hol11]. Updating [NS12]. Upgrading [CS10c]. upon [BTV14]. Upper [LU19, LSW17, NC10, PZ19, Yin10]. Urban [HSUG13]. Using [CSC14, CIM14, CDCL18, DM11, DLP16, DNP17, GRT17, HLZ17, Jus10, LTM16, LRL17, LKR^{+}11, MSS11b, MCK15, MBC^{+}13, NE16, PVCG11, PM17, RESA10, RPPF15, Tos17, WLL11, ZSHL15].

V [BS15g, RK12]. V3.0 [Kad14]. Vacua [BHNY15, Bis19]. Valency [LZHS19]. Valency-Based [LZHS19]. Validity [DP17, LLS17, WG18]. Valleys [DRCV19]. Value [AFFR17, FLTV11, Fed17, FFT11, FDR12, LFW12, MSB13, RL11, YK13]. Valued [BCF17, BN15, LM17, MWY16, PR15a]. Values [BP18, EHR12, FPR11, JS17, LNT13, ST16a]. Vandermonde [TF12]. Vanishing [BGMS14]. Vanishingly [LK18]. Vapor [Ber12, Ono11]. Variable [CJW17, CGL12, CG12a, GL13, GM15, GI19, Ryt12]. Variables [ABA14, BHS13, CM18, IiS17, Ken10, KV17, MCG12, MMST13, Moh11, WFK11a, WFK11b]. Variance [BvdHK19, CG19, DLP16, HSZ19, HM16, Ito19b, MN12, Thä11]. Variances [Kan12]. variate [Kie17b]. Variation [MW10, Ply15]. Variational [Ari11, BP12b, FdHM14, FD16, RLCMRT10, dHO13]. Variations [Qia10, YBF^{+}17]. Variety [BMC17]. Varying [BDL10, FGJ14, RL11]. Vastness [ALS18]. Vector [BCF17, DTW19, Kel19, ZJ14]. Vectors [CH11]. Vehicles [LH11]. Velocities [BO11, LSY18, Pal11, PR19]. Velocity [BCMP15, BV16, Ber18, FS14a, FS14b, GM13, HMRW13, Luk14, LMN16, MPL^{+}16, PRD12, PR19, RE13, YKS16]. Verblunsky [DMY13]. Vernon [Aur18]. Version [BAS18, Ber19, CC14]. Versus [Bai10, BK17a, CCFR18, Grm17, IST15, ME14, MDP^{+}18, Mob13, TKK15, WTM19, dSLPV17, BvE11]. Vertebrate [Per19]. Vertex [CS16c, CPS19, Cug17, FK18a, HF12a, KS13, SS11a]. Vertices [DEK^{+}15]. Very [BCM12, DUU15]. Vessel [TAG10]. VI [SS11a]. Via [Chu16, FC17, FCLK14, Kra16, ABR18, AK19, ACR18, BU18, BCM12, CGN16, DG19b, Ghe10, HHT10, IiS15a, JRS15, KOT11, KO15, LK19, LN15, MS19b, MMST13, Naz18, RW12, SZ19]. Vibrations [Kac13, OD11]. Vicinity [BP18]. Vicious [CB16, KMS14, Sch12a, SMCF13]. Vicsek [BAC13]. View [BC15a, BPR14, Fis11, Fis12, Leb11, KRBN10]. Viewpoint [HiS19]. Views [CG11b]. Villain [CTH^{+}11]. Villani [CDTA10, LS12]. Violation [GS11a]. Viot [Mar15]. Viral [LH17]. Virial [Jan12, Jan15b, Pro17, Tat13]. Viscoelastic [Pos16]. Viscosity [Mat12]. Visibility [BA19, IK10]. Vision [PF17]. Vlasov [BP19, CCM16, CCM17, CP10a, CH14, DF16, DRS18, DL17, FL15a, FKK10, Fow19, HR18, Kie14, Laf19, Lan16, Tri17, Zha10b]. Void [Shl17]. Volatility [AG15b]. Voltage [NML^{+}11]. Volume [CPV10, CP10a, CFS18, Coq15, GKLS17a, GKLS17b, KM19c, MP18a, PT15]. Volumes [PVCG11, Szá12]. Voronoi [KD19]. Vortex [Esl17, MS19a]. Vortices [KW12, Ser14]. Vorticity [TJ15]. Voter [BV15, HMN15, Mob13, MLCPS13, PT14a, SM15, SV10, Xue12, Xue15]. Voting [Gal13, Lee18].

W [Sta11]. Waals [TN21, TN18]. Wagner [Hal19a, CN14, Hal19b, KSY13]. Waiting [GW12, MNBC19, OW11, ZGL15, ZL19]. Waiting-Time [MNBC19]. Waiting-Times [OW11]. Wake [Hep18]. Wake-Sleep [Hep18]. Walk [AKM13, BFKR10, BBBP11, BSW17, BL19, BL11a, BS15c, BD11b, Caë10, CS16b, CMVW11, Dai17, DGK^{+}11, FvdH13, GPS13, GMM18, GMM19, Kar10, Ken15a, Ken15b, Ken19b, Mar15, Mis16, Nán11, Per13, PZ17, SS18a, Sco11, XD18]. Walkers [CD14a, DRS10, HJ17a, KMS14, Sch12a, SMCF13]. Walking [BDP19]. Walks [APdM^{+}18, AO19, APSS12, ADGPP17, AT12, ABF16, BBP17, BGLZ19, BCL19, BCLL16, BR13, CRS14, Caë11, CP15a, CLM15, CCR19, Cli10, Cli18, CG17, CMM14, CCFR18, DKY19, DMM14, Gil15, HIK^{+}18, JL17a, JSV19, Jav15, JM10, JM14, Kat15, Ken12, Ken16, KKSY19, KY13, Kos13, KT19, LMT15, LS15b, LS16b, LMM16, MSZ12, PS10a, Pel14, RW11, SWB10, SH16, SV16, WZL^{+}14, XY13, Zuc11a]. Wall [ELO11, Sam15, Sam16, ST18a, SSB15, TAG10, YKS16]. Wall-Adjacent [YKS16]. Walls [CFLV19, ST14, TJ15]. Walrasian [BVL16]. wandering [AC14a]. Wang [KBSM16, WLL11]. Wannier [MP14]. Ward [AW18]. Warren [FF14]. Wasserstein [Hau16]. Water [RRW11, RVB16, SvHM^{+}11, VCT11, WP11a, Dzu11]. Watson [BGMS14, BZ13, SL12, Zuc11b]. Watson-Like [SL12]. Watts [KV16b]. Wave [BDG^{+}14, BP19, BD15a, Kad13, KKA19, Liu15a, Man11, MC10, PT14b, YM11]. Waveguides [BV11, KKR18]. Wavelet [Osi16]. Waves [MS17]. Way [TGP12]. Weak [BTV14, BV11, CE12, Coq15, CSS15, CG17, FRT15, KV15, KV16a, LKD12, LC19, MQW18, MAPS11, PR15a, Ryt12, Var12, Zha10b]. Weak-Disorder [BV11]. Weakly [BSW17, BBC15, CL16, LK14, MOT14a, MU15, MAPS11, Naz18, O’C12a, SS10, Sim11, Tri17, Ven14]. Wealth [DLR14, ZGH18]. Web [PXX15, Sch10]. Weber [AG12b]. Wedge [CMS10]. Weeks [Fan17a]. Wegner [TV15a]. Weight [BvdHK19, BM11, MP18b]. Weight-Distances [BvdHK19]. Weighted [AK18a, BCCD18, BDL11, CS10a, DLY18, SX10a, SX10b]. Weighted-Set [SX10a, SX10b]. Weights [Cal15, CJW17, FK18a, Ker10, XP17]. Weiss [AT18, Col14, CF19b, EK10, Fed14, GNS18, GV15, LM12b, OK14, OEA18, PFR13, SZ18]. Well [ACG15, FS11b, Han18, HJ17b, KID^{+}11, LTR17, MS19c]. Well-Posedness [Han18, HJ17b]. Wentzell [dSLPV17]. were [Bri19]. Wetting [AB14, ADMS11, CDH15, Zho18]. Weyl [AGR19, CP19, Mas14, TV16]. Where [Hol11]. Which [Coq15]. White [Ngu18]. White-Noise [Ngu18]. Whole [FL15a, MZ19]. Whole-Plane [MZ19]. Whose [LK18, RGL11]. Wide [Lee18]. Widom [Sod17, AD14, BDES19, DH19, FKKO15, Fri17, Sod09, Zoc18]. Wiener [Jia14, NP16]. Wigner [AF14b, BP12a, BPZ13, He19, JLMG11, KS12, PWZ16, PRS12]. Wilson [BKPW14, Bré14, Gin14a, Gin14b, Pes14, Uka15, Weg14]. Wind [Tro10]. Window [LK14]. Windows [LH17]. Winfree [Ko19]. Wires [BBR12, KMS19b]. Wishart [Kum19, RKGZ12]. Within [MN15, MSB13, Pat11, TP19]. Without [BV16, Ber18, CC19, Gal15, MWY16, Pic10, RM16a, Tri14, Vid17]. Witness [AP14b]. Witnesses [HA19]. Witting [HY19b]. Wolf [CTH^{+}11]. Word [AWM13]. Words [HM16, Mol14]. Work [Bat17, Dav11, LTR17, SB15a]. World [APRT17, ALAF18, GS17a, KV16b, Sha12]. Worm [CGHT16]. WR [MSS15]. Writhing [Mar11a, SS11b]. Wulff [Sch13b]. WYD [Han13]. WYD-like [Han13].

Xenon [GLM^{+}15]. XY [BLU16, Cra11, GL16a, LMC19, MS11a, PL13, SML19]. XY-Models [MS11a]. XYZ [HA19].

Y. [Sta11]. Ya. [Szá12]. Yakov [Kha17]. Yang [KS13, MR19]. Yard [BJM15]. Yard-Sale [BJM15]. Yau [Ano19b]. YBE [Yam14]. Years [Zia10, Zuc11b]. Yesterday [Sin10a]. Yor [IS16, Jan15a]. Young [Fun14, MSLT16]. Yttria [MW10]. Yttria-Stabilized [MW10]. Yuhjtman [Pro17]. Yukawa [KID^{+}11, KM19a]. Yule [LPS16].

Zamolodchikov [KS13]. Zealot [PT14a]. Zener [CFK13]. Zernike [CG11a, CG15]. Zero [AP12, ANSW16, BL18, BDP19, BDES19, BP12b, CL10, CG10b, CGP17, DSZ17, GNS18, Gon14, IY14, Ito19b, JRS15, Kem11, KS14, KMO16, LL16, LLM19, MD10, MC19, SLM12, Sta15, Thi16, WW16]. Zero-Range [CG10b, CGP17, Gon14]. Zero-Temperature [ANSW16, GNS18, WW16]. Zero-Variance [Ito19b]. Zeros [BCF10, DMS12, FZ11b, JZ10, LSS19, SS11a]. Zeroth [GS11a]. Zeta [AW18]. ZGB [SM12a]. Zhang [Ano15d, CG17, FSS13, TS12, TA16, XTL14]. Ziff [SM12a, SM12b]. Zipf [Lee18]. Zirconia [MW10]. Zwanzig [WL13].

References

Adams:2010:HER

[AB10] Fred C. Adams and Anthony M. Bloch. Hill’s equation with random forcing parameters: Determination of growth rates through random matrices. Journal of Statistical Physics, 139 (1):139–158, April 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9931-x.

Albano:2014:CTW

[AB14] Ezequiel V. Albano and Kurt Binder. Critical and tricritical wetting in the two-dimensional Blume–Capel model. Journal of Statistical Physics, 157(3):436–455, November 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1091-y.

Angelini:2017:RSM

[AB17] Maria Chiara Angelini and Giulio Biroli. Real space Migdal– Kadanoff renormalisation of glassy systems: Recent results and a critical assessment. Journal of Statistical Physics, 167 (3–4):476–498, May 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Alexander:2018:GTC

[AB18a] Kenneth S. Alexander and Quentin Berger. Geodesics toward corners in first passage percolation. Journal of Statistical Physics, 172(4):1029–1056, August 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Aurzada:2018:PPT

[AB18b] Frank Aurzada and Micha Buck. Persistence probabilities of two-sided (integrated) sums of correlated stationary Gaussian sequences. Journal of Statistical Physics, 170(4):784–799, February 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Akemann:2019:HTC

[AB19] Gernot Akemann and Sung-Soo Byun. The high temperature crossover for general 2D Coulomb gases. Journal of Statistical Physics, 175(6):1043–1065, June 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Angeletti:2014:GLD

[ABA14] Florian Angeletti, Eric Bertin, and Patrice Abry. General limit distributions for sums of random variables with a matrix product representation. Journal of Statistical Physics, 157(6):1255–1283, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1111-y.

Agliari:2010:DSD

[ABC10] Elena Agliari, Raffaella Burioni, and Pierluigi Contucci. A diffusive strategic dynamics for social systems. Journal of Statistical Physics, 139(3):478–491, May 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-9948-1.

Avena:2016:CRW

[ABF16] Luca Avena, Oriane Blondel, and Alessandra Faggionato. A class of random walks in reversible dynamic environments: Antisymmetry and applications to the East model. Journal of Statistical Physics, 165(1):1–23, October 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1596-7.

Argentieri:2015:CPT

[ABFP15] Giuseppe Argentieri, Fabio Benatti, Roberto Floreanini, and Marco Pezzutto. Complete positivity and thermodynamics in a driven open quantum system. Journal of Statistical Physics, 159(5):1127–1153, June 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1210-4.

Aoki:2017:SBC

[ABH^{+}17] Kazuo Aoki, Céline Baranger, Masanari Hattori, Shingo Kosuge, Giorgio Martalò, Julien Mathiaud, and Luc Mieussens. Slip boundary conditions for the compressible Navier–Stokes equations. Journal of Statistical Physics, 169(4):744–781, November 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Asch:2012:DLC

[ABJ12] Joachim Asch, Olivier Bourget, and Alain Joye. Dynamical localization of the Chalker–Coddington model far from transition. Journal of Statistical Physics, 147(1):194–205, April 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0477-y.

Agliari:2017:NNR

[ABLT17] Elena Agliari, Adriano Barra, Chiara Longo, and Daniele Tantari. Neural networks retrieving Boolean patterns in a sea of Gaussian ones. Journal of Statistical Physics, 168(5):1085– 1104, September 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Adams:2016:RPT

[ABMP16] Stefan Adams, Raimundo Briceño, Brian Marcus, and Ronnie Pavlov. Representation and poly-time approximation for pressure of Z^{2} lattice models in the non-uniqueness region. Journal of Statistical Physics, 162(4):1031–1067, February 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1433-4.

Abramson:2013:SSB

[Abr13] Joshua Abramson. Structure of shocks in Burgers turbulence with Lévy noise initial data. Journal of Statistical Physics, 152(3):541–568, August 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0777-x.

Abramov:2017:LOR

[Abr17] Rafail V. Abramov. Leading order response of statistical averages of a dynamical system to small stochastic perturbations. Journal of Statistical Physics, 166(6):1483–1508, March 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-017-1721-2.

Addario-Berry:2018:MTB

[ABR18] Louigi Addario-Berry and Matthew I. Roberts. Mixing time bounds via bottleneck sequences. Journal of Statistical Physics, 173(3–4):845–871, November 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Androulakis:2012:DDS

[ABS12] George Androulakis, Jean Bellissard, and Christian Sadel. Dissipative dynamics in semiconductors at low temperature. Journal of Statistical Physics, 147(2):448–486, April 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0454-5.

Azencott:2010:ASS

[ABT10] R. Azencott, A. Beri, and I. Timofeyev. Adaptive subsampling for parametric estimation of Gaussian diffusions. Journal of Statistical Physics, 139(6):1066–1089, June 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-9975-y.

Azencott:2011:PES

[ABT11] R. Azencott, A. Beri, and I. Timofeyev. Parametric estimation of stationary stochastic processes under indirect observability. Journal of Statistical Physics, 144(1):150–170, July 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0253-4.

Appuhamillage:2014:SDC

[ABT^{+}14] Thilanka A. Appuhamillage, Vrushali A. Bokil, Enrique A. Thomann, Edward C. Waymire, and Brian D. Wood. Skew disperson and continuity of local time. Journal of Statistical Physics, 156(2):384–394, July 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1010-2.

Alves:2014:IPM

[AC14a] José F. Alves and Maria Carvalho. Invariant probability measures and non-wandering sets for impulsive semiflows. Journal of Statistical Physics, 157(6):1097–1113, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1101-0.

Auffinger:2014:FEC

[AC14b] Antonio Auffinger and Wei-Kuo Chen. Free energy and complexity of spherical bipartite models. Journal of Statistical Physics, 157(1):40–59, October 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1073-0.

Andreucci:2019:FFP

[ACCG19] D. Andreucci, E. N. M. Cirillo, M. Colangeli, and D. Gabrielli. Fick and Fokker–Planck diffusion law in inhomogeneous media. Journal of Statistical Physics, 174(2):469–493, January 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Abadi:2015:PWS

[ACG15] Miguel Abadi, Liliam Cardeño, and Sandro Gallo. Potential well spectrum and hitting time in renewal processes. Journal of Statistical Physics, 159(5):1087–1106, June 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1216-y.

Avena:2018:RFN

[ACGM18] Luca Avena, Fabienne Castell, Alexandre Gaudillière, and Clothilde Mélot. Random forests and networks analysis. Journal of Statistical Physics, 173(3–4):985–1027, November 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-018-2124-8.pdf.

Abert:2015:MMB

[ACH15] Miklós Abért, Péter Csikvári, and Tamás Hubai. Matching measure, benjamini–Schramm convergence and the monomer– dimer free energy. Journal of Statistical Physics, 161(1):16–34, October 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1309-7.

Ayyer:2011:EPD

[ACL^{+}11] A. Ayyer, E. A. Carlen, J. L. Lebowitz, P. K. Mohanty, D. Mukamel, and E. R. Speer. Erratum to: Phase Diagram of the ABC Model on an Interval. Journal of Statistical Physics, 144(4):920–921, August 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-011-0287-7.pdf.

Antenucci:2014:CSH

[ACL14] F. Antenucci, A. Crisanti, and L. Leuzzi. Critical study of hierarchical lattice renormalization group in magnetic ordered and quenched disordered systems: Ising and Blume–Emery– Griffiths models. Journal of Statistical Physics, 155(5):909– 931, June 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-0977-z.

Alvarado:2011:GRE

[ACM11] Sotero Alvarado, Jorge Alberto Calvo, and Kenneth C. Millett. The generation of random equilateral polygons. Journal of Statistical Physics, 143(1):102–138, April 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0164-4.

Alberici:2015:MFM

[ACM15] Diego Alberici, Pierluigi Contucci, and Emanuele Mingione. A mean–field monomer–dimer model with randomness: Exact solution and rigorous results. Journal of Statistical Physics, 160(6):1721–1732, September 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1306-x.

Ayala:2018:QBG

[ACR18] Mario Ayala, Gioia Carinci, and Frank Redig. Quantitative Boltzmann–Gibbs principles via orthogonal polynomial duality. Journal of Statistical Physics, 171(6):980–999, June 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-018-2060-7.pdf.

Arguin:2011:SRS

[AD11] Louis-Pierre Arguin and Michael Damron. Short-range spin glasses and random overlap structures. Journal of Statistical Physics, 143(2):226–250, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0177-z.

Allez:2014:TWH

[AD14] Romain Allez and Laure Dumaz. Tracy–Widom at high temperature. Journal of Statistical Physics, 156(6):1146–1183, September 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1058-z.

Allez:2015:RMN

[AD15a] Romain Allez and Laure Dumaz. Random matrices in nonconfining potentials. Journal of Statistical Physics, 160(3): 681–714, August 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1258-1.

Alm:2015:FPP

[AD15b] Sven Erick Alm and Maria Deijfen. First passage percolation on Z^{2}: A simulation study. Journal of Statistical Physics, 161(3):657–678, November 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1356-0.

Alberts:2010:BDR

[ADC10] Tom Alberts and Hugo Duminil-Copin. Bridge decomposition of restriction measures. Journal of Statistical Physics, 140 (3):467–493, August 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9999-3.

Agazzi:2018:GCR

[ADE18] A. Agazzi, A. Dembo, and J.-P. Eckmann. On the geometry of chemical reaction networks: Lyapunov function and large deviations. Journal of Statistical Physics, 172(2):321–352, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Antown:2018:OLR

[ADF18] Fadi Antown, Davor Dragicević, and Gary Froyland. Optimal linear responses for Markov chains and stochastically perturbed dynamical systems. Journal of Statistical Physics, 170 (6):1051–1087, March 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Aurzada:2017:RWB

[ADGPP17] Frank Aurzada, Alexis Devulder, Nadine Guillotin-Plantard, and Françoise Pène. Random walks and branching processes in correlated Gaussian environment. Journal of Statistical Physics, 166(1):1–23, January 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1677-7.

Arsuaga:2012:EAR

[ADH12] J. Arsuaga, Y. Diao, and K. Hinson. The effect of angle restriction on the topological characteristics of minicircle networks. Journal of Statistical Physics, 146(2):434–445, January 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0386-5.

Alexander:2010:LIM

[ADMS10] Kenneth S. Alexander, François Dunlop, and Salvador Miracle-Solé. Layering in the Ising model. Journal of Statistical Physics, 141(2):217–241, October 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-010-0042-
5.

Alexander:2011:LWT

[ADMS11] Kenneth S. Alexander, François Dunlop, and Salvador Miracle-Solé. Layering and wetting transitions for an SOS interface. Journal of Statistical Physics, 142(3):524–576, February 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0127-9.

Attal:2014:EBQ

[ADP14] S. Attal, J. Deschamps, and C. Pellegrini. Entanglement of bipartite quantum systems driven by repeated interactions. Journal of Statistical Physics, 154(3):819–837, February 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-0917-y.

Accardi:2018:QQ

[ADR18] Luigi Accardi, Ameur Dhahri, and Habib Rebei. C^{∗}-quadratic quantization. Journal of Statistical Physics, 172(5):1187–1209, September 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Apollonio:2019:CMD

[ADS^{+}19] Valentina Apollonio, Roberto D’Autilia, Benedetto Scoppola, Elisabetta Scoppola, and Alessio Troiani. Criticality of measures on 2-d Ising configurations: From square to hexagonal graphs. Journal of Statistical Physics, 177(5):1009–1021, December 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Albeverio:2017:ECB

[ADU17] Sergio Albeverio, Francesco C. De Vecchi, and Stefania Ugolini. Entropy chaos and Bose–Einstein condensation. Journal of Statistical Physics, 168(3):483–507, August 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Adams:2016:PTD

[AE16] Stefan Adams and Michael Eyers. Phase transitions in Delaunay Potts models. Journal of Statistical Physics, 162(1):162– 185, January 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1393-8.

Agazzi:2014:CHB

[AEG14] A. Agazzi, J.-P. Eckmann, and G. M. Graf. The colored Hofstadter butterfly for the honeycomb lattice. Journal of Statistical Physics, 156(3):417–426, August 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0992-0.

Ajanki:2016:LSS

[AEK16] Oskari H. Ajanki, László Erdős, and Torben Krüger. Local spectral statistics of Gaussian matrices with correlated entries. Journal of Statistical Physics, 163(2):280–302, April 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-016-1479-y.pdf.

Aharony:2019:TNG

[AEWCD19] A. Aharony, O. Entin-Wohlman, D. Chowdhury, and S. Dattagupta. Is telegraph noise a good model for the environment of mesoscopic systems? Journal of Statistical Physics, 175(3–4):704–724, May 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Aharony:2014:RCI

[AEWI14] A. Aharony, O. Entin-Wohlman, and Y. Imry. Renormalization of competing interactions and superconductivity on small scales. Journal of Statistical Physics, 157(4–5):979–989, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1100-1.

Araujo:2012:GSF

[AF12] M. T. Araujo and E. Drigo Filho. A general solution of the Fokker–Planck equation. Journal of Statistical Physics, 146(3):610–619, February 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0411-8.

Amini:2014:BPP

[AF14a] Hamed Amini and Nikolaos Fountoulakis. Bootstrap percolation in power-law random graphs. Journal of Statistical Physics, 155(1):72–92, April 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-0946-6.

Ammari:2014:WMA

[AF14b] Zied Ammari and Marco Falconi. Wigner measures approach to the classical limit of the Nelson model: Convergence of dynamics and ground state energy. Journal of Statistical Physics, 157(2):330–362, October 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1079-7.

Araujo:2014:FPE

[AF14c] M. T. Araujo and E. Drigo Filho. Fokker–Planck equation for a metastable time dependent potential. Journal of Statistical Physics, 156(4):766–774, August 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1031-x.

Alastuey:2016:FMS

[AF16] Angel Alastuey and Riccardo Fantoni. Fourth moment sum rule for the charge correlations of a two-component classical plasma. Journal of Statistical Physics, 163(4):887–913, May 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1512-1.

Afanasiev:2016:CFC

[Afa16] Ie. Afanasiev. On the correlation functions of the characteristic polynomials of the sparse Hermitian random matrices. Journal of Statistical Physics, 163(2):324–356, April 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1486-z.

Afanasiev:2019:CFC

[Afa19] Ie. Afanasiev. On the correlation functions of the characteristic polynomials of non–Hermitian random matrices with independent entries. Journal of Statistical Physics, 176(6):1561– 1582, September 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Anjos:2016:ASF

[AFCA16] R. C. Anjos, G. B. Freitas, and C. H. Coimbra-Araújo. Analytical solutions of the Fokker–Planck equation for generalized Morse and Hulthén potentials. Journal of Statistical Physics, 162(2):387–396, January 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1414-7.

Azevedo:2017:EVL

[AFFR17] Davide Azevedo, Ana Cristina Moreira Freitas, Jorge Milhazes Freitas, and Fagner B. Rodrigues. Extreme value laws for dynamical systems with countable extremal sets. Journal of Statistical Physics, 167(5):1244–1261, June 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Avron:2012:ARL

[AFG12] J. E. Avron, M. Fraas, and G. M. Graf. Adiabatic response for Lindblad dynamics. Journal of Statistical Physics, 148(5): 800–823, September 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0550-6.

Armendariz:2015:FCG

[AFGL15] Inés Armendáriz, Pablo A. Ferrari, Pablo Groisman, and Florencia Leonardi. Finite cycle Gibbs measures on permutations of Z^{d}. Journal of Statistical Physics, 158(6):1213–1233, March 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1169-6.

Aoyama:2010:ECS

[AFI^{+}10] Hideaki Aoyama, Y. Fujiwara, Y. Ikeda, H. Iyetomi, and W. Souma. Econophysics and companies: statistical life and death in complex business networks. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-19149-1. xxvi + 234 pp. LCCN HB131 .E268

2010.  URL http://assets.cambridge.org/97805211/
91494/cover/9780521191494.jpg;   http://www.loc.gov/
catdir/enhancements/fy1009/2010018296-b.html;   http:
//www.loc.gov/catdir/enhancements/fy1009/2010018296-
d.html;   http://www.loc.gov/catdir/enhancements/fy1009/
2010018296-t.html.

Ayyer:2019:PDM

[AFR19] Arvind Ayyer, Caley Finn, and Dipankar Roy. The phase diagram for a multispecies left-permeable asymmetric exclusion process. Journal of Statistical Physics, 174(3):605–621, February 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Altshuler:2013:SAR

[AFS^{+}13] Yaniv Altshuler, Michael Fire, Erez Shmueli, Yuval Elovici, Alfred Bruckstein, Alex (Sandy) Pentland, and David Lazer. The social amplifier — reaction of human communities to emergencies. Journal of Statistical Physics, 152(3):399–418, August 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-013-0759-z.pdf.

Afzali:2012:ABE

[Afz12] R. Afzali. Application of Boltzmann equation on spin diffusion of ferromagnetic superfluid ^{3}He: Near critical temperature. Journal of Statistical Physics, 147(4):808–820, June 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0489-7.

Alonso:2011:RCS

[AG11] Ricardo J. Alonso and Irene M. Gamba. A revision on classical solutions to the Cauchy Boltzmann problem for soft potentials. Journal of Statistical Physics, 143(4):740–746, May 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0205-z.

Abdel-Gawad:2012:TUM

[AG12a] H. I. Abdel-Gawad. Towards a unified method for exact solutions of evolution equations. An application to reaction diffusion equations with finite memory transport. Journal of Statistical Physics, 147(3):506–518, May 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0467-0.

Apte:2012:NDA

[AG12b] Pankaj A. Apte and Arvind K. Gautam. Nonmonotonic dependence of the absolute entropy on temperature in supercooled Stillinger–Weber silicon. Journal of Statistical Physics, 149(3):551–567, November 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0594-7.

Affleck:2014:SCM

[AG14] Ian Affleck and Domenico Giuliano. Screening clouds and Majorana fermions. Journal of Statistical Physics, 157(4–5): 666–691, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1056-1.

Allahverdyan:2015:AIB

[AG15a] Armen E. Allahverdyan and Aram Galstyan. Active inference for binary symmetric hidden Markov models. Journal of Statistical Physics, 161(2):452–466, October 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1321-y.

Azencott:2015:AML

[AG15b] Robert Azencott and Yutheeka Gadhyan. Accuracy of maximum likelihood parameter estimators for Heston stochastic volatility SDE. Journal of Statistical Physics, 159(2):393– 420, April 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1120-x.

Agoritsas:2016:DIF

[AGGL^{+}16] Elisabeth Agoritsas, Reinaldo García-García, Vivien Lecomte, Lev Truskinovsky, and Damien Vandembroucq. Driven interfaces: From flow to creep through model reduction. Journal of Statistical Physics, 164(6):1394–1428, September 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1588-7.

Angeli:2019:RES

[AGJP19] Letizia Angeli, Stefan Grosskinsky, Adam M. Johansen, and Andrea Pizzoferrato. Rare event simulation for stochastic dynamics in continuous time. Journal of Statistical Physics, 176(5):1185–1210, September 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-019-02340-
1.pdf.

Aurell:2012:RSL

[AGMM^{+}12] Erik Aurell, Krzysztof Gawȩdzki, Carlos Mejía-Monasterio, Roya Mohayaee, and Paolo Muratore-Ginanneschi. Refined second law of thermodynamics for fast random processes. Journal of Statistical Physics, 147(3):487–505, May 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0478-x.

Azencott:2018:LDG

[AGO18] Robert Azencott, Brett Geiger, and William Ott. Large deviations for Gaussian diffusions with delay. Journal of Statistical Physics, 170(2):254–285, January 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Abreu:2019:HAP

[AGR19] Luís Daniel Abreu, Karlheinz Gröchenig, and José Luis Romero. Harmonic analysis in phase space and finite Weyl– Heisenberg ensembles. Journal of Statistical Physics, 174(5): 1104–1136, March 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-019-02226-
2.pdf.

Alves:2017:NTL

[AHDV17] C. Alves, M. R. Hilário, B. N. B. De Lima, and D. Valesin. A note on truncated long-range percolation with heavy tails on oriented graphs. Journal of Statistical Physics, 169(5): 972–980, December 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Akama:2012:RFM

[AI12] Yohji Akama and Shinji Iizuka. Random fields on model sets with localized dependency and their diffraction. Journal of Statistical Physics, 149(3):478–495, November 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0588-5.

Auffinger:2019:SDG

[AJ19] Antonio Auffinger and Aukosh Jagannath. On spin distributions for generic p-spin models. Journal of Statistical Physics, 174(2):316–332, January 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Aghamohammadi:2014:LMS

[AK14a] Amir Aghamohammadi and Mohammad Khorrami. Lattice models solvable through the full interval method on links. Journal of Statistical Physics, 157(6):1320–1330, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1097-5.

Arguin:2014:MAR

[AK14b] Louis-Pierre Arguin and Nicola Kistler. Microcanonical analysis of the random energy model in a random magnetic field. Journal of Statistical Physics, 157(1):1–16, October 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1072-1.

Atwal:2016:LQS

[AK16] Gurinder S. Atwal and Justin B. Kinney. Learning quantitative sequence–function relationships from massively parallel experiments. Journal of Statistical Physics, 162(5):1203– 1243, March 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-015-1398-3.pdf.

Albert:2017:OCS

[AK17] Leif Albert and Michael K.-H. Kiessling. Order and chaos in some deterministic infinite trigonometric products. Journal of Statistical Physics, 168(3):595–619, August 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Adriaans:2018:WDS

[AK18a] Erwin Adriaans and Júlia Komjáthy. Weighted distances in scale-free configuration models. Journal of Statistical Physics, 173(3–4):1082–1109, November 2018. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-018-1957-5.pdf.

Avron:2018:TS

[AK18b] J. E. Avron and O. Kenneth. Teleportation for septuagenarians. Journal of Statistical Physics, 172(2):555–561, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Aurzada:2019:PEP

[AK19] Frank Aurzada and Marvin Kettner. Persistence exponents via perturbation theory: AR(1)-processes. Journal of Statistical Physics, 177(4):651–665, November 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Agarwal:2019:SCB

[AKD19] Sanaa Agarwal, Manas Kulkarni, and Abhishek Dhar. Some connections between the classical Calogero–Moser model and the log-gas. Journal of Statistical Physics, 176(6):1463–1479, September 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Amitai:2013:KDP

[AKH13] A. Amitai, I. Kupka, and D. Holcman. Kinetics of diffusing polymer encounter in confined cellular microdomains. Journal of Statistical Physics, 153(6):1107–1131, December 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0876-8.

Avrachenkov:2018:MFA

[AKL18] Konstantin Avrachenkov, Arun Kadavankandy, and Nelly Litvak. Mean field analysis of personalized PageRank with implications for local graph clustering. Journal of Statistical Physics, 173(3–4):895–916, November 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Alberts:2013:SPR

[AKM13] Tom Alberts, Michael J. Kozdron, and Robert Masson. Some partial results on the convergence of loop-erased random walk to SLE(2) in the natural parametrization. Journal of Statistical Physics, 153(1):119–141, October 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0816-7.

Alberts:2014:CDR

[AKQ14] Tom Alberts, Konstantin Khanin, and Jeremy Quastel. The continuum directed random polymer. Journal of Statistical Physics, 154(1–2):305–326, January 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0872-z.

Aldous:2018:FIO

[AL18] David Aldous and Xiang Li. A framework for imperfectly observed networks. Journal of Statistical Physics, 173(3–4): 1303–1320, November 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Aquino:2018:PMO

[ALAF18] Édio O. Aquino, F. W. S. Lima, Ascânio D. Araújo, and Raimundo N. Costa Filho. Potts model in one-dimension on directed small-world networks. Journal of Statistical Physics, 171(6):1112–1121, June 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Alberici:2016:CEA

[Alb16] Diego Alberici. A cluster expansion approach to the Heilmann–Lieb liquid crystal model. Journal of Statistical Physics, 162(3):761–791, February 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-015-1421-
8.

Aletti:2018:FPF

[ALM18] Giacomo Aletti, Nikolai Leonenko, and Ely Merzbach. Fractional Poisson fields and martingales. Journal of Statistical Physics, 170(4):700–730, February 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Avdeeva:2014:NEK

[ALS14] Maria Avdeeva, Dong Li, and Yakov G. Sinai. New Erdős– Kac type theorems for signed measures on square-free integers. Journal of Statistical Physics, 154(1–2):327–333, January 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0885-7.

Ahmad:2018:PAG

[ALS18] Mohd Ali Khameini Ahmad, Lingmin Liao, and Mansoor Saburov. Periodic p-adic Gibbs measures of q-state Potts model on Cayley trees I: The chaos implies the vastness of the set of p-adic Gibbs measures. Journal of Statistical Physics, 171(6):1000–1034, June 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Atapour:2010:NEC

[AM10] Mahshid Atapour and Neal Madras. On the number of entangled clusters. Journal of Statistical Physics, 139(1):1–26, April 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-9941-8.

Antonov:2012:IRB

[AM12] N. V. Antonov and A. V. Malyshev. Inertial-range behavior of a passive scalar field in a random shear flow: Renormalization group analysis of a simple model. Journal of Statistical Physics, 146(1):33–55, January 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0399-0.

Akimoto:2014:PDS

[AM14] Takuma Akimoto and Tomoshige Miyaguchi. Phase diagram in stored–energy-driven Lévy flight. Journal of Statistical Physics, 157(3):515–530, November 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-014-1084-x.pdf.

Alberici:2018:TPM

[AM18] Diego Alberici and Emanuele Mingione. Two populations mean-field monomer–dimer model. Journal of Statistical Physics, 171(1):96–105, April 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Abbott:2019:SLD

[AM19] Michael C. Abbott and Benjamin B. Machta. A scaling law from discrete to continuous solutions of channel capacity problems in the low-noise limit. Journal of Statistical Physics, 176(1):214–227, July 2019. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-019-02296-2.pdf.

Amini:2010:BPL

[Ami10] Hamed Amini. Bootstrap percolation in living neural networks. Journal of Statistical Physics, 141(3):459–475, November 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-0056-z.

Accardi:2014:QMC

[AMS14] Luigi Accardi, Farrukh Mukhamedov, and Mansoor Saburov. On quantum Markov chains on Cayley tree III: Ising model. Journal of Statistical Physics, 157(2):303–329, October 2014.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1083-y.

Amati:2019:MEF

[AMS19] Graziano Amati, Hugues Meyer, and Tanja Schilling. Memory effects in the Fermi–Pasta–Ulam model. Journal of Statistical Physics, 174(1):219–257, January 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Agliari:2018:NCM

[AMT18] Elena Agliari, Danila Migliozzi, and Daniele Tantari. Nonconvex multi-species Hopfield models. Journal of Statistical Physics, 172(5):1247–1269, September 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Alekseev:2017:TSL

[AMW17] Oleg Alekseev and Mark Mineev-Weinstein. Theory of stochastic Laplacian growth. Journal of Statistical Physics, 168(1):68–91, July 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Arkeryd:2015:BCI

[AN15] Leif Arkeryd and Anne Nouri. Bose condensates in interaction with excitations: A two-component space-dependent model close to equilibrium. Journal of Statistical Physics, 160(1):209–238, July 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1229-6.

Azadi:2019:SMS

[AN19] Amir Azadi and David R. Nelson. Statistical mechanics of specular reflections from fluctuating membranes and interfaces. Journal of Statistical Physics, 175(3–4):578–597, May 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Andre:2019:RMI

[And19] Morgan André. A result of metastability for an infinite system of spiking neurons. Journal of Statistical Physics, 177(5):984– 1008, December 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Anonymous:2010:P

[Ano10] Anonymous. Preface. Journal of Statistical Physics, 138(1–3): 1, February 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-010-9922-y.pdf.

Anonymous:2011:Pa

[Ano11a] Anonymous. Preface. Journal of Statistical Physics, 142(6): 1123, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-011-0188-9.pdf.

Anonymous:2011:Pb

[Ano11b] Anonymous. Preface. Journal of Statistical Physics, 144 (2):217–218, July 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0270-3.

Anonymous:2011:Pc

[Ano11c] Anonymous. Preface. Journal of Statistical Physics, 145 (3):507, November 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-011-0380-y.
pdf.

Anonymous:2011:PSIa

[Ano11d] Anonymous. Preface to special issue on statistical mechanics and computational physics. Journal of Statistical Physics, 144(3):443–444, August 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0303-y.

Anonymous:2011:PSIb

[Ano11e] Anonymous. Preface to special issue on statistical mechanics and computational physics. Journal of Statistical Physics, 145(4):785–786, November 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0395-4.

Anonymous:2012:SMCa

[Ano12a] Anonymous. 106th Statistical Mechanics Conference Rutgers University, Busch Campus, Hill Center, Room 114, Sunday, Tuesday, December 18-20, 2011. Journal of Statistical Physics, 147(1):224–230, April 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0440-y.

Anonymous:2012:SMCb

[Ano12b] Anonymous. 107th Statistical Mechanics Conference Rutgers University, Busch Campus, Hill Center, Room 114, Sunday– Tuesday, May 6–8, 2012. Journal of Statistical Physics, 148 (2):387–391, August 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0515-9.

Anonymous:2014:F

[Ano14a] Anonymous. Foreword. Journal of Statistical Physics, 157 (4–5):609, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-014-1117-5.
pdf.

Anonymous:2014:OLI

[Ano14b] Anonymous. Oscar E. Lanford III (Jan 6, 1940–Nov 16, 2013). Journal of Statistical Physics, 155(3):419–420, May 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-014-0931-0.pdf.

Anonymous:2014:P

[Ano14c] Anonymous. Preface. Journal of Statistical Physics, 154(1–2): 1, January 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-013-0906-6.pdf.

Anonymous:2015:SMC

[Ano15a] Anonymous. 110th Statistical Mechanics Conference. Journal of Statistical Physics, 161(2):518–520, October 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1305-y.

Anonymous:2015:BJM

[Ano15b] Anonymous. Bernard Jancovici (1930–2013): In memoriam. Journal of Statistical Physics, 160(1):1–3, July 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1213-1.

Anonymous:2015:P

[Ano15c] Anonymous. Preface. Journal of Statistical Physics, 158 (3):513, February 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-015-1191-3.
pdf.

Anonymous:2015:SIK

[Ano15d] Anonymous. Special issue on Kardar–Parisi–Zhang growth dynamics. Journal of Statistical Physics, 160(4):793, August 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-015-1317-7.pdf.

Anonymous:2016:SMC

[Ano16a] Anonymous. 111th–113th Statistical Mechanics Conference publication lists. Journal of Statistical Physics, 162(1):267– 275, January 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1397-4.

Anonymous:2016:HRM

[Ano16b] Anonymous. Howard Reiss: In memoriam. Journal of Statistical Physics, 164(5):1027–1028, September 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1586-9.

Anonymous:2017:Pa

[Ano17a] Anonymous. Preface. Journal of Statistical Physics, 166 (3–4):457, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-017-1729-7.
pdf.

Anonymous:2017:Pb

[Ano17b] Anonymous. Preface. Journal of Statistical Physics, 167(3– 4):417–419, May 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Anonymous:2018:SMC

[Ano18a] Anonymous. 119th Statistical Mechanics Conference Rutgers University, Busch Campus, Hill Center Room 114, Sunday– Tuesday, May 6–8, 2018 Guests of Honor: Michael Aizenman, Paul Chaikin, and Jennifer Chayes. Journal of Statistical Physics, 172(4):1181–1184, August 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Anonymous:2018:E

[Ano18b] Anonymous. Editorial. Journal of Statistical Physics, 172 (5):1186, September 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-018-2148-0.
pdf.

Anonymous:2018:LJL

[Ano18c] Anonymous. Letter from Joel Lebowitz. Journal of Statistical Physics, 172(5):1185, September 2018. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-018-2153-3.pdf.

Anonymous:2018:PSI

[Ano18d] Anonymous. Preface for the special issue dedicated to J. Fröhlich, T. Spencer and H. Spohn. Journal of Statistical Physics, 172(2):313, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-018-2101-2.
pdf.

Anonymous:2019:SMCa

[Ano19a] Anonymous. 120th Statistical Mechanics Conference Rutgers University, Busch Campus, Hill Center Room 114, Sunday– Tuesday, December 16-18, 2018 Guests of Honor: Daniel Fisher, David Huse, and Mehran Kardar. Journal of Statistical Physics, 174(5):953–957, March 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Anonymous:2019:SMCb

[Ano19b] Anonymous. 121st Statistical Mechanics Conference Rutgers University, Busch Campus, Hill Center Room 114, Sunday, May 12, 2019–Tuesday May 14, 2019 Guests of Honor: Curtis Callan, Nigel Goldenfeld, and Horng-Tzer Yau. Journal of Statistical Physics, 176(3):505–509, August 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Abraham:2018:CPS

[ANS18] Douglas Abraham, Charles M. Newman, and Senya Shlosman. A continuum of pure states in the Ising model on a halfplane. Journal of Statistical Physics, 172(2):611–626, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Arguin:2014:FBI

[ANSW14] L.-P. Arguin, C. M. Newman, D. L. Stein, and J. Wehr. Fluctuation bounds for interface free energies in spin glasses. Journal of Statistical Physics, 156(2):221–238, July 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1009-8.

Arguin:2016:ZTF

[ANSW16] L.-P. Arguin, C. M. Newman, D. L. Stein, and J. Wehr. Zerotemperature fluctuations in short-range spin glasses. Journal of Statistical Physics, 163(5):1069–1078, June 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1516-x.

Arista:2019:LEW

[AO19] Jonas Arista and Neil O’Connell. Loop-erased walks and random matrices. Journal of Statistical Physics, 177(3):528–567, November 2019. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-019-02378-1.pdf.

Alastuey:2010:ASS

[AP10] A. Alastuey and J. Piasecki. Approach to a stationary state in an external field. Journal of Statistical Physics, 139(6):991– 1012, June 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-9976-x.

Ancliff:2011:SCS

[AP11] Mark Ancliff and Jeong-Man Park. Spin coherent state representation of the Crow–Kimura and eigen models of quasispecies theory. Journal of Statistical Physics, 143(4):636– 656, May 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0207-x.

Alexakis:2012:CEZ

[AP12] A. Alexakis and F. Pétrélis. Critical exponents in zero dimensions. Journal of Statistical Physics, 149(4):738–753, November 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0615-6.

Akemann:2014:IAK

[AP14a] G. Akemann and M. J. Phillips. The interpolating Airy kernels for the β = 1 and β = 4 elliptic Ginibre ensembles. Journal of Statistical Physics, 155(3):421–465, May 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-0962-6.

Alves:2014:WTM

[AP14b] Rogério Gomes Alves and Aldo Procacci. Witness trees in the Moser–Tardos algorithmic Lovász local lemma and Penrose trees in the hard-core lattice gas. Journal of Statistical Physics, 156(5):877–895, September 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1054-3.

Austin:2018:GML

[AP18] Tim Austin and Moumanti Podder. Gibbs measures over locally tree-like graphs and percolative entropy over infinite regular trees. Journal of Statistical Physics, 170(5):932–951, March 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Adorisio:2018:EES

[APdM^{+}18] Matteo Adorisio, Alberto Pezzotta, Clélia de Mulatier, Cristian Micheletti, and Antonio Celani. Exact and efficient sampling of conditioned walks. Journal of Statistical Physics, 170 (1):79–100, January 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Akopyan:2015:IMG

[APR15] Arseniy V. Akopyan, Sergey A. Pirogov, and Aleksandr N. Rybko. Invariant measures of genetic recombination processes. Journal of Statistical Physics, 160(1):163–167, July 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1238-5.

Agliari:2017:PTM

[APRT17] Elena Agliari, Angelica Pachon, Pablo M. Rodriguez, and Flavia Tavani. Phase transition for the Maki–Thompson rumour model on a small-world network. Journal of Statistical Physics, 169(4):846–875, November 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Alves:2012:PIG

[APS12] Rogério G. Alves, Aldo Procacci, and Remy Sanchis. Percolation on infinite graphs and isoperimetric inequalities. Journal of Statistical Physics, 149(5):831–845, November 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0644-1.

Attal:2012:OQR

[APSS12] S. Attal, F. Petruccione, C. Sabot, and I. Sinayskiy. Open quantum random walks. Journal of Statistical Physics, 147 (4):832–852, June 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0491-0.

Albi:2019:BGH

[APZ19] Giacomo Albi, Lorenzo Pareschi, and Mattia Zanella. Boltzmann games in heterogeneous consensus dynamics. Journal of Statistical Physics, 175(1):97–125, April 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Aristoff:2011:DTG

[AR11a] David Aristoff and Charles Radin. Dilatancy transition in a granular model. Journal of Statistical Physics, 143(2):215– 225, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0180-4.

Aristoff:2011:RS

[AR11b] David Aristoff and Charles Radin. Rigidity in solids. Journal of Statistical Physics, 144(6):1247–1255, September 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0304-x.

Arafa:2011:SST

[Ara11a] A. A. M. Arafa. Series solutions of time-fractional host– parasitoid systems. Journal of Statistical Physics, 145(5): 1357–1367, December 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0339-z.

Aranson:2011:RDR

[Ara11b] Igor Aranson. Rashmi Desai and Raymond Kapral: Dynamics of Self-Organized and Self-Assembled Structures. Journal of Statistical Physics, 142(1):220–222, January 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0105-7.

Ariel:2015:ODP

[ARBJ15] Gil Ariel, Oren Rimer, and Eshel Ben-Jacob. Order-disorder phase transition in heterogeneous populations of self-propelled particles. Journal of Statistical Physics, 158(3):579–588, February 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1095-7.

Arinshteyn:2011:VPT

[Ari11] Eduard A. Arinshteyn. Variational principle in the theory of partial distribution functions. Journal of Statistical Physics, 144(4):831–845, August 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0275-y.

Arkin:2010:CPT

[Ark10] Handan Arkin. Comparison of the parallel tempering algorithm and multicanonical method as applied to coarsegrained off-lattice models for folding heteropolymers. Journal of Statistical Physics, 139(2):326–332, April 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-9942-7.

Arkeryd:2013:LTK

[Ark13] Leif Arkeryd. On low temperature kinetic theory; spin diffusion, Bose Einstein condensates, anyons. Journal of Statistical Physics, 150(6):1063–1079, March 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0695-y.

Albeverio:2017:EP

[ARS17a] S. Albeverio, B. Rüdiger, and P. Sundar. The Enskog process. Journal of Statistical Physics, 167(1):90–122, April 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-017-1743-9.

Alves:2017:LCS

[ARS17b] Caio Alves, Rodrigo Ribeiro, and Rémy Sanchis. Large communities in a scale-free network. Journal of Statistical Physics, 166(1):137–149, January 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1676-8.

Akin:2011:NSL

[ART11] H. Akin, U. A. Rozikov, and S. Temir. A new set of limiting Gibbs measures for the Ising model on a Cayley tree. Journal of Statistical Physics, 142(2):314–321, January 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-0106-6.

Azencott:2015:PEA

[ART15] Robert Azencott, Peng Ren, and Ilya Timofeyev. Parametric estimation from approximate data: Non–Gaussian diffusions. Journal of Statistical Physics, 161(5):1276–1298, December 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1379-6.

Artemov:2019:CAS

[Art19a] A. N. Artemov. Cumulant analysis of the statistical properties of a deterministically thermostated harmonic oscillator. Journal of Statistical Physics, 174(5):992–1010, March 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Artigue:2019:BTG

[Art19b] Alfonso Artigue. Billiards and toy gravitons. Journal of Statistical Physics, 175(2):213–232, April 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ardebili:2016:GSP

[AS16] S. Bahareh Seyedein Ardebili and Reza Sepehrinia. Ground state properties of Ising chain with random monomer-dimer couplings. Journal of Statistical Physics, 163(3):568–575, May 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1499-7.

Asad:2013:EER

[Asa13] Jihad H. Asad. Exact evaluation of the resistance in an infinite face-centered cubic network. Journal of Statistical Physics, 150(6):1177–1182, March 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0716-x.

Akimoto:2015:DBT

[ASA15] Takuma Akimoto, Soya Shinkai, and Yoji Aizawa. Distributional behavior of time averages of non-L^{1} observables in onedimensional intermittent maps with infinite invariant measures. Journal of Statistical Physics, 158(2):476–493, January 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1138-0.

Astrauskas:2012:ETS

[Ast12] A. Astrauskas. Extremal theory for spectrum of random discrete Schrödinger operator. II. Distributions with heavy tails. Journal of Statistical Physics, 146(1):98–117, January 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0402-9.

Astrauskas:2013:ETS

[Ast13] A. Astrauskas. Extremal theory for spectrum of random discrete Schrödinger operator. III. Localization properties. Journal of Statistical Physics, 150(5):889–907, March 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0669-5.

Avena:2012:CAF

[AT12] L. Avena and P. Thomann. Continuity and anomalous fluctuations in random walks in dynamic random environments: Numerics, phase diagrams and conjectures. Journal of Statistical Physics, 147(6):1041–1067, July 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0502-1.

Andreis:2018:CSL

[AT18] Luisa Andreis and Daniele Tovazzi. Coexistence of stable limit cycles in a generalized Curie–Weiss model with dissipation. Journal of Statistical Physics, 173(1):163–181, October 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Albeverio:2015:DTG

[AU15] Sergio Albeverio and Stefania Ugolini. A Doob h-transform of the Gross–Pitaevskii Hamiltonian. Journal of Statistical Physics, 161(2):486–508, October 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-015-1337-
3.

Aumann:2015:SBR

[Aum15] Simon Aumann. Spontaneous breaking of rotational symmetry with arbitrary defects and a rigidity estimate. Journal of Statistical Physics, 160(1):168–208, July 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1234-9.

Aurell:2018:GEE

[Aur18] Erik Aurell. Global estimates of errors in quantum computation by the Feynman–Vernon formalism. Journal of Statistical Physics, 171(5):745–767, June 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-018-2037-6.
pdf.

Argun:2016:BSM

[AV16] Aykut Argun and Giovanni Volpe. Better stability with measurement errors. Journal of Statistical Physics, 163(6):1477– 1485, June 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1518-8.

Abdelmalik:2016:MCA

[AvB16] M. R. A. Abdelmalik and E. H. van Brummelen. Moment closure approximations of the Boltzmann equation based on φ-divergences. Journal of Statistical Physics, 164(1):77–104, July 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-016-1529-5.pdf.

Aizenman:2017:PCF

[AVW17] Michael Aizenman, Manuel Laínz Valcázar, and Simone Warzel. Pfaffian correlation functions of planar dimer covers. Journal of Statistical Physics, 166(3–4):1078–1091, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1684-8.

Aizenman:2018:KWF

[AW18] Michael Aizenman and Simone Warzel. Kac–Ward formula and its extension to order–disorder correlators through a graph zeta function. Journal of Statistical Physics, 173(6): 1755–1778, December 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Aquino:2016:KSR

[AWE16] Gerardo Aquino, Ned S. Wingreen, and Robert G. Endres. Know the single-receptor sensing limit? Think again. Journal of Statistical Physics, 162(5):1353–1364, March 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-015-1412-9.pdf.

Altmann:2013:ITW

[AWM13] Eduardo G. Altmann, Zakary L. Whichard, and Adilson E. Motter. Identifying trends in word frequency dynamics. Journal of Statistical Physics, 151(1–2):277–288, April 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0699-7.

Arita:2010:EQP

[AY10] Chikashi Arita and Daichi Yanagisawa. Exclusive queueing process with discrete time. Journal of Statistical Physics, 141(5):829–847, December 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0075-9.

Aspelmeier:2011:IDS

[AZ11] T. Aspelmeier and A. Zippelius. The integrated density of states of the random graph Laplacian. Journal of Statistical Physics, 144(4):759–773, August 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-011-0271-2.pdf.

Azad:2011:ADP

[Aza11] Elahe Zohoorian Azad. Analysis of a drop–push model for percolation and coagulation. Journal of Statistical Physics, 143 (6):1055–1073, June 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0236-5.

ben-Avraham:2010:HGJ

[bA10] Daniel ben Avraham. H. Gould and J. Tobochnik: Statistical and Thermal Physics. Journal of Statistical Physics, 140(5):1022–1023, September 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0035-4.

ben-Avraham:2012:PBM

[bA12] Daniel ben Avraham. L. Peliti: Statistical Mechanics in a Nutshell. Journal of Statistical Physics, 146(4):878–879, February 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0420-2.

Birch:2014:SLD

[BA14] Annika Birch and Tomaso Aste. Systemic losses due to counterparty risk in a stylized banking system. Journal of Statistical Physics, 156(5):998–1024, September 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-014-1040-9.pdf.

Bordbar:2019:TII

[BA19] Pejman Bordbar and Sodeif Ahadpour. Type-II intermittency from Markov binary block visibility graph perspective. Journal of Statistical Physics, 177(3):399–414, November 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Baber:2012:SCC

[Bab12] John Baber. Scaled correlations of critical points of random sections on Riemann surfaces. Journal of Statistical Physics, 148(2):250–279, August 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0533-7.

Baglietto:2013:CNS

[BAC13] Gabriel Baglietto, Ezequiel V. Albano, and Julián Candia. Complex network structure of flocks in the standard vicsek model. Journal of Statistical Physics, 153(2):270–288, October 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0827-4.

Baeyens:2011:SMH

[Bae11] Bruno Baeyens. Statistical mechanics of hard spheres: Application of the improved scaled particle theory to the hard sphere crystal. Journal of Statistical Physics, 145(6):1640– 1648, December 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0321-9.

Baird:2010:MGB

[Bai10] Jim Baird. Moshe Gitterman: Chemistry Versus Physics: Chemical Reactions Near Critical Points. Journal of Statistical Physics, 141(3):607–608, November 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0061-2.

Baker:2010:PPL

[Bak10] George A. Baker, Jr. A personal perspective on the last halfcentury of critical phenomena. Journal of Statistical Physics, 138(1–3):8–19, February 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9880-4.

Ball:2014:RCI

[Bal14] Neville Ball. Rigorous confidence intervals on critical thresholds in 3 dimensions. Journal of Statistical Physics, 156 (3):574–585, August 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1018-7.

Baladi:2017:QUA

[Bal17] Viviane Baladi. The quest for the ultimate anisotropic Banach space. Journal of Statistical Physics, 166(3–4): 525–557, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1663-0. See

correction [Bal18].

Baladi:2018:CQU

[Bal18] Viviane Baladi. Correction to: The Quest for the Ultimate Anisotropic Banach Space. Journal of Statistical Physics, 170(6):1242–1247, March 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-018-1976-2.

pdf. See [Bal17].

Bandyopadhyay:2010:DCM

[Ban10] Malay Bandyopadhyay. Dissipative cyclotron motion of a charged quantum-oscillator and third law. Journal of Statistical Physics, 140(3):603–618, August 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-9998-4.

Bao:2017:NMT

[Bao17] J.-D. Bao. Non-Markovian two-time correlation dynamics and nonergodicity. Journal of Statistical Physics, 168(3):561–572, August 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Barre:2012:CMR

[Bar12] J. Barré. Combinatorial models of rigidity and renormalization. Journal of Statistical Physics, 146(2):359–377, January 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0394-5.

Bargigli:2014:SEE

[Bar14] Leonardo Bargigli. Statistical ensembles for economic networks. Journal of Statistical Physics, 155(4):810–825, May 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-0968-0.

Barton:2016:NFH

[Bar16] Gabriel Barton. Near-field heat flow between two quantum oscillators. Journal of Statistical Physics, 165(6):1153–1180, December 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-016-1654-1.pdf.

Ben-Ari:2016:SME

[BAS16] Iddo Ben-Ari and Rinaldo B. Schinazi. A stochastic model for the evolution of a quasispecies. Journal of Statistical Physics, 162(2):415–425, January 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1405-8.

Ben-Ari:2018:LVB

[BAS18] Iddo Ben-Ari and Roger W. C. Silva. On a local version of the Bak–Sneppen model. Journal of Statistical Physics, 173 (2):362–380, October 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Batterman:2017:PIK

[Bat17]  Robert W. Batterman.  Philosophical implications of

Kadanoff’s work on the renormalization group. Journal of Statistical Physics, 167(3–4):559–574, May 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Baxter:2011:OKC

[Bax11] R. J. Baxter. Onsager and Kaufman’s calculation of the spontaneous magnetization of the Ising model. Journal of Statistical Physics, 145(3):518–548, November 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0213-z.

Baxter:2012:OKC

[Bax12] R. J. Baxter. Onsager and Kaufman’s calculation of the spontaneous magnetization of the Ising model: II. Journal of Statistical Physics, 149(6):1164–1167, December 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0658-8.

Bakhtin:2011:OSB

[BB11] Yuri Bakhtin and Leonid Bunimovich. The optimal sink and the best source in a Markov chain. Journal of Statistical Physics, 143(5):943–954, June 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0223-x.

Bertini:2013:DPF

[BB13] Lorenzo Bertini and Paolo Buttà. Drift of phase fluctuations in the ABC model. Journal of Statistical Physics, 152(1):15– 36, July 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0757-1.

Benzi:2015:HIT

[BB15] Roberto Benzi and Luca Biferale. Homogeneous and isotropic turbulence: A short survey on recent developments. Journal of Statistical Physics, 161(6):1351–1365, December 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1323-9;   http://link.springer.com/
content/pdf/10.1007/s10955-015-1323-9.pdf.

Bradde:2017:PMR

[BB17] Serena Bradde and William Bialek. PCA meets RG. Journal of Statistical Physics, 167(3–4):462–475, May 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Baryshnikov:2011:TDQ

[BBBP11] Yuliy Baryshnikov, Wil Brady, Andrew Bressler, and Robin Pemantle. Two-dimensional quantum random walk. Journal of Statistical Physics, 142(1):78–107, January 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0098-2.

Bock:2015:SPW

[BBC15] Wolfgang Bock, Jinky B. Bornales, and Cresente O. Cabahug. Scaling properties of weakly self-avoiding fractional Brownian motion in one dimension. Journal of Statistical Physics, 161(5):1155–1162, December 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1368-9.

Barre:2018:DLD

[BBC18] J. Barré, C. Bernardin, and R. Chetrite. Density large deviations for multidimensional stochastic hyperbolic conservation laws. Journal of Statistical Physics, 170(3):466–491, February 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Binder:2011:MCM

[BBD^{+}11] Kurt Binder, Benjamin Block, Subir K. Das, Peter Virnau, and David Winter. Monte Carlo methods for estimating interfacial free energies and line tensions. Journal of Statistical Physics, 144(3):690–729, August 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0226-7.

Bannink:2019:ILC

[BBGS19] Tom Bannink, Harry Buhrman, András Gilyén, and Mario Szegedy. The interaction light cone of the discrete Bak– Sneppen, contact and other local processes. Journal of Statistical Physics, 176(6):1500–1525, September 2019. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-019-02351-y.pdf.

Berrada:2011:GSC

[BBH11] K. Berrada, M. El Baz, and Y. Hassouni. Generalized spin coherent states: Construction and some physical properties. Journal of Statistical Physics, 142(3):510–523, February 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0124-z.

Brotto:2017:MDD

[BBK17] Tommaso Brotto, Guy Bunin, and Jorge Kurchan. A model with Darwinian dynamics on a rugged landscape. Journal of Statistical Physics, 166(3–4):1065–1077, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1637-2.

Bertola:2012:SRH

[BBLP12] M. Bertola, R. Buckingham, S. Y. Lee, and V. Pierce. Spectra of random Hermitian matrices with a small-rank external source: The critical and near-critical regimes. Journal of Statistical Physics, 146(3):475–518, February 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0409-2.

Bertola:2013:SRH

[BBLP13] M. Bertola, R. Buckingham, S. Y. Lee, and V. Pierce. Spectra of random Hermitian matrices with a small-rank external source: The supercritical and subcritical regimes. Journal of Statistical Physics, 153(4):654–697, November 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0845-2.

Baiesi:2010:NLR

[BBMW10] Marco Baiesi, Eliran Boksenbojm, Christian Maes, and Bram Wynants. Nonequilibrium linear response for Markov dynamics, II: Inertial dynamics. Journal of Statistical Physics, 139(3):492–505, May 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9951-6.

Bardet:2017:PTE

[BBP17] Ivan Bardet, Denis Bernard, and Yan Pautrat. Passage times, exit times and Dirichlet problems for open quantum walks. Journal of Statistical Physics, 167(2):173–204, April 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-017-1749-3.

Bachmann:2012:DQW

[BBR12] Sven Bachmann, Maximilian Butz, and Wojciech De Roeck. Disordered quantum wires: Microscopic origins of the DMPK theory and Ohm’s law. Journal of Statistical Physics, 148 (1):164–189, July 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0517-7.

Bouarnouna:2019:MRT

[BBR^{+}19] Kaoutar Bouarnouna, Abdelkader Boutra, Karim Ragui, Nabila Labsi, and Youb Khaled Benkahla. Multiplerelaxation-time lattice Boltzmann model for flow and convective heat transfer in channel with porous media. Journal of Statistical Physics, 174(5):972–991, March 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Berestycki:2011:SNC

[BBS11] Julien Berestycki, Nathanaël Berestycki, and Jason Schweinsberg. Survival of near-critical branching Brownian motion. Journal of Statistical Physics, 143(5):833–854, June 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0224-9.

Bauerschmidt:2014:SLC

[BBS14] Roland Bauerschmidt, David C. Brydges, and Gordon Slade. Scaling limits and critical behaviour of the 4-dimensional ncomponent |φ|^{4} spin model. Journal of Statistical Physics, 157(4–5):692–742, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1060-5.

Bauerschmidt:2015:RGM

[BBS15a] Roland Bauerschmidt, David C. Brydges, and Gordon Slade. A renormalisation group method. III. perturbative analysis. Journal of Statistical Physics, 159(3):492–529, May 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1165-x.

Bologna:2015:SDF

[BBS^{+}15b] M. Bologna, M. T. Beig, A. Svenkeson, P. Grigolini, and B. J. West. Spectral decomposition of a Fokker–Planck equation at criticality. Journal of Statistical Physics, 160(2):466–476, July 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1262-5.

Bishop:2015:LTR

[BBW15] Michael Bishop, Vita Borovyk, and Jan Wehr. Lifschitz tails for random Schrödinger operator in Bernoulli distributed potentials. Journal of Statistical Physics, 160(1):151–162, July 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1242-9.

Bissacot:2010:PTF

[BC10] Rodrigo Bissacot and Leandro Cioletti. Phase transition in ferromagnetic Ising models with non-uniform external magnetic fields. Journal of Statistical Physics, 139(5):769–778, June 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-9961-4.

Biskup:2011:AMC

[BC11] Marek Biskup and Nicholas Crawford. Absence of magnetism in continuous-spin systems with long-range antialigning forces. Journal of Statistical Physics, 144(4):731–748, August 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0274-z.

Barletti:2012:DIQ

[BC12] Luigi Barletti and Carlo Cintolesi. Derivation of isothermal quantum fluid equations with Fermi–Dirac and Bose– Einstein statistics. Journal of Statistical Physics, 148(2):353– 386, August 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0535-5.

Bauer:2014:AFM

[BC14a] M. Bauer and F. Cornu. Affinity and fluctuations in a mesoscopic noria. Journal of Statistical Physics, 155(4):703– 736, May 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-0969-z.

Bo:2014:EPS

[BC14b] Stefano Bo and Antonio Celani. Entropy production in stochastic systems with fast and slow time-scales. Journal of Statistical Physics, 154(5):1325–1351, March 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0922-1.

Barato:2015:FVL

[BC15a] Andre C. Barato and Raphael Chetrite. A formal view on level 2.5 large deviations and fluctuation relations. Journal of Statistical Physics, 160(5):1154–1172, September 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1283-0.

Barnett:2015:CMI

[BC15b] Nix Barnett and James P. Crutchfield. Computational mechanics of input–output processes: Structured transformations and the ϵ-transducer. Journal of Statistical Physics, 161(2):404–451, October 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1327-5.

Bo:2016:DCC

[BC16] Stefano Bo and Antonio Celani. Detecting concentration changes with cooperative receptors. Journal of Statistical Physics, 162(5):1365–1382, March 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-015-1354-

Budanur:2017:UMR

[BC17] Nazmi Burak Budanur and Predrag Cvitanović. Unstable manifolds of relative periodic orbits in the symmetry-reduced state space of the Kuramoto–Sivashinsky system. Journal of Statistical Physics, 167(3–4):636–655, May 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bethuelsen:2018:OSC

[BC18] Stein Andreas Bethuelsen and Diana Conache. One-sided continuity properties for the schonmann projection. Journal of Statistical Physics, 172(4):1147–1163, August 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bomfim:2019:LRC

[BC19] Thiago Bomfim and Armando Castro. Linear response, and consequences for differentiability of statistical quantities and multifractal analysis. Journal of Statistical Physics, 174(1): 135–159, January 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Barton:2016:EPF

[BCC^{+}16] John P. Barton, Arup K. Chakraborty, Simona Cocco, Hugo Jacquin, and Rémi Monasson. On the entropy of protein families. Journal of Statistical Physics, 162(5):1267–1293, March 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1441-4.

Bhamidi:2018:WER

[BCCD18] Shankar Bhamidi, Suman Chakraborty, Skyler Cranmer, and Bruce Desmarais. Weighted exponential random graph models: Scope and large network limits. Journal of Statistical Physics, 173(3–4):704–735, November 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bonetto:2014:PCT

[BCE^{+}14] F. Bonetto, E. A. Carlen, R. Esposito, J. L. Lebowitz, and R. Marra. Propagation of chaos for a thermostated kinetic model. Journal of Statistical Physics, 154(1–2):265–285, January 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0861-2.

Billiot:2010:CZP

[BCF10] J.-M. Billiot, F. Corset, and E. Fontenas. On complex zeros of the q-Potts partition function for a self-dual family of graphs. Journal of Statistical Physics, 139(6):960–971, June 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-9977-9.

Bessaih:2017:MFL

[BCF17] Hakima Bessaih, Michele Coghi, and Franco Flandoli. Mean field limit of interacting filaments and vector valued nonlinear PDEs. Journal of Statistical Physics, 166(5):1276–1309, March 2017. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1706-6.

Bessaih:2019:MFL

[BCF19] Hakima Bessaih, Michele Coghi, and Franco Flandoli. Mean field limit of interacting filaments for 3D Euler equations. Journal of Statistical Physics, 174(3):562–578, February 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Benatti:2017:LLM

[BCFS17] F. Benatti, F. Carollo, R. Floreanini, and J. Surace. Longlived mesoscopic entanglement between two damped infinite harmonic chains. Journal of Statistical Physics, 168(3):620– 651, August 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Barato:2012:GCE

[BCHM12] Andre Cardoso Barato, Raphaël Chetrite, Haye Hinrichsen, and David Mukamel. A Gallavotti–Cohen–Evans– Morriss like symmetry for a class of Markov jump processes. Journal of Statistical Physics, 146(2):294–313, January 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0389-2.

Ban:2019:DIT

[BCHW19] Jung-Chao Ban, Chih-Hung Chang, Nai-Zhu Huang, and Yu-Liang Wu. Decidability of irreducible tree shifts of finite type.

Journal of Statistical Physics, 177(6):1043–1062, December 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Behnejad:2015:ECP

[BCJ15] Hassan Behnejad, Hashem Cheshmpak, and Asma Jamali. The extended crossover Peng–Robinson equation of state for describing the thermodynamic properties of pure fluids. Journal of Statistical Physics, 158(2):372–385, January 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1134-4.

Bandopadhyay:2017:RBM

[BCJ17] Swarnali Bandopadhyay, Debasish Chaudhuri, and A. M. Jayannavar. Rotational Brownian motion: Trajectory, reversibility and stochastic entropy. Journal of Statistical Physics, 168(3):549–560, August 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Berlyand:2019:CAS

[BCJP19] Leonid Berlyand, Robert Creese, Pierre-Emmanuel Jabin, and Mykhailo Potomkin. Continuum approximations to systems of correlated interacting particles. Journal of Statistical Physics, 174(4):808–829, February 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bonetto:2012:SSS

[BCKL12] Federico Bonetto, Nikolai Chernov, Alexey Korepanov, and Joel L. Lebowitz. Spatial structure of stationary nonequilibrium states in the thermostatted periodic Lorentz gas. Journal of Statistical Physics, 146(6):1221–1243, March 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0444-7.

Bae:2019:KDH

[BCKY19] Hyeong-Ohk Bae, Seung-Yeon Cho, Jeongho Kim, and Seok-Bae Yun. A kinetic description for the herding behavior in financial market. Journal of Statistical Physics, 176(2):398– 424, July 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Binder:2010:CSC

[BCL10a] I. Binder, L. Chayes, and H. K. Lei. On convergence to SLE_{6} I: Conformal invariance for certain models of the bondtriangular type. Journal of Statistical Physics, 141(2):359– 390, October 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-010-0052-3.pdf.

Binder:2010:CSI

[BCL10b] I. Binder, L. Chayes, and H. K. Lei. On convergence to SLE_{6} II: Discrete approximations and extraction of Cardy’s formula for general domains. Journal of Statistical Physics, 141 (2):391–408, October 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-010-0053-2.
pdf.

Beltran:2019:CRW

[BCL19] J. Beltrán, E. Chavez, and C. Landim. From coalescing random walks on a torus to Kingman’s coalescent. Journal of Statistical Physics, 177(6):1172–1206, December 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bianchi:2016:RWO

[BCLL16] Alessandra Bianchi, Giampaolo Cristadoro, Marco Lenci, and Marilena Ligabò. Random walks in a one–dimensional Lévy random environment. Journal of Statistical Physics, 163(1): 22–40, April 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1469-0.

Balint:2010:HTI

[BCM10] András Bálint, Federico Camia, and Ronald Meester. The high temperature Ising model on the triangular lattice is a critical Bernoulli percolation model. Journal of Statistical Physics, 139(1):122–138, April 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-010-9930-y.
pdf.

Butta:2012:TET

[BCM12] Paolo Buttà, Guido Cavallaro, and Carlo Marchioro. Time evolution of two dimensional systems with infinitely many particles mutually interacting via very singular forces. Journal of Statistical Physics, 147(2):412–423, April 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0461-6.

Benedetto:2016:EDO

[BCM16] Dario Benedetto, Emanuele Caglioti, and Umberto Montemagno. Exponential dephasing of oscillators in the kinetic Kuramoto model. Journal of Statistical Physics, 162(4):813–823, February 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1426-3.

Barre:2015:MIP

[BCMP15] Julien Barré, Raphaël Chétrite, Massimiliano Muratori, and Fernando Peruani. Motility-induced phase separation of active particles in the presence of velocity alignment. Journal of Statistical Physics, 158(3):589–600, February 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1008-9.

Bellouquid:2012:RBB

[BCNS12] Abdelghani Bellouquid, Juan Calvo, Juanjo Nieto, and Juan Soler. On the relativistic BGK–Boltzmann model: Asymptotics and hydrodynamics. Journal of Statistical Physics, 149 (2):284–316, October 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0600-0.

Bertini:2011:DBA

[BCP11] Lorenzo Bertini, Nicoletta Cancrini, and Gustavo Posta. On the dynamical behavior of the ABC model. Journal of Statistical Physics, 144(6):1284–1307, September 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0294-8.

Benettin:2013:FPU

[BCP13] G. Benettin, H. Christodoulidi, and A. Ponno. The Fermi– Pasta–Ulam problem and its underlying integrable dynamics. Journal of Statistical Physics, 152(2):195–212, July 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0760-6.

Braides:2018:ABG

[BCPS18] Andrea Braides, Andrea Causin, Andrey Piatnitski, and Margherita Solci. Asymptotic behaviour of ground states for mixtures of ferromagnetic and antiferromagnetic interactions in a dilute regime. Journal of Statistical Physics, 171(6):1096– 1111, June 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Barucca:2018:TIA

[BCS18] Paolo Barucca, Guido Caldarelli, and Tiziano Squartini. Tackling information asymmetry in networks: A new entropybased ranking index. Journal of Statistical Physics, 173(3–4): 1028–1044, November 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Barreira:2013:MAA

[BCW13] Luis Barreira, Yongluo Cao, and Juan Wang. Multifractal analysis of asymptotically additive sequences. Journal of Statistical Physics, 153(5):888–910, December 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0853-2.

Braides:2016:CMI

[BCY16] Andrea Braides, Marco Cicalese, and Nung Kwan Yip. Crystalline motion of interfaces between patterns. Journal of Statistical Physics, 165(2):274–319, October 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1609-6.

Bodineau:2011:PFA

[BD11a] T. Bodineau and B. Derrida. Phase fluctuations in the ABC model. Journal of Statistical Physics, 145(3):745–762, November 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0315-7.

Brunet:2011:BRW

[BD11b] Éric Brunet and Bernard Derrida. A branching random walk seen from the tip. Journal of Statistical Physics, 143(3):420– 446, May 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0185-z.

Brunet:2015:EST

[BD15a] Éric Brunet and Bernard Derrida. An exactly solvable travelling wave equation in the Fisher–KPP class. Journal of Statistical Physics, 161(4):801–820, November 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1350-6.

Bullara:2015:CEL

[BD15b] D. Bullara and Y. De Decker. Chemical equilibrium on low dimensional supports: Connecting the microscopic mechanism to the macroscopic observations. Journal of Statistical Physics, 161(1):210–226, October 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1314-x.

Berglund:2016:IDM

[BD16a] Nils Berglund and Sébastien Dutercq. Interface dynamics of a metastable mass-conserving spatially extended diffusion. Journal of Statistical Physics, 162(2):334–370, January 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1415-6.

Blanchet:2016:TIB

[BD16b] Adrien Blanchet and Pierre Degond. Topological interactions in a Boltzmann-type framework. Journal of Statistical Physics, 163(1):41–60, April 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-016-1471-6.
pdf.

Blanchet:2017:KMT

[BD17] Adrien Blanchet and Pierre Degond. Kinetic models for topological nearest-neighbor interactions. Journal of Statistical Physics, 169(5):929–950, December 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-017-1882-z.pdf.

Bauer:2014:IRC

[BDDH14] Gernot Bauer, Dirk-André Deckert, Detlef Dürr, and Günter Hinrichs. On irreversibility and radiation in classical electrodynamics of point particles. Journal of Statistical Physics, 154(1–2):610–622, January 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0837-2.

Basor:2019:MSW

[BDES19] E. Basor, J. Dubail, T. Emig, and R. Santachiara. Modified Szegő–Widom asymptotics for block Toeplitz matrices with zero modes. Journal of Statistical Physics, 174(1):28–39, January 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bedingham:2014:MDR

[BDG^{+}14] Daniel Bedingham, Detlef Dürr, GianCarlo Ghirardi, Sheldon Goldstein, Roderich Tumulka, and Nino Zanghì. Matter density and relativistic models of wave function collapse. Journal of Statistical Physics, 154(1–2):623–631, January 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0814-9.

Bodineau:2010:DSD

[BDL10] T. Bodineau, B. Derrida, and J. L. Lebowitz. A diffusive system driven by a battery or by a smoothly varying field. Journal of Statistical Physics, 140(4):648–675, August 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-0012-y.

Britton:2011:WCM

[BDL11] Tom Britton, Maria Deijfen, and Fredrik Liljeros. A weighted configuration model and inhomogeneous epidemics. Journal of Statistical Physics, 145(5):1368–1384, December 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0343-3.

Beznea:2016:SEF

[BDL16] Lucian Beznea, Madalina Deaconu, and Oana Lupascu. Stochastic equation of fragmentation and branching processes related to avalanches. Journal of Statistical Physics, 162(4): 824–841, February 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1432-5.

Blass:2013:ABF

[BdlL13] Timothy Blass and Rafael de la Llave. The analyticity breakdown for Frenkel–Kontorova models in quasi-periodic media: Numerical explorations. Journal of Statistical Physics, 150 (6):1183–1200, March 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0718-8.

Balazs:2019:ZRR

[BDP19] Márton Balázs, Lewis Duffy, and Dimitri Pantelli. q-zero range has random walking shocks. Journal of Statistical Physics, 174(5):958–971, March 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s10955-018-
02218-8.pdf.

Bach:2014:SDP

[BdSPMS14] Volker Bach, Walter de Siqueira Pedra, Marco Merkli, and Israel Michael Sigal. Suppression of decoherence by periodic forcing. Journal of Statistical Physics, 155(6):1271–1298, June 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-0952-8.

Birmpa:2017:LDM

[BDT17] P. Birmpa, N. Dirr, and D. Tsagkarogiannis. Large deviations for the macroscopic motion of an interface. Journal of Statistical Physics, 166(5):1163–1192, March 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-017-1720-3.pdf.

Barreira:2012:LFC

[BDV12] Luis Barreira, Davor Dragicević, and Claudia Valls. Lyapunov functions and cone families. Journal of Statistical Physics, 148(1):137–163, July 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0524-8.

Bleher:2017:TEC

[BDY17] Pavel Bleher, Alfredo Deaño, and Maxim Yattselev. Topological expansion in the complex cubic log-gas model: Onecut case. Journal of Statistical Physics, 166(3–4):784–827, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1621-x.

Briant:2016:CPH

[BE16] Marc Briant and Amit Einav. On the Cauchy problem for the homogeneous Boltzmann–Nordheim equation for bosons: Local existence, uniqueness and creation of moments. Journal of Statistical Physics, 163(5):1108–1156, June 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1517-9.

Beck:2010:DAK

[Bec10] József Beck. Deterministic approach to the kinetic theory of gases. Journal of Statistical Physics, 138(1–3):160–269, February 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-009-9871-5.

Beck:2011:HFE

[Bec11] Thomas L. Beck. Hydration free energies by energetic partitioning of the potential distribution theorem. Journal of Statistical Physics, 145(2):335–354, October 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0298-4.

Barmpalias:2015:TPD

[BELP15] George Barmpalias, Richard Elwes, and Andy Lewis-Pye. Tipping points in 1-dimensional Schelling models with switching agents. Journal of Statistical Physics, 158(4):806–852, February 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1141-5.

Barmpalias:2016:USS

[BELP16] George Barmpalias, Richard Elwes, and Andrew Lewis-Pye. Unperturbed Schelling segregation in two or three dimensions. Journal of Statistical Physics, 164(6):1460–1487, September 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1589-6.

Barmpalias:2018:MPO

[BELP18] George Barmpalias, Richard Elwes, and Andrew Lewis-Pye. Minority population in the one-dimensional Schelling model of segregation. Journal of Statistical Physics, 173(5):1408–1458, December 2018. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-018-2146-2.pdf.

Bleher:2018:PST

[BEP18] Pavel Bleher, Brad Elwood, and Drazen Petrović. The Pfaffian sign theorem for the dimer model on a triangular lattice. Journal of Statistical Physics, 171(3):400–426, May 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bernhoff:2012:HSP

[Ber12] Niclas Bernhoff. Half-space problem for the discrete Boltzmann equation: Condensing vapor flow in the presence of a non-condensable gas. Journal of Statistical Physics, 147 (6):1156–1181, July 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0513-y.

Berger:2014:PMR

[Ber14] Quentin Berger. Pinning model in random correlated environment: Appearance of an infinite disorder regime. Journal of Statistical Physics, 155(3):544–570, May 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0965-3.

Bernhoff:2015:HSP

[Ber15] Niclas Bernhoff. Half-space problems for a linearized discrete quantum kinetic equation. Journal of Statistical Physics, 159(2):358–379, April 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1190-4.

Bernhoff:2018:DVM

[Ber18] Niclas Bernhoff. Discrete velocity models for polyatomic molecules without nonphysical collision invariants. Journal of Statistical Physics, 172(3):742–761, August 2018. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-018-2063-4.pdf.

Bertoin:2019:VHS

[Ber19] Jean Bertoin. A version of Herbert A. Simon’s model with slowly fading memory and its connections to branching processes. Journal of Statistical Physics, 176(3):679–691, August 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Betz:2014:RPR

[Bet14] Volker Betz. Random permutations of a regular lattice. Journal of Statistical Physics, 155(6):1222–1248, June 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-0945-7.

Barletti:2010:DLT

[BF10] Luigi Barletti and Giovanni Frosali. Diffusive limit of the twoband k · p model for semiconductors. Journal of Statistical Physics, 139(2):280–306, April 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9940-9.

Barreto:2011:DFL

[BF11] Stephen Dias Barreto and Francesco Fidaleo. Disordered fermions on lattices and their spectral properties. Journal of Statistical Physics, 143(4):657–684, May 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0197-8.

Bessaih:2012:IMG

[BF12] Hakima Bessaih and Benedetta Ferrario. Invariant measures of Gaussian type for 2D turbulence. Journal of Statistical Physics, 149(2):259–283, October 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0601-z.

Ballesteros:2016:IAI

[BFFS16] M. Ballesteros, M. Fraas, J. Fröhlich, and B. Schubnel. Indirect acquisition of information in quantum mechanics. Journal of Statistical Physics, 162(4):924–958, February 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1410-y.

Bandtlow:2010:AFF

[BFKP10] Oscar F. Bandtlow, Jan Fiala, Peter Kleban, and Thomas Prellberg. Asymptotics of the Farey fraction spin chain free energy at the critical point. Journal of Statistical Physics, 138(1–3):447–464, February 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9900-4.

Balazs:2010:RWS

[BFKR10] Márton Balázs, György Farkas, Péter Kovács, and Attila Rákos. Random walk of second class particles in product shock measures. Journal of Statistical Physics, 139(2):252– 279, April 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-9933-8.

Brightwell:2018:SMB

[BFL18] Graham Brightwell, Marianne Fairthorne, and Malwina J. Luczak. The supermarket model with bounded queue lengths in equilibrium. Journal of Statistical Physics, 173(3–4): 1149–1194, November 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-018-2044-7.
pdf.

Butta:2010:SMB

[BFM10] Paolo Buttà, Giorgio Ferrari, and Carlo Marchioro. Speedy motions of a body immersed in an infinitely extended medium. Journal of Statistical Physics, 140(6):1182–1194, September 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-0036-3.

Brzezniak:2011:CIP

[BFNZ11] Z. Brzeźniak, F. Flandoli, M. Neklyudov, and B. Zegarliński. Conservative interacting particles system with anomalous rate of ergodicity. Journal of Statistical Physics, 144(6):1171–1185, September 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0327-3.

Bissacot:2010:CCE

[BFP10] Rodrigo Bissacot, Roberto Fernández, and Aldo Procacci. On the convergence of cluster expansions for polymer gases. Journal of Statistical Physics, 139(4):598–617, May 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-9956-1.

Bessaih:2010:SAS

[BFT10] Hakima Bessaih, Franco Flandoli, and Edriss S. Titi. Stochastic attractors for Shell phenomenological models of turbulence. Journal of Statistical Physics, 140(4):688–717, August 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-010-0010-0.pdf.

Barreira:2010:IMI

[BFVZ10] Luis Barreira, Meng Fan, Claudia Valls, and Jimin Zhang. Invariant manifolds for impulsive equations and nonuniform polynomial dichotomies. Journal of Statistical Physics, 141 (1):179–200, October 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0043-4.

Barral:2011:ELD

[BG11] J. Barral and P. Gonçalves. On the estimation of the large deviations spectrum. Journal of Statistical Physics, 144(6):

1256–1283, September 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0296-6.

Bostan:2012:ISM

[BG12] Mihai Bostan and Irene M. Gamba. Impact of strong magnetic fields on collision mechanism for transport of charged particles. Journal of Statistical Physics, 148(5):856–895, September 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0560-4.

Bloznelis:2014:PAA

[BG14] Mindaugas Bloznelis and Friedrich Götze. Preferred attachment in affiliation networks. Journal of Statistical Physics, 156(4):800–821, August 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1033-8.

Bachmann:2017:MFL

[BG17a] Sven Bachmann and François Genoud. Mean-field limit and phase transitions for nematic liquid crystals in the continuum. Journal of Statistical Physics, 168(4):746–771, August 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bunimovich:2017:TPM

[BG17b] Leonid A. Bunimovich and Alexander Grigo. Transport processes from mechanics: Minimal and simplest models. Journal of Statistical Physics, 166(3–4):750–764, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1701-y.

Bricmont:2019:DTQ

[BG19] Jean Bricmont and Sheldon Goldstein. Diagnosing the trouble with quantum mechanics. Journal of Statistical Physics, 175 (3–4):690–703, May 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Barra:2010:RSA

[BGG10] Adriano Barra, Giuseppe Genovese, and Francesco Guerra. The replica symmetric approximation of the analogical neural network. Journal of Statistical Physics, 140(4):784–796, August 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-0020-y.

Benfatto:2015:KEF

[BGJ15a] Giuseppe Benfatto, Giovanni Gallavotti, and Ian Jauslin. Kondo effect in a fermionic hierarchical model. Journal of Statistical Physics, 161(5):1203–1230, December 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1378-7.

Bernardin:2015:NDS

[BGJ^{+}15b] Cédric Bernardin, Patrícia Gonçalves, Milton Jara, Makiko Sasada, and Marielle Simon. From normal diffusion to superdiffusion of energy in the evanescent flip noise limit. Journal of Statistical Physics, 159(6):1327–1368, June 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1235-8.

Bertini:2012:TTN

[BGJLL12] Lorenzo Bertini, Davide Gabrielli, Giovanni Jona-Lasinio, and Claudio Landim. Thermodynamic transformations of nonequilibrium states. Journal of Statistical Physics, 149(5):773–802, November 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0624-5.

Bachmann:2010:TOC

[BGL10] S. Bachmann, G. M. Graf, and G. B. Lesovik. Time ordering and counting statistics. Journal of Statistical Physics, 138(1–3):333–350, February 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9885-z.

Benichou:2011:MFP

[BGL^{+}11] O. Bénichou, D. S. Grebenkov, P. E. Levitz, C. Loverdo, and R. Voituriez. Mean first-passage time of surface-mediated diffusion in spherical domains. Journal of Statistical Physics, 142(4):657–685, February 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0138-6.

Bernardin:2014:ENE

[BGL14] Cédric Bernardin, Patrícia Gonçalves, and Claudio Landim. Entropy of non-equilibrium stationary measures of boundary driven TASEP. Journal of Statistical Physics, 154(1–2):378– 420, January 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0882-x.

Beaud:2013:SCT

[BGLL13] V. Beaud, G. M. Graf, A. V. Lebedev, and G. B. Lesovik. Statistics of charge transport and modified time ordering. Journal of Statistical Physics, 153(2):177–210, October 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0815-8.

Belton:2019:QSC

[BGLZ19] Alexander C. R. Belton, Michał Gnacik, J. Martin Lindsay, and Ping Zhong. Quasifree stochastic cocycles and quantum random walks. Journal of Statistical Physics, 176(1): 1–39, July 2019. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-019-02273-9.pdf.

Berestycki:2014:GWT

[BGMS14] Nathanaël Berestycki, Nina Gantert, Peter Mörters, and Nadia Sidorova. Galton–Watson trees with vanishing martingale limit. Journal of Statistical Physics, 155(4):737–762, May 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-0975-1.

Bahsoun:2016:RAD

[BGN16a] Wael Bahsoun, Stefano Galatolo, and Isaia Nisoli. Rigorous approximation of diffusion coefficients for expanding maps. Journal of Statistical Physics, 163(6):1486–1503, June 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1523-y;   http://link.springer.com/
content/pdf/10.1007/s10955-016-1523-y.pdf.

Bouchet:2016:PCQ

[BGN16b] Freddy Bouchet, Krzysztof Gawȩdzki, and Cesare Nardini. Perturbative calculation of quasi-potential in nonequilibrium diffusions: A mean-field example. Journal of Statistical Physics, 163(5):1157–1210, June 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1503-2.

Balint:2017:LMP

[BGN^{+}17] Péter Bálint, Thomas Gilbert, Péter Nándori, Domokos Szász, and Imre Péter Tóth. On the limiting Markov process of energy exchanges in a rarely interacting ball–piston gas. Journal of Statistical Physics, 166(3–4):903–925, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1598-5.

Bertini:2010:DAM

[BGP10] Lorenzo Bertini, Giambattista Giacomin, and Khashayar Pakdaman. Dynamical aspects of mean field plane rotators and the Kuramoto model. Journal of Statistical Physics, 138(1– 3):270–290, February 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9908-9.

Benaych-Georges:2015:PSM

[BGP15a] Florent Benaych-Georges and Sandrine Péché. Poisson statistics for matrix ensembles at large temperature. Journal of Statistical Physics, 161(3):633–656, November 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1340-8.

Bobylev:2015:SPL

[BGP15b] Alexander Bobylev, Irene Gamba, and Irina Potapenko. On some properties of the Landau kinetic equation. Journal of Statistical Physics, 161(6):1327–1338, December 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1311-0;   http://link.springer.com/
content/pdf/10.1007/s10955-015-1311-0.pdf.

Buric:2011:ICD

[BGTV11] Nikola Burić, Ines Grozdanović, Kristina Todorović, and Nebojsa Vasović. Influence of coupling delay on noise induced coherent oscillations in excitable systems. Journal of Statistical Physics, 145(1):175–186, October 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0316-6.

Bouchet:2016:LDF

[BGTVE16] Freddy Bouchet, Tobias Grafke, Tomás Tangarife, and Eric Vanden-Eijnden. Large deviations in fast–slow systems. Journal of Statistical Physics, 162(4):793–812, February 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1449-4.

Bobylev:2017:RRL

[BGZ17] Alexander Bobylev, Irene M. Gamba, and Chenglong Zhang. On the rate of relaxation for the Landau kinetic equation and related models. Journal of Statistical Physics, 168(3):535– 548, August 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bashan:2011:CEC

[BH11] Amir Bashan and Shlomo Havlin. The combined effect of connectivity and dependency links on percolation of networks. Journal of Statistical Physics, 145(3):686–695, November 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0333-5.

Bhattacharjee:2015:SCF

[Bha15] Jayanta K. Bhattacharjee. Self-consistent field theory for the convective turbulence in a Rayleigh–Benard system in the infinite Prandtl number limit. Journal of Statistical Physics, 160(6):1519–1528, September 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1292-z.

Bernardi:2012:ESP

[BHF^{+}12] Stefano Bernardi, J. S. Hansen, Federico Frascoli, B. D. Todd, and Carl P. Dettmann. Ergodicity of a single particle confined in a nanopore. Journal of Statistical Physics, 148(6): 1156–1169, September 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0577-8.

Bleher:2012:NND

[BHJ^{+}12] Pavel M. Bleher, Youkow Homma, Lyndon L. Ji, Roland K. W. Roeder, and Jeffrey D. Shen. Nearest neighbor distances on a circle: Multidimensional case. Journal of Statistical Physics, 146(2):446–465, January 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0367-8.

Borge-Holthoefer:2013:EIS

[BHMGM13] Javier Borge-Holthoefer, Sandro Meloni, Bruno Gonçalves, and Yamir Moreno. Emergence of influential spreaders in modified rumor models. Journal of Statistical Physics, 151 (1–2):383–393, April 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0595-6.

Bachmann:2015:PVB

[BHNY15] Sven Bachmann, Eman Hamza, Bruno Nachtergaele, and Amanda Young. Product vacua and boundary state models in d-dimensions. Journal of Statistical Physics, 160(3):636– 658, August 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1260-7.

Banavar:2012:PSS

[BHS^{+}12] Jayanth R. Banavar, T. X. Hoang, F. Seno, A. Trovato, and A. Maritan. Protein sequence and structure: Is one more fundamental than the other? Journal of Statistical Physics, 148(4):637–646, September 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0456-3.

Barato:2013:RMI

[BHS13] Andre C. Barato, David Hartich, and Udo Seifert. Rate of mutual information between coarse-grained non–Markovian variables. Journal of Statistical Physics, 153(3):460–478, November 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0834-5.

Bovier:2018:HLL

[BIM18] Anton Bovier, Dmitry Ioffe, and Patrick Müller. The hydrodynamic limit for local mean-field dynamics with unbounded spins. Journal of Statistical Physics, 172(2):434–457, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Brak:2011:IFA

[BIP11] R. Brak, G. K. Iliev, and T. Prellberg. An infinite family of adsorption models and restricted Łukasiewicz paths. Journal of Statistical Physics, 145(3):669–685, November 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0306-8.

Birrell:2018:EAL

[Bir18] Jeremiah Birrell. Entropy anomaly in Langevin–Kramers dynamics with a temperature gradient, matrix drag, and magnetic field. Journal of Statistical Physics, 173(6):1549–1586, December 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bishop:2019:SGT

[Bis19] Michael Bishop. Spectral gaps for the two-species product vacua and boundary states models on the d-dimensional lattice. Journal of Statistical Physics, 175(2):418–455, April 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Barrera:2016:ACS

[BJ16] G. Barrera and M. Jara. Abrupt convergence for stochastic small perturbations of one dimensional dynamical systems. Journal of Statistical Physics, 163(1):113–138, April 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1468-1.

Baity-Jesi:2017:ESF

[BJGL^{+}17] Marco Baity-Jesi, Carl P. Goodrich, Andrea J. Liu, Sidney R. Nagel, and James P. Sethna. Emergent SO(3) symmetry of the frictionless shear jamming transition. Journal of Statistical Physics, 167(3–4):735–748, May 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Boghosian:2015:TBE

[BJM15] Bruce M. Boghosian, Merek Johnson, and Jeremy A. Marcq. An H theorem for Boltzmann’s equation for the yard-sale model of asset exchange. Journal of Statistical Physics, 161 (6):1339–1350, December 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-015-1316-
8;   http://link.springer.com/content/pdf/10.1007/
s10955-015-1316-8.pdf.

Benoist:2017:ESO

[BJP17] Tristan Benoist, Vojkan Jaksić, and Claude-Alain Pillet. Energy statistics in open harmonic networks. Journal of Statistical Physics, 168(5):1016–1030, September 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bollobas:2010:CMR

[BJR10] Béla Bollobás, Svante Janson, and Oliver Riordan. The cut metric, random graphs, and branching processes. Journal of Statistical Physics, 140(2):289–335, July 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-9982-z.

Bermolen:2017:SLG

[BJS17] Paola Bermolen, Matthieu Jonckheere, and Jaron Sanders. Scaling limits and generic bounds for exploration processes. Journal of Statistical Physics, 169(5):989–1018, December 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bradford:2011:ATM

[BK11] Kyle Bradford and Yevgeniy Kovchegov. Adiabatic times for Markov chains and applications. Journal of Statistical Physics, 143(5):955–969, June 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0219-6.

Barany:2015:ACB

[BK15] Balázs Bárány and István Kolossváry. On the absolute continuity of the Blackwell measure. Journal of Statistical Physics, 159(1):158–171, April 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1176-7.

Benjamini:2017:CVD

[BK17a] Itai Benjamini and Gady Kozma. Continuous versus discrete spins in the hyperbolic plane. Journal of Statistical Physics, 168(1):11–14, July 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Berglund:2017:MSR

[BK17b] Nils Berglund and Christian Kuehn. Model spaces of regularity structures for space–fractional SPDEs. Journal of Statistical Physics, 168(2):331–368, July 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Baba:2018:EPA

[BK18] Hiroya Baba and Makoto Katori. Excursion processes associated with elliptic combinatorics. Journal of Statistical Physics, 171(6):1035–1066, June 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Berns:2015:MJD

[BKK15] Christoph Berns, Yuri Kondratiev, and Oleksandr Kutoviy. Markov jump dynamics with additive intensities in continuum: State evolution and mesoscopic scaling. Journal of Statistical Physics, 161(4):876–901, November 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1365-z.

Bernardin:2012:HSB

[BKLL12] C. Bernardin, V. Kannan, J. L. Lebowitz, and J. Lukkarinen. Harmonic systems with bulk noises. Journal of Statistical Physics, 146(4):800–831, February 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0416-3.

Broadbridge:2019:RSH

[BKLO19] Phil Broadbridge, Alexander D. Kolesnik, Nikolai Leonenko, and Andriy Olenko. Random spherical hyperbolic diffusion. Journal of Statistical Physics, 177(5):889–916, December 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Baake:2015:DTP

[BKM15] Michael Baake, Holger Kösters, and Robert V. Moody. Diffraction theory of point processes: Systems with clumping and repulsion. Journal of Statistical Physics, 159(4):915– 936, May 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1178-5.

Bunin:2013:CSB

[BKP13] Guy Bunin, Yariv Kafri, and Daniel Podolsky. Cusp singularities in boundary-driven diffusive systems. Journal of Statistical Physics, 152(1):112–135, July 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0752-6.

Bowler:2014:PRK

[BKPW14] Ken Bowler, Richard Kenway, Stuart Pawley, and David Wallace. Personal reflections on Kenneth Wilson at Princeton and Edinburgh. Journal of Statistical Physics, 157(4–5):639–643, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1057-0.

Balazs:2012:FBE

[BKS12] Márton Balázs, Júlia Komjáthy, and Timo Seppäläinen. Fluctuation bounds in the exponential bricklayers process. Journal of Statistical Physics, 147(1):35–62, April 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0470-5.

Beltran:2010:TMC

[BL10a] J. Beltrán and C. Landim. Tunneling and metastability of continuous time Markov chains. Journal of Statistical Physics, 140(6):1065–1114, September 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0030-9.

Bernardin:2010:ESN

[BL10b] Cédric Bernardin and Claudio Landim. Entropy of stationary nonequilibrium measures of boundary driven symmetric simple exclusion processes. Journal of Statistical Physics, 141(6):1014–1038, December 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0082-x.

Bodineau:2010:CLD

[BL10c] T. Bodineau and M. Lagouge. Current large deviations in a driven dissipative model. Journal of Statistical Physics, 139(2):201–218, April 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9934-7.

Britton:2010:DRN

[BL10d] Tom Britton and Mathias Lindholm. Dynamic random networks in dynamic populations. Journal of Statistical Physics, 139(3):518–535, May 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9952-5.

Berger:2011:EDF

[BL11a] Quentin Berger and Hubert Lacoin. The effect of disorder on the free-energy for the random walk pinning model: Smoothing of the phase transition and low temperature asymptotics. Journal of Statistical Physics, 142(2):322–341, January 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-0110-x.

Bolhuis:2011:RBP

[BL11b] Peter G. Bolhuis and Wolfgang Lechner. On the relation between projections of the reweighted path ensemble. Journal of Statistical Physics, 145(4):841–859, November 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-011-0324-6.pdf.

Beltran:2012:TMC

[BL12] J. Beltrán and C. Landim. Tunneling and metastability of continuous time Markov chains II, the nonreversible case. Journal of Statistical Physics, 149(4):598–618, November 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0617-4.

Boon:2015:NTA

[BL15] Jean Pierre Boon and James F. Lutsko. Nonlinear theory of anomalous diffusion and application to fluorescence correlation spectroscopy. Journal of Statistical Physics, 161 (6):1366–1378, December 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-015-1315-
9;   http://link.springer.com/content/pdf/10.1007/
s10955-015-1315-9.pdf.

Baik:2016:FFE

[BL16a] Jinho Baik and Ji Oon Lee. Fluctuations of the free energy of the spherical Sherrington–Kirkpatrick model. Journal of Statistical Physics, 165(2):185–224, October 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1610-0.

Baik:2016:TRS

[BL16b] Jinho Baik and Zhipeng Liu. TASEP on a ring in subrelaxation time scale. Journal of Statistical Physics, 165(6): 1051–1085, December 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1665-y.

Boon:2017:TDM

[BL17] Jean Pierre Boon and James F. Lutsko. Temporal diffusion: From microscopic dynamics to generalised Fokker–Planck and fractional equations. Journal of Statistical Physics, 166(6): 1441–1454, March 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-017-1716-z.

Bakhtin:2018:ZTL

[BL18] Yuri Bakhtin and Liying Li. Zero temperature limit for directed polymers and inviscid limit for stationary solutions of stochastic Burgers equation. Journal of Statistical Physics, 172(5):1358–1397, September 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bercu:2019:MDE

[BL19] Bernard Bercu and Lucile Laulin. On the multi-dimensional elephant random walk. Journal of Statistical Physics, 175(6): 1146–1163, June 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Blank:2010:MPD

[Bla10] Michael Blank. Metric properties of discrete time exclusion type processes in continuum. Journal of Statistical Physics, 140(1):170–197, July 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9983-y.

Bardoscia:2013:SCG

[BLL^{+}13] Marco Bardoscia, Giancarlo De Luca, Giacomo Livan, Matteo Marsili, and Claudio J. Tessone. The social climbing game. Journal of Statistical Physics, 151(3–4):440–457, May 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0693-0.

Benois:2013:HTR

[BLM13] O. Benois, C. Landim, and M. Mourragui. Hitting times of rare events in Markov chains. Journal of Statistical Physics, 153(6):967–990, December 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0875-9.

Biskup:2013:TRC

[BLRVR13] M. Biskup, O. Louidor, A. Rozinov, and A. Vandenberg-Rodes. Trapping in the random conductance model. Journal of Statistical Physics, 150(1):66–87, January 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0688-2.

Barton:2011:PDG

[BLS11] J. Barton, J. L. Lebowitz, and E. R. Speer. Phase diagram of a generalized ABC model on the interval. Journal of Statistical Physics, 145(3):763–784, November 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0279-7.

Boldrighini:2017:CSS

[BLS17] Carlo Boldrighini, Dong Li, and Yakov G. Sinai. Complex singular solutions of the 3-d Navier–Stokes equations and related real solutions. Journal of Statistical Physics, 167(1):1– 13, April 2017. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-017-1730-1.

Bassetti:2011:KMR

[BLT11] Federico Bassetti, Lucia Ladelli, and Giuseppe Toscani. Kinetic models with randomly perturbed binary collisions. Journal of Statistical Physics, 142(4):686–709, February 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0136-8.

Bodineau:2012:FSS

[BLT12] Thierry Bodineau, Vivien Lecomte, and Cristina Toninelli. Finite size scaling of the dynamical free-energy in a kinetically constrained model. Journal of Statistical Physics, 147(1):1– 17, April 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0458-1.

Benassi:2016:CIQ

[BLU16] Costanza Benassi, Benjamin Lees, and Daniel Ueltschi. Correlation inequalities for the quantum XY model. Journal of Statistical Physics, 164(5):1157–1166, September 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-016-1580-2.pdf.

Bonetto:2014:KMC

[BLV14] Federico Bonetto, Michael Loss, and Ranjini Vaidyanathan. The Kac model coupled to a thermostat. Journal of Statistical Physics, 156(4):647–667, August 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0999-6.

Baik:2018:FPT

[BLW18] Jinho Baik, Ji Oon Lee, and Hao Wu. Ferromagnetic to paramagnetic transition in spherical spin glass. Journal of Statistical Physics, 173(5):1484–1522, December 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bouchet:2014:LDL

[BLZ14] Freddy Bouchet, Jason Laurie, and Oleg Zaboronski. Langevin dynamics, large deviations and instantons for the quasigeostrophic model and two-dimensional Euler equations. Journal of Statistical Physics, 156(6):1066–1092, September 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1052-5.

Benfatto:2011:DWN

[BM11] G. Benfatto and V. Mastropietro. Drude weight in non solvable quantum spin chains. Journal of Statistical Physics, 143(2):251–260, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0181-3.

Braga:2012:RGA

[BM12a] Gastão A. Braga and Jussara M. Moreira. Renormalization group analysis of nonlinear diffusion equations with time dependent coefficients and marginal perturbations. Journal of Statistical Physics, 148(2):280–295, August 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0539-1.

Brydges:2012:CCF

[BM12b] David C. Brydges and P. K. Mitter. On the convergence to the continuum of finite range lattice covariances. Journal of Statistical Physics, 147(4):716–727, June 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0492-z.

Butta:2016:DIC

[BM16] Paolo Buttà and Carlo Marchioro. Dynamics of infinite classical anharmonic crystals. Journal of Statistical Physics, 164 (3):680–692, August 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1540-x.

Broman:2017:PTU

[BM17] Erik Broman and Ronald Meester. Phase transition and uniqueness of levelset percolation. Journal of Statistical Physics, 167(6):1376–1400, June 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s10955-017-
1782-2.pdf.

Balint:2018:SPF

[BM18] Péter Bálint and Ian Melbourne. Statistical properties for flows with unbounded roof function, including the Lorenz attractor. Journal of Statistical Physics, 172(4):1101–1126, August 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Boyd:2017:LEC

[BMC17] Alexander B. Boyd, Dibyendu Mandal, and James P. Crutchfield. Leveraging environmental correlations: The thermodynamics of requisite variety. Journal of Statistical Physics, 167 (6):1555–1585, June 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bonachela:2012:PDN

[BML12] Juan A. Bonachela, Miguel A. Muñoz, and Simon A. Levin. Patchiness and demographic noise in three ecological examples. Journal of Statistical Physics, 148(4):724–740, September 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0506-x.

Baldasso:2017:EPS

[BMNS17] Rangel Baldasso, Otávio Menezes, Adriana Neumann, and Rafael R. Souza. Exclusion process with slow boundary. Journal of Statistical Physics, 167(5):1112–1142, June 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bosco:2010:ERC

[BMR10] G. G. Bosco, F. P. Machado, and Thomas Logan Ritchie. Exponential rates of convergence in the ergodic theorem: A constructive approach. Journal of Statistical Physics, 139(3):367– 374, May 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-9945-4.

Bufetov:2013:ESM

[BMSS13] Alexander Bufetov, Sevak Mkrtchyan, Maria Shcherbina, and Alexander Soshnikov. Entropy and the Shannon– McMillan–Breiman theorem for beta random matrix ensembles. Journal of Statistical Physics, 152(1):1–14, July 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0761-5.

Biamonte:2015:TNC

[BMT15] Jacob D. Biamonte, Jason Morton, and Jacob Turner. Tensor network contractions for #SAT. Journal of Statistical Physics, 160(5):1389–1404, September 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1276-z.

Bachmann:2014:GPC

[BN14] Sven Bachmann and Bruno Nachtergaele. On gapped phases with a continuous symmetry and boundary operators. Journal of Statistical Physics, 154(1–2):91–112, January 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0850-5.

Boumaza:2015:LTC

[BN15] Hakim Boumaza and Hatem Najar. Lifshitz tails for continuous matrix-valued Anderson models. Journal of Statistical Physics, 160(2):371–396, July 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1255-4.

Barletti:2018:QTC

[BN18] Luigi Barletti and Claudia Negulescu. Quantum transmission conditions for diffusive transport in graphene with steep potentials. Journal of Statistical Physics, 171(4):696–726, May 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ben-Naim:2013:RC

[BNHRV13] E. Ben-Naim, N. W. Hengartner, S. Redner, and F. Vazquez. Randomness in competitions. Journal of Statistical Physics, 151(3–4):458–474, May 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0648-x.

Basile:2014:DLT

[BNP14] G. Basile, A. Nota, and M. Pulvirenti. A diffusion limit for a test particle in a random distribution of scatterers. Journal of Statistical Physics, 155(6):1087–1111, June 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-0940-z.

Bouchet:2013:KTJ

[BNT13] Freddy Bouchet, Cesare Nardini, and Tomás Tangarife. Kinetic theory of jet dynamics in the stochastic barotropic and 2D Navier–Stokes equations. Journal of Statistical Physics, 153(4):572–625, November 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0828-3.

Balazs:2016:CSR

[BNTT16] Márton Balázs, Attila László Nagy, Bálint Tóth, and István Tóth. Coexistence of shocks and rarefaction fans: Complex phase diagram of a simple hyperbolic particle system. Journal of Statistical Physics, 165(1):115–125, October 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1600-2.

Bonachela:2011:UBC

[BNXL11] Juan A. Bonachela, Carey D. Nadell, João B. Xavier, and Simon A. Levin. Universality in bacterial colonies. Journal of Statistical Physics, 144(2):303–315, July 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0179-x.

Bishop:2016:SGE

[BNY16] Michael Bishop, Bruno Nachtergaele, and Amanda Young. Spectral gap and edge excitations of d-dimensional PVBS models on half–spaces. Journal of Statistical Physics, 162 (6):1485–1521, March 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1457-4.

Bernardin:2011:TPC

[BO11] Cédric Bernardin and Stefano Olla. Transport properties of a chain of anharmonic oscillators with random flip of velocities. Journal of Statistical Physics, 145(5):1224–1255, December 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0385-6.

Basile:2014:EDH

[BO14] Giada Basile and Stefano Olla. Energy diffusion in harmonic system with conservative noise. Journal of Statistical Physics, 155(6):1126–1142, June 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0908-4.

Beghin:2016:PPS

[BO16] Luisa Beghin and Enzo Orsingher. Population processes sampled at random times. Journal of Statistical Physics, 163(1): 1–21, April 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1475-2.

Bonanno:2015:CAS

[Bon15] Claudio Bonanno. A complexity approach to the soliton resolution conjecture. Journal of Statistical Physics, 160(5): 1432–1448, September 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1297-7.

Bothner:2018:SNS

[Bot18] Thomas Bothner. A short note on the scaling function constant problem in the two-dimensional Ising model. Journal of Statistical Physics, 170(4):672–683, February 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bouchaud:2013:CCS

[Bou13a] Jean-Philippe Bouchaud. Crises and collective socio-economic phenomena: Simple models and challenges. Journal of Statistical Physics, 151(3–4):567–606, May 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0687-3.

Bourgain:2013:LBL

[Bou13b] J. Bourgain. A lower bound for the Lyapounov exponents of the random Schrödinger operator on a strip. Journal of Statistical Physics, 153(1):1–9, October 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0821-x.

Bourgain:2018:HP

[Bou18] J. Bourgain. On a homogenization problem. Journal of Statistical Physics, 172(2):314–320, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Benettin:2011:TSE

[BP11a] G. Benettin and A. Ponno. Time-scales to equipartition in the Fermi–Pasta–Ulam problem: Finite-size effects and thermodynamic limit. Journal of Statistical Physics, 144(4):793– 812, August 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0277-9.

Bievre:2011:EGE

[BP11b] S. De Bièvre and P. E. Parris. Equilibration, generalized equipartition, and diffusion in dynamical Lorentz gases. Journal of Statistical Physics, 142(2):356–385, January 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-0109-3.

Bai:2012:LBE

[BP12a] Z. D. Bai and G. M. Pan. Limiting behavior of eigenvectors of large Wigner matrices. Journal of Statistical Physics, 146(3):519–549, February 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0407-4.

Braides:2012:VPP

[BP12b] Andrea Braides and Andrey Piatnitski. Variational problems with percolation: Dilute spin systems at zero temperature. Journal of Statistical Physics, 149(5):846–864, November 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0628-1.

Burghoff:2015:SAD

[BP15] Toralf Burghoff and Ilya Pavlyukevich. Spectral analysis for a discrete metastable system driven by Lévy flights. Journal of Statistical Physics, 161(1):171–196, October 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1313-y.

Boers:2016:MFL

[BP16] Niklas Boers and Peter Pickl. On mean field limits for dynamical systems. Journal of Statistical Physics, 164(1):1–16, July 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1351-5.

Barsuk:2018:BSA

[BP18] Alexandr A. Barsuk and Florentin Paladi. Bifurcation and stability analysis of the equilibrium states in thermodynamic systems in a small vicinity of the equilibrium values of parameters. Journal of Statistical Physics, 171(2):361–381, April 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bobylev:2019:LWA

[BP19] A. V. Bobylev and I. F. Potapenko. Long wave asymptotics for the Vlasov–Poisson–Landau kinetic equation. Journal of Statistical Physics, 175(1):1–18, April 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Binder:2010:PSC

[BPDH10] Kurt Binder, Sanjay Puri, Subir K. Das, and Jürgen Horbach. Phase separation in confined geometries. Journal of Statistical Physics, 138(1–3):51–84, February 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-9924-9.

Borile:2014:TAH

[BPF^{+}14] Claudio Borile, Paolo Dai Pra, Markus Fischer, Marco Formentin, and Amos Maritan. Time to absorption for a heterogeneous neutral competition model. Journal of Statistical Physics, 156(1):119–130, July 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-0989-8.

Benettin:2018:FPU

[BPP18] G. Benettin, S. Pasquali, and A. Ponno. The Fermi–Pasta– Ulam problem and its underlying integrable dynamics: An approach through Lyapunov exponents. Journal of Statistical Physics, 171(4):521–542, May 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Biondo:2013:BRR

[BPR13] Alessio Emanuele Biondo, Alessandro Pluchino, and Andrea Rapisarda. The beneficial role of random strategies in social and financial systems. Journal of Statistical Physics, 151 (3–4):607–622, May 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0691-2.

Bonomi:2014:PVE

[BPR14] Andrea Bonomi, Eleonora Perversi, and Eugenio Regazzini. Probabilistic view of explosion in an inelastic Kac model. Journal of Statistical Physics, 154(5):1292–1324, March 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-0921-2.

Brezin:2014:CRB

[BPRT14] E. Brezin, G. Parisi, and F. Ricci-Tersenghi. The crossover region between long-range and short-range interactions for the critical exponents. Journal of Statistical Physics, 157 (4–5):855–868, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1081-0.

Barany:2012:SMP

[BPS12] B. Bárány, M. Pollicott, and K. Simon. Stationary measures for projective transformations: The Blackwell and Furstenberg measures. Journal of Statistical Physics, 148(3):393–421, August 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0541-7.

Bao:2013:CLT

[BPZ13] Zhigang Bao, Guangming Pan, and Wang Zhou. Central limit theorem for partial linear eigenvalue statistics of Wigner matrices. Journal of Statistical Physics, 150(1):88–129, January 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0663-y.

Bachmann:2010:AMS

[BR10] S. Bachmann and W. De Roeck. From the Anderson model on a strip to the DMPK equation and random matrix theory. Journal of Statistical Physics, 139(4):541–564, May 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-010-9947-2.pdf.

Barrat:2011:PIQ

[BR11] Jean-Louis Barrat and David Rodney. Portable implementation of a quantum thermal bath for molecular dynamics simulations. Journal of Statistical Physics, 144(3):679–689, August 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0193-z.

Boivin:2013:EHM

[BR13] Daniel Boivin and Clément Rau. Existence of the harmonic measure for random walks on graphs and in random environments. Journal of Statistical Physics, 150(2):235–263, January 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0685-5.

Bhatnagar:2016:STS

[BR16] Nayantara Bhatnagar and Dana Randall. Simulated tempering and swapping on mean-field models. Journal of Statistical Physics, 164(3):495–530, August 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1526-8.

Braides:2014:ENE

[Bra14] Andrea Braides. An example of non-existence of plane-like minimizers for an almost-periodic Ising system. Journal of Statistical Physics, 157(2):295–302, October 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1069-9.

Brezin:2010:ISF

[Bré10] Edouard Brézin. Introduction to statistical field theory. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-19303-6 (hardcover). x + 166 pp. LCCN QC175.16.P5

B74 2010. URL http://assets.cambridge.org/97805211/
93030/cover/9780521193030.jpg.

Brezin:2014:WRG

[Bré14] E. Brézin. Wilson’s renormalization group: A paradigmatic shift. Journal of Statistical Physics, 157(4–5):644–650, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-0971-5.

Brinkman:2019:HLW

[Bri19] W. F. Brinkman. How lucky we were: in remembrance of Pierre Hohenberg. Journal of Statistical Physics, 175 (3–4):518–520, May 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-018-2171-1.
pdf.

Baccelli:2018:MQN

[BRSV18] F. Baccelli, A. Rybko, S. Shlosman, and A. Vladimirov. Metastability of queuing networks with mobile servers. Journal of Statistical Physics, 173(3–4):1227–1251, November 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bertozzi:2015:CSO

[BRSW15] Andrea L. Bertozzi, Jesus Rosado, Martin B. Short, and Li Wang. Contagion shocks in one dimension. Journal of Statistical Physics, 158(3):647–664, February 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1019-6.

Bruneau:2014:MPO

[Bru14] Laurent Bruneau. Mixing properties of the one-atom maser. Journal of Statistical Physics, 155(5):888–908, June 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-0982-2.

Bapst:2011:LTB

[BS11] Victor Bapst and Guilhem Semerjian. Lifshitz tails on the Bethe lattice: A combinatorial approach. Journal of Statistical Physics, 145(1):51–92, October 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0319-3.

Belitsky:2013:MSS

[BS13a] V. Belitsky and G. M. Schütz. Microscopic structure of shocks and antishocks in the ASEP conditioned on low current. Journal of Statistical Physics, 152(1):93–111, July 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0758-0.

Bernard:2013:OES

[BS13b] Étienne Bernard and Francesco Salvarani. Optimal estimate of the spectral gap for the degenerate Goldstein–Taylor model. Journal of Statistical Physics, 153(2):363–375, October 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0825-6. See correction [BS20].

Belitsky:2015:QAS

[BS15a] V. Belitsky and G. M. Schütz. Quantum algebra symmetry of the ASEP with second-class particles. Journal of Statistical Physics, 161(4):821–842, November 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1363-1.

Bjornberg:2015:SPR

[BS15b] Jakob E. Björnberg and Sigurdur Örn Stefánsson. On site percolation in random quadrangulations of the halfplane. Journal of Statistical Physics, 160(2):336–356, July 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1256-3.

Bjornberg:2015:RWR

[BS15c] Jakob E. Björnberg and Sigurdur Örn Stefánsson. Random walk on random infinite looptrees. Journal of Statistical Physics, 158(6):1234–1261, March 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-014-1174-
9.

Brydges:2015:RGMa

[BS15d] David C. Brydges and Gordon Slade. A renormalisation group method. I. Gaussian integration and normed algebras. Journal of Statistical Physics, 159(3):421–460, May 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1163-z.

Brydges:2015:RGMb

[BS15e] David C. Brydges and Gordon Slade. A renormalisation group method. II. approximation by local polynomials. Journal of Statistical Physics, 159(3):461–491, May 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1164-y.

Brydges:2015:RGMc

[BS15f] David C. Brydges and Gordon Slade. A renormalisation group method. IV. stability analysis. Journal of Statistical Physics, 159(3):530–588, May 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1166-9.

Brydges:2015:RGMd

[BS15g] David C. Brydges and Gordon Slade. A renormalisation group method. V. A single renormalisation group step. Journal of Statistical Physics, 159(3):589–667, May 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1167-8.

Bhattacharya:2016:HTA

[BS16] Bhaswar B. Bhattacharya and Subhabrata Sen. High temperature asymptotics of orthogonal mean-field spin glasses. Journal of Statistical Physics, 162(1):63–80, January 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1406-7.

Buckley:2017:FIA

[BS17] Jeremiah Buckley and Mikhail Sodin. Fluctuations of the increment of the argument for the Gaussian entire function. Journal of Statistical Physics, 168(2):300–330, July 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bryc:2019:MPA

[BS19] Włodzimierz Bryc and Marcin ’Swieca. On matrix product ansatz for asymmetric simple exclusion process with open boundary in the singular case. Journal of Statistical Physics, 177(2):252–284, October 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bernard:2020:COE

[BS20] Étienne Bernard and Francesco Salvarani. Correction to: Optimal Estimate of the Spectral Gap for the Degenerate Goldstein–Taylor Model. Journal of Statistical Physics, 181 (4):1470–1471, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-020-02631-

y.pdf. See [BS13b].

Byshkin:2016:APM

[BSM^{+}16] Maksym Byshkin, Alex Stivala, Antonietta Mira, Rolf Krause, Garry Robins, and Alessandro Lomi. Auxiliary parameter MCMC for exponential random graph models. Journal of Statistical Physics, 165(4):740–754, November 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1650-5.

Buchholz:2014:MCL

[BSS14] Simon Buchholz, Chiara Saffirio, and Benjamin Schlein. Multivariate central limit theorem in quantum dynamics. Journal of Statistical Physics, 154(1–2):113–152, January 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0897-3.

Bates:2019:RSB

[BSS19] Erik Bates, Leila Sloman, and Youngtak Sohn. Replica symmetry breaking in multi-species Sherrington–Kirkpatrick model. Journal of Statistical Physics, 174(2):333–350, January 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bessa:2018:SGP

[BSV18] Mário Bessa, César M. Silva, and Helder Vilarinho. Stretching generic Pesin’s entropy formula. Journal of Statistical Physics, 173(5):1523–1546, December 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bauerschmidt:2017:FDW

[BSW17] Roland Bauerschmidt, Gordon Slade, and Benjamin C. Wallace. Four-dimensional weakly self-avoiding walk with contact self-attraction. Journal of Statistical Physics, 167(2):317– 350, April 2017. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-017-1754-6.

Batakis:2011:MAI

[BT11] Athanasios Batakis and Benoît Testud. Multifractal analysis of inhomogeneous Bernoulli products. Journal of Statistical Physics, 142(5):1105–1120, March 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0147-5.

Barbier:2012:FIS

[BT12] Matthieu Barbier and Emmanuel Trizac. Field induced stationary state for an accelerated tracer in a bath. Journal of Statistical Physics, 149(2):317–341, October 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0591-x.

Barbier:2014:POR

[BT14] Matthieu Barbier and Emmanuel Trizac. Polydispersity and optimal relaxation in the hard sphere fluid. Journal of Statistical Physics, 154(5):1365–1390, March 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0929-7.

Beringer:2019:CMR

[BT19] Dorottya Beringer and Ádám Timár. Controllability, matching ratio and graph convergence. Journal of Statistical Physics, 174(5):1080–1103, March 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s10955-019-
02225-3.pdf.

Beuman:2014:GDW

[BTV14] Thomas H. Beuman, Ari M. Turner, and Vincenzo Vitelli. Geometrical detection of weak non–Gaussianity upon coarse– graining. Journal of Statistical Physics, 157(3):571–581, November 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1088-6.

Bjornberg:2018:CTH

[BU18] Jakob E. Björnberg and Daniel Ueltschi. Critical temperature of Heisenberg models on regular trees, via random loops. Journal of Statistical Physics, 173(5):1369–1385, December 2018. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-018-2154-2.pdf.

Bucciantini:2016:SSA

[Buc16] Leda Bucciantini. Stationary state after a quench to the Lieb– Liniger from rotating BECs. Journal of Statistical Physics, 164(3):621–644, August 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-016-1535-7.
pdf.

Bunimovich:2014:FSS

[Bun14] Leonid A. Bunimovich. Fine structure of sticky sets in mushroom billiards. Journal of Statistical Physics, 154(1–2):421– 431, January 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0898-2.

Burkhardt:2011:HCS

[Bur11] Theodore W. Burkhardt. Harmonically confined, semiflexible polymer in a channel: Response to a stretching force and spatial distribution of the endpoints. Journal of Statistical Physics, 145(6):1472–1484, December 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0360-2.

Burkhardt:2017:OTR

[Bur17] Theodore W. Burkhardt. Occupation time of a randomly accelerated particle on the positive half axis: Results for the first five moments. Journal of Statistical Physics, 169(4):730–743, November 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Burkhardt:2019:TPS

[Bur19] Theodore W. Burkhardt. Tagged-particle statistics in singlefile motion with random-acceleration and Langevin dynamics. Journal of Statistical Physics, 177(5):806–824, December 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Businger:2018:SRS

[Bus18] Silvia Businger. The shark random swim. Journal of Statistical Physics, 172(3):701–717, August 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Borisov:2011:LLS

[BV11] Denis Borisov and Ivan Veselić. Low lying spectrum of weakdisorder quantum waveguides. Journal of Statistical Physics, 142(1):58–77, January 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0099-1.

Bricmont:2015:IMB

[BV15] Jean Bricmont and Hanne Van Den Bosch. Intermediate model between majority voter PCA and its mean field model. Journal of Statistical Physics, 158(5):1090–1099, March 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1037-4.

Bernhoff:2016:DVM

[BV16] Niclas Bernhoff and Mirela Vinerean. Discrete velocity models for mixtures without nonphysical collision invariants. Journal of Statistical Physics, 165(2):434–453, October 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-016-1624-7.pdf.

Baroni:2019:TFW

[BvdHK19] Enrico Baroni, Remco van der Hofstad, and Júlia Komjáthy. Tight fluctuations of weight-distances in random graphs with infinite-variance degrees. Journal of Statistical Physics, 174 (4):906–934, February 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-018-2213-8.
pdf.

Baake:2011:CPD

[BvE11] Michael Baake and Aernout van Enter. Close-packed dimers on the line: Diffraction versus dynamical spectrum. Journal of Statistical Physics, 143(1):88–101, April 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-011-0163-5.pdf.

Bargigli:2016:STW

[BVL16] Leonardo Bargigli, Stefano Viaggiu, and Andrea Lionetto. A statistical test of Walrasian equilibrium by means of complex networks theory. Journal of Statistical Physics, 165(2):351– 370, October 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1599-4.

Baake:2010:PDS

[BW10] Michael Baake and Tom Ward. Planar dynamical systems with pure Lebesgue diffraction spectrum. Journal of Statistical Physics, 140(1):90–102, July 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9984-x.

Barat:2012:CDP

[BW12a] János Barát and Ian M. Wanless. A cube dismantling problem related to bootstrap percolation. Journal of Statistical Physics, 149(4):754–770, November 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0622-7.

Bishop:2012:GSE

[BW12b] Michael Bishop and Jan Wehr. Ground state energy of the one-dimensional discrete random Schrödinger operator with Bernoulli potential. Journal of Statistical Physics, 147(3):529– 541, May 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0480-3.

Bryc:2017:ASE

[BW17] Wlodek Bryc and Jacek Wesolowski. Asymmetric simple exclusion process with open boundaries and quadratic harnesses. Journal of Statistical Physics, 167(2):383–415, April 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-017-1747-5.

Baake:2018:LDU

[BW18] Ellen Baake and Anton Wakolbinger. Lines of descent under selection. Journal of Statistical Physics, 172(1):156–174, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bae:2019:PMH

[ByChL^{+}19] Hyeong-Ohk Bae, Seung yeon Cho, Sang hyeok Lee, Jane Yoo, and Seok-Bae Yun. A particle model for the herding phenomena induced by dynamic market signals. Journal of Statistical Physics, 177(2):365–398, October 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bourgade:2019:RBM

[BYYY19] P. Bourgade, F. Yang, H.-T. Yau, and J. Yin. Random band matrices in the delocalized phase, II: Generalized resolvent estimates. Journal of Statistical Physics, 174(6):1189–1221, March 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bertacchi:2013:RPR

[BZ13] Daniela Bertacchi and Fabio Zucca. Rumor processes in random environment on N and on Galton–Watson trees. Journal of Statistical Physics, 153(3):486–511, November 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0843-4.

Canova:2018:RIA

[CA18] Gabriel A. Canova and Jeferson J. Arenzon. Risk and interaction aversion: Screening mechanisms in the Prisoner’s Dilemma Game. Journal of Statistical Physics, 172(1):279– 292, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Cajahuaringa:2019:NFE

[CA19] Samuel Cajahuaringa and Alex Antonelli. Nonequilibrium free energy methods applied to magnetic systems: The degenerate Ising model. Journal of Statistical Physics, 175(5):1006–1021, June 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Caceres:2014:PTS

[Các14] Manuel O. Cáceres. Passage time statistics in exponential distributed time-delay models: Noisy asymptotic dynamics. Journal of Statistical Physics, 156(1):94–118, July 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-0993-z.

Caer:2010:PRW

[Caë10] Gérard Le Caër. A Pearson random walk with steps of uniform orientation and Dirichlet distributed lengths. Journal of Statistical Physics, 140(4):728–751, August 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0015-8.

Caer:2011:NFS

[Caë11] Gérard Le Caër. A new family of solvable Pearson–Dirichlet random walks. Journal of Statistical Physics, 144(1):23– 45, July 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0245-4.

Christoforou:2013:CCC

[CAG^{+}13] Evgenia Christoforou, Antonio Fernández Anta, Chryssis Georgiou, Miguel A. Mosteiro, and Angel Sánchez. Crowd computing as a cooperation problem: An evolutionary approach. Journal of Statistical Physics, 151(3–4):654–672, May 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0661-0.

Calder:2015:DLP

[Cal15] Jeff Calder. Directed last passage percolation with discontinuous weights. Journal of Statistical Physics, 158(4):903–949, February 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1146-0.

Cameron:2013:CFC

[Cam13] M. K. Cameron. Computing Freidlin’s cycles for the overdamped Langevin dynamics. Application to the Lennard-Jones-38 cluster. Journal of Statistical Physics, 152(3):493– 518, August 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0770-4.

Can:2017:CBA

[Can17] Van Hao Can. Critical behavior of the annealed Ising model on random regular graphs. Journal of Statistical Physics, 169(3): 480–503, November 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Carstens:2011:NCI

[Car11] Sebastian Carstens. Non-coexistence of infinite clusters in two-dimensional dependent site percolation. Journal of Statistical Physics, 144(6):1223–1237, September 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0329-1.

Cinkir:2011:RCA

[CAS11] Zubeyir Cinkir, Hasan Akin, and Irfan Siap. Reversibility of 1D cellular automata with periodic boundary over finite fields Z_{p}. Journal of Statistical Physics, 143(4):807–823, May 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0202-2.

Castellana:2014:RRH

[Cas14] Michele Castellana. Rigorous results for hierarchical models of structural glasses. Journal of Statistical Physics, 157(2):219– 233, October 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1085-9.

Castillo:2016:GAD

[CB16] Isaac Pérez Castillo and Denis Boyer. A generalised Airy distribution function for the accumulated area swept by N vicious Brownian paths. Journal of Statistical Physics, 162 (6):1587–1607, March 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1467-2.

Castellana:2014:FEB

[CBG14] Michele Castellana, Adriano Barra, and Francesco Guerra. Free-energy bounds for hierarchical spin models. Journal of Statistical Physics, 155(2):211–222, April 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0951-9.

Chang:2014:ABV

[CC14] Shu-Chiuan Chang and Lung-Chi Chen. Asymptotic behavior for a version of directed percolation on the triangular lattice. Journal of Statistical Physics, 155(3):500–522, May 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-0961-7.

Campbell:2018:SHI

[CC18] Michael Campbell and David Carfì. Speculative and hedging interaction model in oil and U.S. dollar markets-phase transition. Journal of Statistical Physics, 170(1):165–205, January 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Chanphana:2019:GSP

[CC19] R. Chanphana and P. Chatraphorn. Generalized survival probabilities in height fluctuations of limited mobility growth models with and without up–down symmetry. Journal of Statistical Physics, 176(4):932–945, August 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Chanphana:2015:EPS

[CCD15] R. Chanphana, P. Chatraphorn, and C. Dasgupta. Effects of patterned substrate on thin films simulated by family model. Journal of Statistical Physics, 160(2):397–408, July 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1251-8.

Calvo:2010:GCL

[CCEF10] Iván Calvo, Juan C. Cuchí, José G. Esteve, and Fernando Falceto. Generalized central limit theorem and renormalization group. Journal of Statistical Physics, 141(3):409–421, November 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-0065-y.

Csaki:2018:TDA

[CCFR18] Endre Csáki, Miklós Csörgő, Antónia Földes, and Pál Révész. Two-dimensional anisotropic random walks: Fixed versus random column configurations for transport phenomena. Journal of Statistical Physics, 171(5):822–841, June 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Cao:2014:DCT

[CCG14a] Jiarui Cao, Paul Chleboun, and Stefan Großkinsky. Dynamics of condensation in the totally asymmetric inclusion process. Journal of Statistical Physics, 155(3):523–543, May 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-0966-2.

Cao:2014:CDP

[CCG14b] Man Cao, Yang Chen, and James Griffin. Continuous and discrete Painlevé equations arising from the gap probability distribution of the finite n Gaussian unitary ensembles. Journal of Statistical Physics, 157(2):363–375, October 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1076-x.

Chernyak:2010:NES

[CCGT10] Vladimir Y. Chernyak, Michael Chertkov, David A. Goldberg, and Konstantin Turitsyn. Non-equilibrium statistical physics of currents in queuing networks. Journal of Statistical Physics, 140(5):819–845, September 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0018-5.

Carrillo:2014:CTD

[CCH^{+}14] José A. Carrillo, Young-Pil Choi, Seung-Yeal Ha, Moon-Jin Kang, and Yongduck Kim. Contractivity of transport distances for the kinetic Kuramoto equation. Journal of Statistical Physics, 156(2):395–415, July 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1005-z.

Chiarini:2016:ESG

[CCH16] Alberto Chiarini, Alessandra Cipriani, and Rajat Subhra Hazra. Extremes of some Gaussian random interfaces. Journal of Statistical Physics, 165(3):521–544, November 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1634-5.

Costin:2019:IOF

[CCL19] Ovidiu Costin, Rodica D. Costin, and Joel L. Lebowitz. Ionization by an oscillating field: Resonances and photons. Journal of Statistical Physics, 175(3–4):681–689, May 2019. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Caprino:2016:TEI

[CCM16] Silvia Caprino, Guido Cavallaro, and Carlo Marchioro. Time evolution of an infinitely extended Vlasov system with singular mutual interaction. Journal of Statistical Physics, 162(2):426– 456, January 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1403-x.

Caprino:2017:MSV

[CCM17] Silvia Caprino, Guido Cavallaro, and Carlo Marchioro. On the magnetic shield for a Vlasov–Poisson plasma. Journal of Statistical Physics, 169(6):1066–1097, December 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Cassandro:2016:HAS

[CCP16] M. Cassandro, M. Colangeli, and E. Presutti. Highly anisotropic scaling limits. Journal of Statistical Physics, 162 (4):997–1030, February 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1437-0.

Chazottes:2017:CIT

[CCR17] J.-R. Chazottes, P. Collet, and F. Redig. On concentration inequalities and their applications for Gibbs measures in lattice systems. Journal of Statistical Physics, 169(3):504–546, November 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Cirillo:2019:TQM

[CCR19] E. N. M. Cirillo, M. Colangeli, and L. Rondoni. Transport in quantum multi-barrier systems as random walks on a lattice. Journal of Statistical Physics, 176(3):692–709, August 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Castillo:2014:RPO

[CD14a] Isaac Pérez Castillo and Thomas Dupic. Reunion probabilities of N one-dimensional random walkers with mixed boundary conditions. Journal of Statistical Physics, 156(3):606–616, August 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1017-8.

Clark:2014:BPM

[CD14b] Jeremy Thane Clark and Loïc Dubois. A Brownian particle in a microscopic periodic potential. Journal of Statistical Physics, 155(2):323–391, April 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-0949-3.

Chen:2016:TPS

[CD16] Hanlin Chen and Hanyuan Deng. Tutte polynomial of scalefree networks. Journal of Statistical Physics, 163(4):714– 732, May 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1465-4.

Chen:2018:GOP

[CDCL18] Xiangping Chen, Haiming Duan, Biaobing Cao, and Mengqiu Long. The global optimization of Pt_{13} cluster using the first-principle molecular dynamics with the quenching technique. Journal of Statistical Physics, 171(3):427–433, May 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Cioletti:2014:DBT

[CDdS14] L. M. Cioletti, C. C. Y. Dorea, and S. Vasconcelos da Silva. Diffusive–ballistic transition in random polymers with drift and repulsive long-range interactions. Journal of Statistical Physics, 156(4):760–765, August 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-014-1021-
z.

Coon:2012:FCC

[CDG12] Justin Coon, Carl P. Dettmann, and Orestis Georgiou. Full connectivity: Corners, edges and faces. Journal of Statistical Physics, 147(4):758–778, June 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0493-y.

Cavagna:2015:FTN

[CDG^{+}15] Andrea Cavagna, Lorenzo Del Castello, Irene Giardina, Tomas Grigera, Asja Jelic, Stefania Melillo, Thierry Mora, Leonardo Parisi, Edmondo Silvestri, Massimiliano Viale, and Aleksandra M. Walczak. Flocking and turning: a new model for selforganized collective motion. Journal of Statistical Physics, 158(3):601–627, February 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1119-3.

Collet:2015:WTR

[CDH15] P. Collet, F. Dunlop, and T. Huillet. Wetting transitions for a random line in long–range potential. Journal of Statistical Physics, 160(6):1545–1622, September 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1296-8.

Cuff:2012:GDM

[CDL^{+}12] P. Cuff, J. Ding, O. Louidor, E. Lubetzky, Y. Peres, and A. Sly. Glauber dynamics for the mean-field Potts model. Journal of Statistical Physics, 149(3):432–477, November 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0599-2.

Calleja:2010:CBA

[CdlL10] Renato Calleja and Rafael de la Llave. Computation of the breakdown of analyticity in statistical mechanics models: Numerical results and a renormalization group explanation. Journal of Statistical Physics, 141(6):940–951, December 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-0085-7.

Couto:2013:TCH

[CdLS13] Rodrigo G. Couto, Bernardo N. B. de Lima, and Rémy Sanchis. Truncated connectivities in a highly supercritical anisotropic percolation model. Journal of Statistical Physics, 153(5):739–750, December 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0864-z.

Colangeli:2017:PMS

[CDP17] M. Colangeli, A. De Masi, and E. Presutti. Particle models with self sustained current. Journal of Statistical Physics, 167 (5):1081–1111, June 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Collet:2010:SMF

[CDS10] Francesca Collet, Paolo Dai Pra, and Elena Sartori. A simple mean field model for social interactions: Dynamics, fluctuations, criticality. Journal of Statistical Physics, 139(5):820– 858, June 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-9964-1.

Ching:2017:FTB

[CDS17] Emily S. C. Ching, On-Yu Dung, and Olga Shishkina. Fluctuating thermal boundary layers and heat transfer in turbulent Rayleigh–Bénard convection. Journal of Statistical Physics, 167(3–4):626–635, May 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Caracciolo:2019:ASO

[CDS19] Sergio Caracciolo, Matteo D’Achille, and Gabriele Sicuro. Anomalous scaling of the optimal cost in the one-dimensional random assignment problem. Journal of Statistical Physics, 174(4):846–864, February 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Collins:2010:SSR

[CDTA10] Benoît Collins, Kenneth J. Dykema, and Francisco Torres-Ayala. Sum-of-squares results for polynomials related to the Bessis–Moussa–Villani conjecture. Journal of Statistical Physics, 139(5):779–799, June 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9959-y.

Cozzi:2017:PIL

[CDV17] Matteo Cozzi, Serena Dipierro, and Enrico Valdinoci. Planelike interfaces in long-range Ising models and connections with nonlocal minimal surfaces. Journal of Statistical Physics, 167 (6):1401–1451, June 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Cao:2012:WLA

[CE12] Zhenwei Cao and Alexander Elgart. The weak localization for the alloy-type Anderson model on a cubic lattice. Journal of Statistical Physics, 148(6):1006–1039, September 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0562-2.

Carnio:2014:APT

[CE14] Brett N. Carnio and Janet A. W. Elliott. Analytical proof that there is no effect of confinement or curvature on the Maxwell– Boltzmann collision frequency. Journal of Statistical Physics, 156(4):668–685, August 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1028-5.

Cortes:2019:BEC

[CE19] E. Cortés and M. Escobedo. On a Boltzmann equation for Compton scattering from non relativistic electrons at low density. Journal of Statistical Physics, 175(5):819–878, June 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Chinellato:2015:DRN

[CEB^{+}15] David D. Chinellato, Irving R. Epstein, Dan Braha, Yaneer Bar-Yam, and Marcus A. M. de Aguiar. Dynamical response of networks under external perturbations: Exact results. Journal of Statistical Physics, 159(2):221–230, April 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1189-x.

Collevecchio:2018:CTH

[CEGW18] Andrea Collevecchio, Eren Metin Elçi, Timothy M. Garoni, and Martin Weigel. On the coupling time of the heat-bath process for the Fortuin–Kasteleyn random-cluster model. Journal of Statistical Physics, 170(1):22–61, January 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Carlen:2018:ASS

[CEL^{+}18] E. A. Carlen, R. Esposito, J. L. Lebowitz, R. Marra, and C. Mouhot. Approach to the steady state in kinetic models with thermal reservoirs at different temperatures. Journal of Statistical Physics, 172(2):522–543, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Centola:2013:SMS

[Cen13] Damon Centola. A simple model of stability in critical mass dynamics. Journal of Statistical Physics, 151(1–2):238–253, April 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0679-3.

Chandra:2019:LSM

[CES19] Ajay Chandra, Dirk Erhard, and Hao Shen. Local solution to the multi-layer KPZ equation. Journal of Statistical Physics, 175(6):1080–1106, June 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Cerrai:2011:SMA

[CF11] Sandra Cerrai and Mark Freidlin. Small mass asymptotics for a charged particle in a magnetic field and long-time influence of small perturbations. Journal of Statistical Physics, 144(1):101–123, July 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0238-3.

Cerrai:2015:LDL

[CF15] Sandra Cerrai and Mark Freidlin. Large deviations for the Langevin equation with strong damping. Journal of Statistical Physics, 161(4):859–875, November 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1346-2.

Chen:2016:FLI

[CF16] Zhen-Qing Chen and Wai-Tong Louis Fan. Fluctuation limit for interacting diffusions with partial annihilations through membranes. Journal of Statistical Physics, 164(4):890–936, August 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1567-z.

Chafai:2019:SCL

[CF19a] Djalil Chafaï and Grégoire Ferré. Simulating Coulomb and log-gases with hybrid Monte Carlo algorithms. Journal of Statistical Physics, 174(3):692–714, February 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Collet:2019:ELF

[CF19b] Francesca Collet and Marco Formentin. Effects of local fields in a dissipative Curie–Weiss model: Bautin bifurcation and large self-sustained oscillations. Journal of Statistical Physics, 176(2):478–491, July 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Corsi:2013:LDI

[CFG13] Livia Corsi, Roberto Feola, and Guido Gentile. Lowerdimensional invariant tori for perturbations of a class of nonconvex Hamiltonian functions. Journal of Statistical Physics, 150(1):156–180, January 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0682-8.

Carles:2013:NLZ

[CFK13] Rémi Carles and Clotilde Fermanian-Kammerer. A nonlinear Landau–Zener formula. Journal of Statistical Physics, 152 (4):619–656, August 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0785-x.

Cardin:2017:IML

[CFL17] Franco Cardin, Marco Favretti, and Alberto Lovison. Inertial manifold and large deviations approach to reduced PDE dynamics. Journal of Statistical Physics, 168(5):1000–1015, September 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Cunden:2019:TOP

[CFLV19] Fabio Deelan Cunden, Paolo Facchi, Marilena Ligabò, and Pierpaolo Vivo. Third-order phase transition: Random matrices and screened Coulomb gas with hard walls. Journal of Statistical Physics, 175(6):1262–1297, June 2019. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-019-02281-9.pdf.

Challenger:2014:TIB

[CFM14] Joseph D. Challenger, Duccio Fanelli, and Alan J. McKane. The theory of individual based discrete-time processes. Journal of Statistical Physics, 156(1):131–155, July 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-0990-2.

Chleboun:2017:MLS

[CFMT17] P. Chleboun, A. Faggionato, F. Martinelli, and C. Toninelli. Mixing length scales of low temperature spin plaquettes models. Journal of Statistical Physics, 169(3):441–471, November 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Comets:2015:LRF

[CFN15] Francis Comets, Ryoki Fukushima, and Shuta Nakajima. Limiting results for the free energy of directed polymers in random environment with unbounded jumps. Journal of Statistical Physics, 161(3):577–597, November 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1347-1.

Corwin:2010:LPT

[CFP10] Ivan Corwin, Patrik L. Ferrari, and Sandrine Péché. Limit processes for TASEP with shocks and rarefaction fans. Journal of Statistical Physics, 140(2):232–267, July 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-9995-7.

Chenn:2018:EDV

[CFS18] Ilias Chenn, G. Fournodavlos, and I. M. Sigal. The effective dynamics of the volume preserving mean curvature flow. Journal of Statistical Physics, 172(2):458–476, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Can:2015:EAF

[CFTW15] T. Can, P. J. Forrester, G. Téllez, and P. Wiegmann. Exact and asymptotic features of the edge density profile for the one component plasma in two dimensions. Journal of Statistical Physics, 158(5):1147–1180, March 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1152-2.

Caglioti:2010:BGL

[CG10a] Emanuele Caglioti and François Golse. On the Boltzmann– grad limit for the two dimensional periodic Lorentz gas. Journal of Statistical Physics, 141(2):264–317, October 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0046-1.

Chleboun:2010:FSE

[CG10b] Paul Chleboun and Stefan Großkinsky. Finite size effects and metastability in zero-range condensation. Journal of Statistical Physics, 140(5):846–872, September 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0017-6.

Campanino:2011:OZB

[CG11a] M. Campanino and M. Gianfelice. On the Ornstein–Zernike behaviour for the Bernoulli bond percolation on ±Z^{d}, d ≥ 3, in the supercritical regime. Journal of Statistical Physics, 145(6):1407–1422, December 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0330-8.

Chetrite:2011:TRV

[CG11b] Raphaël Chetrite and Shamik Gupta. Two refreshing views of fluctuation theorems through kinematics elements and exponential martingale. Journal of Statistical Physics, 143(3):543– 584, May 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0184-0.

Chia:2011:SMH

[CG11c] Nicholas Chia and Nigel Goldenfeld. Statistical mechanics of horizontal gene transfer in evolutionary ecology. Journal of Statistical Physics, 142(6):1287–1301, April 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0112-8.

Collet:2012:CIO

[CG12a] Pierre Collet and Antonio Galves. Chains of infinite order, chains with memory of variable length, and maps of the interval. Journal of Statistical Physics, 149(1):73–85, October 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0579-6.

Correggi:2012:FEQ

[CG12b] M. Correggi and A. Giuliani. The free energy of the quantum Heisenberg ferromagnet at large spin. Journal of Statistical Physics, 149(2):234–245, October 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0589-4.

Costin:2013:OCA

[CG13] Ovidiu Costin and Giambattista Giacomin. Oscillatory critical amplitudes in hierarchical models and the Harris function of branching processes. Journal of Statistical Physics, 150(3):471–486, February 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0609-4.

Cenatiempo:2014:RTT

[CG14a] S. Cenatiempo and A. Giuliani. Renormalization theory of a two dimensional Bose gas: Quantum critical point and quasi-condensed state. Journal of Statistical Physics, 157 (4–5):755–829, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1034-7.

Chleboun:2014:CSP

[CG14b] Paul Chleboun and Stefan Großkinsky. Condensation in stochastic particle systems with stationary product measures. Journal of Statistical Physics, 154(1–2):432–465, January 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0844-3.

Coker:2014:STM

[CG14c] Tom Coker and Karen Gunderson. A sharp threshold for a modified bootstrap percolation with recovery. Journal of Statistical Physics, 157(3):531–570, November 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1099-3.

Campanino:2015:OZB

[CG15] M. Campanino and M. Gianfelice. On the Ornstein–Zernike behaviour for the supercritical random-cluster model on Z^{d}, d ≥ 3. Journal of Statistical Physics, 159(6):1456–1476, June 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1222-0.

Chou:2016:HKT

[CG16] Tom Chou and Chris D. Greenman. A hierarchical kinetic theory of birth, death and fission in age-structured interacting populations. Journal of Statistical Physics, 164(1):49– 76, July 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-016-1524-x.pdf.

Corwin:2017:KPZ

[CG17] Ivan Corwin and Yu Gu. Kardar–Parisi–Zhang equation and large deviations for random walks in weak random environments. Journal of Statistical Physics, 166(1):150–168, January 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1693-7.

Ciech:2019:OVD

[CG19] Federico Ciech and Nicos Georgiou. Order of the variance in the discrete Hammersley process with boundaries. Journal of Statistical Physics, 176(3):591–638, August 2019. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-019-02314-3.pdf.

Contucci:2011:IEE

[CGG^{+}11] Pierluigi Contucci, Cristian Giardinà, Claudio Giberti, Giorgio Parisi, and Cecilia Vernia. Interface energy in the Edwards–Anderson model. Journal of Statistical Physics, 142(1):1–10, January 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0100-z.

Carinci:2013:DSM

[CGGR13] Gioia Carinci, Cristian Giardinà, Claudio Giberti, and Frank Redig. Duality for stochastic models of transport. Journal of Statistical Physics, 152(4):657–697, August 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0786-9.

Collevecchio:2016:WPI

[CGHT16] Andrea Collevecchio, Timothy M. Garoni, Timothy Hyndman, and Daniel Tokarev. The worm process for the Ising model is rapidly mixing. Journal of Statistical Physics, 164 (5):1082–1102, September 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1572-2.

Cassandro:2012:POM

[CGL12] M. Cassandro, A. Galves, and E. Löcherbach. Partially observed Markov random fields are variable neighborhood random fields. Journal of Statistical Physics, 147(4):795–807, June 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0488-8.

Cassandro:2017:ITC

[CGL17] M. Cassandro, A. Galves, and E. Löcherbach. Information transmission and criticality in the contact process. Journal of Statistical Physics, 168(6):1180–1190, September 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). See correction [CGL18].

Cassandro:2018:CIT

[CGL18] M. Cassandro, A. Galves, and E. Löcherbach. Correction to: Information transmission and criticality in the contact process. Journal of Statistical Physics, 170(2):436–437, January 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/
pdf/10.1007/s10955-017-1931-7.pdf. See [CGL17].

Crescenzo:2016:ARD

[CGN16] Antonio Di Crescenzo, Virginia Giorno, and Amelia G. Nobile. Analysis of reflected diffusions via an exponential time-based transformation. Journal of Statistical Physics, 163(6):1425– 1453, June 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1525-9.

Chan:2011:ISS

[CGNP11] Y. Chan, A. J. Guttmann, B. G. Nickel, and J. H. H. Perk. The Ising susceptibility scaling function. Journal of Statistical Physics, 145(3):549–590, November 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0212-0.

Chleboun:2017:LCL

[CGP17] Paul Chleboun, Stefan Grosskinsky, and Andrea Pizzoferrato. Lower current large deviations for zero-range processes on a ring. Journal of Statistical Physics, 167(1):64– 89, April 2017. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-017-1740-z.pdf.

Curado:2012:GBD

[CGR12] E. M. F. Curado, J. P. Gazeau, and Ligia M. C. S. Rodrigues. On a generalization of the binomial distribution and its Poisson-like limit. Journal of Statistical Physics, 146 (2):264–280, January 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0383-8.

Carinci:2016:AST

[CGRS16] Gioia Carinci, Cristian Giardinà, Frank Redig, and Tomohiro Sasamoto. Asymmetric stochastic transport models with U_{q}(∫⊓(1, 1)) symmetry. Journal of Statistical Physics, 163 (2):239–279, April 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-016-1473-4.
pdf.

Calef:2015:ENS

[CGS15] Matthew Calef, Whitney Griffiths, and Alexia Schulz. Estimating the number of stable configurations for the generalized Thomson problem. Journal of Statistical Physics, 160(1):239– 253, July 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1245-6.

Chen:2017:MFL

[CGY17] Li Chen, Simone Göttlich, and Qitao Yin. Mean field limit and propagation of chaos for a pedestrian flow model. Journal of Statistical Physics, 166(2):211–229, January 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1679-5.

Costin:2011:GVB

[CH11] O. Costin and M. Huang. Gamow vectors and Borel summability in a class of quantum systems. Journal of Statistical Physics, 144(4):846–871, August 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0276-x.

Choi:2014:GEC

[CH14] Sun-Ho Choi and Seung-Yeal Ha. Global existence of classical solutions to the inelastic Vlasov–Poisson–Boltzmann system. Journal of Statistical Physics, 156(5):948–974, September 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1041-8.

Chatterjee:2012:RCP

[Cha12] Avik P. Chatterjee. A remark concerning percolation thresholds in polydisperse systems of finite-diameter rods. Journal of Statistical Physics, 146(1):244–248, January 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0400-y.

Chair:2014:ERC

[Cha14a] Noureddine Chair. The effective resistance of the Ncycle graph with four nearest neighbors. Journal of Statistical Physics, 154(4):1177–1190, February 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0916-z.

Chatterjee:2014:LBA

[Cha14b] Avik P. Chatterjee. A lattice-based approach to percolation in penetrable sphere systems. Journal of Statistical Physics, 156 (3):586–592, August 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1013-z.

Chatterjee:2015:PTE

[Cha15a] Avik P. Chatterjee. Percolation thresholds and excluded area for penetrable rectangles in two dimensions. Journal of Statistical Physics, 158(1):248–254, January 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1127-3.

Chazelle:2015:AAC

[Cha15b] Bernard Chazelle. An algorithmic approach to collective behavior. Journal of Statistical Physics, 158(3):514–548, February 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1140-6.

Chatterjee:2019:CLT

[Cha19] Sourav Chatterjee. Central Limit Theorem for the free energy of the random field Ising model. Journal of Statistical Physics, 175(1):185–202, April 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Chen:2013:BSM

[Che13] I-Kun Chen. Boundary singularity of moments for the linearized Boltzmann equation. Journal of Statistical Physics, 153(1):93–118, October 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0817-6.

Chekmarev:2014:LTT

[Che14] Sergei F. Chekmarev. Laminar-turbulent transition: The change of the flow state temperature with the Reynolds number. Journal of Statistical Physics, 157(6):1019–1030, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1112-x.

Chevallier:2018:SSS

[Che18] Julien Chevallier. Stimulus sensitivity of a spiking neural network model. Journal of Statistical Physics, 170(4):800–808, February 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Chhita:2012:HFC

[Chh12] Sunil Chhita. The height fluctuations of an off-critical dimer model on the square grid. Journal of Statistical Physics, 148(1):67–88, July 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0529-3.

Cox:2019:MSN

[CHHK19] Alexander M. G. Cox, Simon C. Harris, Emma L. Horton, and Andreas E. Kyprianou. Multi-species neutron transport equation. Journal of Statistical Physics, 176(2):425– 455, July 2019. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-019-02307-2.pdf.

Chen:2015:DCS

[CHHS15] Wei-Kuo Chen, Hsi-Wei Hsieh, Chii-Ruey Hwang, and Yuan-Chung Sheu. Disorder chaos in the spherical mean-field model. Journal of Statistical Physics, 160(2):417–429, July 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1264-3.

Chulaevsky:2016:ESL

[Chu16] Victor Chulaevsky. Exponential scaling limit of the singleparticle Anderson model via adaptive feedback scaling. Journal of Statistical Physics, 162(3):603–614, February 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1438-z.

Chulaevsky:2018:LAR

[Chu18] Victor Chulaevsky. Localization in asymmetric random displacements models with infinite range of interaction. Journal of Statistical Physics, 172(4):1164–1180, August 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ciesla:2017:SPN

[Cie17] Michal Cieśla. Scaling properties of the number of random sequential adsorption iterations needed to generate saturated random packing. Journal of Statistical Physics, 166(1):39–44, January 2017. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-016-1673-y.pdf.

Chari:2014:SKE

[CIM14] S. Siva Nasarayya Chari, Ramarao Inguva, and K. P. N. Murthy. Study of kinetic equation for non-ideal gases using lattice Boltzmann method: Interparticle forces from BBGKY hierarchy. Journal of Statistical Physics, 157(1):113–123, October 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1075-y.

Camia:2018:NED

[CJN18] Federico Camia, Jianping Jiang, and Charles M. Newman. A note on exponential decay in the random field Ising model. Journal of Statistical Physics, 173(2):268–284, October 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Cao:2017:STF

[CJW17] Shaohua Cao, Jianguo Jiang, and Jichun Wu. Solving timefractional advection–dispersion equation by variable weights particle tracking method. Journal of Statistical Physics, 168 (6):1248–1258, September 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Cheng:2010:ZES

[CL10] Wen-Chiao Cheng and Bing Li. Zero entropy systems. Journal of Statistical Physics, 140(5):1006–1021, September 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-0019-4.

Calogero:2013:MSU

[CL13a] F. Calogero and F. Leyvraz. A macroscopic system with undamped periodic compressional oscillations. Journal of Statistical Physics, 151(5):922–937, June 2013. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-013-0741-
9.

Carinci:2013:LDT

[CL13b] Gioia Carinci and Stephan Luckhaus. Langevin dynamics with a tilted periodic potential. Journal of Statistical Physics, 151(5):870–895, June 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0721-0.

Calogero:2014:MBP

[CL14] F. Calogero and F. Leyvraz. Many-body problem with quadratic and/or inversely-quadratic potentials in one- and more-dimensional spaces: Some retrospective remarks. Journal of Statistical Physics, 155(4):658–665, May 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0973-3.

Cioletti:2015:PTO

[CL15] Leandro Cioletti and Artur O. Lopes. Phase transitions in one-dimensional translation invariant systems: A Ruelle operator approach. Journal of Statistical Physics, 159(6):1424– 1455, June 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1202-4.

Chavez:2016:CHL

[CL16] E. Chavez and C. Landim. A correction to the hydrodynamic limit of boundary driven weakly asymmetric exclusion processes in a quasi-static time scale. Journal of Statistical Physics, 163(5):1079–1107, June 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1514-z.

Chu:2018:ABK

[CL18] Weiqi Chu and Xiantao Li. On the asymptotic behavior of the kernel function in the generalized Langevin equation: A one-dimensional lattice model. Journal of Statistical Physics, 170(2):378–398, January 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Cai:2019:SHB

[CL19] Shuzhe Cai and Xuguang Lu. The spatially homogeneous Boltzmann equation for Bose–Einstein particles: Rate of strong convergence to equilibrium. Journal of Statistical Physics, 175(2):289–350, April 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Clark:2013:SDR

[Cla13] Jeremy Thane Clark. Suppressed dispersion for a randomly kicked quantum particle in a Dirac comb. Journal of Statistical Physics, 150(5):940–1015, March 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0708-x.

Clark:2019:HTS

[Cla19] Jeremy Thane Clark. High-temperature scaling limit for directed polymers on a hierarchical lattice with bond disorder. Journal of Statistical Physics, 174(6):1372–1403, March 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Clisby:2010:EIP

[Cli10] Nathan Clisby. Efficient implementation of the pivot algorithm for self-avoiding walks. Journal of Statistical Physics, 140(2):349–392, July 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9994-8.

Clisby:2018:MCS

[Cli18] Nathan Clisby. Monte Carlo study of four-dimensional selfavoiding walks of up to one billion steps. Journal of Statistical Physics, 172(2):477–492, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Cheng:2018:MDL

[CLL18] Lingyan Cheng, Ruinan Li, and Wei Liu. Moderate deviations for the Langevin equation with strong damping. Journal of Statistical Physics, 170(5):845–861, March 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Cao:2019:EDR

[CLL19] Yu Cao, Jianfeng Lu, and Yulong Lu. Exponential decay of Rényi divergence under Fokker–Planck equations. Journal of Statistical Physics, 176(5):1172–1184, September 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Casini:2015:SHR

[CLM15] Emanuele Casini, Gérard Le Caër, and Andrea Martinelli. Short hyperuniform random walks. Journal of Statistical Physics, 160(1):254–273, July 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1244-7.

Chen:2018:CEA

[CLMK18] Chen Chen, Like Li, Renwei Mei, and James F. Klausner. Chapman–Enskog analyses on the Gray lattice Boltzmann equation method for fluid flow in porous media. Journal of Statistical Physics, 171(3):493–520, May 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Climenhaga:2017:GAC

[CLP17] Vaughn Climenhaga, Stefano Luzzatto, and Yakov Pesin. The geometric approach for constructing Sinai–Ruelle–Bowen measures. Journal of Statistical Physics, 166(3–4):467–493, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1608-7.

Corberi:2019:ODP

[CLP19] Federico Corberi, Eugenio Lippiello, and Paolo Politi. One dimensional phase-ordering in the Ising model with space decaying interactions. Journal of Statistical Physics, 176(3):510– 540, August 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Caflisch:2011:AAS

[CLS11a] Russel E. Caflisch, Maria Carmela Lombardo, and Marco Sammartino. Asymptotic analysis of a slightly rarefied gas with nonlocal boundary conditions. Journal of Statistical Physics, 143(4):725–739, May 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0204-0.

Chen:2011:RCT

[CLS11b] Li Chen, Ji Oon Lee, and Benjamin Schlein. Rate of convergence towards Hartree dynamics. Journal of Statistical Physics, 144(4):872–903, August 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-011-0283-
y.

Cheng:2017:SGS

[CLSW17] Xinyu Cheng, Dong Li, David Shirokoff, and Brian Wetton. On the spectral gap of a square distance matrix. Journal of Statistical Physics, 166(3–4):1029–1035, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1685-7.

Crane:2019:DSL

[CLT19] Edward Crane, Sean Ledger, and Bálint Tóth. Diffusion and superdiffusion in lattice models for colliding particles with stored momentum. Journal of Statistical Physics, 177(6): 1240–1262, December 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-019-02419-
9.pdf.

Carter:2015:EHO

[CLTC15] Jason A. Carter, Christine H. Lind, M. Phuong Truong, and Eva-Maria S. Collins. To each his own: Reproductive strategies and success of three common planarian species: Schmidtea mediterranea, Dugesia japonica, and Dugesia tigrina. Journal of Statistical Physics, 161(1):250–272, October 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.

1007/s10955-015-1310-1. See corrections [CLTC23].

Carter:2023:CEH

[CLTC23] Jason A. Carter, Christine H. Lind, M. Phuong Truong, and Eva-Maria S. Collins. Correction: To each his own. Journal of Statistical Physics, 190(1):??, January 2023. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s10955-022-03009-y. See [CLTC15].

Comtet:2013:LEP

[CLTT13] Alain Comtet, Jean-Marc Luck, Christophe Texier, and Yves Tourigny. The Lyapunov exponent of products of random 2 × 2 matrices close to the identity. Journal of Statistical Physics, 150(1):13–65, January 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0674-8.

Chereji:2011:SMN

[CM11] Razvan V. Chereji and Alexandre V. Morozov. Statistical mechanics of nucleosomes constrained by higher-order chromatin structure. Journal of Statistical Physics, 144(2):379–404, July 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0214-y.

Chetrite:2012:QFR

[CM12a] R. Chetrite and K. Mallick. Quantum fluctuation relations for the Lindblad master equation. Journal of Statistical Physics, 148(3):480–501, August 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0557-z.

Cocco:2012:ACE

[CM12b] S. Cocco and R. Monasson. Adaptive cluster expansion for the inverse Ising problem: Convergence, algorithm and tests. Journal of Statistical Physics, 147(2):252–314, April 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0463-4.

Constable:2018:EFV

[CM18] George W. A. Constable and Alan J. McKane. Exploiting fast-variables to understand population dynamics and evolution. Journal of Statistical Physics, 172(1):3–43, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-017-1900-1.pdf.

Carinci:2014:SHL

[CMGP14] Gioia Carinci, Anna De Masi, Cristian Giardinà, and Errico Presutti. Super-hydrodynamic limit in interacting particle systems. Journal of Statistical Physics, 155(5):867–887, June 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-0984-0.

Crescenzo:2014:ARR

[CMM14] Antonio Di Crescenzo, Claudio Macci, and Barbara Martinucci. Asymptotic results for random walks in continuous time with alternating rates. Journal of Statistical Physics, 154(5):1352–1364, March 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-0928-8.

Coletti:2016:IUQ

[CMM16] Cristian F. Coletti, Daniel Miranda, and Filipe Mussini. Invariance under quasi-isometries of subcritical and supercritical behavior in the Boolean model of percolation. Journal of Statistical Physics, 162(3):685–700, February 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1422-7.

Cunden:2018:FFC

[CMO18] Fabio Deelan Cunden, Francesco Mezzadri, and Neil O’Connell. Free fermions and the classical compact groups. Journal of Statistical Physics, 171(5):768–801, June 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/content/pdf/10.1007/
s10955-018-2029-6.pdf.

Cassandro:2017:PSL

[CMP17] Marzio Cassandro, Immacolata Merola, and Pierre Picco. Phase separation for the long range one-dimensional Ising model. Journal of Statistical Physics, 167(2):351–382, April 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-017-1722-1.

Cox:2010:CPW

[CMS10] J. Theodore Cox, Nevena Marić, and Rinaldo Schinazi. Contact process in a wedge. Journal of Statistical Physics, 139 (3):506–517, May 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9950-7.

Contucci:2013:FPD

[CMS13] Pierluigi Contucci, Emanuele Mingione, and Shannon Starr. Factorization properties in d-dimensional spin glasses. Rigorous results and some perspectives. Journal of Statistical Physics, 151(5):809–829, June 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0730-z.

Costa:2018:LGM

[CMSV18] M. Costa, M. Menshikov, V. Shcherbakov, and M. Vachkovskaia. Localisation in a growth model with interaction. Journal of Statistical Physics, 171(6):1150–1175, June 2018. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-018-2055-4.pdf.

Cassandro:2011:SLR

[CMV11] Marzio Cassandro, Immacolata Merola, and Maria Eulalia Vares. Study of a long range perturbation of a one-dimensional Kac model. Journal of Statistical Physics, 142(3):487–509, February 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0123-0.

Cunden:2016:LDR

[CMV16] Fabio Deelan Cunden, Francesco Mezzadri, and Pierpaolo Vivo. Large deviations of radial statistics in the twodimensional one-component plasma. Journal of Statistical Physics, 164(5):1062–1081, September 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-016-1577-x.pdf.

Comets:2011:RWB

[CMVW11] Francis Comets, Mikhail V. Menshikov, Stanislav Volkov, and Andrew R. Wade. Random walk with barycentric selfinteraction. Journal of Statistical Physics, 143(5):855–888, June 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0218-7.

Carlen:2015:PCT

[CMW15] Eric Carlen, Dawan Mustafa, and Bernt Wennberg. Propagation of chaos for the thermostatted Kac master equation. Journal of Statistical Physics, 158(6):1341–1378, March 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1155-z.

Cirillo:2013:RHE

[CN13] Emilio N. M. Cirillo and Francesca R. Nardi. Relaxation height in energy landscapes: An application to multiple metastable states. Journal of Statistical Physics, 150(6):1080– 1114, March 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0717-9.

Conlon:2014:GSL

[CN14] Joseph G. Conlon and Barbara Niethammer. On global stability for Lifschitz–Slyozov–Wagner like equations. Journal of Statistical Physics, 154(5):1251–1291, March 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0927-9.

Cirillo:2015:MGD

[CNS15] Emilio N. M. Cirillo, Francesca R. Nardi, and Julien Sohier. Metastability for general dynamics with rare transitions: Escape time and critical configurations. Journal of Statistical Physics, 161(2):365–403, October 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-015-1334-
6;   http://link.springer.com/content/pdf/10.1007/
s10955-015-1334-6.pdf.

Carney:2017:CPL

[CNZ17] Meagan Carney, Matthew Nicol, and Hong-Kun Zhang. Compound Poisson law for hitting times to periodic orbits in two-dimensional hyperbolic systems. Journal of Statistical Physics, 169(4):804–823, November 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Cancrini:2017:EDF

[CO17] Nicoletta Cancrini and Stefano Olla. Ensemble dependence of fluctuations: Canonical microcanonical equivalence of ensembles. Journal of Statistical Physics, 168(4):707–730, August 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Cohen:2009:NED

[Coh09] E. G. D. Cohen. Non existence of a density expansion of the transport coefficients in a moderately dense gas. Journal of Statistical Physics, 137(5–6):878–889, December 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-009-9838-6. See erratum [Coh10].

Cohen:2010:ENE

[Coh10] E. G. D. Cohen. Erratum to: Non Existence of a Density Expansion of the Transport Coefficients in a Moderately Dense Gas. Journal of Statistical Physics, 138(1–3):551, February 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/
pdf/10.1007/s10955-009-9851-9.pdf. See [Coh09].

Collet:2014:MLB

[Col14] Francesca Collet. Macroscopic limit of a bipartite Curie–Weiss model: A dynamical approach. Journal of Statistical Physics, 157(6):1301–1319, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1105-9.

Contucci:2010:SSR

[Con10] Pierluigi Contucci. Stochastic stability: A review and some perspectives. Journal of Statistical Physics, 138(1–3):543–550, February 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-009-9887-x.

Conrady:2011:SLT

[Con11] Florian Conrady. Space as a low-temperature regime of graphs. Journal of Statistical Physics, 142(4):898–917, February 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0135-9.

Contucci:2013:DPB

[Con13] Pierluigi Contucci. Dmitry Panchenko: The Sherrington– Kirkpatrick Model. Journal of Statistical Physics, 153(3): 551–552, November 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0851-4.

Concetti:2018:FRS

[Con18] Francesco Concetti. The full replica symmetry breaking in the Ising spin glass on random regular graph. Journal of Statistical Physics, 173(5):1459–1483, December 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Coquille:2015:EDS

[Coq15] Loren Coquille. Examples of DLR states which are not weak limits of finite volume Gibbs measures with deterministic boundary conditions. Journal of Statistical Physics, 159(4):958–971, May 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1211-3.

Cortez:2016:UPC

[Cor16] Roberto Cortez. Uniform propagation of chaos for Kac’s 1D particle system. Journal of Statistical Physics, 165(6): 1102–1113, December 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1674-x.

Chayes:2010:MVE

[CP10a] L. Chayes and V. Panferov. The McKean–Vlasov equation in finite volume. Journal of Statistical Physics, 138(1–3): 351–380, February 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-009-9913-z.
pdf.

Chernyak:2010:LID

[CP10b] V. G. Chernyak and A. P. Polikarpov. Light induced drift and heat transfer of one-component gas in a capillary. Journal of Statistical Physics, 140(3):504–517, August 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-0001-1.

Cahoy:2012:FBP

[CP12] Dexter O. Cahoy and Federico Polito. On a fractional binomial process. Journal of Statistical Physics, 146(3):646–662, February 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0408-3.

Chafai:2014:NSO

[CP14] Djalil Chafaï and Sandrine Péché. A note on the second order universality at the edge of Coulomb gases on the plane. Journal of Statistical Physics, 156(2):368–383, July 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1007-x.

Carbone:2015:HOQ

[CP15a] Raffaella Carbone and Yan Pautrat. Homogeneous open quantum random walks on a lattice. Journal of Statistical Physics, 160(5):1125–1153, September 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1261-6.

Corwin:2015:PNI

[CP15b] Ivan Corwin and Leonid Petrov. The q-PushASEP: A new integrable model for traffic in 1 + 1 dimension. Journal of Statistical Physics, 160(4):1005–1026, August 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1218-9.

Cassandro:2016:NFO

[CP16] M. Cassandro and E. Presutti. A note on first order phase transitions in the continuum. Journal of Statistical Physics, 162(4):994–996, February 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1435-2.

Camargo:2017:TFT

[CP17a] Darcy Camargo and Serguei Popov. Total flooding time and rumor propagation on graphs. Journal of Statistical Physics, 166(6):1558–1571, March 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-017-1731-0.

Chen:2017:TCS

[CP17b] Wei-Kuo Chen and Dmitry Panchenko. Temperature chaos in some spherical mixed p-spin models. Journal of Statistical Physics, 166(5):1151–1162, March 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-016-1709-
3.

Carati:2018:CTF

[CP18] A. Carati and A. Ponno. Chopping time of the FPU α-model. Journal of Statistical Physics, 170(5):883–894, March 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Can:2019:PPR

[CP19] Van Hao Can and Viet-Hung Pham. Persistence probability of random Weyl polynomial. Journal of Statistical Physics, 176 (1):262–277, July 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Correggi:2011:CRS

[CPRY11] M. Correggi, F. Pinsker, N. Rougerie, and J. Yngvason. Critical rotational speeds in the Gross–Pitaevskii theory on a disc with Dirichlet boundary conditions. Journal of Statistical Physics, 143(2):261–305, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0182-2.

Colomo:2019:ACF

[CPS19] F. Colomo, A. G. Pronko, and A. Sportiello. Arctic curve of the free-fermion six-vertex model in an L-shaped domain. Journal of Statistical Physics, 174(1):1–27, January 2019. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Comets:2010:KGF

[CPSV10] Francis Comets, Serguei Popov, Gunter M. Schütz, and Marina Vachkovskaia. Knudsen gas in a finite random tube: Transport diffusion and first passage properties. Journal of Statistical Physics, 140(5):948–984, September 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0023-8.

Campellone:2010:RMF

[CPV10] Matteo Campellone, Giorgio Parisi, and Miguel Angel Virasoro. Replica method and finite volume corrections. Journal of Statistical Physics, 138(1–3):29–39, February 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-009-9891-1.

Comets:2013:LPP

[CQR13] Francis Comets, Jeremy Quastel, and Alejandro F. Ramírez. Last passage percolation and traveling fronts. Journal of Statistical Physics, 152(3):419–451, August 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0779-8.

Corwin:2015:RFP

[CQR15] Ivan Corwin, Jeremy Quastel, and Daniel Remenik. Renormalization fixed point of the KPZ universality class. Journal of Statistical Physics, 160(4):815–834, August 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1243-8.

Chen:2011:BLH

[CR11a] Thomas Chen and Igor Rodnianski. Boltzmann limit for a homogeneous Fermi gas with dynamical Hartree–Fock interactions in a random medium. Journal of Statistical Physics, 142(5):1000–1051, March 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0145-7.

Chung:2011:ECT

[CR11b] Yoo Jin Chung and Steven W. Rick. The effects of charge transfer interactions on the properties of ice ih. Journal of Statistical Physics, 145(2):355–364, October 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0255-2.

Camiola:2014:HMC

[CR14] V. D. Camiola and V. Romano. Hydrodynamical model for charge transport in graphene. Journal of Statistical Physics, 157(6):1114–1137, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1102-z.

Collet:2016:SSF

[CR16] Francesca Collet and Wioletta Ruszel. Synchronization and spin-flop transitions for a mean-field XY model in random field. Journal of Statistical Physics, 164(3):645–666, August 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-016-1557-1.pdf.

Cohen:2017:PSP

[CR17] Yossi Cohen and Daniel H. Rothman. Path selection in a Poisson field. Journal of Statistical Physics, 167(3–4):703– 712, May 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Crawford:2011:RFI

[Cra11] Nicholas Crawford. On random field induced ordering in the classical XY model. Journal of Statistical Physics, 142(1):11– 42, January 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-0094-6.

Crane:2013:PPM

[Cra13] Harry Crane. Permanental partition models and Markovian Gibbs structures. Journal of Statistical Physics, 153(4): 698–726, November 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0855-0.

Coronel:2015:SDG

[CRL15] Daniel Coronel and Juan Rivera-Letelier. Sensitive dependence of Gibbs measures at low temperatures. Journal of Statistical Physics, 160(6):1658–1683, September 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1288-8.

Crommelin:2012:ESD

[Cro12] Daan Crommelin. Estimation of space-dependent diffusions and potential landscapes from non-equilibrium data. Journal of Statistical Physics, 149(2):220–233, October 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0597-4.

Coletti:2012:SSM

[CRS12] Cristian F. Coletti, Pablo M. Rodríguez, and Rinaldo B. Schinazi. A spatial stochastic model for rumor transmission. Journal of Statistical Physics, 147(2):375–381, April 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0469-y.

Cabezas:2014:NEP

[CRS14] M. Cabezas, L. T. Rolla, and V. Sidoravicius. Nonequilibrium phase transitions: Activated random walks at criticality. Journal of Statistical Physics, 155(6):1112–1125, June 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0909-3.

Connaughton:2013:NEP

[CRTZ13] Colm Connaughton, R. Rajesh, Roger Tribe, and Oleg Zaboronski. Non-equilibrium phase diagram for a model with coalescence, evaporation and deposition. Journal of Statistical Physics, 152(6):1115–1144, September 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0800-2.

Chevillard:2013:GMC

[CRV13] Laurent Chevillard, Rémi Rhodes, and Vincent Vargas. Gaussian multiplicative chaos for symmetric isotropic matrices. Journal of Statistical Physics, 150(4):678–703, February 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0697-9.

Carvalho:2017:SAE

[CRV17] Maria Carvalho, Fagner B. Rodrigues, and Paulo Varandas. Semigroup actions of expanding maps. Journal of Statistical Physics, 166(1):114–136, January 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1697-3.

Chang:2010:WGC

[CS10a] Shu-Chiuan Chang and Robert Shrock. Weighted graph colorings. Journal of Statistical Physics, 138(1–3):496–542, February 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-009-9882-2.

Chatterjee:2010:DRS

[CS10b] Sakuntala Chatterjee and Gunter M. Schütz. Determinant representation for some transition probabilities in the TASEP with second class particles. Journal of Statistical Physics, 140(5):900–916, September 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0022-9.

Chernov:2010:ULE

[CS10c] N. Chernov and N. Simányi. Upgrading the local ergodic theorem for planar semi-dispersing billiards. Journal of Statistical Physics, 139(3):355–366, May 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9927-6.

Chang:2012:SER

[CS12] Shu-Chiuan Chang and Robert Shrock. Some exact results on bond percolation. Journal of Statistical Physics, 149(4): 676–700, November 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0616-5.

Chang:2015:EPF

[CS15] Shu-Chiuan Chang and Robert Shrock. Exact partition functions for the q-state Potts model with a generalized magnetic field on lattice strip graphs. Journal of Statistical Physics, 161(4):915–932, November 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1357-z.

Chen:2016:ERM

[CS16a] Ying Chen and Christopher A. Schuh. Elasticity of random multiphase materials: Percolation of the stiffness tensor. Journal of Statistical Physics, 162(1):232–241, January 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1387-6.

Chino:2016:QCP

[CS16b] Yuki Chino and Akira Sakai. The quenched critical point for self-avoiding walk on random conductors. Journal of Statistical Physics, 163(4):754–764, May 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1477-0.

Colomo:2016:ACS

[CS16c] F. Colomo and A. Sportiello. Arctic curves of the sixvertex model on generic domains: The tangent method. Journal of Statistical Physics, 164(6):1488–1523, September 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1590-0.

Chang:2019:SEG

[CS19] Shu-Chiuan Chang and Robert Shrock. Study of exponential growth constants of directed heteropolygonal Archimedean lattices. Journal of Statistical Physics, 174(6):1288–1315, March 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Chang:2017:RPM

[CSAS17] Chih-Hung Chang, Jing-Yi Su, Hasan Akin, and Ferhat Sah. Reversibility problem of multidimensional finite cellular automata. Journal of Statistical Physics, 168(1):208–231, July 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Chen:2011:OCE

[CSC11] Liwei Chen, Guozhen Su, and Jincan Chen. Oscillating Casimir effect of a trapped Fermi gas. Journal of Statistical Physics, 143(3):523–531, May 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0189-8.

Chang:2014:PIA

[CSC14] Joshua C. Chang, Van M. Savage, and Tom Chou. A pathintegral approach to Bayesian inference for inverse problems using the semiclassical approximation. Journal of Statistical Physics, 157(3):582–602, November 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1059-y.

Capanna:2019:TIM

[CSL19] Monia Capanna and Nahuel Soprano-Loto. Turing instability in a model with two interacting Ising lines: Non-equilibrium fluctuations. Journal of Statistical Physics, 174(2):365–403, January 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Corwin:2015:SWL

[CSS15] Ivan Corwin, Timo Seppäläinen, and Hao Shen. The strict– weak lattice polymer. Journal of Statistical Physics, 160(4): 1027–1053, August 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1267-0.

Celani:2011:MFA

[CSV11] A. Celani, T. S. Shimizu, and M. Vergassola. Molecular and functional aspects of bacterial chemotaxis. Journal of Statistical Physics, 144(2):219–240, July 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0251-6.

Coletti:2010:IMD

[CT10] Cristian F. Coletti and Pierre Tisseur. Invariant measures and decay of correlations for a class of ergodic probabilistic cellular automata. Journal of Statistical Physics, 140(1):103–121, July 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-9985-9.

Cabana:2013:LDD

[CT13] Tanguy Cabana and Jonathan Touboul. Large deviations, dynamics and phase transitions in large stochastic and disordered neural networks. Journal of Statistical Physics, 153 (2):211–269, October 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0818-5.

Carmi:2010:DFA

[CTB10] Shai Carmi, Lior Turgeman, and Eli Barkai. On distributions of functionals of anomalous diffusion paths. Journal of Statistical Physics, 141(6):1071–1092, December 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0086-6.

Chen:2011:CII

[CTH^{+}11] Yuling Chen, Gang Tang, Kui Han, Hui Xia, Dapeng Hao, Zhipeng Xun, and Rongji Wen. Conformal invariance of isoheight lines of the (2 + 1)-dimensional Wolf–Villain surfaces. Journal of Statistical Physics, 143(3):501–508, May 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0187-x.

Carro:2013:RNI

[CTM13] Adrián Carro, Raúl Toral, and Maxi San Miguel. The role of noise and initial conditions in the asymptotic solution of a bounded confidence, continuous-opinion model. Journal of Statistical Physics, 151(1–2):131–149, April 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0635-2.

Comtet:2010:PRM

[CTT10] Alain Comtet, Christophe Texier, and Yves Tourigny. Products of random matrices and generalised quantum point scatterers. Journal of Statistical Physics, 140(3):427–466, August 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-0005-x.

Comtet:2011:SQM

[CTT11] Alain Comtet, Christophe Texier, and Yves Tourigny. Supersymmetric quantum mechanics with Lévy disorder in one dimension. Journal of Statistical Physics, 145(5):1291–1323, December 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0351-3.

Cugliandolo:2017:ASI

[Cug17] Leticia F. Cugliandolo. Artificial spin-ice and vertex models. Journal of Statistical Physics, 167(3–4):499–514, May 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Cioletti:2016:GRI

[CV16] Leandro Cioletti and Roberto Vila. Graphical representations for Ising and Potts models in general external fields. Journal of Statistical Physics, 162(1):81–122, January 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1396-5.

Cameron:2014:FCN

[CVE14] Maria Cameron and Eric Vanden-Eijnden. Flows in complex networks: Theory, algorithms, and application to Lennard-Jones cluster rearrangement. Journal of Statistical Physics, 156(3):427–454, August 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-0997-8.

Coquille:2018:ADS

[CvELR18] Loren Coquille, Aernout C. D. van Enter, Arnaud Le Ny, and Wioletta M. Ruszel. Absence of Dobrushin states for 2 d long-range Ising models. Journal of Statistical Physics, 172(5):1210–1222, September 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Chen:2017:LBD

[CWD17] Yao Chen, Xudong Wang, and Weihua Deng. Localization and ballistic diffusion for the tempered fractional Brownian– Langevin motion. Journal of Statistical Physics, 169(1):18– 37, October 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Chen:2018:NSA

[CYZ18] Jianyu Chen, Yun Yang, and Hong-Kun Zhang. Nonstationary almost sure invariance principle for hyperbolic systems with singularities. Journal of Statistical Physics, 172(6): 1499–1524, September 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Chernov:2013:ECS

[CZZ13] Nikolai Chernov, Hong-Kun Zhang, and Pengfei Zhang. Electrical current in Sinai billiards under general small forces. Journal of Statistical Physics, 153(6):1065–1083, December 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0870-1.

Delgosha:2018:LBH

[DA18] Payam Delgosha and Venkat Anantharam. Load balancing in hypergraphs. Journal of Statistical Physics, 173(3–4):546–625, November 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Dai:2017:EDS

[Dai17] Yan Dai. The exit distribution for smart kinetic walk with symmetric and asymmetric transition probability. Journal of Statistical Physics, 166(6):1455–1463, March 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-017-1735-9.

Daligault:2011:RKT

[Dal11] Jérôme Daligault. Renormalized kinetic theory of classical fluids in and out of equilibrium. Journal of Statistical Physics, 143(6):1189–1246, June 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0228-5.

deArruda:2013:INP

[dAPdA^{+}13] Guilherme F. de Arruda, Thomas Kauê Dal’Maso Peron, Marinho Gomes de Andrade, Jorge Alberto Achcar, and Francisco Aparecido Rodrigues. The influence of network properties on the synchronization of Kuramoto oscillators quantified by a Bayesian regression analysis. Journal of Statistical Physics, 152(3):519–533, August 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0775-z.

Prado:2011:LET

[dAPS11] Fernando Pigeard de Almeida Prado and Gunter M. Schütz. Loss of ergodicity in the transition from annealed to quenched disorder in a finite kinetic Ising model. Journal of Statistical Physics, 142(5):984–999, March 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-011-0143-
9.

Degond:2013:HHB

[DARM^{+}13] P. Degond, C. Appert-Rolland, M. Moussaïd, J. Pettré, and G. Theraulaz. A hierarchy of heuristic-based models of crowd dynamics. Journal of Statistical Physics, 152(6):1033–1068, September 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0805-x.

Davydov:2011:INE

[Dav11] Alexander Davydov. Inequalities for non-equilibrium fluctuations of work. Journal of Statistical Physics, 142(2):394–402, January 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0119-9.

deBernardini:2019:IDI

[dBGP19] Diego F. de Bernardini, Christophe Gallesco, and Serguei Popov. An improved decoupling inequality for random interlacements. Journal of Statistical Physics, 177(6):1216–1239, December 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

deBernardini:2015:RFR

[dBP15] Diego F. de Bernardini and Serguei Popov. Russo’s formula for random interlacements. Journal of Statistical Physics, 160(2):321–335, July 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1268-z.

deCrouy-Chanel:2019:RKD

[dCCS19] Marthe de Crouy-Chanel and Damien Simon. Random knots in 3-dimensional 3-colour percolation: Numerical results and conjectures. Journal of Statistical Physics, 176(3):574–590, August 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

deCandia:2011:DHC

[dCFC11] A. de Candia, A. Fierro, and A. Coniglio. Dynamical heterogeneities in the crossover region from gel-like to glassylike behavior. Journal of Statistical Physics, 145(3):652–660, November 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0314-8.

Dekking:2010:CFP

[DD10] Michel Dekking and Henk Don. Correlated fractal percolation and the Palis conjecture. Journal of Statistical Physics, 139(2):307–325, April 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-009-9907-x.
pdf.

Dharmaraja:2015:CTE

[DD15] Selvamuthu Dharmaraja and Antonio Di Crescenzo. A continuous-time Ehrenfest model with catastrophes and its jump-diffusion approximation. Journal of Statistical Physics, 161(2):326–345, October 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1336-4.

Dessalles:2018:ESS

[DDC18] Renaud Dessalles, Maria D’Orsogna, and Tom Chou. Exact steady-state distributions of multispecies birth–deathimmigration processes: Effects of mutations and carrying capacity on diversity. Journal of Statistical Physics, 173(1):182– 221, October 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Dellacherie:2015:DSF

[DDF15] Stéphane Dellacherie, François Dubois, and Stéphan Fauve. Discrete simulation of fluid dynamics. Journal of Statistical Physics, 161(6):1325–1326, December 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1407-6;   http://link.springer.com/
content/pdf/10.1007/s10955-015-1407-6.pdf.

DeRoeck:2017:SDP

[DDHS17] Wojciech De Roeck, Abhishek Dhar, François Huveneers, and Marius Schütz. Step density profiles in localized chains. Journal of Statistical Physics, 167(5):1143–1163, June 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Das:2014:HCF

[DDN14] Suman G. Das, Abhishek Dhar, and Onuttom Narayan. Heat conduction in the α–β Fermi–Pasta–Ulam chain. Journal of Statistical Physics, 154(1–2):204–213, January 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0871-0.

Debbasch:2012:CDP

[DEF12] F. Debbasch, D. Espaze, and V. Foulonneau. Can diffusions propagate? Journal of Statistical Physics, 149(1):37–49, October 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0580-0.

Damron:2015:CDF

[DEK^{+}15] M. Damron, S. M. Eckner, H. Kogan, C. M. Newman, and V. Sidoravicius. Coarsening dynamics on Z^{d} with frozen vertices. Journal of Statistical Physics, 160(1):60–72, July 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1247-4.

dellErba:2012:SMS

[del12] Matías Germán dell’Erba. Statistical mechanics of a simplified bipartite matching problem: An analytical treatment. Journal of Statistical Physics, 146(6):1263–1273, March 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0447-4.

Desai:2011:NPN

[Des11] Rashmi C. Desai. Noëlle pottier: Nonequilibrium statistical physics, linear irreversible processes. Journal of Statistical Physics, 142(2):439–440, January 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0114-6.

Dettmann:2012:NHM

[Det12] Carl P. Dettmann. New horizons in multidimensional diffusion: The Lorentz gas and the Riemann hypothesis. Journal of Statistical Physics, 146(1):181–204, January 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0397-2.

Dettmann:2018:ICR

[Det18] Carl P. Dettmann. Isolation and connectivity in random geometric graphs with self-similar intensity measures. Journal of Statistical Physics, 172(3):679–700, August 2018. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-018-2059-0.pdf.

DeVille:2019:SSQ

[DeV19] Lee DeVille. Synchronization and stability for quantum Kuramoto. Journal of Statistical Physics, 174(1):160–187, January 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Deveaux:2010:POM

[DF10] Vincent Deveaux and Roberto Fernández. Partially ordered models. Journal of Statistical Physics, 141(3):476–516, November 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-0063-0.

Degond:2016:ALT

[DF16] Pierre Degond and Francis Filbet. On the asymptotic limit of the three dimensional Vlasov–Poisson system for large magnetic field: Formal derivation. Journal of Statistical Physics, 165(4):765–784, November 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-016-1645-2.pdf.

DeLuca:2017:CPN

[DF17] Lucia De Luca and Gero Friesecke. Classification of particle numbers with unique Heitmann–Radin minimizer. Journal of Statistical Physics, 167(6):1586–1592, June 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Daletskii:2018:NEP

[DF18] Alexei Daletskii and Dmitri Finkelshtein. Non-equilibrium particle dynamics with unbounded number of interacting neighbors. Journal of Statistical Physics, 173(6):1639–1659, December 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Decelle:2018:TRB

[DFF18] A. Decelle, G. Fissore, and C. Furtlehner. Thermodynamics of restricted Boltzmann machines and related learning dynamics. Journal of Statistical Physics, 172(6):1576–1608, September 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Di:2017:MSG

[DFL17] Yana Di, Yuwei Fan, and Ruo Li. 13-moment system with global hyperbolicity for quantum gas. Journal of Statistical Physics, 167(5):1280–1302, June 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Degond:2014:LSP

[DFR14] Pierre Degond, Amic Frouvelle, and Gaël Raoul. Local stability of perfect alignment for a spatially homogeneous kinetic model. Journal of Statistical Physics, 157(1):84–112, October 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1062-3.

Derrida:2014:LPC

[DG14] Bernard Derrida and Giambattista Giacomin. Log-periodic critical amplitudes: A perturbative approach. Journal of Statistical Physics, 154(1–2):286–304, January 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0774-0.

Destainville:2015:EAS

[DG15a] Nicolas Destainville and Suresh Govindarajan. Estimating the asymptotics of solid partitions. Journal of Statistical Physics, 158(4):950–967, February 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1147-z.

Dubedat:2015:AHC

[DG15b] Julien Dubédat and Reza Gheissari. Asymptotics of height change on toroidal Temperleyan dimer models. Journal of Statistical Physics, 159(1):75–100, April 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1181-x.

Dudzinski:2015:CDS

[DG15c] Marcin Dudziński and Przemyslaw Górka. The crystalline dynamics of spiral-shaped curves. Journal of Statistical Physics, 160(2):409–416, July 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-015-1254-5.
pdf.

DeCarlo:2017:GSN

[DG17a] Leonardo De Carlo and Davide Gabrielli. Gibbsian stationary non-equilibrium states. Journal of Statistical Physics, 168(6): 1191–1222, September 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

DeCarlo:2017:TAL

[DG17b] Leonardo De Carlo and Davide Gabrielli. Totally asymmetric limit for models of heat conduction. Journal of Statistical Physics, 168(3):508–534, August 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Das:2019:DCM

[DG19a] Suddhasattwa Das and Dimitrios Giannakis. Delay-coordinate maps and the spectra of Koopman operators. Journal of Statistical Physics, 175(6):1107–1145, June 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

DiFrancesco:2019:ACA

[DG19b] Philippe Di Francesco and Emmanuel Guitter. The Arctic curve for Aztec rectangles with defects via the tangent method. Journal of Statistical Physics, 176(3):639–678, August 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Dommers:2018:LDA

[DGGvdH18] Sander Dommers, Cristian Giardinà, Claudio Giberti, and Remco van der Hofstad. Large deviations for the annealed Ising model on inhomogeneous random graphs: Spins and degrees. Journal of Statistical Physics, 173(3–4):1045–1081, November 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Dyhr:2011:SAW

[DGK^{+}11] Ben Dyhr, Michael Gilbert, Tom Kennedy, Gregory F. Lawler, and Shane Passon. The self-avoiding walk spanning a strip. Journal of Statistical Physics, 144(1):1–22, July 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0258-z.

Delattre:2016:NDM

[DGL16] Sylvain Delattre, Giambattista Giacomin, and Eric Luçon. A note on dynamical models on random graphs and Fokker– Planck equations. Journal of Statistical Physics, 165(4): 785–798, November 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1652-3.

DeMasi:2015:HLI

[DGLP15] A. De Masi, A. Galves, E. Löcherbach, and E. Presutti. Hydrodynamic limit for interacting neurons. Journal of Statistical Physics, 158(4):866–902, February 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1145-1.

Dommers:2010:IMP

[DGvdH10] Sander Dommers, Cristian Giardinà, and Remco van der Hofstad. Ising models on power-law random graphs. Journal of Statistical Physics, 141(4):638–660, November 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-010-0067-9.pdf.

Dereudre:2019:PTC

[DH19] David Dereudre and Pierre Houdebert. Phase transition for continuum Widom–Rowlinson model with random radii. Journal of Statistical Physics, 174(1):56–76, January 2019. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Durr:2011:STF

[DHK11] Detlef Dürr, Günter Hinrichs, and Martin Kolb. On a stochastic Trotter formula with application to spontaneous localization models. Journal of Statistical Physics, 143(6):1096– 1119, June 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0235-6.

Demircigil:2017:MAM

[DHL^{+}17] Mete Demircigil, Judith Heusel, Matthias Löwe, Sven Upgang, and Franck Vermet. On a model of associative memory with huge storage capacity. Journal of Statistical Physics, 168 (2):288–299, July 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

denHollander:2011:KDT

[dHNT11] F. den Hollander, F. R. Nardi, and A. Troiani. Kawasaki dynamics with two types of particles: Stable/ metastable configurations and communication heights. Journal of Statistical Physics, 145(6):1423–1457, December 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-011-0370-0.pdf.

denHollander:2012:KDT

[dHNT12] F. den Hollander, F. R. Nardi, and A. Troiani. Kawasaki dynamics with two types of particles: Critical droplets.

Journal of Statistical Physics, 149(6):1013–1057, December 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0637-0.

denHollander:2013:CPV

[dHO13] F. den Hollander and A. A. Opoku. Copolymer with pinning: Variational characterization of the phase diagram. Journal of Statistical Physics, 152(5):846–893, September 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0747-3.

DalBorgo:2018:APD

[DHR18] M. Dal Borgo, E. Hovhannisyan, and A. Rouault. Asymptotic properties of the density of particles in β-ensembles. Journal of Statistical Physics, 170(3):439–465, February 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Damron:2018:AET

[DHS18] Michael Damron, Jack Hanson, and Philippe Sosoe. Arm events in two-dimensional invasion percolation. Journal of Statistical Physics, 173(5):1321–1352, December 2018. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Dostoglou:2019:EMC

[DHX19] S. Dostoglou, A. Hughes, and Jianfei Xue. On entropy minimization and convergence. Journal of Statistical Physics, 177 (3):485–505, November 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Datsyuk:2013:SPC

[DI13] Vitaly V. Datsyuk and Iryna V. Ivanytska. Statistical properties of conduction electrons in an isolated metal nanosphere. Journal of Statistical Physics, 152(5):969–978, September 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0798-5.

Diaconis:2011:MMT

[Dia11] Persi Diaconis. The mathematics of mixing things up. Journal of Statistical Physics, 144(3):445–458, August 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0284-x.

Ding:2011:MEM

[DJRZ11] Jiu Ding, Congming Jin, Noah H. Rhee, and Aihui Zhou. A maximum entropy method based on piecewise linear functions for the recovery of a stationary density of interval mappings. Journal of Statistical Physics, 145(6):1620–1639, December 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0366-9.

Durhuus:2010:SDC

[DJW10] Bergfinnur Durhuus, Thordur Jonsson, and John F. Wheater. On the spectral dimension of causal triangulations. Journal of Statistical Physics, 139(5):859–881, June 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-9968-x.

Desai:2009:DSO

[DK09] Rashmi C. Desai and Raymond Kapral. Dynamics of selforganized and self-assembled structures. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-88361-X (hardcover). xiv + 328 pp. LCCN Q172.5.C45 D47 2009.

Deviren:2010:DPT

[DK10] Bayram Deviren and Mustafa Keskin. Dynamic phase transitions and compensation temperatures in a mixed spin-3/2 and spin-5/2 Ising system. Journal of Statistical Physics, 140(5):934–947, September 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0025-6.

Daletskii:2014:PTQ

[DKKP14] Alexei Daletskii, Yuri Kondratiev, Yuri Kozitsky, and Tanja Pasurek. A phase transition in a quenched amorphous ferromagnet. Journal of Statistical Physics, 156(1):156–176, July 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-014-0996-9.pdf.

Dhar:2019:TPC

[DKLS19] Abhishek Dhar, Aritra Kundu, Joel L. Lebowitz, and Jasen A. Scaramazza. Transport properties of the classical Toda chain: Effect of a pinning potential. Journal of Statistical Physics, 175(6):1298–1310, June 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Danino:2018:ESS

[DKS18] Matan Danino, David A. Kessler, and Nadav M. Shnerb. Environmental stochasticity and the speed of evolution. Journal of Statistical Physics, 172(1):126–142, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Dhahri:2019:QMC

[DKY19] Ameur Dhahri, Chul Ki Ko, and Hyun Jae Yoo. Quantum Markov chains associated with open quantum random walks. Journal of Statistical Physics, 176(5):1272–1295, September 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Duan:2012:HLB

[DL12] Renjun Duan and Wei-Xi Li. Hypocoercivity for the linear Boltzmann equation with confining forces. Journal of Statistical Physics, 148(2):306–324, August 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0545-3.

Duan:2017:VPB

[DL17] Renjun Duan and Shuangqian Liu. The Vlasov–Poisson– Boltzmann system for a disparate mass binary mixture. Journal of Statistical Physics, 169(3):614–684, November 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Dechant:2011:SFP

[DLBK11] A. Dechant, E. Lutz, E. Barkai, and D. A. Kessler. Solution of the Fokker–Planck equation with a logarithmic potential. Journal of Statistical Physics, 145(6):1524–1545, December 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0363-z.

delaLlave:2011:DTA

[dlLL11] Rafael de la Llave and Alejandro Luque. Differentiability at the tip of Arnold tongues for Diophantine rotations: Numerical studies and renormalization group explanations. Journal of Statistical Physics, 143(6):1154–1188, June 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0233-8.

delaLlave:2011:CLS

[dlLOP11] Rafael de la Llave, Arturo Olvera, and Nikola P. Petrov. Combination laws for scaling exponents and relation to the geometry of renormalization operators. Journal of Statistical Physics, 143(5):889–920, June 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0211-1.

delaLlave:2016:REC

[dlLSZ16] Rafael de la Llave, Xifeng Su, and Lei Zhang. Resonant equilibrium configurations in quasi-periodic media: Perturbative expansions. Journal of Statistical Physics, 162(6):1522–1538, March 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1464-5.

Ding:2016:CRC

[DLLX16] Mingnan Ding, Yihao Liang, Bing-Sui Lu, and Xiangjun Xing. Charge renormalization and charge oscillation in asymmetric primitive model of electrolytes. Journal of Statistical Physics, 165(5):970–989, December 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1644-3.

deLima:2014:MSL

[dLP14] Bernardo N. B. de Lima and Aldo Procacci. The Mayer series of the Lennard-Jones gas: Improved bounds for the convergence radius. Journal of Statistical Physics, 157(3):422–435, November 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1094-8.

Duncan:2016:VRU

[DLP16] A. B. Duncan, T. Lelièvre, and G. A. Pavliotis. Variance reduction using nonreversible Langevin samplers. Journal of Statistical Physics, 163(3):457–491, May 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-016-1491-2.pdf.

deLima:2015:RMB

[dLPS15] Bernardo N. B. de Lima, Aldo Procacci, and Rémy Sanchis. A remark on monotonicity in Bernoulli bond percolation. Journal of Statistical Physics, 160(5):1244–1248, September 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1284-z.

Degond:2014:EDW

[DLR14] Pierre Degond, Jian-Guo Liu, and Christian Ringhofer. Evolution of the distribution of wealth in an economic environment driven by local Nash equilibria. Journal of Statistical Physics, 154(3):751–780, February 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0888-4.

Dinaburg:2010:NSS

[DLS10] Efim Dinaburg, Dong Li, and Yakov G. Sinai. Navier–Stokes system on the unit square with no slip boundary condition. Journal of Statistical Physics, 141(2):342–358, October 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-0051-4.

DeMuse:2018:PTE

[DLY18] Ryan DeMuse, Danielle Larcomb, and Mei Yin. Phase transitions in edge-weighted exponential random graphs: Neardegeneracy and universality. Journal of Statistical Physics, 171(1):127–144, April 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Desvillettes:2010:SAA

[DM10] Laurent Desvillettes and Julien Mathiaud. Some aspects of the asymptotics leading from gas-particles equations towards multiphase flows equations. Journal of Statistical Physics, 141(1):120–141, October 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0044-3.

Degond:2011:MMS

[DM11] Pierre Degond and Sébastien Motsch. A macroscopic model for a system of swarming agents using curvature control. Journal of Statistical Physics, 143(4):685–714, May 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0201-3.

Das:2012:FTF

[DM12] Shankar P. Das and Gene F. Mazenko. Field theoretic formulation of kinetic theory: Basic development. Journal of Statistical Physics, 149(4):643–675, November 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0610-y.

Das:2013:NKT

[DM13] Shankar P. Das and Gene F. Mazenko. Newtonian kinetic theory and the ergodic–nonergodic transition. Journal of Statistical Physics, 152(1):159–194, July 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0755-3.

Demaerel:2018:DRC

[DM18a] Thibaut Demaerel and Christian Maes. Death and resurrection of a current by disorder, interaction or periodic driving. Journal of Statistical Physics, 173(1):99–119, October 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Derrida:2018:FSC

[DM18b] Bernard Derrida and Peter Mottishaw. Finite size corrections to the Parisi overlap function in the GREM. Journal of Statistical Physics, 172(2):592–610, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

DiCrescenzo:2014:ARR

[DMM14] Antonio Di Crescenzo, Claudio Macci, and Barbara Martinucci. Asymptotic results for random walks in continuous time with alternating rates. Journal of Statistical Physics, 154(5):1352–1364, March 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-0928-8.

deMaere:2012:EDC

[dMP12] Augustin de Maere and Lise Ponselet. Exponential decay of correlations for strongly coupled Toom probabilistic cellular automata. Journal of Statistical Physics, 147(3):634–652, May 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0487-9.

Duan:2017:NNL

[DMP17] Jun Tao Duan, Heinrich Matzinger, and Ionel Popescu. Nonnormal limiting distribution for optimal alignment scores of strings in binary alphabets. Journal of Statistical Physics, 168(5):1056–1084, September 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

delMolino:2016:RGE

[dMPTW16] Luis Carlos García del Molino, Khashayar Pakdaman, Jonathan Touboul, and Gilles Wainrib. The real Ginibre ensemble with k = O(n) real eigenvalues. Journal of Statistical Physics, 163(2):303–323, April 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1485-0.

Dinh:2012:EZH

[DMS12] Tien-Cuong Dinh, George Marinescu, and Viktoria Schmidt. Equidistribution of zeros of holomorphic sections in the noncompact setting. Journal of Statistical Physics, 148(1):113– 136, July 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0526-6.

deMenibus:2017:SOC

[dMS17a] Benjamin Hellouin de Menibus and Mathieu Sablik. Selforganisation in cellular automata with coalescent particles:

Qualitative and quantitative approaches. Journal of Statistical Physics, 167(5):1180–1220, June 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Dettmann:2017:UHT

[DMS17b] Carl P. Dettmann, Jens Marklof, and Andreas Strömbergsson. Universal hitting time statistics for integrable flows. Journal of Statistical Physics, 166(3–4):714–749, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-016-1604-y.pdf.

Damanik:2013:OPU

[DMY13] David Damanik, Paul Munger, and William N. Yessen. Orthogonal polynomials on the unit circle with Fibonacci Verblunsky coefficients, II. Applications. Journal of Statistical Physics, 153(2):339–362, October 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-013-0830-
9.

Degond:2010:CMM

[DNBS10] Pierre Degond, Laurent Navoret, Richard Bon, and David Sanchez. Congestion in a macroscopic model of self-driven particles modeling gregariousness. Journal of Statistical Physics, 138(1–3):85–125, February 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9879-x.

Duncan:2017:UPU

[DNP17] A. B. Duncan, N. Nüsken, and G. A. Pavliotis. Using perturbed underdamped Langevin dynamics to efficiently sample from probability distributions. Journal of Statistical Physics, 169(6):1098–1131, December 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-017-1906-8.pdf.

Dagrosa:2014:GRT

[DO14a] Eduardo Dagrosa and Aleksander L. Owczarek. Generalizing ribbons and the twist of the lattice ribbon. Journal of Statistical Physics, 155(2):392–417, April 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0941-y.

DOvidio:2014:CCR

[D’O14b] Mirko D’Ovidio. Coordinates changed random fields on the sphere. Journal of Statistical Physics, 154(4):1153–1176, February 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0911-9.

DeGregorio:2015:RRF

[DO15a] Alessandro De Gregorio and Enzo Orsingher. Reflecting random flights. Journal of Statistical Physics, 160(6):1483–1506, September 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1293-y.

DeMasi:2015:QSH

[DO15b] Anna De Masi and Stefano Olla. Quasi-static hydrodynamic limits. Journal of Statistical Physics, 161(5):1037–1058, December 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1383-x.

Dereich:2018:LNF

[DO18] Steffen Dereich and Marcel Ortgiese. Local neighbourhoods for first-passage percolation on the configuration model. Journal of Statistical Physics, 173(3–4):485–501, November 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Dodson:2015:MGP

[Dod15] C. T. J. Dodson. A model for Gaussian perturbations of graphene. Journal of Statistical Physics, 161(4):933–941, November 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1362-2.

Duran-Olivencia:2016:DDF

[DOGK16] Miguel A. Durán-Olivencia, Benjamin D. Goddard, and Serafim Kalliadasis. Dynamical density functional theory for orientable colloids including inertia and hydrodynamic interactions. Journal of Statistical Physics, 164(4):785–809, August 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1545-5.

Duran-Olivencia:2019:IRF

[DOGKP19] Miguel A. Durán-Olivencia, Rishabh S. Gvalani, Serafim Kalliadasis, and Grigorios A. Pavliotis. Instability, rupture and fluctuations in thin liquid films: Theory and computations. Journal of Statistical Physics, 174(3):579–604, February 2019. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-018-2200-0.pdf.

deOliveira:2018:PSS

[dOP18] César R. de Oliveira and Mariane Pigossi. Point spectrum and SULE for time-periodic perturbations of discrete 1D Schrödinger operators with electric fields. Journal of Statistical Physics, 173(1):140–162, October 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

DeMasi:2019:NFL

[DOP19] Anna De Masi, Stefano Olla, and Errico Presutti. A note on Fick’s Law with phase transitions. Journal of Statistical Physics, 175(1):203–211, April 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Duarte:2015:HLS

[DOR15] Aline Duarte, Guilherme Ost, and Andrés A. Rodríguez. Hydrodynamic limit for spatially structured interacting neurons. Journal of Statistical Physics, 161(5):1163–1202, December 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1366-y.

Dorofeyev:2016:DST

[Dor16] Illarion Dorofeyev. Dynamics and stationarity of two coupled arbitrary oscillators interacting with separate reservoirs. Journal of Statistical Physics, 162(1):218–231, January 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1394-7.

Dybalski:2014:CSM

[DP14] W. Dybalski and A. Pizzo. Coulomb scattering in the massless Nelson model I. Foundations of two-electron scattering. Journal of Statistical Physics, 154(1–2):543–587, January 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0857-y.

DeSantis:2015:ODI

[DP15] Emilio De Santis and Mauro Piccioni. One-dimensional infinite memory imitation models with noise. Journal of Statistical Physics, 161(2):346–364, October 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1335-5.

DeBievre:2017:RDV

[DP17] S. De Bièvre and P. E. Parris. A rigourous demonstration of the validity of Boltzmann’s scenario for the spatial homogenization of a freely expanding gas and the equilibration of the Kac ring. Journal of Statistical Physics, 168(4):772–793, August 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

DeSantis:2019:IPR

[DP19] Emilio De Santis and Mauro Piccioni. Infinite paths on a random environment of Z^{2} with bounded and recurrent sums. Journal of Statistical Physics, 176(5):1088–1114, September 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Dharmaraja:2017:MCM

[DPT17] Selvamuthu Dharmaraja, Puneet Pasricha, and Paola Tardelli. Markov chain model with catastrophe to determine mean time to default of credit risky assets. Journal of Statistical Physics, 169(4):876–888, November 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Derrida:2013:FSC

[DR13] Bernard Derrida and Martin Retaux. Finite size corrections to the large deviation function of the density in the one dimensional symmetric simple exclusion process.

Journal of Statistical Physics, 152(5):824–845, September 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0797-6.

Derrida:2014:DTP

[DR14] Bernard Derrida and Martin Retaux. The depinning transition in presence of disorder: A toy model. Journal of Statistical Physics, 156(2):268–290, July 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1006-y.

Duque:2019:APC

[DRCV19] Frank Duque, Alejandro Roldán-Correa, and Leon A. Valencia. Accessibility percolation with crossing valleys on n-ary trees. Journal of Statistical Physics, 174(5):1027–1037, March 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Dickman:2010:ARW

[DRS10] Ronald Dickman, Leonardo T. Rolla, and Vladas Sidoravicius. Activated random walkers: Facts, conjectures and challenges. Journal of Statistical Physics, 138(1–3):126–142, February 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-009-9918-7.pdf.

Dietler:2018:HDR

[DRS18] Elia Dietler, Simone Rademacher, and Benjamin Schlein. From Hartree dynamics to the relativistic Vlasov equation. Journal of Statistical Physics, 172(2):398–433, July 2018. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Deijfen:2018:TDD

[DRT18] Maria Deijfen, Sebastian Rosengren, and Pieter Trapman. The tail does not determine the size of the giant. Journal of Statistical Physics, 173(3–4):736–745, November 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-018-2071-4.pdf.

Derrida:2016:LDB

[DS16] Bernard Derrida and Zhan Shi. Large deviations for the branching Brownian motion in presence of selection or coalescence. Journal of Statistical Physics, 163(6):1285–1311, June 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1522-z.

Dolgopyat:2017:TDL

[DS17] Dmitry Dolgopyat and Omri Sarig. Temporal distributional limit theorems for dynamical systems. Journal of Statistical Physics, 166(3–4):680–713, February 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1689-3.

Derrida:2019:LDCa

[DS19a] Bernard Derrida and Tridib Sadhu. Large deviations conditioned on large deviations I: Markov chain and Langevin equation. Journal of Statistical Physics, 176(4):773–805, August 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Derrida:2019:LDCb

[DS19b] Bernard Derrida and Tridib Sadhu. Large deviations conditioned on large deviations II: Fluctuating hydrodynamics. Journal of Statistical Physics, 177(1):151–182, October 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

deSimoi:2017:FSP

[dSLPV17] Jacopo de Simoi, Carlangelo Liverani, Christophe Poquet, and Denis Volk. Fast–slow partially hyperbolic systems versus Freidlin–Wentzell random systems. Journal of Statistical Physics, 166(3–4):650–679, February 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1628-3.

DelRe:2014:AFM

[DSPC14] E. DelRe, S. Di Sabatino, P. Di Porto, and B. Crosignani. Anomalous fluctuations in the motion of partitioning objects. Journal of Statistical Physics, 156(2):291–300, July 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1001-3.

deSantana:2015:EPT

[dSRT15] L. H. de Santana, A. D. Ramos, and A. Toom. Eroders on a plane with three states at a point. Part I: Deterministic. Journal of Statistical Physics, 159(5):1175–1195, June 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1226-9.

DaiPra:2015:FML

[DSS15] Paolo Dai Pra, Benedetto Scoppola, and Elisabetta Scoppola. Fast mixing for the low temperature 2D Ising model through irreversible parallel dynamics. Journal of Statistical Physics, 159(1):1–20, April 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1180-y.

Dirr:2017:HLC

[DSZ17] Nicolas Dirr, Marios G. Stamatakis, and Johannes Zimmer. Hydrodynamic limit of condensing two-species zero range processes with sub-critical initial profiles. Journal of Statistical Physics, 168(4):794–825, August 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-017-1827-6.pdf.

Duki:2018:SRF

[DT18] Solomon Fekade Duki and Mesfin Asfaw Taye. Stochastic resonance and first arrival time for excitable systems. Journal of Statistical Physics, 171(5):878–896, June 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Dimarco:2019:KMA

[DT19] Giacomo Dimarco and Giuseppe Toscani. Kinetic modeling of alcohol consumption. Journal of Statistical Physics, 177(5): 1022–1042, December 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

DePolsi:2019:CIV

[DTW19] Gonzalo De Polsi, Matthieu Tissier, and Nicolás Wschebor. Conformal invariance and vector operators in the O(N) model. Journal of Statistical Physics, 177(6):1089–1130, December 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Dudynski:2013:SPL

[Dud13] Marek Dudyński. Spectral properties of the linearized Boltzmann operator in L^{p} for 1 ≤ p ≤ ∞. Journal of Statistical Physics, 153(6):1084–1106, December 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-013-0873-y.pdf.

DeRosis:2015:RVL

[DUU15] Alessandro De Rosis, Francesco Ubertini, and Stefano Ubertini. The role of very low–Reynolds hydrodynamics on the transfer of information among active agents. Journal of Statistical Physics, 161(6):1390–1403, December 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1312-z;   http://link.springer.com/
content/pdf/10.1007/s10955-015-1312-z.pdf.

Dommers:2010:DPA

[DvdHH10] Sander Dommers, Remco van der Hofstad, and Gerard Hooghiemstra. Diameters in preferential attachment models. Journal of Statistical Physics, 139(1):72–107, April 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-010-9921-z.pdf.

Dhara:2016:GRS

[DvLM16] Souvik Dhara, Johan S. H. van Leeuwaarden, and Debankur Mukherjee. Generalized random sequential adsorption on Erdős–Rényi random graphs. Journal of Statistical Physics, 164(5):1217–1232, September 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-016-1583-z.pdf.

Dhara:2018:CMF

[DvLM18] Souvik Dhara, Johan S. H. van Leeuwaarden, and Debankur Mukherjee. Corrected mean-field model for random sequential adsorption on random geometric graphs. Journal of Statistical Physics, 173(3–4):872–894, November 2018. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-018-2019-8.pdf.

daVeiga:2015:SCL

[dVO15] Paulo A. Faria da Veiga and Michael O’Carroll. Stability and clustering for lattice many–body quantum Hamiltonians with multiparticle potentials. Journal of Statistical Physics, 161(3):712–720, November 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1343-5.

deWoul:2010:PGF

[dWL10] Jonas de Woul and Edwin Langmann. Partially gapped fermions in 2D. Journal of Statistical Physics, 139(6):1033– 1065, June 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-9971-2.

deWoul:2014:GIC

[dWL14] Jonas de Woul and Edwin Langmann. Gauge invariance, correlated fermions, and photon mass in 2 + 1 dimensions. Journal of Statistical Physics, 154(3):877–894, February 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0912-8.

Deng:2016:ETA

[DWTW16] Weihua Deng, Wanli Wang, Xinchun Tian, and Yujiang Wu. Effects of the tempered aging and the corresponding Fokker– Planck equation. Journal of Statistical Physics, 164(2):377– 398, July 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1547-3.

Dong:2014:EES

[DXZ14] Zhao Dong, Lihu Xu, and Xicheng Zhang. Exponential ergodicity of stochastic Burgers equations driven by α-stable processes. Journal of Statistical Physics, 154(4):929–949, February 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0881-y.

Dykman:2014:MGB

[Dyk14] Mark Dykman. Moshe Gitterman, Phase Transitions: Modern Applications. Journal of Statistical Physics, 156(5):1025–1026, September 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1055-2.

Dymov:2015:NSM

[Dym15] A. Dymov. Nonequilibrium statistical mechanics of Hamiltonian rotators with alternated spins. Journal of Statistical Physics, 158(4):968–1006, February 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1144-2.

Dembo:2015:MOU

[DZ15] Amir Dembo and Ofer Zeitouni. Matrix optimization under random external fields. Journal of Statistical Physics, 159 (6):1306–1326, June 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1228-7.

deZarate:2011:HFL

[dZS11] J. M. Ortiz de Zárate and J. V. Sengers. Hydrodynamic fluctuations in laminar fluid flow. I. Fluctuating Orr–Sommerfeld equation. Journal of Statistical Physics, 144(4):774–792, August 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0256-1.

deZarate:2013:HFL

[dZS13] José M. Ortiz de Zárate and Jan V. Sengers. Hydrodynamic fluctuations in laminar fluid flow. II. Fluctuating Squire equation. Journal of Statistical Physics, 150(3):540–558, February 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0495-9.

Dzubiella:2011:HIG

[Dzu11] Joachim Dzubiella. How interface geometry dictates Water’s thermodynamic signature in hydrophobic association. Journal of Statistical Physics, 145(2):227–239, October 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0217-8.

Elkin:2012:EFS

[EAL12] Michael Elkin, Ingemar Andre, and David B. Lukatsky. Energy fluctuations shape free energy of nonspecific biomolecular interactions. Journal of Statistical Physics, 146(4):870–877, February 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0421-1.

Enting:2011:SAK

[EC11] I. G. Enting and N. Clisby. Series analysis of a Kosterlitz– Thouless transition: The 6-state planar Potts model. Journal of Statistical Physics, 145(3):696–712, November 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0322-8.

Eyink:2015:SSA

[ED15] Gregory L. Eyink and Theodore D. Drivas. Spontaneous stochasticity and anomalous dissipation for Burgers equation. Journal of Statistical Physics, 158(2):386–432, January 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1135-3.

Erdos:2013:LED

[EF13] László Erdős and Brendan Farrell. Local eigenvalue density for general MANOVA matrices. Journal of Statistical Physics, 152(6):1003–1032, September 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0807-8.

Eldon:2018:GPS

[EF18] Bjarki Eldon and Fabian Freund. Genealogical properties of subsamples in highly fecund populations. Journal of Statistical Physics, 172(1):175–207, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Egli:2011:ALT

[EFO11] Daniel Egli, Jürg Fröhlich, and Hans-Rudolf Ott. Anderson localization triggered by spin disorder — with an application to Eu_{x}Ca_{1−x}B_{6}. Journal of Statistical Physics, 143(5):970– 989, June 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0216-9.

Erdos:2012:ALB

[EH12] László Erdős and David Hasler. Anderson localization at band edges for random magnetic fields. Journal of Statistical Physics, 146(5):900–923, March 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-012-0445-
6.

Eshkabilov:2012:NUG

[EHR12] Y. K. Eshkabilov, F. H. Haydarov, and U. A. Rozikov. Nonuniqueness of Gibbs measure for models with uncountable set of spin values on a Cayley tree. Journal of Statistical Physics, 147(4):779–794, June 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0494-x.

Einav:2012:CEC

[Ein12] Amit Einav. A counter example to Cercignani’s conjecture for the d dimensional Kac model. Journal of Statistical Physics, 148(6):1076–1103, September 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0565-z.

Edlund:2010:RCA

[EJ10] E. Edlund and M. Nilsson Jacobi. Renormalization of cellular automata and self-similarity. Journal of Statistical Physics, 139(6):972–984, June 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9974-z.

Ermolaev:2010:LTD

[EK10] Victor Ermolaev and Christof Külske. Low-temperature dynamics of the Curie–Weiss model: Periodic orbits, multiple histories, and loss of Gibbsianness. Journal of Statistical Physics, 141(5):727–756, December 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-010-0074-x.pdf.

Esfahani:2019:ITH

[EK19] Reihaneh Kouhi Esfahani and Norbert Kern. Interpreting traffic on a highway with on/off ramps in the light of TASEP. Journal of Statistical Physics, 177(4):588–607, November 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ertas:2012:MDP

[EKD12] Mehmet Ertas, Mustafa Keskin, and Bayram Deviren. Multicritical dynamic phase diagrams and dynamic hysteresis loops in a mixed spin-2 and spin-5/2 Ising ferrimagnetic system with repulsive biquadratic coupling: Glauber dynamic approach. Journal of Statistical Physics, 146(6):1244–1262, March 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0441-x.

Ellis:2012:CGF

[EL12] Richard S. Ellis and Jingran Li. Conditional Gaussian fluctuations and refined asymptotics of the spin in the phasecoexistence region. Journal of Statistical Physics, 149(5): 803–830, November 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0620-9.

Essam:2011:MLF

[ELO11] J. W. Essam, H. Lonsdale, and A. L. Owczarek. Mean length of finite clusters in directed compact percolation near a damp wall. Journal of Statistical Physics, 145(3):639–646, November 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0317-5.

Elskens:2012:GCS

[Els12] Yves Elskens. Gaussian convergence for stochastic acceleration of particles in the dense spectrum limit. Journal of Statistical Physics, 148(3):591–605, August 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0546-2.

Elsey:2015:SST

[Els15] Matt Elsey. Self-similarity in two-phase curvature flow. Journal of Statistical Physics, 159(4):905–914, May 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1203-3.

Erignoux:2018:SSB

[ELX18] C. Erignoux, C. Landim, and T. Xu. Stationary states of boundary driven exclusion processes with nonreversible boundary dynamics. Journal of Statistical Physics, 171(4): 599–631, May 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Escobedo:2010:SBA

[EM10] Miguel Escobedo and Stéphane Mischler. Scalings for a ballistic aggregation equation. Journal of Statistical Physics, 141(3):422–458, November 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0060-3.

Eckhoff:2018:NCP

[EMO18] Maren Eckhoff, Peter Mörters, and Marcel Ortgiese. Near critical preferential attachment networks have small giant components. Journal of Statistical Physics, 173(3–4):663–703, November 2018. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-018-2054-5.pdf.

El-Nabulsi:2018:PIF

[EN18] Rami Ahmad El-Nabulsi. Path integral formulation of fractionally perturbed Lagrangian oscillators on fractal. Journal of Statistical Physics, 172(6):1617–1640, September 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Einstein:2014:DSI

[EP14] T. L. Einstein and Alberto Pimpinelli. Dynamical scaling implications of Ferrari, Prähofer, and Spohn’s remarkable spatial scaling results for facet–edge fluctuations. Journal of Statistical Physics, 155(6):1178–1190, June 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0981-3.

Elgart:2017:LBP

[EPS17] A. Elgart, L. Pastur, and M. Shcherbina. Large block properties of the entanglement entropy of free disordered fermions. Journal of Statistical Physics, 166(3–4):1092–1127, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1656-z.

Erignoux:2018:HLB

[Eri18] C. Erignoux. Hydrodynamic limit of boundary driven exclusion processes with nonreversible boundary dynamics. Journal of Statistical Physics, 172(5):1327–1357, September 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Eroglu:2014:MAL

[Ero14] Sertac Eroglu. Menzerath–Altmann law: Statistical mechanical interpretation as applied to a linguistic organization. Journal of Statistical Physics, 157(2):392–405, October 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1078-8.

Eliazar:2012:SFC

[ES12] Iddo I. Eliazar and Michael F. Shlesinger. Stochastic flow cascades. Journal of Statistical Physics, 146(1):1–24, January 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0340-6.

E:2013:ERG

[ES13] Weinan E and Hao Shen. Exact renormalization group analysis of turbulent transport by the shear flow. Journal of Statistical Physics, 153(4):553–571, November 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0854-1.

Esler:2017:SPV

[Esl17] J. G. Esler. Statistics of point vortex turbulence in nonneutral flows and in flows with translational and rotational symmetries. Journal of Statistical Physics, 169(6):1045–1065, December 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

El-Showk:2014:SIM

[ESPP^{+}14] Sheer El-Showk, Miguel F. Paulos, David Poland, Slava Rychkov, David Simmons-Duffin, and Alessandro Vichi. Solving the 3D Ising model with the conformal bootstrap II. c-Minimization and precise critical exponents. Journal of Statistical Physics, 157(4–5):869–914, December 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1042-7.

Evenbly:2011:TNS

[EV11] G. Evenbly and G. Vidal. Tensor network states and geometry. Journal of Statistical Physics, 145(4):891–918, November 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0237-4.

Evenbly:2014:AER

[EV14] G. Evenbly and G. Vidal. Algorithms for entanglement renormalization: Boundaries, impurities and interfaces. Journal of Statistical Physics, 157(4–5):931–978, December 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0983-1.

Evans:2016:NES

[Eva16] Josephine Evans. Non-equilibrium steady states in Kac’s model coupled to a thermostat. Journal of Statistical Physics, 164(5):1103–1121, September 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-016-1581-1.
pdf.

Evans:2016:TNS

[EWSR16] Denis J. Evans, Stephen R. Williams, Debra J. Searles, and Lamberto Rondoni. On typicality in nonequilibrium steady states. Journal of Statistical Physics, 164(4):842–857, August 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1563-3.

Francois:2016:CAL

[FAB16] Paul François and Grégoire Altan-Bonnet. The case for absolute ligand discrimination: Modeling information processing and decision by immune T cells. Journal of Statistical Physics, 162(5):1130–1152, March 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-015-1444-1.pdf.

Fujie:2013:MCA

[FAM13] Ryo Fujie, Kazuyuki Aihara, and Naoki Masuda. A model of competition among more than two languages. Journal of Statistical Physics, 151(1–2):289–303, April 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-012-0613-8.pdf.

Fantoni:2016:EER

[Fan16a] Riccardo Fantoni. Erratum to: Exact results for one dimensional fluids through functional integration. Journal of Statistical Physics, 163(6):1504, June 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-016-1519-
7;   http://link.springer.com/content/pdf/10.1007/
s10955-016-1519-7.pdf. See [Fan16b].

Fantoni:2016:ERO

[Fan16b] Riccardo Fantoni. Exact results for one dimensional fluids through functional integration. Journal of Statistical Physics, 163(5):1247–1267, June 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1510-3. See

erratum [Fan16a].

Fantoni:2017:AWC

[Fan17a] Riccardo Fantoni. Andersen–Weeks–Chandler perturbation theory and one-component sticky–hard-spheres. Journal of Statistical Physics, 168(3):652–665, August 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Fantoni:2017:ODFa

[Fan17b] Riccardo Fantoni. One-dimensional fluids with positive potentials. Journal of Statistical Physics, 166(5):1334–1342, March 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1707-5.

Faruk:2015:UST

[Far15] Mir Mehedi Faruk. Unified statistical thermodynamics of quantum gases trapped under generic power law potential in d dimension and their equivalence in d = 1. Journal of Statistical Physics, 161(3):679–687, November 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1344-4.

Feng:2011:IAF

[FBE^{+}11] Zhengzheng Feng, Adam Bymaster, Chris Emborsky, Deepti Ballal, Bennett Marshall, Kai Gong, Alejandro Garcia, Kenneth R. Cox, and Walter G. Chapman. Insights into associating fluid properties and microstructure from classical density functional theory. Journal of Statistical Physics, 145(2):467– 480, October 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0263-2.

Farnell:2019:NCM

[FBR19] D. J. J. Farnell, R. F. Bishop, and J. Richter. Noncoplanar model states in quantum magnetism applications of the high-order coupled cluster method. Journal of Statistical Physics, 176(1):180–213, July 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s10955-019-
02297-1.pdf.

Fedele:2011:SLM

[FC11] Micaela Fedele and Pierluigi Contucci. Scaling limits for multi-species statistical mechanics mean-field models. Journal of Statistical Physics, 144(6):1186–1205, September 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0334-4.

Fok:2013:IAD

[FC13] Pak-Wing Fok and Tom Chou. Identifiability of age-dependent branching processes from extinction probabilities and number distributions. Journal of Statistical Physics, 152(4):769–786, August 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0781-1.

Forman:2017:MAP

[FC17] Yakir Forman and Maria Cameron. Modeling aggregation processes of Lennard-Jones particles via stochastic networks. Journal of Statistical Physics, 168(2):408–433, July 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Frapolli:2015:SHB

[FCK15] Nicolò Frapolli, Shyam S. Chikatamarla, and Ilya Karlin. Simulations of heated bluff-bodies with the multi-speed entropic lattice Boltzmann method. Journal of Statistical Physics, 161(6):1434–1452, December 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-015-1373-
z;   http://link.springer.com/content/pdf/10.1007/
s10955-015-1373-z.pdf.

Fushing:2014:BUB

[FCLK14] Hsieh Fushing, Chen Chen, Shan-Yu Liu, and Patrice Koehl. Bootstrapping on undirected binary networks via statistical mechanics. Journal of Statistical Physics, 156(5):823–842, September 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1043-6.

Fennell:2011:PMA

[FD11] Christopher J. Fennell and Ken A. Dill. Physical modeling of aqueous solvation. Journal of Statistical Physics, 145(2):209– 226, October 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0232-9.

Fyodorov:2014:TTL

[FD14] Yan V. Fyodorov and Pierre Le Doussal. Topology trivialization and large deviations for the minimum in the simplest random optimization. Journal of Statistical Physics, 154(1–2):466–490, January 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0838-1.

Furtlehner:2016:CBC

[FD16] Cyril Furtlehner and Aurélien Decelle. Cycle-based cluster variational method for direct and inverse inference. Journal of Statistical Physics, 164(3):531–574, August 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1566-0.

Fernandez:2014:VDG

[FdHM14] R. Fernández, F. den Hollander, and J. Martínez. Variational description of Gibbs–non–Gibbs dynamical transitions for spin-flip systems with a Kac-type interaction. Journal of Statistical Physics, 156(2):203–220, July 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1004-0.

Fyodorov:2012:CFF

[FDR12] Yan V. Fyodorov, Pierre Le Doussal, and Alberto Rosso. Counting function fluctuations and extreme value threshold in multifractal patterns: The case study of an ideal 1/f noise. Journal of Statistical Physics, 149(5):898–920, November 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0623-6.

Federbush:2013:AED

[Fed13] Paul Federbush. Asymptotic expansions for λ_{d} of the dimer and monomer–dimer problems. Journal of Statistical Physics, 150(3):487–490, February 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0540-8.

Fedele:2014:RMC

[Fed14] Micaela Fedele. Rescaled magnetization for critical bipartite Curie–Weiss models. Journal of Statistical Physics, 155 (2):223–236, April 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-0960-8.

Federbush:2017:MCE

[Fed17] Paul Federbush. A mysterious cluster expansion associated to the expectation value of the permanent of 0–1 matrices. Journal of Statistical Physics, 167(6):1489–1495, June 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Fernandez:2014:BEE

[Fer14] Bastien Fernandez. Breaking of ergodicity in expanding systems of globally coupled piecewise affine circle maps. Journal of Statistical Physics, 154(4):999–1029, February 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0903-9.

Ferrari:2018:FGD

[Fer18] P. L. Ferrari. Finite GUE distribution with cut-off at a shock. Journal of Statistical Physics, 172(2):505–521, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ferrari:2010:PCB

[FF10] Patrik L. Ferrari and René Frings. On the partial connection between random matrices and interacting particle systems. Journal of Statistical Physics, 141(4):613–637, November 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-0070-1.

Federbush:2011:AER

[FF11a] Paul Federbush and Shmuel Friedland. An asymptotic expansion and recursive inequalities for the monomer–dimer problem. Journal of Statistical Physics, 143(2):306–325, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0170-6.

Ferrari:2011:FTC

[FF11b] Patrik L. Ferrari and René Frings. Finite time corrections in KPZ growth models. Journal of Statistical Physics, 144 (6):1123–1150, September 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0318-4.

Fleming:2011:IPS

[FF11c] Benjamin J. Fleming and Peter J. Forrester. Interlaced particle systems and tilings of the Aztec diamond. Journal of Statistical Physics, 142(3):441–459, February 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0121-2.

Ferrari:2014:PGM

[FF14] Patrik L. Ferrari and René Frings. Perturbed GUE minor process and Warren’s process with drifts. Journal of Statistical Physics, 154(1–2):356–377, January 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0887-5.

Freitas:2011:EVL

[FFT11] Ana Cristina Moreira Freitas, Jorge Milhazes Freitas, and Mike Todd. Extreme value laws in dynamical systems for non-smooth observations. Journal of Statistical Physics, 142 (1):108–126, January 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0096-4.

Frank:2011:LLN

[FG11] Jason Frank and Georg A. Gottwald. The Langevin limit of the Nosé-Hoover–Langevin thermostat. Journal of Statistical Physics, 143(4):715–724, May 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0203-1.

Franco:2012:PSE

[FG12] Tertuliano Franco and Pablo Groisman. A particle system with explosions: Law of large numbers for the density of particles and the blow-up time. Journal of Statistical Physics, 149(4):629–642, November 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0621-8.

Falkovich:2014:THP

[FG14] Gregory Falkovich and Krzysztof Gawȩdzki. Turbulence on hyperbolic plane: The fate of inverse cascade. Journal of Statistical Physics, 156(1):10–54, July 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0995-x.

Ferrari:2018:PTI

[FGGL18] P. A. Ferrari, A. Galves, I. Grigorescu, and E. Löcherbach. Phase transition for infinite systems of spiking neurons. Journal of Statistical Physics, 172(6):1564–1575, September 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Farago:2014:FDR

[FGJ14] Oded Farago and Niels Grønbech-Jensen. Fluctuation– dissipation relation for systems with spatially varying friction. Journal of Statistical Physics, 156(6):1093–1110, September 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1045-4.

Franco:2014:OTE

[FGN14] Tertuliano Franco, Patrícia Gonçalves, and Adriana Neumann. Occupation time of exclusion processes with conductances. Journal of Statistical Physics, 156(5):975–997, September 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1039-2.

Florescu:2015:HDF

[FGP15] Laura Florescu, Shirshendu Ganguly, and Yuval Peres. Heat diffusion with frozen boundary. Journal of Statistical Physics, 161(3):521–531, November 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1342-6.

Freidlin:2011:PGL

[FH11] Mark Freidlin and Wenqing Hu. On perturbations of generalized Landau–Lifshitz dynamics. Journal of Statistical Physics, 144(5):978–1008, September 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0289-5.

Freidlin:2013:DNR

[FH13] Mark Freidlin and Wenqing Hu. On diffusion in narrow random channels. Journal of Statistical Physics, 152(1):136– 158, July 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0763-3.

Fidaleo:2015:HAI

[Fid15] Francesco Fidaleo. Harmonic analysis on inhomogeneous amenable networks and the Bose–Einstein condensation. Journal of Statistical Physics, 160(3):715–759, August 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1263-4.

Filatov:2016:MIC

[Fil16] Denis M. Filatov. A method for identification of critical states of open stochastic dynamical systems based on the analysis of acceleration. Journal of Statistical Physics, 165(4):681–692, November 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1641-6.

Fisher:2011:CDP

[Fis11] Michael E. Fisher. Cyril Domb: A personal view and appreciation. Journal of Statistical Physics, 145(3):510–517, November 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/
pdf/10.1007/s10955-011-0381-x.pdf. See [Fis12].

Fisher:2012:ECD

[Fis12] Michael E. Fisher. Erratum to: Cyril Domb: A Personal View and Appreciation. Journal of Statistical Physics, 146 (4):883, February 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-012-0433-x.

pdf. See [Fis11].

Fort:2018:CEA

[FJLS18] G. Fort, B. Jourdain, T. Lelièvre, and G. Stoltz. Convergence and efficiency of adaptive importance sampling techniques with partial biasing. Journal of Statistical Physics, 171 (2):220–268, April 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Fleurke:2010:MPS

[FK10] S. R. Fleurke and C. Külske. Multilayer parking with screening on a random tree. Journal of Statistical Physics, 139(3):417– 431, May 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-9935-6.

Frohlich:2011:MDT

[FK11] Jürg Fröhlich and Antti Knowles. A microscopic derivation of the time-dependent Hartree–Fock equation with Coulomb two-body interaction. Journal of Statistical Physics, 145 (1):23–50, October 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0311-y.

Franke:2012:EAT

[FK12] Jasper Franke and Joachim Krug. Evolutionary accessibility in tunably rugged fitness landscapes. Journal of Statistical Physics, 148(4):706–723, September 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0507-9.

Freidlin:2017:MDM

[FK17] M. Freidlin and L. Koralov. Metastable distributions of Markov chains with rare transitions. Journal of Statistical Physics, 167(6):1355–1375, June 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Foss:2018:LPR

[FK18a] Sergey Foss and Takis Konstantopoulos. Limiting properties of random graph models with vertex and edge weights. Journal of Statistical Physics, 173(3–4):626–643, November 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-018-2080-3.pdf.

Freidlin:2018:FPR

[FK18b] M. Freidlin and L. Koralov. Front propagation for reaction– diffusion equations in composite structures. Journal of Statistical Physics, 172(6):1663–1681, September 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Friesen:2018:SAP

[FK18c] Martin Friesen and Yuri Kondratiev. Stochastic averaging principle for spatial birth-and-death evolutions in the continuum. Journal of Statistical Physics, 171(5):842–877, June 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Finkelshtein:2010:VSS

[FKK10] Dmitri Finkelshtein, Yuri Kondratiev, and Oleksandr Kutoviy. Vlasov scaling for stochastic dynamics of continuous systems. Journal of Statistical Physics, 141(1):158–178, October 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-0038-1.

Finkelshtein:2015:DWR

[FKKO15] Dmitri Finkelshtein, Yuri Kondratiev, Oleksandr Kutoviy, and Maria João Oliveira. Dynamical Widom–Rowlinson model and its mesoscopic limit. Journal of Statistical Physics, 158 (1):57–86, January 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1124-6.

Fan:2016:MRK

[FKLL16] Yuwei Fan, Julian Koellermeier, Jun Li, and Ruo Li. Model reduction of kinetic equations by operator projection. Journal of Statistical Physics, 162(2):457–486, January 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1384-9.

Formentin:2012:MMF

[FKR12] M. Formentin, C. Külske, and A. Reichenbachs. Metastates in mean-field models with random external fields generated by Markov chains. Journal of Statistical Physics, 146(2):314–329, January 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0391-8.

Farias:2012:SAR

[FL12] C. F. Farias and Marcelo M. Leite. Susceptibility amplitude ratio for generic competing systems. Journal of Statistical Physics, 148(5):972–980, September 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0558-y.

Field:2013:RRS

[FL13] Laurence S. Field and Gregory F. Lawler. Reversed radial SLE and the Brownian loop measure. Journal of Statistical Physics, 150(6):1030–1062, March 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0729-5.

Forrester:2014:LCL

[FL14] Peter J. Forrester and Joel L. Lebowitz. Local central limit theorem for determinantal point processes. Journal of Statistical Physics, 157(1):60–69, October 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1071-2.

Fan:2015:GST

[FL15a] Yingzhe Fan and Yuanjie Lei. Global solutions and time decay of the non-cutoff Vlasov–Maxwell–Boltzmann system in the whole space. Journal of Statistical Physics, 161(5):1059–1097, December 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1380-0.

Forrester:2015:RDR

[FL15b] Peter J. Forrester and Dang-Zheng Liu. Raney distributions and random matrix theory. Journal of Statistical Physics, 158(5):1051–1082, March 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1150-4.

Friedman:2016:SEE

[FL16a] B. A. Friedman and G. C. Levine. Scaling of entanglement entropy for the Heisenberg model on clusters joined by point contacts. Journal of Statistical Physics, 165(4):727–739, November 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1640-7.

Fyodorov:2016:MPM

[FL16b] Yan V. Fyodorov and Pierre Le Doussal. Moments of the position of the maximum for GUE characteristic polynomials and for log-correlated Gaussian processes. Journal of Statistical Physics, 164(1):190–240, July 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-016-1536-6.pdf.

Foxall:2018:CTF

[FL18] Eric Foxall and Hanbaek Lyu. Clustering in the three and four color cyclic particle systems in one dimension. Journal of Statistical Physics, 171(3):470–483, May 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). See correction [FL20].

Foxall:2020:CCT

[FL20] Eric Foxall and Hanbaek Lyu. Correction to: Clustering in the Three and Four Color Cyclic Particle Systems in One Dimension. Journal of Statistical Physics, 181(5): 2011–2014, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-020-02651-

8.pdf. See [FL18].

Fey:2010:GRE

[FLP10] Anne Fey, Lionel Levine, and Yuval Peres. Growth rates and explosions in sandpiles. Journal of Statistical Physics, 138(1–3):143–159, February 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-009-9899-6.
pdf.

Furtlehner:2012:ODP

[FLS12] Cyril Furtlehner, Jean-Marc Lasgouttes, and Maxim Samsonov. One-dimensional particle processes with Acceleration/ braking asymmetry. Journal of Statistical Physics, 147 (6):1113–1144, July 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0521-y.

Faranda:2011:NCB

[FLTV11] Davide Faranda, Valerio Lucarini, Giorgio Turchetti, and Sandro Vaienti. Numerical convergence of the block-maxima approach to the generalized extreme value distribution. Journal of Statistical Physics, 145(5):1156–1180, December 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0234-7.

Foss:2018:CST

[FM18] Sergey Foss and Masakiyo Miyazawa. Customer sojourn time in GI/GI/1 feedback queue in the presence of heavy tails. Journal of Statistical Physics, 173(3–4):1195–1226, November 2018. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-018-2079-9.pdf.

Family:2011:DSL

[FMAG11] Fereydoon Family, K. I. Mazzitello, C. M. Arizmendi, and H. E. Grossniklaus. Dynamic scaling of lipofuscin deposition in aging cells. Journal of Statistical Physics, 144(2):332– 343, July 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0178-y.

Fontes:2014:PTL

[FMM^{+}14] Luiz Renato Fontes, Domingos H. U. Marchetti, Immacolata Merola, Errico Presutti, and Maria Eulalia Vares. Phase transitions in layered systems. Journal of Statistical Physics, 157(3):407–421, November 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1090-z.

Fontes:2015:LSM

[FMM^{+}15] Luiz Renato Fontes, Domingos H. U. Marchetti, Immacolata Merola, Errico Presutti, and Maria Eulalia Vares. Layered systems at the mean field critical temperature. Journal of Statistical Physics, 161(1):91–122, October 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1307-9.

Fytas:2018:RRD

[FMMPS18] Nikolaos G. Fytas, Víctor Martín-Mayor, Marco Picco, and Nicolas Sourlas. Review of recent developments in the randomfield Ising model. Journal of Statistical Physics, 172(2):665– 672, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Fortunato:2013:E

[FMR13] Santo Fortunato, Michael Macy, and Sidney Redner. Editorial. Journal of Statistical Physics, 151(1–2):1–8, April 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-013-0703-2.pdf.

Fu:2013:GMC

[FN13] Feng Fu and Martin A. Nowak. Global migration can lead to stronger spatial selection than local migration. Journal of Statistical Physics, 151(3–4):637–653, May 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0631-6.

Ferrari:2015:SFF

[FN15a] Patrik L. Ferrari and Peter Nejjar. Shock fluctuations in flat TASEP under critical scaling. Journal of Statistical Physics, 160(4):985–1004, August 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1208-y.

Fyodorov:2015:RMA

[FN15b] Y. V. Fyodorov and A. Nock. On random matrix averages involving half-integer powers of GOE characteristic polynomials. Journal of Statistical Physics, 159(4):731–751, May 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-015-1209-x.pdf.

Ferreira:2018:SGM

[FO18] Giovane Ferreira and Krerley Oliveira. Sequential Gibbs measures and factor maps. Journal of Statistical Physics, 172(3): 833–853, August 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Fortunato:2011:RCS

[For11] Santo Fortunato. Reuven Cohen and Shlomo Havlin: Complex Networks. Journal of Statistical Physics, 142(3):640–641, February 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0129-7.

Forrester:2013:LEP

[For13] Peter J. Forrester. Lyapunov exponents for products of complex Gaussian random matrices. Journal of Statistical Physics, 151(5):796–808, June 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0735-7.

Fowler:2019:PTB

[Fow19] A. C. Fowler. Phase transition in the Boltzmann–Vlasov equation. Journal of Statistical Physics, 174(5):1011–1026, March 2019. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-019-02222-6.pdf.

Faggionato:2011:GCT

[FP11] Alessandra Faggionato and Daniele Di Pietro. Gallavotti– Cohen-type symmetry related to cycle decompositions for Markov chains and biochemical applications. Journal of Statistical Physics, 143(1):11–32, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-011-0161-

Franceschetti:2011:SIC

[FPR11] Massimo Franceschetti, Mathew D. Penrose, and Tom Rosoman. Strict inequalities of critical values in continuum percolation. Journal of Statistical Physics, 142(3):460–486, February 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-011-0122-1.pdf.

Frohlich:2017:CPC

[FR17] Jürg Fröhlich and Pierre-François Rodríguez. On cluster properties of classical ferromagnets in an external magnetic field. Journal of Statistical Physics, 166(3–4):828–840, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1556-2.

Franzosi:2011:MED

[Fra11] Roberto Franzosi. Microcanonical entropy and dynamical measure of temperature for systems with two first integrals. Journal of Statistical Physics, 143(4):824–830, May 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0200-4.

Fradkin:2017:DOT

[Fra17a] Eduardo Fradkin. Disorder operators and their descendants. Journal of Statistical Physics, 167(3–4):427–461, May 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Franceschetti:2017:QLE

[Fra17b] Massimo Franceschetti. Quantum limits on the entropy of bandlimited radiation. Journal of Statistical Physics, 169(2): 374–394, October 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Freidlin:2014:SPD

[Fre14] M. Freidlin. On stochastic perturbations of dynamical systems with a “Rough” symmetry. hierarchy of Markov chains. Journal of Statistical Physics, 157(6):1031–1045, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1110-z.

Freire:2015:AMD

[Fre15] Marcelo Ventura Freire. Application of moderate deviation techniques to prove Sinai theorem on RWRE. Journal of Statistical Physics, 160(2):357–370, July 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1266-1.

Freidlin:2016:TLP

[Fre16] Mark Freidlin. Thermostat-like perturbations of an oscillator. Journal of Statistical Physics, 164(1):130–141, July 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1531-y.

Freitas:2017:ABE

[Fre17] Pedro Freitas. Asymptotic behaviour of extremal averages of Laplacian eigenvalues. Journal of Statistical Physics, 167(6): 1511–1518, June 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Friesen:2017:NED

[Fri17] Martin Friesen. Non-equilibrium dynamics for a Widom– Rowlinson type model with mutations. Journal of Statistical Physics, 166(2):317–353, January 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-016-1699-
1.

Foias:2015:CTA

[FRT15] Ciprian Foias, Ricardo M. S. Rosa, and Roger M. Temam. Convergence of time averages of weak solutions of the threedimensional Navier–Stokes equations. Journal of Statistical Physics, 160(3):519–531, August 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-015-1248-
3.

Forrester:2011:GPI

[FS11a] Peter J. Forrester and Christopher D. Sinclair. A generalized plasma and interpolation between classical random matrix ensembles. Journal of Statistical Physics, 143(2):326–345, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0173-3.

Fukuizumi:2011:BSN

[FS11b] Reika Fukuizumi and Andrea Sacchetti. Bifurcation and stability for nonlinear Schrödinger equations with double well potential in the semiclassical limit. Journal of Statistical Physics, 145(6):1546–1594, December 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0356-y.

Faggionato:2014:DKM

[FS14a] Alessandra Faggionato and Vittoria Silvestri. Discrete kinetic models for molecular motors: Asymptotic velocity and Gaussian fluctuations. Journal of Statistical Physics, 157 (6):1062–1096, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1106-8.

Faupin:2014:MPV

[FS14b] Jérémy Faupin and Israel Michael Sigal. Minimal photon velocity bounds in non-relativistic quantum electrodynamics. Journal of Statistical Physics, 154(1–2):58–90, January 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0862-1.

Fukuizumi:2014:SSN

[FS14c] Reika Fukuizumi and Andrea Sacchetti. Stationary states for nonlinear Schrödinger equations with periodic potentials. Journal of Statistical Physics, 156(4):707–738, August 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1023-x.

Fantoni:2017:ODFb

[FS17] Riccardo Fantoni and Andrés Santos. One-dimensional fluids with second nearest-neighbor interactions. Journal of Statistical Physics, 169(6):1171–1201, December 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Fatkullin:2018:LSG

[FS18] Ibrahim Fatkullin and Valeriy Slastikov. Limit shapes for Gibbs ensembles of partitions. Journal of Statistical Physics, 172(6):1545–1563, September 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Faulhuber:2019:EPH

[FS19] Markus Faulhuber and Stefan Steinerberger. An extremal property of the hexagonal lattice. Journal of Statistical Physics, 177(2):285–298, October 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ferrari:2013:CKP

[FSS13] Patrik L. Ferrari, Tomohiro Sasamoto, and Herbert Spohn. Coupled Kardar–Parisi–Zhang equations in one dimension. Journal of Statistical Physics, 153(3):377–399, November 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0842-5.

Farfan:2010:DLD

[FSV10] Jonathan Farfan, Alexandre B. Simas, and Fábio J. Valentim. Dynamical large deviations for a boundary driven stochastic lattice gas model with many conserved quantities. Journal of Statistical Physics, 139(4):658–685, May 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-9957-0.

Ferre:2018:ASL

[FT18a] Grégoire Ferré and Hugo Touchette. Adaptive sampling of large deviations. Journal of Statistical Physics, 172(6):1525– 1544, September 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). See correction [FT18b].

Ferre:2018:CAS

[FT18b] Grégoire Ferré and Hugo Touchette. Correction to: Adaptive sampling of large deviations. Journal of Statistical Physics, 173(5):1547, December 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-018-2133-7.

pdf. See [FT18a].

Franco:2019:HLS

[FT19a] Tertuliano Franco and Mariana Tavares. Hydrodynamic limit for the SSEP with a slow membrane. Journal of Statistical Physics, 175(2):233–268, April 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Funaki:2019:MMC

[FT19b] Tadahisa Funaki and Kenkichi Tsunoda. Motion by mean curvature from Glauber–Kawasaki dynamics. Journal of Statistical Physics, 177(2):183–208, October 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Funaki:2014:EEU

[Fun14] Tadahisa Funaki. Equivalence of ensembles under inhomogeneous conditioning and its applications to random Young diagrams. Journal of Statistical Physics, 154(1–2):588–609, January 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0841-6.

Felderhof:2014:MNV

[FvB14] Ubbo Felderhof and Henk van Beijeren. In memoriam Nico van Kampen. Journal of Statistical Physics, 154(3):656–659, February 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-013-0902-x.pdf.

Fitzner:2013:NBR

[FvdH13] Robert Fitzner and Remco van der Hofstad. Non-backtracking random walk. Journal of Statistical Physics, 150(2):264–284, January 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0684-6.

Fanelli:2014:RHI

[FVV14] Luca Fanelli, Luis Vega, and Nicola Visciglia. Relativistic Hardy inequalities in magnetic fields. Journal of Statistical Physics, 154(3):866–876, February 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0915-0. See erratum [FVV15].

Fanelli:2015:ERH

[FVV15] Luca Fanelli, Luis Vega, and Nicola Visciglia. Erratum to: Relativistic Hardy Inequalities in Magnetic Fields. Journal of Statistical Physics, 158(6):1413–1414, March 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-014-1171-z.pdf. See [FVV14].

Freidlin:2012:PMC

[FW12] Mark Freidlin and Matthias Weber. Perturbations of the motion of a charged particle in a noisy magnetic field. Journal of Statistical Physics, 147(3):565–581, May 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0484-z.

Flierl:2015:CAI

[FW15] Glenn R. Flierl and Nicholas W. Woods. Copepod aggregations: Influences of physics and collective behavior. Journal of Statistical Physics, 158(3):665–698, February 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1162-0.

Fu:2017:DSM

[FW17] Jingcheng Fu and Jianliang Wu. A deep stochastic model for detecting community in complex networks. Journal of Statistical Physics, 166(2):230–243, January 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1681-y.

Fatriansyah:2016:ECF

[FY16] Jaka Fajar Fatriansyah and Yusril Yusuf. The effect of the chirality on the fluctuation of liquid crystal. Journal of Statistical Physics, 165(1):86–93, October 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1606-9.

Fyodorov:2019:SGM

[Fyo19] Yan V. Fyodorov. A spin glass model for reconstructing nonlinearly encrypted signals corrupted by noise. Journal of Statistical Physics, 175(5):789–818, June 2019. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-018-02217-9.pdf.

Feres:2010:SBL

[FZ10] Renato Feres and Hong-Kun Zhang. The spectrum of the billiard Laplacian of a family of random billiards. Journal of Statistical Physics, 141(6):1039–1054, December 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-0079-5.

Fan:2011:EIF

[FZ11a] Shuangli Fan and Fan Zhong. Evidences of the instability fixed points of first-order phase transitions. Journal of Statistical Physics, 143(6):1136–1153, June 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0225-8.

Feng:2011:LDZ

[FZ11b] Renjie Feng and Steve Zelditch. Large deviations for zeros of P (φ)_{2} random polynomials. Journal of Statistical Physics, 143(4):619–635, May 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0206-y.

Fang:2012:STI

[FZ12] Ming Fang and Ofer Zeitouni. Slowdown for time inhomogeneous branching Brownian motion. Journal of Statistical Physics, 149(1):1–9, October 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0581-z.

Galam:2013:DOV

[Gal13] Serge Galam. The drastic outcomes from voting alliances in three-party democratic voting (1990 → 2013). Journal of Statistical Physics, 151(1–2):46–68, April 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0641-4.

Gallavotti:2014:RGD

[Gal14] Giovanni Gallavotti. Renormalization group and divergences. Journal of Statistical Physics, 157(4–5):743–754, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1027-6.

Galvan:2015:RBM

[Gal15] Bruno Galvan. Relativistic Bohmian mechanics without a preferred foliation. Journal of Statistical Physics, 161(5): 1268–1275, December 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1369-8.

Gallavotti:2017:ADR

[Gal17a] Giovanni Gallavotti. About David Ruelle, after his 80th birthday. Journal of Statistical Physics, 166(3–4):458–462, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1533-9.

Gallavotti:2017:RML

[Gal17b] Giovanni Gallavotti. Random matrices and Lyapunov coefficients regularity. Journal of Statistical Physics, 166(3–4): 558–574, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1429-0.

Gandolfi:2018:RCC

[Gan18] Alberto Gandolfi. Random-cluster correlation inequalities for Gibbs fields. Journal of Statistical Physics, 173(2):249–267, October 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Gao:2018:APH

[Gao18] Peng Gao. Averaging principle for the higher order nonlinear Schrödinger equation with a random fast oscillation. Journal of Statistical Physics, 171(5):897–926, June 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Garrigue:2019:HKT

[Gar19] Louis Garrigue. Hohenberg–Kohn theorems for interactions, spin and temperature. Journal of Statistical Physics, 177(3): 415–437, November 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Gaspard:2016:GDM

[Gas16] Pierre Gaspard. Growth and dissolution of macromolecular Markov chains. Journal of Statistical Physics, 164(1):17– 48, July 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1532-x.

Gora:2016:SSM

[GBL16] Pawel Góra, Abraham Boyarsky, and Zhenyang Li. Singular SRB measures for a non 1–1 map of the unit square. Journal of Statistical Physics, 165(2):409–433, October 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1620-y.

Giner-Baldo:2017:PSN

[GBTL17] Jordi Giner-Baldó, Peter J. Thomas, and Benjamin Lindner. Power spectrum of a noisy system close to a heteroclinic orbit. Journal of Statistical Physics, 168(2):447–469, July 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Geloun:2017:ODG

[GC17] Joseph Ben Geloun and Francesco Caravelli. On an orderingdependent generalization of the Tutte polynomial. Journal of Statistical Physics, 168(5):1105–1124, September 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Garlaschelli:2019:SPO

[GdHMZ19] D. Garlaschelli, F. den Hollander, J. M. Meylahn, and B. Zeegers. Synchronization of phase oscillators on the hierarchical lattice. Journal of Statistical Physics, 174(1):188–218, January 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Garlaschelli:2018:CSB

[GdHR18] Diego Garlaschelli, Frank den Hollander, and Andrea Roccaverde. Covariance structure behind breaking of ensemble equivalence in random graphs. Journal of Statistical Physics, 173(3–4):644–662, November 2018. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-018-2114-x.pdf.

Ginzburg:2010:OSA

[GdK10] Irina Ginzburg, Dominique d’Humières, and Alexander Kuzmin. Optimal stability of advection–diffusion lattice Boltzmann models with two relaxation times for Positive/ negative equilibrium. Journal of Statistical Physics, 139(6):1090– 1143, June 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-9969-9.

Gerschenfeld:2010:AFL

[GDL10] A. Gerschenfeld, B. Derrida, and J. L. Lebowitz. Anomalous Fourier’s law and long range correlations in a 1D non-momentum conserving mechanical model. Journal of Statistical Physics, 141(5):757–766, December 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0076-8.

Gautie:2019:NCB

[GDMS19] Tristan Gautié, Pierre Le Doussal, Satya N. Majumdar, and Grégory Schehr. Non-crossing Brownian paths and Dyson Brownian motion under a moving boundary. Journal of Statistical Physics, 177(5):752–805, December 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Gong:2011:IDT

[GG11a] Min Gong and Feng Guo. Influence of delayed-time and colored-noise in a bistable system subject to asymmetric dichotomous noise and colored noise. Journal of Statistical Physics, 142(2):386–393, January 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-010-0115-
5.

Gravner:2011:ODE

[GG11b] Janko Gravner and David Griffeath. The one-dimensional exactly 1 cellular automaton: Replication, periodicity, and chaos from finite seeds. Journal of Statistical Physics, 142(1):168– 200, January 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-010-0103-9.pdf.

Gallo:2018:DOD

[GG18] Sandro Gallo and Nancy L. Garcia. Discrete one-dimensional coverage process on a renewal process. Journal of Statistical Physics, 173(2):381–397, October 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Giles:2016:CSR

[GGD16] Alexander P. Giles, Orestis Georgiou, and Carl P. Dettmann. Connectivity of soft random geometric graphs over annuli. Journal of Statistical Physics, 162(4):1068–1083, February 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/
pdf/10.1007/s10955-015-1436-1.pdf.

Gallo:2014:RPD

[GGJR14] Sandro Gallo, Nancy L. Garcia, Valdivino Vargas Junior, and Pablo M. Rodríguez. Rumor processes on N and discrete renewal processes. Journal of Statistical Physics, 155(3):591– 602, May 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-0959-1.

Galves:2010:PSC

[GGP10] Antonio Galves, Nancy L. Garcia, and Clémentine Prieur. Perfect simulation of a coupling achieving the d̄-distance between ordered pairs of binary chains of infinite order. Journal of Statistical Physics, 141(4):669–682, November 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-0071-0.

Giardina:2015:QCL

[GGvdHP15] Cristian Giardinà, Claudio Giberti, Remco van der Hofstad, and Maria Luisa Prioriello. Quenched central limit theorems for the Ising model on random graphs. Journal of Statistical Physics, 160(6):1623–1657, September 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1302-1.

Golden:2016:EGN

[GH16a] Kenneth I. Golden and Joshuah T. Heath. Erratum to: Generalized nonlinear fluctuation–dissipation relation for the one– component plasma. Journal of Statistical Physics, 163(1): 210, April 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-016-1454-7.pdf.

Golden:2016:GNF

[GH16b] Kenneth I. Golden and Joshuah T. Heath. Generalized nonlinear fluctuation–dissipation relation for the one-component plasma. Journal of Statistical Physics, 162(1):199–217, January 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1395-6.

Gherardi:2010:ESS

[Ghe10] Marco Gherardi. Exact sampling of self-avoiding paths via discrete Schramm–Loewner evolution. Journal of Statistical Physics, 140(6):1115–1129, September 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0031-8.

Gripon:2016:CSS

[GHLV16] Vincent Gripon, Judith Heusel, Matthias Löwe, and Franck Vermet. A comparative study of sparse associative memories. Journal of Statistical Physics, 164(1):105–129, July 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1530-z.

Glatt-Holtz:2017:UEN

[GHMR17] Nathan Glatt-Holtz, Jonathan C. Mattingly, and Geordie Richards. On unique ergodicity in nonlinear stochastic partial differential equations. Journal of Statistical Physics, 166 (3–4):618–649, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1605-x.

Gandolfo:2013:NPT

[GHRR13] D. Gandolfo, F. H. Haydarov, U. A. Rozikov, and J. Ruiz. New phase transitions of the Ising model on Cayley trees. Journal of Statistical Physics, 153(3):400–411, November 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0836-3.

Grassberger:2017:PMC

[GHS17] Peter Grassberger, Marcelo R. Hilário, and Vladas Sidoravicius. Percolation in media with columnar disorder. Journal of Statistical Physics, 168(4):731–745, August 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Gyongyosi:2019:SDM

[GI19] Laszlo Gyongyosi and Sandor Imre. Subcarrier domain of multicarrier continuous–variable quantum key distribution. Journal of Statistical Physics, 177(5):960–983, December 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-019-02404-2.pdf.

Gilbert:2015:FIB

[Gil15] Michael James Gilbert. The fixed irreducible bridge ensemble for self-avoiding walks. Journal of Statistical Physics, 159(2):321–335, April 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1137-1.

Ginsparg:2014:EKG

[Gin14a] Paul Ginsparg. Erratum to: Kenneth G. Wilson: Renormalized after-dinner anecdotes. Journal of Statistical Physics, 157(4–5):1017, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-014-1128-2.

pdf. See [Gin14b].

Ginsparg:2014:KGW

[Gin14b] Paul Ginsparg. Kenneth G. Wilson: Renormalized afterdinner anecdotes. Journal of Statistical Physics, 157(4–5): 610–624, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1092-x. See

erratum [Gin14a].

Gitterman:2012:NTB

[Git12] M. Gitterman. New type of Brownian motion. Journal of Statistical Physics, 146(1):239–243, January 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0398-1.

Gitterman:2014:SOR

[Git14] M. Gitterman. Stability of an oscillator with random mass. Journal of Statistical Physics, 157(2):376–379, October 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1080-1.

Gallavotti:2015:KEH

[GJ15a] Giovanni Gallavotti and Ian Jauslin. Kondo effect in the hierarchical s-d model. Journal of Statistical Physics, 161 (5):1231–1235, December 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1370-2.

Goncalves:2015:ERK

[GJ15b] Patrícia Gonçalves and Milton Jara. The Einstein relation for the KPZ equation. Journal of Statistical Physics, 158(6):1262– 1270, March 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1158-9.

Giuliani:2016:PFM

[GJL16] Alessandro Giuliani, Ian Jauslin, and Elliott H. Lieb. A Pfaffian formula for monomer-dimer partition functions. Journal of Statistical Physics, 163(2):211–238, April 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1484-1.

Gabrielli:2010:DCR

[GJMS10] A. Gabrielli, M. Joyce, B. Marcos, and F. Sicard. A dynamical classification of the range of pair interactions. Journal of Statistical Physics, 141(6):970–989, December 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-0090-x.

Goncalves:2017:SOB

[GJS17] Patrícia Gonçalves, Milton Jara, and Marielle Simon. Second order Boltzmann–Gibbs principle for polynomial functions and applications. Journal of Statistical Physics, 166(1):90– 113, January 2017. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1686-6.

Goldstein:2017:ETI

[GKLS17a] S. Goldstein, T. Kuna, J. L. Lebowitz, and E. R. Speer. Erratum to: Translation Invariant Extensions of Finite Volume Measures. Journal of Statistical Physics, 166(3– 4):783, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-016-1711-9.

pdf. See [GKLS17b].

Goldstein:2017:TIE

[GKLS17b] S. Goldstein, T. Kuna, J. L. Lebowitz, and E. R. Speer. Translation invariant extensions of finite volume measures. Journal of Statistical Physics, 166(3–4):765–782, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1595-8. See erratum [GKLS17a].

Giardina:2011:SRE

[GKLT11] Cristian Giardina, Jorge Kurchan, Vivien Lecomte, and Julien Tailleur. Simulating rare events in dynamical processes. Journal of Statistical Physics, 145(4):787–811, November 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0350-4.

Grinfeld:2012:RDM

[GKW12] Michael Grinfeld, Philip A. Knight, and Andrew R. Wade. Rank-driven Markov processes. Journal of Statistical Physics, 146(2):378–407, January 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0368-7.

Galves:2013:ISI

[GL13] A. Galves and E. Löcherbach. Infinite systems of interacting chains with memory of variable length — a stochastic model for biological neural nets. Journal of Statistical Physics, 151(5):896–921, June 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0733-9.

Gallavotti:2014:ENE

[GL14] Giovanni Gallavotti and Valerio Lucarini. Equivalence of non-equilibrium ensembles and representation of friction in turbulent flows: The Lorenz 96 model. Journal of Statistical Physics, 156(6):1027–1065, September 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1051-6.

Gebert:2016:PLR

[GL16a] Martin Gebert and Marius Lemm. On polynomial Lieb– Robinson bounds for the XY chain in a decaying random field. Journal of Statistical Physics, 164(3):667–679, August 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1558-0.

Grimmett:2016:CSH

[GL16b] Geoffrey R. Grimmett and Zhongyang Li. Critical surface of the hexagonal polygon model. Journal of Statistical Physics, 163(4):733–753, May 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1497-9.

Ghosh:2017:ENR

[GL17a] Subhro Ghosh and Joel Lebowitz. Erratum to: Number Rigidity in Superhomogeneous Random Point Fields. Journal of Statistical Physics, 166(3–4):1028, February 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-016-1712-8.pdf. See [GL17b].

Ghosh:2017:NRS

[GL17b] Subhro Ghosh and Joel Lebowitz. Number rigidity in superhomogeneous random point fields. Journal of Statistical Physics, 166(3–4):1016–1027, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1633-6. See

erratum [GL17a].

Gora:2012:HAN

[GLBP12] Pawel Góra, Zhenyang Li, Abraham Boyarsky, and Harald Proppe. Harmonic averages and new explicit constants for invariant densities of piecewise expanding maps of the interval. Journal of Statistical Physics, 146(4):850–863, February 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0425-x.

Gora:2014:TMH

[GLBP14] Pawel Góra, Zhenyang Li, Abraham Boyarsky, and Harald Proppe. Toward a mathematical holographic principle. Journal of Statistical Physics, 156(4):775–799, August 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1029-4.

Garrabos:2015:CES

[GLM^{+}15] Y. Garrabos, C. Lecoutre, S. Marre, R. Guillaument, D. Beysens, and I. Hahn. Crossover equation of state models applied to the critical behavior of xenon. Journal of Statistical Physics, 158(6):1379–1412, March 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1157-x.

Garrabos:2016:CCFb

[GLM^{+}16] Yves Garrabos, Carole Lecoutre, Samuel Marre, Bernard LeNeindre, and Inseob Hahn. Critical crossover functions for simple fluids: Towards the crossover modelling uniqueness. Journal of Statistical Physics, 165(3):471–506, November 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1627-4.

Gawedzki:2018:FTU

[GLM18] Krzysztof Gawedzki, Edwin Langmann, and Per Moosavi. Finite-time universality in nonequilibrium CFT. Journal of Statistical Physics, 172(2):353–378, July 2018. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-018-2025-x.pdf.

Garrabos:2016:CCFa

[GLML16] Yves Garrabos, Carole Lecoutre, Samuel Marre, and Bernard LeNeindre. Critical crossover functions for simple fluids: Nonanalytical scaling determination of the Ising-like crossover parameter. Journal of Statistical Physics, 164(3):575–615, August 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1554-4.

Galves:2010:PSI

[GLO10] A. Galves, E. Löcherbach, and E. Orlandi. Perfect simulation of infinite range Gibbs measures and coupling with their finite range approximations. Journal of Statistical Physics, 138(1–3):476–495, February 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9881-3.

Grahovac:2016:ISO

[GLST16] Danijel Grahovac, Nikolai N. Leonenko, Alla Sikorskii, and Irena Tesnjak. Intermittency of superpositions of Ornstein– Uhlenbeck type processes. Journal of Statistical Physics, 165 (2):390–408, October 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1616-7.

Grahovac:2015:SPE

[GLT15] Danijel Grahovac, Nikolai N. Leonenko, and Murad S. Taqqu. Scaling properties of the empirical structure function of linear fractional stable motion and estimation of its parameters. Journal of Statistical Physics, 158(1):105–119, January 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1126-4.

Grosskinsky:2012:LPP

[GLU12] Stefan Großkinsky, Alexander A. Lovisolo, and Daniel Ueltschi. Lattice permutations and Poisson–Dirichlet distribution of cycle lengths. Journal of Statistical Physics, 146 (6):1105–1121, March 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0450-9.

Gholizade:2010:FEG

[GM10] H. Gholizade and D. Momeni. Ferromagnetism of electron gas. Journal of Statistical Physics, 141(6):957–969, December 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-0091-9.

Gustavsson:2013:LEP

[GM13] K. Gustavsson and B. Mehlig. Lyapunov exponents for particles advected in compressible random velocity fields at small and large Kubo numbers. Journal of Statistical Physics, 153 (5):813–827, December 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-013-0848-z.
pdf.

Garcia:2015:SPC

[GM15] Nancy L. Garcia and Lucas Moreira. Stochastically perturbed chains of variable memory. Journal of Statistical Physics, 159 (5):1107–1126, June 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1227-8.

Gantert:2018:RSB

[GMM18] Nina Gantert, Matthias Meiners, and Sebastian Müller. Regularity of the speed of biased random walk in a one-dimensional percolation model. Journal of Statistical Physics, 170(6):1123– 1160, March 2018. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-018-1982-4.pdf.

Gantert:2019:ERR

[GMM19] Nina Gantert, Matthias Meiners, and Sebastian Müller. Einstein relation for random walk in a one-dimensional percolation model. Journal of Statistical Physics, 176(4):737–772, August 2019. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-019-02319-y.pdf.

Grabsch:2017:TLSb

[GMT17a] Aurélien Grabsch, Satya N. Majumdar, and Christophe Texier. Truncated linear statistics associated with the eigenvalues of random matrices II. Partial sums over proper time delays for chaotic quantum dots. Journal of Statistical Physics, 167 (6):1452–1488, June 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Grabsch:2017:TLSa

[GMT17b] Aurélien Grabsch, Satya N. Majumdar, and Christophe Texier. Truncated linear statistics associated with the top eigenvalues of random matrices. Journal of Statistical Physics, 167(2):234–259, April 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-017-1755-5.

Gherardi:2013:DLE

[GN13] Marco Gherardi and Alessandro Nigro. q-deformed Loewner evolution. Journal of Statistical Physics, 152(3):452–472, August 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0771-3.

Giorno:2019:CSC

[GN19] Virginia Giorno and Amelia G. Nobile. On the construction of a special class of time-inhomogeneous diffusion processes. Journal of Statistical Physics, 177(2):299–323, October 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Gonzalez-Navarrete:2016:PTF

[GNP16] Manuel González-Navarrete and Eugene Pechersky. Phase transition in ferromagnetic Ising model with a cell-board external field. Journal of Statistical Physics, 162(1):139–161, January 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1392-9.

Galatolo:2018:DCQ

[GNP18] Stefano Galatolo, Isaia Nisoli, and Maria Jose Pacifico. Decay of correlations, quantitative recurrence and logarithm law for contracting Lorenz attractors. Journal of Statistical Physics, 170(5):862–882, March 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Gordon:2013:EBD

[GNPS13] Mirta B. Gordon, Jean-Pierre Nadal, Denis Phan, and Viktoriya Semeshenko. Entanglement between demand and supply in markets with bandwagon goods. Journal of Statistical Physics, 151(3–4):494–522, May 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0660-1.

Gheissari:2018:ZTD

[GNS18] Reza Gheissari, Charles M. Newman, and Daniel L. Stein. Zero-temperature dynamics in the dilute Curie–Weiss model. Journal of Statistical Physics, 172(4):1009–1028, August 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Gutkin:2011:SPB

[GO11] B. Gutkin and V. A. Osipov. Spectral problem of blockrectangular hierarchical matrices. Journal of Statistical Physics, 143(1):72–87, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0162-6.

Gutkin:2013:CPO

[GO13] Boris Gutkin and Vladimir Osipov. Clustering of periodic orbits and ensembles of truncated unitary matrices. Journal of Statistical Physics, 153(6):1049–1064, December 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0859-9.

Grothaus:2011:SAF

[GOdSS11] Martin Grothaus, Maria João Oliveira, José Luís da Silva, and Ludwig Streit. Self-avoiding fractional Brownian motion — the Edwards model. Journal of Statistical Physics, 145(6):1513–1523, December 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0344-2.

Goldman:2010:HMC

[Gol10] Carla Goldman. A hopping mechanism for cargo transport by molecular motors on crowded microtubules. Journal of Statistical Physics, 140(6):1167–1181, September 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0037-2.

Goncalves:2014:AZR

[Gon14] Patrícia Gonçalves. On the asymmetric zero-range in the rarefaction fan. Journal of Statistical Physics, 154(4):1074–1095, February 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0892-8.

Goodman:2012:EGP

[Goo12] Jesse Goodman. Exponential growth of ponds in invasion percolation on regular trees. Journal of Statistical Physics, 147(5):919–941, June 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-012-0509-7.
pdf.

Garra:2014:FKG

[GOP14] Roberto Garra, Enzo Orsingher, and Federico Polito. Fractional Klein–Gordon equations and related stochastic processes. Journal of Statistical Physics, 155(4):777–809, May 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-0976-0.

Garoni:2011:DCB

[GOPS11] Timothy M. Garoni, Giovanni Ossola, Marco Polin, and Alan D. Sokal. Dynamic critical behavior of the Chayes– Machta algorithm for the random-cluster model, I. Two dimensions. Journal of Statistical Physics, 144(3):459–518, August 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0267-y.

Gorbachev:2018:NEQ

[Gor18] Yuriy E. Gorbachev. Non-equilibrium quasi-chemical nucleation model. Journal of Statistical Physics, 171(2):288–344, April 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Gould:2014:LMS

[Gou14] Harvey Gould. Leonard M. Sander: Equilibrium Statistical Physics. Journal of Statistical Physics, 156(1):201–202, July 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-0994-y.

Gough:2015:SNC

[Gou15] John E. Gough. Symplectic noise and the classical analog of the Lindblad generator. Journal of Statistical Physics, 160 (6):1709–1720, September 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1303-0.

Gallavotti:2010:FTI

[GP10] G. Gallavotti and E. Presutti. Frictionless thermostats and intensive constants of motion. Journal of Statistical Physics, 139(4):618–629, May 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9949-0.

Granath:2016:DLR

[GPD16] Mats Granath and Alvaro Perez-Diaz. Diffusion and localization of relative strategy scores in the minority game. Journal of Statistical Physics, 165(1):94–114, October 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-016-1607-8.pdf.

Gautam:2017:DIC

[GPGA17] Arvind Kumar Gautam, Nandlal Pingua, Aashish Goyal, and Pankaj A. Apte. Dynamical instability causes the demise of a supercooled tetrahedral liquid. Journal of Statistical Physics, 168(6):1302–1318, September 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Garcia-Perciante:2017:HFR

[GPMEA17] A. L. García-Perciante, A. R. Méndez, and E. Escobar-Aguilar. Heat flux for a relativistic dilute bidimensional gas. Journal of Statistical Physics, 167(1):123–134, April 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.

1007/s10955-017-1742-x. See correction [GPMEA18].

Garcia-Perciante:2018:CHF

[GPMEA18] A. L. García-Perciante, A. R. Méndez, and E. Escobar-Aguilar. Correction to: Heat Flux for a Relativistic Dilute Bidimensional Gas. Journal of Statistical Physics, 172 (6):1682, September 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-018-2120-z.

pdf. See [GPMEA17].

Garcia-Perciante:2015:SPS

[GPMSBBSV15] A. L. Garcia-Perciante, H. Mondragon-Suarez, D. Brun-Battistini, and A. Sandoval-Villalbazo. On the stability problem in special relativistic thermodynamics: Implications of the Chapman–Enskog formalism. Journal of Statistical Physics, 160(3):760–769, August 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1272-3.

Giorgilli:2012:EAI

[GPP12] Antonio Giorgilli, Simone Paleari, and Tiziano Penati. Extensive adiabatic invariants for nonlinear chains. Journal of Statistical Physics, 148(6):1106–1134, September 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0568-9.

Gallesco:2013:LRW

[GPS13] Christophe Gallesco, Serguei Popov, and Gunter M. Schütz. Localization for a random walk in slowly decreasing random potential. Journal of Statistical Physics, 150(2):285–298, January 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0671-y.

Garcia-Perez:2018:SCS

[GPSB18] Guillermo García-Pérez, M. Ángeles Serrano, and Marián Boguñá. Soft communities in similarity space. Journal of Statistical Physics, 173(3–4):775–782, November 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ge:2017:MFN

[GQ17] Hao Ge and Hong Qian. Mathematical formalism of nonequilibrium thermodynamics for nonlinear chemical reaction systems with general rate law. Journal of Statistical Physics, 166(1):190–209, January 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1678-6.

Gabetta:2012:CCC

[GR12] Ester Gabetta and Eugenio Regazzini. Complete characterization of convergence to equilibrium for an inelastic Kac model. Journal of Statistical Physics, 147(5):1007–1019, June 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0505-y.

Grassberger:2013:TDS

[Gra13] Peter Grassberger. Two-dimensional SIR epidemics with long range infection. Journal of Statistical Physics, 153(2):289–311, October 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0824-7.

Gregorio:2012:RFN

[Gre12] Alessandro De Gregorio. On random flights with nonuniformly distributed directions. Journal of Statistical Physics, 147(2):382–411, April 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0471-4.

Grebenkov:2016:AGA

[Gre16] Denis S. Grebenkov. Anomalous growth of aging populations. Journal of Statistical Physics, 163(2):440–455, April 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1488-x.

Grigo:2019:RDH

[Gri19] Alexander Grigo. A rigorous derivation of Haff’s Law for a periodic two-disk fluid. Journal of Statistical Physics, 176(4): 806–835, August 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Grmela:2010:HTM

[Grm10] Miroslav Grmela. Hamiltonian and thermodynamic modeling of quantum turbulence. Journal of Statistical Physics, 141(2):318–341, October 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0048-z.

Grmela:2017:EDM

[Grm17] Miroslav Grmela. Externally driven macroscopic systems: Dynamics versus thermodynamics. Journal of Statistical Physics, 166(2):282–316, January 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1694-6.

Groenevelt:2019:OSD

[Gro19] Wolter Groenevelt. Orthogonal stochastic duality functions from Lie algebra representations. Journal of Statistical Physics, 174(1):97–119, January 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s10955-018-
2178-7.pdf.

Gandolfo:2017:BCT

[GRR17] D. Gandolfo, M. M. Rahmatullaev, and U. A. Rozikov. Boundary conditions for translation-invariant Gibbs measures of the Potts model on Cayley trees. Journal of Statistical Physics, 167(5):1164–1179, June 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Gandolfo:2013:FEI

[GRRR13] D. Gandolfo, M. M. Rakhmatullaev, U. A. Rozikov, and J. Ruiz. On free energies of the Ising model on the Cayley tree. Journal of Statistical Physics, 150(6):1201–1217, March 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0713-0.

Gandolfo:2012:MPG

[GRS12] Daniel Gandolfo, Jean Ruiz, and Senya Shlosman. A manifold of pure Gibbs states of the Ising model on a Cayley tree. Journal of Statistical Physics, 148(6):999–1005, September 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0574-y.

Gandolfo:2017:MFA

[GRT17] Daniel Gandolfo, Roger Rodriguez, and Henry C. Tuckwell. Mean field analysis of large-scale interacting populations of stochastic conductance-based spiking neurons using the Klimontovich method. Journal of Statistical Physics, 166 (5):1310–1333, March 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1702-x.

Giardina:2010:CII

[GRV10] C. Giardinà, F. Redig, and K. Vafayi. Correlation inequalities for interacting particle systems with duality. Journal of Statistical Physics, 141(2):242–263, October 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0055-0.

Grosskinsky:2011:CIP

[GRV11] Stefan Großkinsky, Frank Redig, and Kiamars Vafayi. Condensation in the inclusion process and related models. Journal of Statistical Physics, 142(5):952–974, March 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-011-0151-9.pdf.

Gaveau:2011:VZL

[GS11a] B. Gaveau and L. S. Schulman. Violation of the zeroth law of thermodynamics for a non-ergodic interaction. Journal of Statistical Physics, 145(6):1458–1471, December 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0362-0.

Gitterman:2011:SRH

[GS11b] M. Gitterman and I. Shapiro. Stochastic resonance in a harmonic oscillator with random mass subject to asymmetric dichotomous noise. Journal of Statistical Physics, 144(1):139– 149, July 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0240-9.

Grisi:2011:CSP

[GS11c] Rafael M. Grisi and Gunter M. Schütz. Current symmetries for particle systems with several conservation laws. Journal of Statistical Physics, 145(6):1499–1512, December 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0341-5.

Genovese:2012:SBH

[GS12] Giuseppe Genovese and Sergio Simonella. On the stationary BBGKY hierarchy for equilibrium states. Journal of Statistical Physics, 148(1):89–112, July 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0525-7.

Golosovsky:2013:TTI

[GS13] Michael Golosovsky and Sorin Solomon. The transition towards immortality: Non-linear autocatalytic growth of citations to scientific papers. Journal of Statistical Physics, 151 (1–2):340–354, April 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0714-z.

Guggiola:2015:MCS

[GS15] Alberto Guggiola and Guilhem Semerjian. Minimal contagious sets in random regular graphs. Journal of Statistical Physics, 158(2):300–358, January 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1136-2.

Gaveau:2017:DDM

[GS17a] Bernard Gaveau and Lawrence S. Schulman. Decoherence, the density matrix, the thermal state and the classical world. Journal of Statistical Physics, 169(5):889–901, December 2017. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-017-1901-0.pdf.

Goldenfeld:2017:ETP

[GS17b] Nigel Goldenfeld and Hong-Yan Shih. Erratum to: Turbulence as a Problem in Non-equilibrium Statistical Mechanics. Journal of Statistical Physics, 167(3–4):595, May 2017.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-016-1713-7.pdf. See [GS17c].

Goldenfeld:2017:TPN

[GS17c] Nigel Goldenfeld and Hong-Yan Shih. Turbulence as a problem in non-equilibrium statistical mechanics. Journal of Statistical Physics, 167(3–4):575–594, May 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). See erratum [GS17b].

Gaveau:2019:DPT

[GS19] B. Gaveau and L. S. Schulman. Decoherence and phase transitions in quantum dynamics. Journal of Statistical Physics, 174(4):800–807, February 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-018-2214-7.
pdf.

Gaudilliere:2011:PTC

[GSSV11] Alexandre Gaudillière, Benedetto Scoppola, Elisabetta Scoppola, and Massimiliano Viale. Phase transitions for the cavity approach to the clique problem on random graphs. Journal of Statistical Physics, 145(5):1127–1155, December 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0336-2.

Gould:2010:STP

[GT10] Harvey Gould and Jan Tobochnik. Statistical and Thermal Physics: with Computer Applications. Princeton University Press, Princeton, NJ, USA, 2010. ISBN 0-691-13744-7. xvii + 511 pp. LCCN QC174.8 .G68 2010.

Garibaldi:2012:DSG

[GT12] Eduardo Garibaldi and Philippe Thieullen. Description of some ground states by Puiseux techniques. Journal of Statistical Physics, 146(1):125–180, January 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0357-x.

Garibaldi:2015:EDG

[GT15a] Eduardo Garibaldi and Philippe Thieullen. An ergodic description of ground states. Journal of Statistical Physics, 158 (2):359–371, January 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1139-z.

Grimaldo:2015:RAT

[GT15b] Johnny Alejandro Mora Grimaldo and Gabriel Téllez. Relations among two methods for computing the partition function of the two-dimensional one-component plasma. Journal of Statistical Physics, 160(1):4–28, July 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1249-2.

Genovese:2016:NCM

[GT16] Giuseppe Genovese and Daniele Tantari. Non-convex multipartite ferromagnets. Journal of Statistical Physics, 163 (3):492–513, May 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1482-3.

Genovese:2017:OSM

[GT17] Giuseppe Genovese and Daniele Tantari. Overlap synchronisation in multipartite random energy models. Journal of Statistical Physics, 169(6):1162–1170, December 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Grabsch:2014:ODD

[GTT14] Aurélien Grabsch, Christophe Texier, and Yves Tourigny. One-dimensional disordered quantum mechanics and Sinai diffusion with random absorbers. Journal of Statistical Physics, 155(2):237–276, April 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-0957-3.

Goel:2019:SLL

[GTT19] Akshay Goel, Khanh Duy Trinh, and Kenkichi Tsunoda. Strong law of large numbers for Betti numbers in the thermodynamic regime. Journal of Statistical Physics, 174(4): 865–892, February 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Goldstein:2012:QFG

[GTZ12] Sheldon Goldstein, Roderich Tumulka, and Nino Zanghì. The quantum formalism and the GRW formalism. Journal of Statistical Physics, 149(1):142–201, October 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0587-6.

Guerra:2013:PCC

[Gue13] Francesco Guerra. Pierluigi Contucci and Cristian Giardinà: Perspectives on Spin Glasses. Journal of Statistical Physics, 151(5):985–986, June 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0736-6.

Gupta:2016:CDT

[Gup16] Arvind Kumar Gupta. Collective dynamics on a two–lane asymmetrically coupled TASEP with mutually interactive Langmuir kinetics. Journal of Statistical Physics, 162(6): 1571–1586, March 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1463-6.

Guttmann:2011:I

[Gut11] Tony Guttmann. Introduction. Journal of Statistical Physics, 145(3):508–509, November 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0379-4. See

erratum [Gut12].

Guttmann:2012:EI

[Gut12] Tony Guttmann. Erratum to: Introduction. Journal of Statistical Physics, 146(5):1104, March 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s10955-012-

0432-y.pdf. See [Gut11].

Gorissen:2012:RDT

[GV12a] Mieke Gorissen and Carlo Vanderzande. Ribosome dwell times and the protein copy number distribution. Journal of Statistical Physics, 148(4):628–636, September 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0452-7.

Gulpinar:2012:EQR

[GV12b] Gul Gulpinar and Erol Vatansever. Effects of the quenched random crystal field on the dynamic spin-1 Blume–Capel model. Journal of Statistical Physics, 146(4):787–799, February 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0423-z.

Gorny:2015:FSN

[GV15] Matthias Gorny and S. R. S. Varadhan. Fluctuations of the self-normalized sum in the Curie–Weiss model of SOC. Journal of Statistical Physics, 160(3):513–518, August 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1274-1.

Garlaschelli:2018:SIJ

[GvdHdHM18] Diego Garlaschelli, Remco van der Hofstad, Frank den Hollander, and Michel Mandjes. Special issue of Journal of Statistical Physics devoted to complex networks. Journal of Statistical Physics, 173(3–4):439–447, November 2018. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-018-2166-y.pdf.

Garavaglia:2017:DPL

[GvdHW17] Alessandro Garavaglia, Remco van der Hofstad, and Gerhard Woeginger. The dynamics of power laws: Fitness and aging in preferential attachment trees. Journal of Statistical Physics, 168(6):1137–1179, September 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-017-1841-8.pdf.

Goto:2018:SEN

[GVJ^{+}18] Hayato Goto, Eduardo Viegas, Henrik Jeldtoft Jensen, Hideki Takayasu, and Misako Takayasu. Smoluchowski equation for networks: Merger induced intermittent giant node formation and degree gap. Journal of Statistical Physics, 172(4):

1086–1100, August 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-018-2073-2.
pdf.

Gajda:2012:GBM

[GW12] Janusz Gajda and Agnieszka Wylomańska. Geometric Brownian motion with tempered stable waiting times. Journal of Statistical Physics, 148(2):296–305, August 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0537-3.

Grant:2015:ADE

[GW15a] John Grant and Michael Wilkinson. Advection-diffusion equation with absorbing boundary. Journal of Statistical Physics, 160(3):622–635, August 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1257-2.

Gunaratnam:2015:MCL

[GW15b] B. Gunaratnam and W. A. Woyczyński. Multiscale conservation laws driven by Lévy stable and Linnik diffusions: Asymptotics, shock creation, preservation and dissolution. Journal of Statistical Physics, 160(1):29–59, July 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1240-y.

Gao:2012:ACE

[GXL12] Ang Gao, Jianbo Xie, and Yueheng Lan. Accelerating cycle expansions by dynamical conjugacy. Journal of Statistical Physics, 146(1):56–66, January 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0369-6.

Halder:2013:SPA

[HA13] Avik Halder and Ansuman Adhikary. Statistical physics analysis of maximum a posteriori estimation for multi-channel hidden Markov models. Journal of Statistical Physics, 150 (4):744–775, February 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0711-2.

Homayoun:2019:EEW

[HA19] T. Homayoun and K. Aghayar. Energy as an entanglement witnesses for one dimensional XYZ Heisenberg lattice: Optimization approach. Journal of Statistical Physics, 176(1): 85–93, July 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Hagendorf:2013:SCD

[Hag13] Christian Hagendorf. Spin chains with dynamical lattice supersymmetry. Journal of Statistical Physics, 150(4):609–657, February 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0709-9.

Haga:2015:NLE

[Hag15] Taiki Haga. Nonequilibrium Langevin equation and effective temperature for particle interacting with spatially extended environment. Journal of Statistical Physics, 159(3):713–729, May 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1195-z.

Halsey:2017:MBC

[Hal17] Thomas C. Halsey. A model for branch competition. Journal of Statistical Physics, 167(3–4):713–725, May 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Halperin:2019:CHM

[Hal19a] Bertrand I. Halperin. Correction to: On the Hohenberg– Mermin–Wagner Theorem and Its Limitations. Journal of Statistical Physics, 175(3–4):530, May 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s10955-019-

02239-x.pdf. See [Hal19b].

Halperin:2019:HMW

[Hal19b] Bertrand I. Halperin. On the Hohenberg–Mermin–Wagner theorem and its limitations. Journal of Statistical Physics, 175(3–4):521–529, May 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). See correction [Hal19a].

Hammond:2011:PSR

[Ham11] Alan Hammond. Phase separation in random cluster models III: Circuit regularity. Journal of Statistical Physics, 142 (2):229–276, January 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0081-y.

Hansen:2013:WLS

[Han13] Frank Hansen. WYD-like skew information measures. Journal of Statistical Physics, 151(5):974–979, June 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0737-5.

Hansen:2014:TFA

[Han14] Frank Hansen. Trace functions with applications in quantum physics. Journal of Statistical Physics, 154(3):807–818, February 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0890-x.

Hansen:2015:GTI

[Han15] Frank Hansen. Golden–Thompson’s inequality for deformed exponentials. Journal of Statistical Physics, 159(6):1300– 1305, June 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1237-6.

Hansen:2016:QED

[Han16] Frank Hansen. Quantum entropy derived from first principles. Journal of Statistical Physics, 165(5):799–808, December 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1651-4.

Hanke:2018:WPI

[Han18] Martin Hanke. Well-posedness of the iterative Boltzmann inversion. Journal of Statistical Physics, 170(3):536–553, February 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Hartmann:2011:GST

[Har11] Alexander K. Hartmann. Ground states of two-dimensional Ising spin glasses: Fast algorithms, recent developments and a ferromagnet-spin glass mixture. Journal of Statistical Physics, 144(3):519–540, August 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0272-1.

Hauray:2016:UCW

[Hau16] Maxime Hauray. Uniform contractivity in Wasserstein metric for the original 1D Kac’s model. Journal of Statistical Physics, 162(6):1566–1570, March 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1476-1.

Haydn:2015:ASI

[Hay15] Nicolai Haydn. The almost sure invariance principle for betamixing measures. Journal of Statistical Physics, 159(2):231– 254, April 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1185-6.

Hagendorf:2010:GFF

[HBB10] Christian Hagendorf, Denis Bernard, and Michel Bauer. The Gaussian free field and SLE_{4} on doubly connected domains. Journal of Statistical Physics, 140(1):1–26, July 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-9980-1.

Helbing:2015:SHL

[HBC^{+}15] Dirk Helbing, Dirk Brockmann, Thomas Chadefaux, Karsten Donnay, Ulf Blanke, Olivia Woolley-Meza, Mehdi Moussaid, Anders Johansson, Jens Krause, Sebastian Schutte, and Matjaz Perc. Saving human lives: What complexity science and information systems can contribute. Journal of Statistical Physics, 158(3):735–781, February 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-014-1024-9.pdf.

Hauert:2014:FTD

[HCI14] Christoph Hauert, Yu-Ting Chen, and Lorens A. Imhof. Fixation times in deme structured, finite populations with rare migration. Journal of Statistical Physics, 156(4):739–759, August 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1022-y.

Hu:2011:QIT

[HCLR11] Bo Hu, Wen Chen, Herbert Levine, and Wouter-Jan Rappel. Quantifying information transmission in eukaryotic gradient sensing and chemotactic response. Journal of Statistical Physics, 142(6):1167–1186, April 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-011-0156-4.pdf.

Hwang:2015:TIR

[HCO15] Zaijong Hwang, Frank Cao, and Maxim Olshanii. Traces of integrability in relaxation of one-dimensional two-mass mixtures. Journal of Statistical Physics, 161(2):467–474, October 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1331-9.

Haug:2017:HOA

[HDP17] Nils Haug, Adri Olde Daalhuis, and Thomas Prellberg. Higher-order Airy scaling in deformed Dyck paths. Journal of Statistical Physics, 166(5):1193–1208, March 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-016-1708-4.pdf.

Hirsch:2015:SAP

[HDS15] Christian Hirsch, Gary Delaney, and Volker Schmidt. Stationary Apollonian packings. Journal of Statistical Physics, 161(1):35–72, October 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1326-6.

He:2014:AAS

[He14] Lingbing He. Asymptotic analysis of the spatially homogeneous Boltzmann equation: Grazing collisions limit. Journal of Statistical Physics, 155(1):151–210, April 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0932-z.

Hryniv:2017:SMM

[HE17] Ostap Hryniv and Antonio Martínez Esteban. Stochastic model of microtubule dynamics. Journal of Statistical Physics, 169(1):203–222, October 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

He:2019:MLS

[He19] Yukun He. Mesoscopic linear statistics of Wigner matrices of mixed symmetry class. Journal of Statistical Physics, 175 (5):932–959, June 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Hurtado:2014:TCN

[HEdPG14] Pablo I. Hurtado, Carlos P. Espigares, Jesús J. del Pozo, and Pedro L. Garrido. Thermodynamics of currents in nonequilibrium diffusive systems: Theory and simulation. Journal of Statistical Physics, 154(1–2):214–264, January 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0894-6.

Helmuth:2016:DRG

[Hel16] Tyler Helmuth. Dimensional reduction for generalized continuum polymers. Journal of Statistical Physics, 165(1):24–43, October 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1597-6.

Henley:2012:POM

[Hen12] Christopher L. Henley. Possible origins of macroscopic leftright asymmetry in organisms. Journal of Statistical Physics, 148(4):741–775, September 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0520-z.

Hepp:2018:WSP

[Hep18] K. Hepp. The wake-sleep ‘phase transition’ at the gate to consciousness. Journal of Statistical Physics, 172(2):562–568, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Herbert:2013:AIS

[Her13] Corentin Herbert. Additional invariants and statistical equilibria for the 2D Euler equations on a spherical domain. Journal of Statistical Physics, 152(6):1084–1114, September 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0809-6.

Hagendorf:2012:EVM

[HF12a] Christian Hagendorf and Paul Fendley. The eight-vertex model and lattice supersymmetry. Journal of Statistical Physics, 146(6):1122–1155, March 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-012-0430-
0.

Holcomb:2012:ESJ

[HF12b] Diane Holcomb and Gregorio R. Moreno Flores. Edge scaling of the β-Jacobi ensemble. Journal of Statistical Physics, 149(6):1136–1160, December 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0634-3.

Huepe:2015:SFC

[HFWT15] Cristián Huepe, Eliseo Ferrante, Tom Wenseleers, and Ali Emre Turgut. Scale-free correlations in flocking systems with position-based interactions. Journal of Statistical Physics, 158(3):549–562, February 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-014-1114-
8.

Hsu:2011:RMC

[HG11] Hsiao-Ping Hsu and Peter Grassberger. A review of Monte Carlo simulations of polymers with PERM. Journal of Statistical Physics, 144(3):597–637, August 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0268-x.

Haimi:2013:PGE

[HH13] Antti Haimi and Haakan Hedenmalm. The polyanalytic Ginibre ensembles. Journal of Statistical Physics, 153(1):10–47, October 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0813-x.

Hemelrijk:2015:SFC

[HH15a] Charlotte K. Hemelrijk and Hanno Hildenbrandt. Scalefree correlations, influential neighbours and speed control in flocks of birds. Journal of Statistical Physics, 158(3):563–578, February 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1154-0.

Hirono:2015:JTE

[HH15b] Yuji Hirono and Yoshimasa Hidaka. Jarzynski-type equalities in gambling: Role of information in capital growth. Journal of Statistical Physics, 161(3):721–742, November 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1348-0.

Haeseler:2017:DDS

[HHLP17] Sebastian Haeseler, Xueping Huang, Daniel Lenz, and Felix Pogorzelski. Diffusion on Delone sets. Journal of Statistical Physics, 167(6):1496–1510, June 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Haack:2017:CEM

[HHM17] Jeffrey R. Haack, Cory D. Hauck, and Michael S. Murillo. A conservative, entropic multispecies BGK model. Journal of Statistical Physics, 168(4):826–856, August 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Hu:2010:CDC

[HHT10] Ke Hu, Tao Hu, and Yi Tang. Cascade defense via control of the fluxes in complex networks. Journal of Statistical Physics, 141(3):555–565, November 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0058-x.

Halpin-Healy:2015:KCS

[HHT15] Timothy Halpin-Healy and Kazumasa A. Takeuchi. A KPZ cocktail — shaken, not stirred . . .. Journal of Statistical Physics, 160(4):794–814, August 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1282-1.

Herzog:2016:SML

[HHV16] David P. Herzog, Scott Hottovy, and Giovanni Volpe. The small-mass limit for Langevin dynamics with unbounded coefficients and positive friction. Journal of Statistical Physics, 163(3):659–673, May 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1498-8.

Hairer:2018:TIK

[HI18] Martin Hairer and Massimo Iberti. Tightness of the Ising–Kac model on the two-dimensional torus. Journal of Statistical Physics, 171(4):632–655, May 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ho:2018:SAD

[HIK^{+}18] Choon-Lin Ho, Yusuke Ide, Norio Konno, Etsuo Segawa, and Kentaro Takumi. A spectral analysis of discrete-time quantum walks related to the birth and death chains. Journal of Statistical Physics, 171(2):207–219, April 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Hiura:2018:HDP

[HiS18] Ken Hiura and Shin ichi Sasa. How does pressure fluctuate in equilibrium? Journal of Statistical Physics, 173(2):285–294, October 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Hiura:2019:MRM

[HiS19] Ken Hiura and Shin ichi Sasa. Microscopic reversibility and macroscopic irreversibility: From the viewpoint of algorithmic randomness. Journal of Statistical Physics, 177(5):727–751, December 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Hamza:2017:TFQ

[HJ17a] Eman Hamza and Alain Joye. Thermalization of fermionic quantum walkers. Journal of Statistical Physics, 166(6):1365– 1392, March 2017. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-017-1727-9.

He:2017:WPS

[HJ17b] Lingbing He and Jin-Cheng Jiang. Well-posedness and scattering for the Boltzmann equations: Soft potential with cutoff. Journal of Statistical Physics, 168(2):470–481, July 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Hujo:2011:RFA

[HJR^{+}11] Waldemar Hujo, B. Shadrack Jabes, Varun K. Rana, Charusita Chakravarty, and Valeria Molinero. The rise and fall of anomalies in tetrahedral liquids. Journal of Statistical Physics, 145(2):293–312, October 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-011-0293-
9.

Hotta:2018:HLM

[HK18] Ikkei Hotta and Makoto Katori. Hydrodynamic limit of multiple SLE. Journal of Statistical Physics, 171(1):166–188, April 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ha:2019:EBS

[HK19] Seung-Yeal Ha and Dohyun Kim. Emergent behavior of a second-order lohe matrix model on the unitary group. Journal of Statistical Physics, 175(5):904–931, June 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ha:2019:PKM

[HKLN19] Seung-Yeal Ha, Dohyun Kim, Jaeseung Lee, and Se Eun Noh. Particle and kinetic models for swarming particles on a sphere and stability properties. Journal of Statistical Physics, 174(3): 622–655, February 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Han-Kwan:2016:IMF

[HKN16] Daniel Han-Kwan and Toan T. Nguyen. Instabilities in the mean field limit. Journal of Statistical Physics, 162(6):1639– 1653, March 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1455-6.

Ha:2016:FCP

[HKR16] Seung-Yeal Ha, Hwa Kil Kim, and Sang Woo Ryoo. On the finiteness of collisions and phase-locked states for the Kuramoto model. Journal of Statistical Physics, 163(6):1394– 1424, June 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1528-6.

Ha:2017:EDG

[HKR17] Seung-Yeal Ha, Dongnam Ko, and Sang Woo Ryoo. Emergent dynamics of a generalized Lohe model on some class of Lie groups. Journal of Statistical Physics, 168(1):171–207, July 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). See correction [HKR23a].

Ha:2018:RDL

[HKR18] Seung-Yeal Ha, Dongnam Ko, and Sang Woo Ryoo. On the relaxation dynamics of Lohe oscillators on some Riemannian manifolds. Journal of Statistical Physics, 172(5):1427–1478, September 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). See correction [HKR23b].

Ha:2023:CED

[HKR23a] Seung-Yeal Ha, Dongnam Ko, and Seung-Yeon Ryoo. Correction to: Emergent dynamics of a generalized Lohe model on some class of Lie groups. Journal of Statistical Physics, 190(8): ??, August 2023. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-023-03150-2. See [HKR17].

Ha:2023:CRD

[HKR23b] Seung-Yeal Ha, Dongnam Ko, and Seung-Yeon Ryoo. Correction to: On the relaxation dynamics of Lohe oscillators on some Riemannian manifolds. Journal of Statistical Physics, 190(8):??, August 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-023-03151-1. See

[HKR18].

Hu:2011:CPA

[HKW11] Yueyun Hu, Davar Khoshnevisan, and Marc Wouts. Charged polymers in the attractive regime: A first-order transition from Brownian scaling to four-point localization. Journal of Statistical Physics, 144(5):948–977, September 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0280-1.

Helvik:2014:EEL

[HL14] Torbjørn Helvik and Kristian Lindgren. Expressing the entropy of lattice systems as sums of conditional entropies. Journal of Statistical Physics, 155(4):687–702, May 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0972-4.

Hautphenne:2016:LEB

[HL16] Sophie Hautphenne and Guy Latouche. Lyapunov exponents for branching processes in a random environment: The effect of information. Journal of Statistical Physics, 163(2):393– 410, April 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1474-3.

Hirsch:2018:MHC

[HL18] Christian Hirsch and Günter Last. On maximal hard-core thinnings of stationary particle processes. Journal of Statistical Physics, 170(3):554–583, February 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Harlim:2017:PEM

[HLZ17] John Harlim, Xiantao Li, and He Zhang. A parameter estimation method using linear response statistics. Journal of Statistical Physics, 168(1):146–170, July 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Hauser:2013:LCB

[HM13a] Raphael Hauser and Heinrich Matzinger. Letter change bias and local uniqueness in optimal sequence alignments. Journal of Statistical Physics, 153(3):512–529, November 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0819-4.

Huillet:2013:ODA

[HM13b] Thierry Huillet and Servet Martínez. Occupancy distributions arising in sampling from Gibbs–Poisson abundance models. Journal of Statistical Physics, 153(5):763–800, December 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0865-y.

Houdre:2016:VOA

[HM16] Christian Houdré and Heinrich Matzinger. On the variance of the optimal alignments score for binary random words and an asymmetric scoring function. Journal of Statistical Physics, 164(3):693–734, August 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1549-1.

He:2019:DPR

[HM19] Yukun He and Matteo Marcozzi. Diffusion profile for random band matrices: a short proof. Journal of Statistical Physics, 177(4):666–716, November 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Holmes:2015:VMC

[HMN15] Mark Holmes, Yevhen Mohylevskyy, and Charles M. Newman. The voter model chordal interface in two dimensions. Journal of Statistical Physics, 159(4):937–957, May 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1198-9.

Hart:2012:GAC

[HMO12] Andrew Hart, Servet Martínez, and Felipe Olmos. A Gibbs approach to Chargaff’s second parity rule. Journal of Statistical Physics, 146(2):408–422, January 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0377-6.

Hubsch:2013:RFS

[HMRW13] Florian Hübsch, Nicole Marheineke, Klaus Ritter, and Raimund Wegener. Random field sampling for a simplified model of melt-blowing considering turbulent velocity fluctuations. Journal of Statistical Physics, 150(6):1115–1137, March 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0715-y.

Hassani:2013:TSS

[HMU13] S. Hamed Hassani, Nicolas Macris, and Ruediger Urbanke. Threshold saturation in spatially coupled constraint satisfaction problems. Journal of Statistical Physics, 150(5):807–850, March 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0664-x.

Hottovy:2019:SDC

[HMW19] Scott Hottovy, Austin McDaniel, and Jan Wehr. A small delay and correlation time limit of stochastic differential delay equations with state-dependent colored noise. Journal of Statistical Physics, 175(1):19–46, April 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Hladky:2018:LLU

[HNT18] Jan Hladký, Asaf Nachmias, and Tuan Tran. The local limit of the uniform spanning tree on dense graphs. Journal of Statistical Physics, 173(3–4):502–545, November 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Haydn:2013:CLT

[HNVZ13] Nicolai Haydn, Matthew Nicol, Sandro Vaienti, and Licheng Zhang. Central limit theorems for the shrinking target problem. Journal of Statistical Physics, 153(5):864–887, December 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0860-3.

Hoseinpoor:2016:ALB

[HNZ16] S. Mohammad Hoseinpoor, Narges Nikoofard, and Mostafa Zahedifar. Accuracy limits of the blob model for a flexible polymer confined inside a cylindrical nano-channel. Journal of Statistical Physics, 163(3):593–603, May 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1489-9.

Hogreve:2011:RGS

[Hog11] H. Hogreve. A remark on the ground state energy of bosonic atoms. Journal of Statistical Physics, 144(4):904–908, August 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0265-0.

Holroyd:2011:SCW

[Hol11] Alexander E. Holroyd. Some circumstances where extra updates can delay mixing. Journal of Statistical Physics, 145 (6):1649–1652, December 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0365-x.

Hongo:2019:NHQ

[Hon19] Masaru Hongo. Nonrelativistic hydrodynamics from quantum field theory: (I) normal fluid composed of spinless Schrödinger fields. Journal of Statistical Physics, 174(5):1038–1079, March 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Horbacz:2016:CLT

[Hor16] Katarzyna Horbacz. The central limit theorem for random dynamical systems. Journal of Statistical Physics, 164(6): 1261–1291, September 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-016-1601-1.
pdf.

Howard:2011:IES

[HP11] Jesse J. Howard and B. Montgomery Pettitt. Integral equations in the study of polar and ionic interaction site fluids. Journal of Statistical Physics, 145(2):441–466, October 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0260-5.

Herault:2015:LFF

[HPF15] J. Herault, F. Pétrélis, and S. Fauve. 1/f^{α} low frequency fluctuations in turbulent flows. Journal of Statistical Physics, 161(6):1379–1389, December 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-015-1408-
5;   http://link.springer.com/content/pdf/10.1007/
s10955-015-1408-5.pdf.

Hilhorst:2010:SSF

[HR10] Danielle Hilhorst and Piotr Rybka. Stabilization of solutions to a FitzHugh–Nagumo type system. Journal of Statistical Physics, 138(1–3):291–304, February 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-009-9886-y.

Ha:2016:EOS

[HR16] Seung-Yeal Ha and Sang Woo Ryoo. On the emergence and orbital stability of phase-locked states for the Lohe model. Journal of Statistical Physics, 163(2):411–439, April 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1481-4. See correction [HR23].

Herda:2018:LTB

[HR18] Maxime Herda and L. Miguel Rodrigues. Large-time behavior of solutions to Vlasov–Poisson–Fokker–Planck equations: From evanescent collisions to diffusive limit. Journal of Statistical Physics, 170(5):895–931, March 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ha:2023:CEO

[HR23] Seung-Yeal Ha and Seung-Yeon Ryoo. Correction to: On the emergence and orbital stability of phase-locked states for the Lohe model. Journal of Statistical Physics, 190(8):??, August 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/
10.1007/s10955-023-03148-w. See [HR16].

Huang:2014:NSL

[HRW14] Haiping Huang, Jack Raymond, and K. Y. Michael Wong. The network source location problem: Ground state energy, entropy and effects of freezing. Journal of Statistical Physics, 156(2):301–335, July 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1002-2.

Hunt:2010:IED

[HS10] A. G. Hunt and T. E. Skinner. Incorporation of effects of diffusion into advection-mediated dispersion in porous media. Journal of Statistical Physics, 140(3):544–564, August 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-9992-x.

Heurteaux:2014:MDM

[HS14a] Yanick Heurteaux and Andrzej Stos. On measures driven by Markov chains. Journal of Statistical Physics, 157(6): 1046–1061, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1104-x.

Horowitz:2014:EDQ

[HS14b] Jordan M. Horowitz and Takahiro Sagawa. Equivalent definitions of the quantum nonadiabatic entropy production. Journal of Statistical Physics, 156(1):55–65, July 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-0991-1.

Horbacz:2016:LLN

[HŚ16] Katarzyna Horbacz and Maciej Ślȩczka. Law of large numbers for random dynamical systems. Journal of Statistical Physics, 162(3):671–684, February 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s10955-015-
1423-6.pdf.

Hu:2018:FED

[HS18] Yueyun Hu and Zhan Shi. The free energy in the Derrida– Retaux recursive model. Journal of Statistical Physics, 172 (3):718–741, August 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Hwang:2018:UCI

[HSFK18] Sungmin Hwang, Benjamin Schmiegelt, Luca Ferretti, and Joachim Krug. Universality classes of interaction structures for NK fitness landscapes. Journal of Statistical Physics, 172 (1):226–278, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Hasan:2013:SPU

[HSUG13] Samiul Hasan, Christian M. Schneider, Satish V. Ukkusuri, and Marta C. González. Spatiotemporal patterns of urban human mobility. Journal of Statistical Physics, 151(1–2):304– 318, April 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0645-0.

Hartmann:2019:JEF

[HSZ19] Carsten Hartmann, Christof Schütte, and Wei Zhang. Jarzynski’s equality, fluctuation theorems, and variance reduction: Mathematical analysis and numerical algorithms. Journal of Statistical Physics, 175(6):1214–1261, June 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Humenyuk:2011:EHB

[HT11] Y. A. Humenyuk and M. V. Tokarchuk. Extension of hydrodynamic balance equations for simple fluids. Journal of Statistical Physics, 142(5):1052–1084, March 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0141-y.

Hattori:2015:SOK

[HT15] Masanari Hattori and Shigeru Takata. Second-order Knudsenlayer analysis for the generalized slip-flow theory II: Curvature effects. Journal of Statistical Physics, 161(4):1010–1036, November 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1364-0.

Huveneers:2019:EEC

[HT19] François Huveneers and Elias Theil. Equivalence of ensembles, condensation and glassy dynamics in the Bose–Hubbard Hamiltonian. Journal of Statistical Physics, 177(5):917–935, December 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Hao:2012:SSA

[HTX^{+}12] Da-Peng Hao, Gang Tang, Hui Xia, Kui Han, and Zhi-Peng Xun. Simulation study on the avalanche process of continuous damage fiber bundle model with strong disorder. Journal of Statistical Physics, 146(6):1203–1212, March 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0435-8.

Higuchi:2012:SRT

[HTZ12] Yasunari Higuchi, Masato Takei, and Yu Zhang. Scaling relations for two-dimensional Ising percolation. Journal of Statistical Physics, 148(5):777–799, September 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0561-3.

Huberman:2013:SCA

[Hub13] Bernardo A. Huberman. Social computing and the attention economy. Journal of Statistical Physics, 151(1–2):329– 339, April 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0596-5.

Huillet:2017:RED

[Hui17] Thierry E. Huillet. Random evolutionary dynamics driven by fitness and house-of-cards mutations: Sampling formulae. Journal of Statistical Physics, 168(1):15–42, July 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Huillet:2018:KMM

[Hui18] Thierry E. Huillet. Karlin–McGregor mutational occupancy problem revisited. Journal of Statistical Physics, 171(6):1136– 1149, June 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Huveneers:2012:ETT

[Huv12] François Huveneers. Energy transport through rare collisions. Journal of Statistical Physics, 146(1):73–97, January 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0374-9.

Hagedorn:2014:MRR

[HV14] George A. Hagedorn and Edward F. Valeev. Molecular resonance Raman and Rayleigh scattering stimulated by a short laser pulse. Journal of Statistical Physics, 154(1–2):522–542, January 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0820-y.

Heydenreich:2014:HDI

[HvdHH14] Markus Heydenreich, Remco van der Hofstad, and Tim Hulshof. High-dimensional incipient infinite clusters revisited. Journal of Statistical Physics, 155(5):966–1025, June 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-0979-x.

Hottovy:2012:NID

[HVW12] Scott Hottovy, Giovanni Volpe, and Jan Wehr. Noiseinduced drift in stochastic differential equations with arbitrary friction and diffusion in the Smoluchowski–Kramers limit. Journal of Statistical Physics, 146(4):762–773, February 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0418-9.

Huhn:2013:FEM

[HW13] W. P. Huhn and M. Widom. A free energy model of boron carbide. Journal of Statistical Physics, 150(3):432–441, February 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0642-3.

Han:2019:BLM

[HWZ19] Yong Han, Yuefei Wang, and Michel Zinsmeister. On the Brownian loop measure. Journal of Statistical Physics, 175 (5):987–1005, June 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ha:2015:NII

[HX15a] Seung-Yeal Ha and Qinghua Xiao. Nonlinear instability of the incoherent state for the Kuramoto–Sakaguchi–Fokker–Plank equation. Journal of Statistical Physics, 160(2):477–496, July 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1270-5.

Ha:2015:RST

[HX15b] Seung-Yeal Ha and Qinghua Xiao. A revisiting of the L^{2}-stability theory of the Boltzmann equation near global Maxwellians. Journal of Statistical Physics, 160(2):430–465, July 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1252-7.

Haydn:2016:ETD

[HY16] N. Haydn and F. Yang. Entry times distribution for mixing systems. Journal of Statistical Physics, 163(2):374–392, April 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1487-y.

Haydn:2017:ETD

[HY17] N. Haydn and F. Yang. Entry times distribution for dynamical balls on metric spaces. Journal of Statistical Physics, 167(2):297–316, April 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-017-1745-7.

Hingant:2019:BDP

[HY19a] Erwan Hingant and Romain Yvinec. The Becker–Döring process: Pathwise convergence and phase transition phenomena. Journal of Statistical Physics, 177(3):506–527, November 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Hwang:2019:AWM

[HY19b] Byung-Hoon Hwang and Seok-Bae Yun. Anderson-witting model of the relativistic Boltzmann equation near equilibrium.

Journal of Statistical Physics, 176(4):1009–1045, August 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Hernandez:2013:EDR

[HZ13] Damián G. Hernández and Damián H. Zanette. Evolutionary dynamics of resource allocation in the Colonel Blotto game. Journal of Statistical Physics, 151(3–4):623–636, May 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0659-7.

Heo:2011:DAA

[HZS11] Muyoung Heo, Konstantin B. Zeldovich, and Eugene I. Shakhnovich. Diversity against adversity: How adaptive immune system evolves potent antibodies. Journal of Statistical Physics, 144(2):241–267, July 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0241-8.

Izquierdo:2013:SAU

[II13] Segismundo S. Izquierdo and Luis R. Izquierdo. Stochastic approximation to understand simple simulation models. Journal of Statistical Physics, 151(1–2):254–276, April 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0654-z.

Itami:2015:DSL

[IiS15a] Masato Itami and Shin ichi Sasa. Derivation of Stokes’ Law from Kirkwood’s formula and the Green–Kubo formula via large deviation theory. Journal of Statistical Physics, 161(3): 532–552, November 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1349-z.

Itami:2015:NSM

[IiS15b] Masato Itami and Shin ichi Sasa. Nonequilibrium statistical mechanics for adiabatic piston problem. Journal of Statistical Physics, 158(1):37–56, January 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1115-7.

Itami:2017:UFS

[IiS17] Masato Itami and Shin ichi Sasa. Universal form of stochastic evolution for slow variables in equilibrium systems. Journal of Statistical Physics, 167(1):46–63, April 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-017-1738-6.

Iacobelli:2010:MFT

[IK10] Giulio Iacobelli and Christof Külske. Metastates in finite-type mean-field models: Visibility, invisibility, and random restoration of symmetry. Journal of Statistical Physics, 140(1):27– 55, July 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-010-9979-7.pdf.

Bommarito:2017:MMU

[IK17] Michael J. Bommarito II and Daniel Martin Katz. Measuring and modeling the U.S. regulatory ecosystem. Journal of Statistical Physics, 168(5):1125–1135, September 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-017-1846-3.pdf.

Iorgov:2011:ICE

[IL11] N. Iorgov and O. Lisovyy. Ising correlations and elliptic determinants. Journal of Statistical Physics, 143(1):33–59, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0154-6.

Iacobucci:2010:TCT

[ILOS10] Alessandra Iacobucci, Frédéric Legoll, Stefano Olla, and Gabriel Stoltz. Thermal conductivity of the Toda lattice with conservative noise. Journal of Statistical Physics, 140(2):336– 348, July 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-9996-6.

Ilyin:2012:ASS

[Ily12] Oleg Ilyin. The analytical solutions of 2D stationary Broadwell kinetic model. Journal of Statistical Physics, 146(1):67– 72, January 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0393-6.

Ilyin:2016:ESN

[Ily16] Oleg Ilyin. Exact stationary and non-stationary solutions to inelastic Maxwell model with infinite energy. Journal of Statistical Physics, 165(4):755–764, November 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1643-4.

Imbrie:2016:LSN

[IM16] John Z. Imbrie and Rajinder Mavi. Level spacing for nonmonotone Anderson models. Journal of Statistical Physics, 162(6):1451–1484, March 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1461-8.

Imbrie:2016:MBL

[Imb16] John Z. Imbrie. On many-body localization for quantum spin chains. Journal of Statistical Physics, 163(5):998–1048, June 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1508-x.

Ikhlef:2012:FSL

[IP12] Yacine Ikhlef and Anita K. Ponsaing. Finite-size left-passage probability in percolation. Journal of Statistical Physics, 149 (1):10–36, October 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0573-z.

Iubini:2014:CDS

[IPP14] Stefano Iubini, Antonio Politi, and Paolo Politi. Coarsening dynamics in a simplified DNLS model. Journal of Statistical Physics, 154(4):1057–1073, February 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0896-4.

Imkeller:2010:FET

[IPS10a] Peter Imkeller, Ilya Pavlyukevich, and Michael Stauch. First exit times of non-linear dynamical systems in R^{d} perturbed by multifractal Lévy noise. Journal of Statistical Physics, 141 (1):94–119, October 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0041-6.

Iorgov:2010:SOM

[IPS^{+}10b] N. Iorgov, S. Pakuliak, V. Shadura, Y. Tykhyy, and G. von Gehlen. Spin operator matrix elements in the superintegrable chiral Potts quantum chain. Journal of Statistical Physics, 139(5):743–768, June 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9972-1.

Imamura:2011:CMA

[IS11] Takashi Imamura and Tomohiro Sasamoto. Current moments of 1D ASEP by duality. Journal of Statistical Physics, 142 (5):919–930, March 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0149-3.

Imamura:2013:SCK

[IS13] Takashi Imamura and Tomohiro Sasamoto. Stationary correlations for the 1D KPZ equation. Journal of Statistical Physics, 150(5):908–939, March 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-013-0710-
3.

Imamura:2016:DSO

[IS16] Takashi Imamura and Tomohiro Sasamoto. Determinantal structures in the O’Connell–Yor directed random polymer model. Journal of Statistical Physics, 163(4):675–713, May 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1492-1.

Islambekov:2012:LRB

[IST12] Umar Islambekov, Robert Sims, and Gerald Teschl. Lieb– Robinson bounds for the Toda lattice. Journal of Statistical Physics, 148(3):440–479, August 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-012-0554-

Ioffe:2015:IVE

[IST15] Dmitry Ioffe, Senya Shlosman, and Fabio Lucio Toninelli. Interaction versus entropic repulsion for low temperature Ising polymers. Journal of Statistical Physics, 158(5):1007–1050, March 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1153-1.

Ilyn:2016:SPE

[ISZ16] A. S. Il’yn, V. A. Sirota, and K. P. Zybin. Statistical properties of the T -exponential of isotropically distributed random matrices. Journal of Statistical Physics, 163(4):765–783, May 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1502-3.

Ilyn:2017:IPR

[ISZ17] A. S. Il’yn, V. A. Sirota, and K. P. Zybin. Infinite products of random isotropically distributed matrices. Journal of Statistical Physics, 166(1):24–38, January 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1675-9.

Itoi:2016:GPO

[Ito16] C. Itoi. General properties of overlap operators in disordered quantum spin systems. Journal of Statistical Physics, 163 (6):1339–1349, June 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1527-7.

Itoi:2017:UNR

[Ito17] C. Itoi. Universal nature of replica symmetry breaking in quantum systems with Gaussian disorder. Journal of Statistical Physics, 167(5):1262–1279, June 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Itoi:2018:ARS

[Ito18] C. Itoi. Absence of replica symmetry breaking in the transverse and longitudinal random field Ising model. Journal of Statistical Physics, 170(4):684–699, February 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Itoi:2019:SAP

[Ito19a] C. Itoi. Self-averaging of perturbation Hamiltonian density in perturbed spin systems. Journal of Statistical Physics, 177 (6):1063–1076, December 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Itoi:2019:ZVP

[Ito19b] C. Itoi. Zero-variance of perturbation Hamiltonian density in perturbed spin systems. Journal of Statistical Physics, 176 (3):556–573, August 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Iliev:2015:PDI

[IvRM15] G. K. Iliev, E. J. Janse van Rensburg, and N. Madras. Phase diagram of inhomogeneous percolation with a defect plane. Journal of Statistical Physics, 158(2):255–299, January 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1125-5.

Iommi:2014:ZTL

[IY14] Godofredo Iommi and Yuki Yayama. Zero temperature limits of Gibbs states for almost-additive potentials. Journal of Statistical Physics, 155(1):23–46, April 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0943-9.

Jaffe:2015:SQR

[Jaf15] Arthur Jaffe. Stochastic quantization, reflection positivity, and quantum fields. Journal of Statistical Physics, 161(1):1– 15, October 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1320-z.

Jansen:2012:MVS

[Jan12] Sabine Jansen. Mayer and virial series at low temperature. Journal of Statistical Physics, 147(4):678–706, June 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0490-1.

Janjigian:2015:LDF

[Jan15a] Chris Janjigian. Large deviations of the free energy in the O’Connell–Yor polymer. Journal of Statistical Physics, 160 (4):1054–1080, August 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1269-y.

Jansen:2015:CVE

[Jan15b] S. Jansen. Cluster and virial expansions for the multi-species Tonks gas. Journal of Statistical Physics, 161(5):1299–1323, December 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1367-x.

Janson:2018:EER

[Jan18] Svante Janson. On edge exchangeable random graphs. Journal of Statistical Physics, 173(3–4):448–484, November 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-017-1832-9.pdf.

Javarone:2015:GNG

[Jav15] Marco Alberto Javarone. Gaussian networks generated by random walks. Journal of Statistical Physics, 159(1):108– 119, April 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1175-8.

Joyce:2011:EEF

[JD11] G. S. Joyce and R. T. Delves. On the exact evaluation of the face-centred cubic lattice Green function. Journal of Statistical Physics, 145(3):613–638, November 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0257-0.

Jiang:2014:RBF

[Jia14] Jianping Jiang. A relationship between fixed time Wiener measures and Wiener measures with fixed endpoints. Journal of Statistical Physics, 156(1):177–188, July 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0998-7.

Jin:2018:FMT

[Jin18] Chunyin Jin. Flocking of the Motsch–Tadmor model with a cut-off interaction function. Journal of Statistical Physics, 171 (2):345–360, April 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Jurcisinova:2012:IMP

[JJ12] E. Jurcisinová and M. Jurcisin. The Ising model on pure Husimi lattices: A general formulation and the critical temperatures. Journal of Statistical Physics, 147(6):1077–1093, July 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0518-6.

Jurcisinova:2014:ESA

[JJB14] E. Jurcisinová, M. Jurcisin, and A. Bobák. The exact solution of the anti-ferromagnetic Ising model with multisite interaction on the simplest pure Husimi lattice. Journal of Statistical Physics, 154(4):1096–1112, February 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0913-7.

Jansen:2012:IMA

[JK12] Sabine Jansen and Wolfgang König. Ideal mixture approximation of cluster size distributions at low density. Journal of Statistical Physics, 147(5):963–980, June 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0499-5.

Jiang:2016:BMP

[JL16] Xiaomeng Jiang and Yong Li. Brownian motion on a pseudo sphere in Minkowski space R_{v}^{l}. Journal of Statistical Physics, 165(1):164–183, October 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1574-0.

Jacq:2017:OQR

[JL17a] Thomas S. Jacq and Carlos F. Lardizabal. Open quantum random walks on the half-line: The Karlin–McGregor formula, path counting and Foster’s Theorem. Journal of Statistical Physics, 169(3):547–594, November 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Jitomirskaya:2017:LBL

[JL17b] Svetlana Jitomirskaya and Wencai Liu. A lower bound on the Lyapunov exponent for the generalized Harper’s model. Journal of Statistical Physics, 166(3–4):609–617, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1543-7.

Jahangoshahi:2018:SPF

[JL18] Mohammad Jahangoshahi and Gregory F. Lawler. On the smoothness of the partition function for multiple Schramm– Loewner evolutions. Journal of Statistical Physics, 173(5): 1353–1368, December 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Josserand:2017:TDE

[JLLP17] Christophe Josserand, Martine Le Berre, Thierry Lehner, and Yves Pomeau. Turbulence: Does energy cascade exist? Journal of Statistical Physics, 167(3–4):596–625, May 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Jungel:2011:NDQ

[JLMG11] Ansgar Jüngel, José Luis López, and Jesús Montejo-Gámez. A new derivation of the quantum Navier–Stokes equations in the Wigner–Fokker–Planck approach. Journal of Statistical Physics, 145(6):1661–1673, December 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0388-3.

Joye:2010:DLQ

[JM10] Alain Joye and Marco Merkli. Dynamical localization of quantum walks in random environments. Journal of Statistical Physics, 140(6):1025–1053, September 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0047-0.

Joye:2014:SPQ

[JM14] Alain Joye and Laurent Marin. Spectral properties of quantum walks on rooted binary trees. Journal of Statistical Physics, 155(6):1249–1270, June 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-0950-x.

Johnson:2013:MID

[JMH13] Neil F. Johnson, Pedro Manrique, and Pak Ming Hui. Modeling insurgent dynamics including heterogeneity. Journal of Statistical Physics, 151(3–4):395–413, May 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0706-z.

Junior:2016:DSS

[JMRC16] Valdivino Vargas Junior, Fábio Prates Machado, and Alejandro Roldán-Correa. Dispersion as a survival strategy. Journal of Statistical Physics, 164(4):937–951, August 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1571-3.

Janczura:2013:SMI

[JMSW13] Joanna Janczura, Monika Maciejewska, Andrzej Szczurek, and Agnieszka Wylomańska. Stochastic modeling of indoor air temperature. Journal of Statistical Physics, 152(5):979–994, September 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-013-0794-9.pdf.

Jovanovski:2017:MIM

[Jov17] Oliver Jovanovski. Metastability for the Ising model on the hypercube. Journal of Statistical Physics, 167(1):135–159, April 2017. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-017-1736-8.pdf.

Jain:2018:JSI

[JP18a] Kavita Jain and Luca Peliti. JSP special issue on statistical theory of biological evolution. Journal of Statistical Physics, 172(1):1–2, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-018-2057-2.
pdf.

Jeblick:2018:DTD

[JP18b] Maximilian Jeblick and Peter Pickl. Derivation of the time dependent two dimensional focusing NLS equation. Journal of Statistical Physics, 172(5):1398–1426, September 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Joubaud:2015:LDS

[JPS15] R. Joubaud, G. A. Pavliotis, and G. Stoltz. Langevin dynamics with space–time periodic nonequilibrium forcing. Journal of Statistical Physics, 158(1):1–36, January 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1118-4.

Jaksic:2017:EFT

[JPS17] V. Jaksić, C.-A. Pillet, and A. Shirikyan. Entropic fluctuations in thermally driven harmonic networks. Journal of Statistical Physics, 166(3–4):926–1015, February 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1625-6.

Jenkinson:2018:RCD

[JPV18] O. Jenkinson, M. Pollicott, and P. Vytnova. Rigorous computation of diffusion coefficients for expanding maps. Journal of Statistical Physics, 170(2):221–253, January 2018. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-017-1930-8.pdf.

Jaksic:2014:EFQ

[JPW14] V. Jaksić, C.-A. Pillet, and M. Westrich. Entropic fluctuations of quantum dynamical semigroups. Journal of Statistical Physics, 154(1–2):153–187, January 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0826-5.

Jarai:2015:ACZ

[JRS15] Antal A. Járai, Frank Redig, and Ellen Saada. Approaching criticality via the zero dissipation limit in the Abelian avalanche model. Journal of Statistical Physics, 159(6):1369– 1407, June 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1231-z.

Jonsson:2011:CNT

[JS11] Thordur Jonsson and Sigurdur Örn Stefánsson. Condensation in nongeneric trees. Journal of Statistical Physics, 142(2):277– 313, January 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-0104-8.

Jana:2017:DSV

[JS17] Indrajit Jana and Alexander Soshnikov. Distribution of singular values of random band matrices; Marchenko–Pastur law and more. Journal of Statistical Physics, 168(5):964–985, September 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Jafarizadeh:2010:EER

[JSJ10] S. Jafarizadeh, R. Sufiani, and M. A. Jafarizadeh. Evaluation of effective resistances in pseudo-distance-regular resistor networks. Journal of Statistical Physics, 139(1):177–199, April 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-009-9909-8.

Janson:2019:LTB

[JSV19] Svante Janson, Vadim Shcherbakov, and Stanislav Volkov. Long term behaviour of a reversible system of interacting random walks. Journal of Statistical Physics, 175(1):71–96, April 2019. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-019-02244-0.pdf.

Just:2010:TMG

[Jus10] Wolfram Just. Toom’s model with Glauber rates: An exact solution using elementary methods. Journal of Statistical Physics, 139(6):985–990, June 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9973-0.

Jones:2019:ACG

[JV19] N. G. Jones and R. Verresen. Asymptotic correlations in gapped and critical topological phases of 1D quantum systems. Journal of Statistical Physics, 175(6):1164–1213, June 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-019-02257-9.pdf.

Janssen:2019:CCC

[JvLS19] A. J. E. M. Janssen, Johan S. H. van Leeuwaarden, and Seva Shneer. Counting cliques and cycles in scale-free inhomogeneous random graphs. Journal of Statistical Physics, 175 (1):161–184, April 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-019-02248-
w.pdf.

Jiang:2018:STP

[JY18] Zhuozhuo Jiang and Weigen Yan. Some two-point resistances of the Sierpinski gasket network. Journal of Statistical Physics, 172(3):824–832, August 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ji:2011:SMS

[JYZ11] Yong-Yun Ji, Wei-Qi Yi, and Lin-Xi Zhang. Simple model study of phase transition properties of isolated and aggregated protein. Journal of Statistical Physics, 142(5):975–983, March 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0148-4.

Jin:2010:ZJP

[JZ10] Xian’an Jin and Fuji Zhang. Zeros of the Jones polynomial for multiple crossing-twisted links. Journal of Statistical Physics, 140(6):1054–1064, September 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0027-4.

Khlyupin:2017:RPT

[KA17] Aleksey Khlyupin and Timur Aslyamov. Random process theory approach to geometric heterogeneous surfaces: Effective fluid–solid interaction. Journal of Statistical Physics, 167(6): 1519–1545, June 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kac:2013:CTL

[Kac13] Mark Kac. On certain Toeplitz-like matrices and their relation to the problem of lattice vibrations. Journal of Statistical Physics, 151(5):785–795, June 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0675-7.

Kadanoff:2013:SWF

[Kad13] Leo P. Kadanoff. Slippery wave functions. Journal of Statistical Physics, 152(5):805–823, September 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0795-8.

Kadanoff:2014:RGS

[Kad14] Leo P. Kadanoff. Reflections on Gibbs: From statistical physics to the amistad V3.0. Journal of Statistical Physics, 156 (1):1–9, July 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1000-4.

Kaneko:2012:ERP

[Kan12] Kunihiko Kaneko. Evolution of robustness and plasticity under environmental fluctuation: Formulation in terms of phenotypic variances. Journal of Statistical Physics, 148(4):687–705, September 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-012-0563-1.pdf.

Kanai:2014:ROB

[Kan14] Masahiro Kanai. Realization of the open-boundary totally asymmetric simple exclusion process on a ring. Journal of Statistical Physics, 157(2):282–294, October 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1089-5.

Kargin:2010:CTQ

[Kar10] Vladislav Kargin. Continuous-time quantum walk on integer lattices and homogeneous trees. Journal of Statistical Physics, 140(2):393–408, July 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9991-y.

Kardar:2011:EBB

[Kar11] Mehran Kardar. Edouard Brézin: Introduction to Statistical Field Theory. Journal of Statistical Physics, 142(5):1121– 1122, March 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0152-8.

Kargin:2012:EST

[Kar12] V. Kargin. On eigenvalues of the sum of two random projections. Journal of Statistical Physics, 149(2):246–258, October 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0592-9.

Kargin:2014:PRP

[Kar14a] V. Kargin. On Pfaffian random point fields. Journal of Statistical Physics, 154(3):681–704, February 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0900-z.

Kargin:2014:LLE

[Kar14b] Vladislav Kargin. On the largest Lyapunov exponent for products of Gaussian matrices. Journal of Statistical Physics, 157 (1):70–83, October 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1077-9.

Kargin:2018:GPF

[Kar18] Vladislav Kargin. A 3D Ginibre point field. Journal of Statistical Physics, 171(6):1067–1095, June 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Katori:2012:DPS

[Kat12a] Makoto Katori. Determinantal process starting from an orthogonal symmetry is a Pfaffian process. Journal of Statistical Physics, 146(2):249–263, January 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0372-y.

Katori:2012:RTR

[Kat12b] Makoto Katori. Reciprocal time relation of noncolliding Brownian motion with drift. Journal of Statistical Physics, 148(1): 38–52, July 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0527-5.

Katori:2012:SPM

[Kat12c] Makoto Katori. Survival probability of mutually Killing Brownian motions and the O’Connell process. Journal of Statistical Physics, 147(1):206–223, April 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0472-3.

Katori:2012:SCB

[Kat12d] Makoto Katori. System of complex Brownian motions associated with the O’Connell process. Journal of Statistical Physics, 149(3):411–431, November 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0602-y.

Katori:2015:DMC

[Kat15] Makoto Katori. Determinantal martingales and correlations of noncolliding random walks. Journal of Statistical Physics, 159(1):21–42, April 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1179-4.

Kautz:2011:CSP

[Kau11] Richard Kautz. Chaos: the science of predictable random motion. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-959457-0 (hardcover), 0-19-959458-9 (paperback). xiii + 369 pp. LCCN QA279.2 .K38 2011.

Kawai:2016:HOF

[Kaw16] Reiichiro Kawai. Higher order fractional stable motion: Hyperdiffusion with heavy tails. Journal of Statistical Physics, 165(1):126–152, October 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1602-0.

Kuchler:2019:ESB

[KBB19] Christian Küchler, Gregory Bewley, and Eberhard Bodenschatz. Experimental study of the bottleneck in fully developed turbulence. Journal of Statistical Physics, 175(3–4):617– 639, May 2019. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-019-02251-1.pdf.

Keller:2013:NCP

[KBLL13] Corina Keller, Florian Berger, Steffen Liepelt, and Reinhard Lipowsky. Network complexity and parametric simplicity for cargo transport by two molecular motors. Journal of Statistical Physics, 150(2):205–234, January 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-012-0662-z.pdf.

Kalyan:2016:JDS

[KBSM16] M. Suman Kalyan, R. Bharath, V. S. S. Sastry, and K. P. N. Murthy. Joint density of states calculation employing Wang– Landau algorithm. Journal of Statistical Physics, 163(1):197– 209, April 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1472-5.

Kawagoe:2018:PBI

[KC18] Daisuke Kawagoe and I.-Kun Chen. Propagation of boundaryinduced discontinuity in stationary radiative transfer. Journal of Statistical Physics, 170(1):127–140, January 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kohler:2019:TIU

[KC19] Tamara Kohler and Toby Cubitt. Translationally invariant universal classical Hamiltonians. Journal of Statistical Physics, 176(1):228–261, July 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s10955-019-
02295-3.pdf.

Krstulovic:2013:ERD

[KCB13] G. Krstulovic, M. Cencini, and J. Bec. Effective rates in dilute reaction–advection systems for the Annihilation process A + A → ⊘. Journal of Statistical Physics, 153(3):530–550, November 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0823-8.

Koufos:2019:DCA

[KD19] Konstantinos Koufos and Carl P. Dettmann. Distribution of cell area in bounded Poisson Voronoi tessellations with application to secure local connectivity. Journal of Statistical Physics, 176(5):1296–1315, September 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Keskin:2010:MSI

[KE10] Mustafa Keskin and Mehmet Ertas. Mixed-spin Ising model in an oscillating magnetic field and compensation temperature. Journal of Statistical Physics, 139(2):333–344, April 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-9943-6.

Kels:2019:EZI

[Kel19] Andrew P. Kels. Extended Z-invariance for integrable vector and face models and multi-component integrable quad equations. Journal of Statistical Physics, 176(6):1375–1408, September 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kemppainen:2010:SS

[Kem10] Antti Kemppainen. Stationarity of SLE. Journal of Statistical Physics, 139(1):108–121, April 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9929-4.

Kempton:2011:ZTL

[Kem11] Tom Kempton. Zero temperature limits of Gibbs equilibrium states for countable Markov shifts. Journal of Statistical Physics, 143(4):795–806, May 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0195-x.

Kennedy:2010:RGM

[Ken10] Tom Kennedy. Renormalization group maps for Ising models in lattice-gas variables. Journal of Statistical Physics, 140 (3):409–426, August 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0002-0.

Kennedy:2012:TFL

[Ken12] Tom Kennedy. Transforming fixed-length self-avoiding walks into radial SLE_{8/3}. Journal of Statistical Physics, 146(2):281– 293, January 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0406-5.

Kennedy:2015:CIP

[Ken15a] Tom Kennedy. Conformal invariance predictions for the threedimensional self-avoiding walk. Journal of Statistical Physics, 158(6):1195–1212, March 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1183-8.

Kennedy:2015:SKS

[Ken15b] Tom Kennedy. The smart kinetic self-avoiding walk and Schramm Loewner evolution. Journal of Statistical Physics, 160(2):302–320, July 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1271-4.

Kennedy:2016:FOC

[Ken16] Tom Kennedy. The first order correction to the exit distribution for some random walks. Journal of Statistical Physics, 164(1):174–189, July 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1534-8.

Kennedy:2019:CIL

[Ken19a] Tom Kennedy. Conformal invariance of the loop-erased percolation explorer. Journal of Statistical Physics, 177(1):1–19, October 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kennedy:2019:NIR

[Ken19b] Tom Kennedy. A non-intersecting random walk on the Manhattan lattice and SLE_{6}. Journal of Statistical Physics, 174 (1):77–96, January 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kerl:2010:SCT

[Ker10] John Kerl. Shift in critical temperature for random spatial permutations with cycle weights. Journal of Statistical Physics, 140(1):56–75, July 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9988-6.

Kerstein:2013:HPS

[Ker13] Alan R. Kerstein. Hierarchical parcel-swapping representation of turbulent mixing. Part 1. Formulation and scaling properties. Journal of Statistical Physics, 153(1):142–161, October 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-013-0811-z.pdf.

Khanin:2017:MJY

[Kha17] Konstantin Khanin. Mathematical journey of Yakov Sinai. Journal of Statistical Physics, 166(3–4):463–466, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1668-8.

Khalil:2019:MDA

[Kha19] Nagi Khalil. Mesoscopic description of the adiabatic piston: Kinetic equations and H-theorem. Journal of Statistical Physics, 176(5):1138–1160, September 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kalyuzhnyi:2011:NFF

[KID^{+}11] Y. V. Kalyuzhnyi, C. R. Iacovella, H. Docherty, M. Holovko, and P. T. Cummings. Network forming fluids: Yukawa squarewell m-point model. Journal of Statistical Physics, 145(2): 481–506, October 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0264-1.

Kiessling:2011:EHL

[Kie11] Michael K.-H. Kiessling. Elliott H. Lieb & Robert Seiringer: The Stability of Matter in Quantum Mechanics. Journal of Statistical Physics, 144(4):909–917, August 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0259-y.

Kiessling:2013:OCS

[Kie13] Michael K.-H. Kiessling. Order and chaos in some trigonometric series: Curious adventures of a statistical mechanic. Journal of Statistical Physics, 150(3):572–600, February 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0578-7.

Kiessling:2014:MFV

[Kie14] Michael K.-H. Kiessling. The microscopic foundations of Vlasov theory for jellium-like Newtonian N-body systems. Journal of Statistical Physics, 155(6):1299–1328, June 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-0934-x.

Kiessling:2017:BRJ

[Kie17a] Michael K.-H. Kiessling. Book review of Jean Bricmont’s “Making Sense of Quantum Mechanics”. Journal of Statistical Physics, 169(3):685–695, November 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kiessling:2017:HRE

[Kie17b] Michael Karl-Heinz Kiessling. Heuristic relative entropy principles with complex measures: Large-degree asymptotics of a family of multi-variate normal random polynomials. Journal of Statistical Physics, 169(1):63–106, October 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kim:2012:NED

[Kim12] Ilki Kim. Non-equilibrium dynamics in the quantum Brownian oscillator and the second law of thermodynamics. Journal of Statistical Physics, 146(1):217–238, January 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0375-8.

Kuroda:2013:ACE

[KiMM13] Koji Kuroda, Jun ichi Maskawa, and Joshin Murai. Application of the cluster expansion to a mathematical model of the long memory phenomenon in a financial market. Journal of Statistical Physics, 152(4):706–723, August 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0783-z.

Kirov:2012:NTD

[Kir12] Mikhail V. Kirov. New two-dimensional ice models. Journal of Statistical Physics, 149(5):865–877, November 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0632-5.

Kaiser:2017:ACE

[KJZ17] Marcus Kaiser, Robert L. Jack, and Johannes Zimmer. Acceleration of convergence to equilibrium in Markov chains by breaking detailed balance. Journal of Statistical Physics, 168(2):259–287, July 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-017-1805-z.
pdf.

Kaiser:2018:CSO

[KJZ18] Marcus Kaiser, Robert L. Jack, and Johannes Zimmer. Canonical structure and orthogonality of forces and currents in irreversible Markov chains. Journal of Statistical Physics, 170 (6):1019–1050, March 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-018-1986-0.
pdf.

Komech:2014:EES

[KK14] A. Komech and E. Kopylova. On eigenfunction expansion of solutions to the Hamilton equations. Journal of Statistical Physics, 154(1–2):503–521, January 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0846-1.

Kim:2015:BMR

[KK15] Changho Kim and George Em Karniadakis. Brownian motion of a Rayleigh particle confined in a channel: A generalized Langevin equation approach. Journal of Statistical Physics, 158(5):1100–1125, March 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1160-2.

Komech:2016:LSC

[KK16] A. Komech and E. Kopylova. On the linear stability of crystals in the Schrödinger–Poisson model. Journal of Statistical Physics, 165(2):246–273, October 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-016-1613-x.pdf.

Kosuge:2019:KMP

[KKA19] Shingo Kosuge, Hung-Wen Kuo, and Kazuo Aoki. A kinetic model for a polyatomic gas with temperature-dependent specific heats and its application to shock-wave structure. Journal of Statistical Physics, 177(2):209–251, October 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kosmrlj:2012:ICD

[KKC12] Andrej Kosmrlj, Mehran Kardar, and Arup K. Chakraborty. The influence of T cell development on pathogen specificity and autoreactivity. Journal of Statistical Physics, 149(2):203– 219, October 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0403-8.

Kaminaga:2012:NAD

[KKN12] M. Kaminaga, M. Krishna, and S. Nakamura. A note on the analyticity of density of states. Journal of Statistical Physics, 149(3):496–504, November 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0603-x.

Kirsch:2018:LTR

[KKR18] Werner Kirsch, David Krejcirík, and Georgi Raikov. Lifshits tails for randomly twisted quantum waveguides. Journal of Statistical Physics, 171(3):383–399, May 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ko:2019:CLT

[KKSY19] Chul Ki Ko, Norio Konno, Etsuo Segawa, and Hyun Jae Yoo. Central limit theorems for open quantum random walks on the crystal lattices. Journal of Statistical Physics, 176(3):710– 735, August 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kanter:2011:NMS

[KKV^{+}11] Ido Kanter, Evi Kopelowitz, Roni Vardi, Meital Zigzag, Dana Cohen, and Wolfgang Kinzel. Nonlocal mechanism for synchronization of time delay networks. Journal of Statistical Physics, 145(3):713–733, November 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0361-1.

Karlstrom:2011:DOM

[KL11] Gunnar Karlström and Per Linse. Dipolar order in molecular fluids: I. Toward an understanding. Journal of Statistical Physics, 145(2):410–417, October 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-011-0353-1.pdf.

Koskinen:2013:LSG

[KL13] Johan Koskinen and Alessandro Lomi. The local structure of globalization. Journal of Statistical Physics, 151(3–4):523– 548, May 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0732-x.

Kessler:2015:SSL

[KL15] David A. Kessler and Herbert Levine. Scaling solution in the large population limit of the general asymmetric stochastic Luria–Delbrück evolution process. Journal of Statistical Physics, 158(4):783–805, February 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1143-3.

Kotecky:2019:SLR

[KL19] Roman Kotecký and Benjamin Lees. Staggered longrange order for diluted quantum spin models. Journal of Statistical Physics, 175(5):972–986, June 2019. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-019-02263-x.pdf.

Klages:2011:RKB

[Kla11] Rainer Klages. Richard Kautz: Chaos — The Science of Predictable Random Motion. Journal of Statistical Physics, 144 (4):918–919, August 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0282-z.

Krajenbrink:2019:DBC

[KLACTL19] Alexandre Krajenbrink, Bertrand Lacroix-A-Chez-Toine, and Pierre Le Doussal. Distribution of Brownian coincidences. Journal of Statistical Physics, 177(1):119–150, October 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Klemm:2013:SCN

[Kle13] Konstantin Klemm. Searchability of central nodes in networks. Journal of Statistical Physics, 151(3–4):707–719, May 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0727-7.

Kleeman:2015:PIF

[Kle15] Richard Kleeman. A path integral formalism for nonequilibrium Hamiltonian statistical systems. Journal of Statistical Physics, 158(6):1271–1297, March 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1149-x.

Kager:2013:SLA

[KLM13] Wouter Kager, Marcin Lis, and Ronald Meester. The signed loop approach to the Ising model: Foundations and critical point. Journal of Statistical Physics, 152(2):353–387, July 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0767-z.

Kyprianou:2017:LFH

[KLM17] Andreas Kyprianou, Francis Lane, and Peter Mörters. The largest fragment of a homogeneous fragmentation process. Journal of Statistical Physics, 166(5):1226–1246, March 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-017-1714-1.

Kraaij:2018:DGG

[KLMP18] Richard Kraaij, Alexandre Lazarescu, Christian Maes, and Mark Peletier. Deriving GENERIC from a generalized fluctuation symmetry. Journal of Statistical Physics, 170(3):492–508, February 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kabluchko:2019:FMI

[KLS19] Zakhar Kabluchko, Matthias Löwe, and Kristina Schubert. Fluctuations of the magnetization for Ising models on dense Erdős–Rényi random graphs. Journal of Statistical Physics, 177(1):78–94, October 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Klumpp:2011:PBT

[Klu11] Stefan Klumpp. Pausing and backtracking in transcription under dense traffic conditions. Journal of Statistical Physics, 142(6):1252–1267, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0120-3.

Kirkpatrick:2013:AMF

[KM13] Kay Kirkpatrick and Elizabeth Meckes. Asymptotics of the mean-field Heisenberg model. Journal of Statistical Physics, 152(1):54–92, July 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0753-5.

Kupiainen:2017:RGK

[KM17] Antti Kupiainen and Matteo Marcozzi. Renormalization of generalized KPZ equation. Journal of Statistical Physics, 166(3–4):876–902, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1636-3.

Kwon:2018:SCF

[KM18] Younghak Kwon and Georg Menz. Strict convexity of the free energy of the canonical ensemble under decay of correlations. Journal of Statistical Physics, 172(4):927–979, August 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kroschinsky:2019:MST

[KM19a] Wilhelm Kroschinsky and Domingos H. U. Marchetti. On the Mayer series of two-dimensional Yukawa gas at inverse temperature in the interval of collapse. Journal of Statistical Physics, 177(2):324–364, October 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kumar:2019:DPM

[KM19b] Sameer Kumar and Shradha Mishra. Dynamics of a particle moving in one dimensional Lorentz lattice gas. Journal of Statistical Physics, 176(5):1161–1171, September 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kwon:2019:DCU

[KM19c] Younghak Kwon and Georg Menz. Decay of correlations and uniqueness of the infinite-volume Gibbs measure of the canonical ensemble of 1 d-lattice systems. Journal of Statistical Physics, 176(4):836–872, August 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kessler:2014:DAU

[KMB14] David A. Kessler, Shlomi Medalion, and Eli Barkai. The distribution of the area under a Bessel excursion and its moments. Journal of Statistical Physics, 156(4):686–706, August 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1032-9.

Kimiagar:2011:MPE

[KMKT11] S. Kimiagar, M. Sadegh Movahed, S. Khorram, and M. Reza Rahimi Tabar. Markov properties of electrical discharge current fluctuations in plasma. Journal of Statistical Physics, 143(1):148–167, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0171-5.

Kirsch:2011:RBO

[KMM11] Werner Kirsch, Bernd Metzger, and Peter Müller. Random block operators. Journal of Statistical Physics, 143(6):1035– 1054, June 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0230-y.

Kuniba:2016:IGM

[KMO16] Atsuo Kuniba, Shouya Maruyama, and Masato Okado. Inhomogeneous generalization of a multispecies totally asymmetric zero range process. Journal of Statistical Physics, 164(4):952– 968, August 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1555-3.

Kundu:2014:MDT

[KMS14] Anupam Kundu, Satya N. Majumdar, and Grégory Schehr. Maximal distance travelled by N vicious walkers till their survival. Journal of Statistical Physics, 157(1):124–157, October 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1064-1.

Krapivsky:2015:TPS

[KMS15] P. L. Krapivsky, Kirone Mallick, and Tridib Sadhu. Tagged particle in single-file diffusion. Journal of Statistical Physics, 160(4):885–925, August 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1291-0.

Kilanowski:2019:CFR

[KMS19a] Humbert Philip Kilanowski, Peter March, and Marko Samara. Convergence of the freely rotating chain to the Kratky–Porod model of semi-flexible polymers. Journal of Statistical Physics, 174(6):1222–1238, March 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Koma:2019:QCQ

[KMS19b] Tohru Koma, Toru Morishita, and Taro Shuya. Quantization of conductance in quasi-periodic quantum wires. Journal of Statistical Physics, 174(5):1137–1160, March 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kessebohmer:2019:DRT

[KMSS19] M. Kesseböhmer, A. Mosbach, T. Samuel, and M. Steffens. Diffraction of return time measures. Journal of Statistical Physics, 174(3):519–535, February 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kalmykov:2010:MEP

[KMTC10] Yuri P. Kalmykov, Bernard P. J. Mulligan, Serguey V. Titov, and William T. Coffey. Master equation in phase space for a spin in an arbitrarily directed uniform external field. Journal of Statistical Physics, 141(3):589–606, November 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-0059-9.

Klein:2013:BMA

[KN13] Abel Klein and Son T. Nguyen. The bootstrap multiscale analysis for the multi-particle Anderson model. Journal of Statistical Physics, 151(5):938–973, June 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0734-8.

Kirkpatrick:2016:AMF

[KN16] Kay Kirkpatrick and Tayyab Nawaz. Asymptotics of meanfield O(N) models. Journal of Statistical Physics, 165(6): 1114–1140, December 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1667-9.

Komatsu:2011:ENN

[KNiST11] Teruhisa S. Komatsu, Naoko Nakagawa, Shin ichi Sasa, and Hal Tasaki. Entropy and nonlinear nonequilibrium thermodynamic relation for heat conducting steady states. Journal of Statistical Physics, 142(1):127–153, January 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0095-5.

Komatsu:2015:EET

[KNiST15] Teruhisa S. Komatsu, Naoko Nakagawa, Shin ichi Sasa, and Hal Tasaki. Exact equalities and thermodynamic relations for nonequilibrium steady states. Journal of Statistical Physics, 159(6):1237–1285, June 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1221-1.

Kenmoe:2015:EPS

[KNK15] G. Djuidjé Kenmoé, Y. J. Wadop Ngouongo, and T. C. Kofané. Effect of the potential shape on the stochastic resonance processes. Journal of Statistical Physics, 161(2):475– 485, October 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1333-7.

Kharmiani:2019:AHD

[KNPF19] Soroush Fallah Kharmiani, Hamid Niazmand, and Mohammad Passandideh-Fard. An alternative high-density ratio pseudo-potential lattice Boltzmann model with surface tension adjustment capability. Journal of Statistical Physics, 175 (1):47–70, April 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Krug:2018:LST

[KNSS18] J. Krug, R. A. Neiss, A. Schadschneider, and J. Schmidt. Logarithmic superdiffusion in two dimensional driven lattice gases. Journal of Statistical Physics, 172(2):493–504, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kovchegov:2015:RMG

[KO15] Yevgeniy Kovchegov and Peter T. Otto. Rapid mixing of Glauber dynamics of Gibbs ensembles via aggregate path coupling and large deviations methods. Journal of Statistical Physics, 161(3):553–576, November 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1345-3.

Ko:2019:PSW

[Ko19] Dongnam Ko. Practical synchronization of winfree oscillators in a random environment. Journal of Statistical Physics, 174 (6):1263–1287, March 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Koibuchi:2010:FHM

[Koi10] Hiroshi Koibuchi. Flat histogram Monte Carlo simulations of triangulated fixed-connectivity surface models. Journal of Statistical Physics, 140(4):676–687, August 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0011-z.

Koike:2018:MRB

[Koi18] Kai Koike. Motion of a rigid body in a special Lorentz gas: Loss of memory effect. Journal of Statistical Physics, 172(3): 795–823, August 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kolesnik:2014:PLE

[Kol14] Alexander D. Kolesnik. Probability law for the Euclidean distance between two planar random flights. Journal of Statistical Physics, 154(4):1124–1152, February 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0901-y.

Kolesnik:2017:ART

[Kol17] Alexander D. Kolesnik. Asymptotic relation for the transition density of the three-dimensional Markov random flight on small time intervals. Journal of Statistical Physics, 166 (2):434–452, January 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1692-8.

Korobov:2016:RRS

[Kor16] A. Korobov. Reversible reshaping of supported metal nanoislands under reaction conditions in a minimalistic lattice model. Journal of Statistical Physics, 163(3):576–592, May 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1494-z.

Kosmrlj:2011:TSC

[Kos11] Andrej Kosmrlj. Thymic selection of T cells as diffusion with intermittent traps. Journal of Statistical Physics, 142 (6):1277–1286, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0144-8.

Kosygina:2013:CSR

[Kos13] Elena Kosygina. Crossing speeds of random walks among ‘sparse’ or ‘spiky’ Bernoulli potentials on integers. Journal of Statistical Physics, 152(2):213–236, July 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0765-1.

Kozachenko:2018:EFS

[KOSV18] Yu. Kozachenko, E. Orsingher, L. Sakhno, and O. Vasylyk. Estimates for functionals of solutions to higher-order heat-type equations with random initial conditions. Journal of Statistical Physics, 172(6):1641–1662, September 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kovchegov:2011:MTM

[KOT11] Yevgeniy Kovchegov, Peter T. Otto, and Mathew Titus. Mixing times for the mean-field Blume–Capel model via aggregate path coupling. Journal of Statistical Physics, 144(5): 1009–1027, September 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0286-8.

Kozhan:2017:RON

[Koz17] Rostyslav Kozhan. Rank one non-Hermitian perturbations of Hermitian β-ensembles of random matrices. Journal of Statistical Physics, 168(1):92–108, July 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-017-1792-0.pdf.

Kolesnik:2011:RED

[KP11] Alexander D. Kolesnik and Mark A. Pinsky. Random evolutions are driven by the hyperparabolic operators. Journal of Statistical Physics, 142(4):828–846, February 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0131-0.

Kalinay:2012:PSR

[KP12] Pavol Kalinay and Jerome K. Percus. Phase space reduction of the one-dimensional Fokker–Planck (Kramers) equation. Journal of Statistical Physics, 148(6):1135–1155, September 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0570-2.

Kao:2018:LCM

[KP18] Ta-Chu Kao and Mason A. Porter. Layer communities in multiplex networks. Journal of Statistical Physics, 173(3–4): 1286–1302, November 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kang:2018:EMB

[KPR18] Mihyun Kang, Angelica Pachon, and Pablo M. Rodríguez. Evolution of a modified binomial random graph by agglomeration. Journal of Statistical Physics, 170(3):509–535, February 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kerner:2019:BEC

[KPS19] Joachim Kerner, Maximilian Pechmann, and Wolfgang Spitzer. On Bose–Einstein condensation in the Luttinger–Sy model with finite interaction strength. Journal of Statistical Physics, 174(6):1346–1371, March 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kondratiev:2016:QCC

[KPZ16] Yuri Kondratiev, Sergey Pirogov, and Elena Zhizhina. A quasispecies continuous contact model in a critical regime. Journal of Statistical Physics, 163(2):357–373, April 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1480-5.

Koch:2010:MQP

[KR10] Hans Koch and Charles Radin. Modelling quasicrystals at positive temperature. Journal of Statistical Physics, 138(1– 3):465–475, February 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9896-9.

Kuelske:2015:ETI

[KR15] C. Kuelske and U. A. Rozikov. Extremality of translation– invariant phases for a three-state SOS-model on the binary tree. Journal of Statistical Physics, 160(3):659–680, August 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1279-9.

Kappen:2016:AIS

[KR16] H. J. Kappen and H. C. Ruiz. Adaptive importance sampling for control and inference. Journal of Statistical Physics, 162(5):1244–1266, March 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-016-1446-7.
pdf.

Kraaij:2016:LDF

[Kra16] Richard Kraaij. Large deviations for finite state Markov jump processes with mean-field interaction via the comparison principle for an associated Hamilton–Jacobi equation. Journal of Statistical Physics, 164(2):321–345, July 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-016-1542-8.pdf.

Krapivsky:2010:KVS

[KRBN10] Pavel L. Krapivsky, Sidney Redner, and E. (Eli) Ben-Naim. A kinetic view of statistical physics. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-85103-3. xv + 488 pp. LCCN QC174.86.N65 K73

2010.  URL http://assets.cambridge.org/97805218/
51039/cover/9780521851039.jpg;   http://www.cambridge.
org/9780521851039;   http://www.loc.gov/catdir/enhancements/
fy1012/2010033852-b.html;   http://www.loc.gov/catdir/
enhancements/fy1012/2010033852-d.html;   http://www.
loc.gov/catdir/enhancements/fy1012/2010033852-t.html.

Kulske:2014:DTI

[KRK14] C. Külske, U. A. Rozikov, and R. M. Khakimov. Description of the translation-invariant splitting Gibbs measures for the Potts model on a Cayley tree. Journal of Statistical Physics, 156(1):189–200, July 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-0986-y.

Kenyon:2017:MSP

[KRRS17] Richard Kenyon, Charles Radin, Kui Ren, and Lorenzo Sadun. Multipodal structure and phase transitions in large constrained graphs. Journal of Statistical Physics, 168(2):233– 258, July 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kruger:2012:CES

[Krü12] Helge Krüger. Concentration of eigenvalues for skew-shift Schrödinger operators. Journal of Statistical Physics, 149 (6):1096–1111, December 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0650-3.

Komorowski:2012:LTL

[KS12] Tomasz Komorowski and Lukasz Stȩpień. Long time, large scale limit of the Wigner transform for a system of linear oscillators in one dimension. Journal of Statistical Physics, 148(1): 1–37, July 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-012-0528-4.pdf.

Khachatryan:2013:SYB

[KS13] S. Khachatryan and A. Sedrakyan. On the solutions of the Yang–Baxter equations with general inhomogeneous eightvertex R-matrix: Relations with Zamolodchikov’s tetrahedral algebra. Journal of Statistical Physics, 150(1):130–155, January 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0666-8.

Kraemer:2014:ZDO

[KS14] Atahualpa S. Kraemer and David P. Sanders. Zero density of open paths in the Lorentz mirror model for arbitrary mirror probability. Journal of Statistical Physics, 156(5):908–916, September 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1038-3.

Kulske:2018:NRP

[KS18] C. Külske and P. Schriever. Non-robust phase transitions in the generalized clock model on trees. Journal of Statistical Physics, 170(1):1–21, January 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kurushina:2019:HDO

[KS19] S. E. Kurushina and E. A. Shapovalova. How disorder originates and grows inside order. Journal of Statistical Physics, 176(1):69–84, July 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kirakosyan:2011:FGL

[KSH11] Zara Kirakosyan, David B. Saakian, and Chin-Kun Hu. Finite genome length corrections for the mean fitness and gene probabilities in evolution models. Journal of Statistical Physics, 144(1):198–212, July 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0254-3.

Kruger:2016:SPU

[KSM16] Benedikt Krüger, Ella M. Schmidt, and Klaus Mecke. Spectral properties of unimodular lattice triangulations. Journal of Statistical Physics, 163(3):514–543, May 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1493-0.

Kanazawa:2015:ADL

[KSSH15] Kiyoshi Kanazawa, Tomohiko G. Sano, Takahiro Sagawa, and Hisao Hayakawa. Asymptotic derivation of Langevin-like equation with non–Gaussian noise and its analytical solution. Journal of Statistical Physics, 160(5):1294–1335, September 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1286-x.

Kelbert:2013:MWT

[KSY13] M. Kelbert, Yu. Suhov, and A. Yambartsev. A Mermin– Wagner theorem for Gibbs states on Lorentzian triangulations. Journal of Statistical Physics, 150(4):671–677, February 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0698-8.

Katori:2011:NSB

[KT11a] Makoto Katori and Hideki Tanemura. Noncolliding squared Bessel processes. Journal of Statistical Physics, 142(3): 592–615, February 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0117-y.

Krasnyuk:2011:STL

[KT11b] Igor B. Krasnyuk and Roman M. Taranets. The spatialtemporal lamellar structures in the confined ideal polymer blends. Journal of Statistical Physics, 145(6):1485–1498, December 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0378-5.

Kim:2012:ASC

[KT12] Woo Kyun Kim and Ellad B. Tadmor. An analytical selfconsistent solution for the free energy of a 1-D chain of atoms including anharmonic effects. Journal of Statistical Physics, 148(5):951–971, September 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0559-x.

Kari:2015:SMS

[KT15] Jarkko Kari and Siamak Taati. Statistical mechanics of surjective cellular automata. Journal of Statistical Physics, 160 (5):1198–1243, September 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-015-1281-2.
pdf.

Kuang:2017:GHM

[KT17] Yangyu Kuang and Huazhong Tang. Globally hyperbolic moment model of arbitrary order for one-dimensional special relativistic Boltzmann equation. Journal of Statistical Physics, 167(5):1303–1353, June 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kubota:2019:GFS

[KT19] Naoki Kubota and Masato Takei. Gaussian fluctuation for superdiffusive elephant random walks. Journal of Statistical Physics, 177(6):1157–1171, December 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kauf:2010:SIC

[KTJ10] Peter Kauf, Manuel Torrilhon, and Michael Junk. Scaleinduced closure for approximations of kinetic equations. Journal of Statistical Physics, 141(5):848–888, December 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-0073-y.

Kolomeisky:2011:HIC

[KU11] Anatoly B. Kolomeisky and Karthik Uppulury. How interactions control molecular transport in channels. Journal of Statistical Physics, 142(6):1268–1276, April 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0069-7.

Kuan:2013:ADT

[Kua13] Jeffrey Kuan. Asymptotics of a discrete-time particle system near a reflecting boundary. Journal of Statistical Physics, 150 (2):398–411, January 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0681-9.

Kumar:2019:RSE

[Kum19] Santosh Kumar. Recursion for the smallest eigenvalue density of β-Wishart–Laguerre ensemble. Journal of Statistical Physics, 175(1):126–149, April 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kuo:2015:EEM

[Kuo15] Hung-Wen Kuo. Equilibrating effect of Maxwell-type boundary condition in highly rarefied gas. Journal of Statistical Physics, 161(3):743–800, November 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1355-1.

Kuo:2017:ABR

[Kuo17] Hung-Wen Kuo. Asymptotic behavior for Rayleigh problem based on kinetic theory. Journal of Statistical Physics, 166(5):1247–1275, March 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-017-1717-y.

Kurchan:2018:QBC

[Kur18] Jorge Kurchan. Quantum bound to chaos and the semiclassical limit. Journal of Statistical Physics, 171(6):965–979, June 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kierkels:2015:TET

[KV15] A. H. M. Kierkels and J. J. L. Velázquez. On the transfer of energy towards infinity in the theory of weak turbulence for the nonlinear Schrödinger equation. Journal of Statistical Physics, 159(3):668–712, May 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1194-0.

Kierkels:2016:SSS

[KV16a] A. H. M. Kierkels and J. J. L. Velázquez. On self-similar solutions to a kinetic equation arising in weak turbulence theory for the nonlinear Schrödinger equation. Journal of Statistical Physics, 163(6):1350–1393, June 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1505-0.

Komjathy:2016:FPP

[KV16b] Júlia Komjáthy and Viktória Vadon. First passage percolation on the Newman–Watts small world model. Journal of Statistical Physics, 162(4):959–993, February 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-015-1442-3.pdf.

Kifer:2017:TPR

[KV17] Yuri Kifer and S. R. S. Varadhan. Tails of polynomials of random variables and stable limits for nonconventional sums. Journal of Statistical Physics, 166(3–4):575–608, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1561-5.

Kiessling:2012:OEP

[KW12] Michael K.-H. Kiessling and Yu Wang. Onsager’s ensemble for point vortices with random circulations on the sphere. Journal of Statistical Physics, 148(5):896–932, September 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0552-4.

Kucherenko:2015:LPE

[KW15] Tamara Kucherenko and Christian Wolf. Localized pressure and equilibrium states. Journal of Statistical Physics, 160 (6):1529–1544, September 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1289-7.

Kirr:2014:DSS

[KWZ14] Eduard Kirr, Mark Wilkinson, and Arghir Zarnescu. Dynamic statistical scaling in the Landau-de Gennes theory of nematic liquid crystals. Journal of Statistical Physics, 155(4):625– 657, May 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-0970-6.

Krikun:2012:PTI

[KY12] Maxim Krikun and Anatoly Yambartsev. Phase transition for the Ising model on the critical Lorentzian triangulation. Journal of Statistical Physics, 148(3):422–439, August 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0548-0.

Konno:2013:LTO

[KY13] Norio Konno and Hyun Jae Yoo. Limit theorems for open quantum random walks. Journal of Statistical Physics, 150 (2):299–319, January 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0668-6.

Kobayashi:2016:FRE

[KYA16] Tetsuya J. Kobayashi, Ryo Yokota, and Kazuyuki Aihara. Feedback regulation and its efficiency in biochemical networks. Journal of Statistical Physics, 162(5):1425–1449, March 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-015-1443-2.pdf.

Lafleche:2019:PMS

[Laf19] Laurent Lafleche. Propagation of moments and semiclassical limit from Hartree to Vlasov equation. Journal of Statistical Physics, 177(1):20–60, October 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lambert:2019:IDP

[Lam19] Gaultier Lambert. Incomplete determinantal processes: From random matrix to Poisson statistics. Journal of Statistical Physics, 176(6):1343–1374, September 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Langmann:2010:LM

[Lan10] Edwin Langmann. A 2D Luttinger model. Journal of Statistical Physics, 141(1):17–52, October 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0029-2.

Lan:2011:NCG

[Lan11] Yueheng Lan. Novel computation of the growth rate of generalized random Fibonacci sequences. Journal of Statistical Physics, 142(4):847–861, February 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0132-z.

Landy:2013:LCS

[Lan13] Jonathan Landy. Local constraints satisfied by realizable correlation functions. Journal of Statistical Physics, 152(4):698– 705, August 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0788-7.

Landy:2015:ESC

[Lan15] Jonathan Landy. Eigenspace stability conditions in mean– field replica theories. Journal of Statistical Physics, 159(1): 62–74, April 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1172-y.

Lancellotti:2016:TAE

[Lan16] Carlo Lancellotti. Time-asymptotic evolution of spatially uniform Gaussian Vlasov fluctuation fields. Journal of Statistical Physics, 163(4):868–886, May 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1504-1.

Lanchier:2017:RPB

[Lan17] Nicolas Lanchier. Rigorous proof of the Boltzmann–Gibbs distribution of money on connected graphs. Journal of Statistical Physics, 167(1):160–172, April 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-017-1744-8.

Langer:2019:STC

[Lan19] J. S. Langer. Statistical thermodynamics of crystal plasticity. Journal of Statistical Physics, 175(3–4):531–541, May 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Laurencot:2018:UMC

[Lau18] Philippe Laurençot. Uniqueness of mass-conserving selfsimilar solutions to Smoluchowski’s coagulation equation with inverse power law kernels. Journal of Statistical Physics, 171 (3):484–492, May 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lee:2015:SMU

[LBB15] Edward D. Lee, Chase P. Broedersz, and William Bialek. Statistical mechanics of the US Supreme Court. Journal of Statistical Physics, 160(2):275–301, July 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1253-6.

Li:2013:NCO

[LBW^{+}13] Qian Li, Lidia A. Braunstein, Huijuan Wang, Jia Shao, H. Eugene Stanley, and Shlomo Havlin. Non-consensus opinion models on complex networks. Journal of Statistical Physics, 151(1–2):92–112, April 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0625-4.

Loomis:2019:SWO

[LC19] Samuel P. Loomis and James P. Crutchfield. Strong and weak optimizations in classical and quantum models of stochastic processes. Journal of Statistical Physics, 176(6):1317–1342, September 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lin:2015:SRF

[LCZW15] Li-Feng Lin, Cong Chen, Su-Chuan Zhong, and Hui-Qi Wang. Stochastic resonance in a fractional oscillator with random mass and random frequency. Journal of Statistical Physics, 160(2):497–511, July 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1265-2.

Li:2015:FAN

[LE15] Qianxiao Li and Weinan E. The free action of nonequilibrium dynamics. Journal of Statistical Physics, 161(2):300–325, October 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1339-1.

Lamstaes:2017:ABR

[LE17] Catherine Lamstaes and Jens Eggers. Arrested bubble rise in a narrow tube. Journal of Statistical Physics, 167(3–4):656– 682, May 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lebowitz:2010:PSM

[Leb10] Joel L. Lebowitz. Program of the 103rd Statistical Mechanics Conference. Journal of Statistical Physics, 140(4):812–817, August 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-0008-7.

Lebowitz:2011:PKS

[Leb11] Joel L. Lebowitz. Pavel L. Krapivsky, Sidney Redner and Eli Ben-Naim: A Kinetic View of Statistical Physics. Journal of Statistical Physics, 143(4):831–832, May 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0198-7.

Lebowitz:2012:EPS

[Leb12a] Joel L. Lebowitz. Erratum to: Program of the 105th Statistical Mechanics Conference, Rutgers University, Busch Campus, Hill Center, Room 114, Sunday, Monday, Tuesday, May 8– 10, 2011. Journal of Statistical Physics, 146(4):882, February 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/
pdf/10.1007/s10955-012-0434-9.pdf. See [Leb12c].

Lebowitz:2012:P

[Leb12b] Joel L. Lebowitz. Preface. Journal of Statistical Physics, 148(4):607, September 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-012-0566-y.
pdf.

Lebowitz:2012:PSM

[Leb12c] Joel L. Lebowitz. Program of the 105th Statistical Mechanics Conference, Rutgers University, Busch Campus, Hill Center, Room 114, Sunday, Monday, Tuesday, May 8–10, 2011. Journal of Statistical Physics, 146(3):663–668, February 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0387-4. See erratum [Leb12a].

Lebowitz:2013:P

[Leb13a] Joel Lebowitz. Preface. Journal of Statistical Physics, 150 (3):413, February 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-013-0696-x.
pdf.

Lebowitz:2013:PSMa

[Leb13b] Joel L. Lebowitz. Program of the 108th Statistical Mechanics Conference Rutgers University, Busch Campus, Hill Center, Room 114 Sunday, Monday and Tuesday December 16– December 18, 2012. Journal of Statistical Physics, 152(4):799– 803, August 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0792-y.

Lebowitz:2013:PSMb

[Leb13c] Joel L. Lebowitz. Program of the 109th Statistical Mechanics Conference Rutgers University, Busch Campus, Hill Center Sunday, May 12, 2013–Tuesday, May 14, 2013. Journal of Statistical Physics, 152(5):997–1001, September 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0793-x.

Lebowitz:2014:PSM

[Leb14] Joel L. Lebowitz. Program of the 111th Statistical Mechanics Conference Rutgers University, Busch Campus, Hill Center Sunday, May 11, 2014–Tuesday, May 13, 2014. Journal of Statistical Physics, 157(3):603–607, November 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1074-z.

Leble:2016:LCR

[Leb16] Thomas Leblé. Logarithmic, Coulomb and Riesz energy of point processes. Journal of Statistical Physics, 162(4): 887–923, February 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1425-4.

Lebowitz:2019:SID

[Leb19] Joel Lebowitz. Special issue: Dedicated to the memory of Pierre Hohenberg. Journal of Statistical Physics, 175 (3–4):513–514, May 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-019-02285-
5.pdf.

Lee:2010:DPP

[Lee10] Eunghyun Lee. Distribution of a Particle’s position in the ASEP with the alternating initial condition. Journal of Statistical Physics, 140(4):635–647, August 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-010-0014-9.pdf.

Lee:2011:TPB

[Lee11] Eunghyun Lee. Transition probabilities of the Bethe ansatz solvable interacting particle systems. Journal of Statistical Physics, 142(4):643–656, February 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s10955-011-
0139-5.pdf.

Lee:2012:CDM

[Lee12] Eunghyun Lee. The current distribution of the multiparticle hopping asymmetric diffusion model. Journal of Statistical Physics, 149(1):50–72, October 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0582-y.

Lee:2018:PIB

[Lee18] Edward D. Lee. Partisan intuition belies strong, institutional consensus and wide Zipf’s Law for voting blocs in US Supreme Court. Journal of Statistical Physics, 173(6):1722–1733, December 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lee:2019:TDR

[Lee19] Jinyeop Lee. On the time dependence of the rate of convergence towards Hartree dynamics for interacting bosons. Journal of Statistical Physics, 176(2):358–381, July 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lefevere:2013:MDH

[Lef13] Raphaël Lefevere. Macroscopic diffusion from a Hamilton-like dynamics. Journal of Statistical Physics, 151(5):861–869, June 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0738-4.

Lehle:2013:STS

[Leh13] Bernd Lehle. Stochastic time series with strong, correlated measurement noise: Markov analysis in N dimensions.

Journal of Statistical Physics, 152(6):1145–1169, September 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0803-z.

Leplaideur:2012:FCS

[Lep12] Renaud Leplaideur. Flatness is a criterion for selection of maximizing measures. Journal of Statistical Physics, 147(4):728– 757, June 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0497-7.

Leplaideur:2015:CBG

[Lep15] Renaud Leplaideur. Chaos: Butterflies also generate phase transitions. Journal of Statistical Physics, 161(1):151–170, October 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1299-5.

Lucarini:2012:UBE

[LFW12] Valerio Lucarini, Davide Faranda, and Jeroen Wouters. Universal behaviour of extreme value statistics for selected observables of dynamical systems. Journal of Statistical Physics, 147(1):63–73, April 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0468-z.

Lucarini:2014:TGT

[LFWK14] Valerio Lucarini, Davide Faranda, Jeroen Wouters, and Tobias Kuna. Towards a general theory of extremes for observables of chaotic dynamical systems. Journal of Statistical Physics, 154(3):723–750, February 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-013-0914-6.pdf.

Lee:2011:ADS

[LH11] Chi-Lun Lee and Chia-Ling Huang. Asymptotic dynamics of self-driven vehicles in a closed boundary. Journal of Statistical Physics, 144(4):813–825, August 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0262-3.

Li:2013:TSC

[LH13a] Yan Li and Dong Han. A two-stage contact process on scalefree networks. Journal of Statistical Physics, 153(2):312–324, October 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0832-7.

Luposchainsky:2013:EPC

[LH13b] David Luposchainsky and Haye Hinrichsen. Entropy production in continuous phase space systems. Journal of Statistical Physics, 153(5):828–841, December 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0863-0.

Lagache:2017:ENE

[LH17] T. Lagache and D. Holcman. Extended narrow escape with many windows for analyzing viral entry into the cell nucleus. Journal of Statistical Physics, 166(2):244–266, January 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1691-9.

Lafhal:2019:FTP

[LHZ^{+}19] A. Lafhal, N. Hachem, H. Zahir, M. El Bouziani, M. Madani, and A. Alrajhi. Finite temperature phase diagrams of the mixed spin-1 and spin-2 Blume–Capel model by renormalization group approach. Journal of Statistical Physics, 174 (1):40–55, January 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Li:2012:BCP

[Li12] Dong Li. Bifurcation of critical points for solutions of the 2D Euler and 2D quasi-geostrophic equations. Journal of Statistical Physics, 149(1):92–107, October 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0583-x.

Li:2019:MTM

[Li19] Zhongyang Li. Mixing time of Markov chains for the 1– 2 model. Journal of Statistical Physics, 176(6):1526–1560, September 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Liaw:2013:AES

[Lia13] Constanze Liaw. Approach to the extended states conjecture. Journal of Statistical Physics, 153(6):1022–1038, December 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0879-5.

Liang:2018:MMB

[Lia18] Song Liang. A mechanical model of Brownian motion for one massive particle including slow light particles. Journal of Statistical Physics, 170(2):286–350, January 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Liao:2019:RSE

[Lia19] Gang Liao. Regularity of SRB entropy for geometric Lorenz attractors. Journal of Statistical Physics, 174(3):536–547, February 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Liechty:2012:NBM

[Lie12] Karl Liechty. Nonintersecting Brownian motions on the halfline and discrete Gaussian orthogonal polynomials. Journal of Statistical Physics, 147(3):582–622, May 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0485-y.

Lima:2016:LTA

[Lim16] Paulo C. Lima. Low temperature analysis of correlation functions of the Blume–Emery–Griffiths model at the antiquadrupolar-disordered interface. Journal of Statistical Physics, 165(3):645–660, November 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-016-1631-
8.

Lis:2017:PIM

[Lis17] Marcin Lis. The planar Ising model and total positivity. Journal of Statistical Physics, 166(1):72–89, January 2017.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-016-1690-x.pdf.

Liu:2015:SWP

[Liu15a] Gi-Ren Liu. Stochastic wave propagation in Maxwell’s equations. Journal of Statistical Physics, 158(5):1126–1146, March 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1148-y.

Liu:2015:PTG

[Liu15b] Hanze Liu. Painlevé test, generalized symmetries, Bäcklund transformations and exact solutions to the third-order Burgers’ equations. Journal of Statistical Physics, 158(2):433–446, January 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1130-8.

Liu:2018:TKS

[LJN18] Wenjian Liu, Sreenivasa Rao Jammalamadaka, and Ning Ning. The tightness of the Kesten–Stigum reconstruction bound of symmetric model with multiple mutations. Journal of Statistical Physics, 170(3):617–641, February 2018. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Linse:2011:DOM

[LK11] Per Linse and Gunnar Karlström. Dipolar order in molecular fluids: II. Molecular influence. Journal of Statistical Physics, 145(2):418–440, October 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s10955-011-
0352-2.pdf.

Lam:2014:SPI

[LK14] Khanh-Dang Nguyen Thu Lam and Jorge Kurchan. Stochastic perturbation of integrable systems: A window to weakly chaotic systems. Journal of Statistical Physics, 156(4):619– 646, August 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1030-y.

Leonel:2018:ICD

[LK18] Edson D. Leonel and Célia M. Kuwana. An investigation of chaotic diffusion in a family of Hamiltonian mappings whose angles diverge in the limit of vanishingly action. Journal of Statistical Physics, 170(1):69–78, January 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Loos:2019:FPE

[LK19] Sarah A. M. Loos and Sabine H. L. Klapp. Fokker–Planck equations for time–delayed systems via Markovian embedding. Journal of Statistical Physics, 177(1):95–118, October 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lin:2012:FFL

[LKD12] Yen Ting Lin, Hyejin Kim, and Charles R. Doering. Features of fast living: On the weak selection for longevity in degenerate birth–death processes. Journal of Statistical Physics, 148(4):647–663, September 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0479-9.

Lukas:2011:HMC

[LKR^{+}11] C. Lukas, P. Kollmannsberger, D. Ruffoni, P. Roschger, P. Fratzl, and R. Weinkamer. The heterogeneous mineral content of bone — using stochastic arguments and simulations to overcome experimental limitations. Journal of Statistical Physics, 144(2):316–331, July 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0209-8.

Liepelt:2010:ISC

[LL10] Steffen Liepelt and Reinhard Lipowsky. Impact of slip cycles on the operation modes and efficiency of molecular motors. Journal of Statistical Physics, 141(1):1–16, October 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-010-0050-5.pdf.

Lahbabi:2013:EDK

[LL13] Salma Lahbabi and Frédéric Legoll. Effective dynamics for a kinetic Monte–Carlo model with slow and fast time scales. Journal of Statistical Physics, 153(6):931–966, December 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0877-7.

Landim:2016:MTD

[LL16] C. Landim and P. Lemire. Metastability of the two– dimensional Blume–Capel model with zero chemical potential and small magnetic field. Journal of Statistical Physics, 164(2):346–376, July 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1550-8.

Lu:2019:ADI

[LL19] Yan Lü and Hong Lu. Anomalous dynamics of inertial systems driven by colored Lévy noise. Journal of Statistical Physics, 176(4):1046–1056, August 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lin:2010:STA

[LLJH10] Bo-Liang Lin, Jun-Wei Li, Li-Jun Ji, and Yong-Chang Huang. A statistical theory to aggregation in one-dimensional freeway traffic. Journal of Statistical Physics, 141(6):1104–1115, December 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-0083-9.

Li:2017:FSD

[LLL17] Lei Li, Jian-Guo Liu, and Jianfeng Lu. Fractional stochastic differential equations satisfying fluctuation–dissipation theorem. Journal of Statistical Physics, 169(2):316–339, October 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Leggio:2012:MCL

[LLM12] B. Leggio, O. Lychkovskiy, and A. Messina. On the merit of a central limit theorem-based approximation in statistical physics. Journal of Statistical Physics, 146(6):1274–1287, March 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0442-9.

Landim:2019:MTD

[LLM19] C. Landim, P. Lemire, and M. Mourragui. Metastability of the two-dimensional Blume–Capel model with zero chemical potential and small magnetic field on a large torus. Journal of Statistical Physics, 175(2):456–494, April 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lund:2013:ECS

[LLS13] Halvor Lund, Ludvig Lizana, and Ingve Simonsen. Effects of city-size heterogeneity on epidemic spreading in a metapopulation: A reaction–diffusion approach. Journal of Statistical Physics, 151(1–2):367–382, April 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0690-3.

Li:2017:VRC

[LLS17] Qin Li, Jianfeng Lu, and Weiran Sun. Validity and regularization of classical half-space equations. Journal of Statistical Physics, 166(2):398–433, January 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1688-4.

Lopes:2012:DTE

[LM12a] Artur O. Lopes and Jairo K. Mengue. Duality theorems in ergodic transport. Journal of Statistical Physics, 149(5): 921–942, November 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0626-3.

Lowe:2012:MDR

[LM12b] Matthias Löwe and Raphael Meiners. Moderate deviations for random field Curie–Weiss models. Journal of Statistical Physics, 149(4):701–721, November 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0611-x.

Li:2013:NSD

[LM13] Xingjie Helen Li and Govind Menon. Numerical solution of Dyson Brownian motion and a sampling scheme for invariant matrix ensembles. Journal of Statistical Physics, 153(5):

801–812, December 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0858-x.

Leonenko:2017:MCT

[LM17] Nikolai Leonenko and Anatoliy Malyarenko. Matérn class tensor-valued random fields and beyond. Journal of Statistical Physics, 168(6):1276–1301, September 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-017-1847-2.pdf.

Lin:2011:CTH

[LMC11] Lin Lin, Joseph A. Morrone, and Roberto Car. Correlated tunneling in hydrogen bonds. Journal of Statistical Physics, 145(2):365–384, October 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0320-x.

Lima:2019:FFX

[LMC19] A. B. Lima, L. A. S. Mól, and B. V. Costa. The fully frustrated XY model revisited: A new universality class. Journal of Statistical Physics, 175(5):960–971, June 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lebensztayn:2016:RWS

[LMM16] Elcio Lebensztayn, Fábio Prates Machado, and Mauricio Zuluaga Martinez. Random walks systems with finite lifetime on Z. Journal of Statistical Physics, 162(3):727–738, February 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1418-3.

Lukkarinen:2016:HCV

[LMN16] Jani Lukkarinen, Matteo Marcozzi, and Alessia Nota. Harmonic chain with velocity flips: Thermalization and kinetic theory. Journal of Statistical Physics, 165(5):809–844, December 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1647-0.

Lukkarinen:2018:SCC

[LMN18] Jani Lukkarinen, Matteo Marcozzi, and Alessia Nota. Summability of connected correlation functions of coupled lattice fields. Journal of Statistical Physics, 171(2):189–206, April 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Landim:2015:MRR

[LMT15] C. Landim, R. Misturini, and K. Tsunoda. Metastability of reversible random walks in potential fields. Journal of Statistical Physics, 160(6):1449–1482, September 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1298-6.

Lima:2011:REB

[LN11] Paulo C. Lima and Armando G. M. Neves. On the residual entropy of the BEG model at the antiquadrupolar–ferromagnetic coexistence line. Journal of Statistical Physics, 144(4):749– 758, August 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0291-y.

Lanchier:2013:SDH

[LN13] N. Lanchier and J. Neufer. Stochastic dynamics on hypergraphs and the spatial majority rule model. Journal of Statistical Physics, 151(1–2):21–45, April 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0543-5.

Lopes:2015:LDS

[LN15] Artur O. Lopes and Adriana Neumann. Large deviations for stationary probabilities of a family of continuous time Markov chains via Aubry–Mather theory. Journal of Statistical Physics, 159(4):797–822, May 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1205-1.

Liu:2019:IRI

[LN19a] Wenjian Liu and Ning Ning. Information reconstruction on an infinite tree for a 4 × 4-state asymmetric model with community effects. Journal of Statistical Physics, 177(3):438–467, November 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Liu:2019:LDA

[LN19b] Wenjian Liu and Ning Ning. Large degree asymptotics and the reconstruction threshold of the asymmetric binary channels. Journal of Statistical Physics, 174(6):1161–1188, March 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Leimkuhler:2011:CES

[LNP11] Ben Leimkuhler, Emad Noorizadeh, and Oliver Penrose. Comparing the efficiencies of stochastic isothermal molecular dynamics methods. Journal of Statistical Physics, 143(5):921– 942, June 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0210-2.

Latora:2013:SCS

[LNP13a] V. Latora, V. Nicosia, and P. Panzarasa. Social cohesion, structural holes, and a tale of two measures. Journal of Statistical Physics, 151(3–4):745–764, May 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0722-z.

Lelievre:2013:ONR

[LNP13b] T. Lelièvre, F. Nier, and G. A. Pavliotis. Optimal nonreversible linear drift for the convergence to equilibrium of a diffusion. Journal of Statistical Physics, 152(2):237–274, July 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0769-x.

Liu:2019:SSS

[LNP19] Jian-Guo Liu, B. Niethammer, and Robert L. Pego. Selfsimilar spreading in a merging–splitting model of animal group size. Journal of Statistical Physics, 175(6):1311–1330, June 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lancia:2012:EDC

[LNS12a] Carlo Lancia, Francesca R. Nardi, and Benedetto Scoppola. Entropy-driven cutoff phenomena. Journal of Statistical Physics, 149(1):108–141, October 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-012-0584-
9.

Li:2012:ECD

[LNS^{+}12b] P. Li, L. R. Nie, C. Z. Shu, S. Hu, and Q. Shao. Effect of correlated dichotomous noises on stochastic resonance in a linear system. Journal of Statistical Physics, 146(6):1184– 1202, March 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0427-8.

Lopes:2013:TFC

[LNT13] Artur Lopes, Adriana Neumann, and Philippe Thieullen. A thermodynamic formalism for continuous time Markov chains with values on the Bernoulli space: Entropy, pressure and large deviations. Journal of Statistical Physics, 152(5):894–933, September 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0796-7.

Li:2016:LTE

[LNY16] Yao Li, Péter Nándori, and Lai-Sang Young. Local thermal equilibrium for certain stochastic models of heat transport. Journal of Statistical Physics, 163(1):61–91, April 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1466-3.

Lindgren:2017:ATE

[LO17] Kristian Lindgren and Eckehard Olbrich. The approach towards equilibrium in a reversible Ising dynamics model: An information-theoretic analysis based on an exact solution. Journal of Statistical Physics, 168(4):919–935, August 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-017-1833-8.pdf.

Lelarge:2018:RBC

[LO18] Marc Lelarge and Mendes Oulamara. Replica bounds by combinatorial interpolation for diluted spin systems. Journal of Statistical Physics, 173(3–4):917–940, November 2018.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Los:2017:IIC

[Los17] Victor F. Los. Influence of initial correlations on evolution of a subsystem in a heat bath and polaron mobility. Journal of Statistical Physics, 168(4):857–872, August 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Levin:2016:MEP

[LP16] David A. Levin and Yuval Peres. Mixing of the exclusion process with small bias. Journal of Statistical Physics, 165 (6):1036–1050, December 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1664-z.

Latorre:2013:CER

[LPK13] J. C. Latorre, G. A. Pavliotis, and P. R. Kramer. Corrections to Einstein’s relation for Brownian motion in a tilted periodic potential. Journal of Statistical Physics, 150(4):776–803, February 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0692-1.

Landim:2012:MGD

[LPS12] C. Landim, R. D. Portugal, and B. F. Svaiter. A Markovian growth dynamics on rooted binary trees evolving according to the Gompertz curve. Journal of Statistical Physics, 148 (3):565–578, August 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0549-z.

Lansky:2016:GNY

[LPS16] Petr Lansky, Federico Polito, and Laura Sacerdote. Generalized nonlinear Yule models. Journal of Statistical Physics, 165(3):661–679, November 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1630-9.

Loeffen:2019:ETN

[LPS19] R. Loeffen, P. Patie, and M. Savov. Extinction time of non– Markovian self-similar processes, persistence, annihilation of jumps and the Fréchet distribution. Journal of Statistical Physics, 175(5):1022–1041, June 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lv:2012:FFP

[LQR12] Longjin Lv, Weiyuan Qiu, and Fuyao Ren. Fractional Fokker–Planck equation with space and time dependent drift and diffusion. Journal of Statistical Physics, 149(4):619– 628, November 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0618-3. See

comment [MGZ14].

Liu:2017:UTO

[LQY17] Qinghui Liu, Yanhui Qu, and Xiao Yao. Unbounded trace orbits of Thue–Morse Hamiltonian. Journal of Statistical Physics, 166(6):1509–1557, March 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-017-1726-
x.

Lundholm:2015:AFA

[LR15] Douglas Lundholm and Nicolas Rougerie. The average field approximation for almost bosonic extended anyons. Journal of Statistical Physics, 161(5):1236–1267, December 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1382-y.

Lanchier:2018:RRD

[LR18] Nicolas Lanchier and Stephanie Reed. Rigorous results for the distribution of money on connected graphs. Journal of Statistical Physics, 171(4):727–743, May 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lanchier:2019:RRD

[LR19] Nicolas Lanchier and Stephanie Reed. Rigorous results for the distribution of money on connected graphs (models with debts). Journal of Statistical Physics, 176(5):1115–1137, September 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lucarini:2017:PCC

[LRL17] Valerio Lucarini, Francesco Ragone, and Frank Lunkeit. Predicting climate change using response theory: Global averages and spatial patterns. Journal of Statistical Physics, 166 (3–4):1036–1064, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-016-1506-z.
pdf.

Lieb:2018:RLL

[LRY18] Elliott H. Lieb, Nicolas Rougerie, and Jakob Yngvason. Rigidity of the Laughlin liquid. Journal of Statistical Physics, 172(2):544–554, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-018-2082-1.
pdf.

Lieb:2010:SMQ

[LS10]  Elliott H. Lieb and Robert Seiringer.  The stability

of matter in quantum mechanics. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-19118-1 (hardcover). xv + 293 pp. LCCN QC173.4.T48 L543

2010.  URL http://assets.cambridge.org/97805211/
91180/cover/9780521191180.jpg.

Lieb:2012:FIB

[LS12] Elliott H. Lieb and Robert Seiringer. Further implications of the Bessis–Moussa–Villani conjecture. Journal of Statistical Physics, 149(1):86–91, October 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0585-8.

Lancia:2013:ENE

[LS13] Carlo Lancia and Benedetto Scoppola. Equilibrium and nonequilibrium Ising models by means of PCA. Journal of Statistical Physics, 153(4):641–653, November 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0847-0.

Li:2014:ARG

[LS14a] Dong Li and Yakov G. Sinai. An application of the renormalization group method to stable limit laws. Journal of Statistical Physics, 157(4–5):915–930, December 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1098-4.

Lieb:2014:ETD

[LS14b] Elliott H. Lieb and Robert Seiringer. Equivalence of two definitions of the effective mass of a polaron. Journal of Statistical Physics, 154(1–2):51–57, January 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0791-z.

LaCour:2015:GCL

[LS15a] Brian R. La Cour and William C. Schieve. A general conditional large deviation principle. Journal of Statistical Physics, 161(1):123–130, October 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-015-1328-4.
pdf.

Lardizabal:2015:CQC

[LS15b] Carlos F. Lardizabal and Rafael R. Souza. On a class of quantum channels, open random walks and recurrence. Journal of Statistical Physics, 159(4):772–796, May 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1217-x.

Landim:2016:MNR

[LS16a] C. Landim and I. Seo. Metastability of non-reversible, mean-field Potts model with three spins. Journal of Statistical Physics, 165(4):693–726, November 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1638-1.

Lardizabal:2016:OQR

[LS16b] Carlos F. Lardizabal and Rafael R. Souza. Open quantum random walks: Ergodicity, hitting times, Gambler’s ruin and potential theory. Journal of Statistical Physics, 164(5): 1122–1156, September 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1578-9.

Lima-Santos:2016:TLM

[LS16c] A. Lima-Santos. On the U_{q}[osp(1|2)] Temperley–Lieb model. Journal of Statistical Physics, 165(5):953–969, December 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1648-z.

Leppanen:2017:NFD

[LS17] Juho Leppänen and Mikko Stenlund. A note on the finitedimensional distributions of dispersing billiard processes. Journal of Statistical Physics, 168(1):128–145, July 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ludwig:2013:LSA

[LSBS13] Andreas W. W. Ludwig, Hermann Schulz-Baldes, and Michael Stolz. Lyapunov spectra for all ten symmetry classes of quasi-one-dimensional disordered systems of non-interacting fermions. Journal of Statistical Physics, 152(2):275–304, July 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0764-2.

Liu:2019:IPF

[LSS19] Jingcheng Liu, Alistair Sinclair, and Piyush Srivastava. The Ising partition function: Zeros and deterministic approximation. Journal of Statistical Physics, 174(2):287–315, January 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lohmann:2017:CTP

[LSW17] Martin Lohmann, Gordon Slade, and Benjamin C. Wallace. Critical two-point function for long-range O(n) models below the upper critical dimension. Journal of Statistical Physics, 169(6):1132–1161, December 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lega:2018:CTS

[LSY18] Joceline Lega, Sunder Sethuraman, and Alexander L. Young. On collisions times of ‘self-sorting’ interacting particles in one-dimension with random initial positions and velocities. Journal of Statistical Physics, 170(6):1088–1122, March 2018.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lemm:2019:AMH

[LSY19] Marius Lemm, Anders W. Sandvik, and Sibin Yang. The AKLT model on a hexagonal chain is gapped. Journal of Statistical Physics, 177(6):1077–1088, December 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lafuerza:2010:GAM

[LT10a] Luis F. Lafuerza and Raul Toral. On the Gaussian approximation for master equations. Journal of Statistical Physics, 140(5):917–933, September 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0024-7.

Linshiz:2010:CRE

[LT10b] Jasmine S. Linshiz and Edriss S. Titi. On the convergence rate of the Euler-α, an inviscid second-grade complex fluid, model to the Euler equations. Journal of Statistical Physics, 138(1–3):305–332, February 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-009-9916-9.pdf.

LaCour:2016:CSA

[LTM16] Brian R. La Cour, James E. Troupe, and Hans M. Mark. Classical simulated annealing using quantum analogues. Journal of Statistical Physics, 164(4):772–784, August 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-016-1570-4.pdf.

Lin:2017:WDD

[LTR17] Henry W. Lin, Max Tegmark, and David Rolnick. Why does deep and cheap learning work so well? Journal of Statistical Physics, 168(6):1223–1247, September 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lu:2012:BSS

[Lu12] Xuguang Lu. On backward solutions of the spatially homogeneous Boltzmann equation for Maxwelian molecules. Journal of Statistical Physics, 147(5):991–1006, June 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0501-2.

Lu:2013:BEB

[Lu13] Xuguang Lu. The Boltzmann equation for Bose–Einstein particles: Condensation in finite time. Journal of Statistical Physics, 150(6):1138–1176, March 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-013-0725-
9.

Lu:2014:BEB

[Lu14] Xuguang Lu. The Boltzmann equation for Bose–Einstein particles: Regularity and condensation. Journal of Statistical Physics, 156(3):493–545, August 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s10955-014-
1026-7.pdf.

Lu:2016:LTC

[Lu16] Xuguang Lu. Long time convergence of the Bose–Einstein condensation. Journal of Statistical Physics, 162(3):652–670, February 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1427-2.

Lebensztayn:2019:NUB

[LU19] Elcio Lebensztayn and Jaime Utria. A new upper bound for the critical probability of the frog model on homogeneous trees. Journal of Statistical Physics, 176(1):169–179, July 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lucarini:2012:SPD

[Luc12] Valerio Lucarini. Stochastic perturbations to dynamical systems: A response theory approach. Journal of Statistical Physics, 146(4):774–786, February 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-012-0422-

Lucarini:2016:ROM

[Luc16] Valerio Lucarini. Response operators for Markov processes in a finite state space: Radius of convergence and link to the response theory for axiom a systems. Journal of Statistical Physics, 162(2):312–333, January 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-015-1409-
4;   http://link.springer.com/content/pdf/10.1007/
s10955-015-1409-4.pdf.

Lucon:2017:QLD

[Luç17] Eric Luçon. Quenched large deviations for interacting diffusions in random media. Journal of Statistical Physics, 166 (6):1405–1440, March 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-017-1719-9.

Lucarini:2018:REL

[Luc18] Valerio Lucarini. Revising and extending the linear response theory for statistical mechanical systems: Evaluating observables as predictors and predictands. Journal of Statistical Physics, 173(6):1698–1721, December 2018. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-018-2151-5.pdf.

Lukkarinen:2014:THP

[Luk14] Jani Lukkarinen. Thermalization in harmonic particle chains with velocity flips. Journal of Statistical Physics, 155(6):1143– 1177, June 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-0930-1.

Lowe:2011:HMS

[LV11] Matthias Löwe and Franck Vermet. The Hopfield model on a sparse Erdős–Renyi graph. Journal of Statistical Physics, 143(1):205–214, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0167-1.

Lenells:2019:SFG

[LV19] Jonatan Lenells and Fredrik Viklund. Schramm’s formula and the Green’s function for multiple SLE. Journal of Statistical Physics, 176(4):873–931, August 2019. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-019-02325-0.pdf.

Lu:2019:MCA

[LVE19] Jianfeng Lu and Eric Vanden-Eijnden. Methodological and computational aspects of parallel tempering methods in the infinite swapping limit. Journal of Statistical Physics, 174(3): 715–733, February 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

LeClair:2018:IPB

[LW18] Joshua S. LeClair and Lindi M. Wahl. The impact of population bottlenecks on microbial adaptation. Journal of Statistical Physics, 172(1):114–125, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lim:2019:HCG

[LW19] Soon Hoe Lim and Jan Wehr. Homogenization for a class of generalized Langevin equations with an application to thermophoresis. Journal of Statistical Physics, 174(3):656–691, February 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Liang:2012:FFP

[LWL^{+}12] Jin-Rong Liang, Jun Wang, Long-Jin Lǔ, Hui Gu, Wei-Yuan Qiu, and Fu-Yao Ren. Fractional Fokker–Planck equation and Black–Scholes formula in composite–diffusive regime. Journal of Statistical Physics, 146(1):205–216, January 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0396-3.

Lim:2018:SML

[LWL^{+}18] Soon Hoe Lim, Jan Wehr, Aniello Lampo, Miguel Ángel García-March, and Maciej Lewenstein. On the small mass limit of quantum Brownian motion with inhomogeneous damping and diffusion. Journal of Statistical Physics, 170(2): 351–377, January 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lin:2018:QPE

[LWW18] Yu-Chu Lin, Haitao Wang, and Kung-Chien Wu. Quantitative pointwise estimate of the solution of the linearized Boltzmann equation. Journal of Statistical Physics, 171(5):927–964, June 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Liao:2018:GED

[LWY18] Jie Liao, Qianrong Wang, and Xiongfeng Yang. Global existence and decay rates of the solutions near Maxwellian for nonlinear Fokker–Planck equations. Journal of Statistical Physics, 173(1):222–241, October 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Li:2017:NSP

[LX17] Yao Li and Hui Xu. Numerical simulation of polynomial-speed convergence phenomenon. Journal of Statistical Physics, 169 (4):697–729, November 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Liao:2019:TPT

[LXHAA19] Yunhua Liao, Xiaoliang Xie, Yaoping Hou, and M. A. Aziz-Alaoui. Tutte polynomials of two self-similar network models. Journal of Statistical Physics, 174(4):893–905, February 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lin:2010:NSS

[LY10] Kevin K. Lin and Lai-Sang Young. Nonequilibrium steady states for certain Hamiltonian models. Journal of Statistical Physics, 139(4):630–657, May 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9958-z.

Li:2013:ENS

[LY13] Yao Li and Lai-Sang Young. Existence of nonequilibrium steady state for a simple model of heat conduction. Journal of Statistical Physics, 152(6):1170–1193, September 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0801-1.

Luo:2016:SAL

[LY16] Lan Luo and Hongjun Yu. Spectrum analysis of the linearized relativistic Landau equation. Journal of Statistical Physics, 163(4):914–935, May 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1501-4.

Li:2016:SLS

[LYT16] Shuli Li, Weigen Yan, and Tao Tian. The spectrum and Laplacian spectrum of the dice lattice. Journal of Statistical Physics, 164(2):449–462, July 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1552-6.

Lagro:2017:PFT

[LYX17] Matthew Lagro, Wei-Shih Yang, and Sheng Xiong. A Perron– Frobenius type of theorem for quantum operations. Journal of Statistical Physics, 169(1):38–62, October 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Loutsenko:2011:CML

[LYZ11] Igor Loutsenko, Oksana Yermolayeva, and Michel Zinsmeister. On a competitive model of Laplacian growth. Journal of Statistical Physics, 145(4):919–931, November 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0278-8.

Lefevere:2010:HST

[LZ10a] Raphaël Lefevere and Lorenzo Zambotti. Hot scatterers and tracers for the transfer of heat in collisional dynamics. Journal of Statistical Physics, 139(4):686–713, May 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-9962-3.

Liu:2010:SUP

[LZ10b] Dang-Zheng Liu and Da-Sheng Zhou. Some universal properties for restricted trace Gaussian orthogonal, unitary and symplectic ensembles. Journal of Statistical Physics, 140(2):268– 288, July 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-9993-9.

Lu:2011:BEB

[LZ11] Xuguang Lu and Xiangdong Zhang. On the Boltzmann equation for 2D Bose–Einstein particles. Journal of Statistical Physics, 143(5):990–1019, June 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0221-z.

Lam:2015:DST

[LZ15] Pui-Man Lam and Yi Zhen. Dynamic scaling theory of the forced translocation of a semi-flexible polymer through a nanopore. Journal of Statistical Physics, 161(1):197–209, October 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1322-x.

Liu:2019:VBT

[LZHS19] Jia-Bao Liu, Jing Zhao, Hailang He, and Zehui Shao. Valencybased topological descriptors and structural property of the generalized Sierpiński networks. Journal of Statistical Physics, 177(6):1131–1147, December 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Macdonald:2010:DCC

[Mac10] Brian Macdonald. Density of complex critical points of a real random SO(m+1) polynomial. Journal of Statistical Physics, 141(3):517–531, November 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0057-y.

Machta:2013:DSC

[Mac13] Jonathan Machta. Daniel Stein and Charles Newman: Spin Glasses and Complexity. Journal of Statistical Physics, 152 (5):995–996, September 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0802-0.

Maes:2014:SFD

[Mae14] Christian Maes. On the second fluctuation–dissipation theorem for nonequilibrium baths. Journal of Statistical Physics, 154(3):705–722, February 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0904-8.

Maldonado:2012:FBC

[Mal12] Cesar Maldonado. Fluctuation bounds for chaos plus noise in dynamical systems. Journal of Statistical Physics, 148(3):548– 564, August 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0553-3.

Mansour:2011:TWF

[Man11] M. B. A. Mansour. On traveling wave fronts in a bacterial growth model with density-dependent diffusion and chemotaxis. Journal of Statistical Physics, 143(1):197–204, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0159-1.

Mulansky:2011:SWC

[MAPS11] M. Mulansky, K. Ahnert, A. Pikovsky, and D. L. Shepelyansky. Strong and weak chaos in weakly nonintegrable manybody Hamiltonian systems. Journal of Statistical Physics, 145(5):1256–1274, December 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0335-3.

Marko:2011:SLW

[Mar11a] John F. Marko. Scaling of linking and writhing numbers for spherically confined and topologically equilibrated flexible polymers. Journal of Statistical Physics, 142(6):1353– 1370, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0172-4.

Martinez:2011:HPF

[Mar11b] M. A. Martínez. Hagen–Poiseuille flow solutions in gradtype equations. Journal of Statistical Physics, 142(4):710–725, February 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0142-x.

Maric:2015:FVP

[Mar15] Nevena Marić. Fleming–Viot particle system driven by a random walk on N. Journal of Statistical Physics, 160(3):548– 560, August 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1275-0.

Marshall:2016:ASE

[Mar16] J. S. Marshall. Analytical solutions for an escape problem in a disc with an arbitrary distribution of exit holes along its boundary. Journal of Statistical Physics, 165(5):920–952, December 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1653-2.

Margarint:2018:CHP

[Mar18] Vlad Margarint. Convergence in high probability of the quantum diffusion in a random band matrix model. Journal of Statistical Physics, 172(3):781–794, August 2018. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-018-2065-2.pdf.

Mastromatteo:2013:BII

[Mas13] Iacopo Mastromatteo. Beyond inverse Ising model: Structure of the analytical solution. Journal of Statistical Physics, 150(4):658–670, February 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0707-y.

Mastropietro:2014:WSP

[Mas14] Vieri Mastropietro. Weyl semimetallic phase in an interacting lattice system. Journal of Statistical Physics, 157(4–5): 830–854, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1020-0.

Mascali:2016:NFT

[Mas16] Giovanni Mascali. A new formula for thermal conductivity based on a hierarchy of hydrodynamical models. Journal of Statistical Physics, 163(5):1268–1284, June 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1509-9.

Matsuoka:2012:GKF

[Mat12] Hiroshi Matsuoka. Green–Kubo formulas with symmetrized correlation functions for quantum systems in steady states:

The shear viscosity of a fluid in a steady shear flow. Journal of Statistical Physics, 148(5):933–950, September 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0556-0.

Matsui:2015:MSA

[Mat15] Chihiro Matsui. Multi-state asymmetric simple exclusion processes. Journal of Statistical Physics, 158(1):158–191, January 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1121-9.

Mays:2013:RQS

[May13] Anthony Mays. A real quaternion spherical ensemble of random matrices. Journal of Statistical Physics, 153(1):48–69, October 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0808-7.

Mora:2011:BSP

[MB11] Thierry Mora and William Bialek. Are biological systems poised at criticality? Journal of Statistical Physics, 144 (2):268–302, July 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0229-4.

Mulder:2018:NGC

[MB18] Daan Mulder and Ginestra Bianconi. Network geometry and complexity. Journal of Statistical Physics, 173(3–4):783–805, November 2018. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-018-2115-9.pdf.

Musmeci:2013:BTP

[MBC^{+}13] Nicolò Musmeci, Stefano Battiston, Guido Caldarelli, Michelangelo Puliga, and Andrea Gabrielli. Bootstrapping topological properties and systemic risk of complex networks using the fitness model. Journal of Statistical Physics, 151(3–4):720– 734, May 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0720-1.

Maiocchi:2014:ATF

[MBC14] A. Maiocchi, D. Bambusi, and A. Carati. An averaging theorem for FPU in the thermodynamic limit. Journal of Statistical Physics, 155(2):300–322, April 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0958-2.

Marathe:2012:DSD

[MBGK12] Rahul Marathe, Veronika Bierbaum, David Gomez, and Stefan Klumpp. Deterministic and stochastic descriptions of gene expression dynamics. Journal of Statistical Physics, 148(4):608–627, September 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0459-0.

Mukhamedov:2016:PTQ

[MBS16] Farrukh Mukhamedov, Abdessatar Barhoumi, and Abdessatar Souissi. Phase transitions for quantum Markov chains associated with Ising type models on a Cayley tree. Journal of Statistical Physics, 163(3):544–567, May 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1495-y.

Mishra:2016:SAD

[MBWC16] Shradha Mishra, Sanchari Bhattacharya, Benjamin Webb, and E. G. D. Cohen. Subdiffusion, anomalous diffusion and propagation of a particle moving in random and periodic media. Journal of Statistical Physics, 162(4):855–868, February 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1448-5.

Mounaix:2010:WLD

[MC10] Philippe Mounaix and Pierre Collet. Wave localization does not affect the breakdown of a Schrödinger-Type amplifier driven by the square of a Gaussian field. Journal of Statistical Physics, 139(3):466–477, May 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9946-3.

Mounaix:2011:LAB

[MC11] Philippe Mounaix and Pierre Collet. Linear amplifier breakdown and concentration properties of a Gaussian field given that its L^{2}-norm is large. Journal of Statistical Physics, 143(1):139–147, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0165-3.

Marzen:2016:PRD

[MC16] Sarah E. Marzen and James P. Crutchfield. Predictive rate-distortion for infinite-order Markov processes. Journal of Statistical Physics, 163(6):1312–1338, June 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1520-1. See correction [MC21].

Marzen:2017:ICA

[MC17a] Sarah Marzen and James P. Crutchfield. Informational and causal architecture of continuous-time renewal processes. Journal of Statistical Physics, 168(1):109–127, July 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Marzen:2017:SRC

[MC17b] Sarah E. Marzen and James P. Crutchfield. Structure and randomness of continuous-time, discrete-event processes. Journal of Statistical Physics, 169(2):303–315, October 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Mei:2019:MPS

[MC19] Zhongtao Mei and Jaeyoon Cho. Matrix product solution of the stationary state of two-species open zero range processes. Journal of Statistical Physics, 175(1):150–160, April 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Marzen:2021:CPR

[MC21] Sarah E. Marzen and James P. Crutchfield. Correction to: Predictive Rate-Distortion for Infinite-Order Markov Processes. Journal of Statistical Physics, 182(2):??, February 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/
pdf/10.1007/s10955-021-02698-1.pdf. See [MC16].

Maiocchi:2012:SET

[MCG12] Alberto Mario Maiocchi, Andrea Carati, and Antonio Giorgilli. A series expansion for the time autocorrelation of dynamical variables. Journal of Statistical Physics, 148 (6):1054–1071, September 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0575-x.

Moqaddam:2015:SDC

[MCK15] A. Mazloomi Moqaddam, S. S. Chikatamarla, and I. V. Karlin. Simulation of droplets collisions using two-phase entropic lattice Boltzmann method. Journal of Statistical Physics, 161(6):1420–1433, December 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-015-1329-
3;   http://link.springer.com/content/pdf/10.1007/
s10955-015-1329-3.pdf.

Majumdar:2010:ZTP

[MD10] Kingshuk Majumdar and Trinanjan Datta. Zero temperature phases of the frustrated J_{1}–J_{2} antiferromagnetic spin-1/2 Heisenberg model on a simple cubic lattice. Journal of Statistical Physics, 139(4):714–726, May 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-9967-y.

Maes:2013:LTA

[MdG13] Christian Maes and Winny O’Kelly de Galway. A low temperature analysis of the boundary driven Kawasaki process. Journal of Statistical Physics, 153(6):991–1007, December 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0878-6.

Mihelich:2018:MKS

[MDP^{+}18] M. Mihelich, B. Dubrulle, D. Paillard, Q. Kral, and D. Faranda. Maximum Kolmogorov–Sinai entropy versus minimum mixing time in Markov chains. Journal of Statistical Physics, 170(1):62–68, January 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Maynar:2018:EEC

[MdSB18] P. Maynar, M. I. García de Soria, and J. Javier Brey. The Enskog equation for confined elastic hard spheres. Journal of Statistical Physics, 170(5):999–1018, March 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Machta:2011:MCM

[ME11] J. Machta and R. S. Ellis. Monte Carlo methods for rough free energy landscapes: Population annealing and parallel tempering. Journal of Statistical Physics, 144(3):541–553, August 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0249-0.

Melendez:2014:GJE

[ME14] M. Meléndez and P. Español. Gibbs–Jaynes entropy versus relative entropy. Journal of Statistical Physics, 155(1):93– 105, April 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-0954-6.

Mermin:2014:EMK

[Mer14] N. David Mermin. Early memories of Ken. Journal of Statistical Physics, 157(4–5):625–627, December 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1025-8.

Mermin:2019:MAP

[Mer19] N. David Mermin. Memories of and about Pierre. Journal of Statistical Physics, 175(3–4):515–517, May 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Montessori:2015:TDL

[MFLA15] A. Montessori, G. Falcucci, M. La Rocca, and S. Ansumali. Three-dimensional lattice pseudo-potentials for multiphase flow simulations at high density ratios. Journal of Statistical Physics, 161(6):1404–1419, December 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s10955-015-1318-6;   http://link.springer.com/
content/pdf/10.1007/s10955-015-1318-6.pdf.

Midha:2017:RIC

[MG17] Tripti Midha and Arvind Kumar Gupta. Role of interactions and correlations on collective dynamics of molecular motors along parallel filaments. Journal of Statistical Physics, 169 (4):824–845, November 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Metivier:2019:BTD

[MG19] David Métivier and Shamik Gupta. Bifurcations in the timedelayed Kuramoto model of coupled oscillators: Exact results. Journal of Statistical Physics, 176(2):279–298, July 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Martinez-Gonzalez:2013:PBB

[MGAPQH13] J. A. Martínez-González, J. C. Armas-Pérez, and J. Quintana-H. Phase behavior of bow-shaped hard particles in two dimensions. Journal of Statistical Physics, 150(3):559–571, February 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0606-7.

Muratore-Ginanneschi:2013:HRC

[MGMMP13] Paolo Muratore-Ginanneschi, Carlos Mejía-Monasterio, and Luca Peliti. Heat release by controlled continuous-time Markov jump processes. Journal of Statistical Physics, 150 (1):181–203, January 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0676-6.

Magdziarz:2014:CFF

[MGZ14] Marcin Magdziarz, Janusz Gajda, and Tomasz Zorawik. Comment on Fractional Fokker–Planck Equation with Space and Time Dependent Drift and Diffusion. Journal of Statistical Physics, 154(5):1241–1250, March 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-014-0919-9.pdf. See [LQR12].

Mulken:2017:IDS

[MHD17] Oliver Mülken, Sarah Heinzelmann, and Maxim Dolgushev. Information dimension of stochastic processes on networks: Relating entropy production to spectral properties. Journal of Statistical Physics, 167(5):1233–1243, June 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Miao:2011:ESL

[Mia11] Dong Miao. Eigenvalue statistics for lattice Hamiltonian with off-diagonal disorder. Journal of Statistical Physics, 143(3):509–522, May 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0190-2.

Mielke:2018:PFH

[Mie18] Andreas Mielke. Pair formation of hard core bosons in flat band systems. Journal of Statistical Physics, 171(4):679–695, May 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Mihailescu:2010:PMM

[Mih10] Eugen Mihailescu. Physical measures for multivalued inverse iterates near hyperbolic repellors. Journal of Statistical Physics, 139(5):800–819, June 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9960-5.

Mihailescu:2011:LGD

[Mih11] Eugen Mihailescu. Local geometry and dynamical behavior on folded basic sets. Journal of Statistical Physics, 142(1):154– 167, January 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-0097-3.

Mihailescu:2019:HLE

[Mih19] Eugen Mihailescu. Hyperbolic lifts and estimates for overlap numbers. Journal of Statistical Physics, 177(3):468–484, November 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Misumi:2015:DLR

[Mis15] Jun Misumi. The diameter of a long-range percolation cluster on pre-Sierpinski gasket. Journal of Statistical Physics, 158(5):1083–1089, March 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1170-0.

Misumi:2016:MTR

[Mis16] Jun Misumi. The mixing time of a random walk on a longrange percolation cluster in pre-Sierpinski gasket. Journal of Statistical Physics, 165(1):153–163, October 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1611-z.

Minami:2018:MEM

[MiS18] Yuki Minami and Shin ichi Sasa. The most effective model for describing the universal behavior of a noisy Kuramoto– Sivashinsky equation as a paradigmatic model. Journal of Statistical Physics, 173(1):120–139, October 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Misumi:2019:DLR

[Mis19] Jun Misumi. The diameter of a long-range percolation cluster on generalized pre-Sierpinski carpet and regular tree. Journal of Statistical Physics, 174(2):276–286, January 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Mitter:2016:FRD

[Mit16] P. K. Mitter. On a finite range decomposition of the resolvent of a fractional power of the Laplacian. Journal of Statistical Physics, 163(5):1235–1246, June 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1507-y. See erratum [Mit17a].

Mitter:2017:EFR

[Mit17a] P. K. Mitter. Erratum to: On a Finite Range Decomposition of the Resolvent of a Fractional Power of the Laplacian. Journal of Statistical Physics, 166(2):453–455, January 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-016-1687-5.pdf. See [Mit16].

Mitter:2017:FRD

[Mit17b] P. K. Mitter. On a finite range decomposition of the resolvent of a fractional power of the Laplacian II. The torus. Journal of Statistical Physics, 168(5):986–999, September 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Miyao:2012:GSP

[Miy12a] Tadahiro Miyao. Ground state properties of the SSH model. Journal of Statistical Physics, 149(3):519–550, November 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0598-3.

Miyao:2012:NOD

[Miy12b] Tadahiro Miyao. Note on the one-dimensional Holstein– Hubbard model. Journal of Statistical Physics, 147(2):436– 447, April 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0466-1.

Miyao:2013:MPE

[Miy13] Tadahiro Miyao. Monotonicity of the polaron energy II: General theory of operator monotonicity. Journal of Statistical Physics, 153(1):70–92, October 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0812-y.

Miyao:2016:LRC

[Miy16a] Tadahiro Miyao. Long-range charge order in the extended Holstein–Hubbard model. Journal of Statistical Physics, 165 (2):225–245, October 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1617-6.

Miyao:2016:QGI

[Miy16b] Tadahiro Miyao. Quantum Griffiths inequalities. Journal of Statistical Physics, 164(2):255–303, July 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1546-4.

Miyao:2019:SFM

[Miy19] Tadahiro Miyao. Stability of ferromagnetism in many-electron systems. Journal of Statistical Physics, 176(5):1211–1271, September 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Mayo:2019:LCL

[MK19] Jackson R. Mayo and Alan R. Kerstein. Log-correlated largedeviation statistics governing Huygens fronts in turbulence. Journal of Statistical Physics, 176(2):456–477, July 2019. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Mao:2015:MDS

[ML15] Qian Mao and K. H. Luo. Molecular dynamics simulation of sintering dynamics of many TiO_{2} nanoparticles. Journal of Statistical Physics, 160(6):1696–1708, September 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1304-z.

Moretti:2013:MFA

[MLCPS13] Paolo Moretti, Suyu Liu, Claudio Castellano, and Romualdo Pastor-Satorras. Mean-field analysis of the q-voter model on networks. Journal of Statistical Physics, 151(1–2):113– 130, April 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0704-1.

Mehta:2016:LAS

[MLS16] Pankaj Mehta, Alex H. Lang, and David J. Schwab. Landauer in the age of synthetic biology: Energy consumption and information processing in biochemical networks. Journal of Statistical Physics, 162(5):1153–1166, March 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1431-6.

McDonald:2012:PMP

[MM12] Leslie M. McDonald and Iain Moffatt. On the Potts model partition function in an external field. Journal of Statistical Physics, 146(6):1288–1302, March 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0449-2.

Mera:2013:ECM

[MM13] Maria Eugenia Mera and Manuel Morán. Error covariance matrix estimation of noisy and dynamically coupled time series. Journal of Statistical Physics, 150(2):375–397, January 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0683-7.

Meidinger:2014:DLS

[MM14] David Meidinger and Vladimir Mitev. Dynamic lattice supersymmetry in }↕ (n|m) spin chains. Journal of Statistical Physics, 156(6):1199–1220, September 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1049-0.

Mathiaud:2016:FPM

[MM16] J. Mathiaud and L. Mieussens. A Fokker–Planck model of the Boltzmann equation with correct Prandtl number. Journal of Statistical Physics, 162(2):397–414, January 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1404-9.

Mathiaud:2017:FPM

[MM17a] J. Mathiaud and L. Mieussens. A Fokker–Planck model of the Boltzmann equation with correct Prandtl number for polyatomic gases. Journal of Statistical Physics, 168(5):1031–1055, September 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Mittnenzweig:2017:EGS

[MM17b] Markus Mittnenzweig and Alexander Mielke. An entropic gradient structure for Lindblad equations and couplings of quantum systems to macroscopic models. Journal of Statistical Physics, 167(2):205–233, April 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-017-1756-4.

Mellet:2015:AET

[MMA15] Antoine Mellet and Sara Merino-Aceituno. Anomalous energy transport in FPU-β chain. Journal of Statistical Physics, 160 (3):583–621, August 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1273-2.

Mazzone:2015:SMI

[MMM15] Valerio Mazzone, Simone Melchionna, and Umberto Marini Bettolo Marconi. Steric modulation of ionic currents in DNA translocation through nanopores. Journal of Statistical Physics, 158(5):1181–1194, March 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-014-1159-
8.

Moore:2018:REP

[MMR18] Sam Moore, Peter Mörters, and Tim Rogers. A re-entrant phase transition in the survival of secondary infections on networks. Journal of Statistical Physics, 171(6):1122–1135, June 2018. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-018-2050-9.pdf.

Martis:2013:EGS

[MMST13] Stephen Martis, Étienne Marcotte, Frank H. Stillinger, and Salvatore Torquato. Exotic ground states of directional pair potentials via collective-density variables. Journal of Statistical Physics, 150(3):414–431, February 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0619-2.

Martin-Mayor:2011:TMC

[MMSY11] V. Martin-Mayor, B. Seoane, and D. Yllanes. Tethered Monte Carlo: Managing rugged free-energy landscapes with a Helmholtz-potential formalism. Journal of Statistical Physics, 144(3):554–596, August 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0261-4.

Mancini:2016:TOD

[MMW16] F. Mancini, M. Marsili, and A. M. Walczak. Tradeoffs in delayed information transmission in biochemical networks. Journal of Statistical Physics, 162(5):1088–1129, March 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1332-8.

Matic:2012:SVB

[MN12] Ivan Matic and James Nolen. A sublinear variance bound for solutions of a random Hamilton–Jacobi equation. Journal of Statistical Physics, 149(2):342–361, October 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0590-y.

Maes:2014:NEC

[MN14a] Christian Maes and Karel Netocný. A nonequilibrium extension of the Clausius heat theorem. Journal of Statistical Physics, 154(1–2):188–203, January 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0822-9.

Menz:2014:DCL

[MN14b] Georg Menz and Robin Nittka. Decay of correlations in 1D lattice systems of continuous spins and long-range interaction. Journal of Statistical Physics, 156(2):239–267, July 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1011-1.

Maes:2015:RGP

[MN15] Christian Maes and Karel Netocný. Revisiting the Glansdorff– Prigogine criterion for stability within irreversible thermodynamics. Journal of Statistical Physics, 159(6):1286–1299, June 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1239-4.

Marcozzi:2016:DLL

[MN16a] M. Marcozzi and A. Nota. Derivation of the linear Landau equation and linear Boltzmann equation from the Lorentz model with magnetic field. Journal of Statistical Physics, 162(6):1539–1565, March 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1453-8.

Merchan:2016:SPI

[MN16b] Lina Merchan and Ilya Nemenman. On the sufficiency of pairwise interactions in maximum entropy models of networks. Journal of Statistical Physics, 162(5):1294–1308, March 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1456-5.

McHardy:2019:IWT

[MNBC19] Isaias McHardy, Marco Nizama, Adrian A. Budini, and Manuel O. Cáceres. Intermittent waiting-time noises through embedding processes. Journal of Statistical Physics, 177(4): 608–625, November 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Mayer:2012:DMF

[MNS12] Jörg Mayer, Hong-Viet V. Ngo, and Heinz Georg Schuster. Dynamical mean-field equations for a neural network with spike timing dependent plasticity. Journal of Statistical Physics, 148(4):677–686, September 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0486-x.

McLeod:2011:ASS

[MNV11] J. B. McLeod, B. Niethammer, and J. J. L. Velázquez. Asymptotics of self-similar solutions to coagulation equations with product kernel. Journal of Statistical Physics, 144(1):76– 100, July 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0239-2.

Mobilia:2013:CVP

[Mob13] Mauro Mobilia. Commitment versus persuasion in the threeparty constrained voter model. Journal of Statistical Physics, 151(1–2):69–91, April 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0656-x.

Mohamed:2011:EAG

[Moh11] Nader M. A. Mohamed. Efficient algorithm for generating Maxwell random variables. Journal of Statistical Physics, 145(6):1653–1660, December 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0364-y.

Mohazzabi:2017:BCD

[Moh17] Pirooz Mohazzabi. Binary collision density in a non-ideal gas as a function of particle density, collision diameter, and temperature. Journal of Statistical Physics, 169(2):362–373, October 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Moll:2014:DRN

[Mol14] Markus Moll. Diffraction of random noble means words. Journal of Statistical Physics, 156(6):1221–1236, September 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1047-2.

Molchan:2017:IBE

[Mol17] G. Molchan. The inviscid Burgers equation with fractional Brownian initial data: The dimension of regular Lagrangian points. Journal of Statistical Physics, 167(6):1546–1554, June 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Molchan:2018:PEG

[Mol18] G. Molchan. Persistence exponents for Gaussian random fields of fractional Brownian motion type. Journal of Statistical Physics, 173(6):1587–1597, December 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Monasson:2012:LSA

[Mon12a] Remi Monasson. Lorenzo Saitta, Attilio Giordana, Antoine Cornuéjols: Phase Transitions in Machine Learning. Journal of Statistical Physics, 149(6):1161–1163, December 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0655-y.

Monnai:2012:MRN

[Mon12b] Takaaki Monnai. Microscopic reversibility for nonequilibrium classical open systems. Journal of Statistical Physics, 149 (6):1058–1068, December 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0643-2.

Montanari:2015:FOC

[Mon15] Andrea Montanari. Finding one community in a sparse graph. Journal of Statistical Physics, 161(2):273–299, October 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1338-2.

Morawetz:2011:ABS

[Mor11] Klaus Morawetz. Asymmetric Bethe–Salpeter equation for pairing and condensation. Journal of Statistical Physics, 143(3):482–500, May 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0186-y.

Mori:2012:MAE

[Mor12] Takashi Mori. Microcanonical analysis of exactness of the mean-field theory in long-range interacting systems. Journal of Statistical Physics, 147(5):1020–1040, June 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0511-0.

Mori:2015:QEN

[Mor15] Takashi Mori. Quasi-equilibrium nonadditivity. Journal of Statistical Physics, 159(1):172–194, April 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1177-6.

Moriya:2018:EBL

[Mor18] Hajime Moriya. Ergodicity breaking and localization of the Nicolai supersymmetric fermion lattice model. Journal of Statistical Physics, 172(5):1270–1290, September 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Markiv:2014:CDK

[MOT14a] B. Markiv, I. Omelyan, and M. Tokarchuk. Consistent description of kinetics and hydrodynamics of weakly nonequilibrium processes in simple liquids. Journal of Statistical Physics, 155(5):843–866, June 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-0980-4.

Mottram:2014:PCF

[Mot14b] Edward Mottram. Percolation with constant freezing. Journal of Statistical Physics, 155(5):932–965, June 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0985-z.

Mounaix:2015:QDP

[Mou15] Philippe Mounaix. Quasi-deterministic properties of random Gaussian fields constrained by a large quadratic form. Journal of Statistical Physics, 160(3):561–582, August 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1278-x.

Movassagh:2016:EA

[Mov16] Ramis Movassagh. Eigenvalue attraction. Journal of Statistical Physics, 162(3):615–643, February 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1424-5.

Magdziarz:2011:OPS

[MOW11] Marcin Magdziarz, Sebastian Orzel, and Aleksander Weron. Option pricing in subdiffusive Bachelier model. Journal of Statistical Physics, 145(1):187–203, October 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-011-0310-z.pdf.

Mehats:2010:IPQ

[MP10] Florian Méhats and Olivier Pinaud. An inverse problem in quantum statistical physics. Journal of Statistical Physics, 140(3):565–602, August 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0003-z.

Morais:2013:CPC

[MP13] Thiago Morais and Aldo Procacci. Continuous particles in the canonical ensemble as an abstract polymer gas. Journal of Statistical Physics, 151(5):830–849, June 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0731-y.

Monaco:2014:TIE

[MP14] Domenico Monaco and Gianluca Panati. Topological invariants of eigenvalue intersections and decrease of Wannier functions in graphene. Journal of Statistical Physics, 155(6):1027– 1071, June 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-0918-x.

Mallick:2018:BBL

[MP18a] Kirone Mallick and Sylvain Prolhac. Brownian bridges for late time asymptotics of KPZ fluctuations in finite volume. Journal of Statistical Physics, 173(2):322–361, October 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Mastropietro:2018:CDW

[MP18b] Vieri Mastropietro and Marcello Porta. Canonical Drude weight for non-integrable quantum spin chains. Journal of Statistical Physics, 172(2):379–397, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Manacorda:2016:LMG

[MPL^{+}16] Alessandro Manacorda, Carlos A. Plata, Antonio Lasanta, Andrea Puglisi, and Antonio Prados. Lattice models for granular-like velocity fields: Hydrodynamic description. Journal of Statistical Physics, 164(4):810–841, August 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1575-z.

Moirogiannis:2017:RCM

[MPM17] Dimitrios Moirogiannis, Oreste Piro, and Marcelo O. Magnasco. Renormalization of collective modes in large-scale neural dynamics. Journal of Statistical Physics, 167(3–4):543– 558, May 2017. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-017-1753-7.pdf.

Mora:2016:JSI

[MPR16] Thierry Mora, Luca Peliti, and Olivier Rivoire. JSP special issue on information processing in living systems. Journal of Statistical Physics, 162(5):1085–1087, March 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-016-1459-2.pdf.

Morais:2014:LJT

[MPS14] Thiago Morais, Aldo Procacci, and Benedetto Scoppola. On Lennard-Jones type potentials and hard-core potentials with an attractive tail. Journal of Statistical Physics, 157(1):17– 39, October 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1067-y.

Mainini:2019:LCL

[MPSS19] Edoardo Mainini, Paolo Piovano, Bernd Schmidt, and Ulisse Stefanelli. N^{3/4} law in the cubic lattice. Journal of Statistical Physics, 176(6):1480–1499, September 2019. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-019-02350-z.pdf.

Morales:2019:FSR

[MPT19] Javier Morales, Jan Peszek, and Eitan Tadmor. Flocking with short-range interactions. Journal of Statistical Physics, 176 (2):382–397, July 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Masi:2011:CRS

[MPTV11] A. De Masi, E. Presutti, D. Tsagkarogiannis, and M. E. Vares. Current reservoirs in the simple exclusion process. Journal of Statistical Physics, 144(6):1151–1170, September 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0326-4.

Masi:2012:NES

[MPTV12] Anna De Masi, Errico Presutti, Dimitrios Tsagkarogiannis, and Maria Eulalia Vares. Non-equilibrium stationary states in the symmetric simple exclusion with births and deaths. Journal of Statistical Physics, 147(3):519–528, May 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0481-2.

Mischler:2018:WSC

[MQW18] S. Mischler, C. Quiñinao, and Q. Weng. Weak and strong connectivity regimes for a general time elapsed neuron network model. Journal of Statistical Physics, 173(1):77–98, October 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Mucha:2012:NMS

[MR12] Piotr Bogusław Mucha and Piotr Rybka. A note on a model system with sudden directional diffusion. Journal of Statistical Physics, 146(5):975–988, March 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0446-5.

Modarresi:2013:NSA

[MR13] N. Modarresi and S. Rezakhah. A new structure for analyzing discrete scale invariant processes: Covariance and spectra. Journal of Statistical Physics, 153(1):162–176, October 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0799-4.

Marino:2019:EPR

[MR19] Marcos Mariño and Tomás Reis. Exact perturbative results for the Lieb–Liniger and Gaudin–Yang models. Journal of Statistical Physics, 177(6):1148–1156, December 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Machado:2018:CCH

[MRCJ18] Fábio P. Machado, Alejandro Roldán-Correa, and Valdivino V. Junior. Colonization and collapse on homogeneous trees. Journal of Statistical Physics, 173(5):1386–1407, December 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Mansour:2019:BDP

[MRR19] Toufik Mansour, Reza Rastegar, and Alexander Roitershtein. On ballistic deposition process on a strip. Journal of Statistical Physics, 177(4):626–650, November 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Menon:2010:KTL

[MS10] Govind Menon and Ravi Srinivasan. Kinetic theory and Lax equations for shock clustering and Burgers turbulence. Journal of Statistical Physics, 140(6):1195–1223, September 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-0028-3.

Maes:2011:RSD

[MS11a] Christian Maes and Senya Shlosman. Rotating states in driven clock- and XY-models. Journal of Statistical Physics, 144 (6):1238–1246, September 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0325-5.

Muscato:2011:ETM

[MS11b] O. Muscato and V. Di Stefano. An energy transport model describing heat generation and conduction in silicon semiconductors. Journal of Statistical Physics, 144(1):171–197, July 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0247-2.

Meerschaert:2012:FDM

[MS12a] Mark M. Meerschaert and Peter Straka. Fractional dynamics at multiple times. Journal of Statistical Physics, 149(5): 878–886, November 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0638-z.

Meester:2012:RSO

[MS12b] Ronald Meester and Anish Sarkar. Rigorous self-organised criticality in the modified Bak–Sneppen model. Journal of Statistical Physics, 149(5):964–968, November 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0627-2.

Mountford:2012:DPH

[MS12c] Thomas S. Mountford and Aidan Sudbury. Deposition processes with hardcore behaviour. Journal of Statistical Physics, 146(4):687–700, February 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0415-4.

Marklof:2014:PLD

[MS14] Jens Marklof and Andreas Strömbergsson. Power-law distributions for the free path length in Lorentz gases. Journal of Statistical Physics, 155(6):1072–1086, June 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0935-9.

Morita:2016:PTT

[MS16] Satoshi Morita and Sei Suzuki. Phase transition of twodimensional Ising models on the honeycomb and related lattices with striped random impurities. Journal of Statistical Physics, 162(1):123–138, January 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-015-1400-
0.

Mendl:2017:SRW

[MS17] Christian B. Mendl and Herbert Spohn. Shocks, rarefaction waves, and current fluctuations for anharmonic chains. Journal of Statistical Physics, 166(3–4):841–875, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1626-5.

Maestrini:2019:ENT

[MS19a] Davide Maestrini and Hayder Salman. Entropy of negative temperature states for a point vortex gas. Journal of Statistical Physics, 176(4):981–1008, August 2019. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-019-02329-w.pdf.

Manzo:2019:ERF

[MS19b] F. Manzo and E. Scoppola. Exact results on the first hitting via conditional strong quasi-stationary times and applications to metastability. Journal of Statistical Physics, 174(6):1239– 1262, March 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Montero:2019:TWR

[MS19c] Ana M. Montero and Andrés Santos. Triangle-well and ramp interactions in one-dimensional fluids: A fully analytic exact solution. Journal of Statistical Physics, 175(2):269–288, April 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

McCalla:2013:ESV

[MSB13] Scott G. McCalla, Martin B. Short, and P. Jeffrey Brantingham. The effects of sacred value networks within an evolutionary, adversarial game. Journal of Statistical Physics, 151 (3–4):673–688, May 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0678-4.

Mandjes:2018:QDE

[MSB18] Michel Mandjes, Nicos J. Starreveld, and René Bekker. Queues on a dynamically evolving graph. Journal of Statistical Physics, 173(3–4):1124–1148, November 2018. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-018-2036-7.pdf.

Montino:2016:TSI

[MSLT16] Alessandro Montino, Nahuel Soprano-Loto, and Dimitrios Tsagkarogiannis. Thermodynamics for spatially inhomogeneous magnetization and Young–Gibbs measures. Journal of Statistical Physics, 164(6):1318–1353, September 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-016-1592-y.pdf.

Maruvka:2011:SDG

[MSS^{+}11a] Yosef E. Maruvka, Nadav M. Shnerb, Sorin Solomon, Gur Yaari, and David A. Kessler. Slicing and dicing the genome: A statistical physics approach to population genetics. Journal of Statistical Physics, 142(6):1302–1316, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-0113-7.

Mehta:2011:SMT

[MSS11b] Pankaj Mehta, David J. Schwab, and Anirvan M. Sengupta. Statistical mechanics of transcription-factor binding site discovery using hidden Markov models. Journal of Statistical Physics, 142(6):1187–1205, April 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0102-x.

Mazel:2015:CWM

[MSS15] A. Mazel, Y. Suhov, and I. Stuhl. A classical WR model with q particle types. Journal of Statistical Physics, 159(5):1040– 1086, June 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1219-8.

Marchetti:2010:OPO

[MSV10] D. H. U. Marchetti, V. Sidoravicius, and M. E. Vares. Oriented percolation in one-dimensional 1/|x − y|^{2} percolation models. Journal of Statistical Physics, 139(6):941–959, June 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-9966-z. See commentary [MSV13].

Marchetti:2013:COP

[MSV13] D. H. U. Marchetti, V. Sidoravicius, and M. E. Vares. Commentary to: Oriented Percolation in One-Dimension 1/|x−y|^{2} Percolation Models. Journal of Statistical Physics, 150(4): 804–805, February 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-013-0702-3.

pdf. See [MSV10].

Magdziarz:2012:LPL

[MSZ12] Marcin Magdziarz, Władysław Szczotka, and Piotr Zebrowski. Langevin picture of Lévy walks and their extensions. Journal of Statistical Physics, 147(1):74–96, April 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0465-2.

Morikuni:2011:QJS

[MT11a] Yohei Morikuni and Hal Tasaki. Quantum Jarzynski–Sagawa– Ueda relations. Journal of Statistical Physics, 143(1):1–10, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0153-7.

Motsch:2011:NMS

[MT11b] Sebastien Motsch and Eitan Tadmor. A new model for selforganized dynamics and its flocking behavior. Journal of Statistical Physics, 144(5):923–947, September 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0285-9.

Mehdipour:2016:SMH

[MT16] P. Mehdipour and A. Tahzibi. SRB measures and homoclinic relation for endomorphisms. Journal of Statistical Physics, 163(1):139–155, April 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1458-3.

Maes:2017:MDD

[MT17] Christian Maes and Thimothée Thiery. Midpoint distribution of directed polymers in the stationary regime: Exact result through linear response. Journal of Statistical Physics, 168 (5):937–963, September 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Maldonado:2018:CLR

[MTVU18] C. Maldonado, L. Trejo-Valencia, and E. Ugalde. Constantlength random substitutions and Gibbs measures. Journal of Statistical Physics, 171(2):269–287, April 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Mihailescu:2013:EPC

[MU13] Eugen Mihailescu and Mariusz Urbański. Entropy production for a class of inverse SRB measures. Journal of Statistical Physics, 150(5):881–888, March 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0672-x.

Mayer:2015:CAR

[MU15] Volker Mayer and Mariusz Urbański. Countable alphabet random subshifts of finite type with weakly positive transfer operator. Journal of Statistical Physics, 160(5):1405–1431, September 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1287-9.

Mihailescu:2016:OFM

[MU16] Eugen Mihailescu and Mariusz Urbański. Overlap functions for measures in conformal iterated function systems. Journal of Statistical Physics, 162(1):43–62, January 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1389-4.

Magnen:2018:SLK

[MU18] Jacques Magnen and Jérémie Unterberger. The scaling limit of the KPZ equation in space dimension 3 and higher. Journal of Statistical Physics, 171(4):543–598, May 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Muller:2011:SAS

[Mül11] Marcus Müller. Studying amphiphilic self-assembly with soft coarse-grained models. Journal of Statistical Physics, 145(4): 967–1016, November 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-011-0302-z.
pdf.

Mulken:2016:EQT

[Mül16] Oliver Mülken. Enhanced quantum transport in multiplex networks. Journal of Statistical Physics, 162(3):644–651, February 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1434-3.

Murugan:2016:BID

[MV16] Arvind Murugan and Suriyanarayanan Vaikuntanathan. Biological implications of dynamical phases in non-equilibrium networks. Journal of Statistical Physics, 162(5):1183–1202, March 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1445-0.

Morozov:2019:SIP

[MvS19] Alexander Morozov and Wim van Saarloos. Subcritical instabilities in plane Poiseuille flow of an Oldroyd-B fluid. Journal of Statistical Physics, 175(3–4):554–577, May 2019. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-019-02268-6.pdf.

Mebane:2010:GMS

[MW10] D. S. Mebane and J. H. Wang. A general method of solution for the cluster variation method in ionic solids, with application to diffusionless transitions in yttria-stabilized zirconia. Journal of Statistical Physics, 139(4):727–742, May 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-9963-2.

Marchetti:2012:ALT

[MW12a] Domingos H. U. Marchetti and Walter F. Wreszinski. Anderson-like transition for a class of random sparse models in d ≥= 2 dimensions. Journal of Statistical Physics, 146(5):885– 899, March 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0439-4.

Martinelli:2012:GDQ

[MW12b] Fabio Martinelli and Marc Wouts. Glauber dynamics for the quantum Ising model in a transverse field on a regular tree. Journal of Statistical Physics, 146(5):1059–1088, March 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0436-7.

Morimoto:2016:MVS

[MWY16] Yoshinori Morimoto, Shuaikun Wang, and Tong Yang. Measure valued solutions to the spatially homogeneous Boltzmann equation without angular cutoff. Journal of Statistical Physics, 165(5):866–906, December 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1655-0.

Mackey:2019:MER

[MZ19] Benjamin Mackey and Dapeng Zhan. Multipoint estimates for radial and whole-plane SLE. Journal of Statistical Physics, 175(5):879–903, June 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Nakano:2014:LSO

[Nak14] Fumihiko Nakano. Level statistics for one-dimensional Schrödinger operators and Gaussian beta ensemble. Journal of Statistical Physics, 156(1):66–93, July 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0987-x.

Nakano:2017:FDS

[Nak17] Fumihiko Nakano. Fluctuation of density of states for 1 d Schrödinger operators. Journal of Statistical Physics, 166 (6):1393–1404, March 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-017-1728-8.

Nakajima:2019:POP

[Nak19] Shuta Nakajima. On properties of optimal paths in firstpassage percolation. Journal of Statistical Physics, 174(2): 259–275, January 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Nandori:2011:RPS

[Nán11] Péter Nándori. Recurrence properties of a special type of heavy-tailed random walk. Journal of Statistical Physics, 142 (2):342–355, January 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0116-4.

Nandori:2016:LEI

[Nán16] Péter Nándori. Local equilibrium in inhomogeneous stochastic models of heat transport. Journal of Statistical Physics, 164(2):410–437, July 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1551-7.

Naze:2018:WCD

[Naz18] Pierre Nazé. From weakly chaotic dynamics to deterministic subdiffusion via copula modeling. Journal of Statistical Physics, 171(3):434–448, May 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Noble:2013:RCD

[NBB13] Charleston Noble, James P. Bagrow, and Dirk Brockmann. The role of caretakers in disease dynamics. Journal of Statistical Physics, 152(4):787–798, August 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0787-8.

Nerattini:2014:OPC

[NBK14] R. Nerattini, J. S. Brauchart, and M. K.-H. Kiessling. Optimal N-point configurations on the sphere: “Magic” numbers and Smale’s 7th problem. Journal of Statistical Physics, 157(6):1138–1206, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1107-7.

Novick-Cohen:2010:UBC

[NC10] Amy Novick-Cohen. Upper bounds for coarsening for the deep quench obstacle problem. Journal of Statistical Physics, 141 (1):142–157, October 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0040-7.

Norouzi:2016:CNE

[NE16] Ali Norouzi and Javad Abolfazli Esfahani. Capturing nonequilibrium effects of micro/nano scale gaseous flow using a novel lattice Boltzmann model. Journal of Statistical Physics, 162(3):712–726, February 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1420-9.

Neirotti:2012:LUC

[Nei12] J. P. Neirotti. Learning in ultrametric committee machines. Journal of Statistical Physics, 149(5):887–897, November 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0636-1.

Neumaier:2014:ARC

[Neu14] Arnold Neumaier. Analytic representation of critical equations of state. Journal of Statistical Physics, 155(3):603–624, May 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-0967-1.

Nguyen:2010:SAS

[NG10] Binh T. Nguyen and Denis S. Grebenkov. A spectral approach to survival probabilities in porous media. Journal of Statistical Physics, 141(3):532–554, November 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0054-1.

Nguyen:2017:BPG

[Ngu17] Dinh-Thi Nguyen. On blow-up profile of ground states of boson stars with external potential. Journal of Statistical Physics, 169(2):395–422, October 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Nguyen:2018:SML

[Ngu18] Hung D. Nguyen. The small-mass limit and white-noise limit of an infinite dimensional generalized Langevin equation. Journal of Statistical Physics, 173(2):411–437, October 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Nguyen:2019:FLS

[Ngu19a] Thu Dang Thien Nguyen. Fick Law and sticky Brownian motions. Journal of Statistical Physics, 174(2):494–518, January 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Nguyen:2019:PMR

[Ngu19b] Thu Dang Thien Nguyen. Particle model for the reservoirs in the simple symmetric exclusion process. Journal of Statistical Physics, 175(2):402–417, April 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Nigro:2015:LIM

[Nig15] Alessandro Nigro. Lattice integrals of motion of the Ising model on the strip. Journal of Statistical Physics, 159(2):380– 392, April 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1186-0.

Nakagawa:2019:GTH

[NiS19a] Naoko Nakagawa and Shin ichi Sasa. Global thermodynamics for heat conduction systems. Journal of Statistical Physics, 177(5):825–888, December 2019. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-019-02393-2.pdf.

Nakano:2019:SME

[NiS19b] Hiroyoshi Nakano and Shin ichi Sasa. Statistical mechanical expressions of slip length. Journal of Statistical Physics, 176 (2):312–357, July 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Nakano:2019:SCP

[NiS19c] Hiroyoshi Nakano and Shin ichi Sasa. Surface critical phenomena of a free Bose gas with enhanced hopping at the surface. Journal of Statistical Physics, 174(4):762–799, February 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Neha:2019:ASD

[NKK19] Neha, Divya Katyal, and Rama Kant. Anomalous stretching dynamics of tagged monomer of branched polymer in layered random flows. Journal of Statistical Physics, 177(5):936–959, December 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Nath:2015:HAE

[NKR15] Trisha Nath, Joyjit Kundu, and R. Rajesh. High-activity expansion for the columnar phase of the hard rectangle gas. Journal of Statistical Physics, 160(5):1173–1197, September 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1285-y.

Nie:2011:CNI

[NML^{+}11] L. R. Nie, D. C. Mei, X. M. Lv, X. X. Sun, and P. Li. Colored noise induced net voltage of the underdamped Josephson junction. Journal of Statistical Physics, 143(3):532–542, May 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0160-8.

Nadal:2011:SDQ

[NMV11] Celine Nadal, Satya N. Majumdar, and Massimo Vergassola. Statistical distribution of quantum entanglement for a random bipartite state. Journal of Statistical Physics, 142(2):403–438, January 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-0108-4.

Novak:2015:LTH

[Nov15] Jonathan Novak. Lozenge tilings and Hurwitz numbers. Journal of Statistical Physics, 161(2):509–517, October 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1330-x.

Napiorkowski:2012:BCL

[NP12] Marek Napiórkowski and Jaroslaw Piasecki. The bulk correlation length and the range of thermodynamic Casimir forces at Bose–Einstein condensation. Journal of Statistical Physics, 147(6):1145–1155, July 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-012-0522-x.
pdf.

Napiorkowski:2014:RBC

[NP14] Marek Napiórkowski and Jaroslaw Piasecki. On the relation between Casimir forces and bulk correlations. Journal of Statistical Physics, 156(6):1136–1145, September 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-014-1050-7.pdf.

Nekovar:2016:FTA

[NP16] S. Nekovar and G. Pruessner. A field-theoretic approach to the Wiener sausage. Journal of Statistical Physics, 163(3):604– 641, May 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-016-1483-2.pdf.

Nakano:2010:BTD

[NS10] Fumihiko Nakano and Taizo Sadahiro. A bijection theorem for domino tilings with diagonal impurities. Journal of Statistical Physics, 139(4):565–597, May 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9954-3.

Nagahata:2011:SGM

[NS11] Yukio Nagahata and Makiko Sasada. Spectral gap for multispecies exclusion processes. Journal of Statistical Physics, 143(2):381–398, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0176-0.

Nardi:2012:SAS

[NS12] F. R. Nardi and C. Spitoni. Sharp asymptotics for stochastic dynamics with parallel updating rule. Journal of Statistical Physics, 146(4):701–718, February 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0413-6.

Naeij:2016:PMU

[NS16] Hamid Reza Naeij and Afshin Shafiee. Position-momentum uncertainty relation for an open macroscopic quantum system. Journal of Statistical Physics, 165(6):1141–1152, December 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1666-x.

Nachtergaele:2012:QHO

[NSS12] Bruno Nachtergaele, Robert Sims, and Günter Stolz. Quantum harmonic oscillator systems with disorder. Journal of Statistical Physics, 149(6):969–1012, December 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0652-1.

Nandori:2012:CLT

[NSV12] Péter Nándori, Domokos Szász, and Tamás Varjú. A central limit theorem for time-dependent dynamical systems. Journal of Statistical Physics, 146(6):1213–1220, March 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0451-8.

Nyczka:2013:AII

[NSW13] Piotr Nyczka and Katarzyna Sznajd-Weron. Anticonformity or independence? — Insights from statistical physics. Journal of Statistical Physics, 151(1–2):174–202, April 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0701-4.

Napolitano:2016:IMR

[NT16] George M. Napolitano and Tatyana S. Turova. The Ising model on the random planar causal triangulation: Bounds on the critical line and magnetization properties. Journal of Statistical Physics, 162(3):739–760, February 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1430-7.

Nakajima:2017:EEP

[NT17] Satoshi Nakajima and Yasuhiro Tokura. Excess entropy production in quantum system: Quantum master equation approach. Journal of Statistical Physics, 169(5):902–928, December 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Nakano:2018:GBE

[NT18] Fumihiko Nakano and Khanh Duy Trinh. Gaussian beta ensembles at high temperature: Eigenvalue fluctuations and bulk statistics. Journal of Statistical Physics, 173(2):295–321, October 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Niethammer:2016:URS

[NTV16] B. Niethammer, S. Throm, and J. J. L. Velázquez. A uniqueness result for self-similar profiles to Smoluchowski’s coagulation equation revisited. Journal of Statistical Physics, 164(2):399–409, July 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1553-5.

Niethammer:2014:USS

[NV14] B. Niethammer and J. J. L. Velázquez. Uniqueness of self-similar solutions to Smoluchowski’s coagulation equations for kernels that are close to constant. Journal of Statistical Physics, 157(1):158–181, October 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1070-3.

Nagar:2011:TRM

[NVL11] Apoorva Nagar, Angelo Valleriani, and Reinhard Lipowsky. Translation by ribosomes with mRNA degradation: Exclusion processes on aging tracks. Journal of Statistical Physics, 145(5):1385–1404, December 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-011-0347-z.
pdf.

Nota:2019:KDR

[NWL19] Alessia Nota, Raphael Winter, and Bertrand Lods. Kinetic description of a Rayleigh gas with annihilation. Journal of Statistical Physics, 176(6):1434–1462, September 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ny:2013:AGP

[Ny13] Arnaud Le Ny. Almost Gibbsianness and parsimonious description of the decimated 2 d–Ising model. Journal of Statistical Physics, 152(2):305–335, July 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0773-1.

Nardi:2016:HTA

[NZB16] F. R. Nardi, A. Zocca, and S. C. Borst. Hitting time asymptotics for hard-core interactions on grids. Journal of Statistical Physics, 162(2):522–576, January 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-015-1391-
x;   http://link.springer.com/content/pdf/10.1007/
s10955-015-1391-x.pdf.

Ouellette:2011:PLT

[OBX11] Nicholas T. Ouellette, Eberhard Bodenschatz, and Haitao Xu. Path lengths in turbulence. Journal of Statistical Physics, 145 (1):93–101, October 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0323-7.

OCarroll:2012:APT

[O’C12a] Michael O’Carroll. Analytic perturbation theory for bound states in the transfer matrix spectrum of weakly correlated lattice ferromagnetic spin systems. Journal of Statistical Physics, 146(4):864–869, February 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0428-7.

OMalley:2012:RRG

[OC12b] Daniel O’Malley and John H. Cushman. Random renormalization group operators applied to stochastic dynamics. Journal of Statistical Physics, 149(5):943–950, November 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0630-7.

OMalley:2012:RGC

[OC12c] Daniel O’Malley and John H. Cushman. A renormalization group classification of nonstationary and/or infinite second moment diffusive processes. Journal of Statistical Physics, 146(5):989–1000, March 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0448-3.

Obuchi:2015:LPR

[OCM15] Tomoyuki Obuchi, Simona Cocco, and Rémi Monasson. Learning probabilities from random observables in high dimensions: The maximum entropy distribution and others. Journal of Statistical Physics, 161(3):598–632, November 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1341-7.

Orsingher:2011:VFV

[OD11] Enzo Orsingher and Mirko D’Ovidio. Vibrations and fractional vibrations of rods, plates and Fresnel pseudo-processes. Journal of Statistical Physics, 145(1):143–174, October 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0309-5.

Opoku:2018:CCW

[OEA18] Alex Akwasi Opoku, Kwame Owusu Edusei, and Richard Kwame Ansah. A conditional Curie–Weiss model for stylized multigroup binary choice with social interaction. Journal of Statistical Physics, 171(1):106–126, April 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Omidvar:2018:SOS

[OF18] Hamed Omidvar and Massimo Franceschetti. Self-organized segregation on the grid. Journal of Statistical Physics, 170(4): 748–783, February 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ohkubo:2010:DIP

[Ohk10] Jun Ohkubo. Duality in interacting particle systems and Boson representation. Journal of Statistical Physics, 139(3):454– 465, May 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-009-9910-2.

Ohkubo:2014:LAD

[Ohk14] Jun Ohkubo. Lie algebraic discussions for time-inhomogeneous linear birth–death processes with immigration. Journal of Statistical Physics, 157(2):380–391, October 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1068-x.

Ohta:2014:JTK

[OiS14] Hiroki Ohta and Shin ichi Sasa. Jamming transition in kinetically constrained models with reflection symmetry. Journal of Statistical Physics, 155(5):827–842, June 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0978-y.

Olivier:2014:MAC

[OK14] Gerrit Olivier and Michael Kastner. Microcanonical analysis of the Curie–Weiss anisotropic quantum Heisenberg model in a magnetic field. Journal of Statistical Physics, 157(3):456–473, November 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1093-9.

Orlandini:2011:CRE

[OMC11] Sergio Orlandini, Simone Meloni, and Giovanni Ciccotti. Combining rare events techniques: Phase change in Si nanoparticles. Journal of Statistical Physics, 145(4):812–830, November 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0390-9.

Ong:2014:PSC

[Ong14] Darren C. Ong. Purely singular continuous spectrum for CMV operators generated by subshifts. Journal of Statistical Physics, 155(4):763–776, May 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-0974-2.

Ono:2011:RTL

[Ono11] Kotaro Ono. Rigorous treatment of the liquid–vapor transition in a polydisperse system with Kac interaction. Journal of Statistical Physics, 142(1):43–57, January 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0101-y.

Osada:2018:DAD

[OO18] Hirofumi Osada and Shota Osada. Discrete approximations of determinantal point processes on continuous spaces: Tree representations and tail triviality. Journal of Statistical Physics, 170(2):421–435, January 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-017-1928-2.
pdf.

Orsingher:2012:CRS

[OP12] Enzo Orsingher and Federico Polito. Compositions, random sums and continued random fractions of Poisson and fractional Poisson processes. Journal of Statistical Physics, 148(2):233– 249, August 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0534-6.

Orsingher:2010:FNL

[OPS10] Enzo Orsingher, Federico Polito, and Ludmila Sakhno. Fractional non-linear, linear and sublinear death processes. Journal of Statistical Physics, 141(1):68–93, October 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0045-2.

Opoku:2015:CHL

[OR15] Alex Opoku and Frank Redig. Coupling and hydrodynamic limit for the inclusion process. Journal of Statistical Physics, 160(3):532–547, August 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1277-y.

Osmanovic:2016:NIO

[OR16] Dino Osmanović and Yitzhak Rabin. Neighborhood identity ordering and quenched to annealed transition in random bond models. Journal of Statistical Physics, 162(1):186–198, January 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1401-z.

Oliveira:2019:IDR

[OR19] Roberto I. Oliveira and Guilherme H. Reis. Interacting diffusions on random graphs with diverging average degrees: Hydrodynamics and large deviations. Journal of Statistical Physics, 176(5):1057–1087, September 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Orsingher:2016:MAR

[ORS16] Enzo Orsingher, Costantino Ricciuti, and Francesco Sisti. Motion among random obstacles on a hyperbolic space. Journal of Statistical Physics, 162(4):869–886, February 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1450-y.

ORourke:2015:PIE

[ORSV15] Sean O’Rourke, David Renfrew, Alexander Soshnikov, and Van Vu. Products of independent elliptic random matrices. Journal of Statistical Physics, 160(1):89–119, July 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1246-5.

Ott:2015:OLE

[ORW15] William Ott, Mauricio A. Rivas, and James West. Observing Lyapunov exponents of infinite-dimensional dynamical systems. Journal of Statistical Physics, 161(5):1098–1111, December 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1376-9.

Osipov:2016:WAS

[Osi16] V. Al. Osipov. Wavelet analysis on symbolic sequences and two-fold de Bruijn sequences. Journal of Statistical Physics, 164(1):142–165, July 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1537-5.

Ostrovsky:2016:BBD

[Ost16] Dmitry Ostrovsky. On Barnes beta distributions and applications to the maximum distribution of the 2D Gaussian free field. Journal of Statistical Physics, 164(6):1292–1317, September 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1591-z.

Otwinowski:2011:SES

[OTNN11] Jakub Otwinowski, Sorin Tanase-Nicola, and Ilya Nemenman. Speeding up evolutionary search by small fitness fluctuations.

Journal of Statistical Physics, 144(2):367–378, July 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-011-0199-6.pdf.

Ozogany:2015:MEM

[OV15] Katalin Ozogány and Tamás Vicsek. Modeling the emergence of modular leadership hierarchy during the collective motion of herds made of harems. Journal of Statistical Physics, 158(3):628–646, February 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1131-7.

OMalley:2014:MID

[OVC14] D. O’Malley, V. V. Vesselinov, and J. H. Cushman. A method for identifying diffusive trajectories with stochastic models. Journal of Statistical Physics, 156(5):896–907, September 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1035-6.

Orzel:2011:CSA

[OW11] Sebastian Orzel and Agnieszka Wylomańska. Calibration of the subdiffusive arithmetic Brownian motion with tempered stable waiting-times. Journal of Statistical Physics, 143(3):447–454, May 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0191-1.

Pah:2010:SPC

[Pah10] Chin Hee Pah. Single polygon counting on Cayley tree of order 3. Journal of Statistical Physics, 140(1):198–207, July 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-9989-5.

Pallard:2011:AVS

[Pal11] Christophe Pallard. About the velocities of self-interacting particles in the 1-D torus. Journal of Statistical Physics, 144 (4):826–830, August 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0295-7.

Panchenko:2012:SKM

[Pan12] Dmitry Panchenko. The Sherrington–Kirkpatrick model: An overview. Journal of Statistical Physics, 149(2):362–383, October 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0586-7.

Panchenko:2014:SRA

[Pan14] Dmitry Panchenko. Structure of 1-RSB asymptotic Gibbs measures in the diluted p-spin models. Journal of Statistical Physics, 155(1):1–22, April 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-0955-5.

Panchenko:2016:SFR

[Pan16] Dmitry Panchenko. Structure of finite-RSB asymptotic Gibbs measures in the diluted spin glass models. Journal of Statistical Physics, 162(1):1–42, January 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1385-8.

Parisi:2017:MSB

[Par17] Giorgio Parisi. The marginally stable Bethe lattice spin glass revisited. Journal of Statistical Physics, 167(3–4):515–542, May 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Patrick:2011:DWS

[Pat11] A. E. Patrick. A droplet within the spherical model. Journal of Statistical Physics, 142(5):1085–1104, March 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0150-x.

Patrick:2017:SMC

[Pat17] A. E. Patrick. Spherical model on a Cayley tree: Large deviations. Journal of Statistical Physics, 166(1):45–71, January 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1696-4.

Pavan:2011:GEA

[Pav11] V. Pavan. General entropic approximations for canonical systems described by kinetic equations. Journal of Statistical Physics, 142(4):792–827, February 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0130-1.

Premnath:2011:TDC

[PB11] Kannan N. Premnath and Sanjoy Banerjee. On the threedimensional central moment lattice Boltzmann method. Journal of Statistical Physics, 143(4):747–794, May 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0208-9.

Pernici:2019:PCI

[PC19] Mario Pernici and Giovanni M. Cicuta. Proof of a conjecture on the infinite dimension limit of a unifying model for random matrix theory. Journal of Statistical Physics, 175(2):384–401, April 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Perret:2015:CFM

[PCM15] Anthony Perret, Alain Comtet, and Satya N. Majumdar. On certain functionals of the maximum of Brownian motion and their applications. Journal of Statistical Physics, 161(5):1112–1154, December 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1377-8.

Pigolotti:2018:SSM

[PCMM18] Simone Pigolotti, Massimo Cencini, Daniel Molina, and Miguel A. Muñoz. Stochastic spatial models in ecology: A statistical physics approach. Journal of Statistical Physics, 172(1):44–73, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Prado:2017:DLD

[PdOC17] Roberto A. Prado, César R. de Oliveira, and Silas L. Carvalho. Dynamical localization for discrete Anderson Dirac operators. Journal of Statistical Physics, 167(2):260–296, April 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-017-1746-6.

Pimentel:2017:SFH

[PdS17] Leandro P. R. Pimentel and Marcio W. A. de Souza. Shock fluctuations for the Hammersley process. Journal of Statistical Physics, 166(1):169–189, January 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1695-5.

Polettini:2019:EFR

[PE19] Matteo Polettini and Massimiliano Esposito. Effective fluctuation and response theory. Journal of Statistical Physics, 176(1):94–168, July 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Peliti:2011:SMN

[Pel11] L. (Luca) Peliti. Statistical mechanics in a nutshell. In a nutshell. Princeton University Press, Princeton, NJ, USA, 2011. ISBN 0-691-14529-6 (hardcover). xvi + 398 pp. LCCN QC174.8 .P4513 2011.

Pellegrini:2014:CTO

[Pel14] Clément Pellegrini. Continuous time open quantum random walks and non–Markovian Lindblad master equations. Journal of Statistical Physics, 154(3):838–865, February 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0910-x.

Percus:2010:TPD

[Per10] J. K. Percus. On tagged particle dynamics in highly confined fluids. Journal of Statistical Physics, 138(1–3):40–50, February 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-009-9917-8.

Percus:2013:RWF

[Per13] Jerome K. Percus. A random walk to fundamental measure theory — a mini-review at a personal level. Journal of Statistical Physics, 150(3):601–608, February 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0572-0.

Pernici:2017:ENM

[Per17] Mario Pernici. 1/n expansion for the number of matchings on regular graphs and monomer–dimer entropy. Journal of Statistical Physics, 168(3):666–679, August 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Percus:2019:SRP

[Per19] Jerome K. Percus. Some remarks on pattern formation in vertebrate embryology. Journal of Statistical Physics, 175(3– 4):725–728, May 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Peskin:2014:KWS

[Pes14] Michael E. Peskin. Ken Wilson: Solving the strong interactions. Journal of Statistical Physics, 157(4–5):651–665, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1048-1.

Peters:2010:SDP

[Pet10] Hjalmar Peters. Statistics of distinguishable particles and resolution of the Gibbs paradox of the first kind. Journal of Statistical Physics, 141(5):785–828, December 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-010-0077-7.pdf.

Parzani:2017:TDV

[PF17] Céline Parzani and Francis Filbet. On a three dimensional vision based collision avoidance model. Journal of Statistical Physics, 168(3):680–706, August 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Pra:2013:CWM

[PFR13] Paolo Dai Pra, Markus Fischer, and Daniele Regoli. A Curie–Weiss model with dissipation. Journal of Statistical Physics, 152(1):37–53, July 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0756-2.

Peters:2010:UUC

[PG10] Ole Peters and Michelle Girvan. Universality under conditions of self-tuning. Journal of Statistical Physics, 141(1):53–59, October 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-0039-0.

Pajor-Gyulai:2012:ETJ

[PGS12] Zs. Pajor-Gyulai and D. Szász. Energy transfer and joint diffusion. Journal of Statistical Physics, 146(5):1001–1025, March 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0426-9.

Pickl:2010:DTD

[Pic10] Peter Pickl. Derivation of the time dependent Gross– Pitaevskii equation without positivity condition on the interaction. Journal of Statistical Physics, 140(1):76–89, July 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-9981-0.

Pimentel:2018:LBA

[Pim18] Leandro P. R. Pimentel. Local behaviour of Airy processes. Journal of Statistical Physics, 173(6):1614–1638, December 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Pirjol:2014:EHT

[Pir14] Dan Pirjol. Emergence of heavy-tailed distributions in a random multiplicative model driven by a Gaussian stochastic process. Journal of Statistical Physics, 154(3):781–806, February 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0889-3.

Piraino:2018:PGS

[Pir18a] Mark Piraino. Projections of Gibbs states for Hölder potentials. Journal of Statistical Physics, 170(5):952–961, March 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Pirner:2018:BMG

[Pir18b] Marlies Pirner. A BGK model for gas mixtures of polyatomic molecules allowing for slow and fast relaxation of the temperatures. Journal of Statistical Physics, 173(6):1660–1687, December 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Parker:2010:MET

[PK10] Matt Parker and Alex Kamenev. Mean extinction time in predator–prey model. Journal of Statistical Physics, 141 (2):201–216, October 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0049-y.

Paik:2011:PLB

[PK11] Kyungrock Paik and Praveen Kumar. Power-law behavior in geometric characteristics of full binary trees. Journal of Statistical Physics, 142(4):862–878, February 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0125-y.

Prabhakaran:2019:LPF

[PKBW19] Prasanth Prabhakaran, Alexei Krekhov, Eberhard Bodenschatz, and Stephan Weiss. Leidenfrost pattern formation and boiling. Journal of Statistical Physics, 175(3–4):598– 616, May 2019. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-019-02283-7.pdf.

Palchykov:2013:CRS

[PKDK13] Vasyl Palchykov, János Kertész, Robin Dunbar, and Kimmo Kaski. Close relationships: A study of mobile communication records. Journal of Statistical Physics, 151(3–4):735–744, May 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0705-0.

Pakter:2013:NED

[PL13] Renato Pakter and Yan Levin. Non-equilibrium dynamics of an infinite range XY model in an external field. Journal of Statistical Physics, 150(3):531–539, February 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0576-9.

Plyukhin:2015:ABM

[Ply15] A. V. Plyukhin. Autonomous Brownian motor driven by nonadiabatic variation of internal parameters. Journal of Statistical Physics, 159(2):294–311, April 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1196-y.

Poozesh:2017:FSA

[PM17] Amin Poozesh and Masoud Mirzaei. Flow simulation around cambered airfoil by using conformal mapping and intermediate domain in lattice Boltzmann method. Journal of Statistical Physics, 166(2):354–367, January 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1657-y.

Passemier:2015:ALS

[PMC15] Damien Passemier, Matthew R. McKay, and Yang Chen. Asymptotic linear spectral statistics for spiked Hermitian random matrices. Journal of Statistical Physics, 160(1):120– 150, July 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1233-x.

Pollicott:2017:NTO

[Pol17] Mark Pollicott. A nonlinear transfer operator theorem. Journal of Statistical Physics, 166(3–4):516–524, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-016-1646-1.pdf.

Posnansky:2016:VPH

[Pos16] Oleg Posnansky. Viscoelastic properties of a hierarchical model of soft biological tissue: Two-dimensional and threedimensional cases. Journal of Statistical Physics, 164(5): 1043–1061, September 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1576-y.

Percus:2014:RBM

[PP14] Jerome K. Percus and Ora E. Percus. Reinforced Brownian motion: A prototype. Journal of Statistical Physics, 156(5):917–931, September 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1036-5.

Praprotnik:2011:SPP

[PPK11] Matej Praprotnik, Simon Poblete, and Kurt Kremer. Statistical physics problems in adaptive resolution computer simulations of complex fluids. Journal of Statistical Physics, 145(4):946–966, November 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0312-x.

Povolotsky:2011:GGF

[PPS11] A. M. Povolotsky, V. B. Priezzhev, and G. M. Schütz. Generalized Green functions and current correlations in the TASEP. Journal of Statistical Physics, 142(4):754–791, February 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0133-y.

Pachon:2016:RGA

[PPS16] Angelica Pachon, Federico Polito, and Laura Sacerdote. Random graphs associated to some discrete and continuous time preferential attachment models. Journal of Statistical Physics, 162(6):1608–1638, March 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1462-7.

Perversi:2015:CWC

[PR15a] Eleonora Perversi and Eugenio Regazzini. Characterization of weak convergence of probability-valued solutions of general one-dimensional kinetic equations. Journal of Statistical Physics, 159(4):823–852, May 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1200-6.

Popov:2015:DIP

[PR15b] Serguei Popov and Balázs Ráth. On decoupling inequalities and percolation of excursion sets of the Gaussian free field. Journal of Statistical Physics, 159(2):312–320, April 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1187-z.

Prasad:2019:ABV

[PR19] V. V. Prasad and R. Rajesh. Asymptotic behavior of the velocity distribution of driven inelastic gas with scalar velocities: Analytical results. Journal of Statistical Physics, 176(6): 1409–1433, September 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Pogorui:2011:IRM

[PRD11] Anatoliy A. Pogorui and Ramón M. Rodríguez-Dagnino. Isotropic random motion at finite speed with K-Erlang distributed direction alternations. Journal of Statistical Physics, 145(1):102–112, October 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0328-2.

Pogorui:2012:RMU

[PRD12] Anatoliy A. Pogorui and Ramón M. Rodríguez-Dagnino. Random motion with uniformly distributed directions and random velocity. Journal of Statistical Physics, 147(6):1216–1225, July 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0516-8.

Prolhac:2015:ANB

[Pro15] Sylvain Prolhac. Asymptotics for the norm of Bethe eigenstates in the periodic totally asymmetric exclusion process. Journal of Statistical Physics, 160(4):926–964, August 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1230-0.

Procacci:2017:CRP

[Pro17] Aldo Procacci. A correction to a remark in a paper by Procacci and Yuhjtman: New Lower Bounds for the Convergence Radius of the Virial Series. Journal of Statistical Physics, 168(6):1353–1362, September 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Pizzo:2012:FME

[PRS12] Alessandro Pizzo, David Renfrew, and Alexander Soshnikov. Fluctuations of matrix entries of regular functions of Wigner matrices. Journal of Statistical Physics, 146(3):550–591, February 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0404-7.

Porta:2017:MFE

[PRSS17] Marcello Porta, Simone Rademacher, Chiara Saffirio, and Benjamin Schlein. Mean field evolution of fermions with Coulomb interaction. Journal of Statistical Physics, 166(6): 1345–1364, March 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-017-1725-y.

Paulin:2010:LPR

[PS10a] Daniel Paulin and Domokos Szász. Locally perturbed random walks with unbounded jumps. Journal of Statistical Physics, 141(6):1116–1130, December 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0078-6.

Preis:2010:SPS

[PS10b] Tobias Preis and H. Eugene Stanley. Switching phenomena in a system with no switches. Journal of Statistical Physics, 138(1–3):431–446, February 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9914-y.

Popkov:2011:TPD

[PS11] V. Popkov and G. M. Schütz. Transition probabilities and dynamic structure function in the ASEP conditioned on strong flux. Journal of Statistical Physics, 142(3):627–639, February 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0137-7.

Perret:2014:NEE

[PS14] Anthony Perret and Grégory Schehr. Near-extreme eigenvalues and the first gap of Hermitian random matrices. Journal of Statistical Physics, 156(5):843–876, September 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1044-5.

Pigolotti:2016:PCP

[PS16] Simone Pigolotti and Pablo Sartori. Protocols for copying and proofreading in template-assisted polymerization. Journal of Statistical Physics, 162(5):1167–1182, March 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1399-2.

Pastur:2018:ASP

[PS18] L. Pastur and V. Slavin. The absence of the selfaveraging property of the entanglement entropy of disordered free fermions in one dimension. Journal of Statistical Physics, 170 (2):207–220, January 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Pasricha:2019:MMD

[PS19] Puneet Pasricha and Dharmaraja Selvamuthu. A Markov modulated dynamic contagion process with application to credit risk. Journal of Statistical Physics, 175(2):495–511, April 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Provencher:2012:GED

[PSAPR12] Guillaume Provencher, Yvan Saint-Aubin, Paul A. Pearce, and Jørgen Rasmussen. Geometric exponents of dilute loop models. Journal of Statistical Physics, 147(2):315–350, April 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0464-3.

Pastor-Satorras:2018:ELR

[PSC18] Romualdo Pastor-Satorras and Claudio Castellano. Eigenvector localization in real networks and its implications for epidemic spreading. Journal of Statistical Physics, 173(3–4): 1110–1123, November 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Porto:2013:PLR

[PSCD13] P. Di Porto, S. Di Sabatino, B. Crosignani, and E. DelRe. Power-law relaxation and cumulative information. Journal of Statistical Physics, 153(3):479–485, November 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0840-7.

Park:2010:SEL

[PSK10] Su-Chan Park, Damien Simon, and Joachim Krug. The speed of evolution in large asexual populations. Journal of Statistical Physics, 138(1–3):381–410, February 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-009-9915-x.

Paul:2011:RQA

[PSS11] Gerald Paul, Jia Shao, and H. Eugene Stanley. The random quadratic assignment problem. Journal of Statistical Physics, 145(3):734–744, November 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0308-6.

Pra:2012:SGM

[PSS12] Paolo Dai Pra, Benedetto Scoppola, and Elisabetta Scoppola. Sampling from a Gibbs measure with pair interaction by means of PCA. Journal of Statistical Physics, 149(4):722–737, November 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0612-9.

Popkov:2015:UCT

[PSS15] V. Popkov, J. Schmidt, and G. M. Schütz. Universality classes in two-component driven diffusive systems. Journal of Statistical Physics, 160(4):835–860, August 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1241-x.

Procacci:2016:PCA

[PSS16] Aldo Procacci, Benedetto Scoppola, and Elisabetta Scoppola. Probabilistic cellular automata for low-temperature 2-d Ising model. Journal of Statistical Physics, 165(6):991–1005, December 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1661-2.

Pinski:2012:LTP

[PST12] F. J. Pinski, A. M. Stuart, and F. Theil. Γ-limit for transition paths of maximal probability. Journal of Statistical Physics, 146(5):955–974, March 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0443-8.

Pra:2013:SIT

[PST13] Paolo Dai Pra, Elena Sartori, and Marco Tolotti. Strategic interaction in trend-driven dynamics. Journal of Statistical Physics, 152(4):724–741, August 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-013-0784-
y.

Perry:2018:CAP

[PSVG18] Logan Perry, Mahendra Duwal Shrestha, Michael D. Vose, and Sergey Gavrilets. Collective action problem in heterogeneous groups with punishment and foresight. Journal of Statistical Physics, 172(1):293–312, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Patterson:2017:KED

[PSW17] Robert I. A. Patterson, Sergio Simonella, and Wolfgang Wagner. A kinetic equation for the distribution of interaction clusters in rarefied gases. Journal of Statistical Physics, 169(1): 126–167, October 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Pacifico:2010:TFC

[PT10] Maria José Pacifico and Mike Todd. Thermodynamic formalism for contracting Lorenz flows. Journal of Statistical Physics, 139(1):159–176, April 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9939-2.

Palombi:2014:SDM

[PT14a] Filippo Palombi and Simona Toti. Stochastic dynamics of the multi-state voter model over a network based on interacting cliques and zealot candidates. Journal of Statistical Physics, 156(2):336–367, July 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1003-1.

Pandya:2014:STE

[PT14b] Viraj Pandya and Roderich Tumulka. Spin and the thermal equilibrium distribution of wave functions. Journal of Statistical Physics, 154(1–2):491–502, January 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0849-y.

Pulvirenti:2015:FVC

[PT15] Elena Pulvirenti and Dimitrios Tsagkarogiannis. Finite volume corrections and decay of correlations in the canonical ensemble. Journal of Statistical Physics, 159(5):1017–1039, June 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1207-z.

Patel:2011:QDF

[PVCG11] Amish J. Patel, Patrick Varilly, David Chandler, and Shekhar Garde. Quantifying density fluctuations in volumes of all shapes and sizes using indirect umbrella sampling. Journal of Statistical Physics, 145(2):265–275, October 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0269-9.

Pumir:2013:MSA

[PW13] Alain Pumir and Michael Wilkinson. A model for the shapes of advected triangles. Journal of Statistical Physics, 152(5): 934–953, September 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0789-6.

Pan:2016:FLE

[PWZ16] Guangming Pan, Shaochen Wang, and Wang Zhou. Fluctuations of linear eigenvalues statistics for Wigner matrices: Edge case. Journal of Statistical Physics, 165(3):507–520, November 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1618-5.

Peng:2015:TPP

[PXX15] Junhao Peng, Jian Xiong, and Guoai Xu. Tutte polynomial of pseudofractal scale-free Web. Journal of Statistical Physics, 159(5):1196–1215, June 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1225-x.

Pirjol:2015:GRL

[PZ15] Dan Pirjol and Lingjiong Zhu. On the growth rate of a linear stochastic recursion with Markovian dependence. Journal of Statistical Physics, 160(5):1354–1388, September 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1280-3.

Piatnitski:2017:SLS

[PZ17] A. Piatnitski and E. Zhizhina. Scaling limit of symmetric random walk in high-contrast periodic environment. Journal of Statistical Physics, 169(3):595–613, November 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Procaccia:2019:SHM

[PZ19] Eviatar B. Procaccia and Yuan Zhang. Stationary harmonic measure and DLA in the upper half plane. Journal of Statistical Physics, 176(4):946–980, August 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Qi:2012:FDD

[QD12] Yan Qi and An Du. Ferroelectricity driven by Dzyaloshinskii– Moriya interaction in an anisotropic Heisenberg antiferromagnetic chain. Journal of Statistical Physics, 147(4):821– 831, June 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0504-z.

Qian:2010:CBT

[Qia10] Hong Qian. Cellular biology in terms of stochastic nonlinear biochemical dynamics: Emergent properties, isogenetic variations and chemical system inheritability. Journal of Statistical Physics, 141(6):990–1013, December 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0093-7.

Qiao:2016:LAF

[QLCL16] Yu Qiao, Xuejiao Liu, Minxin Chen, and Benzhuo Lu. A local approximation of fundamental measure theory incorporated into three dimensional Poisson–Nernst-Planck equations to account for hard sphere repulsion among ions. Journal of Statistical Physics, 163(1):156–174, April 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1470-7.

Quastel:2013:SAP

[QR13] Jeremy Quastel and Daniel Remenik. Supremum of the Airy_{2} process minus a parabola on a half line. Journal of Statistical Physics, 150(3):442–456, February 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0633-4.

Quastel:2015:ODK

[QS15] Jeremy Quastel and Herbert Spohn. The one-dimensional KPZ equation and its universality class. Journal of Statistical Physics, 160(4):965–984, August 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-015-1250-
9.

Quinodoz:2011:MM

[QTDS11] Sofia Quinodoz, Michael A. Thomas, Jörn Dunkel, and Eva-Maria Schötz. The more the merrier? Journal of Statistical Physics, 142(6):1324–1336, April 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0157-3. See erratum [QTDS12].

Quinodoz:2012:EMM

[QTDS12] Sofia Quinodoz, Michael A. Thomas, Jörn Dunkel, and Eva-Maria Schötz. Erratum to: The More the Merrier? Journal of Statistical Physics, 149(4):771, November 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-012-0608-5.pdf. See [QTDS11].

Rabin:2011:FLA

[Rab11] Yitzhak Rabin. Fundamental limitation on applicability of statistical methods to study of living organisms and other complex systems. Journal of Statistical Physics, 144(1):213– 216, July 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0252-5.

Rador:2017:DNN

[Rad17] Tonguç Rador. Dynamics of nearest-neighbour competitions on graphs. Journal of Statistical Physics, 169(2):265–302, October 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Rastegin:2011:SGP

[Ras11] Alexey E. Rastegin. Some general properties of unified entropies. Journal of Statistical Physics, 143(6):1120–1135, June 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0231-x.

Rastegin:2012:RCS

[Ras12] Alexey E. Rastegin. Relations for certain symmetric norms and anti-norms before and after partial trace. Journal of Statistical Physics, 148(6):1040–1053, September 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0569-8.

Razbin:2018:EFK

[Raz18] Mohammadhosein Razbin. Elasticity of a filament with kinks. Journal of Statistical Physics, 170(3):642–651, February 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Rupprecht:2012:KAS

[RBGV12] J.-F. Rupprecht, O. Bénichou, D. S. Grebenkov, and R. Voituriez. Kinetics of active surface-mediated diffusion in spherically symmetric domains. Journal of Statistical Physics, 147(5):891–918, June 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0496-8.

Rupprecht:2015:ETD

[RBGV15] J.-F. Rupprecht, O. Bénichou, D. S. Grebenkov, and R. Voituriez. Exit time distribution in spherically symmetric twodimensional domains. Journal of Statistical Physics, 158 (1):192–230, January 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1116-6.

Rabadan:2018:SMS

[RBM^{+}18] Raul Rabadan, Gyan Bhanot, Sonia Marsilio, Nicholas Chiorazzi, Laura Pasqualucci, and Hossein Khiabanian. On statistical modeling of sequencing noise in high depth data to assess tumor evolution. Journal of Statistical Physics, 172(1):143– 155, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Rogers:2011:ITA

[RBR11] David M. Rogers, Thomas L. Beck, and Susan B. Rempe. An information theory approach to nonlinear, nonequilibrium thermodynamics. Journal of Statistical Physics, 145(2):385– 409, October 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0358-9.

Rey-Bellet:2016:ICR

[RBS16a] Luc Rey-Bellet and Konstantinos Spiliopoulos. Improving the convergence of reversible samplers. Journal of Statistical Physics, 164(3):472–494, August 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1565-1.

Rolland:2016:CTR

[RBS16b] Joran Rolland, Freddy Bouchet, and Eric Simonnet. Computing transition rates for the 1-D stochastic Ginzburg–Landau– Allen–Cahn equation for finite-amplitude noise with a rare event algorithm. Journal of Statistical Physics, 162(2):277– 311, January 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1417-4.

Ramola:2017:SRG

[RC17a] Kabir Ramola and Bulbul Chakraborty. Stress response of granular systems. Journal of Statistical Physics, 169(1):1–17, October 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Riechers:2017:FWD

[RC17b] Paul M. Riechers and James P. Crutchfield. Fluctuations when driving between nonequilibrium steady states. Journal of Statistical Physics, 168(4):873–918, August 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Reddy:2016:ISO

[RCV16] Gautam Reddy, Antonio Celani, and Massimo Vergassola. Infomax strategies for an optimal balance between exploration and exploitation. Journal of Statistical Physics, 163(6):1454– 1476, June 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1521-0.

Ramos:2018:ODN

[RdAB18] Marlon Ramos, Marcus A. M. de Aguiar, and Dan Braha. Opinion dynamics on networks under correlated disordered external perturbations. Journal of Statistical Physics, 173 (1):54–76, October 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Roy:2015:TPD

[RDNS15] Anjan Roy, Abhishek Dhar, Onuttom Narayan, and Sanjib Sabhapandit. Tagged particle diffusion in one–dimensional systems with Hamiltonian dynamics — II. Journal of Statistical Physics, 160(1):73–88, July 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1232-y.

Ribeiro:2013:OUL

[RE13] Bruno V. Ribeiro and Yves Elskens. Ornstein–Uhlenbeck limit for the velocity process of an N-particle system interacting stochastically. Journal of Statistical Physics, 153(4):626–640, November 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0835-4.

Reichl:2009:MCS

[Rei09] Linda E. Reichl. A modern course in statistical physics. John Wiley, New York, NY, USA, third edition, 2009. ISBN 3-527-40782-0. xvi + 411 pp. LCCN QC174.8 .R44 2009.

URL http://www.loc.gov/catdir/enhancements/fy1009/
2010287506-b.html;   http://www.loc.gov/catdir/enhancements/
fy1009/2010287506-d.html;   http://www.loc.gov/catdir/
enhancements/fy1009/2010287506-t.html.

Rading:2011:SSD

[RELV11] M. M. Rading, T. A. Engel, R. Lipowsky, and A. Valleriani. Stationary size distributions of growing cells with binary and multiple cell division. Journal of Statistical Physics, 145(1):1– 22, October 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-011-0305-9.pdf.

Renger:2018:FLD

[Ren18] D. R. M. Renger. Flux large deviations of independent and reacting particle systems, with implications for macroscopic fluctuation theory. Journal of Statistical Physics, 172(5):1291– 1326, September 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Rida:2010:EBM

[RESA10] S. Z. Rida, A. M. A. El-Sayed, and A. A. M. Arafa. Effect of bacterial memory dependent growth by using fractional derivatives reaction–diffusion chemotactic model. Journal of Statistical Physics, 140(4):797–811, August 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0007-8.

Ruiz:2018:LBA

[RF18] Diego Ruiz and Jorge Finke. Lyapunov-based anomaly detection in highly-clustered networks. Journal of Statistical Physics, 172(4):1127–1146, August 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Reichl:2017:MHM

[RG17] L. E. Reichl and Erich D. Gust. Microscopic hydrodynamic modes in a binary hard sphere mixture. Journal of Statistical Physics, 168(1):43–67, July 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Roy:2018:FBM

[RG18] Subhadeep Roy and Sanchari Goswami. Fiber bundle model under heterogeneous loading. Journal of Statistical Physics, 170(6):1197–1214, March 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Remacle:2011:WEM

[RGL11] F. Remacle, T. G. Graeber, and R. D. Levine. Whose entropy: A maximal entropy analysis of phosphorylation signaling. Journal of Statistical Physics, 144(2):429–442, July 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0215-x.

Rivoire:2016:IME

[Riv16] Olivier Rivoire. Informations in models of evolutionary dynamics. Journal of Statistical Physics, 162(5):1324–1352, March 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1381-z.

Recher:2012:HBF

[RK12] Christian Recher and Heiner Kohler. From hardcore bosons to free fermions with Painlevé V. Journal of Statistical Physics, 147(3):542–564, May 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0482-1.

Recher:2012:SAW

[RKGZ12] Christian Recher, Mario Kieburg, Thomas Guhr, and Martin R. Zirnbauer. Supersymmetry approach to Wishart correlation matrices: Exact results. Journal of Statistical Physics, 148(6):981–998, September 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0567-x.

Rivoire:2011:VIP

[RL11] Olivier Rivoire and Stanislas Leibler. The value of information for populations in varying environments. Journal of Statistical Physics, 142(6):1124–1166, April 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0166-2.

Ramos:2017:CTP

[RL17] A. D. Ramos and A. Leite. Convergence time and phase transition in a non-monotonic family of probabilistic cellular automata. Journal of Statistical Physics, 168(3):573–594, August 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Rizzo:2010:RCV

[RLCMRT10] Tommaso Rizzo, Alejandro Lage-Castellanos, Roberto Mulet, and Federico Ricci-Tersenghi. Replica cluster variational method. Journal of Statistical Physics, 139(3):375–416, May 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-9938-3.

Rolland:2011:PFT

[RM11] Joran Rolland and Paul Manneville. Pattern fluctuations in transitional plane Couette flow. Journal of Statistical Physics, 142(3):577–591, February 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0126-x.

Reiss:2016:STO

[RM16a] Howard Reiss and José A. Manzanares. Statistical thermodynamics of an ”Open” hard sphere system on the equilibrium fluid isotherm: Study of properties of the freezing transition without direct involvement of the equilibrium solid phase. Journal of Statistical Physics, 164(5):1029–1042, September 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1585-x.

Rezakhah:2016:DCT

[RM16b] Saeid Rezakhah and Yasaman Maleki. Discretization of continuous time discrete scale invariant processes: Estimation and spectra. Journal of Statistical Physics, 164(2):438–448, July 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1541-9.

Rohwer:2015:OFT

[RMN15] C. M. Rohwer and K. K. Müller-Nedebock. Operator formalism for topology–conserving crossing dynamics in planar knot diagrams. Journal of Statistical Physics, 159(1):120–157, April 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1161-1.

Ramezanali:2019:CBU

[RMS19] Mohammad Ramezanali, Partha P. Mitra, and Anirvan M. Sengupta. Critical behavior and universality classes for an algorithmic phase transition in sparse reconstruction. Journal of Statistical Physics, 175(3–4):764–788, May 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Roy:2013:TPD

[RNDS13] Anjan Roy, Onuttom Narayan, Abhishek Dhar, and Sanjib Sabhapandit. Tagged particle diffusion in one-dimensional gas with Hamiltonian dynamics. Journal of Statistical Physics, 150(5):851–866, March 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0673-9.

Raghavan:2019:HSL

[ROS19] Bharath Venkatesh Raghavan and Martin Ostoja-Starzewski. On the hydrodynamic stability of a Lennard-Jones molecular fluid. Journal of Statistical Physics, 177(1):61–77, October 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Renner:2012:GSM

[RP12] C. Renner and J. Peinke. A generalization of scaling models of turbulence. Journal of Statistical Physics, 146(1):25–32, January 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0345-1.

Rocca:2015:NSD

[RPPF15] M. C. Rocca, A. R. Plastino, A. Plastino, and G. L. Ferri. New solution of diffusion–advection equation for cosmic-ray transport using ultradistributions. Journal of Statistical Physics, 161(4):986–1009, November 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1359-x.

Redig:2012:ASM

[RRS12] F. Redig, W. M. Ruszel, and E. Saada. The Abelian sandpile model on a random binary tree. Journal of Statistical Physics, 147(4):653–677, June 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-012-0498-6.
pdf.

Remsing:2011:DCW

[RRW11] Richard C. Remsing, Jocelyn M. Rodgers, and John D. Weeks. Deconstructing classical water models at interfaces and in bulk. Journal of Statistical Physics, 145(2):313–334, October 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0299-3.

Roeck:2013:DST

[RS13] Wojciech De Roeck and Dominique Spehner. Derivation of some translation-invariant Lindblad equations for a quantum Brownian particle. Journal of Statistical Physics, 150(2):320– 352, January 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0649-9.

Rams:2014:DPF

[RS14] Michal Rams and Károly Simon. The dimension of projections of fractal percolations. Journal of Statistical Physics, 154(3):633–655, February 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0886-6.

Radin:2015:SEA

[RS15a] Charles Radin and Lorenzo Sadun. Singularities in the entropy of asymptotically large simple graphs. Journal of Statistical Physics, 158(4):853–865, February 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1151-3.

Roy:2015:EPG

[RS15b] Parna Roy and Parongama Sen. Exit probability in generalised kinetic Ising model. Journal of Statistical Physics, 159(4):893–904, May 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1201-5.

Redig:2018:FDS

[RS18] Frank Redig and Federico Sau. Factorized duality, stationary product measures and generating functions. Journal of Statistical Physics, 172(4):980–1008, August 2018. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-018-2090-1.pdf.

Radin:2019:PGM

[RS19] Charles Radin and Harry L. Swinney. Phases of granular matter. Journal of Statistical Physics, 175(3–4):542–553, May 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Romer:2010:RPP

[RSB10] Rudolf A. Römer and Hermann Schulz-Baldes. The random phase property and the Lyapunov spectrum for disordered multi-channel systems. Journal of Statistical Physics, 140(1):122–153, July 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9986-8.

Rocha:2011:SO

[RST11] Andréa V. Rocha, Alexandre B. Simas, and André Toom. Substitution operators. Journal of Statistical Physics, 143(3):585– 618, May 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0192-0.

Redon:2016:EAM

[RST16] Stephane Redon, Gabriel Stoltz, and Zofia Trstanova. Error analysis of modified Langevin dynamics. Journal of Statistical Physics, 164(4):735–771, August 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1544-6.

Rougerie:2014:QHP

[RSY14] N. Rougerie, S. Serfaty, and J. Yngvason. Quantum Hall phases and plasma analogy in rotating trapped Bose gases. Journal of Statistical Physics, 154(1–2):2–50, January 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0766-0.

Ramos:2011:PTD

[RT11a] A. D. Ramos and A. Toom. Phase transitions in the dynamics of slow random monads. Journal of Statistical Physics, 145(5):1324–1342, December 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0373-x.

Ramos:2011:TRM

[RT11b] A. D. Ramos and A. Toom. Trajectories in random monads. Journal of Statistical Physics, 142(1):201–219, January 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-0087-5.

Ramos:2012:MDT

[RT12] A. D. Ramos and A. Toom. Moments and distributions of trajectories in slow random monads. Journal of Statistical Physics, 147(3):623–633, May 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0483-0.

Ramola:2014:FRM

[RT14] Kabir Ramola and Christophe Texier. Fluctuations of random matrix products and 1D Dirac equation with random mass. Journal of Statistical Physics, 157(3):497–514, November 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1082-z.

Rousseau:2015:HTP

[RT15] Jérôme Rousseau and Mike Todd. Hitting times and periodicity in random dynamics. Journal of Statistical Physics, 161(1):131–150, October 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1325-7.

Robert:2016:DRN

[RT16] Philippe Robert and Jonathan Touboul. On the dynamics of random neuronal networks. Journal of Statistical Physics, 165(3):545–584, November 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1622-9.

Ricciuti:2017:SMM

[RT17a] Costantino Ricciuti and Bruno Toaldo. Semi-Markov models and motion in heterogeneous media. Journal of Statistical Physics, 169(2):340–361, October 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Roy:2017:PCG

[RT17b] Rahul Roy and Hideki Tanemura. Percolation clusters as generators for orientation ordering. Journal of Statistical Physics, 168(6):1259–1275, September 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ruelle:2014:NES

[Rue14] David Ruelle. Non-equilibrium statistical mechanics of turbulence. Journal of Statistical Physics, 157(2):205–218, October 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1086-8.

Ruelle:2016:GDB

[Rue16] David Ruelle. A generalized detailed balance relation. Journal of Statistical Physics, 164(3):463–471, August 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1564-2.

Ruelle:2017:THT

[Rue17] David Ruelle. A theory of hydrodynamic turbulence based on non-equilibrium statistical mechanics. Journal of Statistical Physics, 169(6):1039–1044, December 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ruelle:2018:GCP

[Rue18] David Ruelle. Graph-counting polynomials for oriented graphs. Journal of Statistical Physics, 173(2):243–248, October 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Renaud:2016:ESM

[RVB16] A. Renaud, A. Venaille, and F. Bouchet. Equilibrium statistical mechanics and energy partition for the shallow water model. Journal of Statistical Physics, 163(4):784–843, May 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1496-x.

Rebeschini:2014:CTG

[RvH14] Patrick Rebeschini and Ramon van Handel. Comparison theorems for Gibbs measures. Journal of Statistical Physics, 157 (2):234–281, October 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1087-7.

Reddy:2018:CLT

[RVY18] Tulasi Ram Reddy, Sreekar Vadlamani, and D. Yogeshwaran. Central limit theorem for exponentially quasi-local statistics of spin models on Cayley graphs. Journal of Statistical Physics, 173(3–4):941–984, November 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Rychlewski:2011:SAW

[RW11] G. Rychlewski and S. G. Whittington. Self-avoiding walks and polymer adsorption: Low temperature behaviour. Journal of Statistical Physics, 145(3):661–668, November 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0290-z.

Redig:2012:GNG

[RW12] Frank Redig and Feijia Wang. Gibbs-non-Gibbs transitions via large deviations: Computable examples. Journal of Statistical Physics, 147(6):1094–1112, July 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-012-0523-9.pdf.

Redig:2014:HLT

[RW14] Frank Redig and Feijia Wang. Hamiltonian and Lagrangian for the trajectory of the empirical distribution and the empirical measure of Markov processes. Journal of Statistical Physics, 157(1):182–204, October 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1063-2.

Remsing:2019:IDB

[RW19] Richard C. Remsing and John D. Weeks. The influence of distant boundaries on the solvation of charged particles. Journal of Statistical Physics, 175(3–4):743–763, May 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ryals:2012:NSS

[RY12] Brian Ryals and Lai-Sang Young. Nonequilibrium steady states of some simple 1-D mechanical chains. Journal of Statistical Physics, 146(5):1089–1103, March 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0437-6.

Ryter:2012:EPW

[Ryt12] Dietrich Ryter. The exit problem at weak noise, the twovariable quasipotential, and the Kramers problem. Journal of Statistical Physics, 149(6):1069–1085, December 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0646-z.

Ren:2017:ISP

[RZ17] Jingli Ren and Liying Zhang. Inhomogeneous site percolation on an irregular Bethe lattice with random site distribution. Journal of Statistical Physics, 168(2):394–407, July 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Sakagawa:2012:FEG

[Sak12] Hironobu Sakagawa. On the free energy of a Gaussian membrane model with external potentials. Journal of Statistical Physics, 147(1):18–34, April 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0475-0.

Sakai:2018:HOP

[Sak18] Akira Sakai. Hyperscaling for oriented percolation in 1 + 1 space–time dimensions. Journal of Statistical Physics, 171 (3):462–469, May 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Samaj:2013:TTC

[Sam13] Ladislav Samaj. Thermodynamics of two-component log-gases with alternating charges. Journal of Statistical Physics, 152 (4):599–618, August 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0780-2.

Samaj:2015:CIN

[Sam15] Ladislav Samaj. Counter-ions near a charged wall: Exact results for disc and planar geometries. Journal of Statistical Physics, 161(1):227–249, October 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1308-8.

Samaj:2016:AFA

[Sam16] Ladislav Samaj. Amplitude function of asymptotic correlations along charged wall in Coulomb fluids. Journal of Statistical Physics, 164(2):304–320, July 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1548-2.

Samaj:2017:FSE

[Sam17] Ladislav Samaj. Finite-size effects in non-neutral two– dimensional Coulomb fluids. Journal of Statistical Physics, 168(2):434–446, July 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Samaj:2018:LFS

[Sam18] Ladislav Samaj. Logarithmic finite-size correction in nonneutral two-component plasma on sphere. Journal of Statistical Physics, 173(1):42–53, October 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Samaj:2019:FMC

[Sam19] Ladislav Samaj. Fourth moment of the charge density induced around a guest charge in two-dimensional jellium. Journal of Statistical Physics, 175(6):1066–1079, June 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Sander:2012:KJB

[San12] Leonard M. Sander. Kurt Jacobs: Stochastic Processes for Physicists. Journal of Statistical Physics, 146(4):880–881, February 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0419-8.

Sander:2013:ESP

[San13] Leonard M. (Leonard Michael) Sander. Equilibrium statistical physics: with computer simulations in Python. CreateSpace Independent Publishing Platform, ??, ??, USA, 2013. ISBN 1-4910-6651-2. xii + 321 pp. LCCN QC174.8 .S36 2013.

Sancho:2018:SPA

[San18] J. M. Sancho. Statistical physics approach to thermophoresis of colloids. Journal of Statistical Physics, 172(6):1609–1616, September 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Sollich:2012:NPI

[SB12] Peter Sollich and Adriano Barra. Notes on the polynomial identities in random overlap structures. Journal of Statistical Physics, 147(2):351–374, April 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0476-z.

Sachdeva:2014:ASG

[SB14a] Himani Sachdeva and Mustansir Barma. Analytical study of giant fluctuations and temporal intermittency. Journal of Statistical Physics, 154(4):950–987, February 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0899-1. See erratum [SB14b].

Sachdeva:2014:EAS

[SB14b] Himani Sachdeva and Mustansir Barma. Erratum to: Analytical Study of Giant Fluctuations and Temporal Intermittency in an Aggregation Model. Journal of Statistical Physics, 156(3):617, August 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-014-0948-4.

pdf. See [SB14a].

Seleznev:2015:TTE

[SB15a] V. D. Seleznev and O. Buchina. H-theorem and thermodynamic efficiency of the radiation work inducing a chemically nonequilibrium state of matter. Journal of Statistical Physics, 159(6):1477–1494, June 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1224-y.

Spricer:2015:CMP

[SB15b] Kristoffer Spricer and Tom Britton. The configuration model for partially directed graphs. Journal of Statistical Physics, 161(4):965–985, November 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-015-1360-
4;   http://link.springer.com/content/pdf/10.1007/
s10955-015-1360-4.pdf.

Selley:2016:MFC

[SB16] Fanni Sélley and Péter Bálint. Mean-field coupling of identical expanding circle maps. Journal of Statistical Physics, 164 (4):858–889, August 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1568-y.

Sausset:2010:DSF

[SBK10a] F. Sausset, G. Biroli, and J. Kurchan. Do solids flow? Journal of Statistical Physics, 140(4):718–727, August 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0006-9.

Schneider:2010:TSP

[SBK10b] Johannes J. Schneider, Thomas Bukur, and Antje Krause. Traveling salesman problem with clustering. Journal of Statistical Physics, 141(5):767–784, December 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0080-z.

Shao:2010:GLE

[SC10] Zhi-Gang Shao and Tao Chen. Game of life on the equal degree random lattice. Journal of Statistical Physics, 141(6):952–956, December 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-0092-8.

Singh:2011:FTM

[ScCdP11] Sadanand Singh, Chi cheng Chiu, and Juan J. de Pablo. Flux tempered metadynamics. Journal of Statistical Physics, 145(4):932–945, November 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0301-0.

Schreiber:2010:PWR

[Sch10] Tomasz Schreiber. Polygonal Web representation for higher order correlation functions of consistent polygonal Markov fields in the plane. Journal of Statistical Physics, 140(4):752– 783, August 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-010-0016-7.pdf.

Schick:2011:MFE

[Sch11] M. Schick. Membrane fusion: the emergence of a new paradigm. Journal of Statistical Physics, 142(6):1317–1323, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-0111-9.

Schehr:2012:EVW

[Sch12a] Grégory Schehr. Extremes of N vicious walkers for large N: Application to the directed polymer and KPZ interfaces. Journal of Statistical Physics, 149(3):385–410, November 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0593-8.

Schlemm:2012:AFD

[Sch12b] Eckhard Schlemm. Asymptotic fitness distribution in the Bak–Sneppen model of biological evolution with four species.

Journal of Statistical Physics, 148(2):191–203, August 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0538-2.

Schwarzenberger:2012:UAI

[Sch12c] Fabian Schwarzenberger. Uniform approximation of the integrated density of states for long-range percolation Hamiltonians. Journal of Statistical Physics, 146(6):1156–1183, March 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0431-z.

Schinazi:2013:ICP

[Sch13a] Rinaldo B. Schinazi. An inhomogeneous contact process model for speciation. Journal of Statistical Physics, 151(5):980– 984, June 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0739-3.

Schmidt:2013:GSS

[Sch13b] Bernd Schmidt. Ground states of the 2D sticky disc model: Fine properties and N^{3/4} law for the deviation from the asymptotic Wulff shape. Journal of Statistical Physics, 153 (4):727–738, November 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0852-3.

Schruben:2013:NKT

[Sch13c] Dale Schruben. Novel kinetic theory of the classical isotropic oscillator gas, the flexible Shell model. Journal of Statistical Physics, 153(5):911–930, December 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0839-0.

Schinazi:2015:DRD

[Sch15] Rinaldo B. Schinazi. Does random dispersion help survival? Journal of Statistical Physics, 159(1):101–107, April 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1173-x.

Scoppola:2011:ESC

[Sco11] Benedetto Scoppola. Exact solution for a class of random walk on the hypercube. Journal of Statistical Physics, 143(3):413– 419, May 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0194-y.

Straka:2018:EFB

[SCSS18] Mika J. Straka, Guido Caldarelli, Tiziano Squartini, and Fabio Saracco. From ecology to finance (and back?): A review on entropy-based null models for the analysis of bipartite networks. Journal of Statistical Physics, 173(3–4):1252–1285, November 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Shi:2012:RNI

[SCY^{+}12] Jianghong Shi, Tianqi Chen, Ruoshi Yuan, Bo Yuan, and Ping Ao. Relation of a new interpretation of stochastic differential equations to Ito process. Journal of Statistical Physics, 148 (3):579–590, August 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0532-8.

Salahshoor:2018:NIP

[SdlL18] Hossein Salahshoor and Rafael de la Llave. A numerical investigation of the pinning phenomenon in quasi-periodic Frenkel Kontrova model under an external force. Journal of Statistical Physics, 173(2):398–410, October 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Santamaria:2016:SCM

[SdlPRA16] Ruben Santamaria, Antonio Alvarez de la Paz, Luke Roskop, and Ludwik Adamowicz. Statistical contact model for confined molecules. Journal of Statistical Physics, 164(4):1000–1025, August 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1569-x.

Sepehrinia:2013:ACL

[Sep13] Reza Sepehrinia. Anomalies in conductance and localization length of disordered ladders. Journal of Statistical Physics, 153(6):1039–1048, December 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0869-7.

Serfaty:2014:GLV

[Ser14] Sylvia Serfaty. Ginzburg–Landau vortices, Coulomb gases, and renormalized energies. Journal of Statistical Physics, 154(3):660–680, February 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0891-9.

Sahoo:2015:OLG

[SG15] Shaon Sahoo and Soumya Kanti Ganguly. Optimal linear Glauber model. Journal of Statistical Physics, 159(2):336– 357, April 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1188-y.

Sowinski:2017:IDP

[SG17] Damian Sowinski and Marcelo Gleiser. Information dynamics at a phase transition. Journal of Statistical Physics, 167(5): 1221–1232, June 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Saitta:2011:PTM

[SGC11] L. (Lorenza) Saitta, Attilio Giordana, and Antoine Cornuejols. Phase transitions in machine learning. Cambridge University Press, Cambridge, UK, 2011. ISBN 0-521-76391-6 (hardcover). xv + 383 pp. LCCN Q324.4 .S25 2011.

Stratford:2015:LCC

[SGL15] K. Stratford, A. Gray, and J. S. Lintuvuori. Large colloids in cholesteric liquid crystals. Journal of Statistical Physics, 161(6):1496–1507, December 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-015-1411-
x;   http://link.springer.com/content/pdf/10.1007/
s10955-015-1411-x.pdf.

Shahmansoori:2019:NMD

[SGS19] Nasim Shahmansoori, Farhad Taher Ghahramani, and Afshin Shafiee. Non-Markovian dynamics of macroscopic quantum systems in interaction with non-equilibrium environments. Journal of Statistical Physics, 176(3):541–555, August 2019.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Salgado-Garcia:2013:ESE

[SGU13] R. Salgado-García and E. Ugalde. Exact scaling in the expansion-modification system. Journal of Statistical Physics, 153(5):842–863, December 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0866-x.

Sizhuk:2012:KOP

[SH12] Andrii Sizhuk and Philip Hemmer. Kinetics and optical properties of the strongly driven gas medium of interacting atoms. Journal of Statistical Physics, 147(1):132–180, April 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0457-2.

Shimada:2016:FDS

[SH16] Hirohiko Shimada and Shinobu Hikami. Fractal dimensions of self-avoiding walks and Ising high-temperature graphs in 3D conformal bootstrap. Journal of Statistical Physics, 165 (6):1006–1035, December 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1658-x.

Sharipov:2010:CSL

[Sha10] Felix Sharipov. Comments on ‘Symmetry of the Linearized Boltzmann Equation’ by S. Takata. Journal of Statistical Physics, 139(3):536–537, May 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9955-2. See

[Tak09, Tak10a].

Shang:2012:DCC

[Sha12] Yilun Shang. Distinct clusterings and characteristic path lengths in dynamic small-world networks with identical limit degree distribution. Journal of Statistical Physics, 149(3): 505–518, November 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0605-8.

Shang:2018:FPF

[Sha18] Yilun Shang. False positive and false negative effects on network attacks. Journal of Statistical Physics, 170(1):141–164, January 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Shcherbina:2011:ULE

[Shc11] T. Shcherbina. On universality of local edge regime for the deformed Gaussian unitary ensemble. Journal of Statistical Physics, 143(3):455–481, May 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0196-9.

Shcherbina:2014:ULR

[Shc14] Tatyana Shcherbina. Universality of the local regime for the block band matrices with a finite number of blocks. Journal of Statistical Physics, 155(3):466–499, May 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0964-4.

Shcherbina:2015:FER

[Shc15] M. Shcherbina. On fluctuations of eigenvalues of random band matrices. Journal of Statistical Physics, 161(1):73–90, October 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1324-8.

Shen:2015:MAS

[She15] Shuang Shen. Multifractal analysis of some inhomogeneous multinomial measures with distinct analytic Olsen’s b and B functions. Journal of Statistical Physics, 159(5):1216– 1235, June 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1223-z.

Shiraishi:2013:ASS

[Shi13] Naoto Shiraishi. Anomalous system size dependence of large deviation functions for local empirical measure. Journal of Statistical Physics, 152(2):336–352, July 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0768-y.

Shirley:2015:DER

[Shi15] Christopher Shirley. Decorrelation estimates for random discrete Schrödinger operators in dimension one and applications to spectral statistics. Journal of Statistical Physics, 158(6):1298–1340, March 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1168-7.

Shizgal:2016:PSF

[Shi16] Bernie D. Shizgal. Pseudospectral solution of the Fokker– Planck equation with equilibrium bistable states: the eigenvalue spectrum and the approach to equilibrium. Journal of Statistical Physics, 164(6):1379–1393, September 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1594-9.

Shlosman:2017:CV

[Shl17] Senya Shlosman. Crystals in the void. Journal of Statistical Physics, 169(3):472–479, November 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Simas:2010:HLB

[Sim10] Alexandre B. Simas. Hydrodynamic limit for a boundary driven stochastic lattice gas model with many conserved quantities. Journal of Statistical Physics, 139(2):219–251, April 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-9932-9.

Simon:2011:BAW

[Sim11] Damien Simon. Bethe ansatz for the weakly asymmetric simple exclusion process and phase transition in the current distribution. Journal of Statistical Physics, 142(5):931–951, March 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0146-6.

Simonella:2014:ECF

[Sim14] Sergio Simonella. Evolution of correlation functions in the hard sphere dynamics. Journal of Statistical Physics, 155 (6):1191–1221, June 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0905-7.

Sinai:2010:CTY

[Sin10a] Y. G. Sinai. Chaos theory yesterday, today and Tomorrow. Journal of Statistical Physics, 138(1–3):2–7, February 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-009-9912-0.

Sinai:2010:S

[Sin10b] Ya. G. (Yakov Grigor’evich) Sinai. Selecta. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 0-387-87869-6 (vol. 1), 1-4419-6204-2 (vol. 2). ???? pp. LCCN QA611.5 .S6127 2010.

Site:2011:KFI

[Sit11] L. Delle Site. Kinetic functional of interacting electrons: A numerical procedure and its statistical interpretation. Journal of Statistical Physics, 144(3):663–678, August 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0220-0.

Samaj:2010:CCS

[SJ10] Ladislav Samaj and Bernard Jancovici. Charge and current sum rules in quantum media coupled to radiation II. Journal of Statistical Physics, 139(3):432–453, May 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-9936-5.

Shi:2011:SAC

[SJHW11] Qi-Hong Shi, Rui Jiang, Mao-Bin Hu, and Qing-Song Wu. Strong asymmetric coupling of two parallel exclusion processes. Journal of Statistical Physics, 142(3):616–626, February 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0128-8.

Schmiegelt:2014:EAM

[SK14] B. Schmiegelt and J. Krug. Evolutionary accessibility of modular fitness landscapes. Journal of Statistical Physics, 154(1–2):334–355, January 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0868-8.

Straetmans:2019:MCS

[SK19] John Straetmans and Evgeniy Khain. Modeling cell size dynamics in a confined nonuniform dense cell culture. Journal of Statistical Physics, 176(2):299–311, July 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Savir:2016:BTF

[SKT16] Yonatan Savir, Jacob Kagan, and Tsvi Tlusty. Binding of transcription factors adapts to resolve information–energy tradeoff. Journal of Statistical Physics, 162(5):1383–1394, March 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1388-5.

Santis:2012:DPS

[SL12] Emilio De Santis and Andrea Lissandrelli. Developments in perfect simulation of Gibbs measures through a new result for the extinction of Galton–Watson-like processes. Journal of Statistical Physics, 147(2):231–251, April 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0473-2.

Seri:2011:RHE

[SLdEC11] Marcello Seri, Marco Lenci, Mirko degli Esposti, and Giampaolo Cristadoro. Recurrence and higher ergodic properties for quenched random Lorentz tubes in dimension bigger than two. Journal of Statistical Physics, 144(1):124–138, July 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0244-5.

Schulman:2012:SPZ

[SLM12] L. S. Schulman, J. M. Luck, and Anita Mehta. Spectral properties of zero temperature dynamics in a model of a compacting granular column. Journal of Statistical Physics, 146 (5):924–954, March 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0429-6.

Scoppola:2015:BPN

[SLM15] Benedetto Scoppola, Carlo Lancia, and Riccardo Mariani. On the blockage problem and the non-analyticity of the current for parallel TASEP on a ring. Journal of Statistical Physics, 161(4):843–858, November 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1352-4.

Sirbu:2013:ODD

[SLST13] Alina Sîrbu, Vittorio Loreto, Vito D. P. Servedio, and Francesca Tria. Opinion dynamics with disagreement and modulated information. Journal of Statistical Physics, 151 (1–2):218–237, April 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0724-x.

Sinha:2012:EDS

[SM12a] I. Sinha and A. K. Mukherjee. Effect of droplet size on the first order Ziff–Gulari–Barshad (ZGB) phase transition. Journal of Statistical Physics, 147(4):707–715, June 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0508-8.

Sinha:2012:FOP

[SM12b] I. Sinha and A. K. Mukherjee. First-order phase transition in a modified Ziff–Gulari–Barshad model with self-oscillating reactant coverages. Journal of Statistical Physics, 146(4):669–686, February 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0414-5.

Spyridis:2014:PLD

[SM14] Paul Spyridis and Gene F. Mazenko. Power-law decay and the ergodic–nonergodic transition in simple fluids. Journal of Statistical Physics, 154(4):1030–1056, February 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0907-5.

Slowinski:2015:PDM

[SM15] Piotr Slowiński and Robert S. MacKay. Phase diagrams of majority voter probabilistic cellular automata. Journal of Statistical Physics, 159(1):43–61, April 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1156-y.

Schehr:2013:RPV

[SMCF13] Grégory Schehr, Satya N. Majumdar, Alain Comtet, and Peter J. Forrester. Reunion probability of N vicious walkers: Typical and large fluctuations for large N. Journal of Statistical Physics, 150(3):491–530, February 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0614-7.

Smerlak:2018:NSC

[Sme18] Matteo Smerlak. Natural selection as coarsening. Journal of Statistical Physics, 172(1):105–113, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Saghafi:2019:MNM

[SML19] Z. Saghafi, S. Mahdavifar, and E. Hosseini Lapasar. Markovian and non–Markovian dynamics in the one-dimensional transverse-field XY model. Journal of Statistical Physics, 176 (2):492–504, July 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Stein:2013:SGC

[SN13] Daniel L. Stein and Charles M. (Charles Michael) Newman. Spin glasses and complexity. Primers in complex systems. Princeton University Press, Princeton, NJ, USA, 2013. ISBN 0-691-14733-7 (paperback). xviii + 317 pp. LCCN QC176.8.S68 S74 2013.

Sodin:2009:TWL

[Sod09] Sasha Sodin. The Tracy–Widom law for some sparse random matrices. Journal of Statistical Physics, 136(5):834–841, September 2009. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-009-9813-2. See erratum [Sod17].

Sodin:2011:EAS

[Sod11] Sasha Sodin. An estimate for the average spectral measure of random band matrices. Journal of Statistical Physics, 144(1):

46–59, July 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0243-6.

Sodin:2017:ETW

[Sod17] Sasha Sodin. Erratum to: The Tracy–Widom Law for Some Sparse Random Matrices. Journal of Statistical Physics, 166(5):1343, March 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-017-1715-0.

pdf. See [Sod09].

Soret:2018:EDD

[Sor18] Émilie Soret. Equilibration and diffusion for a dynamical Lorentz gas. Journal of Statistical Physics, 172(3):762–780, August 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Sourlas:2018:EUT

[Sou18] Nicolas Sourlas. The ϵ expansion and universality in three dimensions. Journal of Statistical Physics, 172(2):673–677, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Santis:2013:PSA

[SP13] Emilio De Santis and Mauro Piccioni. Perfect simulation of autoregressive models with infinite memory. Journal of Statistical Physics, 150(6):1017–1029, March 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0719-7.

Salisbury:2016:OPR

[SP16] Jared M. Salisbury and Stephanie E. Palmer. Optimal prediction in the retina and natural motion statistics. Journal of Statistical Physics, 162(5):1309–1323, March 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1439-y.

Summy:2018:FMH

[SP18] D. P. Summy and D. I. Pullin. On the five-moment Hamburger maximum entropy reconstruction. Journal of Statistical Physics, 172(3):854–879, August 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Spohn:2014:NFH

[Spo14] Herbert Spohn. Nonlinear fluctuating hydrodynamics for anharmonic chains. Journal of Statistical Physics, 154(5):1191– 1227, March 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-0933-y.

Simkin:2013:MTF

[SR13] M. V. Simkin and V. P. Roychowdhury. A mathematical theory of fame. Journal of Statistical Physics, 151(1–2):319– 328, April 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0677-5.

Sajadi:2019:RPA

[SR19] Farkhondeh Alsadat Sajadi and Rahul Roy. On rumour propagation among sceptics. Journal of Statistical Physics, 174(4): 935–952, February 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Sasamoto:2010:CRW

[SS10] Tomohiro Sasamoto and Herbert Spohn. The crossover regime for the weakly asymmetric simple exclusion process. Journal of Statistical Physics, 140(2):209–231, July 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-9990-z.

Salas:2011:TMP

[SS11a] Jesús Salas and Alan D. Sokal. Transfer matrices and partition-function zeros for antiferromagnetic Potts models VI. Square lattice with extra-vertex boundary conditions. Journal of Statistical Physics, 144(5):1028–1122, September 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0292-x.

Samuel:2011:TWD

[SS11b] Joseph Samuel and Supurna Sinha. Tops and writhing DNA. Journal of Statistical Physics, 143(2):399–412, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0175-1.

Spohn:2015:NFH

[SS15] Herbert Spohn and Gabriel Stoltz. Nonlinear fluctuating hydrodynamics in one dimension: The case of two conserved fields. Journal of Statistical Physics, 160(4):861–884, August 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1214-0.

Shcherbina:2016:TMA

[SS16] Mariya Shcherbina and Tatyana Shcherbina. Transfer matrix approach to 1d random band matrices: Density of states. Journal of Statistical Physics, 164(6):1233–1260, September 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1593-x.

Sanchis:2017:DCA

[SS17] Rémy Sanchis and Roger W. C. Silva. Dimensional crossover in anisotropic percolation on Z^{d+s}. Journal of Statistical Physics, 169(5):981–988, December 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). See [SS18b].

Salvi:2018:RWP

[SS18a] Michele Salvi and François Simenhaus. Random walk on a perturbation of the infinitely-fast mixing interchange process. Journal of Statistical Physics, 171(4):656–678, May 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Sanchis:2018:CDC

[SS18b] Rémy Sanchis and Roger W. C. Silva. Correction to: Dimensional Crossover in Anisotropic Percolation on Z^{d+s}. Journal of Statistical Physics, 172(3):904, August 2018. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-018-2064-3.pdf. See [SS17].

Shcherbina:2018:URB

[SS18c] Mariya Shcherbina and Tatyana Shcherbina. Universality for 1 d random band matrices: Sigma-model approximation. Journal of Statistical Physics, 172(2):627–664, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Shiraishi:2019:FRB

[SS19] Naoto Shiraishi and Keiji Saito. Fundamental relation between entropy production and heat current. Journal of Statistical Physics, 174(2):433–468, January 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Scagliarini:2015:MSS

[SSB15] Andrea Scagliarini, Mauro Sbragaglia, and Massimo Bernaschi. Mesoscopic simulation study of wall roughness effects in micro-channel flows of dense emulsions. Journal of Statistical Physics, 161(6):1482–1495, December 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1374-y;   http://link.springer.com/
content/pdf/10.1007/s10955-015-1374-y.pdf.

Streib:2014:BAM

[SSBS14] Noah Streib, Amanda Streib, Isabel Beichl, and Francis Sullivan. A binomial approximation method for the Ising model. Journal of Statistical Physics, 156(3):593–605, August 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1016-9.

Suess:2015:HEO

[SSE15] D. Suess, W. T. Strunz, and A. Eisfeld. Hierarchical equations for open system dynamics in fermionic and bosonic environments. Journal of Statistical Physics, 159(6):1408–1423, June 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1236-7.

Sahoo:2012:SLT

[SSR12] Shaon Sahoo, Jean-Pascal Sutter, and S. Ramasesha. Study of low temperature magnetic properties of a single chain magnet with alternate isotropic and non-collinear anisotropic units. Journal of Statistical Physics, 147(1):181–193, April 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0460-7.

Sinclair:2014:AAT

[SST14] Alistair Sinclair, Piyush Srivastava, and Marc Thurley. Approximation algorithms for two-state anti-ferromagnetic spin systems on bounded degree graphs. Journal of Statistical Physics, 155(4):666–686, May 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-0947-5.

Simione:2015:EGS

[SST15] Robert Simione, Dejan Slepcev, and Ihsan Topaloglu. Existence of ground states of nonlocal-interaction energies. Journal of Statistical Physics, 159(4):972–986, May 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1215-z.

Saito:2011:ECR

[ST11a] Keiji Saito and Hal Tasaki. Extended Clausius relation and entropy for nonequilibrium steady states in heat conducting quantum systems. Journal of Statistical Physics, 145 (5):1275–1290, December 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0346-0.

Sartori:2011:NFS

[ST11b] Pablo Sartori and Yuhai Tu. Noise filtering strategies in adaptive biochemical signaling networks. Journal of Statistical Physics, 142(6):1206–1217, April 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0169-z.

Shinault:2011:ACA

[ST11c] Gregory Shinault and Craig A. Tracy. Asymptotics for the covariance of the Airy_{2} process. Journal of Statistical Physics, 143(1):60–71, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-011-0155-5.
pdf.

Samaj:2014:CIB

[ST14] Ladislav Samaj and Emmanuel Trizac. Counter-ions between or at asymmetrically charged walls: 2D free-fermion point.

Journal of Statistical Physics, 156(5):932–947, September 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1053-4.

Salazar:2016:EEC

[ST16a] R. Salazar and G. Téllez. Exact energy computation of the one component plasma on a sphere for even values of the coupling parameter. Journal of Statistical Physics, 164(4):969– 999, August 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1562-4.

Strungaru:2016:DTA

[ST16b] Nicolae Strungaru and Venta Terauds. Diffraction theory and almost periodic distributions. Journal of Statistical Physics, 164(5):1183–1216, September 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1579-8.

Sarna:2018:SWB

[ST18a] Neeraj Sarna and Manuel Torrilhon. On stable wall boundary conditions for the Hermite discretization of the linearised Boltzmann equation. Journal of Statistical Physics, 170(1): 101–126, January 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Scoppola:2018:GMF

[ST18b] Benedetto Scoppola and Alessio Troiani. Gaussian mean field lattice gas. Journal of Statistical Physics, 170(6):1161–1176, March 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Stanley:2011:AFI

[Sta11] H. Eugene Stanley. Y. Aoyama, Y. Fujiwara, Y. Ikeda, H. Iyetomi, W. Souma: Econophysics and Companies: Statistical Life and Death in Complex Business Networks. Journal of Statistical Physics, 145(1):204–205, October 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0300-1.

Stauffer:2013:BRS

[Sta13] Dietrich Stauffer. A biased review of sociophysics. Journal of Statistical Physics, 151(1–2):9–20, April 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0604-9.

Stamatakis:2015:HLM

[Sta15] Marios Georgios Stamatakis. Hydrodynamic limit of mean zero condensing zero range processes with sub-critical initial profiles. Journal of Statistical Physics, 158(1):87–104, January 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1113-9.

Sausset:2010:BPK

[STBT10] François Sausset, Cristina Toninelli, Giulio Biroli, and Gilles Tarjus. Bootstrap percolation and kinetically constrained models on hyperbolic lattices. Journal of Statistical Physics, 138(1–3):411–430, February 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9903-1.

Stenlund:2010:SPC

[Ste10] Mikko Stenlund. A strong pair correlation bound implies the CLT for Sinai billiards. Journal of Statistical Physics, 140(1):154–169, July 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9987-7.

Steinberg:2019:RTN

[Ste19] Victor Steinberg. Role of thermal noise in dynamics of nonequilibrium systems: Macro-, meso- and microscopic. Journal of Statistical Physics, 175(3–4):664–680, May 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Suga:2010:AML

[Sug10] Shinsuke Suga. An accurate multi-level finite difference scheme for 1D diffusion equations derived from the lattice Boltzmann method. Journal of Statistical Physics, 140(3): 494–503, August 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0004-y.

Sun:2018:FCL

[Sun18] Wen Sun. A functional central limit theorem for the Becker– Döring model. Journal of Statistical Physics, 171(1):145–165, April 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Swart:2010:NAR

[SV10] Jan M. Swart and Karel Vrbenský. Numerical analysis of the rebellious voter model. Journal of Statistical Physics, 140(5):873–899, September 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0021-x.

Shcherbakov:2015:LTB

[SV15] Vadim Shcherbakov and Stanislav Volkov. Long term behaviour of locally interacting birth-and-death processes. Journal of Statistical Physics, 158(1):132–157, January 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1122-8.

Szabo:2016:ARR

[SV16] Réka Szabó and Bálint Vető. Ages of records in random walks. Journal of Statistical Physics, 165(6):1086–1101, December 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1671-0.

Sedlmeier:2011:WDI

[SvHM^{+}11] Felix Sedlmeier, Yann von Hansen, Liang Mengyu, Dominik Horinek, and Roland R. Netz. Water dynamics at interfaces and solutes: Disentangling free energy and diffusivity contributions. Journal of Statistical Physics, 145(2):240–252, October 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0338-0.

Silva:2011:GKE

[SVRL11] C. A. B. Silva, Aurea R. Vasconcellos, J. Galvão Ramos, and Roberto Luzzi. Generalized kinetic equation for farfrom-equilibrium many-body systems. Journal of Statistical Physics, 143(5):1020–1034, June 2011. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-011-0222-
y.

Starr:2012:ATI

[SVW12] Shannon Starr, Brigitta Vermesi, and Ang Wei. About thinning invariant partition structures. Journal of Statistical Physics, 148(2):325–344, August 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-012-0544-
4.

Sauerwine:2011:ALR

[SW11] Ben Sauerwine and Michael Widom. Arrhenius lifetimes of RNA structures from free energy landscapes. Journal of Statistical Physics, 142(6):1337–1352, April 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0174-2.

Sosoe:2012:LSL

[SW12] Philippe Sosoe and Percy Wong. Local semicircle law in the bulk for Gaussian β-ensemble. Journal of Statistical Physics, 148(2):204–232, August 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0536-4.

Szehr:2015:PTP

[SW15] Oleg Szehr and Michael M. Wolf. Perturbation theory for parent Hamiltonians of matrix product states. Journal of Statistical Physics, 159(4):752–771, May 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1204-2.

Scher:2010:TEE

[SWB10] Harvey Scher, Karen Willbrand, and Brian Berkowitz. Transport equation evaluation of coupled continuous time random walks. Journal of Statistical Physics, 141(6):1093–1103, December 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-0088-4.

Sedivy:2018:DDS

[SWK^{+}18] Ondrej Sedivý, Daniel Westhoff, Jaromír Kopecek, Carl E. Krill III, and Volker Schmidt. Data-driven selection of tessellation models describing polycrystalline microstructures. Journal of Statistical Physics, 172(5):1223–1246, September 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Spettl:2014:PRG

[SWKS14] Aaron Spettl, Thomas Werz, Carl E. Krill III, and Volker Schmidt. Parametric representation of 3D grain ensembles in polycrystalline microstructures. Journal of Statistical Physics, 154(4):913–928, February 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0893-7.

Shrock:2010:ERP

[SX10a] Robert Shrock and Yan Xu. Exact results on Potts model partition functions in a generalized external field and weightedset graph colorings. Journal of Statistical Physics, 141(6): 909–939, December 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0089-3.

Shrock:2010:WSG

[SX10b] Robert Shrock and Yan Xu. Weighted-set graph colorings. Journal of Statistical Physics, 139(1):27–61, April 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-9937-4.

Shcherbakov:2012:EDR

[SY12] Vadim Shcherbakov and Anatoly Yambartsev. On equilibrium distribution of a reversible growth model. Journal of Statistical Physics, 148(1):53–66, July 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0530-x.

Sisko:2013:GUI

[SYZ13] V. Sisko, A. Yambartsev, and S. Zohren. Growth of uniform infinite causal triangulations. Journal of Statistical Physics, 150(2):353–374, January 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0665-9.

Stefansson:2012:SDT

[SZ12] Sigurdur Ö. Stefánsson and Stefan Zohren. Spectral dimension of trees with a unique infinite spine. Journal of Statistical Physics, 147(5):942–962, June 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0510-1.

Shamis:2018:CWM

[SZ18] Mira Shamis and Ofer Zeitouni. The Curie–Weiss model with complex temperature: Phase transitions. Journal of Statistical Physics, 172(2):569–591, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Shahidi:2019:TI

[SZ19] Farruh Shahidi and Agnieszka Zelerowicz. Thermodynamics via inducing. Journal of Statistical Physics, 175(2):351–383, April 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Szasz:2012:YGS

[Szá12] Domokos Szász. Ya. G. Sinai: Selecta, Volumes I and II. Journal of Statistical Physics, 146(6):1303–1305, March 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0453-6.

Sewran:2015:NHM

[SZS15] Sashwin Sewran, Konstantin G. Zloshchastiev, and Alessandro Sergi. Non-Hamiltonian modeling of squeezing and thermal disorder in driven oscillators. Journal of Statistical Physics, 159(2):255–273, April 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1193-1.

Tsuji:2012:DLP

[TA12] Tetsuro Tsuji and Kazuo Aoki. Decay of a linear pendulum in a free-molecular gas and in a special Lorentz gas. Journal of Statistical Physics, 146(3):620–645, February 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0412-7.

Takeuchi:2016:CSR

[TA16] Kazumasa A. Takeuchi and Takuma Akimoto. Characteristic sign renewals of Kardar–Parisi–Zhang fluctuations. Journal of Statistical Physics, 164(5):1167–1182, September 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1582-0.

Tsuji:2010:RFM

[TAG10] Tetsuro Tsuji, Kazuo Aoki, and François Golse. Relaxation of a free-molecular gas to equilibrium caused by interaction with vessel wall. Journal of Statistical Physics, 140(3):518– 543, August 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-9997-5.

Taggi:2015:CPC

[Tag15] Lorenzo Taggi. Critical probabilities and convergence time of percolation probabilistic cellular automata. Journal of Statistical Physics, 159(4):853–892, May 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1199-8.

Takata:2009:SLB

[Tak09] Shigeru Takata. Symmetry of the linearized Boltzmann equation and its application. Journal of Statistical Physics, 136 (4):751–784, August 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-009-9793-2. See

comments [Sha10].

Takata:2010:RCS

[Tak10a] Shigeru Takata. Reply to the comments of Sharipov on ‘Symmetry of the Linearized Boltzmann Equation’. Journal of Statistical Physics, 139(3):538–539, May 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-9953-4. See [Tak09, Sha10].

Takata:2010:SUL

[Tak10b] Shigeru Takata. Symmetry of the unsteady linearized Boltzmann equation in a fixed bounded domain. Journal of Statistical Physics, 140(5):985–1005, September 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0009-6.

Takata:2015:TMS

[Tak15] Shigeru Takata. A toy-model study of the grazing collisions in the kinetic theory. Journal of Statistical Physics, 160(3):770– 792, August 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1259-0.

Takahasi:2016:RPT

[Tak16] Hiroki Takahasi. Removal of phase transition in the Chebyshev quadratic and thermodynamics for Hénon-Like maps near the first bifurcation. Journal of Statistical Physics, 164 (6):1354–1378, September 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1584-y.

Tanaka:2018:FHM

[Tan18] Akinori Tanaka. Ferromagnetism in the Hubbard model with a gapless nearly-flat band. Journal of Statistical Physics, 170 (2):399–420, January 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Tasaki:2016:TTE

[Tas16] Hal Tasaki. Typicality of thermal equilibrium and thermalization in isolated macroscopic quantum systems. Journal of Statistical Physics, 163(5):937–997, June 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1511-2.

Tasaki:2018:LSM

[Tas18a] Hal Tasaki. Lieb–Schultz–Mattis theorem with a local twist for general one-dimensional quantum systems. Journal of Statistical Physics, 170(4):653–671, February 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Tasaki:2018:LEB

[Tas18b] Hal Tasaki. On the local equivalence between the canonical and the microcanonical ensembles for quantum spin systems. Journal of Statistical Physics, 172(4):905–926, August 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Tasaki:2019:LRO

[Tas19] Hal Tasaki. Long-range order, ”Tower” of states, and symmetry breaking in lattice quantum systems. Journal of Statistical Physics, 174(4):735–761, February 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Tate:2013:VEB

[Tat13] Stephen James Tate. Virial expansion bounds. Journal of Statistical Physics, 153(2):325–338, October 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0831-8.

Tauber:2010:LRB

[Täu10] Uwe C. Täuber. Linda E. Reichl: A Modern Course in Statistical Physics. Journal of Statistical Physics, 141(3):609–611, November 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-0062-1.

Tautenhahn:2011:LCA

[Tau11] Martin Tautenhahn. Localization criteria for Anderson models on locally finite graphs. Journal of Statistical Physics, 144(1): 60–75, July 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0248-1.

Taylor:2016:MEM

[Tay16] Jamie M. Taylor. Maximum entropy methods as the bridge between microscopic and macroscopic theory. Journal of Statistical Physics, 164(6):1429–1459, September 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1587-8.

Taye:2017:IBH

[Tay17] Mesfin Asfaw Taye. Irreversible Brownian heat engine. Journal of Statistical Physics, 169(2):423–440, October 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Toscani:2013:KMT

[TBD13] Giuseppe Toscani, Carlo Brugna, and Stefano Demichelis. Kinetic models for the trading of goods. Journal of Statistical Physics, 151(3–4):549–566, May 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0653-0.

Travers:2011:ASF

[TC11a] Nicholas F. Travers and James P. Crutchfield. Asymptotic synchronization for finite-state sources. Journal of Statistical Physics, 145(5):1202–1223, December 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0349-x.

Travers:2011:ESF

[TC11b] Nicholas F. Travers and James P. Crutchfield. Exact synchronization for finite-state sources. Journal of Statistical Physics, 145(5):1181–1201, December 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0342-4.

Talukdar:2019:TSP

[TD19] Dhritiman Talukdar and Kishore Dutta. Transition of spatial patterns in an interacting Turing system. Journal of Statistical Physics, 174(2):351–364, January 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Telcs:2010:DLP

[Tel10] A. Telcs. Diffusive limits on the Penrose tiling. Journal of Statistical Physics, 141(4):661–668, November 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0072-z.

Temmel:2014:SCU

[Tem14] Christoph Temmel. Sufficient conditions for uniform bounds in abstract polymer systems and explorative partition schemes. Journal of Statistical Physics, 157(6):1225–1254, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1108-6.

Terauds:2013:IPP

[Ter13] Venta Terauds. The inverse problem of pure point diffraction — examples and open questions. Journal of Statistical Physics, 152(5):954–968, September 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0790-0.

Tellez:2012:EVP

[TF12] Gabriel Téllez and Peter J. Forrester. Expanded Vandermonde powers and sum rules for the two-dimensional onecomponent plasma. Journal of Statistical Physics, 148(5): 824–855, September 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0551-5.

Tawfik:2019:ASF

[TFES19] Ashraf M. Tawfik, Horst Fichtner, A. Elhanbaly, and Reinhard Schlickeiser. An analytical study of fractional Klein– Kramers approximations for describing anomalous diffusion of energetic particles. Journal of Statistical Physics, 174(4): 830–845, February 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Tessone:2013:HBT

[TGGS13] Claudio J. Tessone, Antonios Garas, Beniamino Guerra, and Frank Schweitzer. How big is too big? Critical shocks for systemic failure cascades. Journal of Statistical Physics, 151 (3–4):765–783, May 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0723-y.

Tejedor:2012:OWM

[TGP12] Alejandro Tejedor, Javier B. Gomez, and Amalio F. Pacheco. One-way Markov process approach to repeat times of large earthquakes in faults. Journal of Statistical Physics, 149(5): 951–963, November 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0629-0.

Takata:2012:ATT

[TH12] Shigeru Takata and Masanari Hattori. Asymptotic theory for the time-dependent behavior of a slightly rarefied gas over a smooth solid boundary. Journal of Statistical Physics, 147 (6):1182–1215, July 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0512-z.

Thale:2011:ACS

[Thä11] Christoph Thäle. Arak–Clifford–Surgailis tessellations. Basic properties and variance of the total edge length. Journal of Statistical Physics, 144(6):1329–1339, September 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0331-7.

Thiery:2016:SMT

[Thi16] Thimothée Thiery. Stationary measures for two dual families of finite and zero temperature models of directed polymers on the square lattice. Journal of Statistical Physics, 165(1):44– 85, October 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-016-1603-z.

Thompson:2011:CSI

[Tho11] Colin J. Thompson. Colouring solutions of the ice problem. Journal of Statistical Physics, 145(3):647–651, November 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0246-3.

Thomas:2012:FLS

[Tho12] Drew M. Thomas. A not-so-fundamental limitation on studying complex systems with statistics: Comment on Rabin (2011). Journal of Statistical Physics, 149(6):1168–1171, December 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0647-y.

Throm:2018:TBS

[Thr18] Sebastian Throm. Tail behaviour of self-similar profiles with infinite mass for Smoluchowski’s coagulation equation. Journal of Statistical Physics, 170(6):1215–1241, March 2018. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Tian:2014:PEF

[Tia14] Xueting Tian. Pesin’s entropy formula for systems between C^{1} and C^{1+α}. Journal of Statistical Physics, 156(6):1184–1198, September 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1065-0.

Tidriri:2016:CSN

[Tid16] Moulay Tidriri. Construction and study of new measures associated with a given pair of functions. Journal of Statistical Physics, 162(2):577–602, January 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1390-y.

Toth:2015:VGR

[TJ15] Gábor Tóth and Imre M. Jánosi. Vorticity generation by rough walls in 2D decaying turbulence. Journal of Statistical Physics, 161(6):1508–1518, December 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1375-x;   http://link.springer.com/
content/pdf/10.1007/s10955-015-1375-x.pdf.

Tada:2016:TNG

[TK16] Yasuhiro Tada and Tohru Koma. Two no-go theorems on superconductivity. Journal of Statistical Physics, 165(3): 455–470, November 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1629-2.

Tarasov:2015:GCV

[TKK15] S. V. Tarasov, Vl. V. Kocharovsky, and V. V. Kocharovsky. Grand canonical versus canonical ensemble: Universal structure of statistics and thermodynamics in a critical region of Bose–Einstein condensation of an ideal gas in arbitrary trap. Journal of Statistical Physics, 161(4):942–964, November 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1361-3.

Torney:2013:DAR

[TLC13] Colin J. Torney, Simon A. Levin, and Iain D. Couzin. Decision accuracy and the role of spatial interaction in opinion dynamics. Journal of Statistical Physics, 151(1–2):203–217, April 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0700-5.

Tantet:2018:RCA

[TLD18] Alexis Tantet, Valerio Lucarini, and Henk A. Dijkstra. Resonances in a chaotic attractor crisis of the Lorenz flow. Journal of Statistical Physics, 170(3):584–616, February 2018. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Taillefumier:2010:FAF

[TM10] Thibaud Taillefumier and Marcelo O. Magnasco. A fast algorithm for the first-passage times of Gauss–Markov processes with Hölder continuous boundaries. Journal of Statistical Physics, 140(6):1130–1156, September 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-010-0033-6.

Tikhonov:2018:IRT

[TM18] Mikhail Tikhonov and Remi Monasson. Innovation rather than improvement: A solvable high-dimensional model highlights the limitations of scalar fitness. Journal of Statistical Physics, 172(1):74–104, July 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Tchernookov:2013:PIN

[TN13] Martin Tchernookov and Ilya Nemenman. Predictive information in a nonequilibrium critical model. Journal of Statistical Physics, 153(3):442–459, November 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0833-6.

Takata:2018:SKM

[TN18] Shigeru Takata and Takashi Noguchi. A simple kinetic model for the phase transition of the van der Waals fluid. Journal of Statistical Physics, 172(3):880–903, August 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). See correction [TN21].

Takata:2021:CSK

[TN21] Shigeru Takata and Takashi Noguchi. Correction to: A Simple Kinetic Model for the Phase Transition of the van der Waals Fluid. Journal of Statistical Physics, 185(1):??, October 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/
10.1007/s10955-021-02827-w. See [TN18].

Torikai:2012:ESP

[Tor12] Masashi Torikai. Equation of state for parallel rigid spherocylinders. Journal of Statistical Physics, 148(2):345–352, August 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0542-6.

Toscani:2016:EIS

[Tos16] Giuseppe Toscani. Entropy inequalities for stable densities and strengthened central limit theorems. Journal of Statistical Physics, 165(2):371–389, October 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1619-4.

Tossounian:2017:EKM

[Tos17] H. Tossounian. Equilibration in the Kac model using the GTW metric d_{2}. Journal of Statistical Physics, 169(1):168–186, October 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Touboul:2012:LDS

[Tou12] Jonathan Touboul. Limits and dynamics of stochastic neuronal networks with random heterogeneous delays. Journal of Statistical Physics, 149(4):569–597, November 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0607-6.

Touboul:2014:SEN

[Tou14] Jonathan Touboul. Spatially extended networks with singular multi-scale connectivity patterns. Journal of Statistical Physics, 156(3):546–573, August 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-014-1015-
x.

Touchette:2015:ENE

[Tou15] Hugo Touchette. Equivalence and nonequivalence of ensembles: Thermodynamic, macrostate, and measure levels. Journal of Statistical Physics, 159(5):987–1016, June 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1212-2.

Touchette:2018:AEP

[Tou18] Hugo Touchette. Asymptotic equivalence of probability measures and stochastic processes. Journal of Statistical Physics, 170(5):962–978, March 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Terborg:2015:ENU

[TP15] Heinrich Terborg and Luis A. Pérez. Effects of nonuniform occupancy on selective transport through nanochannels. Journal of Statistical Physics, 158(2):494–512, January 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1132-6.

Tilles:2019:CRT

[TP19] Paulo F. C. Tilles and Sergei V. Petrovskii. On the consistency of the reaction-telegraph process within finite domains. Journal of Statistical Physics, 177(4):569–587, November 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Thomas:2012:SM

[TQS12] Michael A. Thomas, Sofia Quinodoz, and Eva-Maria Schötz. Size matters! Journal of Statistical Physics, 148(4):664–676, September 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0514-x.

Tristani:2014:ECE

[Tri14] Isabelle Tristani. Exponential convergence to equilibrium for the homogeneous Boltzmann equation for hard potentials without cut-off. Journal of Statistical Physics, 157(3):474–496, November 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1066-z.

Tristani:2017:LDL

[Tri17] Isabelle Tristani. Landau damping for the linearized Vlasov Poisson equation in a weakly collisional regime. Journal of Statistical Physics, 169(1):107–125, October 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Troubetzkoy:2010:TRE

[Tro10] Serge Troubetzkoy. Typical recurrence for the Ehrenfest wind– tree model. Journal of Statistical Physics, 141(1):60–67, October 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-0026-5.

Takeuchi:2012:EGD

[TS12] Kazumasa A. Takeuchi and Masaki Sano. Evidence for geometry-dependent universal fluctuations of the Kardar– Parisi–Zhang interfaces in liquid-crystal turbulence. Journal of Statistical Physics, 147(5):853–890, June 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0503-0.

Thompson:2019:LBP

[TS19] Jeffrey P. Thompson and Isaac C. Sanchez. A lower bound on the partition function for a strictly neutral charge-symmetric system. Journal of Statistical Physics, 177(6):1207–1215, December 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Tunc:2013:EAS

[TSS13] Ilker Tunc, Maxim S. Shkarayev, and Leah B. Shaw. Epidemics in adaptive social networks with temporary link deactivation. Journal of Statistical Physics, 151(1–2):355–366, April 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0667-7.

Tellez:2012:TDO

[TT12] Gabriel Téllez and Emmanuel Trizac. A two-dimensional one component plasma and a test charge: Polarization effects and effective potential. Journal of Statistical Physics, 146(4):832–849, February 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0424-y.

Tanaka:2016:MFS

[TT16] Akinori Tanaka and Hal Tasaki. Metallic ferromagnetism supported by a single band in a multi-band Hubbard model. Journal of Statistical Physics, 163(5):1049–1068, June 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1515-y.

Takata:2017:GDF

[TT17a] Shigeru Takata and Satoshi Taguchi. Gradient divergence of fluid-dynamic quantities in rarefied gases on smooth boundaries. Journal of Statistical Physics, 168(6):1319–1352, September 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Thu:2017:CFT

[TT17b] Nguyen Van Thu and Luong Thi Theu. Casimir force of twocomponent Bose–Einstein condensates confined by a parallel plate geometry. Journal of Statistical Physics, 168(1):1–10, July 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Teixeira:2017:EP

[TU17] Augusto Teixeira and Daniel Ungaretti. Ellipses percolation. Journal of Statistical Physics, 168(2):369–393, July 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Turkington:2013:OPD

[Tur13] Bruce Turkington. An optimization principle for deriving nonequilibrium statistical models of Hamiltonian dynamics. Journal of Statistical Physics, 152(3):569–597, August 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0778-9.

Toth:2012:SBS

[TV12] Bálint Tóth and Benedek Valkó. Superdiffusive bounds on self-repellent Brownian polymers and diffusion in the curl of the Gaussian free field in d = 2. Journal of Statistical Physics, 147(1):113–131, April 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0462-5.

Taufer:2015:CWE

[TV15a] Matthias Täufer and Ivan Veselić. Conditional Wegner estimate for the standard random breather potential. Journal of Statistical Physics, 161(4):902–914, November 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1358-y.

Turova:2015:BPG

[TV15b] Tatyana S. Turova and Thomas Vallier. Bootstrap percolation on a graph with random and local connections. Journal of Statistical Physics, 160(5):1249–1276, September 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1294-x.

Tautenhahn:2016:SIN

[TV16] Martin Tautenhahn and Ivan Veselić. Sampling inequality for L^{2}-norms of eigenfunctions, spectral projectors, and Weyl sequences of Schrödinger operators. Journal of Statistical Physics, 164(3):616–620, August 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-016-1560-
6.

Toth:2013:OMC

[TVP13] Bálint Tóth, Tamás Vicsek, and Gergely Palla. Overlapping modularity at the critical point of k-clique percolation. Journal of Statistical Physics, 151(3–4):689–706, May 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0640-5.

Tracy:2010:FAT

[TW10] Craig A. Tracy and Harold Widom. Formulas for ASEP with two-sided Bernoulli initial condition. Journal of Statistical Physics, 140(4):619–634, August 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-010-0013-x.pdf.

Teufl:2011:RSN

[TW11] Elmar Teufl and Stephan Wagner. Resistance scaling and the number of spanning trees in self-similar lattices. Journal of Statistical Physics, 142(4):879–897, February 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0140-z.

Tracy:2013:BGA

[TW13a] Craig A. Tracy and Harold Widom. The Bose gas and asymmetric simple exclusion process on the half-line. Journal of Statistical Physics, 150(1):1–12, January 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0686-4.

Tracy:2013:ASE

[TW13b] Craig A. Tracy and Harold Widom. On the asymmetric simple exclusion process with multiple species. Journal of Statistical Physics, 150(3):457–470, February 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0531-9.

Tracy:2014:SSE

[TW14] Craig A. Tracy and Harold Widom. On the singularities in the susceptibility expansion for the two-dimensional Ising model. Journal of Statistical Physics, 156(6):1125–1135, September 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1061-4.

Toppaladoddi:2017:SMC

[TW17] Srikanth Toppaladoddi and J. S. Wettlaufer. Statistical mechanics and the climatology of the Arctic sea ice thickness distribution. Journal of Statistical Physics, 167(3–4):683– 702, May 2017. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-016-1704-8.pdf.

tenWolde:2016:FLC

[tWBOM16] Pieter Rein ten Wolde, Nils B. Becker, Thomas E. Ouldridge, and Andrew Mugler. Fundamental limits to cellular sensing. Journal of Statistical Physics, 162(5):1395–1424, March 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-015-1440-5.pdf.

Takayasu:2014:GCL

[TWT14] Misako Takayasu, Hayafumi Watanabe, and Hideki Takayasu. Generalised central limit theorems for growth rate distribution of complex systems. Journal of Statistical Physics, 155(1):47– 71, April 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-0956-4.

Tyomkyn:2012:NAS

[Tyo12] Mykhaylo Tyomkyn. A note on the Abelian sandpile in ±Z^{d}. Journal of Statistical Physics, 148(6):1072–1075, September 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0564-0.

Tamura:2016:DOS

[TZ16] Hiroshi Tamura and Valentin A. Zagrebnov. Dynamics of an open system for repeated harmonic perturbation. Journal of Statistical Physics, 163(4):844–867, May 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1500-5.

Tzioufas:2013:RCT

[Tzi13] A. Tzioufas. Rates of convergence for the three state contact process in one dimension. Journal of Statistical Physics, 152(2):388–398, July 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0762-4.

Tzioufas:2018:CLT

[Tzi18] Achillefs Tzioufas. The central limit theorem for supercritical oriented percolation in two dimensions. Journal of Statistical Physics, 171(5):802–821, June 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Usui:2016:PTM

[UK16] Satoshi Usui and Hiroshi Koibuchi. Parallel tempering Monte Carlo simulations of spherical fixed–connectivity model for polymerized membranes. Journal of Statistical Physics, 162 (3):701–711, February 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1428-1.

Ukawa:2015:KWL

[Uka15] Akira Ukawa. Kenneth Wilson and lattice QCD. Journal of Statistical Physics, 160(5):1081–1124, September 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1197-x.

Uribe:2016:ERE

[UV16] F. J. Uribe and R. M. Velasco. Einstein relation for electrons in an electric field. Journal of Statistical Physics, 162(1):242– 266, January 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1386-7.

Vanicat:2017:ESI

[Van17] M. Vanicat. Exact solution to integrable open multispecies SSEP and macroscopic fluctuation theory. Journal of Statistical Physics, 166(5):1129–1150, March 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1705-7.

Varandas:2012:NUS

[Var12] Paulo Varandas. Non-uniform specification and large deviations for weak Gibbs measures. Journal of Statistical Physics, 146(2):330–358, January 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0392-7.

Vauchelet:2010:DLT

[Vau10] Nicolas Vauchelet. Diffusive limit of a two dimensional kinetic system of partially quantized particles. Journal of Statistical Physics, 139(5):882–914, June 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9970-3.

Vatansever:2012:SSS

[VAY^{+}12] E. Vatansever, B. O. Aktas, Y. Yüksel, U. Akinci, and H. Polat. Stationary state solutions of a bond diluted kinetic Ising model: An effective-field theory analysis. Journal of Statistical Physics, 147(6):1068–1076, July 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0519-5.

Venaille:2011:SPD

[VB11] Antoine Venaille and Freddy Bouchet. Solvable phase diagrams and ensemble inequivalence for two-dimensional and geophysical turbulent flows. Journal of Statistical Physics, 143(2):346–380, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0168-0.

vonBrecht:2013:SRG

[vBKBS13] James von Brecht, Theodore Kolokolnikov, Andrea L. Bertozzi, and Hui Sun. Swarming on random graphs. Journal of Statistical Physics, 151(1–2):150–173, April 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-012-0680-x.pdf.

vonBrecht:2015:SRG

[vBSB15] James H. von Brecht, Benny Sudakov, and Andrea L. Bertozzi. Swarming on random graphs II. Journal of Statistical Physics, 158(3):699–734, February 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-0923-0.

Vaitheeswaran:2011:HII

[VCT11] S. Vaitheeswaran, Jie Chen, and D. Thirumalai. Hydrophobic and ionic-interactions in bulk and confined water with implications for collapse and folding of proteins. Journal of Statistical Physics, 145(2):276–292, October 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-011-0313-9.

vanderHofstad:2017:WSF

[vdHK17] Remco van der Hofstad and Júlia Komjáthy. When is a scalefree graph ultra-small? Journal of Statistical Physics, 169 (2):223–264, October 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-017-1864-1.
pdf.

vanderHofstad:2018:CTC

[vdHKvL18] Remco van der Hofstad, Sandra Kliem, and Johan S. H. van Leeuwaarden. Cluster tails for critical power-law inhomogeneous random graphs. Journal of Statistical Physics, 171(1): 38–95, April 2018. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-018-1978-0.pdf.

vanderHoorn:2018:SME

[vdHLK18] Pim van der Hoorn, Gabor Lippner, and Dmitri Krioukov. Sparse maximum–entropy random graphs with a given powerlaw degree distribution. Journal of Statistical Physics, 173(3– 4):806–844, November 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

vanderHofstad:2018:TCC

[vdHvLS18] Remco van der Hofstad, Johan S. H. van Leeuwaarden, and Clara Stegehuis. Triadic closure in configuration models with unbounded degree fluctuations. Journal of Statistical Physics, 173(3–4):746–774, November 2018. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-018-1952-x.pdf.

vanEnter:2011:USS

[vEdG11] Aernout C. D. van Enter and Ellis de Groote. An ultrametric state space with a dense discrete overlap distribution: Paperfolding sequences. Journal of Statistical Physics, 142(2):223– 228, January 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-010-0107-5.pdf.

vanEnter:2016:TLR

[vEdLV16] A. C. D. van Enter, B. N. B. de Lima, and D. Valesin. Truncated long-range percolation on oriented graphs. Journal of Statistical Physics, 164(1):166–173, July 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1538-4.

vanEnter:2012:MTA

[vEF12] Aernout C. D. van Enter and Anne Fey. Metastability thresholds for anisotropic bootstrap percolation in three dimensions. Journal of Statistical Physics, 147(1):97–112, April 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-012-0455-4.pdf.

vanEnter:2019:NOD

[vEKRS19] Aernout C. D. van Enter, Bruno Kimura, Wioletta Ruszel, and Cristian Spitoni. Nucleation for one-dimensional long-range Ising models. Journal of Statistical Physics, 174(6):1327–1345, March 2019. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-019-02238-y.pdf.

Venegeroles:2014:QUC

[Ven14] Roberto Venegeroles. Quantitative universality for a class of weakly chaotic systems. Journal of Statistical Physics, 154(4):988–998, February 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0895-5.

Vucelja:2012:FIC

[VFT12] Marija Vucelja, Gregory Falkovich, and Konstantin S. Turitsyn. Fractal iso-contours of passive scalar in two-dimensional smooth random flows. Journal of Statistical Physics, 147 (2):424–435, April 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0474-1.

vanGinkel:2016:DSD

[vGRS16] Bart van Ginkel, Frank Redig, and Federico Sau. Duality and stationary distributions of the ”Immediate exchange model” and its generalizations. Journal of Statistical Physics, 163(1):92–112, April 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-016-1478-z.
pdf.

vanHemmen:2013:TFS

[vHL13] J. Leo van Hemmen and André Longtin. Temperature fluctuations for a system in contact with a heat bath. Journal of Statistical Physics, 153(6):1132–1142, December 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0867-9.

Vidybida:2015:AEN

[Vid15] Alexander K. Vidybida. Activity of excitatory neuron with delayed feedback stimulated with Poisson stream is non–Markov. Journal of Statistical Physics, 160(6):1507–1518, September 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1301-2.

Vidybida:2017:OSL

[Vid17] Alexander K. Vidybida. Output stream of leaky integrateand-fire neuron without diffusion approximation. Journal of Statistical Physics, 166(2):267–281, January 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1698-2.

Viehland:2016:UMT

[Vie16] Larry A. Viehland. Uniform moment theory for charged particle motion in gases. 2. Second approximation. Journal of Statistical Physics, 163(1):175–196, April 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1460-9.

vonKleist:2018:SAF

[vKSZ18] Max von Kleist, Christof Schütte, and Wei Zhang. Statistical analysis of the first passage path ensemble of jump processes. Journal of Statistical Physics, 170(4):809–843, February 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Vollmayr-Lee:2012:AKF

[VL12] Ben Vollmayr-Lee. Alex Kamenev: Field theory of nonequilibrium systems. Journal of Statistical Physics, 148(6): 1104–1105, September 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0571-1.

Veerman:2014:RDD

[VP14] J. J. P. Veerman and F. J. Prieto. On rank driven dynamical systems. Journal of Statistical Physics, 156(3):455–472, August 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1012-0.

Veerman:2015:ERD

[VP15] J. J. P. Veerman and F. J. Prieto. Erratum to: On rank driven dynamical systems. Journal of Statistical Physics, 161(5):1324, December 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-015-1372-
0;   http://link.springer.com/content/pdf/10.1007/
s10955-015-1372-0.pdf.

Vandebroek:2017:GLE

[VV17] Hans Vandebroek and Carlo Vanderzande. On the generalized Langevin equation for a Rouse bead in a nonequilibrium bath. Journal of Statistical Physics, 167(1):14–28, April 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-017-1734-x.

Vroylandt:2019:NED

[VV19] Hadrien Vroylandt and Gatien Verley. Non-equivalence of dynamical ensembles and emergent non-ergodicity. Journal of Statistical Physics, 174(2):404–432, January 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Velazquez:2018:TPC

[VW18] Juan J. L. Velázquez and Raphael Winter. The two-particle correlation function for systems with long–range interactions. Journal of Statistical Physics, 173(1):1–41, October 2018. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Vakeroudis:2011:MFR

[VYH11] S. Vakeroudis, M. Yor, and D. Holcman. The mean first rotation time of a planar polymer. Journal of Statistical Physics, 143(6):1074–1095, June 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0227-6.

Wang:2012:LER

[Wan12] Dong Wang. The largest eigenvalue of real symmetric, Hermitian and Hermitian self-dual random matrix models with rank one external source, Part I. Journal of Statistical Physics, 146(4):719–761, February 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0417-x.

Watanabe:2019:PBT

[Wat19] Haruki Watanabe. A proof of the Bloch theorem for lattice models. Journal of Statistical Physics, 177(4):717–726, November 2019. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-019-02386-1.pdf. See cor-

rection [Wat20].

Watanabe:2020:CPB

[Wat20] Haruki Watanabe. Correction to: A Proof of the Bloch Theorem for Lattice Models. Journal of Statistical Physics, 178(6):1515, March 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-020-02502-

6.pdf. See [Wat19].

Wang:2011:LSD

[WBL11] Jihang Wang, Dusan Bratko, and Alenka Luzar. Lengthscale dependence of hydration free energy: Effect of solute charge. Journal of Statistical Physics, 145(2):253–264, October 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0337-1.

Wittmer:2011:SFS

[WCX^{+}11] J. P. Wittmer, A. Cavallo, H. Xu, J. E. Zabel, P. Polińska, N. Schulmann, H. Meyer, J. Farago, A. Johner, S. P. Obukhov, and J. Baschnagel. Scale-free static and dynamical correlations in melts of monodisperse and Flory-distributed homopolymers. Journal of Statistical Physics, 145(4):1017–1126, November 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0354-0.

Webb:2011:EAF

[Web11] Christian Webb. Exact asymptotics of the freezing transition of a logarithmically correlated random energy model. Journal of Statistical Physics, 145(6):1595–1619, December 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0359-8.

Wegner:2014:MKG

[Weg14] Franz J. Wegner. In memory of Kenneth G. Wilson. Journal of Statistical Physics, 157(4–5):628–638, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-014-0988-9.pdf.

Wegner:2017:MLP

[Weg17] Franz J. Wegner. In memory of Leo P. Kadanoff. Journal of Statistical Physics, 167(3–4):420–426, May 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Weisbuch:2013:SDS

[Wei13] Gérard Weisbuch. Sustainable development and spatial inhomogeneities. Journal of Statistical Physics, 151(3–4):475– 493, May 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0639-y.

Wei:2016:LRD

[Wei16] Ran Wei. On the long–range directed polymer model. Journal of Statistical Physics, 165(2):320–350, October 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1612-y.

Wei:2018:FEC

[Wei18] Ran Wei. Free energy of the Cauchy directed polymer model at high temperature. Journal of Statistical Physics, 172(4): 1057–1085, August 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Wennberg:2012:FPL

[Wen12] Bernt Wennberg. Free path lengths in quasi crystals. Journal of Statistical Physics, 147(5):981–990, June 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0500-3.

Werness:2012:POS

[Wer12] Brent Morehouse Werness. The parafermionic observable in SLE. Journal of Statistical Physics, 149(6):1112–1135, December 2012. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-012-0657-9.

Williams:2011:NUS

[WES11] Stephen R. Williams, Denis J. Evans, and Debra J. Searles. Nonequilibrium umbrella sampling and the functional crooks fluctuation theorem. Journal of Statistical Physics, 145(4):831–840, November 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0281-0.

Wergen:2011:CBR

[WFK11a] Gregor Wergen, Jasper Franke, and Joachim Krug. Correlations between record events in sequences of random variables with a linear trend. Journal of Statistical Physics, 144 (6):1206–1222, September 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0307-7. See

erratum [WFK11b].

Wergen:2011:ECB

[WFK11b] Gregor Wergen, Jasper Franke, and Joachim Krug. Erratum to: Correlations Between Record Events in Sequences of Random Variables with a Linear Trend. Journal of Statistical Physics, 145(5):1405–1406, December 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-011-0376-7.pdf. See [WFK11a].

Wormell:2018:VLR

[WG18] Caroline L. Wormell and Georg A. Gottwald. On the validity of linear response theory in high-dimensional deterministic dynamical systems. Journal of Statistical Physics, 172(6): 1479–1498, September 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Wang:2019:DRP

[WG19] Michael Wang and Alexander Y. Grosberg. Dynamical response of passive and active particles to time-periodic mechanical forcing. Journal of Statistical Physics, 175(3–4):640– 663, May 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Wang:2011:SSM

[WGLE11] Chi-Jen Wang, Xiaofang Guo, Da-Jiang Liu, and J. W. Evans. Schloegl’s second model for autocatalysis on a cubic lattice: Mean-field-type discrete reaction–diffusion equation analysis. Journal of Statistical Physics, 144(6):1308–1328, September 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0288-6.

Wu:2014:AOG

[WHC14] Sheng-Jhih Wu, Chii-Ruey Hwang, and Moody T. Chu. Attaining the optimal Gaussian diffusion acceleration. Journal of Statistical Physics, 155(3):571–590, May 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0963-5.

Widom:2017:FEM

[Wid17] M. Widom. Frequency estimate for multicomponent crystalline compounds. Journal of Statistical Physics, 167(3–4):

726–734, May 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Wilkinson:2010:PTS

[Wil10] Michael Wilkinson. Perturbation theory for a stochastic process with Ornstein–Uhlenbeck noise. Journal of Statistical Physics, 139(2):345–353, April 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-9944-5.

Wilding:2011:ASE

[Wil11] Nigel B. Wilding. Accurate simulation estimates of phase behavior in ternary mixtures with prescribed composition. Journal of Statistical Physics, 144(3):652–662, August 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0250-7.

Webster:2018:SGH

[WK18] Jason R. Webster and Michael Kastner. Subexponentially growing Hilbert space and nonconcentrating distributions in a constrained spin model. Journal of Statistical Physics, 171 (3):449–461, May 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Wouters:2013:MLD

[WL13] Jeroen Wouters and Valerio Lucarini. Multi-level dynamical systems: Connecting the Ruelle response theory and the Mori– Zwanzig approach. Journal of Statistical Physics, 151(5):850– 860, June 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0726-8.

Wang:2017:NHK

[WLEC17] Chu Wang, Qianxiao Li, Weinan E, and Bernard Chazelle. Noisy Hegselmann–Krause systems: Phase transition and the 2R-conjecture. Journal of Statistical Physics, 166(5):1209– 1225, March 2017. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-017-1718-x.

Weinreich:2018:IHO

[WLJH18] Daniel M. Weinreich, Yinghong Lan, Jacob Jaffe, and Robert B. Heckendorn. The influence of higher-order epistasis on biological fitness landscape topography. Journal of Statistical Physics, 172(1):208–225, July 2018. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-018-1975-3.pdf.

Wust:2011:UBC

[WLL11] T. Wüst, Y. W. Li, and D. P. Landau. Unraveling the beautiful complexity of simple lattice model polymers and proteins using Wang–Landau sampling. Journal of Statistical Physics, 144(3):638–651, August 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0266-z.

Weeks:2011:ISI

[WP11a] John D. Weeks and Lawrence R. Pratt. Introduction to special issue on water and associated liquids. Journal of Statistical Physics, 145(2):207–208, October 2011. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-011-0382-9.pdf.

Wilkinson:2011:SOU

[WP11b] Michael Wilkinson and Alain Pumir. Spherical Ornstein– Uhlenbeck processes. Journal of Statistical Physics, 145(1): 113–142, October 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0332-6.

Weaver:2015:RCA

[WPB15] Iain S. Weaver and Adam Prügel-Bennett. Renormalisation of 2D cellular automata with an absorbing state. Journal of Statistical Physics, 159(2):211–220, April 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1142-4.

Wang:2010:DSM

[WQ10] Jia-Zeng Wang and Min Qian. Discrete stochastic modeling for epidemics in networks. Journal of Statistical Physics, 140(6):1157–1166, September 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0034-5.

Wreszinski:2012:GSE

[Wre12] Walter F. Wreszinski. The ground state energy per site of the quantum and classical Edwards–Anderson spin glass in the thermodynamic limit. Journal of Statistical Physics, 146 (1):118–124, January 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0401-x.

Wreszinski:2017:ENE

[Wre17] Walter F. Wreszinski. Equilibrium and non-equilibrium properties of superfluids and superconductors. Journal of Statistical Physics, 169(4):782–803, November 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Weber:2019:DVS

[WTM19] Dylan Weber, Ryan Theisen, and Sebastien Motsch. Deterministic versus stochastic consensus dynamics on graphs. Journal of Statistical Physics, 176(1):40–68, July 2019. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Wu:2014:EAT

[Wu14] Xintian Wu. Efficient algorithms for the two-dimensional Ising model with a surface field. Journal of Statistical Physics, 157(6):1284–1300, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1109-5.

Wu:2015:PDL

[Wu15] Kung-Chien Wu. Pointwise description for the linearized Fokker–Planck–Boltzmann model. Journal of Statistical Physics, 160(5):1277–1293, September 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1206-0.

Wu:2018:PAE

[Wu18] Hao Wu. Polychromatic arm exponents for the critical planar FK–Ising model. Journal of Statistical Physics, 170(6):1177– 1196, March 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Wehr:2016:UTC

[WW16] Jan Wehr and Aramian Wasielak. Uniqueness of translationcovariant zero-temperature metastate in disordered Ising ferromagnets. Journal of Statistical Physics, 162(2):487–494, January 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1413-8.

Wirtz:2016:CJC

[WWKK16] Tim Wirtz, Daniel Waltner, Mario Kieburg, and Santosh Kumar. The correlated Jacobi and the correlated Cauchy– Lorentz ensembles. Journal of Statistical Physics, 162(2):495– 521, January 2016. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1416-5.

Wang:2015:AEP

[WX15] Ran Wang and Lihu Xu. Asymptotics of the entropy production rate for d-dimensional Ornstein–Uhlenbeck processes. Journal of Statistical Physics, 160(5):1336–1353, September 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1295-9.

Wang:2016:ASE

[WXX16] Feng-Yu Wang, Jie Xiong, and Lihu Xu. Asymptotics of sample entropy production rate for stochastic differential equations. Journal of Statistical Physics, 163(5):1211–1234, June 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1513-0.

Wu:2016:AAT

[WYG16] Lei Wu, Xiongfeng Yang, and Yan Guo. Asymptotic analysis of transport equation in annulus. Journal of Statistical Physics, 165(3):585–644, November 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1623-8.

Wei:2012:CPI

[WZ12] Jinbo Wei and Xianwen Zhang. On the Cauchy problem for the inelastic Boltzmann equation with external force. Journal of Statistical Physics, 146(3):592–609, February 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0410-9.

Wu:2014:AEI

[WZIG14] Xintian Wu, Ru Zheng, Nickolay Izmailian, and Wenan Guo. Accurate expansions of internal energy and specific heat of critical two-dimensional Ising model with free boundaries. Journal of Statistical Physics, 155(1):106–150, April 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-0942-x.

Wang:2014:HMC

[WZL^{+}14] Jun Wang, Ji Zhou, Long-Jin Lv, Wei-Yuan Qiu, and Fu-Yao Ren. Heterogeneous memorized continuous time random walks in an external force fields. Journal of Statistical Physics, 156(6):1111–1124, September 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1046-3.

Xu:2018:LWM

[XD18] Pengbo Xu and Weihua Deng. Lévy walk with multiple internal states. Journal of Statistical Physics, 173(6):1598–1613, December 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Xu:2013:ALO

[XJZY13] Yong Xu, Xiaoqin Jin, Huiqing Zhang, and Tingting Yang. The availability of logical operation induced by dichotomous noise for a nonlinear bistable system. Journal of Statistical Physics, 152(4):753–768, August 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0782-0.

Xu:2015:PTB

[XLL^{+}15] Yong Xu, Yongge Li, Juanjuan Li, Jing Feng, and Huiqing Zhang. The phase transition in a bistable Duffing system driven by Lévy noise. Journal of Statistical Physics, 158 (1):120–131, January 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1129-1.

Xue:2017:CPR

[XP17] Xiaofeng Xue and Yu Pan. Contact processes with random recovery rates and edge weights on complete graphs. Journal of Statistical Physics, 169(5):951–971, December 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Xia:2014:NCE

[XTL14] Hui Xia, Gang Tang, and Yueheng Lan. Nonuniversality of critical exponents in a fractional quenched Kardar–Parisi– Zhang equation. Journal of Statistical Physics, 154(5):1228– 1240, March 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-0920-3.

Xia:2012:UBD

[XTpXpH12] Hui Xia, Gang Tang, Zhi peng Xun, and Da peng Hao. Universal behaviour of (2 + 1)-dimensional stochastic equations for epitaxial growth processes. Journal of Statistical Physics, 149(6):1086–1095, December 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0651-2.

Xue:2012:CDP

[Xue12] Xiaofeng Xue. Critical density points for threshold voter models on homogeneous trees. Journal of Statistical Physics, 146 (2):423–433, January 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0405-6.

Xue:2015:FLT

[Xue15] Xiaofeng Xue. Fluid limit of threshold voter models on tori. Journal of Statistical Physics, 159(2):274–293, April 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1192-2.

Xue:2016:CRS

[Xue16a] Xiaofeng Xue. Convergence rates for subcritical threshold-one contact processes on lattices. Journal of Statistical Physics, 162(2):371–386, January 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-015-1419-2.

Xue:2016:PTL

[Xue16b] Xiaofeng Xue. Phase transition for the large-dimensional contact process with random recovery rates on open clusters. Journal of Statistical Physics, 165(5):845–865, December 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1660-3.

Xiong:2013:OQR

[XY13] Sheng Xiong and Wei-Shih Yang. Open quantum random walks with decoherence on coins with n degrees of freedom. Journal of Statistical Physics, 152(3):473–492, August 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0772-2.

Xu:2017:OSD

[XZ17] Yi-Zhi Xu and Hai-Jun Zhou. Optimal segmentation of directed graph and the minimum number of feedback arcs. Journal of Statistical Physics, 169(1):187–202, October 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Xu:2019:TMT

[XZ19] Xining Xu and Yunxin Zhang. Theoretical model of transcription based on torsional mechanics of DNA template. Journal of Statistical Physics, 174(6):1316–1326, March 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Yaginuma:2016:SSI

[Yag16] K. Yaginuma. A stochastic system with infinite interacting components to model the time evolution of the membrane potentials of a population of neurons. Journal of Statistical Physics, 163(3):642–658, May 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1490-3.

Yamamoto:2013:ESG

[Yam13] Ken Yamamoto. Emergence of the Sierpiński gasket in coindividing problems. Journal of Statistical Physics, 152(3):534– 540, August 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0776-y.

Yamazaki:2014:NIM

[Yam14] Masahito Yamazaki. New integrable models from the Gauge/ YBE correspondence. Journal of Statistical Physics, 154(3): 895–911, February 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0884-8.

Yamazaki:2017:SHM

[Yam17] Kazuo Yamazaki. Stochastic Hall–magneto-hydrodynamics system in three and two and a half dimensions. Journal of Statistical Physics, 166(2):368–397, January 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-016-1683-9.

Yang:2011:SBL

[Yan11] Xiongfeng Yang. The solutions for the boundary layer problem of Boltzmann equation in a half-space. Journal of Statistical Physics, 143(1):168–196, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0158-2.

Yano:2015:QFP

[Yan15] Ryosuke Yano. On quantum Fokker–Planck equation. Journal of Statistical Physics, 158(1):231–247, January 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1123-7.

Yarmola:2014:SEM

[Yar14] Tatiana Yarmola. Sub-exponential mixing of open systems with particle–disk interactions. Journal of Statistical Physics, 156(3):473–492, August 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1014-y.

Yin:2014:CRS

[YB14] Yanqing Yin and Zhidong Bai. Convergence rates of the spectral distributions of large random quaternion self-dual Hermitian matrices. Journal of Statistical Physics, 157(6): 1207–1224, December 2014. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1096-6.

Yolcu:2017:GFR

[YBF^{+}17] Cem Yolcu, Antoine Bérut, Gianmaria Falasco, Artyom Petrosyan, Sergio Ciliberto, and Marco Baiesi. A general fluctuation–response relation for noise variations and its application to driven hydrodynamic experiments. Journal of Statistical Physics, 167(1):29–45, April 2017. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-017-1732-z.

Yedidia:2011:MPA

[Yed11] Jonathan S. Yedidia. Message-passing algorithms for inference and optimization. Journal of Statistical Physics, 145(4): 860–890, November 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0384-7.

Yoshida:2014:TRC

[YH14] Hiroaki Yoshida and Hidemitsu Hayashi. Transmissionreflection coefficient in the lattice Boltzmann method. Journal of Statistical Physics, 155(2):277–299, April 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-0953-7.

Yin:2010:FED

[Yin10] Jun Yin. Free energies of dilute Bose gases: Upper bound. Journal of Statistical Physics, 141(4):683–726, November 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-0066-x.

Yin:2013:CPE

[Yin13] Mei Yin. Critical phenomena in exponential random graphs. Journal of Statistical Physics, 153(6):1008–1021, December 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0874-x.

Yin:2016:DIN

[Yin16] Mei Yin. A detailed investigation into near degenerate exponential random graphs. Journal of Statistical Physics, 164(1):241–253, July 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1539-3.

Yasseri:2013:VPC

[YK13] Taha Yasseri and János Kertész. Value production in a collaborative environment. Journal of Statistical Physics, 151 (3–4):414–439, May 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0728-6.

Yasuoka:2016:WAV

[YKS16] Haruka Yasuoka, Masayuki Kaneda, and Kazuhiko Suga. Wall-adjacent velocity profiles of nano-scale gas flows. Journal of Statistical Physics, 165(5):907–919, December 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1662-1.

Yang:2011:SSW

[YM11] Jaw-Yen Yang and Bagus Putra Muljadi. Simulation of shock wave diffraction over 90^{∘} sharp corner in gases of arbitrary statistics. Journal of Statistical Physics, 145(6):1674–1688, December 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0355-z.

Yoo:2010:FMS

[Yoo10] Jisang Yoo. On factor maps that send Markov measures to Gibbs measures. Journal of Statistical Physics, 141(6): 1055–1070, December 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-010-0084-8.

Young:2017:GSM

[You17] Lai-Sang Young. Generalizations of SRB measures to nonautonomous, random, and infinite dimensional systems. Journal of Statistical Physics, 166(3–4):494–515, February 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-016-1639-0.

Yuzbashyan:2013:QIS

[YS13] Emil A. Yuzbashyan and B. Sriram Shastry. Quantum integrability in systems with finite number of levels. Journal of Statistical Physics, 150(4):704–721, February 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0689-9.

Yuge:2013:GEE

[YSSH13] Tatsuro Yuge, Takahiro Sagawa, Ayumu Sugita, and Hisao Hayakawa. Geometrical excess entropy production in nonequilibrium quantum systems. Journal of Statistical Physics, 153(3):412–441, November 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0829-2.

Yuhjtman:2015:SES

[Yuh15] Sergio A. Yuhjtman. A sensible estimate for the stability constant of the Lennard-Jones potential. Journal of Statistical Physics, 160(6):1684–1695, September 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1300-3.

Yamamoto:2010:TSS

[YY10] Ken Yamamoto and Yoshihiro Yamazaki. Topological selfsimilarity on the random binary-tree model. Journal of Statistical Physics, 139(1):62–71, April 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-010-9928-
5.

Yan:2011:DBT

[YYZ11] Weigen Yan, Yeong-Nan Yeh, and Fuji Zhang. Dimers belonging to three orientations on plane honeycomb lattices. Journal of Statistical Physics, 145(5):1343–1356, December 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0348-y.

Zhao:2013:PIN

[ZB13] Kun Zhao and Ginestra Bianconi. Percolation on interdependent networks with a fraction of antagonistic interactions.

Journal of Statistical Physics, 152(6):1069–1083, September 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-013-0806-9.

Zhang:2011:RFE

[ZBVE11] Qi Zhang, Jasna Brujić, and Eric Vanden-Eijnden. Reconstructing free energy profiles from nonequilibrium relaxation trajectories. Journal of Statistical Physics, 144(2):344– 366, July 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0242-7.

Zamparo:2019:MRT

[ZDG19] Marco Zamparo, Luca Dall’Asta, and Andrea Gamba. On the mean residence time in stochastic lattice–gas models. Journal of Statistical Physics, 174(1):120–134, January 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Zia:2011:MTP

[ZDS11] R. K. P. Zia, J. J. Dong, and B. Schmittmann. Modeling translation in protein synthesis with TASEP: A tutorial and recent developments. Journal of Statistical Physics, 144(2):405– 428, July 2011. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-011-0183-1.

Zhang:2011:MLP

[ZF11] Yunxin Zhang and Michael E. Fisher. Measuring the limping of processive motor proteins. Journal of Statistical Physics, 142(6):1218–1251, April 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0118-x.

Zhu:2018:IEW

[ZGH18] Yi Zhu, Aziz Guergachi, and Huaxiong Huang. The impacts of entrepreneurship on wealth distribution. Journal of Statistical Physics, 173(6):1734–1754, December 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Zhang:2013:FPT

[ZGL13] Yun-Xiu Zhang, Hui Gu, and Jin-Rong Liang. Fokker– Planck type equations associated with subordinated processes controlled by tempered α-stable processes. Journal of Statistical Physics, 152(4):742–752, August 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0754-4.

Zhang:2015:ESP

[ZGL15] Yun-Xiu Zhang, Hui Gu, and Jin-Rong Liang. Equivalence of subordinated processes with tempered α-stable waiting times and fractional Fokker–Planck equations in space and time dependent fields. Journal of Statistical Physics, 159(6):1495– 1503, June 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-014-1184-7.

Zhan:2010:RSC

[Zha10a] Dapeng Zhan. Reversibility of some chordal SLE(κ; ρ) traces. Journal of Statistical Physics, 139(6):1013–1032, June 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-9978-8.

Zhang:2010:CPV

[Zha10b] Xianwen Zhang. On the Cauchy problem of the Vlasov– Poisson-BGK system: Global existence of weak solutions. Journal of Statistical Physics, 141(3):566–588, November 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-010-0064-z.

Zhang:2010:MTF

[Zha10c] Yong Zhang. Moments for tempered fractional advection– diffusion equations. Journal of Statistical Physics, 139(5):915– 939, June 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-9965-0.

Zhan:2012:RPA

[Zha12a] Dapeng Zhan. Restriction properties of annulus SLE. Journal of Statistical Physics, 146(5):1026–1058, March 2012. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-012-0438-5.

Zhang:2012:IKI

[Zha12b] Pan Zhang. Inference of kinetic Ising model on sparse graphs. Journal of Statistical Physics, 148(3):502–512, August 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0547-1.

Zhang:2012:DFF

[Zha12c] Yu Zhang. A derivative formula for the free energy function. Journal of Statistical Physics, 146(2):466–473, January 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0371-z.

Zhang:2013:TAK

[Zha13] Yunxin Zhang. Theoretical analysis of kinesin KIF1A transport along microtubule. Journal of Statistical Physics, 152 (6):1207–1221, September 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0804-y.

Zhang:2014:MSI

[Zha14] Zuhe Zhang. Merrifield–Simmons index and its entropy of the 4–8–8 lattice. Journal of Statistical Physics, 154(4):1113–1123, February 2014. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0883-9.

Zhang:2015:LUB

[Zha15] Lun Zhang. Local universality in biorthogonal Laguerre ensembles. Journal of Statistical Physics, 161(3):688–711, November 2015. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-015-1353-3.

Zhelezov:2013:PRO

[Zhe13] Dmitry Zhelezov. On a property of random-oriented percolation in a quadrant. Journal of Statistical Physics, 153(5): 751–762, December 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-013-0856-z.

Zhou:2017:EEI

[Zho17] S. Zhou. Effective electrostatic interactions between two overall neutral surfaces with quenched charge heterogeneity over atomic length scale. Journal of Statistical Physics, 169(5): 1019–1037, December 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Zhou:2018:WTN

[Zho18] Shiqi Zhou. Wetting transition of nonpolar neutral molecule system on a neutral and atomic length scale roughness substrate. Journal of Statistical Physics, 170(5):979–998, March 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Zhang:2016:RBD

[ZHRB16] Xiaojun Zhang, Zheng He, and Lez Rayman-Bacchus. Random birth-and-death networks. Journal of Statistical Physics, 162(4):842–854, February 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-016-1447-6.

Zhu:2017:ASC

[Zhu17] Lingjiong Zhu. Asymptotic structure of constrained exponential random graph models. Journal of Statistical Physics, 166(6):1464–1482, March 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-017-1733-y.

Zhao:2015:SMM

[ZHZ15] Jin-Hua Zhao, Yusupjan Habibulla, and Hai-Jun Zhou. Statistical mechanics of the minimum dominating set problem. Journal of Statistical Physics, 159(5):1154–1174, June 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-015-1220-2.

Zia:2010:TFY

[Zia10] R. K. P. Zia. Twenty five years after KLS: A celebration of non-equilibrium statistical mechanics. Journal of Statistical Physics, 138(1–3):20–28, February 2010. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-009-9884-0.pdf.

Zinn-Justin:2014:RVM

[ZJ14] Jean Zinn-Justin. Random vector and matrix theories: A renormalization group approach. Journal of Statistical Physics, 157(4–5):990–1016, December 2014. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-014-1103-y.

Zhang:2013:CGR

[ZL13] Chenfei Zhang and Yueheng Lan. Computation of growth rates of random sequences with multi-step memory. Journal of Statistical Physics, 150(4):722–743, February 2013. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-013-0712-1.

Zhang:2019:FRA

[ZL19] Hong Zhang and Guo-Hua Li. Fluid reactive anomalous transport with random waiting time depending on the preceding jump length. Journal of Statistical Physics, 174(3):548–561, February 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Zhang:2013:FFK

[ZLL13] Hong Zhang, Guo-Hua Li, and Mao-Kang Luo. Fractional Feynman–Kac equation with space-dependent anomalous exponent. Journal of Statistical Physics, 152(6):1194–1206, September 2013. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-013-0810-0.

Zhang:2019:MS

[ZMD^{+}19] Weiting Zhang, Vanessa Mayr, Bertrand Ducos, Martin Distel, and David Bensimon. A model of somitogenesis. Journal of Statistical Physics, 175(3–4):729–742, May 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Zocca:2018:LTB

[Zoc18] Alessandro Zocca. Low-temperature behavior of the multicomponent Widom–Rowlison model on finite square lattices. Journal of Statistical Physics, 171(1):1–37, April 2018. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Zhao:2015:SED

[ZP15] Yun Zhao and Yakov Pesin. Scaled entropy for dynamical systems. Journal of Statistical Physics, 158(2):447–475, January 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-014-1133-5. See erratum [ZP16].

Zhao:2016:ESE

[ZP16] Yun Zhao and Yakov Pesin. Erratum to: Scaled Entropy for Dynamical Systems. Journal of Statistical Physics, 162 (6):1654–1660, March 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-016-1451-x.

pdf. See [ZP15].

Zhang:2015:PSS

[ZSHL15] Ting Zhang, Baochang Shi, Changsheng Huang, and Hong Liang. Pore-scale study of miscible displacements in porous media using lattice Boltzmann method. Journal of Statistical Physics, 161(6):1453–1481, December 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s10955-015-1371-1;   http://link.springer.com/
content/pdf/10.1007/s10955-015-1371-1.pdf.

Zernik:2018:EME

[ZST18] Amitai Netser Zernik, Tomer M. Schlank, and Ran J. Tessler. Exact maximum-entropy estimation with Feynman diagrams. Journal of Statistical Physics, 170(4):731–747, February 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Zucca:2011:SEA

[Zuc11a] Fabio Zucca. Survival, extinction and approximation of discrete-time branching random walks. Journal of Statistical Physics, 142(4):726–753, February 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/s10955-011-0134-
x.

Zucker:2011:YWI

[Zuc11b] I. J. Zucker. 70+ years of the Watson integrals. Journal of Statistical Physics, 145(3):591–612, November 2011.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0273-0.

Zeng:2010:CNE

[ZW10] Chunhua Zeng and Hua Wang. Colored noise enhanced stability in a tumor cell growth system under immune response. Journal of Statistical Physics, 141(5):889–908, December 2010. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
article/10.1007/s10955-010-0068-8.

Zhou:2012:RGP

[ZW12] Haijun Zhou and Chuang Wang. Region graph partition function expansion and approximate free energy landscapes: Theory and some numerical results. Journal of Statistical Physics, 148(3):513–547, August 2012. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-012-0555-1.

Zhong:2013:SRL

[ZWGM13] Suchuan Zhong, Kun Wei, Shilong Gao, and Hong Ma. Stochastic resonance in a linear fractional Langevin equation. Journal of Statistical Physics, 150(5):867–880, March 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-012-0670-z.

Zhong:2015:TNI

[ZWGM15] Suchuan Zhong, Kun Wei, Shilong Gao, and Hong Ma. Trichotomous noise induced resonance behavior for a fractional oscillator with random mass. Journal of Statistical Physics, 159(1):195–209, April 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-014-1182-9.

Zhao:2018:ISR

[ZYZ^{+}18] Jian Zhao, Jianhua Yang, Jingling Zhang, Chengjin Wu, and Dawen Huang. Improving the stochastic resonance in a bistable system with the bounded noise excitation. Journal of Statistical Physics, 173(6):1688–1697, December 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).